A Complete Bibliography of Publications in The Journal of Supercomputing: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

26 March 2024 Version 3.03

Title word cross-reference

(1|1) [187]. (s, t) [1591]. 2 [191, 231, 368, 397, 470, 552, 827, 894, 923, 952, 1120, 1126, 1441, 1805, 1806, 1928, 2307]. 3 [151, 277, 294, 378, 587, 698, 953, 1032, 1076, 1106, 1126, 1441, 1513, 1526, 1532, 1549, 1573, 1614, 1668, 1720, 1729, 1871, 1946, 2147, 2203, 2208, 2246, 2355, 2359, 2360]. 4 [736, 848, 1229]. ^{∗} [615]. _{3} [579]. A∗ [19]. β [1075]. C [1137]. d [1722]. K [50, 56, 216, 298, 412, 730, 770, 771, 867, 973, 999, 1255, 1368, 1430, 1721, 1769, 1830]. DS_{spirit} [1244]. μ [981]. N [216, 353, 1255, 1365, 1651, 1803]. P [836, 1607, 2329]. Q [1894]. QR [2032]. s [1255]. ε [775].

-ary [216, 1255]. -based [669]. -body [1651]. -centroid [187]. -core [56]. -coverage [771]. -Cube [353]. -D [231, 1229, 1946]. -dimensional [1365]. -direct [1255]. -disjoint [848]. -Disjoint-path-coverable [1120]. -divergence [1075]. -Epidemic [2329]. -fuzzy [775]. -gram [1803]. -harmonic [770]. -indirect [1255]. -learning-based [1894]. -Means [50, 867, 973, 1137, 1430, 1721, 1769, 1830]. -median [1607]. -NN [412]. -pairwise [730]. -paths [1591]. -SAT [378]. -Set [999]. -tree [216].

/manycore [1719].

10 [1968]. 128 [906].

2 [1729]. 2D [1189].

3 [902]. 3D [698, 1442, 1463]. 3D-HEVC [1463]. 3D-NoC [698]. 3DyRM [943].

4.0 [2164].

6 [1704]. 600JD [2356].

7T [1861].

80 [906]. 802.15.4 [350, 391]. 802.15.4/ZigBee [391]. 802.16e [238].

AA [1987]. AA-DBSCAN [1987]. ABC [2349]. ABC-FCM [2349]. ability [217, 1791]. abnormal [1989]. abstract [725]. Abstraction [197, 240, 388, 1909]. abuse [541]. Academic [408]. accelerate [188, 1412, 2155]. Accelerated [440, 821, 1184, 1320, 1439, 1454, 1624, 1920, 1969]. Accelerating [174, 222, 443, 605, 627, 894, 951, 1014, 1099, 1103, 1153, 1159, 1409, 1421, 1503, 2049, 2133, 2157, 2257, 2322]. Acceleration [1621, 2035, 2043, 2060, 2076]. accelerator [329, 936, 2036, 2229, 2324]. accelerators [1235, 1639, 1712]. accelerometers [1330]. access [49, 71, 83, 115, 288, 358, 373, 438, 481, 728, 740, 852, 899, 944, 1331, 1539, 1677, 1748, 1877, 1943, 2112, 2328]. accessed [1216]. accesses [2302]. Accessing [18, 959, 1407, 1565, 1817]. accident [1327, 1328]. accompaniment [1419]. accountability [1061, 1922]. accountable [590]. accounting [482]. accumulative [2075]. accuracy [602, 1074, 1083, 1312]. accurate [177, 1734]. ACE [1596]. achieve [247]. achieved [2237]. Achieving [241, 541, 893, 1263, 1984]. ACK [236]. ACK-based [236]. Acknowledgment [1897]. Acknowledgment-based [1897]. ACO [1644, 2024]. acoustic [172, 471, 1932]. acoustic-scattering [471]. acquired [1697]. acquisition [1355, 1382, 1808]. across [1147, 2324]. ACT [669]. action [821]. actions [1162]. Active [12, 369, 787, 1326, 1611, 1999, 2050]. active-set [2050]. Activities [64]. Activity [659, 1352, 2092, 2177, 2241]. Actor [1715]. acyclic [750]. ad [30, 47–49, 100, 204, 205, 213, 314, 345, 346, 406, 497, 502, 667, 716, 728, 1089, 1132, 1262, 1273, 1530, 1561, 1750, 1762, 1897, 1904, 1939]. ad-hoc [497, 667]. Adaboost [2337]. Adaboost-based [2337]. adaptable [1776]. adaptation [974]. adaptations [972]. Adapting [681, 1396]. Adaptive [8, 22, 35, 38, 39, 49, 128, 182, 183, 216, 243, 269, 293, 298, 441, 466, 522, 593, 624, 632, 661, 669, 761, 789–791, 793, 798, 803, 809, 923, 1008, 1027, 1038, 1047, 1096, 1168, 1172, 1213, 1225, 1230, 1239, 1282, 1368, 1475, 1479, 1522, 1651, 1682, 1757, 1790, 1816, 1844, 1867, 1930, 1956, 1971, 1987, 2088, 2206, 2323]. adaptivity [1239]. adder [1292, 1741, 1869]. adders [2098]. addition [1680]. additional [462]. additive [1759, 1812]. Addressing [1054]. ADITHE [157]. adjacent [239]. adjusting [792]. adjustment [72, 425, 493, 1563]. Admissible [91]. Admission [60, 435, 1226, 1895]. advance [244, 287, 779]. Advanced [565, 620, 677, 718, 887, 989, 1298, 1560, 2005]. advancement [1703]. advancements [497]. Advances [302, 344, 489, 789, 1121, 1344, 1537, 2002]. advection [1445]. advection-dispersion [1445]. adversarial [2089]. adversaries [92]. ADVS [514]. aerial [594, 1876]. aerial-image [594]. aerodynamic [1153, 2118]. AES [735, 2300]. AF [1965]. affine [982]. affordable [2054]. after [1891]. against [12, 92, 1095, 1143, 1336, 1550, 1650, 1687, 1884, 1887, 1934, 2141, 2222]. age [2177]. Agent [23, 28, 214, 245, 372, 439, 590, 1009, 1313, 1887, 2066]. Agent-based [28, 439, 590, 1313, 2066]. agents [284, 516, 827, 1298, 2357]. aggregated [371]. Aggregating [2011]. aggregation [517, 1065, 2178]. agile [2237]. agreement [390, 430, 672, 1138, 1177, 1190, 1283, 1740, 1857, 1868, 1923]. AGVs [2265]. ahead [1133]. AHP [573]. AI [1872, 2193]. AIC [177]. AIC-based [177]. aided [98, 700, 1262]. aiming [1435]. air [1900]. Aircraft [936]. airspace [1167]. ALBR [248]. alerts [387]. algebra [156]. algebraic [710, 725]. algebras [198]. Algorithm [52, 79, 142, 167, 224, 228, 321, 362, 368, 473, 604, 609, 618, 658, 694, 696, 698, 763, 770, 771, 785, 830, 858, 886, 894, 907, 923, 942, 952, 998, 1015, 1027, 1044, 1049, 1050, 1072, 1080, 1180, 1266, 1299, 1320, 1348, 1353, 1385, 1394, 1408, 1420, 1424, 1430, 1437, 1449, 1461, 1507, 1582, 1607, 1632, 1642, 1656, 1658, 1667, 1673, 1680, 1762, 1769, 1772, 1807, 1813, 1832, 1858, 1904, 1941, 2015, 2017, 2018, 2025, 2029, 2035, 2048, 2049, 2067, 2088, 2096, 2108, 2109, 2112, 2115, 2122, 2236, 2239, 2248, 2267, 2272, 2273, 2277, 2300, 2304, 2370, 2374]. algorithm [34, 60, 166, 185, 237, 239, 248, 311, 313, 341, 345, 359, 371, 383, 394, 458, 539, 552, 565, 582, 583, 629, 639, 644, 707, 720, 741, 750, 805, 862, 888, 957, 970, 975, 1039, 1119, 1154, 1160, 1169, 1189, 1261, 1304, 1306– 1308, 1310, 1422, 1448, 1464, 1475, 1481, 1492, 1506, 1512, 1539, 1583, 1584, 1591, 1592, 1597, 1601, 1640, 1662, 1664, 1665, 1692, 1693, 1702, 1721, 1722, 1726, 1750, 1755, 1770, 1794, 1831, 1903, 1905, 1943, 1970, 2024, 2039, 2041, 2050, 2063, 2111, 2144, 2162, 2173, 2242, 2253, 2259, 2262, 2268, 2309, 2327, 2341, 2349, 2355]. algorithm [14, 88, 230, 401, 423, 529, 713, 1046, 1102, 1333, 1363, 1443, 1621, 1634, 1655, 1814, 1819, 1847, 1930, 2020, 2121, 2265, 2269, 2275, 2276]. algorithm-based [886, 2088]. Algorithmic [1098, 1631]. Algorithms [52, 70, 109, 133, 170, 182, 187, 216, 346, 361, 370, 397, 438, 459, 462, 493, 602, 642, 657, 699, 701, 726, 748, 758, 766, 802, 823, 825, 831, 857, 865, 897, 898, 938, 1004, 1033, 1056, 1077, 1137, 1184, 1239, 1253, 1259, 1322, 1342, 1344, 1373, 1416, 1446, 1484, 1503, 1504, 1533, 1574, 1587, 1626, 1688, 1732, 1733, 1784, 1900, 1991, 2003, 2060, 2077, 2110, 2114, 2189, 2213, 2245, 2261, 2284, 2294, 2297, 2312, 2362, 2365, 2370]. aliasing [379]. Aligning [317]. alignment [161, 267, 457, 476, 602, 1333, 1403, 1484, 1621, 1624, 2028, 2051, 2096]. alignment-free [1484]. all-optical [1243, 1455]. all-pairs-shortest-path [1186]. all-port [368]. all-reduce [1785]. All-to-all [827, 1112]. Alliance [314]. Alliance-based [314]. allocate [1132]. allocating [1827]. allocation [72, 76, 100, 116, 184, 282, 315, 434, 457, 507, 519, 528, 588, 654, 655, 692, 855, 875, 897, 1002, 1040, 1119, 1142, 1168, 1227, 1228, 1259, 1270, 1285, 1337, 1390, 1447, 1476, 1559, 1604, 1636, 1642, 1682, 1745, 1787, 1847, 1895, 1902, 1999, 2187, 2202, 2289, 2315, 2321, 2369]. Almanac [1466, 1467]. Almost [612, 1717]. ALOHA [888]. alone [1383]. alternate [1192]. alternating [73, 202, 318]. alternative [428]. alternatives [191, 2186]. ALU [2233]. ambient [676]. Amdahl [637]. among [842]. AMR [1172]. AMRC [710]. analog [883]. analog/digital [883]. Analysing [464]. Analysis [43, 104, 142, 206, 214, 216, 248, 290, 391, 398, 409, 415, 472, 476, 492, 545, 573, 611, 642, 646, 659, 687, 703, 725, 745, 812, 817, 842, 848, 866, 875, 904, 928, 939, 941, 963, 979, 1011, 1081, 1089, 1124, 1138, 1141, 1161, 1193, 1203, 1245, 1249, 1281, 1299, 1303, 1305, 1319, 1327, 1335, 1338, 1350, 1360, 1370, 1424, 1430, 1432, 1459, 1511, 1539, 1540, 1545, 1551, 1610, 1619, 1690, 1695, 1700, 1741, 1748, 1768, 1769, 1815, 1829, 1859, 1873, 1972, 2001, 2030, 2041, 2043, 2050, 2081, 2082, 2092, 2094, 2098, 2104, 2119, 2147, 2156, 2159, 2195, 2215, 2266, 2272, 2367]. analysis [107, 272, 543, 665, 862, 911, 918, 1257, 1258, 1328, 1593, 1638, 1839, 2160]. analytic [460, 519, 719, 1288]. Analytical [9, 233, 715, 801, 1000, 1106, 1201, 1476, 1707, 2074, 2242]. analytics [1326, 1554, 1892, 2187]. analyze [56]. analyzer [806, 1127]. Analyzing [53, 162, 263, 480, 632, 1064, 1084, 1359, 1945, 2297]. Ancilla [1257, 1258, 1277]. Ancilla-input [1277]. Android [742, 1220, 1222, 1953]. Android^{T M} [939]. Anemone [1715]. ANFIS [2241]. ANG [2010]. Angle [502, 1106]. Anisotropic [1722]. ANN [1900]. annealing [1506, 1762, 1775, 2229]. annotations [1910]. Anomaly [50, 186, 290, 491, 555, 1079, 1557, 1929, 2212, 2279]. anomaly-based [491]. anonymity [351]. Anonymous [30, 96, 667, 903, 908, 1164, 1185, 1283, 1295, 1727, 1817, 1857, 2348]. ant [458, 569, 840, 1644, 1954, 2248, 2265]. ant-stigmergy [458]. ant-swarm [569]. Anti [352, 888, 1222, 1842]. anti-collision [888]. Anti-debugging [1222]. anti-MITM [352]. Anti-reversible [1842]. AnyNoC [1853]. AODV [48]. AP [560]. Apache [1601, 1605, 2305]. aperiodic [1069]. aperture [1232]. APGAS [1711]. API [295, 1909]. app [539, 1289]. appeal [2083]. applicability [1484]. Application [30, 139, 221, 226, 311, 367, 382, 391, 468, 482, 506, 576, 616, 621, 709, 777, 797, 880, 940, 972, 978, 1025, 1033, 1093, 1167, 1173, 1180, 1219, 1329, 1336, 1357, 1364, 1373, 1401, 1454, 1464, 1494, 1510, 1546, 1575, 1582, 1597, 1619, 1696, 1731, 1778, 1834, 1846, 1852, 1863, 2189, 2218, 2261, 2268, 2269, 2272, 2275, 2276, 2360]. Application-aware [1025, 1494, 2218]. application-level [139, 940, 978, 1401]. application-oriented [1778]. application-specific [226, 311, 1510, 1852]. Applications [1, 4, 6, 18, 32, 35, 69, 174, 177, 190, 207, 240, 244, 257, 285, 336, 385, 409, 421, 437, 489, 508, 509, 537, 607, 610, 620, 651, 712, 714, 718, 778, 811, 815, 826, 832, 889, 920, 971, 992, 1008, 1019, 1060, 1071, 1096, 1097, 1140, 1181, 1206, 1275, 1339, 1340, 1356, 1368, 1388, 1391, 1401, 1417, 1418, 1461, 1495, 1513, 1516, 1520, 1566, 1576, 1586, 1588, 1596, 1627, 1659, 1670, 1688, 1752, 1760, 1765, 1781, 1839, 1842, 1861, 1901, 1940, 1946, 1947, 1952, 2000, 2066, 2085, 2116, 2139, 2193, 2197, 2228, 2274, 2283, 2289, 2316, 2323, 2362]. applications/jobs [832]. applied [469, 1437, 2068]. Applying [1412, 2292]. approach [8, 50, 94, 137, 211, 328, 348, 365, 414, 424, 425, 455, 465, 473, 483, 488, 510, 555, 593, 629, 673, 726, 733, 734, 744, 800, 830, 868, 872, 932, 954, 1053, 1065, 1068, 1069, 1078, 1090, 1104, 1126, 1176, 1181, 1206, 1211, 1247, 1249, 1250, 1314, 1321, 1329, 1375, 1402, 1451, 1502, 1510, 1563, 1643, 1649, 1675, 1722, 1749, 1818, 1873, 1886, 1946, 1954, 2013, 2054, 2069, 2094, 2099– 2101, 2109, 2119, 2120, 2138, 2149, 2155, 2191, 2195, 2198, 2247, 2249, 2252, 2258, 2287, 2293, 2299, 2334, 2335]. approaches [168, 170, 250, 476, 609, 798, 935, 1406, 1631, 2044, 2224, 2286]. Appropriate [1707, 2012]. approximate [114, 726, 922, 1317, 1353, 1651, 1732, 1733, 1987, 2271]. Approximation [60, 752, 854, 855, 1940, 2122, 2280]. approximation-based [60]. Approximative [319]. apps [539, 1222]. Apriori [1703, 2010]. Apriori-based [1703]. APS [1776]. APSP [950]. APT [1540, 2211]. arbitrary [309, 359]. ARC [326]. Architectural [1100, 2325]. Architecture [20, 86, 121, 141–143, 150, 218, 258, 265, 283, 329, 336, 376, 513, 525, 566, 639, 669, 698, 743, 766, 785, 786, 796, 815, 841, 869, 872, 898, 925, 932, 967, 1052, 1060, 1091, 1136, 1140, 1158, 1180, 1209, 1221, 1271, 1334, 1377, 1457, 1532, 1556, 1576, 1644, 1672, 1677, 1720, 1795, 1805, 1806, 1810, 1826, 1841, 1849, 1997, 2000, 2053, 2097, 2113, 2166, 2176, 2228, 2248, 2283, 2303]. architectures [53, 130, 158, 163, 174, 182, 324, 360, 375, 397, 808, 836, 863, 908, 919, 936, 948, 951, 1009, 1019, 1033, 1079, 1080, 1115, 1118, 1121, 1153, 1246, 1395, 1415, 1431, 1587, 1614, 1639, 1667, 1718, 1759, 1760, 1798, 1930, 2031, 2313, 2354, 2371]. area [330, 497, 571, 773, 800, 839, 967, 1219, 1453, 1556, 1833, 2233]. areas [1957]. arithmetic [506]. arm [557, 1409, 1431, 1614, 1810, 2311]. ARM-based [1431]. ARP [1290]. arrangement [800, 1585]. Array [501, 763, 1187, 1243, 1455, 1635]. arrays [267, 392, 475, 857, 1041, 1535, 1570, 1988]. arrival [9, 1082, 1304, 1785]. art [563]. arthritis [2247]. artifact [2095]. Artificial [1104, 1245, 1349, 1514, 1650, 1832, 1900, 2114, 2197, 2236, 2253, 2268, 2269, 2327, 2349]. ary [216, 1255]. ASF [1883]. aspect [2119, 2266]. AspectFrameNet [1459]. aspects [1098, 1459]. assembly [630, 873, 1004]. assess [2291]. Assessing [1417, 1911, 2183]. Assessment [639, 891, 989, 1148, 1565, 1921, 2081, 2130, 2216, 2242]. assignation [382]. assignment [325, 406, 893, 1597, 1602]. assisted [241, 410, 707, 723, 765, 968, 979, 1545, 2103]. association [538, 560, 695, 1048, 1227, 1305]. assurance [28]. asymmetric [952]. asymmetrical [194]. asymptotic [2297]. Asynchronous [52, 61, 194, 571, 729, 1134, 1520, 1574, 1775, 1823, 2231, 2348]. Atlas [2015]. atmospheric [117]. Atrak [1625]. ATSDS [1522]. attack [387, 516, 673, 966, 1139, 1550, 1580, 1650, 1934, 2138, 2222]. attacking [673]. attacks [555, 1062, 1290, 1336, 1540, 1616, 1687, 1884, 1887, 2211, 2350]. attention [1921]. attestation [94, 1915]. attitude [1811]. attractive [191]. Attribute [366, 783, 982, 1061, 1918, 2068, 2341]. Attribute-based [366, 1061, 1918]. attribute-hiding [982]. attributes [946, 1886, 2271]. auction [1636, 1822]. auctions [885]. audio [190, 989, 1294, 1403, 1409]. audio-to-score [1403]. auditability [1922]. auditing [1655, 1916, 1931, 1982, 2373]. augment [1093]. augmentation [2149]. augmented [326, 330, 610, 861]. authenticated [665, 672, 738, 844, 914, 1283, 1838, 1868]. Authentication [27, 298, 351, 430, 522, 599, 674, 905, 908, 960, 961, 1138, 1177, 1218, 1347, 1371, 1465, 1486, 1580, 1669, 1687, 1727, 1740, 1804, 1817, 1843, 1918, 1923, 1934, 1944, 1964, 2210, 2268, 2269, 2347]. authorization [27, 965]. Auto [530, 1557, 2364]. auto-encoders [1557]. Auto-tuning [530, 2364]. autocorrelation [195]. autoencoder [1133]. autoencoders [2252]. automata [129, 138, 237, 345, 461, 488, 574, 575, 578– 580, 699, 820, 827, 1027, 1268, 1313, 1365, 1444, 1445, 1741, 1865, 1869, 2251, 2255, 2294, 2308, 2332, 2357]. automata-based [138, 237, 345, 699, 1027, 2294]. Automated [86, 887, 912, 1367, 1457, 1767, 2199]. Automatic [157, 179, 534, 778, 806, 946, 1024, 1081, 1087, 1097, 1279, 1400, 1411, 1435, 1497, 1788, 1803, 1825, 2105, 2131, 2142, 2144, 2227, 2360]. Automatically [85]. automation [1371]. automaton [819, 2356]. Autonomous [245, 746, 1355, 1878]. autoregressive [148]. autotuning [953]. availability [333, 686, 731, 779, 866, 881, 1678, 1777, 1870, 2216, 2291]. availability-aware [779]. available [43, 1530]. AVC [477, 970, 1408, 1414]. Average [839]. averaging [552]. avoidance [571, 1252]. avoiding [2343]. AVX [928]. aware [24, 33, 65, 66, 68, 106, 143, 176, 223, 277, 288, 292, 311, 325, 327, 360, 366, 401, 407, 444, 515, 522, 524, 529, 585, 626, 633, 656, 676, 715, 717, 753, 779, 805, 818, 846, 868, 990, 995, 1003, 1025, 1032, 1034, 1043, 1054, 1061, 1077, 1101, 1109, 1118, 1155, 1200, 1205, 1242, 1253, 1271, 1318, 1461, 1481, 1489, 1494, 1506, 1510, 1577, 1627, 1663, 1664, 1673, 1676, 1696, 1757, 1761, 1783, 1827, 1862, 1864, 1871, 1901, 1905, 1919, 1928, 1939, 1943, 1973, 1979, 1994, 2006, 2099, 2100, 2114, 2128, 2131, 2171, 2209, 2218, 2243, 2258, 2303, 2304, 2309, 2320, 2353]. aware [1665]. awareness [495, 874, 890, 1236, 1454, 2197]. AXC [1774].

BabSang [66]. back [370, 1035, 1349, 1661]. back-pressure [1035]. back-propagation [1349]. backfill [81]. backup [666, 1981]. bag [1773]. bag-of-tasks [1773]. balance [1079, 1236]. balanced [116, 560, 615, 662, 752, 760, 1325, 1796, 1998, 2128, 2169]. Balancing [34, 45, 60, 141, 169, 183, 213, 252, 298, 585, 604, 728, 783, 798, 814, 1135, 1145, 1237, 1351, 1418, 1588, 1641, 1683, 1699, 1791, 1799, 1819, 1898, 1913, 1991, 2037, 2071, 2304, 2336]. band [1020]. bandwidth [33, 43, 401, 402, 598, 655, 1455, 1653, 1827]. bandwidth-aware [33, 401]. bank [1152]. banking [99]. banks [723]. bargaining [425, 1282, 1787]. bargaining-based [1282]. bargaining-driven [425]. barrel [1257, 1258]. Barrier [417, 1640]. base [1298]. based [51, 57, 109, 128, 153, 184, 205, 265, 281, 323, 333, 350, 421, 422, 424, 499, 530, 594, 606, 609, 635, 640, 672, 678, 685, 718, 740, 763, 770, 809, 819, 835, 886, 905, 910, 940, 1011, 1027, 1044, 1051– 1053, 1066, 1080, 1136, 1137, 1163, 1188, 1205, 1231, 1254, 1355, 1373, 1384, 1431, 1507, 1560, 1578, 1594, 1607, 1625, 1626, 1632, 1657, 1695, 1738, 1765, 1796, 1797, 1804, 1843, 1858, 1897, 1908, 1914, 1936, 1947, 1962, 1966, 2030, 2034, 2055, 2075, 2088, 2094, 2112, 2238, 2249, 2254, 2256, 2277, 2284, 2300, 2308, 2317, 2347, 2370]. based [79, 147, 250, 320, 363, 378, 439, 545, 584, 585, 623, 680, 684, 694, 726, 745, 746, 775, 798, 817, 844, 872, 877, 908, 931, 960, 961, 963, 1002, 1009, 1048, 1081, 1124, 1148, 1186, 1189, 1199, 1208, 1223, 1265, 1268, 1270, 1283, 1286, 1331, 1369, 1371, 1375, 1410, 1426, 1462, 1568, 1656, 1694, 1742, 1781, 1790, 1798, 1824, 1826, 1844, 1850, 1859, 1880, 1906, 1918, 2033, 2064, 2073, 2084, 2086, 2091, 2107, 2116, 2134, 2142, 2148, 2151, 2159, 2160, 2170, 2175, 2201, 2208, 2209, 2213, 2226, 2246, 2248, 2270, 2273, 2294, 2298, 2315, 2321, 2331, 2332, 2342, 2344, 2374]. based [3, 71, 91, 97, 105, 149, 177, 194, 236, 304, 315, 334, 345, 383, 493, 544, 557, 590, 591, 626, 669, 690, 697, 699, 731, 741, 743, 864, 875, 884, 903, 958, 1039, 1070, 1090, 1094, 1116, 1149, 1175, 1217, 1290, 1313, 1351, 1370, 1466, 1470, 1487, 1488, 1499, 1513, 1539, 1540, 1542, 1551, 1564, 1579, 1651, 1662, 1663, 1685, 1708, 1734, 1736, 1751, 1758, 1794, 1803, 1816, 1817, 1821, 1851, 1857, 1860, 1866, 1884, 1896, 1932, 1943, 1944, 1964, 1971, 1973, 1976, 1988, 2021, 2051, 2066, 2077, 2114, 2137, 2139, 2144, 2162, 2205, 2220, 2259, 2301, 2306, 2325, 2328, 2334, 2343, 2349, 2351, 2365, 2372, 2375]. based [28, 32, 37, 60, 98, 138, 151, 159, 170, 176, 182, 191, 216, 314, 316, 331, 366, 389, 405, 427, 501, 502, 507, 513, 580, 651, 729, 815, 847, 869, 878, 934, 967, 1017, 1067, 1078, 1099, 1154, 1166, 1177, 1219, 1262, 1282, 1318, 1335, 1337, 1368, 1434, 1471, 1472, 1506, 1533, 1583, 1584, 1592, 1616, 1622, 1636, 1672, 1702, 1703, 1713, 1721, 1746, 1750, 1791, 1800, 1861, 1877, 1899, 1905, 1928, 1939, 1959, 1978, 1985, 2005, 2024, 2046, 2093, 2103, 2133, 2145, 2152, 2161, 2183, 2199, 2200, 2214, 2216, 2223, 2232, 2234, 2253, 2263, 2285, 2289, 2313, 2340, 2341]. based [61, 96, 152, 185, 209, 231, 237, 294, 346, 358, 373, 379, 401, 416, 420, 423, 430, 484, 491, 514, 523, 529, 636, 682, 691, 737, 780, 807, 823, 845, 856, 862, 876, 914, 951, 976, 977, 980, 1033, 1061, 1102, 1168, 1273, 1274, 1296, 1306, 1326, 1329, 1333, 1422, 1435, 1467, 1536, 1554, 1562, 1563, 1597, 1621, 1674, 1710, 1730, 1761, 1764, 1787, 1819, 1846, 1894, 1917, 1921, 1954, 1970, 1975, 2014, 2020, 2063, 2126, 2146, 2178, 2229, 2260, 2264, 2265, 2275, 2337, 2363]. baseline [1963]. basic [51, 1933]. basis [1313, 1820]. bat [1900, 2111]. batch [9, 126, 1360, 1440, 2366]. Batching [1240]. Bayes [780]. Bayesian [573, 640, 1531, 1887, 2261]. BBISS [1273]. BCI [2156]. be [380, 386, 1617, 2237]. beamforming [2369]. bearings [146]. bearings-only [146]. beast [17]. bed [1298]. beds [1835]. bee [1561, 1832, 2114, 2236, 2253, 2327, 2349]. bees [1048]. BeesyCluster [421]. behavior [692, 1146, 1313, 1540, 1550, 1564, 1748, 1811, 1915, 2093]. behavioral [901, 2249]. behaviors [284, 885, 1147, 2094]. belief [1320]. Benchmark [135, 860, 1275, 1372, 2066]. Benchmarking [733, 1383, 1784, 2346]. benefits [802]. benign [416]. Better [1194]. between [101, 114, 338, 539, 735, 814, 1218, 1288, 1371, 1445, 1692, 2286]. Betweenness [104, 2342]. BFCA [1400]. Bi [473, 718]. bi-blending [473]. Bi-level [718]. Biased [45, 150]. BiCG [1589]. Biclique [906]. BICM [930]. Biconjugate [320, 474]. bid [1636, 1822]. bidding [885, 1608]. bidomain [2246]. bidomain-based [2246]. Big [973, 1060, 1140, 1278, 1281, 1327, 1328, 1339, 1341, 1352, 1372, 1376, 1483, 1516, 1554, 1555, 1558, 1572, 1584, 1620, 1625, 1629, 1631, 1633, 1698, 1877, 1880, 1892, 2002, 2007, 2009, 2014, 2051, 2097, 2132, 2153, 2164, 2195, 2258]. Big-Data [1376]. bike [2143]. bike-sharing [2143]. bilinear [91]. billing [664]. billions [1296]. bin [1666]. binaries [221]. Binary [328, 461, 670, 702, 993, 1116, 1224, 1365, 1810, 2008, 2348]. binocular [1734]. binomial [1067]. bio [1030, 2144]. bio-inspired [2144]. bio-medical [1030]. bioinformatics [2154, 2349]. biological [457, 602, 1342, 1484]. biology [2158]. biometric [1817, 1874]. biometric-based [1817]. biometrics [430, 1804]. biometrics-based [430, 1804]. bionics [153]. biosensor [1218]. biosignals [1350]. biosurveillance [542]. birth [1652]. birth-death [1652]. bit [554]. bit-rate [554]. bitmap [774]. bitplane [1320]. bits [915, 2177]. bitwise [2351, 2372]. blackhole [516]. Blacklist [1420]. blades [51]. BLAS [933, 1173, 2324]. BLASTP [1624]. blending [473]. blind [149, 989]. block [175, 317, 371, 596, 608, 663, 689, 754, 958, 1149, 1269, 1306, 1335, 1363, 1477, 1498, 2032, 2062]. block-based [958]. block-oriented [1269]. Block-Toeplitz [596]. block-tridiagonal [1498]. Blockchain [1470, 1885, 2142, 2200, 2333]. Blockchain-based [1470, 2142, 2200]. blockchain-enabled [1885]. blocking [1018, 1035, 1101, 1380, 1899, 2009]. blocks [1999]. Bloom [429]. Blue [1001]. BMCS [1394]. BML [1933]. BOCR [719]. body [571, 1219, 1651, 2219]. Boltzmann [1800]. BonaFide [806]. Boolean [167]. boost [1576]. booting [1978]. both [282]. botnets [2350]. bottleneck [847, 1134]. bound [182, 541, 1318, 1440, 2344, 2366]. boundaries [359]. boundary [478, 951, 1904, 2161]. bounded [1142]. bounds [1716, 1828]. Box [618, 1817]. Box-counting [618]. boxes [138]. BPLG [1394]. BPLG-BMCS [1394]. branch [182, 1287, 1318, 1440, 2366]. branch-and-bound [1440, 2366]. breadth [1930]. breadth-first [1930]. breakthroughs [927]. bricking [2]. bridge [338, 920]. bridges [689]. Bridging [553]. Brillouin [1793]. Broadcast [227, 1080, 1178, 1185, 1194, 1273, 1274, 2326]. Broadcast-based [1080, 1273, 1274]. broadcasting [48, 346, 347, 405, 1112, 2326]. Broker [768]. Brokerage [1797]. Brokerage-based [1797]. brokers [33]. Brownian [600]. BSP [733, 858]. BSS [958]. budget [1390, 2258]. Buffer [54, 566, 803, 1576, 1661, 1674, 2209]. buffered [338]. Buffering [812]. bufferless [338, 781]. builder [2105]. Building [63, 86, 270, 797, 1326, 1428, 1849]. building/shadow [1428]. bulk [402, 1605, 1969]. burst [854, 1599]. bursty [9]. Burton [216]. Bus [1224, 1243, 1455]. business [976, 1854]. BXI [1377]. bypass [1166, 2307]. bytes [735]. Byzantine [1059, 2333].

C

[600, 806, 1598, 1634, 1657, 1911, 2183, 2349]. C-based [2183]. C-DTB-CHR [1657]. C/C [1598]. C2OF2N [1758]. CA [128]. CA-based [128]. Cache [24, 355, 388, 398, 505, 584, 691, 791, 853, 978, 996, 1125, 1172, 1234, 1242, 1259, 1521, 1569, 1603, 1617, 1720, 1723, 1742, 1798, 1978, 1999, 2006, 2121, 2128, 2235, 2245, 2285, 2296, 2303]. cache-coherent [1617]. Cache-conscious [1569]. caches [755, 2177, 2306]. caching [106, 205, 467, 1611]. CAE [236]. Caffe [2033]. CAIM [837]. calculate [1728]. calculating [356]. Calculation [418, 1111, 1324, 1412, 1633, 1790, 2008]. calculations [1759]. calculus [1714]. calibrated [2297]. call [1860]. Camel [1109]. camera [496, 1921]. camera-based [1921]. cameras [1883, 2030]. camouflage [513, 1881]. campus [2245]. can [380, 2237]. cancelable [1711, 1874]. cancellation [2321]. cancellation-based [2321]. candy [1846]. capabilities [1400]. Capability [1214, 1224, 1337, 2328]. Capability-based [1337]. capable [342]. capacity [18, 662, 722, 1469, 1787, 1895, 2306]. capping [703]. CapsNet [1989]. capture [673, 1139]. capturing [1698]. Caputo [813]. CAR [610, 931]. carbon [2356]. card [674]. Cardiac [253, 1437, 2246]. cards [430]. CARE [768, 1351, 1610, 2084]. Carlo [140, 160, 1159]. Cartesian [111, 1452]. carving [1155]. CAS [500]. cascaded [1917]. cascades [1067]. Case [128, 192, 295, 364, 370, 408, 523, 540, 549, 643, 815, 933, 1046, 1085, 1329, 1352, 1491, 1510, 1567, 1648, 1722, 1781, 1782, 1851, 1888, 1900, 1957, 2057, 2156, 2258, 2281]. case-aware [1510]. case-based [523]. Cassandra [1618, 2338]. catastrophe [491]. causal [1768]. cause [2266]. cause-related [2266]. caused [1303]. CAVDO [1858]. CBCH [2238]. CCA [1449]. CCBS [1599]. CCD [2030]. CCDN [2103]. CCFinder [1629]. CCN [1677]. ccNUMA [2354]. CDABC [2327]. CDCSS [1884]. CDMA [883, 1224]. CDN [2103]. CDNs [1578]. cell [51, 258, 386, 603, 612, 1861, 1936, 2255, 2369]. cell-based [51, 1936]. CellStats [51].

Cellular

[129, 138, 461, 462, 492, 574, 575, 578–580, 819, 820, 827, 1268, 1313, 1365, 1444, 1445, 1741, 1865, 1869, 2251, 2255, 2308, 2332, 2356, 2357]. center [150, 327, 684, 832, 869, 980, 1405, 1485, 1491, 1681, 1962, 1992, 2290]. center-biased [150]. centered [800]. centerline [1989]. centers [315, 407, 450, 507, 954, 981, 985, 987, 1108, 1241, 1248, 1506, 1511, 1516, 1597, 1747, 1782, 1829, 1896, 2034, 2099, 2100, 2117, 2291, 2309, 2320, 2341]. Central [364]. centrality [104, 2058]. centralized [1657]. centre [1283, 1410]. centric [1039, 1392, 1677, 1681, 1922, 1992, 2225, 2245]. centroid [187, 847]. centroid-based [847]. certificateless [92]. certified [908]. CF [1976]. CF-CloudOrch [1976]. CFD [920, 1526]. CG [1614]. Chain [649, 1487, 1506, 1581, 1680, 1877]. Chain-based [1506, 1877]. Chained [34, 102]. ChainMail [1154]. chains [371, 720, 736, 1366, 1716]. challenge [1393]. challenges [521, 732, 924, 1198, 1276, 1278, 1352, 1473, 1534, 1876, 1993, 2211, 2244]. Change [1515]. Change-driven [1515]. changes [1541]. changing [706]. channel [406, 438, 835, 988, 1027, 1113, 1135, 1166, 1187, 1409, 1462, 2087]. channel-based [1166]. channel-recommendation [835]. channels [1643, 2326]. chaos [1990, 2014, 2300]. chaotic [227, 1171, 1283, 1817, 1857, 2327]. character [2089]. characteristic [2334]. characteristic-based [2334]. characteristics [659, 1022, 1958]. Characterization [367, 806, 950, 1054, 1057, 1509]. Characterizing [51, 832, 1383]. chassis [954]. Chebyshev [1857, 2214]. check [876, 1554, 1697]. check-ins [1554]. checking [388, 490, 817, 858]. checkpoint [617, 940, 1935, 2023, 2295]. checkpoint-based [940]. checkpoint/restart [617, 2295]. checkpointing [139, 1401, 2192]. checkpointing/restart [2192]. chemistry [1933]. chicken [1764]. children [917]. China [2021]. Chinese [1546]. chip [78, 87, 226, 277, 311, 338, 357, 414, 634, 698, 753, 781, 863, 899, 923, 996, 1033, 1112, 1113, 1124, 1136, 1151, 1157, 1166, 1172, 1180, 1197, 1224, 1236, 1250, 1271, 1306, 1318, 1373, 1490, 1510, 1568, 1664, 1665, 1778, 1798, 1819, 1853, 1863, 2020, 2307, 2317]. chips [1871]. choices [263]. Cholesky [370, 1404]. choose [797]. chooser [631]. choosing [2012]. choreographies [420]. CHR [1657]. chromatic [2221]. chunked [1006]. Churn [656, 1676]. Churn-aware [656, 1676]. Cilk [16, 1713]. cipher [663, 1335]. circuit [225, 1122, 1701, 1852]. circuit-packet [1122]. circuits [372, 1042, 1497, 2339]. circular [1732, 1733]. circumstances [1522]. circumvention [1881]. cities [2330]. citizens [1272, 2084]. city [1083]. CL [2140]. clairvoyant [1756, 1995]. ClaMPP [2139]. class [1589, 2089]. classic [1446]. Classification [50, 460, 461, 594, 780, 823, 859, 1342, 1499, 1575, 1803, 1851, 1873, 1887, 1967, 2012, 2033, 2068, 2133, 2139, 2156, 2247, 2249, 2254, 2261, 2297, 2337]. classification-based [1851]. classified [283, 1297]. classifiers [503]. classifying [1000, 1350]. classroom [1921]. clauses [1024]. clean [1617]. client [844, 961, 1043, 1177, 1291, 1923]. client-server [844, 1177, 1923]. clients [1918]. climate [1831]. clinical [1606, 1802]. cliques [1698, 2225]. clock [1093, 1110, 2197]. clone [1839]. Cloning [765]. Clos [375]. closure [1399]. clothed [2219]. Cloud [302, 303, 384, 419, 441, 455, 494, 519, 534, 536, 549, 563, 663, 679, 680, 684, 725, 740, 804, 815, 832, 855, 878, 892, 897, 924, 979, 995, 1043, 1051, 1052, 1069, 1097, 1108, 1188, 1212, 1226, 1241, 1253, 1267, 1288, 1300, 1301, 1334, 1358, 1430, 1449, 1460, 1511, 1528, 1534, 1537, 1544, 1545, 1586, 1599, 1602, 1609, 1612, 1613, 1619, 1633, 1637, 1642, 1656, 1673, 1684, 1691, 1695, 1781, 1788, 1796, 1797, 1836, 1846, 1890, 1893, 1902, 1905, 1914, 1918, 1922, 1926, 1929, 1946, 1949, 1951, 1962, 1982, 2120, 2141, 2149, 2159, 2160, 2170, 2240, 2248, 2292, 2304, 2323, 2371]. cloud [220, 234, 278, 304, 306, 366, 395, 407, 444, 542, 548, 550, 561, 588, 705, 772, 777, 842, 869, 875, 890, 958, 965, 970, 1017, 1034, 1056, 1062, 1179, 1201, 1204, 1216, 1302, 1314, 1390, 1466, 1467, 1473, 1476, 1481, 1488, 1506, 1522, 1533, 1539, 1551, 1553, 1564, 1575, 1590, 1616, 1628, 1649, 1662, 1682, 1688, 1690, 1700, 1701, 1720, 1743, 1782, 1862, 1866, 1885, 1895, 1896, 1901, 1913, 1916, 1925, 1942, 1944, 1964, 1974, 1976, 1977, 1984, 1994, 2022, 2099, 2100, 2102, 2103, 2114, 2128, 2130, 2131, 2139, 2162, 2171, 2173, 2175, 2205, 2206, 2220, 2244, 2262, 2289, 2301, 2309, 2314, 2335, 2375]. cloud [21, 76, 307, 416, 431, 445, 468, 507, 540, 562, 632, 661, 709, 714, 718, 768, 787, 807, 846, 849, 852, 862, 881, 975, 983, 1003, 1013, 1061, 1119, 1131, 1133, 1150, 1275, 1282, 1323, 1516, 1524, 1531, 1562, 1608, 1622, 1655, 1689, 1745, 1764, 1767, 1783, 1870, 1924, 1950, 1978, 1980, 1985, 2063, 2117, 2127, 2154, 2202, 2263, 2320]. Cloud-assisted [979, 1545]. Cloud-based [416, 815, 1017, 1051, 1488, 1656, 1695, 1846, 2139, 2159, 2160]. cloud-empowered [1537]. cloud-oriented [303]. cloudification [2166]. Cloudlet [1476, 1494, 1590, 1833]. cloudlets [1132]. CloudOrch [1976]. CloudRPS [1551]. Clouds [308, 676, 754, 1073, 1132, 1196, 1228, 1251, 1252, 1295, 1323, 1347, 1450, 1486, 1538, 1552, 1559, 1600, 1636, 1773, 1780, 1931, 1938, 1993, 1995, 2129, 2203, 2315]. CloudSim [2117]. CloudTaint [983]. CLS [2140]. CLUS GPU [1624]. CLUS GPU-BLASTP [1624]. Cluster [3, 54, 60, 81, 85, 116, 167, 184, 262, 335, 365, 411, 418, 669, 688, 710, 782, 818, 907, 910, 1052, 1064, 1189, 1204, 1318, 1425, 1442, 1499, 1505, 1563, 1624, 1657, 1729, 1792, 1831, 1884, 1968, 2126, 2151, 2246, 2260, 2343]. cluster-based [3, 1052, 1318, 1563, 1884, 2126]. clustered [653, 1089, 1237, 1603]. Clustering [30, 50, 146, 203, 314, 331, 741, 752, 770, 847, 867, 918, 973, 1102, 1144, 1168, 1217, 1299, 1410, 1504, 1510, 1627, 1629, 1632, 1679, 1721, 1755, 1803, 1858, 1919, 1967, 2104, 2131, 2169, 2199, 2224, 2236, 2238, 2276, 2312, 2327]. clustering-based [1168, 1410, 2199, 2238]. Clusters [90, 122, 186, 268, 272, 280, 291, 316, 326, 393, 441, 457, 604, 619, 641, 644, 717, 794, 834, 880, 929, 1040, 1063, 1299, 1311, 1383, 1434, 1439, 1736, 1756, 1987, 2156, 2192, 2295]. CMP [124]. CMPs [336, 505, 1125, 1276]. CNN [2033]. CNN-based [2033]. Co [72, 184, 258, 1054, 1252, 1259, 1391, 1787, 2037]. co-allocation [72, 184]. co-execution [2037]. co-located [1787]. co-location [1252, 1391]. Co-occurrence [258]. co-runner [1259]. co-scheduling [1054]. coalition [1144]. coarse [313]. coarse-grain [313]. code [179, 212, 261, 1173, 1247, 1336, 1400, 1758, 1839, 1910, 1911, 2056, 2302, 2343, 2347]. code-based [2347]. codecs [1429]. codes [485, 1153, 1238, 1981]. coding [152, 653, 810, 1269, 1464, 2039, 2215]. CoDiP2P [599]. coefficient [1629, 1803]. Cognitive [245, 1143, 1815, 2014, 2116]. cognitive/reactive [245]. coherence [388, 584, 634, 791, 1023, 1238, 1603]. coherent [410, 1617]. coins [2348]. cold [2011]. collaborative [610, 1308, 1312, 1809, 1943, 1986, 2108, 2205, 2375]. collaborative-based [2205, 2375]. collage [1824]. collage-based [1824]. collection [241, 981, 1382, 2151]. collective [272, 335, 1109, 1785]. collective-aware [1109]. collectives [90]. collector [991]. collision [888]. collisions [2199]. colloidal [1421]. Colony [840, 1561, 1644, 1832, 1954, 2114, 2236, 2248, 2253, 2265, 2327, 2349]. colony-based [2349]. color [147, 1029]. Coloring [1219]. Coloring-based [1219]. colorization [1286]. column [1717]. column-wise [1717]. combination [523, 668, 1332, 2010]. combinatorial [376, 957]. Combined [847, 866, 896, 1114, 1807, 2268, 2269, 2287]. Combining [41, 102, 249, 460, 859, 1305, 1423, 1549, 1905]. Command [2232]. Comment [2373]. Comments [1177, 1486]. commerce [95]. commercial [212, 1726]. commits [1191]. common [1386, 1841]. communicating [2057]. Communication [51, 67, 96, 98, 119, 124, 226, 255, 266, 267, 335, 347, 362, 453, 489, 492, 498, 522, 665, 671, 690, 736, 814, 827, 880, 889, 1021, 1113, 1189, 1198, 1221, 1263, 1273, 1289, 1318, 1333, 1453, 1557, 1574, 1643, 1683, 1705, 1727, 2029, 2074, 2268, 2269, 2313, 2343]. communication-avoiding [2343]. Communication-aware [1318]. communication-based [1189]. communication-efficient [665]. Communication-free [267]. Communication-resource-aware [522]. communications [7, 62, 89, 94, 268, 271, 297, 449, 532, 597, 644, 650, 834, 2087, 2211]. communities [570, 708]. Community [306, 570, 1199, 1223, 1906, 2125, 2224, 2271]. Community-based [1199]. Compact [286, 1378, 1959, 2339]. compaction [1604]. CompactMatrix [706]. Comparative [216, 236, 256, 324, 840, 918, 925, 931, 1102, 1829, 2286]. Comparing [2366]. Comparison [168, 206, 233, 573, 920, 1288, 1388, 1427, 1484, 1771, 1941]. comparison-free [1941]. compatible [895]. compensated [148, 150]. competitive [770, 1640]. CoMPI [207]. compiler [143]. compilers [212, 267]. complement [828]. complete [993]. completion [1064]. complex [631, 708, 734, 741, 980, 1539, 1853, 1904]. complexity [597, 747, 1231, 2344, 2369]. complying [824]. Component [285, 682, 693, 901, 1531, 1547, 2284]. component-based [2284]. components [1357, 1365]. Composable [308]. composite [884, 1648]. composition [107, 425, 901, 1205, 1481, 1901, 2120, 2253]. Compositional [259, 1010]. compound [641]. comprehensive [1130, 1473, 1678, 1777, 2104, 2163, 2211]. compressed [1769]. Compressing [1570]. compression [1092, 1149, 1388, 1723]. Compressive [1694]. computation [47, 109, 157, 183, 188, 300, 310, 319, 363, 415, 670, 852, 882, 1058, 1129, 1132, 1233, 1402, 1494, 1547, 1645, 1646, 1717, 1794, 2041, 2054, 2058, 2075, 2157, 2171]. Computational [29, 125, 131, 198, 317, 469, 476, 583, 692, 712, 721, 732, 787, 874, 919, 994, 1077, 1343, 2027, 2158]. computationally [385, 1586]. computations [77, 98, 222, 334, 386, 393, 564, 1106, 1569, 1611, 1719, 2354]. Compute [525, 1357, 1503, 2006, 2069]. compute-intensive [1357, 1503]. computer [110, 323, 613, 2080, 2085, 2221]. computers [156, 262, 276, 527]. Computing [42, 196, 210, 273, 361, 375, 384, 392, 439, 449, 469, 472, 494, 519, 536, 563, 566, 585, 587, 589, 595, 617, 633, 663, 679, 680, 684, 705, 710, 739, 740, 747, 772, 804, 814, 846, 866, 877, 897, 928, 938, 949, 995, 1002, 1004, 1016, 1040, 1041, 1043, 1050, 1087, 1097, 1159, 1190, 1212, 1216, 1226, 1288, 1356, 1364, 1393, 1413, 1462, 1481, 1511, 1518, 1522, 1528, 1534, 1537, 1560, 1566, 1619, 1633, 1684, 1691, 1728, 1772, 1788, 1792, 1796, 1797, 1824, 1905, 1914, 1929, 1951, 1972, 2042, 2045, 2132, 2142, 2149, 2167, 2194, 2201, 2220, 2270, 2292, 2304, 2344]. computing [7, 32, 37, 46, 82, 89, 104, 108, 155, 171, 220, 234, 278, 304, 316, 331, 352, 387, 395, 407, 436, 445, 468, 477, 507, 518, 537, 540, 548, 551, 553, 555, 565, 590, 599, 632, 681, 687, 714, 769, 787, 805, 842, 869, 875, 878, 927, 937, 965, 970, 975, 1003, 1007, 1021, 1025, 1116, 1131, 1275, 1302, 1307, 1398, 1436, 1457, 1473, 1476, 1514, 1539, 1564, 1575, 1590, 1605, 1639, 1662, 1682, 1688, 1690, 1700, 1710, 1712, 1784, 1924, 1925, 1927, 1942, 1950, 1977, 1984, 1985, 1994, 2027, 2080, 2097, 2102, 2114, 2130, 2153, 2162, 2173, 2205, 2206, 2314, 2328, 2335, 2352, 2375]. computing [61, 76, 85, 86, 96, 116, 209, 293, 294, 302, 426, 434, 490, 508, 556, 709, 773, 807, 825, 849, 852, 912, 924, 974, 1013, 1061, 1119, 1133, 1152, 1168, 1243, 1323, 1455, 1498, 1524, 1599, 1620, 1634, 1689, 1764, 1783, 2004, 2010, 2127, 2146, 2154, 2193]. computing-based [878, 1985]. concealed [1881]. concealment [152, 970]. concealments [1215]. concept [1175]. Concepts [1272]. concerns [307]. concurrency [16, 247, 790, 1182, 1396, 2057]. Concurrent [437, 837, 841, 1776, 1827]. Condition [140, 706, 1010, 1386]. Conditional [442, 615, 1239, 2089]. conduit [1455]. Confidential [1889]. configurable [128, 1985]. configuration [299, 896, 1114]. configurations [1398]. CONFIIT [42]. confirmation [1252]. conflict [1167, 1332]. conformal [1940]. conforming [891]. confronting [2350]. congestion [236, 636, 781, 1045, 1080, 1151, 1236, 1928]. congestion-aware [1928]. Conjugate [320, 952]. conjunction [678]. connect [404]. Connected [38, 47, 74, 233, 765, 857, 863, 1365, 1380, 1495, 1604, 1623]. Connection [736, 912]. Connectivity [25, 154, 194, 547, 559, 711, 745, 1545, 1585]. conquer [1286]. conscious [1569]. consensus [999, 1059, 1164, 2319, 2348]. conservation [125, 879]. consideration [999, 1111, 2110]. considering [282, 559, 562, 692, 842, 1179, 1522, 2205, 2317, 2375]. Consistency [59, 398, 876, 1074, 2051, 2125]. consistency-based [2051]. consolidation [292, 697, 846, 954, 1073, 1241, 1450, 1612, 1613, 1747, 1919, 2019, 2099, 2100, 2323]. constant [1445, 2025]. constant-time [2025]. constrained [80, 194, 242, 434, 488, 1065, 1449, 1522, 1541, 1686, 2078]. constraint [696, 1318, 1943, 2258]. constraints [467, 764, 1390, 1823, 2182]. Constructing [85, 255, 1031, 1487, 2267].

Construction [73, 116, 534, 616, 1968, 1986, 2144]. constructive [1373]. constructs [437]. consumer [2021, 2335]. consumption [101, 206, 607, 717, 784, 921, 1064, 1359, 1520, 1612–1614, 1660, 1681, 1749, 1821, 1870, 2011, 2129, 2284, 2302, 2317, 2331]. contain [416]. container [1781, 1976]. contemporary [1883]. content [79, 97, 934, 964, 1155, 1461, 1649, 1653, 1656, 1677, 2103, 2131, 2245]. content-aware [1155, 1461]. content-based [79]. content-centric [1677, 2245]. Contention [61, 504, 644, 1054]. contention-free [644]. contents [500, 892, 1094]. Context [65, 66, 68, 269, 490, 2171, 2353]. Context-adaptive [269]. Context-aware [65, 66, 68, 2171, 2353]. contextual [1454]. contiguous [1285, 1604]. Continuous [18, 322, 363, 751, 1360, 1902, 2048, 2141, 2199, 2257, 2259]. Contract [679]. Control [49, 60, 71, 78, 233, 236, 288, 435, 578, 737, 740, 761, 777, 781, 790, 810, 852, 890, 944, 962, 1080, 1143, 1145, 1182, 1220, 1226, 1239, 1331, 1529, 1693, 1707, 1877, 1895, 2069, 2095, 2232, 2235]. controlled [926]. controlled-grained [926]. controller [65, 78, 369, 520, 1293]. controls [1010]. Convergence [52, 344, 540, 913, 1213, 1960, 2002]. conversion [1553, 1685]. Convert [404]. Convert-and-Deliver [404]. Converting [915]. convex [1368, 1904, 2109, 2238]. conveying [1482]. convolution [190, 1152]. convolutional [1562, 1990, 2013, 2134]. Cooling [1034]. cooperation [23, 427, 464, 1932, 2080, 2085]. Cooperative [119, 205, 463, 701, 728, 810, 1143, 1145, 1323, 1496, 1561, 1758, 1884, 2079, 2096, 2109, 2148]. coordinate [1521]. Coordinated [36, 305, 1482, 2023]. coprocessors [1443, 1523, 1996]. Copy [484]. core [56, 87, 90, 130, 164, 193, 223, 272, 295, 335, 410, 428, 432, 458, 528, 587, 618, 634, 637, 643, 644, 736, 764, 792, 795, 798, 801, 808, 823, 865, 867, 898, 921, 925, 928, 945, 1023, 1099, 1111, 1115, 1118, 1153, 1197, 1279, 1403, 1427, 1490, 1526, 1532, 1566, 1587, 1648, 1671, 1731, 1759, 1831, 1969, 2015, 2040, 2049, 2076, 2182, 2209, 2303, 2365]. cores [1583]. Corner [920]. corpus [2105]. Correcting [885]. Correction [1247, 1733, 1751, 1777, 1802, 2100, 2123, 2160, 2185, 2251, 2269, 2281, 2351]. corrector [2345]. correlated [758, 1867, 2011]. correlating [2302]. correlation [387, 1645, 1646, 2113]. Corrosion [580]. Corrosion-passivation [580]. coscheduling [436]. Cost [106, 234, 262, 313, 816, 1034, 1043, 1045, 1166, 1253, 1257, 1258, 1553, 1643, 1678, 1728, 1738, 1773, 1777, 1807, 1813, 1814, 1866, 1891, 1938, 1995, 2101, 2165, 2215, 2216, 2304, 2330, 2339]. Cost-aware [106, 1253]. Cost-effective [262, 1643, 1807]. Cost-efficient [1045, 1938, 2304]. cost/performance [1728]. costs [335]. counter [2091]. countermeasure [416, 1216, 2141]. countermeasures [689, 2211]. counters [591]. counting [429, 618]. country [2282]. counts [797]. coupled [257, 305, 887, 1668, 1775]. coupling [1831]. covariance [442]. cover [630, 699, 873, 1998]. coverable [1120]. Coverage [25, 559, 771, 806, 1202, 1640, 1739]. covering [392, 1041, 1736]. covers [2250]. CPS [624]. CPU [472, 541, 618, 640, 641, 794, 948, 1016, 1019, 1020, 1058, 1246, 1402, 1498, 1798, 1830, 2015, 2037, 2044, 2079, 2113, 2190, 2198, 2283, 2315, 2323]. CPU-intensive [2323]. CPU-oriented [2283]. CPU/GPU [640, 641, 1016]. CPU/GPU-based [640]. CPU2006 [212]. CPUs [310, 860, 867, 928, 1256, 1752, 1941]. CRAFT [2266]. crash [199]. crash-recovery [199]. crawling [2282]. created [1216]. created-accessed-modified [1216]. creation [1175]. credential [705, 966]. credit [2021]. CRFF.GP [2175]. criteria [331, 647, 950, 1168, 1241]. criteria-based [331]. critical [693, 889, 1461, 1872, 1874, 1878, 1879]. CRM [1695]. Cross [68, 238, 248, 404, 618, 622, 1348, 1609, 1700, 1741, 2262, 2355]. cross-cloud [2262]. cross-connect [404]. Cross-group [1609]. Cross-layer [68, 238, 622, 1700]. cross-match [1348]. cross-platform [618]. cross-stencil [2355]. cross-wiring [1741]. crossbar [1018]. crossed [1120, 1766]. crossing [1989]. crowd [169, 452, 484, 1313]. crowdphotographing [2204]. crowdsensing [1927]. crowdsourced [1961]. crowdsourcing [1973]. Cryptanalysis [592, 663, 906, 960, 1465]. cryptographic [2141]. cryptographically [1171]. Cryptography [98, 192, 903, 905, 960, 1231, 1364, 1523, 1708, 1751, 1923, 1944, 2036]. cryptosystem [323, 430]. cryptosystems [1675]. CS [1488]. CSF [1888, 2281]. cSketch [1698]. CTL* [858]. CTU [1408]. CU [712, 1464]. cube [310, 353, 1623, 2165]. cube-connected [1623]. cubes [75, 330, 616, 993, 1031, 1120, 1766]. Cubic [34, 838]. CUBLAS [763]. CUBLAS-based [763]. Cuckoo [1266, 2111, 2254]. CUDA [163, 164, 222, 320, 356, 454, 525, 608, 724, 751, 1046, 1103, 1167, 1173, 1362, 1634, 1659, 1752, 2069, 2316]. CUDA-C [1634]. cultural [2309]. cupSODA [821]. curation [1338]. current [1759]. curvature [144]. Curve [430, 905, 960, 1354, 1923, 1944, 2036]. curved [2046]. custom [1363]. customer [99]. customizable [986]. customization [275]. Customized [786]. CWBound [1904]. Cyber [620, 625, 1144, 1469, 1807, 1836, 1876, 1881, 1884]. cyber-physical [1144, 1807]. Cybercrime [1216]. cyberinfrastructure [1210]. cybersecurity [1343, 1856, 1885]. cycles [75, 318, 1478, 1623]. cyclic [1006]. Cyclops64 [104].

D [151, 191, 231, 277, 294, 368, 397, 470, 500, 552, 587, 698, 755, 827, 894, 923, 952, 953, 1032, 1076, 1106, 1126, 1229, 1441, 1513, 1526, 1532, 1549, 1573, 1614, 1668, 1720, 1729, 1805, 1806, 1871, 1928, 1946, 2147, 2203, 2208, 2221, 2246, 2307, 2355, 2359, 2360]. D-CAS [500]. D-network-on-chip [698]. D-NoC [1532]. D-NUCA [755]. D-stacked [1720]. D-Wave [2221]. DaaS [2106]. DADE [2212]. DAG [13, 1253, 2004, 2078]. DAGMap [13]. DAI [71]. Daily [64, 659, 1272]. damage [1694]. Damaged [666, 1958]. dangerous [1564]. DAR [248]. dark [1157, 1174]. DASC [1023]. DASC-DIR [1023]. Data [17, 114, 204, 228, 267, 280, 281, 283, 290, 315, 334, 409, 467, 475, 525, 623, 666, 668, 678, 684, 688, 727, 728, 745, 770, 811, 826, 832, 843, 877, 956, 958, 968, 973, 1051, 1065, 1088, 1108, 1141, 1215, 1218, 1241, 1242, 1248, 1278, 1284, 1341, 1352, 1355, 1382, 1388, 1410, 1441, 1481, 1485, 1519, 1528, 1558, 1582, 1594, 1610, 1617, 1625, 1633, 1694, 1723, 1743, 1799, 1808, 1817, 1844, 1880, 1882, 1910, 1922, 1962, 1991, 2002, 2009, 2011, 2015, 2034, 2106, 2115, 2132, 2147, 2151, 2164, 2182, 2194, 2217, 2236, 2258, 2270, 2272, 2290, 2294, 2302, 2338, 2346, 2365, 2368, 2373]. data [18, 22, 32, 83, 124, 354, 366, 402, 407, 487, 507, 550, 626, 642, 646, 676, 741, 750, 758, 767, 774, 868, 869, 954, 959, 985, 987, 1057, 1060, 1064, 1104, 1106, 1116, 1140, 1144, 1183, 1227, 1230, 1264, 1274, 1327, 1328, 1339, 1359, 1405, 1418, 1433, 1434, 1447, 1453, 1487, 1491, 1506, 1516, 1565, 1569, 1572, 1584, 1611, 1647, 1653, 1663, 1676, 1677, 1681, 1690, 1697, 1702, 1719, 1747, 1754, 1780, 1782, 1829, 1877, 1881, 1892, 1896, 1899, 1973, 1983, 1984, 2007, 2037, 2074, 2097, 2099, 2100, 2117, 2133, 2153, 2178, 2195, 2207, 2259, 2279, 2291, 2309, 2326, 2328, 2341]. data [59, 72, 88, 115, 254, 327, 358, 373, 450, 460, 569, 675, 736, 776, 918, 980, 981, 1005, 1068, 1128, 1244, 1281, 1326, 1372, 1376, 1554, 1555, 1597, 1620, 1631, 1638, 1689, 1725, 1812, 1931, 1981, 1992, 2014, 2051, 2146, 2196, 2212, 2264, 2320, 2363]. data-acquisition [1382]. data-aware [626]. data-centric [1922]. data-collection [1382]. data-dependent [254]. data-flow [1719]. Data-independent [2368]. data-intensive [1594, 1611, 1780]. data-parallel [826, 1418, 2037, 2074]. data-privatization [642]. Data-type [1723]. database [203, 455, 688, 843, 969, 1530, 1618, 1621, 1672, 2106]. database-as-a-service [2106]. databases [1029]. datacenter [1134, 1753]. datacenters [274, 357, 496, 1008, 1225, 1600, 1602]. dataflow [1100]. datapath [219]. dataset [659, 1102]. datasets [2, 415, 867, 1342, 1721, 1769]. DAWA [1650]. DBSCAN [1987]. DC2 [981]. DCSACA [1943]. DDF [1508]. DDoS [746, 1887]. de-interlace [150]. de-synchronization [1687]. Deadline [434, 1069, 1390, 1449, 1522, 1541]. deadline-constrained [1449, 1522, 1541]. deadlines [779]. deadlock [230, 248, 1135]. deadlock-free [230, 1135]. deal [1527]. Dealing [1541]. deallocator [1935]. death [1652]. debugging [1222]. Decentralized [52, 490, 783, 944, 1496, 2200, 2224]. decision [50, 112, 647, 1269, 1540, 1782, 1815, 2034, 2088, 2247]. decision-making [647, 1815]. declarations [916]. Decluster [980]. Decoder [1107, 2314, 2321]. decoding [175, 1920]. decomposition [56, 376, 823, 1569, 2091, 2226]. decomposition-based [823]. decompression [1601]. deconfliction [1167]. Decreasing [2317]. decycling [838]. dedicated [402, 1701]. deduplication [1743]. Deep [134, 1557, 1748, 1792, 1959, 1961, 2152, 2196, 2260]. default [2021]. defend [1095]. Defending [12, 1650]. defense [746]. defined [1107, 1556, 2184, 2185, 2216, 2222, 2287]. definition [1839]. deflection [103, 746, 1113]. deflection-routed [1113]. deformable [762]. deforming [1154]. defragmentation [1001]. degradation [2011]. Degree [488, 527, 884, 1724, 2072]. Delay [69, 194, 614, 638, 696, 810, 856, 962, 1089, 1142, 1653, 1823, 2128, 2245, 2329]. delay- [69]. Delay-bounded [1142]. delay-efficient [1653]. delay-tolerant [810, 856, 2329]. delayed [1735]. delegation [982, 1788, 1994]. deleted [902]. deletion [358, 373, 1896]. Deliver [404]. delivering [929]. delivery [556, 671, 754, 1126, 1151, 1461, 1653, 1656, 1671, 1881, 1981, 2103]. delta [103, 1445]. demand [48, 227, 426, 550, 683, 1791, 1900, 1999, 2081]. demand-driven [426]. Dempster [1730]. Deniability [738]. denial [1062, 2222]. denoising [144, 2013, 2017]. dense [119, 156, 499, 763, 1076, 1389, 1396, 1932]. densest [1325]. densities [1987]. density [80, 1235, 1657]. density- [1657]. Dependability [1606, 1797, 1802]. dependable [13, 681]. dependence [1005, 1244, 1725]. dependencies [114, 2115]. dependency [154, 1111]. dependent [196, 205, 254, 286, 300, 361, 816]. deploying [1211]. Deployment [307, 468, 612, 614, 625, 2330]. depressed [2136]. depth [1816]. Deriving [800]. DES-like [1675]. descent [1521]. description [546, 561, 2164]. descriptor [1029]. descriptors [1573]. Design [63, 78, 84, 87, 90, 93, 187, 263, 297, 312, 325, 355, 467, 510, 512, 584, 614, 664, 681, 698, 700, 848, 889, 904, 963, 990, 996, 1124, 1158, 1180, 1214, 1257, 1258, 1277, 1292, 1297, 1298, 1310, 1367, 1371, 1474, 1514, 1532, 1538, 1560, 1568, 1700, 1704, 1734, 1753, 1760, 1769, 1795, 1810, 1812, 1852, 1854, 1856, 1868, 1894, 1915, 1948, 1965, 2000, 2048, 2098, 2116, 2135, 2181, 2189, 2229, 2255, 2256, 2308, 2339, 2367]. design-aided [700]. design-level [614]. designated [911]. designed [1921, 2140]. Designing [10, 21, 102, 193, 453, 542, 907, 1491, 2066, 2130, 2233]. designs [505, 1869]. desktop [156, 384, 845]. Desktop-as-a-Service [845]. destructiveness [1139]. detect [1284, 1883, 2277]. Detecting [429, 689, 1128, 1500, 2350]. Detection [50, 52, 186, 290, 387, 478, 491, 503, 513, 516, 707, 748, 766, 911, 983, 1045, 1079, 1099, 1167, 1310, 1316, 1336, 1360, 1433, 1468, 1473, 1496, 1540, 1557, 1610, 1616, 1634, 1643, 1730, 1750, 1768, 1803, 1839, 1842, 1890, 1904, 1914, 1917, 1929, 1937, 1958, 1988, 1989, 2090, 2161, 2196, 2212, 2249, 2252, 2260, 2279, 2349]. detection/prevention [1473]. detector [597, 1598, 2047]. detectors [729, 2348]. determination [1464, 2082]. determining [2119]. deterministic [636]. Developing [40, 240, 419, 1096, 1206, 1707, 2085]. Development [464, 472, 543, 641, 939, 971, 992, 1458, 1515, 1536, 1546, 1813, 1814, 2181, 2237]. developments [749]. deviations [1748]. Device [351, 525, 621, 754, 1096, 1175, 1218, 1392, 1454, 1955, 2067, 2112, 2137, 2214, 2237]. device-to-device [2112]. devices [65, 97, 194, 626, 763, 1156, 1314, 1371, 1419, 1465, 1470, 1548, 1654, 1953, 2056, 2149]. DFAs [581]. DFIG [2073]. diagnosability [341, 615]. diagnosis [321, 341, 616, 2208]. diagonal [1006, 1412, 1589]. diagonalization [1582]. dialogue [1860]. diamond [150]. DiamondCandy [2355]. diarization [1291]. dictation [1291]. dictionary [1103]. difference [320, 370, 668]. differencing [662, 1468]. different [217, 445, 836, 936, 1064, 1380, 1548, 1609, 1776, 2016, 2324]. Differential [134, 349, 458, 735, 937, 955, 1445, 2207, 2345]. Diffie [736]. diffusion [24, 658, 813, 1297, 1444, 1722]. DigHR [1768]. digit [1110, 1349]. digit-multiplier [1110]. Digital [344, 500, 883, 887, 909, 989, 1818, 2032, 2053, 2062, 2200]. digraph [191]. digraph-based [191]. digraphs [750]. dimension [1535, 1828]. Dimensional [118, 218, 356, 370, 388, 412, 527, 531, 820, 1070, 1102, 1103, 1152, 1194, 1365, 1407, 1438, 1535, 1570, 1582, 1666, 1988, 2118, 2230, 2356]. Dimensionality [1309, 1828, 1859, 2137, 2349]. dimensions [265, 1811]. DIR [1023]. Direct [612, 724, 1011, 1255, 1501, 2103]. directed [1088]. Direction [149, 633]. Direction-aware [633]. directional [559, 572]. directions [1278]. directive [1908]. directive-based [1908]. directivity [1029]. directories [1238]. directory [1023, 1125, 1999]. disaster [1508, 1786]. disaster-tolerant [1508]. Disclosing [1345]. disclosure [1580, 1934]. DisCoP2P [769]. Discover [538]. Discovering [955, 1348, 1911]. discovery [306, 633, 701, 716, 783, 814, 997, 1025, 1105, 1223, 1254, 1262, 1410, 1658, 1906, 2082, 2322]. Discrete [109, 397, 470, 775, 2188, 2327]. discrete-event [2188]. discretization [837]. discretized [1735]. discriminant [1309, 1859, 2272]. disease [1499, 2208, 2247, 2349]. Disjoint [73, 102, 204, 730, 848, 1120, 1389, 1475, 1998, 2250]. disk [411, 426, 655, 867]. disk-resident [867]. disks [54]. disparity [1956]. Dispatching [211, 322, 448]. dispersion [1445, 1489]. dispersion-aware [1489]. dispersive [1793]. displacement [1423]. display [2, 707, 1824]. displays [2046]. dissemination [79, 243, 569, 623, 1144, 1690, 2338]. dissipation [1741]. distance [48, 114, 319, 405, 716, 839, 882, 1732, 1733, 2014]. distant [815]. distinguishing [217]. distortion [1462]. Distributed [44, 47, 63, 77, 119, 135, 174, 178, 203, 206, 245, 255, 263, 265, 273, 286, 292, 312, 367, 412, 421, 424, 438, 471, 483, 499, 599, 633, 653, 667, 693, 696, 718, 729, 734, 739, 743, 747, 757, 758, 795, 866, 870, 893, 899, 921, 922, 967, 969, 991, 997, 1015, 1060, 1140, 1196, 1202, 1225, 1238, 1240, 1269, 1296, 1364, 1395, 1406, 1447, 1466, 1467, 1496, 1508, 1512, 1542, 1559, 1592, 1611, 1616, 1693, 1714, 1728, 1739, 1750, 1755, 1776, 1827, 1842, 1884, 1893, 1906, 1932, 1943, 2009, 2057, 2096, 2109, 2115, 2120, 2139, 2222, 2248, 2261, 2299, 2319]. distributed [169, 215, 337, 484, 507, 713, 781, 825, 1276, 1482, 1600, 1653, 1721, 1981, 2004, 2282, 2320]. distributed-memory [471]. distribution [26, 32, 265, 377, 409, 418, 482, 656, 1084, 1092, 1194, 1230, 1397, 1492]. divergence [1075]. diversity [690]. divide [1286, 2165]. divide-and-conquer [1286]. divide-and-swap [2165]. divided [1985]. divider [2274]. divisible [339]. division [801]. DLRankSVM [1512]. DMetabench [135]. DNA [109, 323, 955, 1388, 2064]. DNA-based [109, 323]. docking [1731, 2155]. dockless [2143]. Does [1344]. domain [436, 1149, 1311, 1985, 2226]. domain-divided [1985]. domains [965, 1940]. domestic [2086]. dominance [775]. Dominating [47, 765]. Domino [2357]. don’t [1610]. dot [506, 1741, 1865, 1869, 2251, 2255, 2308]. double [418]. downlink [2096]. downloading [500]. downscaling [2026]. DP&TB [634]. DQN [2223]. DQN-based [2223]. dragon [1344]. dragonfly [1038, 1379, 1380, 1858]. DRAM [1280, 1894]. DRDDR [1284]. drive [568]. driven [22, 389, 425, 426, 455, 926, 1515, 1970, 2145]. Driving [1558, 1989, 2094]. drop [1817]. DS [1595]. DSC [1320]. DSDM [2237]. DSDM-Scrum [2237]. DSL [972]. DSP [337]. DT [1725]. DTB [1540, 1657]. DTB-IDS [1540]. Dual [51, 75, 204, 493, 796, 1012, 1103, 1113, 1168, 1250, 1668, 1737]. dual-architecture [796]. dual-based [493]. dual-criteria [1168]. dual-cubes [75]. dual-dictionary [1103]. Dual-mode [1113, 1250, 1737]. duality [877]. duality-based [877]. due [782]. duplex [2369]. duplication [805, 2003, 2344]. duplication-based [2344]. durable [1676]. during [671, 1696]. DVFS [796, 1118, 1203, 1905]. DVFS-enabled [1203]. DVM [2360]. Dynamic [31, 33, 34, 60, 72, 79, 207, 209, 243, 249, 281, 287, 328, 358, 362, 373, 390, 418, 434, 443, 463, 487, 503, 505, 513, 569, 591, 606, 632, 638, 652, 658, 702, 705, 757, 784, 788, 792, 853, 888, 943, 975, 1012, 1042, 1059, 1063, 1161, 1183, 1247, 1269, 1287, 1336, 1337, 1364, 1441, 1451, 1505, 1506, 1509, 1563, 1590, 1597, 1598, 1608, 1612, 1613, 1633, 1654, 1659, 1684, 1764, 1768, 1778, 1796, 1810, 1842, 1915, 1924, 1964, 1999, 2003, 2004, 2071, 2087, 2095, 2099, 2100, 2113, 2158, 2163, 2186, 2187, 2202, 2220, 2234, 2275, 2292, 2338, 2340]. Dynamic-width [1042]. Dynamically [103, 797, 1119, 1225, 1791]. dynamics [122, 575, 600, 642, 712, 732, 1439, 1718, 2351, 2372]. Dynamo [879].

e-health [1347, 1486]. e-learning

[815, 2234]. E-OSched [1898]. e-rental [903]. e-Science [244, 924, 1267]. E2FS [1691]. EA [807]. EAAM [1727]. Early [390, 1914]. Earthquake [28]. EC [2356]. EC-600JD [2356]. ECB4CI [1874]. ECC [1843, 2000]. ECC-based [1843]. ECCN [1677]. ECDSA [2198]. ECG [1963]. echo [1133, 2289]. eco [2004]. eco-hydrological [2004]. economy [389]. Edge [766, 1029, 1337, 1389, 1457, 1475, 1927, 2097, 2328, 2342]. Edge-disjoint [1389, 1475]. edges [1998, 2250]. Editorial [82, 266, 289, 344, 456, 537, 565, 660, 900, 1207, 1837]. Editors [433]. education [1809, 2014]. educational [1212, 1821]. EEG [2156]. EF [224]. EF-MPR [224]. Effect [1720, 1811, 2001, 2199, 2278]. Effective [88, 110, 262, 573, 745, 780, 904, 1265, 1313, 1484, 1643, 1702, 1807, 2195, 2318]. effectively [249]. effects [659, 1361, 1765]. efficiency [241, 275, 408, 673, 709, 897, 979, 1077, 1104, 1145, 1402, 1418, 1422, 1444, 1450, 1464, 1505, 1701, 1780, 1947, 2031, 2034, 2071, 2180, 2288, 2323, 2355]. Efficient [10, 13, 40, 84, 92, 99, 125, 219, 221, 224, 232, 280, 384, 403, 413, 467, 531, 572, 576, 582, 586, 588, 589, 593, 628, 640, 657, 670, 674, 771, 813, 844, 863, 877, 880, 886, 905, 910, 960, 961, 988, 1008, 1040, 1045, 1056, 1058, 1063, 1066, 1075, 1092, 1110, 1123, 1171, 1191, 1248, 1254, 1283, 1314, 1366, 1385, 1386, 1419, 1428, 1512, 1535, 1544, 1561, 1568, 1571, 1577, 1644, 1654, 1675, 1704, 1727, 1735, 1743, 1790, 1833, 1859, 1868, 1886, 1936, 1938, 1966, 2018, 2030, 2061, 2081, 2098, 2115, 2149, 2151, 2161, 2166, 2194, 2201, 2204, 2243, 2304, 2307, 2315]. efficient [21, 69, 90, 194, 230, 269, 273, 276, 278, 292, 313, 315, 338, 359, 383, 394, 400, 402, 406, 453, 460, 552, 625, 653, 665, 769, 782, 811, 859, 870, 876, 914, 929, 996, 1094, 1107, 1118, 1122, 1261, 1274, 1292, 1296, 1308, 1351, 1368, 1477, 1493, 1529, 1548, 1563, 1653, 1655, 1683, 1693, 1708, 1720, 1751, 1760, 1778, 1831, 1917, 1971, 1986, 2020, 2050, 2130, 2168, 2200, 2202, 2220, 2285, 2309, 2325, 2326]. efficient-processing [1368]. Efficiently [1006]. effort [523]. EFuNN [1928]. EGCA [1236]. eigensolvers [1396]. eight [206]. Elastic [419, 714, 983, 1200, 1586, 1691, 2173]. election [1164, 1448, 1750]. electric [381]. electro [376]. electro-optical [376]. electrocardiogram [1963]. electrochemical [2356]. electromagnetic [472, 1996, 2107]. electronic [65, 909, 911, 1221, 1740, 2124]. element [246, 1972]. elementary [1936, 2176]. elephant [1353]. eliminating [307, 424]. elimination [164]. elite [2236]. elite-guided [2236]. Elliptic [430, 905, 960, 1923, 1944, 2036, 2226]. elongation [250]. email [1345]. Embedded [83, 221, 259, 275, 295, 355, 394, 643, 786, 798, 972, 998, 1234, 1287, 1470, 1566, 2189, 2303]. Embedding [993, 1478, 1828, 2018, 2290, 2341]. emergency [69, 904, 1860, 1975]. emergency-efficient [69]. emergent [2059]. Emerging [1062, 1207, 1456, 1720, 2256]. emoticon [2086]. emoticon-based [2086]. emotion [553]. emotion/mood [553]. emotional [1009, 2136]. emotional-based [1009]. empirical [1828]. employing [690, 970, 2086]. empowered [1537]. emulation [1579, 1884]. Enable [1661]. enabled [768, 818, 1203, 1310, 1349, 1624, 1870, 1885, 2287]. Enabling [2, 39, 289, 535, 554, 788, 1007, 1837]. Encoder [93, 941, 1078, 1406, 1414, 1458, 2059, 2308, 2314]. encoders [1432, 1557]. encoding [707, 1163, 1408, 1458, 1844, 1920]. encrypted [1689, 1812, 2106, 2326]. encryption [532, 772, 843, 982, 1149, 1366, 1812, 1838, 2141, 2300]. End [26, 95, 696, 1701, 1805, 1806]. End-To-End [26, 95, 696, 1701, 1805, 1806]. ended [1861]. Energy [10, 21, 101, 125, 194, 206, 224, 273, 276, 278, 315, 360, 380, 394, 400, 402, 403, 406– 408, 411, 593, 607, 755, 782, 784, 796, 846, 863, 870, 880, 886, 897, 910, 921, 979, 984, 985, 995, 1001, 1007, 1008, 1040, 1043, 1062– 1064, 1079, 1107, 1111, 1225, 1359, 1391, 1402, 1404, 1418, 1420, 1450, 1493, 1505, 1506, 1520, 1529, 1563, 1568, 1577, 1612, 1613, 1663, 1675, 1683, 1686, 1736, 1741, 1747, 1749, 1765, 1871, 1947, 1951, 1962, 1966, 2019, 2034, 2063, 2071, 2114, 2151, 2156, 2180, 2220, 2243, 2284, 2302, 2306, 2309, 2315, 2323, 2331]. energy [29, 269, 275, 434, 453, 460, 495, 529, 653, 709, 778, 879, 930, 974, 996, 1145, 1448, 1489, 1681, 1720, 1760, 1780, 1849, 1939, 2020, 2031, 2078, 2099, 2100, 2202, 2285, 2325]. Energy- [360]. Energy-aware [407, 529, 846, 995, 1506, 1663, 1871, 1939, 2099, 2100, 2114, 2309]. energy-efficiency [709]. Energy-Efficient [269, 273, 394, 400, 402, 406, 460, 593, 653, 782, 880, 886, 996, 1008, 1040, 1063, 1107, 1493, 1563, 1568, 1577, 1675, 1683, 1720, 1760, 1966, 2020, 2202, 2220, 2243, 2285, 2315, 2325]. energy-limited [1448]. Energy-oriented [1062]. energy-performance [1489]. energy-proportional [1007]. Energy-saving [985, 1962, 2063]. enforcement [2199]. Enforcing [1822, 1855]. engaging [1283]. engine [467, 953, 1573, 1595, 2212]. EngineCL [2079]. engineering [155, 171, 469, 595, 927]. engines [786, 1525, 1742]. English [2105]. enhance [2095, 2102]. Enhanced [67, 141, 217, 238, 405, 598, 705, 888, 892, 1223, 1236, 1551, 1611, 1874, 1913, 2077, 2214, 2277]. enhancement [365, 742, 1469, 1553, 2126]. enhancer [18]. Enhancing [53, 196, 304, 361, 391, 441, 459, 673, 945, 1077, 1606, 1701, 1802]. Enlargement [1952]. enriched [748, 1887]. ensemble [2289]. enterprise [1329, 1696]. enterprises [2153]. enterprising [424]. entirely [134]. entrepreneurship [1368]. entries [1999]. Entropy [1958, 2060, 2134]. Entropy-based [2134]. environment [85, 86, 115, 245, 261, 270, 287, 326, 351, 367, 395, 421, 434, 487, 494, 519, 543, 551, 553, 555, 556, 588, 652, 680, 687, 736, 768, 772, 777, 818, 855, 878, 913, 958, 969, 970, 975, 1034, 1043, 1053, 1056, 1085, 1119, 1182, 1218, 1247, 1273, 1337, 1390, 1456, 1469, 1470, 1494, 1533, 1539, 1575, 1612, 1613, 1616, 1628, 1637, 1669, 1788, 1795, 1849, 1882, 1893, 1960, 1961, 1972, 1974, 2014, 2092, 2127, 2128, 2130, 2170, 2194, 2214, 2270, 2328, 2340]. environmental [463, 2026]. environments [32, 67, 116, 128, 196, 209, 211, 214, 269, 295, 361, 426, 490, 495, 499, 530, 548, 632, 633, 661, 672, 674, 681, 684, 739, 814, 874, 891, 892, 1025, 1069, 1083, 1097, 1193, 1216, 1217, 1371, 1420, 1522, 1534, 1545, 1564, 1606, 1728, 1776, 1796, 1802, 1804, 1857, 1866, 1923, 1935, 2063, 2086, 2120, 2162, 2171, 2186, 2210, 2240, 2244, 2301]. EPC [1542]. Epidemic [413, 2329]. equation [813, 1427, 2359]. equations [140, 596, 937, 1020, 1178, 1266, 1445, 1498, 1501, 2055, 2345]. equilibrium [486, 491]. equipment [1084]. equipped [1099]. era [1060, 1140, 1376, 1537, 1889]. erasure [810, 1269, 1981, 2215]. ERP [651]. Erratum [373, 447, 992, 1041, 1140, 1258, 1286, 1301, 1328, 1467, 1472, 1613, 1646, 1665, 1806, 1814]. Error [152, 970, 1247, 1281, 1462, 1631, 1937, 2107]. Error-resilient [1462]. ESNemble [2289]. essential [927, 1393]. establishment [1282, 2351, 2372]. estimated [589]. Estimating [1524]. estimation [43, 148, 150, 523, 614, 731, 1082, 1090, 1151, 1235, 1705, 1813, 1814, 1816, 1956, 2043, 2052, 2088, 2121]. estimator [236, 611]. Ethernet [1361, 1968]. Euclid [609]. Euclidean [2122]. evaluate [212]. Evaluating [21, 181, 632, 637, 836, 1130]. Evaluation [234, 264, 293, 342, 365, 419, 479, 562, 601, 695, 719, 840, 931, 933, 969, 1035, 1085, 1104, 1238, 1245, 1285, 1297, 1310, 1413, 1429, 1434, 1531, 1596, 1648, 1659, 1678, 1725, 1777, 1795, 1848, 1854, 1870, 1916, 2066, 2129, 2205, 2316, 2343, 2375]. Evaluations [681, 1771]. evasive [1550]. even [923, 1819]. event [79, 513, 1970, 2188]. event-driven [1970]. events [131, 904]. Eventual [390, 1164]. evictions [1617]. evidence [136, 887, 1730]. Evolution [134, 427, 534, 955, 1666]. Evolutionary [170, 464, 658, 695, 701, 935, 942, 1129, 1416, 1461, 1514, 1547]. evolutionary-mined [695]. evolving [2167]. exa [1599, 2215]. exa-scale [1599, 2215]. exact [506, 865, 873, 2286]. examining [510]. exascale [396, 1376, 1377, 1479, 2318]. exception [423]. exchange [536, 665, 706, 738, 844, 914, 1265, 2150, 2367]. exchanged [839, 2018]. exclusion [729, 1592, 1693]. exclusive [335]. executable [2008]. executed [2180]. executes [29]. executing [1620]. execution [162, 241, 242, 257, 301, 322, 385, 421, 425, 467, 541, 589, 776, 841, 1247, 1253, 1334, 1619, 1767, 1836, 2037, 2057, 2079, 2101, 2106, 2302]. executions [1358]. exhaustive [1626, 1675]. exist [1344]. Expanding [2007]. expansion [662, 668]. experience [2360]. Experiences [27, 2188]. experiment [743, 1298]. Experimental [1303, 1330, 1769, 2362]. experiments [295]. expert [254, 1214].

Explicit [320, 443]. explicitly [105]. exploit [2340]. exploitation [599, 1644]. Exploiting [122, 161, 649, 732, 898, 926, 945, 971, 978, 992, 1142, 1358, 1731, 2155]. exploits [1211]. exploration [1158]. Exploring [1079, 1603, 1645, 1646, 1908, 2038]. exponential [267, 962]. exposing [608]. expression [700, 1567, 2152]. expressway [1703]. Ext2 [902]. Ext2/3 [902]. Extended [353, 1677, 2068, 2307]. Extending [254, 475, 613, 637, 1020, 2177]. extensible [998]. extension [296, 715, 822, 1346, 1459, 1484, 2117]. extensions [1387]. extent [2237]. extra [833]. extraction [147, 283, 359, 988, 1170, 1428, 2131, 2203, 2276]. extrapolated [938]. extreme [246, 289, 290, 1387, 1917]. extreme-scale [1387]. extremely [55, 763]. Eye [1549, 2093]. eye- [1549]. Eye-tracking [2093].

F [271]. F-MPJ [271]. F2N [2325]. fabrication [275]. Face [151, 1086, 1488, 1959, 2108]. facet [682]. facial [1549]. facial-gaze [1549]. facilitates [507]. facilitating [23]. facility [128, 187, 935, 2084]. Facing [1276]. factor [1145, 1312, 1347, 1486, 1804, 1879]. factorization [1404, 1906]. factors [1288, 1701]. fading [492, 1333]. failure [3, 199, 441, 729, 969, 2348]. failure-prone [441, 969]. failures [782, 2129, 2307]. Fair [76, 1822, 1827, 2016]. fake [1531, 2277]. fall [748]. False [265]. family [55, 74, 198, 1192, 1255, 1299, 1384]. fan [404]. fan-out [404]. farm [1491, 2073]. Fast [77, 79, 109, 186, 220, 238, 321, 359, 387, 416, 466, 552, 579, 587, 612, 691, 747, 953, 998, 1068, 1107, 1151, 1152, 1191, 1204, 1414, 1458, 1464, 1634, 1645, 1646, 1680, 1729, 1801, 1813, 1814, 1816, 1826, 1996, 2017, 2032, 2115, 2212, 2351, 2372]. FastFlow [2181]. FastMFDs [2115]. FastNBL [2351, 2372]. fat [1045, 1101, 1378]. fat-tree [1101]. fat-trees [1045, 1378]. Fault [8, 47, 52, 74, 102, 103, 108, 233, 333, 336, 345, 389, 562, 617, 632, 698, 735, 799, 833, 881, 1044, 1130, 1136, 1162, 1185, 1224, 1230, 1385, 1417, 1490, 1532, 1623, 1643, 1772, 1819, 1859, 1864, 1865, 1871, 1903, 1914, 2192, 2206, 2251, 2255, 2333, 2367]. fault-resilient [1490, 1864]. fault-resistant [735]. fault-tolerance [1162, 1224, 2333]. Fault-tolerant [8, 102, 103, 233, 336, 345, 698, 833, 1044, 1136, 1185, 1230, 1385, 1417, 1532, 1623, 1772, 1819, 1865, 1871, 2192, 2206, 2251, 2255, 2367]. faults [1385]. faulty [390, 723, 1998, 2250]. favor [1489]. FCM [2349]. FDMA [654]. FDTD [472, 928]. FEAD [982]. Feasibility [1330, 2001]. feature [748, 931, 1632, 1873, 1926, 1959, 1961, 2090, 2094, 2203, 2236, 2247, 2260]. feature-based [1959]. features [258, 1369, 1488]. federated [305, 1559, 1885]. federation [1773, 2022]. feedback [1095, 1354, 1531]. FEM [472, 2076]. femtocells [1142]. femtolet [1758, 2325]. femtolet-based [1758, 2325]. Feynman [2308]. FFT [481, 1229, 1441]. fiber [883, 1793]. fidelity [662, 1153]. Field [94, 670, 741, 1412, 1413, 1735, 1851, 2048, 2254]. Fieldbus [1852]. File [6, 11, 59, 72, 135, 199, 244, 325, 403, 483, 501, 739, 788, 893, 902, 1215, 1260, 1297, 1341, 1479, 1508, 1776, 1818, 1827, 2008, 2299, 2318]. file-sharing [1260]. files [337, 893, 1530, 1883, 1958]. filesystem [902]. filter [146, 166, 429, 572, 1152, 1208, 1560, 1963, 2254]. filter-based [2254]. filtering [114, 514, 552, 634, 864, 1308, 1312, 1409, 1696, 1880, 1986, 2275]. Financial [1019]. find [1629, 1692]. Finding [70, 699, 708, 865, 1353, 1517, 1591, 1987, 2221]. fine [898]. fine-grained [898]. fingerprint [2142]. Finisterrae [162]. finite [246, 320, 370, 1972]. fire [945, 1077, 1099, 2040]. Firefly [1900]. firewall [1796]. firmware [1470]. first [1930]. Fisher [1859]. fixed [527, 597, 1724, 1732, 1733]. fixed-complexity [597]. fixed-degree [527]. fixed-length [1732, 1733]. FLAPS [1653]. flash [501, 531, 626, 1894, 2350]. flash-based [626]. flaws [1738]. FLDA [1859]. flexibility [2095]. flexible [428, 591, 713, 1455, 1485, 1711, 1848, 1946]. flipped [1809]. floating [506, 1790, 2274]. floating-point [506]. flooding [1265]. Floorplan [311]. Floorplan-aware [311]. FLOPS [1706]. flow [78, 144, 222, 233, 290, 576, 750, 761, 1010, 1015, 1030, 1170, 1239, 1438, 1719, 1997, 2107]. flow-table [1997]. flower [2239]. flows [1639]. fluctuations [1963]. fluid [712, 732, 928, 951, 1070, 1438, 1668]. fluid-structure [1668]. fluidized [1835]. fly [124, 858, 1038, 1474]. flying [1561]. FMM [172, 471, 1415]. fMRI [745]. FOAF [652]. focusing [766, 1298]. Fog [1340, 1448, 1653, 1654, 1758, 1760, 1966, 1976, 2097, 2149, 2325]. Fog-based [1966]. Fog-supported [1448, 1653]. folded [251, 799, 1903, 2321]. follower [935]. food [66, 2042, 2145]. footprint [222]. force [364, 1759]. forecast [975, 1786]. Forecasting [442, 815, 947, 1077, 1821, 2283]. Forensic [909, 1550, 1818, 1883, 1953]. forensics [2200]. forest [945, 1077, 1316, 2040]. Foreword [145]. forgeries [1883]. forgery [966]. fork [1823]. Form [216]. Formal [28, 255, 584, 725, 824, 889, 1619, 1725, 2164]. Formalization [896]. formally [2365]. format [1774]. formation [657, 2022, 2357]. formulas [858]. Formulating [1178]. formulation [2175, 2266]. Fortran [508, 1362]. forum [1369]. forums [1546]. Forward [370, 738]. forwarding [1239, 2329]. FOS [2296]. foundation [550]. Foursquare [1554]. FP [2015]. FP-growth [2015]. FPGA [167, 329, 506, 578, 1009, 1441, 1704, 1771, 1997, 2079, 2184, 2185, 2229]. FPGAs [1790, 2052]. fractional [813, 1027, 2029, 2345]. fractional-order [2029]. Fragmentation [133, 282, 2127]. fragmented [301, 829, 1438, 1447, 2359]. frame [888, 1432]. frameNet [1459]. Framework [23, 31, 63, 96, 293, 337, 382, 494, 512, 545, 565, 573, 686, 688, 702, 712, 719, 742, 767, 777, 787, 824, 835, 846, 983, 1003, 1026, 1127, 1130, 1156, 1190, 1199, 1205, 1233, 1237, 1340, 1369, 1382, 1401, 1474, 1477, 1506, 1538, 1602, 1605, 1618, 1698, 1764, 1767, 1782, 1805, 1806, 1848, 1850, 1854, 1885, 1889, 1892, 1895, 1916, 1922, 1924, 1926, 1948, 2026, 2088, 2127, 2130, 2153, 2164, 2167, 2175, 2200, 2280, 2297]. free [230, 235, 267, 644, 1135, 1439, 1484, 1647, 1941, 2048]. frequency [214, 778, 984, 1280, 1396, 1924]. frequency-scaling [778]. frequent [627, 1105, 1370, 1584, 2010, 2241]. frequently [1192]. friendly [754, 893, 2299]. fruit [1474, 2030]. FS [1848]. FS-IIoTSim [1848]. FSH [352]. FSM [1860]. FSM-based [1860]. full [947, 1741, 1869, 2369]. Fully [102, 599, 1031, 1239, 1380]. Fully-Chained [102]. fully-connected [1380]. Function [197, 800, 1039, 1297, 1445, 1492, 1581, 1718, 1796, 1975, 2008, 2107, 2268, 2269, 2293]. function-based [1975]. function-virtualized [1796]. functional [982, 1715, 2115]. functions [692, 1937, 2045, 2176]. fundamental [77]. FuPerMod [826]. Fusion [933, 1173, 1647, 1688, 1694, 1702, 2147, 2260]. fusion-based [1694, 1702]. Future [62, 400, 495, 565, 1066, 1197, 1198, 1214, 1278, 1942, 2080, 2132, 2142, 2153, 2211]. fuzzing [1851]. Fuzzy [146, 254, 520, 544, 684, 686, 718, 725, 775, 1040, 1137, 1468, 1560, 1561, 1581, 1628, 1650, 1755, 1896, 1914, 1970, 2034, 2213, 2234, 2243, 2349]. fuzzy-based [1896]. fuzzy-bee [1561]. FuzzyCLIPS [254].

G [736, 1572]. G-IK-SVD [1572]. Gabor [1488]. GAER [886]. gain [1142]. Gait [545, 546]. gallbladder [2144]. Game [76, 100, 464, 486, 1089, 1141, 1144– 1146, 1148, 1149, 1219, 1226, 1390, 1531, 1616, 1636, 1964, 2112]. game-based [1964]. game-theoretic [76, 1145, 1146]. games [1787, 1948]. gamma [848, 1192, 2367]. gang [234, 706]. gap [553]. garbage [241, 991, 1257, 1258, 1277]. garbage-output [1277]. gas [1438, 1439]. gate [2308]. gates [1936]. gateway [1808]. gathering [968]. gating [1110]. Gauss [763]. Gaussian [146, 1194, 1389]. gaze [1549]. Gb [1968]. Gbps [93]. gem5 [1055]. Gem5v [1055]. GEMM [2364]. gene [131, 134, 700, 1001]. Gene/Q [1001]. General [331, 357, 1317, 1671, 2046]. generalized [321]. generated [212, 2123]. generating [800]. generation [62, 179, 566, 705, 1171, 1208, 1779, 1825, 1851, 1885, 2089, 2191, 2291]. generations [1197]. generative [2089]. generators [2054, 2174]. Generic [726, 922, 1035, 1909, 2109]. Genetic [167, 311, 361, 462, 583, 602, 886, 957, 1040, 1050, 1077, 1375, 1420, 1437, 1481, 1597, 1607, 1632, 1721, 2048, 2063, 2088, 2175, 2284]. genome [1004]. genomic [543]. geo [1196, 1893]. geo-distributed [1196, 1893]. Geocasting [1466, 1467]. Geocasting-based [1466, 1467]. geographic [1084]. geographically [1600, 2120]. geometric [2155]. Geometry [608, 2046]. Geometry-based [2046]. geophysics [2360]. gesture [557, 1538]. Gillespie [1443]. ginSODA [2361]. GIS [1611]. given [581]. global [425, 691, 793, 803, 861, 1151, 1236, 1536, 1661, 1816, 1831, 2111, 2236, 2239]. glove [1813, 1814]. GMDH [807, 1354]. GMRES [260, 834, 2061, 2343]. goal [315, 613]. goal-oriented [613]. goals [2016]. Golang [2057]. good [2283]. Google [832, 1454]. government [1888, 2281]. GPGPU [172, 452, 471, 575, 576, 937, 1232, 1311, 1752, 1908]. GPGPUs [1123, 1723, 1734]. GPU [132, 190, 253, 324, 364, 370, 417, 440, 442, 459, 466, 472, 594, 618, 641, 707, 818, 819, 821, 828, 834, 841, 843, 894, 928, 930, 942, 946– 948, 1011, 1016, 1019, 1020, 1029, 1048, 1137, 1154, 1155, 1171, 1184, 1205, 1246, 1268, 1310, 1349, 1354, 1375, 1381, 1394, 1402, 1411, 1421, 1423, 1425, 1426, 1439, 1440, 1442, 1454, 1499, 1567, 1573, 1582, 1589, 1607, 1624, 1626, 1675, 1685, 1754, 1771, 1792, 1798, 1800, 1830, 1920, 2033, 2035, 2044, 2047, 2061, 2064, 2065, 2068, 2075, 2079, 2118, 2133, 2166, 2188, 2198, 2257, 2332, 2343, 2366]. GPU [185, 222, 629, 717, 756, 780, 822, 918, 933, 951, 1006, 1058, 1106, 1152, 1308, 1498, 1621, 1630, 1634, 1717, 1748, 1968, 2190]. GPU-accelerated [440, 821, 1184, 1439, 1454, 1920]. GPU-assisted [707]. GPU-based [185, 640, 819, 1048, 1137, 1154, 1205, 1375, 1426, 1607, 1626, 1685, 2064, 2075]. GPU-enabled [818, 1349, 1624]. GPU-likely [841]. GPU-optimized [1573]. GPU-parallelized [2366]. GPU-sorting [1394]. GPUCloudSim [2117]. GPuEGO [942]. GPUs [175, 179, 260, 310, 320, 386, 470, 474, 605, 695, 763, 837, 860, 934, 950, 1127, 1170, 1186, 1362, 1503, 1572, 1668, 1722, 1941, 2045, 2117, 2218, 2297, 2316, 2361]. Gradient [320, 474, 952]. grading [1947]. grain [313]. grained [898, 926]. gram [1803]. Granularity [568, 1493, 1662]. Granularity-based [1662]. graph [346, 486, 800, 898, 1058, 1098, 1585, 1698, 1742, 1744, 1958, 2010, 2138, 2277]. graph-based [346, 1742, 2277]. graphic [1571]. Graphics [132, 160, 188, 525, 627, 631, 712, 732, 762, 1107, 1320, 1322, 1480, 1744, 2246]. Graphs [19, 45, 74, 111, 113, 154, 237, 318, 334, 570, 1452, 1478, 1483, 1591, 1629, 2058, 2167, 2221, 2225]. GRASP [2257]. gravitation [2133]. gravitation-based [2133]. Gravitational [1046]. Gray [258]. greedy [747, 1228, 1521, 2267]. green [274, 399, 401, 449, 450, 563, 874, 875, 877, 878, 887, 890, 1036, 1225, 1491, 1682]. GreenCloud [407]. greener [1597]. greening [2344]. GREMLINS [261]. Grid [8, 13, 20, 32, 33, 35, 43, 57, 71, 79, 101, 108, 115, 193, 196, 211, 213, 243, 249, 261, 282, 287, 305, 342, 361, 387, 389, 392, 439, 487, 498, 583, 590, 664, 681, 683, 685, 686, 692, 718, 731, 743, 768, 775, 811, 866, 872, 997, 1039, 1041, 1050, 1053, 1183, 1221, 1298, 1453, 1478, 1591, 1658]. Grid-based [743]. grids [22, 27, 59, 72, 125, 126, 132, 317, 635, 827, 1507, 2018, 2118]. GridSim [141]. GridSolve [535]. Grossone [577]. Group [73, 202, 314, 318, 427, 532, 541, 833, 888, 1168, 1415, 1609, 1750, 1803, 1857, 1960, 2055, 2373]. group-based [427]. grouped [1603]. Grouping [1670, 2292]. groups [1223]. groupware [40]. growth [2015]. guarantee [362, 1974]. guaranteed [1763, 1833]. guard [1027]. guessing [1554]. Guest [82, 344, 433]. guided [1354, 2236]. Gustafson [637].

H [919, 1013]. H-PC [1013]. H-Phy [919]. H.264 [477, 970, 1408, 1414]. H.264/AVC [477, 970, 1408, 1414]. Hadoop [767, 978, 1156, 1296, 1341, 1383, 1484, 1577, 2005, 2299]. hallucination [1086]. Hama [1605]. Hamiltonian [74, 75, 318, 760, 1591]. Hamiltonianicity [1452]. Hamming [882, 1107, 1732, 1733]. hand [1538, 1891]. Handling [423, 1278, 2135, 2294]. Handoff [495, 691]. handover [220, 238, 352, 876, 1193]. handwritten [2089]. Haralick [258]. hard [1353, 1532]. Hardware [160, 192, 231, 241, 326, 591, 628, 784, 1098, 1100, 1158, 1175, 1191, 1505, 1523, 1527, 1704, 1726, 2036, 2180, 2183, 2249]. hardware-assisted [241]. hardware-level [2249]. hardware/software [628, 1098]. harmful [1128]. harmonic [770]. harmonized [2319]. Harmony [2370]. Hash [483]. Hash/Table [483]. hashing [319, 1618]. Haskell [1714]. having [671]. hazard [576]. HCI [2086]. HD [1458]. HDFS [1068, 1530]. HDMR [356]. head [1657]. healing [366, 734]. health [1091, 1340, 1347, 1486, 1544, 2084, 2263]. healthcare [1218, 1339, 1341, 1882, 2145]. heart [819]. heartbeat [2150]. heat [2219, 2359]. HeDPM [1588]. heliostat [1413, 1435, 2048]. Hellman [736]. help [1192]. hepta [1589]. hepta-diagonal-sparse [1589]. heterogeneities [123]. Heterogeneity [282, 1234, 2110, 2129]. Heterogeneous [14, 91, 106, 117, 157, 166, 174, 178, 181, 183, 184, 186, 215, 270, 280, 360, 365, 367, 386, 414, 477, 530, 589, 603, 604, 613, 629, 640, 688, 747, 782, 805, 826, 929, 948, 949, 963, 1000, 1014, 1016, 1056, 1064, 1073, 1075, 1129, 1147, 1233, 1270, 1307, 1392, 1398, 1415, 1418, 1420, 1422, 1448, 1502, 1505, 1518, 1533, 1588, 1679, 1718, 1749, 1784, 1798, 1893, 1896, 1898, 1907, 1911, 1969, 2027, 2037, 2044, 2074, 2079, 2112, 2148, 2154–2156, 2172, 2179, 2180, 2184–2186, 2189, 2190, 2192, 2209, 2243, 2329, 2340, 2354]. heterogeneous [116, 209, 294, 434, 457, 1096, 1498, 1587]. heuristic [237, 262, 628, 713, 809, 846, 1053, 1188, 1306, 1373, 1510, 1794, 2071, 2148, 2170, 2292, 2293, 2312]. heuristic-based [1188]. Heuristics [125, 196, 206, 367, 747, 1132, 1241, 1747, 2220]. heuristics-based [2220]. HEVC [941, 1014, 1078, 1406, 1408, 1414, 1424, 1426, 1429, 1432, 1458, 1463, 1464, 1971, 2039, 2044]. HFIM [1584]. HHS [1248]. hidden [689, 1768]. hiding [675, 678, 982, 1820, 1844]. Hierarchical [22, 122, 180, 285, 393, 431, 475, 584, 609, 623, 730, 838, 1071, 1119, 1136, 1176, 1446, 1603, 1705, 1724, 1920, 2240, 2333]. hierarchies [105]. hierarchy [519]. High [1, 4, 7, 82, 89, 93, 97, 103, 143, 155, 171, 210, 223, 264, 270, 273, 286, 341, 362, 364, 393, 412, 465, 469, 501, 518, 566, 595, 617, 643, 650, 695, 698, 710, 761, 836, 863, 867, 910, 918, 927, 949, 1004, 1026, 1087, 1088, 1092, 1105, 1132, 1156, 1170, 1197, 1215, 1246, 1287, 1289, 1292, 1299, 1351, 1360, 1361, 1374, 1376, 1393, 1404, 1405, 1407, 1413, 1455, 1464, 1477, 1485, 1490, 1523, 1526, 1532, 1578, 1582, 1675, 1685, 1688, 1712, 1734, 1852, 1855, 1932, 1997, 2000, 2036, 2042, 2053, 2056, 2183, 2193, 2198, 2201, 2205, 2274, 2283, 2306, 2332, 2375]. high [28, 37, 61, 96, 352, 356, 359, 375, 565, 636, 675, 707, 711, 785, 822, 896, 1021, 1046, 1114, 1536, 1710, 1920, 2127, 2154, 2172, 2242]. High-bandwidth [1455]. high-capacity [2306]. high-dimensional [1582]. high-efficiency [1464]. High-efficient [1092, 1292]. high-energy [910]. high-level [822, 1855, 2172, 2183]. high-motion [707]. high-order [286, 393, 1526, 1536]. High-Performance [1, 89, 96, 143, 210, 223, 273, 636, 643, 761, 836, 910, 1004, 1021, 1026, 1046, 1088, 1105, 1132, 1156, 1197, 1246, 1287, 1360, 1376, 1393, 1404, 1477, 1485, 1490, 1523, 1532, 1675, 1688, 1710, 1712, 1852, 1920, 1997, 2000, 2036, 2042, 2053, 2056, 2127, 2154, 2193, 2201, 2205, 2283, 2375]. high-pressure [2042]. high-quality [1351, 1578, 1685]. high-radix [103, 896, 1114]. High-speed [352, 359, 1170, 1361, 1734, 1932, 2198, 2332]. High-throughput [698, 1374, 1405, 2242]. higher [686, 1194, 1535, 1821, 2014]. higher-dimensional [1535]. Highly [14, 77, 620, 747, 893, 1193, 1439, 1530, 1778, 2041, 2059]. Highway [1327, 1328]. histogram [678]. histogramming [1492]. Historical [358, 373, 1198]. history [423, 530, 803]. history-based [530]. HiveQL [1117]. HLMD [2249]. HMM [64]. hoc [30, 47–49, 100, 204, 205, 213, 314, 345, 346, 406, 497, 502, 667, 716, 728, 1089, 1132, 1262, 1273, 1530, 1561, 1750, 1762, 1897, 1904, 1939]. HoL [1101, 1380]. HoL-blocking [1101]. hole [1591]. holistic [510, 1993]. home [65, 659, 1371, 1849]. homogeneous [1679]. homography [546]. homomorphic [1812]. honeypot [516]. hop [119, 398, 560, 1262]. horizontal [131]. Host [807, 963, 1133, 1950, 2314]. host-based [963]. hosting [683]. Hough [1634]. HPC [83, 954, 1040, 1162, 1383, 1391, 1935, 2002, 2006, 2029, 2063, 2066, 2155, 2240, 2256, 2295, 2311]. HPCCS [91]. HRCPS [620]. HReMAS [2028]. HScheduler [1321]. HSIR [2228]. HTTP [1550]. hull [2238]. Human [557, 748, 891, 1352, 1872, 1879, 2080, 2085, 2219]. Human-computer [2080, 2085]. human-robot [557]. Hybrid [34, 57, 88, 108, 144, 180, 193, 245, 249, 268, 503, 641, 642, 660, 661, 669, 707, 722, 743, 770, 828, 855, 883, 900, 942, 955, 957, 962, 1020, 1040, 1058, 1080, 1136, 1188, 1197, 1229, 1289, 1395, 1414, 1508, 1517, 1536, 1584, 1586, 1621, 1632, 1676, 1711, 1721, 1749, 1763, 1792, 1798, 1862, 1863, 1867, 1893, 1894, 1896, 1900, 1949, 2028, 2038, 2158, 2186, 2191, 2192, 2198, 2228, 2348, 2370]. hybrid-classifiers [503]. hybrid-cloud [855]. Hybridizing [2111]. HybridPlan [722]. hydrological [2004]. HYFMGPU [2067]. Hyper [251, 760, 831, 1053, 1363, 2072]. Hyper-Hamiltonian [760]. hyper-heuristic [1053]. hyper-scale [1363]. hyper-star [251]. hyper-torus [831]. hyper-trabeculation [2072]. hypercube [321, 1623, 1667, 2165]. hypercubes [309, 341, 615, 730, 760, 799, 839, 1903, 1998, 2018, 2250]. hypergraph [363, 1622]. hyperheuristic [2158]. hyperheuristics [1731]. hyperparameter [1746]. hyperspectral [186, 934, 1430, 2033, 2052, 2067, 2068]. hypertensive [1013]. hypervisor [446, 447, 1616]. HyperX [648, 1044]. HYSTERY [1749].

I-IoT [1840]. I/O [335, 776, 788, 1022, 1026, 1088, 1134, 1477, 1509, 1663, 2358]. IaaS [588, 754, 816, 1450, 1673, 1914, 2063, 2129]. IBM [104]. ICS [1851]. ICT [913, 1085]. ID [672, 903, 930]. ID-based [672, 903]. ID3 [50]. identification [149, 290, 1294, 1543, 1875, 1961, 2052, 2137, 2228]. identification-issuing [1543]. Identifying [1325, 1531, 1878]. identity [98, 705, 844, 1177]. identity-based [98, 844, 1177]. idle [2201]. IDS [1540, 1887, 1945]. IEEE [238, 350, 391]. II [171, 900, 1517, 2293]. IIoTSim [1848]. IIR [1963]. IIR-LPF [1963]. IK [1572]. illumination [128]. Image [132, 147, 149, 379, 559, 594, 662, 676, 678, 762, 780, 1029, 1137, 1149, 1155, 1286, 1430, 1503, 1548, 1560, 1573, 1634, 1752, 1842, 2012, 2013, 2030, 2068, 2276, 2300]. image-based [2030]. image/video [559]. imagery [934]. images [186, 195, 558, 572, 675, 820, 916, 1148, 1208, 1428, 2033]. imbalanced [1342, 1785]. imbalances [1304]. immediate [390]. immersed [951]. immune [1650]. impact [608, 801, 1617]. impacts [2129]. impaired [1454]. imperfect [1899]. imperialist [770, 1640]. impersonation [1687]. Implementation [63, 72, 185, 199, 222, 258, 265, 368, 372, 458, 467, 506, 531, 594, 597, 664, 681, 687, 735, 751, 756, 763, 812, 823, 828, 829, 930, 942, 947, 950, 990, 1129, 1167, 1268, 1354, 1367, 1428, 1430, 1438, 1439, 1458, 1517, 1571, 1704, 1708, 1751, 1771, 1790, 1810, 1841, 1856, 1915, 1934, 1948, 2050, 2186, 2198, 2246, 2274, 2332, 2343, 2356, 2359]. implementations [601, 617, 918, 1009, 1137, 1229, 1362, 1729]. Implementing [1581, 2188]. Implications [1293]. implicit [246, 1070, 1223, 1439]. improve [58, 177, 721, 978, 1032, 1240, 1556, 1630, 1677, 2071, 2207, 2258]. Improved [149, 236, 432, 646, 833, 888, 902, 959, 1005, 1266, 1504, 1738, 1740, 1764, 1804, 1954, 2024, 2152, 2196, 2247, 2261, 2265]. Improvement [246, 306, 343, 409, 881, 907, 960, 1138, 1501, 1641, 2039, 2305]. improvements [598, 665]. Improving [115, 127, 138, 203, 275, 280, 282, 333, 349, 410, 504, 602, 644, 706, 716, 802, 841, 953, 1075, 1083, 1224, 1291, 1342, 1358, 1395, 1397, 1415, 1774, 1780, 1785, 1906, 1919, 2031, 2034, 2090, 2101, 2138, 2236, 2312, 2328, 2371]. Improvised [1352]. impulse [1560, 1963]. impulsive [962]. In-advance [244]. In-memory [940, 1358]. in-network [593]. In-order [1126]. Inapproximability [2245]. Incentive [895, 1260, 1783, 2298]. Incentive-aware [1783]. including [943]. incomplete [370]. incompressible [222, 1639]. inconsistency [149, 490]. increase [1239]. Increasing [6, 547, 841, 1228, 1312]. Incremental [757, 772, 1502]. Independent [75, 113, 215, 318, 616, 755, 1016, 1031, 1117, 1389, 1452, 1483, 2368]. index [531, 604, 1368, 1428, 2092, 2312]. Indexable [2280]. indexes [774]. indexing [263, 412, 526, 1029, 1892]. Indic [2232]. indicator [540]. Indices [208]. indictment [1822]. indirect [1255]. individual [2137]. indoor [1539]. induced [1801]. industrial [1837, 1848, 1850, 1852, 1856, 1964, 2232]. industry [2042]. inelastic [1793]. inertial [1813, 1814]. Inference [64, 640, 684, 725]. InfiniBand [231, 297, 1381, 1968, 2305]. infinite [1963]. Influence [481, 1753, 2083]. Information [33, 79, 205, 348, 518, 638, 660, 669, 671, 847, 900, 943, 988, 1010, 1066, 1262, 1264, 1265, 1273, 1274, 1297, 1332, 1338, 1343, 1525, 1811, 1820, 1957, 2090, 2116, 2264, 2278]. informed [1658]. Infrastructure [27, 498, 591, 677, 886, 1066, 1244, 1252, 1254, 1298, 1410, 1862, 1872, 1877, 1879, 1929, 2205, 2216, 2310, 2375]. infrastructure-less [886, 1066]. infrastructures [683, 1062, 1593, 1874]. Infrequent [1882]. initialization [30]. initialized [1312]. initiation [1138]. injection [1336]. innovative [547]. input [735, 1080, 1277]. ins [1554]. insecurity [966]. insensitive [1823]. insertion [1863]. Insight [2123]. insights [2310]. inspection [2152, 2196]. inspired [456, 459, 569, 2102, 2144]. instance [383, 1200]. instances [1925, 2101]. institutes [1212]. institutions [1821, 2014]. Instruction [143, 275, 337, 715, 786, 1007, 1706, 2113]. instruction-level [1007]. instructions [928]. instrumentation [702, 1707, 1810]. Integer [1277, 1866, 2043, 2158, 2301]. Integrated [31, 372, 636, 1122, 1402, 1441, 1606, 1628, 1802, 1852, 1930, 1972, 2194, 2201]. Integrating [736, 748, 1573]. integration [550, 1329, 1797, 1967, 2163, 2361]. integrity [221, 892, 1697, 1984, 2212]. Intel [920, 1570, 1614, 1641, 1774, 1990, 2322, 2364]. intelligence [550, 967, 1872, 2197, 2268, 2269]. Intelligent [62, 245, 516, 520, 736, 911, 979, 1298, 1469, 1483, 1562, 1841, 2087, 2234, 2337]. intensity [2081]. intensive [385, 443, 1357, 1503, 1586, 1594, 1611, 1780, 2323]. Inter [571, 644, 649, 1069, 1113, 1424, 1570, 1899]. inter- [1570]. inter-Cloud [1069]. inter-nest [1899]. inter-network [571]. inter-node [644]. inter-operation [649]. inter-router [1113]. interaction [557, 951, 1538, 2075]. interactions [172, 1668]. Interactive [2, 285, 971, 992, 1219, 1395, 1547, 1579, 1668]. intercluster [619]. interconnected [2162]. Interconnection [34, 55, 102, 103, 232, 353, 636, 645, 711, 753, 761, 833, 848, 1018, 1035, 1038, 1178, 1192, 1255, 1285, 1319, 1376, 1380, 1384, 1623, 1724, 2367]. interconnections [637, 1455]. interconnects [359]. interest [147, 785]. interests [2234]. interface [1015, 1754]. interfaces [174, 598, 2232]. interference [288, 334, 571, 697, 1333, 2096, 2278]. interference-aware [288]. interferometric [377]. interlace [150]. intermediate [1064]. intermediate-data [1064]. Internet [415, 451, 708, 1213, 1214, 1326, 1460, 1463, 1470, 1491, 1546, 1580, 1670, 1795, 1808, 1836, 1837, 1839, 1843, 1845, 1847, 1848, 1853, 1855, 1858, 1861, 1878, 1927, 1943, 1959, 2197, 2264, 2346, 2347]. Internet-of-Things [1463]. interoperability [1908]. interoperable [270]. interpolating [678]. interposer [2317]. interposer-based [2317]. interprediction [477]. interpretation [328]. intersection [1693]. Interval [182, 1186, 2253]. Interval-based [1186, 2253]. intra [1424, 1426, 1570, 2039, 2078]. intra-prediction [1426, 2039]. intra-task [1570, 2078]. intrinsics [1713]. introducing [94]. Introduction [374, 433, 1356]. intrusion [503, 669, 680, 1316, 1360, 1473, 1540, 1890, 1917, 2252]. intrusive [1696]. intuitionistic [1970]. inundation [1786]. invariance [1208]. invariant [937]. invasive [139]. inventory [2069]. inverse [188, 1317]. inverses [726, 922, 1651]. inversion [239]. inverted [526]. Investigating [1605, 1957, 2199, 2288]. Investigation [714, 1216, 1883, 1953, 2200, 2313]. IO [197, 308]. IOPro [1026]. IoT [1392, 1456, 1463, 1465, 1469, 1545, 1654, 1669, 1687, 1699, 1760, 1805, 1806, 1840– 1842, 1844, 1849, 1856, 1861, 1887, 1889, 1901, 1934, 1942, 1944, 1948, 1955, 1967, 1976, 2000, 2145, 2163, 2200, 2214, 2217, 2228, 2325, 2350]. IoT-based [2145]. IoT-F2N [2325]. iota [1445]. iota-delta [1445]. IoV [1858]. IP [956, 967, 1306, 1883]. IP-based [967]. IPC [215]. IPS [127]. IPTV [347, 835]. IPv6 [220, 238, 812]. Iran [1491]. iron [2107]. Irregular [105, 230, 306, 1495, 1871]. irrevocability [1527]. IRST [153]. Island [2239]. iso [773, 1620]. iso-area [773]. iso-time [1620]. isolated [1960]. isolation [969, 1684]. isomorphism [1500]. issue [37, 89, 302, 374, 456, 1339, 1652, 2317]. issues [27, 445, 1108, 1528]. issuing [1543]. itemset [627, 1584]. itemsets [2010]. iterated [1322]. iteration [1315, 2069]. iterative [52, 157, 183, 367, 440, 737, 930, 933, 1402, 1571, 1574, 1722, 1774, 1835, 1912, 2224, 2231]. ITK [185].

J2M [1302]. Jaccard [716]. JACK [1574]. Jacobi [386, 1571]. Jacobi PCG [1425]. jam [878]. jamming [1143]. Java [90, 271, 297, 991, 1302]. Jaya [2035, 2077]. JEE [421]. JEE-based [421]. JEM [2043, 2059]. jitter [1303]. Job [13, 33, 38, 81, 115, 184, 196, 257, 361, 389, 436, 487, 613, 713, 1032, 1034, 1064, 1761, 1972, 2201, 2295, 2340, 2370]. job-shop [713]. jobs [126, 779, 794, 832, 1321, 1745, 1756, 1893, 1895]. Johnson [1828]. Join [774, 864, 1163, 1261, 1991]. joins [871]. Joint [611, 954, 1280, 1320, 1462, 1747, 1779, 1926, 2108]. joint-bitplane [1320]. joint-process [611]. JPEG [707]. junction [1555]. just [998]. just-in-time [998]. JVM [2168]. JXTA [40]. JXTA-Overlay [40].

K-DT [1725]. Kahn [428]. Kalman [166, 2275]. Kazahaya [1738]. Kernel [460, 702, 933, 950, 1173, 1235, 1284, 1574, 1830, 1978, 2183, 2212, 2268, 2269]. kernels [837, 1075, 1127, 1281, 1631, 2037, 2074, 2358, 2364]. Ketjenblack [2356]. Key [67, 91, 314, 323, 377, 430, 482, 665, 672, 736, 738, 844, 914, 988, 1138, 1177, 1195, 1283, 1288, 1675, 1740, 1857, 1868, 1923, 2141, 2150]. keying [1838]. keys [908, 1964]. keyword [1689]. kinetics [821]. KL GA [1180]. knapsack [957, 2286]. Knights [920]. KNL [2322, 2364]. kNN [1626]. knowledge [214, 1210, 1311, 1371, 1410, 2248]. knowledge-based [2248]. Korea [573, 1085]. Korean [1338]. KTS [1583]. KVM [1297].

label [2342]. labeling [727, 1365, 1451]. LABILE [350]. LACAV [406]. laceability [760]. lagged [611]. Lagrangian [861]. landform [946]. Landing [1433]. Landmark [2271]. landslides [579]. Language [751, 1387, 1713, 1715, 1746, 2124, 2353]. Language-based [1713]. Languages [296, 2232]. Laplacian [144]. Large [2, 4, 18, 55, 56, 88, 110, 112, 114, 156, 195, 206, 235, 261, 285, 298, 325, 408, 415, 442, 452, 496, 633, 653, 720, 763, 767, 830, 962, 1016, 1018, 1029, 1049, 1062, 1104, 1134, 1141, 1145, 1176, 1230, 1242, 1248, 1255, 1286, 1326, 1375, 1382, 1460, 1511, 1571, 1593, 1742, 1748, 1755, 1831, 1867, 1943, 1961, 2047, 2095, 2115, 2131, 2138, 2169, 2192, 2225, 2270, 2327, 2336, 2344, 2346]. large-capacity [18]. Large-scale [55, 56, 88, 110, 112, 156, 195, 206, 298, 408, 633, 653, 720, 767, 1018, 1104, 1134, 1141, 1145, 1176, 1230, 1248, 1255, 1286, 1326, 1375, 1382, 1460, 1593, 1742, 1755, 1831, 1867, 1943, 1961, 2095, 2115, 2131, 2138, 2169, 2192, 2327, 2344, 2346]. Last [853, 1234, 1242, 1798, 2306]. last-level [853, 1242, 1798, 2306]. Latency [338, 698, 733, 943, 1118, 1166, 1196, 1318, 2167, 2176, 2242, 2313, 2317]. Latency-aware [1118]. latency-constraint [1318]. lateral [2075]. lattice [949, 1231, 1800]. lattice-based [1231]. lattices [1365]. lava [576]. law [637]. layer [68, 163, 238, 622, 626, 656, 694, 1082, 1093, 1212, 1368, 1473, 1700]. layer-based [1368]. layered [20, 365, 1366, 1487]. layers [445]. layout [1413]. layouts [848, 1441]. Lazy [1191]. LBFT [1819]. LBP [1488]. LBVLR [2336]. LDPC [175]. LEACH [2104]. leader [657, 935]. leader-follower [935]. leaders [1164]. leadership [436]. leaf [2349]. Leakage [84, 982, 1111, 1838, 2141]. Leakage-efficient [84]. leakage-resilient [1838]. leakage-tolerating [982]. leaked [1525]. learned [1959]. learners [1610]. Learning [50, 64, 214, 237, 245, 345, 372, 383, 461, 488, 503, 699, 737, 748, 798, 815, 830, 1027, 1040, 1052, 1103, 1143, 1435, 1499, 1748, 1791, 1792, 1796, 1809, 1821, 1873, 1875, 1894, 1917, 1961, 2012–2014, 2021, 2092, 2101, 2132, 2136, 2137, 2234, 2252, 2260, 2270, 2277, 2285, 2294, 2313]. Learning-based [798, 1052, 1796, 2021, 2285]. Least [1159, 1431]. least-square [1159]. Lee [1726]. left [2072]. legacy [508]. legal [348]. Lemma [1828]. length [800, 1732, 1733]. lens [577]. less [886, 1066]. Level [8, 58, 126, 139, 143, 240, 258, 339, 343, 407, 483, 614, 711, 718, 785, 806, 822, 853, 940, 978, 1007, 1057, 1085, 1190, 1234, 1242, 1315, 1357, 1401, 1564, 1759, 1798, 1810, 1844, 1855, 1921, 2077, 2124, 2169, 2172, 2183, 2249, 2306, 2327, 2336, 2337]. Levels [372]. leverage [2149]. Leveraging [300, 598, 621, 1157, 1343, 1527]. lexIcaL [350]. lexicon [2105]. Library [133, 189, 261, 343, 472, 475, 605, 1021, 1074, 1096, 1394, 1574, 1711, 1933, 2174, 2324]. license [558]. LiDAR [1433, 2203]. Lie [198]. life [547, 659, 1272, 1514]. lifelogging [1543]. Lifetime [250, 559, 851, 907, 1202, 1739, 1749, 2177]. Light [149, 377, 381, 416, 702, 1076, 1813, 1814]. light-load [416]. Light-weight [702]. Lightweight [787, 853, 1195, 1284, 1465, 1669, 1670, 1687, 1934, 1974, 2162, 2210]. like [1675]. likelihood [293, 499]. likelihood-based [499]. likely [841]. limit [541]. limitations [919]. Limited [381, 671, 1448, 1604]. Lindenstrauss [1828]. line [99, 766]. Linear [140, 156, 184, 261, 320, 440, 523, 596, 763, 834, 1049, 1184, 1243, 1317, 1425, 1455, 1474, 1480, 1501, 1512, 1571, 1588, 1591, 1725, 1859, 1866, 2003, 2120, 2122, 2226, 2301]. linear-time [1591]. lines [1257, 1258, 1634]. linguistic [2271]. link [350, 1445, 1903, 2307]. link-fault-tolerant [1903]. linkage [2038, 2275]. linkages [708]. linked [1366]. links [154, 248, 1513]. Linux [1284]. List [196, 1002, 1905, 2170]. List-based [1002, 1905, 2170]. Listing [2225]. lists [2351, 2372]. Live [39, 533, 703, 1870, 2262]. live-migration [703]. liveness [1839]. Living [64]. LLC [179, 1661, 2209]. Load [34, 45, 60, 141, 183, 208, 213, 217, 243, 252, 298, 305, 391, 416, 560, 585, 604, 728, 752, 783, 798, 807, 814, 815, 1133, 1188, 1236, 1237, 1351, 1418, 1489, 1588, 1593, 1683, 1699, 1796, 1799, 1819, 1898, 1913, 1950, 1991, 2037, 2071, 2128, 2169, 2304, 2314, 2336]. Load-balanced [560, 1796, 2169]. load-balancing [60, 1819, 1913]. Loadbot [1699]. loading [1590]. loads [339]. Local [262, 1016, 1322, 1434, 1468, 1888, 2281, 2342]. Locality [24, 94, 228, 280, 334, 504, 599, 868, 1032, 1242, 1711, 1899, 2320]. Locality-aware [868, 1032, 1242, 2320]. locality-flexible [1711]. Localization [451, 502, 511, 830, 1496, 2230]. localized [654]. located [1787]. Locating [570]. Location [187, 205, 300, 511, 884, 1011, 1066, 1223, 1252, 1262, 1264, 1274, 1391, 1554, 1643, 2084, 2091]. Location-aided [1262]. Location-based [884, 1011, 1223, 1554, 2084]. location-dependent [300]. location-sensitive [2091]. lock [898, 1979, 2363]. lock-aware [1979]. lock-based [2363]. log [1593, 2142]. log-in [2142]. logarithm [109]. logging [1109]. logic [380, 520, 1116, 1635, 1650, 1755, 1914, 1936, 2162]. logical [2367]. Logically [1960]. Loginson [1593]. logs [423]. Long [557, 1017, 1950, 2314]. long-memory [1017]. long-range [557]. longest [582]. longevity [391]. Lookahead [372]. lookup [1997]. lookups [248]. loop [193, 209, 268, 806, 1641, 1899]. loop-level [806]. loops [1010, 1315]. loosely [257]. Loss [753, 1664, 1665, 1863]. Loss-aware [1664, 1665]. losses [1370]. Low [58, 210, 554, 698, 747, 1023, 1079, 1153, 1166, 1416, 1532, 1553, 1738, 1758, 1813, 1814, 1891, 1963, 2167, 2242, 2296, 2330, 2339, 2369]. low-complexity [2369]. Low-cost [1166, 1553, 1738, 1813, 1814, 2330, 2339]. low-fidelity [1153]. low-latency [698, 2167, 2242]. low-level [58]. low-overhead [210, 1023, 1532]. low-pass [1963]. low-power [1079, 1416, 2296]. Lower [2344]. Lower-bound [2344]. Lowering [1259]. LP [1746, 2120]. LP-based [1746]. LP-WSC [2120]. LPF [1963]. LR [982]. LR-FEAD [982]. LRnLA [2355]. LRU [2177]. LTE [654, 1145, 1270, 1542]. LTE/EPC [1542]. LU [1526]. LuNA [301, 829, 1438, 1447]. Lyapack [605, 1020]. Lyapunov [1020].

M [707]. M-JPEG [707]. M2T [2371]. Machine [64, 123, 303, 468, 697, 703, 721, 748, 846, 881, 975, 990, 1108, 1204, 1252, 1363, 1374, 1499, 1673, 1782, 1783, 1801, 1821, 1873, 1875, 1919, 1924, 1929, 1979, 1995, 2073, 2099– 2101, 2132, 2136– 2138, 2173, 2238, 2248, 2270, 2286, 2309, 2313]. machines [291, 292, 367, 428, 528, 577, 681, 842, 1241, 1286, 1506, 1917, 1920, 2076, 2262, 2323]. Mahout [1672]. MAIDS [1887]. main [1672, 1770, 1894]. maintain [2263]. maintenance [25, 59]. major [1084]. Makara [1344]. makespan [206, 367, 1321]. making [291, 647, 1815, 2034]. malicious [390, 513, 1095, 1730]. malware [983, 1803, 2249]. manage [2299]. managed [105]. Management [28, 31, 40, 54, 59, 91, 200, 269, 274, 303, 305, 314, 333, 384, 411, 443, 483, 512, 517, 518, 520, 556, 624, 636, 638, 652, 785, 792, 845, 872, 878, 909, 912, 1045, 1051, 1052, 1069, 1190, 1195, 1241, 1271, 1410, 1505, 1506, 1531, 1532, 1542, 1581, 1649, 1674, 1684, 1703, 1730, 1743, 1760, 1849, 1891, 1951, 1972, 1973, 1993, 2005, 2121, 2162, 2201, 2209, 2218, 2220, 2240, 2318, 2346]. manager [1218, 2373]. managers [1595]. Managing [299, 688, 764, 768, 1225]. Mandi [536]. MANET [224, 1263–1265, 1274]. MANETs [405, 464, 1650]. maneuvering [153]. manipulation [1068]. manipulator [2232]. manufacturing [878, 1846]. many [164, 528, 587, 634, 643, 795, 798, 921, 1023, 1099, 1118, 1153, 1197, 1427, 1490, 1526, 1532, 1583, 1587, 1731, 1759, 1998, 2004, 2049, 2076, 2254, 2362]. many-core [164, 528, 587, 634, 643, 795, 798, 921, 1023, 1099, 1118, 1153, 1197, 1427, 1490, 1526, 1532, 1587, 1731, 1759, 2049, 2076]. many-cores [1583]. many-field [2254]. many-task [2004, 2362]. many-to-many [1998]. manycore [953, 1079, 1181, 1433, 1568, 1716, 1719]. manycores [1238]. map [91, 409, 1283, 1602, 2178]. map-based [91]. map-reduce [409]. MAP-SDN [1602]. Mapping [178, 409, 529, 568, 576, 644, 696, 1033, 1071, 1180, 1275, 1306, 1318, 1373, 1423, 1495, 1513, 1583, 1587, 1705, 1749, 1825, 1863, 2026, 2189]. mappings [1940]. MapReduce [428, 530, 646, 649, 727, 835, 862, 864, 871, 973, 986, 1064, 1156, 1163, 1261, 1302, 1307, 1321, 1383, 1451, 1499, 1500, 1558, 1625, 1721, 1799, 1895, 1986, 1991, 2294, 2310, 2340]. MapReduce-based [862, 1163, 1625, 1721]. maps [629, 1492, 1817, 1857]. margin [1309, 1573]. maritime [1466, 1467]. market [431, 536, 692, 976, 1608, 2256]. market-oriented [431]. Markov [371, 720, 1506, 1716]. MARTE [1161]. mass [819, 821, 1437]. mass-action [821]. mass-spring [819, 1437]. massive [190, 646, 688, 727, 1808, 2054, 2147, 2361, 2369]. Massively [452, 527]. Master [607, 1189]. Master-based [1189]. master-slave [607]. match [859, 1348]. matching [79, 152, 444, 676, 877, 894, 1337, 1346, 1732, 1733, 1816, 1956, 2008, 2009, 2122, 2280]. mathematical [1567, 1728, 1937, 2027]. Matlab [2045]. matrices [239, 258, 959, 1312, 1589]. Matrix [77, 130, 162, 188, 228, 368, 376, 454, 479, 643, 649, 673, 808, 937, 1115, 1175, 1176, 1246, 1411, 1412, 1439, 1582, 1735, 1906, 1933, 1986, 2045]. matrix-delayed [1735]. matrix-free [1439]. matrix-matrix [228, 1246]. matrix-vector [162, 454]. max [1507]. maximal [229, 2225]. maximally [2341]. maximization [559, 1200, 1390, 1608, 2022]. maximized [696]. Maximizing [796, 851, 948, 1139, 1700]. maximum [80, 293, 499, 1015, 1937]. maximum-density [80]. MBTAC [1737]. MCDM [651, 1628]. MCV [1536]. MCV-based [1536]. MDV [2039]. Mean [144, 552, 2273, 2374]. mean-curvature [144]. Mean-Shift [2273]. Means [50, 770, 867, 973, 1137, 1430, 1721, 1769, 1830, 2017, 2054, 2123, 2295, 2349]. measure [1525, 1622, 1906, 2143]. measured [1017, 1915]. Measurement [704, 733, 806, 984, 1095, 1909, 1958, 2107]. measures [123, 884, 1695, 2058, 2195]. measuring [110]. mechanism [210, 215, 298, 338, 406, 426, 496, 498, 541, 558, 646, 655, 756, 795, 815, 835, 881, 895, 982, 1035, 1101, 1122, 1182, 1262, 1604, 1928, 1974, 2171, 2268, 2269, 2298, 2338, 2344]. mechanisms [61, 233, 349, 410, 617, 723, 1068, 1130, 1606, 1802]. media [3, 129, 288, 518, 544, 1553, 1610, 2119]. median [1607]. Medical [2, 132, 762, 1030, 1084, 1137, 1154, 1817, 2013, 2237]. medicines [2091]. medium [49, 2153]. meet [2027]. meeting [977]. membership [1750]. Membrane [2102]. memories [137, 501, 1160, 1276]. Memory [18, 77, 105, 114, 170, 173, 262, 265, 286, 316, 369, 435, 443, 470, 471, 481, 504, 528, 531, 600, 670, 766, 784, 790, 870, 899, 940, 943, 1006, 1017, 1054, 1152, 1182, 1191, 1227, 1293, 1354, 1358, 1374, 1406, 1441, 1527, 1642, 1652, 1659, 1661, 1672, 1726, 1728, 1748, 1754, 1770, 1791, 1826, 1853, 1867, 1894, 1902, 1935, 1950, 1969, 1978, 2176, 2187, 2190, 2299, 2302, 2314, 2316]. Memory-based [316, 1826]. memory-efficient [670, 870]. menace [1062]. Merge [1256]. mergesort [2288]. merging [219, 465, 549, 2025]. Mesh [14, 38, 142, 191, 231, 246, 438, 516, 560, 857, 863, 923, 1033, 1124, 1136, 1180, 1285, 1306, 1373, 1604, 1651, 1667, 1700, 1992, 2076, 2307]. Mesh-Adaptive [1651]. mesh-based [1033, 1124, 1136, 1306, 1373]. mesh-connected [38, 857, 1604]. Mesh-of-Torus [1992]. mesh-of-tree [1180]. meshes [118, 582, 724, 2345]. message [90, 233, 240, 255, 271, 297, 926, 1015, 1066, 1109, 1520, 1812]. message-driven [926]. message-passing [90, 271, 297, 1520]. messages [448]. messaging [1552]. meta [1747, 2292, 2293, 2312]. meta-heuristic [2292, 2293, 2312]. meta-heuristics [1747]. metacomputing [326]. Metacube [55]. metadata [135, 283, 483, 1240, 1479, 1530, 2005, 2318]. metaheuristic [1039, 1344, 1602, 1644, 2286]. metaheuristics [173, 364, 382, 463, 1046, 2366]. metering [677, 1298]. MeterPU [1909]. method [76, 100, 114, 134, 140, 169, 186, 265, 331, 343, 356, 370, 377, 386, 419, 422, 451, 474, 502, 612, 652, 666, 682, 703, 709, 715, 763, 807, 828, 834, 847, 861, 951, 952, 1076, 1149, 1159, 1264, 1274, 1284, 1291, 1332, 1333, 1498, 1564, 1565, 1571, 1575, 1600, 1614, 1616, 1620, 1633, 1651, 1658, 1701, 1703, 1707, 1758, 1772, 1779, 1800, 1894, 1917, 1971, 1980, 1984, 1988, 2061, 2088, 2134, 2137, 2146–2148, 2152, 2172, 2194, 2199, 2203, 2260, 2264, 2300, 2309, 2329, 2339, 2343]. method-based [2148]. method-level [343]. methodology [228, 577, 766, 808, 896, 1045, 1115, 1246, 1350, 1351, 1623, 2291, 2302]. Methods [50, 161, 301, 317, 320, 337, 466, 564, 689, 789, 1054, 1074, 1352, 1423, 1558, 1628, 1764, 1876, 2046, 2202, 2231, 2237]. methylation [2064]. metric [110, 217, 350, 773, 1261, 2074]. metrics [450, 1075]. metropolitan [1833]. MFiX [1835]. MFReS [1953]. MIC [1443]. micro [671, 860, 1543]. micro-benchmark [860]. micro-movement [1543]. Microarchitectural [920]. microcalcification [2260]. microdata [1886]. microeconomics [2335]. microkernel [1568]. microkernel-based [1568]. microrheology [1421]. Microring [1490, 1864]. microservices [1781]. microservices-based [1781]. middleware [42, 241, 621, 1223, 2217]. migration [118, 591, 703, 787, 792, 795, 899, 940, 1034, 1108, 1172, 1204, 1271, 1285, 1862, 1870, 2173, 2240, 2262, 2295]. migration-replication [1172]. MIH [876]. MIH-based [876]. military [2278]. MIMD [309]. MIMO [597, 930, 1021, 1808, 1965, 2047, 2369]. MIMOPack [1021]. min [1507]. min-max-based [1507]. mined [695]. Minimal [248, 2054, 2115]. minimisation [1001]. minimization [167, 834, 1773, 2168, 2306]. minimize [1321, 2011]. Minimizing [328, 332, 1196, 1253]. minimum [237, 313, 401, 488, 630, 854, 873, 2245]. mining [32, 229, 525, 627, 668, 1048, 1227, 1305, 1311, 1352, 1370, 1481, 1546, 1584, 1703, 1882, 2010, 2115, 2241, 2264, 2272, 2365]. misdetection [2150]. Miss [2209]. Miss-aware [2209]. mission [693]. mission-critical [693]. mitigating [1134]. Mitigation [791, 1361, 1786]. MITM [352]. mix [1081, 1980]. mixed [252, 1795, 2158]. mixed-integer [2158]. mixing [1081]. ML605 [1704]. MLC [2285]. MLN [303]. MLP [112]. MM [615]. MMOFPS [178]. MMSE [166]. MMT [142]. Mobile [48, 49, 67, 98, 101, 112, 204, 205, 220, 238, 269, 284, 314, 345, 351, 352, 439, 468, 492, 494, 497, 544, 610, 666, 667, 672, 705, 714, 772, 777, 787, 812, 830, 844, 854, 892, 908, 914, 931, 944, 956, 958, 967, 976, 1043, 1089, 1094, 1095, 1131, 1132, 1152, 1177, 1193, 1201, 1211, 1222, 1275, 1289, 1314, 1333, 1336, 1419, 1471, 1472, 1476, 1477, 1528, 1553, 1563, 1586, 1590, 1619, 1750, 1824, 1834, 1857, 1897, 1904, 1923, 1939, 1947, 1953–1955, 1994, 2056, 2084, 2142, 2149, 2171, 2205, 2206, 2214, 2256, 2268, 2269, 2326, 2331, 2336, 2337, 2375]. mobile-cloud [705]. Mobility [426, 963, 1542, 1994, 2330]. Mobility-aware [1994]. Möbius [616, 1031]. modal [1330]. mode [1113, 1250, 1737, 1793, 2143]. Model [33, 65, 144, 154, 193, 199, 285, 341, 354, 356, 422, 439, 467, 503, 514, 579, 583, 615, 630, 658, 710, 717, 733, 740, 751, 809, 814, 817, 818, 849, 873, 943, 945, 947, 980, 985, 1024, 1088, 1091, 1099, 1106, 1175, 1179, 1188, 1201, 1212, 1216, 1243, 1390, 1438, 1455, 1462, 1508, 1524, 1531, 1536, 1581, 1663, 1671, 1715, 1728, 1735, 1746, 1763, 1807, 1819, 1860, 1884, 1902, 1975, 1985, 1994, 2004, 2014, 2021, 2092, 2125, 2144, 2151, 2163, 2197, 2219, 2225, 2264, 2325, 2356, 2359]. Model-based [980, 1099, 1663, 2151]. model-directed [1088]. Modeling [28, 107, 117, 233, 348, 388, 421, 454, 460, 607, 632, 697, 725, 832, 919, 984, 1010, 1017, 1040, 1067, 1186, 1363, 1391, 1404, 1442, 1445, 1516, 1526, 1614, 1745, 1746, 1789, 1892, 2082, 2117, 2124, 2205, 2231, 2264, 2360, 2375]. Modelling [9, 44, 235, 452, 1193, 1411, 1615, 2023]. Models [41, 177, 307, 324, 453, 460, 762, 797, 821, 976, 1000, 1313, 1437, 1476, 1594, 1831, 1870, 1908, 2027, 2074, 2180]. Modern [310, 1067, 1269, 1383, 1596, 2031]. modernization [1696]. Modified [356, 370, 572, 1055, 1216, 1316, 1468, 1956, 2248]. modifier [770]. MODM [658]. modular [870, 901, 1281]. modulation [690]. modulator [381]. module [1426]. modulo [247]. moisture [2219]. Molecular [109, 122, 323, 630, 642, 873, 1718, 1731, 2155, 2157, 2351, 2372]. monitor [1562]. monitoring [10, 223, 408, 496, 624, 1367, 1488, 1694, 1807, 1972, 2086, 2212]. monotonic [256]. Monte [140, 160, 1159]. mood [553]. MooseFS [1508]. MOPT [2170]. morphological [1428]. MorphoSys [192]. Mosaic [100]. Mosaic-Net [100]. MoT [232, 1180]. motif [701, 865, 2322]. motifs [745, 955]. Motion [150, 707, 891, 970, 1220, 2043, 2088, 2273, 2374]. motion-compensated [150]. motor [1474]. mountaineering [904]. mouth [2083].

Movement [205, 748, 1543, 1663]. movements [868]. Moving [121, 312, 579, 671, 1468, 1530]. MPDATA [1442, 1614]. MPEG [349]. MPI [4, 164, 193, 197, 207, 239, 268, 270, 272, 356, 437, 641, 642, 918, 940, 1304, 1384, 1401, 1417, 1446, 1492, 1517]. MPI-based [1384]. MPI-IO [197]. MPI/OpenMP [642, 1517]. MPJ [271]. MPR [224]. MPSoC [295, 360, 2243]. MPSoCs [1513]. MR [112]. MRF [185]. MRI [1103]. MSD [1292]. MSNP [2082]. MSP430 [1708, 1751]. MTIDS [1890]. MTPA [121]. MU [1808]. mule [2336]. Multi [23, 65, 87, 102, 119, 124, 142, 153, 164, 168, 184, 223, 224, 277, 287, 295, 335, 349, 383, 398, 410, 432, 436, 458, 463, 531, 560, 618, 621, 637, 658, 674, 694, 775, 783, 823, 828, 865, 867, 908, 928, 945, 946, 1019, 1035, 1097, 1110, 1111, 1133, 1153, 1163, 1181, 1182, 1241, 1270, 1279, 1283, 1300, 1301, 1310, 1366, 1403, 1404, 1409, 1411, 1416, 1425, 1519, 1521, 1529, 1539, 1564, 1566, 1633, 1648, 1781, 1857, 1868, 1890, 1893, 1962, 1965, 1969, 1971, 2015, 2040, 2049, 2067, 2089, 2096, 2128, 2139, 2147, 2149, 2169, 2182, 2189, 2209, 2210, 2274, 2293, 2303, 2327, 2331, 2365]. multi [90, 116, 126, 130, 193, 222, 253, 272, 331, 428, 430, 532, 577, 629, 644, 647, 733, 792, 801, 822, 859, 898, 925, 948, 1056, 1096, 1155, 1168, 1271, 1282, 1308, 1405, 1434, 1474, 1493, 1494, 1499, 1533, 1666, 1671, 1719, 1748, 1782, 1831, 1866, 1887, 2024, 2063, 2077, 2087, 2253, 2260, 2265, 2280, 2301, 2341]. multi- [164, 1153, 1719, 2049]. multi-/manycore [1719]. multi-access [1539]. multi-agent [23, 1887]. Multi-AGVs [2265]. multi-algorithm [629]. multi-attribute [783, 2341]. multi-BSP [733]. Multi-cache [1521]. multi-channel [1409, 2087]. multi-class [2089]. multi-cloud [1056, 1533, 1781, 1866, 2128, 2149, 2301]. multi-cloudlet [1494]. multi-cluster [116, 184]. Multi-CMP [124]. Multi-core [87, 90, 130, 193, 223, 272, 295, 335, 410, 428, 432, 458, 618, 637, 644, 792, 801, 808, 823, 865, 867, 898, 925, 928, 945, 1111, 1115, 1279, 1403, 1566, 1648, 1671, 1831, 1969, 2015, 2040, 2182, 2209, 2303, 2365]. multi-CPU [948, 1019]. multi-criteria [647, 1241]. Multi-device [621, 1096, 2067]. multi-dimensional [531, 1666]. Multi-domain [436]. Multi-environmental [463]. multi-GPU [222, 253, 822, 828, 946, 948, 1019, 1155, 1308, 1310, 1411, 1425, 1499, 1748]. multi-granularity [1493]. multi-hop [119, 398, 560]. multi-instance [383]. multi-layer [694]. multi-layered [1366]. Multi-level [126, 1564, 2077, 2169, 2327]. multi-match [859]. multi-Mesh [142]. multi-mobile [2331]. multi-model [65, 945]. multi-multicast [1405, 1434]. multi-node [331]. Multi-Objective [168, 658, 775, 1282, 1416, 1474, 1529, 1782, 1965, 2063, 2189, 2253, 2293, 2301]. multi-operand [1110]. multi-parameter [1633]. multi-party [2139]. multi-physics [2040]. multi-point [224]. multi-precision [2274]. multi-privileged [532]. multi-processor [1271]. multi-processors [277]. Multi-query [1519]. Multi-queue [1893]. multi-radio [1270]. multi-resource [287]. Multi-scale [2089, 2260]. multi-segment [2280]. multi-sensor [2147]. multi-server [430, 674, 908, 1283, 1857, 1868, 1962]. Multi-Sort [142]. multi-source [1965]. Multi-stage [102, 1035]. Multi-step-ahead [1133]. multi-streaming [349]. multi-tag [2210]. multi-tape [577]. multi-target [153]. Multi-tenant [1890]. multi-threaded [1181, 1182, 1404]. multi-thresholds [2024]. multi-tier [1097]. multi-user [2096]. multi-view [1971]. multi-way [1163]. multi-workflow [1168]. multiagent [214]. multiblock [2118]. multibody [2055]. multicast [26, 231, 401, 404, 448, 465, 609, 622, 802, 1072, 1405, 1434, 1795, 1939]. multicasting [1965]. multichannel [26, 148, 288, 362]. multicomputer [9, 313, 321]. multicomputers [38]. Multicore [5, 122, 136, 137, 158, 163, 182, 268, 291, 299, 365, 393, 457, 587, 596, 597, 629, 721, 733, 789, 794, 803, 853, 932, 946, 951, 986, 1009, 1079, 1121, 1136, 1165, 1174, 1203, 1256, 1397, 1411, 1449, 1519, 1614, 1705, 1716, 1752, 1801, 2047, 2363]. multicore-GPU [629]. multicore/multiprocessor [5]. multicores [1242, 1617, 1898, 2296]. multicriteria [1901, 2034]. Multidimensional [263, 550, 711, 957, 2031, 2240, 2272]. Multidisciplinary [2135]. MultiGPU [606]. multigrid [726, 1639, 2118]. multihop [622]. Multilayer [907, 1178]. Multilevel [687, 1125, 1317, 1437, 1731, 2156]. multilinear [545]. multilingual [1880]. Multimedia [36, 68, 93, 97, 344, 346, 522, 537, 551, 553, 556, 767, 916, 1092, 1456, 1461, 1579, 1649]. multimethod [2158]. Multiobjective [701, 1461, 1832]. multipath [39, 677]. Multiperiod [987]. multiplayer [1089]. Multiple [184, 229, 312, 372, 457, 476, 501, 523, 602, 837, 914, 1047, 1080, 1162, 1164, 1310, 1321, 1323, 1513, 1538, 1722, 1727, 1812– 1814, 1886, 1964, 2016, 2051, 2203, 2259]. multiple-keys [1964]. Multiple-Lookahead-Levels [372]. multiple-workflow [1047, 2016]. multiplexer [1865, 2251]. multiplexing [375, 1142, 1980]. multiplication [228, 368, 376, 454, 643, 649, 670, 808, 1115, 1116, 1176, 1231, 1246, 1735]. multiplicative [1067]. multiplier [1110, 1277]. multipliers [1635]. Multipole [466, 612]. Multipopulation [2077]. Multipopulation-based [2077]. multiprocessing [2184, 2185]. multiprocessor [5, 174, 276, 394, 817, 1012, 1749, 2189, 2354]. multiprocessors [157, 201, 414, 508, 584, 634, 996, 1172, 1604]. multiproduct [1440]. multiresolution [2, 569]. Multiround [1739]. multisensor [1332]. multisensory [1463]. multiserver [1804]. multisource [465]. multisplitting [1049]. Multistage [103, 833, 1178, 1319]. Multitarget [146]. multitasking [2218]. Multitenant [2005]. Multithreaded [53, 467, 639, 1514, 1737]. multithreading [744, 2191]. multivariate [1926]. multiword [175]. Music [1081, 1927, 2136]. musical [1419, 2028]. mutant [1803]. muti [1420]. muti-objective [1420]. Mutual [729, 905, 960, 1218, 1465, 1592, 1687, 1693, 1843, 2116, 2210]. Mutually [75, 318, 1452]. MWC [2149].

naive [780, 1887, 2261]. namespace [2299]. NAND [501, 626, 1894]. nanometric [225]. nanoscale [1823]. nanostructures [1936]. Nash [486]. National [1469]. Natural [143, 208, 1208]. nature [252, 456, 459]. nature-inspired [456, 459]. navigation [972, 2208]. NBC [1887]. NBC-MAIDS [1887]. nD [959]. Near [94, 172, 1548]. Near-reversible [1548]. Nearest [265, 780, 1667]. negative [1104]. Negotiation [86, 692]. Neighbor [780, 1217, 1262, 1265, 1667, 2351, 2372]. neighborhood [346, 1453, 2367]. Neighbors [265]. Nekbone [1362]. nest [1899]. Nested [165, 619, 1315]. nested-parallel [165]. nesting [112]. net [100, 685, 2216]. net-based [2216]. Nets [107, 259, 363, 1161, 1678, 1777]. Network [33, 87, 91, 102, 103, 112, 127, 174, 202, 203, 223, 227, 232, 235, 250, 311, 333, 338, 353, 383, 391, 408, 442, 554, 561, 571, 573, 593, 598, 611, 650, 654, 698, 708, 719, 743, 745, 831, 833, 863, 876, 923, 962, 1018, 1081, 1085, 1086, 1124, 1136, 1166, 1180, 1192, 1197, 1211, 1245, 1248, 1250, 1272, 1288, 1290, 1298, 1305, 1318, 1349, 1354, 1360, 1361, 1373, 1379, 1384, 1485, 1490, 1542, 1550, 1556, 1564, 1677, 1694, 1700, 1757, 1758, 1796, 1815, 1853, 1897, 1900, 1919, 1960, 1971, 2013, 2089, 2102, 2112, 2131, 2148, 2151, 2153, 2165, 2197, 2216, 2252, 2267, 2289, 2307, 2317, 2325]. network [14, 18, 37, 45, 50, 69, 284, 339, 358, 373, 408, 414, 491, 502, 516, 526, 527, 651, 711, 781, 848, 851, 980, 1033, 1067, 1104, 1112, 1201, 1262, 1333, 1335, 1378, 1510, 1565, 1724, 1753, 1778, 1791, 1819, 1871, 1917, 2020, 2242, 2245, 2268, 2269, 2275, 2293, 2367]. network-aware [1919]. Network-based [37, 383, 1564, 2289]. network-on-chip [87, 311, 781, 863, 923, 1033, 1124, 1136, 1180, 1197, 1318, 1373, 1490, 1510, 1819, 2020, 2307, 2317]. network-on-chips [1871]. networked [245, 518, 1008, 1060, 1140, 1747, 1760]. networked-distributed [245]. networking [210, 400, 1811, 1829]. Networks [10, 43, 49, 55, 100, 112, 119, 134, 205, 238, 251, 350, 368, 375, 390, 398, 399, 408, 438, 489, 512, 544, 559, 560, 571, 622, 623, 638, 657, 658, 667, 728, 734, 752, 761, 771, 830, 838, 844, 848, 856, 870, 886, 907, 956, 963, 968, 1028, 1035, 1038, 1065, 1066, 1090, 1094, 1095, 1129, 1139, 1141, 1178, 1192, 1194, 1198, 1202, 1254, 1270, 1324, 1325, 1355, 1376, 1389, 1504, 1623, 1647, 1656, 1657, 1699, 1701, 1727, 1739, 1762, 1833, 1844, 1852, 1863, 1904, 1932, 1966, 1976, 2096, 2104, 2109, 2112, 2134, 2207, 2213, 2224, 2230, 2287, 2298, 2329, 2331, 2369]. networks [4, 9, 25, 30, 34, 39, 41, 47, 48, 68, 73, 78, 220, 230, 231, 233, 288, 314, 324, 345, 346, 387, 401, 402, 406, 428, 496, 497, 511, 514, 517, 520, 522, 547, 569, 636, 645, 647, 652, 653, 656, 699, 741, 749, 753, 810, 851, 914, 976, 1036, 1059, 1113, 1133, 1135, 1143, 1145, 1147, 1151, 1157, 1177, 1195, 1219, 1236, 1255, 1306, 1319, 1345, 1381, 1405, 1448, 1453, 1471, 1472, 1496, 1561– 1563, 1653, 1664, 1665, 1676, 1679, 1681, 1683, 1700, 1702, 1730, 1755, 1763, 1876, 1881, 1939, 1954, 1990, 1992, 2110, 2121, 2126, 2169, 2211, 2222, 2245, 2312, 2336]. networks [6, 154, 226, 1461, 1529]. networks-on-chip [78, 226, 1113, 1151, 1157, 1236, 1306, 1863]. neural [227, 324, 383, 442, 1104, 1129, 1245, 1349, 1354, 1562, 1735, 1762, 1791, 1900, 1990, 2013, 2134, 2242]. neuro [684]. neuro-fuzzy [684]. neutral [512]. Neville [164]. news [1809]. news-in-education [1809]. Newton [737, 2050]. Next [566, 1537, 1554, 1555, 1885, 2291]. Next-generation [1885, 2291]. NFC [903]. NFV [1971, 2248]. nilpotent [198]. NIST [1888, 2281]. NLI [208]. NMF [1906]. NMF-based [1906]. NML [1257, 1258]. NN [412]. NNMF [932]. NNMFPACK [1075]. no [1148, 2172]. no-reference [1148]. no-threading [2172]. NOC [410, 529, 584, 698, 723, 919, 1122, 1189, 1493, 1513, 1532, 1568, 1583, 1643, 1779, 1864, 2218, 2313]. NoC-assisted [410]. NOC-based [529, 1513, 1568, 1583, 2313]. NoCs [191, 1126, 1385, 1928]. Node [6, 204, 331, 513, 644, 673, 782, 1040, 1139, 1389, 1391, 1545, 1904, 1976, 2287, 2312, 2328, 2341]. node-based [1976]. node-disjoint [204]. node-independent [1389]. nodes [1654, 1730, 2011, 2155]. nodule [1573]. Noise [148, 572, 936, 2052]. noise-compensated [148]. noises [1560]. noisy [148]. Non [18, 84, 139, 335, 367, 891, 1018, 1228, 1315, 1350, 1386, 1660, 1696, 1756, 1826, 1904, 2046, 2105, 2109, 2177, 2345, 2373]. non-blocking [1018]. Non-clairvoyant [1756]. non-conforming [891]. non-continuous [18]. non-convex [1904, 2109]. non-English [2105]. non-exclusive [335]. non-increasing [1228]. Non-intrusive [1696]. non-invasive [139]. non-makespan [367]. non-manager [2373]. non-planar [2046]. non-preemptive [1386, 1660]. non-state [84]. non-stationary [1350]. non-uniform [1315, 2345]. non-volatile [1826, 2177]. nondedicated [183]. Nonflat [711]. Nonlinear [158, 189, 242, 339, 737, 1005, 1266, 1722]. nonlocal [2017]. nonnegative [1906]. nonuniform [625]. nonvolatile [2299]. Normal [216]. normalization [1346].

Normalized [631, 1348]. Nornir [428]. NoSQL [1618, 2338]. Note [202, 2374, 2375]. nothing [1736]. Novel [50, 93, 217, 270, 308, 358, 373, 401, 555, 583, 669, 682, 707, 773, 781, 862, 872, 889, 897, 1036, 1039, 1110, 1209, 1221, 1241, 1304, 1306, 1340, 1369, 1372, 1465, 1474, 1556, 1565, 1635, 1674, 1698, 1730, 1851, 1865, 1869, 1940, 2017, 2178, 2202, 2236, 2251, 2255, 2262, 2308, 2312, 2321, 2335]. NP [1353]. NP-hard [1353]. NSGA [1517, 2293]. NSPRING [1346]. NTB [1287]. NUCA [410, 505, 723, 755]. nuclear [1707, 2195]. NUMA [162, 1100, 1361]. number [211, 581, 2174]. numbers [915, 1171, 1511]. Numerical [133, 301, 458, 603, 611, 1130, 1438, 2111]. NVIDIA [454, 950, 1127, 2316]. NVM [1826].

O [335, 776, 788, 1022, 1026, 1088, 1134, 1477, 1509, 1663, 2358]. O2WebCL [1087]. OaaS [549]. OAFPM [2241]. obfuscation [304]. Object [36, 63, 296, 712, 1076, 1326, 1468, 1849, 2146, 2203, 2227]. object-based [1326]. object-oriented [296]. Objective [168, 658, 775, 1000, 1282, 1300, 1301, 1416, 1420, 1474, 1529, 1782, 1965, 2063, 2189, 2253, 2293, 2301]. objectives [588]. objects [614, 2125]. Oblivious [444, 1051]. observations [148, 538]. observer [78]. obtaining [686]. occupation [2233]. occurrence [214, 258]. Ocean [2015]. OCR [2090]. Odd [113, 923, 1591, 1819]. odd-even [923]. odd-sized [1591]. ODE [2361]. OFDM [2116]. off [99, 1282, 1590, 1789]. off-line [99]. off-loading [1590]. offering [351]. Offloading [787, 843, 1132, 1494, 1758, 1834, 2171]. offs [2256]. OGSA [71]. OGSA-DAI [71]. oil [1438]. oil-water-gas [1438]. OLARPBS [1243, 1455]. on-chip [338, 357, 899, 1166, 1224, 1568, 1798]. On-demand [48, 227, 550, 683, 1999, 2081]. On-the-fly [858, 1038]. One [163, 737, 738, 1194, 2250]. one-layer [163]. one-round [738]. one-to-all [1194]. One-to-one [2250]. Online [385, 529, 545, 709, 1057, 1228, 1245, 1403, 1546, 1636, 1637, 1756, 2171]. only [146]. onto [360, 1306, 1373, 1513, 1587]. ontological [348]. Ontology [549, 740, 1091, 1117, 1880, 2145, 2208]. Ontology-based [740, 1880]. ontology-driven [2145]. ontology-independent [1117]. open [212, 1108, 1229, 1384, 1429, 1508, 1528, 2125]. open-source [212, 1229, 1384, 1429, 1508]. OpenACC [601, 1362, 2155]. OpenCF [159]. OpenCL [294, 618, 822, 860, 1087, 1127, 1718, 1771, 2049, 2174, 2223, 2229]. OpenCL-based [2229]. OpenCL-to-WebCL [1087]. OpenCL-written [1771]. OpenFFT [1229]. OpenFlow [1704]. OpenFOAM [1030]. OpenMP [164, 193, 268, 600, 641, 642, 724, 918, 1024, 1517, 1713, 1752, 2155, 2157]. OpenMP/MPI [641]. OpenPOWER [2283]. OpenStack [2201]. operand [1110]. Operating [218, 221, 742, 795, 872, 1057, 1303]. Operation [53, 381, 649]. operational [1678, 1777, 2216]. operations [51, 130, 241, 272, 1240, 1297, 1785, 1969, 2065, 2351, 2372]. operator [1201]. operators [322, 1632, 2065]. opinion [1546]. opinion-mining [1546]. opportunistic [886, 1066, 1757, 1927, 2121]. opportunities [924, 1993]. Optical [61, 374–376, 378– 380, 404, 690, 883, 1243, 1455]. Optimal [81, 112, 113, 242, 248, 299, 309, 362, 395, 465, 629, 638, 656, 757, 816, 818, 890, 1099, 1165, 1208, 1321, 1324, 1415, 1441, 1475, 1666, 1671, 1680, 1717, 1925, 2023, 2029, 2087, 2103, 2271, 2278, 2321]. optimal-register [2321]. optimally [19]. Optimisation [405]. optimise [1402]. Optimization [81, 206, 207, 242, 266, 272, 277, 301, 334, 364, 376, 393, 401, 421, 422, 454, 456, 458, 463, 614, 622, 631, 685, 694, 717, 774, 775, 793, 816, 826, 840, 868, 987, 997, 1048, 1129, 1176, 1202, 1247, 1266, 1300, 1301, 1322, 1341, 1342, 1353, 1416, 1435, 1440, 1461, 1493, 1507, 1526, 1529, 1561, 1589, 1594, 1632, 1644, 1658, 1663, 1674, 1739, 1746, 1764, 1770, 1779, 1781, 1816, 1832, 1866, 1895, 1954, 1965, 1975, 2014, 2035, 2042, 2048, 2051, 2075, 2078, 2082, 2111, 2114, 2135, 2158, 2192, 2220, 2223, 2239, 2244, 2366]. optimization-based [401, 1954]. optimizations [643, 2056, 2176, 2302]. optimize [268, 1157, 1749]. Optimized [68, 220, 523, 707, 956, 973, 1144, 1277, 1307, 1573, 2072, 2159, 2160, 2227, 2241, 2335, 2369]. optimizer [1858]. Optimizing [160, 247, 425, 588, 646, 899, 950, 955, 1114, 1155, 1156, 1162, 1173, 1257, 1258, 1382, 1652, 1899]. optimum [1600]. option [417, 631, 1587]. options [417]. optoelectronic [1667]. orchestration [1781, 1976]. order [286, 393, 959, 1126, 1526, 1536, 2029]. ordered [2001, 2182]. ordering [337]. ore [2107]. organization [20, 2296]. organized [912, 1267, 2178]. organizing [128]. orientation [629, 1813, 1814]. Oriented [35, 57, 296, 303, 431, 435, 526, 561, 613, 712, 1062, 1190, 1259, 1269, 1387, 1507, 1637, 1778, 2283]. originals [1958]. Orthogonal [894, 1309]. OSched [1898]. OSIC [166]. other [839]. OTIS [70, 232, 353, 1667]. OTIS- [353]. OTIS-MOT [232]. OTIS-torus [70]. Otsu [2024]. Outlier [1988, 2279]. output [735, 1277, 1324, 1853]. output-queue [1853]. outsourcing [419, 468, 564]. over-the-counter [2091]. overcoming [1483]. overhead [210, 219, 328, 854, 1023, 1532, 1615]. overheads [332, 1162]. overlapping [708]. Overlay [40, 656, 769, 919]. overlays [415]. oversubscription [794, 2316]. ownership [916]. oxidation [2356].

P [1448]. P-SEP [1448]. P2P [31, 39, 40, 333, 514, 526, 704, 765, 769, 872, 895, 1094, 1254, 1653, 1676, 1763, 1971]. P2P-based [333]. package [185, 1229]. packages [719]. packet [61, 407, 823, 859, 1122, 1151, 2152, 2196, 2254, 2297]. packet-level [407]. packets [78]. packing [81, 1666, 1846]. page [591, 1303]. PageRank [938, 1794]. Paired [1998]. pairing [567, 1708, 1751]. pairing-based [1708, 1751]. pairs [1186]. pairwise [573, 730, 1759]. pancake [1169]. panconnectedness [1120, 1766]. Panconnectivity [111]. panpositionable [1766]. paperless [1295]. paradigm [586, 1528, 1881, 1927]. Paradigmatic [396]. paradigms [1399]. Parallel [3, 5, 70, 109, 132, 136, 150, 158, 168, 170, 180, 195, 207, 212, 229, 239, 258, 286, 294, 300, 310, 320, 363, 367, 368, 376, 415, 417, 437, 452, 456, 470, 480, 485, 525, 600, 616, 701, 726, 727, 763, 813, 826, 834, 857, 932, 935, 938, 941, 1042, 1072, 1159, 1167, 1229, 1232, 1233, 1317, 1349, 1354, 1357, 1364, 1400, 1403, 1407, 1419, 1435, 1436, 1451, 1517, 1667, 1680, 1714, 1729, 1752, 1772, 1775, 1832, 1907, 1912, 2025, 2029, 2038, 2047, 2048, 2055, 2058, 2065, 2106, 2148, 2154, 2174, 2179, 2181, 2219, 2270, 2284, 2352, 2356]. parallel [34, 53, 77, 133, 134, 165, 177, 182, 187, 232, 266, 267, 271, 273, 276, 285, 293, 296, 301, 322, 336, 371, 383, 428, 463, 484, 508, 527, 528, 582, 613, 661, 709, 720, 750, 778, 779, 788, 794, 811, 840, 847, 867, 1002, 1007, 1015, 1016, 1026, 1070, 1078, 1105, 1154, 1168, 1176, 1184, 1204, 1227, 1308, 1316, 1375, 1388, 1406, 1418, 1432, 1492, 1506, 1526, 1569, 1572, 1601, 1605, 1615, 1627, 1639, 1651, 1659, 1709, 1710, 1718, 1719, 1740, 1831, 1986, 2015, 2037, 2050, 2059, 2073, 2074, 2077, 2124, 2158, 2187, 2188, 2198, 2226, 2316, 2321, 2361]. parallel [72, 110, 115, 189, 193, 268, 343, 773, 974, 1071, 1102, 1128, 1243, 1256, 1322, 1455, 1495, 1498, 1536, 1620, 1621, 1634, 1641, 1847, 2172, 2320]. parallelisation [2009]. parallelism [117, 143, 161, 432, 459, 649, 898, 926, 945, 959, 1437, 1648, 1659, 1731, 1793, 1907, 1911, 2156, 2182]. parallelism-aware [143]. parallelisms [122]. Parallelization [105, 356, 393, 471, 642, 1076, 1279, 1429, 1440, 1444, 1570, 1582, 1713, 1716, 1759, 1990, 2039, 2064, 2118, 2360, 2365]. Parallelized [1031, 1311, 1765, 1830, 2266, 2366]. Parallelizing [254, 267, 865, 955]. parameter [144, 1040, 1622, 1633, 2168]. parameterized [173, 388]. parameters [462, 733]. ParaSQL [2007]. Pardis [1714]. parenthesizing [313]. paretic [1891]. Pareto [955]. parity [993, 1635]. parity-preserving [1635]. parking [1845]. parsing [2168]. part [660, 900]. partial [218, 332, 443, 1445]. Particle [146, 631, 774, 775, 816, 1070, 1507, 1632, 1658, 1674, 2114]. Particle-based [1070]. particulars [357]. Partition [246, 1256, 2020, 2223, 2280, 2334]. partition-and-stitch [2280]. partition-based [2020]. partitioned [201, 446, 447, 1368]. partitioning [169, 252, 363, 475, 628, 646, 714, 744, 853, 1098, 1165, 1432, 1618, 1767, 1871, 2074]. party [665, 914, 2139]. Pascal [2316]. pass [756, 1963]. pass-through [756]. passing [90, 240, 271, 297, 1015, 1066, 1520]. passivation [580]. Passive [499, 1580]. password [665, 914, 961]. password-based [914, 961]. past [400, 1066, 2019]. patching [2130]. path [244, 245, 425, 582, 691, 788, 800, 1101, 1120, 1186, 1355, 1359, 1509, 1539, 1928, 1998, 2250, 2265, 2267, 2271, 2278]. pathologies [1037]. Paths [73, 102, 204, 730, 757, 1475, 1478, 1591]. patient [2263]. patients [1013]. pattern [1045, 1370, 1882, 1912, 1988, 2048, 2241, 2279, 2357]. pattern-based [1988]. pattern-free [2048]. patterns [216, 229, 481, 820, 1244, 1348, 1380, 1517, 1785, 1947, 2187]. Payload [1671, 2127]. payment [1295, 1740]. payments [2214]. PC [268, 1013]. PCI [756]. PDEs [2226]. Peacock [986]. peak [1706, 2150]. pedagogy [225]. Peer [11, 41, 42, 305, 403, 517, 533, 548, 585, 599, 656, 1260, 1763, 2298]. peer-guaranteed [1763]. Peer-to-Peer [11, 41, 42, 305, 403, 517, 533, 548, 599, 1260, 2298]. peer-to-peer-based [585]. peerGroup [40]. penetration [1945]. pentadiagonal [239]. people [2086]. Per-packet [1151]. perceptron [1791]. perceptron-based [1791]. perceptual [891, 1148]. PerContRep [964]. perfect [730, 1315]. Perform [610, 880]. Performability [1511]. Performance [1, 32, 43, 44, 143, 209, 223, 272, 342, 364, 365, 375, 398, 410, 454, 469, 476, 479, 504, 562, 566, 573, 595, 617, 639, 695, 761, 836, 841, 842, 863, 866, 875, 910, 920, 921, 928, 931, 933, 939, 956, 969, 974, 978, 1004, 1035, 1087, 1088, 1090, 1186, 1197, 1205, 1240, 1260, 1285, 1287, 1362, 1363, 1376, 1388, 1391, 1393, 1397, 1404, 1411, 1413, 1429, 1432, 1442, 1485, 1516, 1523, 1526, 1532, 1556, 1596, 1659, 1675, 1684, 1695, 1765, 1774, 1834, 1835, 1852, 1854, 1977, 1993, 2019, 2042, 2056, 2074, 2129, 2180, 2236, 2270, 2305, 2316]. performance [4, 7, 37, 82, 89, 93, 96, 97, 104, 155, 171, 177, 182, 203, 210, 234, 248, 264, 270, 273, 275, 282, 299, 362, 441, 465, 481, 501, 518, 565, 568, 608, 643, 650, 687, 710, 818, 862, 867, 897, 925, 927, 949, 989, 1000, 1021, 1026, 1079, 1085, 1105, 1106, 1127, 1132, 1156, 1179, 1201, 1245, 1246, 1291, 1299, 1342, 1360, 1415, 1476, 1477, 1489, 1490, 1652, 1688, 1700, 1701, 1710, 1712, 1728, 1747, 1770, 1780, 1795, 1826, 1848, 1919, 1920, 1926, 1997, 2000, 2011, 2036, 2053, 2066, 2078, 2154, 2201, 2205, 2207, 2223, 2231, 2283, 2312, 2328, 2340, 2343, 2354, 2375]. performance [61, 115, 450, 541, 636, 721, 773, 801, 918, 1046, 1789, 1979, 2127, 2193, 2242]. Performance-aware [1205]. Performance-based [32, 209]. Performance-energy [974]. performance-power-energy [1079]. performances [827]. performing [871]. period [2023]. Periodic [242, 379, 385, 1111, 2285]. periodical [721]. perishable [2069]. permanent [1385]. permissioned [2333]. permutation [317, 1630]. permutations [309, 1192]. perpetual [2267]. person [1961]. personal [308, 426, 1811, 2090]. personalized [835, 1846, 2145]. perspective [357, 1145, 1251, 1951, 2046]. perspectives [521]. pervasive [269, 489, 490, 494, 495, 509, 551, 555, 556, 687, 964, 974]. pessimistic [321]. Peta [1022]. Peta-scale [1022]. petascale [1509, 1789]. Petersen [527, 1028, 1724]. Petersen-torus [527, 1028]. Petri [107, 259, 363, 685, 1161]. PETSc [261]. PFSP [1630]. PGAS [453]. PGSW [872]. PGSW-OS [872]. pharmacosurveillance [1305]. phase [337, 706, 716, 1438]. PHAT [2049]. PHENIC [1864]. phenomenon [1793]. Phi [1311, 1570, 1641, 1644, 1774, 1990, 1996, 2037, 2322]. phone [1294]. phones [610, 931]. Photo [1824]. photograph [1824]. photonic [375, 753, 1197, 1250, 1490, 1664, 1665, 1863, 1864]. photorefractive [381]. PHR [1061]. Phy [919]. phylogenetic [640]. Physical [620, 625, 1082, 1093, 1144, 1807, 1836, 2255]. Physically [891, 1800]. physics [910, 2040]. Pi [1729]. PIC [829]. Piccolo [663]. pipeline [786, 1588, 1722, 2062, 2107]. Pipelined [166, 1243, 1455]. pipelines [1823]. piracy [777]. Pirax [777]. Pit [58]. pivot [1422]. pivot-based [1422]. pixel [152, 678, 1844]. pixel-shifting [678]. pixels [662]. PKDS [736]. PKI [1974]. PL [1905]. PL-DVFS [1905]. placement [22, 88, 327, 524, 533, 758, 861, 879, 954, 975, 998, 1199, 1489, 1656, 1673, 1782, 1833, 1862, 1995, 2128, 2245, 2248, 2286, 2293, 2309, 2320]. plan [1703, 1970]. planar [546, 1940, 2020, 2046]. planner [1860]. planning [245, 257, 722, 1539, 2265]. plant [2195]. plants [1440, 1707, 2366]. plate [558]. Platform [16, 215, 253, 277, 294, 308, 419, 454, 586, 599, 618, 767, 912, 949, 986, 998, 1002, 1168, 1207, 1213, 1214, 1220, 1222, 1308, 1326, 1392, 1422, 1550, 1701, 1792, 1807, 1809, 1824, 1841, 2075, 2085, 2131, 2142]. platforms [180, 200, 222, 386, 587, 826, 840, 845, 921, 1014, 1020, 1136, 1176, 1190, 1233, 1250, 1358, 1416, 1427, 1716, 1840, 1907, 1911, 1967, 2040, 2049, 2074, 2179, 2344]. PLC [1298]. PLS [2272]. pluggable [1875]. plus [1713, 2044]. PMC [341]. point [224, 506, 1324, 1539, 1790, 2203, 2274, 2368]. points [83, 1386, 2012]. polar [2321]. policies [457, 613, 1032, 1064, 1323, 1855, 2069]. policy [264, 304, 740, 890, 1590, 1957, 2263]. policy-based [304]. Pollard [1740]. pollination [2239]. pollutant [603]. polyhedral [334]. polynomial [70, 1231, 2214]. polynomials [1411]. Pool [96]. Pool-based [96]. Pooling [1230]. Popularity [22, 1736]. Popularity-based [1736]. popularity-driven [22]. population [2292]. porous [129]. port [368]. portability [1718]. portable [97, 2324]. positioning [312, 1083]. possession [1638, 1983]. post [275, 1369]. post-fabrication [275]. Potential [749, 1596, 1759]. Power [69, 125, 176, 292, 299, 325, 338, 342, 355, 401, 404, 703, 717, 764, 793, 801, 859, 876, 925, 954, 994, 1073, 1079, 1110, 1118, 1123, 1143, 1145, 1416, 1477, 1493, 1614, 1648, 1660, 1707, 1758, 1761, 1787, 1789, 1821, 1870, 1905, 1975, 2011, 2078, 2113, 2129, 2180, 2195, 2220, 2288, 2296, 2317, 2369]. power- [69]. Power-aware [325, 1761, 1905]. power-capping [703]. power-constrained [2078]. Power-efficient [292, 859, 876, 1110, 1123, 1477]. power-performance [925]. power-saving [342, 1073]. PowerPC [446, 447]. PPC [446, 447]. PPMQSort [1256]. practical [380, 554, 964, 1259, 1462, 1596, 1812]. practice [89]. Pre [486, 776, 1350, 1424, 1558, 2043]. pre-analysis [1424, 2043]. Pre-execution [776]. pre-processing [1350, 1558]. pre-scheduling [486]. precise [1768]. precision [418, 506, 2274]. Preconditioned [320, 440, 1835, 2061]. Preconditioner [370]. preconditioners [158]. predict [1066]. predictable [2331]. Predicting [1017, 1555]. Prediction [131, 168, 205, 329, 385, 471, 731, 807, 936, 945, 984, 1000, 1133, 1147, 1162, 1181, 1188, 1259, 1312, 1327, 1328, 1398, 1426, 1499, 1627, 1834, 1859, 1900, 1914, 1926, 1928, 1950, 2021, 2039, 2040, 2151, 2180, 2289, 2313, 2314]. Predictive [990, 1237, 1598, 2231]. predictor [1287, 2345]. predictor-corrector [2345]. predictors [84, 1160]. preemptible [1925, 2101]. preemption [721]. Preemptive [1386, 1660, 1745]. Preface [15, 120, 279, 302, 469, 850, 1339]. preference [2093]. preference-based [2093]. prefetch [426, 1867]. prefetcher [1754]. Prefetching [176, 776, 803, 1123, 1276, 1776, 1827, 1894, 2005, 2316]. Prefetching-based [2005]. prefix [1042, 1717, 2065]. prefix-sum [1717]. preliminary [601, 2094]. preprocessing [1560]. present [400, 906]. PRESENT-128 [906]. PRESENT-80 [906]. preserve [771]. preserving [84, 664, 668, 1347, 1486, 1635, 1931, 2139, 2347, 2373]. pressure [1035, 2042]. prevent [966]. preventing [1290, 1540]. prevention [680, 1473, 1551, 1786]. pricing [417, 631, 1587, 1636, 2204]. Primal [493]. primary [1884]. prime [1451]. Principal [1363, 1547]. priorities [750]. prioritization [802]. Prioritizing [1628]. Priority [259, 655, 1037, 1161, 1787, 2001, 2308]. Priority-based [1787]. Privacy [67, 99, 304, 366, 444, 509, 510, 664, 668, 676, 1061, 1215, 1223, 1347, 1486, 1534, 1822, 1931, 1973, 2139, 2263, 2347, 2373]. Privacy-aware [444, 1061, 1973]. Privacy-enhanced [67, 1223]. Privacy-preserving [1931, 2347, 2373]. private [882, 1011, 1211, 1565, 1870, 1995]. privatization [642]. Privatizing [1726]. privileged [532]. Proactive [274, 358, 373, 416, 792, 987, 1162, 2034]. Probabilistic [48, 923, 1529, 2009, 2038, 2262, 2264]. probability [1499]. problem [30, 80, 166, 187, 227, 237, 313, 378, 461, 488, 534, 582, 630, 713, 774, 836, 838, 929, 935, 999, 1015, 1186, 1238, 1300, 1301, 1375, 1421, 1475, 1592, 1607, 1667, 1731, 1880, 2022, 2189, 2286]. problem-solving [534]. problems [77, 80, 156, 172, 376, 382, 463, 469, 700, 873, 928, 957, 1353, 1431, 1847, 1996, 2158]. procedure [1435]. procedures [1257, 1258]. Process [200, 214, 422, 428, 519, 534, 541, 611, 644, 719, 725, 990, 1161, 1288, 1304, 1410, 1652, 1714, 1785, 1815, 1836, 2057, 2286, 2320]. process-aware [990]. process-style [2057]. Processes [580, 880, 1483]. Processing [132, 232, 300, 455, 524, 525, 586, 627, 631, 712, 762, 767, 786, 864, 871, 998, 1107, 1163, 1232, 1320, 1322, 1350, 1368, 1480, 1519, 1524, 1558, 1571, 1584, 1594, 1626, 1671, 1744, 1745, 1752, 1790, 1867, 1910, 2032, 2053, 2062, 2097, 2207, 2217, 2246, 2258, 2274, 2276]. processor [121, 143, 280, 282, 286, 299, 379, 603, 612, 853, 857, 989, 1165, 1271, 1280, 1604, 1708, 1737, 1751, 2000, 2322]. processors [127, 136, 188, 247, 275, 277, 337, 342, 382, 390, 446, 447, 637, 732, 792, 796, 801, 803, 984, 1007, 1023, 1174, 1181, 1287, 1409, 1526, 1671, 1720, 1956, 2190, 2256, 2311, 2364]. ProCTA [2334]. producer [2335]. product [111, 162, 506, 1392, 1452, 1459]. production [683]. productive [1000]. productivity [110]. products [479]. Profile [1597, 2109, 2199]. Profile-based [1597]. profiles [327, 1707, 2068, 2371]. profiling [1026, 1244]. profit [1200, 1390, 1608, 2022]. program [29, 641, 1247, 1334, 2181, 2334, 2363]. Programming [137, 165, 170, 184, 193, 215, 268, 289, 296, 301, 315, 324, 437, 586, 606, 700, 829, 1096, 1438, 1447, 1499, 1709, 1710, 1714, 1866, 1908, 2002, 2003, 2073, 2087, 2120, 2172, 2175, 2192, 2301, 2353, 2359]. programming-based [1866, 2301]. programs [58, 301, 428, 443, 926, 940, 974, 1130, 1279, 1598]. progress [385, 732]. progressive [161]. project [1368, 1454, 1695]. projecting [722]. projection [1309]. projections [2046]. prolong [1448]. prone [441, 969]. Proof [910, 1670, 1738, 1982]. PROOF-based [910]. proofs [1371]. propagation [12, 136, 416, 1320, 1349, 2342]. properties [251, 353, 381, 648, 831, 839, 901]. property [333, 1323]. proportional [1007]. proportionality [1736]. proposal [1264, 2330]. protecting [1222]. protection [705, 734, 736, 1707]. protection-key [736]. Protein [168, 354, 1299, 1517, 1621, 1624, 2159, 2160]. Protocol [10, 69, 91, 288, 351, 388, 390, 405, 499, 515, 569, 584, 634, 664, 665, 667, 707, 791, 844, 886, 889, 903, 908, 914, 944, 963, 1138, 1202, 1289, 1347, 1465, 1471, 1472, 1486, 1580, 1669, 1670, 1687, 1690, 1727, 1738–1740, 1750, 1762, 1788, 1843, 1857, 1868, 1915, 1939, 1966, 2102, 2104, 2150, 2169, 2210, 2319, 2333, 2347]. protocols [243, 350, 716, 858, 1657, 1851, 1934]. PROV [1487]. provable [961, 1984]. provably [1817]. provide [2235]. provident [1946]. provider [507, 588, 2034]. providers [1609]. provision [636]. Provisioning [333, 395, 441, 548, 635, 987, 1043, 1053, 1097, 1200, 1251, 1453, 1541, 1586, 1602, 1946, 1949]. provisioning-based [1053]. proxies [36]. proximity [585]. proximity-aware [585]. proxy [567, 772, 812, 966]. pruning [2134]. PS [996, 1125]. PS-Cache [996]. PS3 [418]. pseudo [379]. pseudo-periodic [379]. pseudorandom [1171]. PSML [2124]. PSO [1595, 2161]. PSO-DS [1595]. PSR [807]. Pthreads [222]. Public [323, 908, 1252, 1469, 1608, 1655, 1701, 1890, 1931, 2141, 2262, 2264, 2373]. public-key [323, 2141]. publicly [1638, 1983]. publish [79]. publish/subscribe [79]. publishing [1886]. PUE [1884]. pull [2103]. pulmonary [1573]. punishment [1897]. pupil [1734]. purchase [1147]. pure [2271]. purposes [1840]. pursuit [894]. push [1552, 2103]. push-pull [2103]. pyramid [711]. Python [189, 1712].

Q [1001]. QCA [1936, 2098, 2233]. QoR [2258]. QoS [26, 57, 101, 213, 362, 425, 524, 556, 635, 636, 651, 661, 686, 1003, 1008, 1039, 1251, 1481, 1507, 1612, 1613, 1833, 1862, 1901, 2235, 2253, 2287]. QoS-aware [524, 1003, 1481, 1862, 1901]. QoS-based [57, 651]. QoS-enabled [2287]. QoS-guaranteed [1833]. QR [2062]. QRSF [1003]. quadratic [473]. quAlity [10, 138, 349, 350, 675, 731, 1148, 1351, 1369, 1553, 1578, 1685, 2016, 2022, 2081]. quAlity-based [350]. quality-of-service [731, 2016]. quantifying [123, 2072]. quantile [1245]. quantitative [1525, 2041]. quantum [583, 1277, 1497, 1741, 1865, 1869, 1933, 2221, 2229, 2251, 2255, 2308]. quantum-dot [1741, 1865, 1869, 2251, 2255]. quasi [737, 1823]. quasi-delay-insensitive [1823]. quasi-Newton [737]. quaternary [1497]. queries [322, 929, 1368, 1525, 1626, 2368]. Query [300, 455, 1011, 1163, 1366, 1519, 2106, 2244, 2259, 2271]. querying [688]. queue [250, 1853, 1893]. queue-based [250]. QueueCore [143]. queueless [309]. queues [1080, 1293]. Queuing [235, 849, 1179, 1652, 1671, 1962]. QUICK [1833]. QuickSort [639, 1256]. Quorum [729]. Quorum-based [729].

R&D [573, 1338]. race [17, 1598, 2363]. races [1128, 1284, 1768]. radar [1232, 2116]. radial [1492]. radio [654, 883, 1107, 1143, 1270]. radio-over-fiber [883]. radiosity [160, 180]. radix [103, 896, 1114]. RAID [1920]. RAIDs [1827]. Raising [240, 2355]. RALBA [2304]. RAM [1652, 1720, 1798, 2255, 2285, 2306, 2315]. RAM-based [1798, 2306, 2315]. RAMP [1200]. Random [45, 154, 319, 626, 765, 915, 1316, 1480, 1799, 1932, 2174]. Randomized [30, 855, 2122]. range [493, 557, 671, 783, 1366, 1956, 2368]. ranging [502]. rank [737, 1581]. rank-one [737]. ranking [515, 647, 1002, 1300, 1301, 1369, 2340, 2341]. ranking-based [2340]. RankSVM [1184, 1512]. RANS [2118]. ransomware [1551]. RASC [918]. Raspberry [1729]. rate [93, 256, 554, 2090]. rating [2282]. ratings [1095]. raven [1764]. Raw [1361, 1388, 1820]. RBAC [679]. RCB [1838]. rCUDA [1908]. RDF [1296]. RDMA [2305]. RDT PSO [1658]. re [298, 772, 912, 1225, 1838, 1961]. re-authentication [298]. re-encryption [772]. re-identification [1961]. re-keying [1838]. re-organized [912]. RE-UPS [1225]. reaching [1839]. REACT [2247]. reaction [813, 1444]. reaction-diffusion [813, 1444]. reactive [245, 350, 716, 1938]. read [409, 484, 1776, 1827]. readability [2138]. reader [499, 1788]. reader-to-reader [1788]. Real [44, 54, 190, 259, 342, 394, 411, 496, 558, 578, 782, 817, 853, 862, 889, 989, 991, 995, 1008, 1012, 1029, 1083, 1155, 1203, 1220, 1386, 1403, 1423, 1454, 1463, 1583, 1637, 1660, 1685, 1728, 1747, 1752, 1760, 1782, 1786, 1875, 1905, 1938, 1963, 2003, 2028, 2070, 2085, 2147, 2235]. Real-time [44, 54, 190, 342, 394, 411, 496, 558, 578, 782, 817, 862, 889, 989, 991, 995, 1008, 1012, 1029, 1155, 1203, 1220, 1386, 1403, 1423, 1454, 1463, 1583, 1637, 1660, 1685, 1747, 1752, 1760, 1786, 1875, 1905, 1938, 2003, 2028, 2070, 2085, 2147]. realistic [92, 1313]. reality [610, 1213]. realize [890]. realized [417]. Realizing [1192]. rearrangement [1047]. reasoning [523, 740]. rebalance [1479]. receiver [930]. recharging [2267]. reclaim [1303]. recognition [64, 151, 545, 546, 553, 557, 558, 659, 1086, 1349, 1352, 1488, 1564, 1567, 1940, 1959, 1970, 2090, 2108, 2203, 2241]. Recommendation [66, 539, 835, 884, 1081, 1392, 2093, 2145, 2234]. Recommender [694, 1672, 2091]. Reconfigurable [192, 219, 326, 332, 998, 1042, 1243, 1455, 1826]. Reconfiguration [200, 218, 332, 443, 710, 857, 870, 1110, 1175, 1378, 2307]. Reconstructing [1630]. reconstruction [132, 820, 1076, 1103, 2147]. record [358, 373, 909, 1170, 2009, 2038]. recording [743]. recordings [1294]. records [1615, 2263]. Recovering [1074]. recovery [3, 67, 199, 562, 666, 902, 970, 1093, 1818, 1975, 2109, 2150]. rectangular [1591]. recurrence [2041]. recursive [1317]. RED [349]. redesign [1446]. redistribution [1187]. Redressing [1823]. reduce [211, 409, 1380, 1525, 1785, 1912, 1935]. reduced [197, 404, 805, 2278]. reducers [1799]. Reducing [335, 784, 868, 1034, 1399, 1509, 1660, 1863, 1995, 2101, 2177, 2233, 2238]. Reduction [219, 388, 572, 755, 949, 954, 959, 1006, 1098, 1304, 1309, 1446, 1520, 1828, 1859, 2128, 2137, 2331, 2349]. reduction-based [1859]. reductions [105]. redundancy [424, 1847]. redundant [1434]. Reengineering [1907]. Refactoring [2179]. Reference [307, 1148, 1244, 1953]. references [114, 267]. refinement [422]. Region [105, 1090, 1988, 2161, 2285]. Region-based [105, 1090, 2161]. regional [1831]. Regions [147, 1609]. register [325, 337, 893, 2321]. registration [1249, 1283, 2067, 2228]. regression [700, 1188, 1245, 1354]. regular [851, 2152]. regulated [2092]. regulations [2237]. regulatory [134]. Reinforcement [214, 245, 372, 1143]. reinforcement-learning [372]. rejuvenation [1870]. related [80, 2266]. relation [1768]. relational [354]. relations [539]. relationship [2213]. relationships [965, 1345, 1692]. Relative [346, 1288]. relaxed [140, 938]. relay [224, 1965, 2213]. relaying [1561]. Reliability [6, 225, 277, 360, 448, 696, 715, 717, 818, 848, 1064, 1178, 1200, 1271, 1319, 1453, 1648, 1749, 2101, 2216, 2367]. Reliability-aware [277, 360, 717, 818, 1200]. Reliable [14, 87, 554, 620, 909, 1490, 1529, 1532, 1815, 1928, 1939, 2099, 2100]. relocation [1204]. relying [1845]. remapping [723]. Rematch [14]. Remote [65, 94, 262, 707, 743, 1079, 1381, 1638, 1697, 1705, 1908, 1923, 1983]. remotely [195, 934, 1428, 1697]. removal [1963]. removing [1560]. rendering [724]. Rendezvous [1471, 1472, 1618]. rental [903]. reorder [566]. Reordering [280, 670]. repackaging [1336]. repairing [1597]. repetitive [1451, 1667]. Replacement [36, 505, 782, 1657, 1770]. replacements [1844]. replay [743]. Replica [22, 758, 879, 1737, 2128, 2328]. Replica-aware [2128]. replicas [100, 115]. replicated [199]. Replication [59, 249, 281, 358, 373, 389, 487, 1064, 1172, 1183, 1242, 1374, 1896]. repository [2095]. represent [946]. Representation [151, 356, 1535, 1572, 1825, 2108]. reproduce [1384]. reputation [427, 514, 526, 895, 964, 1095]. reputation-based [514]. reputation-oriented [526]. request [535]. required [1453]. Requirements [512, 515, 722, 1161]. requirements-aware [515]. Research [153, 521, 563, 924, 947, 962, 1108, 1275, 1278, 1333, 1528, 1534, 1575, 1818, 1957, 2019, 2270, 2310]. reservation [244, 287, 402, 718, 779, 2335]. reservations [1228]. resident [867]. residual [1354, 2013]. residual-feedback [1354]. resilience [1130, 1281, 1387, 1631, 1789, 2141]. resilience-oriented [1387]. resiliency [290]. Resilient [1304, 1401, 1417, 1462, 1490, 1838, 1864]. resistant [735]. resizing [1155, 1521, 1548]. resolution [61, 763, 1553, 1685]. resolutions [1548]. Resolving [1480]. Resource [33, 45, 76, 213, 282, 287, 308, 315, 384, 395, 434, 507, 519, 522, 534, 588, 624, 633, 635, 640, 654, 692, 731, 765, 768, 779, 783, 801, 805, 814, 841, 845, 855, 872, 875, 897, 912, 987, 997, 1003, 1025, 1052, 1053, 1069, 1119, 1142, 1241, 1251, 1270, 1337, 1390, 1476, 1559, 1586, 1636, 1649, 1673, 1745, 1747, 1760, 1781, 1926, 1946, 1949, 1980, 2114, 2129, 2202, 2262, 2289, 2304, 2315, 2335]. resource-allocation [1119]. Resource-aware [805, 1673, 2304]. Resource-efficient [640]. resources [31, 83, 278, 441, 686, 718, 809, 842, 866, 1179, 1253, 1323, 1449, 1541, 1658, 1832, 2201]. response [1017, 1963, 1982]. REST [1850]. REST-based [1850]. restart [617, 1417, 2192, 2295]. Restricted [435, 1169, 1585]. Restructuring [508]. results [104, 2064, 2245]. Retraction [2374, 2375]. retrieval [593, 682, 887, 934, 1573]. retweeting [1146]. reuse [247, 2095]. reversals [1169]. Reversibility [662, 916, 1820]. Reversible [429, 668, 1116, 1257, 1258, 1277, 1548, 1635, 1842, 1936, 2098, 2233, 2308, 2339]. Review [450, 804, 1108, 1198, 1473, 1743, 1888, 2019, 2110, 2281, 2310]. Revisiting [346, 566]. reward [1678, 1777, 2216]. RFID [499, 510, 905, 960, 1465, 1669, 1687, 1788, 1843, 2210]. RFSA [197]. rheumatoid [2247]. rho [1740]. rhythm [1331]. rhythm-based [1331]. Riccati [937]. RIP [746]. RIP-based [746]. risk [1581, 1757, 2199, 2282]. risk-aware [1757]. RNA [329]. RNTS [99, 592]. road [558, 1555]. robot [557, 870]. robotic [2085]. Robust [44, 46, 150, 430, 677, 962, 967, 987, 1530, 1804, 1906, 1936]. robustness [1829, 2341]. ROIN [526]. Role [71, 623, 2263]. Role-based [71, 623, 2263].

Rolex [1387]. roofline [943]. roosting [1764]. root [166, 2266]. roots [70]. rough [1622]. round [738]. route [1262, 2278]. routed [48, 368, 1113]. router [1113, 1778]. routers [432, 2245]. routes [1563, 2331]. routine [946]. routines [933]. routing [26, 103, 200, 204, 230, 231, 284, 309, 338, 345, 350, 362, 401, 415, 439, 465, 497, 619, 636, 667, 698, 716, 730, 757, 771, 856, 886, 923, 1038, 1044, 1066, 1135, 1189, 1239, 1250, 1290, 1377, 1385, 1448, 1471, 1472, 1623, 1664, 1665, 1702, 1757, 1762, 1819, 1903, 1932, 1939, 1966, 2020, 2110, 2242, 2248, 2307, 2336]. routings [677]. RPL [2163]. RS [93, 1230]. RS-Pooling [1230]. RSA [323, 736, 1655]. RSEDP [88]. RTD [2214]. RTNSS [1290]. RTS [1161]. Rule [420, 1227, 1305, 1560, 1761, 2234]. Rule-based [420, 1560, 1761, 2234]. rules [255, 461, 538, 695, 1048, 2034]. rules-based [2034]. run [219, 421, 428, 868, 972, 1011, 1522, 1569, 1615]. run-time [219, 421, 428, 868, 972, 1522, 1569, 1615]. runner [1259]. running [1776]. Runtime [218, 332, 453, 693, 710, 714, 778, 788, 899, 1627, 2125, 2175]. RX [186].

S [138]. S-boxes [138]. S2DIO [2307]. SaaS [588, 651, 719, 1329, 1695]. SaaS-based [1329]. Safety [889, 2213]. safety-critical [889]. SAFT [1864]. SAFT-PHENIC [1864]. SAIR [2258]. salesman [1375, 1667]. saliency [147]. salvation [204]. same [2109]. sample [2191]. samples [229]. sampling [379, 688, 1694, 1799]. Sandy [920]. SAPDS [366]. SAPFOR [2360]. SAT [378, 836]. satisfaction [213, 661, 1507]. saturation [1527]. SAVE [2323]. saving [176, 342, 355, 796, 930, 985, 1073, 1420, 1962, 2063]. saving-aware [176]. savings [778]. Scalability [6, 181, 480, 773, 921, 1030, 1181, 1238, 2040]. Scalable [1, 5, 77, 117, 223, 266, 267, 271, 291, 297, 343, 404, 478, 482, 528, 532, 556, 642, 656, 747, 753, 828, 837, 1018, 1049, 1074, 1125, 1293, 1414, 1439, 1447, 1457, 1492, 1501, 1530, 1647, 1652, 1670, 1792, 1796, 1799, 1848, 2041, 2059, 2169, 2246, 2307, 2328]. scale [55, 56, 88, 110, 112, 156, 195, 206, 235, 246, 289, 290, 298, 408, 415, 442, 481, 633, 653, 720, 767, 830, 962, 1018, 1022, 1104, 1134, 1141, 1145, 1176, 1208, 1230, 1242, 1248, 1255, 1286, 1326, 1363, 1375, 1382, 1387, 1460, 1593, 1599, 1720, 1742, 1755, 1831, 1867, 1943, 1961, 2089, 2095, 2115, 2131, 2138, 2169, 2192, 2215, 2260, 2327, 2344, 2346]. scale-invariance [1208]. scale-out [1720]. Scaling [778, 984, 1012, 1174, 1280, 1396, 1620, 1654, 1747, 1829, 1924, 2031]. scaling-out [1829]. Scaling-up [1829]. scan [2065]. Scanless [691]. scanning [876, 1076]. scattering [172, 471, 936, 1398, 1793, 1996]. scenario [764]. scenarios [1398]. scene [1070]. scenes [2067]. schedulability [256, 1386]. schedule [262, 2243]. scheduler [126, 1134, 1761, 1898, 2287, 2340]. scheduler-per-node [2287]. Scheduling [13, 19, 34, 35, 38, 44, 46, 81, 125, 159, 177, 194, 196, 201, 206, 227, 234, 252, 257, 276, 292, 316, 335, 339, 361, 389, 432, 438, 486, 487, 519, 520, 583, 589, 635, 656, 706, 728, 747, 775, 794, 805, 811, 816, 817, 854, 995, 1002, 1003, 1016, 1032, 1037, 1039, 1047, 1050, 1053, 1056, 1069, 1100, 1111, 1131, 1161, 1187, 1219, 1253, 1351, 1357, 1386, 1449, 1502, 1507, 1518, 1522, 1533, 1577, 1595, 1633, 1637, 1649, 1662, 1686, 1749, 1756, 1784, 1808, 1866, 1893, 1905, 1938, 1949, 2016, 2078, 2128, 2159, 2160, 2170, 2173, 2206, 2292, 2301, 2344, 2370]. scheduling [14, 33, 115, 193, 209, 247, 268, 337, 360, 365, 394, 431, 613, 661, 709, 713, 721, 776, 782, 862, 929, 975, 990, 1036, 1054, 1073, 1150, 1168, 1251, 1259, 1307, 1323, 1641, 1764, 1783, 1979, 2003, 2004, 2063, 2362]. schema [1117]. scheme [38, 48, 49, 148, 173, 194, 213, 220, 221, 231, 238, 286, 314, 352, 358, 373, 427, 430, 517, 532, 547, 589, 624, 654, 656, 671, 674, 705, 721, 737, 772, 810, 854, 855, 905, 911, 928, 929, 958, 960, 961, 990, 1012, 1031, 1036, 1081, 1094, 1172, 1195, 1204, 1218, 1222, 1268, 1336, 1371, 1437, 1469, 1525, 1532, 1577, 1609, 1638, 1655, 1674, 1676, 1683, 1699, 1730, 1776, 1804, 1817, 1842, 1897, 1931, 1944, 1964, 1979, 1981, 1990, 2126, 2139, 2178, 2214, 2332, 2345]. schemes [106, 193, 268, 567, 672, 752, 966, 1138, 1177, 1260, 1263, 1427, 1458, 1476, 1956, 2140, 2355]. Schmitt [1861]. scholar [561]. scholar-oriented [561]. Schur [828]. SCIDDICA [579]. Science [155, 171, 244, 469, 595, 924, 927, 1267, 1818, 1957]. Scientific [208, 283, 316, 564, 1198, 1338, 1515, 1595, 1767, 1866, 2301]. SCM [1788]. scoping [1024]. score [1403]. SCOUT [1881]. Scrum [2237]. SCTP [349]. SD [1960]. SDDO [1335]. SDDO-based [1335]. SDN [985, 1270, 1542, 1578, 1602, 1913, 1997]. SDN-based [1270, 1542, 1578]. SDN-enhanced [1913]. SDRAM [1160]. SeaCloudDM [688]. sealed [1636, 1822]. sealed-bid [1822]. seam [1155]. seamless [338, 426, 1619, 1671]. search [5, 41, 150, 412, 597, 628, 783, 885, 929, 1016, 1046, 1322, 1353, 1399, 1422, 1525, 1621, 1630, 1675, 1689, 1742, 1794, 1930, 1956, 2084, 2111, 2236, 2370]. searchable [1366]. searching [11, 444, 765, 868, 1296, 1653]. SEATS [995]. SECM [2121]. secondary [329, 2159, 2160]. SeCRB [768]. secrecy [738]. secret [1263, 1268, 1552, 1580, 1609, 1934, 2332]. section [537, 565, 574, 660, 789, 900, 1060, 1140, 1207, 1456, 1709, 1837]. Secure [92, 98, 347, 377, 498, 500, 511, 542, 564, 567, 593, 674, 677, 736, 739, 824, 844, 901, 904, 909– 911, 958, 963, 988, 1072, 1171, 1209, 1263, 1295, 1331, 1371, 1470, 1609, 1655, 1690, 1805, 1806, 1812, 1817, 1843, 1868, 1944, 1955, 1974, 2036, 2140, 2149, 2210, 2217, 2222]. securing [1874, 1887]. Security [58, 89, 95, 264, 307, 387, 445, 509, 510, 515, 660, 665, 687, 740, 742, 858, 900, 913, 961, 1138, 1214, 1221, 1249, 1289, 1290, 1335, 1343, 1469, 1473, 1528, 1565, 1638, 1669, 1805, 1806, 1855, 1873, 1876, 1879, 1888, 1916, 1977, 1985, 2000, 2102, 2104, 2153, 2195, 2268, 2269, 2281, 2337]. segment [1793, 2280]. segmentation [762, 1137, 1503, 2024, 2144, 2161, 2349]. segregational [2119]. selectable [457]. selected [1659, 2316]. selecting [1101]. selection [100, 213, 224, 280, 298, 684, 719, 774, 788, 809, 825, 1314, 1494, 1547, 1590, 1600, 1622, 1628, 1632, 1873, 1926, 2094, 2186, 2213, 2236, 2247, 2285, 2328, 2368]. Selective [784]. selectors [378]. Self [49, 103, 128, 193, 268, 293, 298, 365, 366, 734, 792, 809, 870, 908, 1267, 1641, 1651, 1757, 1822, 2072, 2092, 2178, 2323]. Self-adaptive [293, 298, 809, 1757, 2323]. self-adjusting [792]. self-certified [908]. self-enforcing [1822]. self-healing [366, 734]. self-optimized [2072]. self-organized [1267, 2178]. self-organizing [128]. self-protection [734]. self-reconfiguration [870]. self-regulated [2092]. self-routing [103]. self-scheduling [193, 268, 365, 1641]. self-similar [49]. Semantic [539, 550, 553, 561, 768, 872, 1337, 1839, 2138]. Semantic-enabled [768]. semantics [1369]. Semi [694, 1816, 1963, 2277, 2321, 2360]. semi-folded [2321]. semi-global [1816]. Semi-real-time [1963]. Semi-sparse [694]. semi-supervised [2277]. SENFIS [6]. senior [2084]. sensed [195, 934, 1428]. sensing [493, 1079, 1813, 1814, 1884, 1889, 1891, 2148, 2204]. sensitive [1886, 2091]. sensitivity [891, 2303]. Sensor [6, 10, 25, 68, 69, 79, 119, 250, 288, 391, 511, 522, 538, 554, 559, 569, 623, 625, 653, 688, 699, 749, 752, 771, 830, 851, 870, 907, 968, 1036, 1065, 1090, 1139, 1195, 1202, 1324, 1335, 1355, 1448, 1471, 1472, 1496, 1504, 1529, 1563, 1647, 1657, 1679, 1683, 1694, 1702, 1730, 1739, 1755, 1837, 1847, 1852, 1854, 1881, 1901, 1932, 1954, 1964, 1966, 2104, 2107, 2109, 2110, 2126, 2147, 2151, 2169, 2207, 2213, 2228, 2230, 2267, 2275, 2312, 2331, 2336]. sensor-cloud [1964]. Sensor-Grid [79]. sensor/robot [870]. sensors [1813, 1814]. sentiment [2105]. sentiments [1459, 2119].

SEP [1448]. separation [989]. Sequence [457, 476, 602, 1004, 1484, 1624, 2051]. sequences [670, 955, 1517]. sequencing [535]. sequential [212, 229, 341, 1146, 1227, 1388, 1582, 2057, 2284, 2339]. serialization [784]. series [877, 1017, 1346, 1348, 1350, 2060, 2322]. Server [98, 199, 299, 430, 482, 484, 533, 674, 844, 880, 908, 954, 1165, 1177, 1209, 1283, 1590, 1652, 1678, 1727, 1777, 1857, 1868, 1923, 1962, 1992]. server-aided [98]. server-centric [1992]. servers [3, 60, 598, 793, 1063, 1134, 1237, 1511, 1926, 1944, 2036]. Service [35, 57, 93, 107, 200, 205, 306, 351, 422, 425, 524, 534, 535, 544, 548, 549, 562, 683, 725, 731, 739, 787, 845, 1043, 1062, 1179, 1190, 1212, 1240, 1254, 1300, 1301, 1314, 1351, 1371, 1481, 1507, 1552, 1622, 1628, 1796, 1841, 1845, 1901, 1943, 1967, 1974, 2016, 2034, 2082, 2106, 2120, 2201, 2214, 2222, 2253, 2328]. service-aware [1943]. service-oriented [57, 1507]. Serviceable [340]. ServiceFlow [57]. services [26, 59, 76, 159, 419, 500, 518, 520, 536, 815, 824, 842, 884, 887, 958, 964, 1011, 1017, 1086, 1213, 1272, 1331, 1491, 1544, 1565, 1579, 1642, 1684, 1797, 1811, 1812, 1836, 1860, 1902, 1985, 2149]. session [1138]. Session Weight [1351]. sessions [465]. Set [47, 197, 275, 715, 765, 786, 999, 1104, 1622, 1706, 1953, 2001, 2050, 2113, 2270]. set-based [1622]. sets [699, 1736]. setup [1122]. several [130, 1398]. SFLA [1315]. SGS [1526]. shadow [1333, 1428, 2161]. Shafer [1730]. shallow [163, 603, 1536]. shallow-water [603, 1536]. shallower [1293]. shape [546, 734, 1474, 1940]. Sharable [11]. shareable [1689]. Shared [54, 170, 173, 411, 470, 600, 1006, 1073, 1152, 1172, 1276, 1576, 1728, 1736, 1798, 1853, 1931, 1978, 1999, 2103, 2303, 2373]. shared-memory [173, 1853]. shared-nothing [1736]. sharing [97, 214, 243, 300, 308, 366, 403, 739, 842, 958, 1061, 1094, 1179, 1210, 1238, 1260, 1263, 1268, 1273, 1609, 1811, 2143, 2332]. sharpness [1573]. ShenZhen [1372]. Shibboleth [71]. Shift [1770, 2273, 2374]. shifters [1257, 1258]. shifting [678]. shifts [396]. ship [1556]. shop [713]. Short [92, 1680, 1950, 2314]. Short-Term [1950, 2314]. shortest [1186, 1539, 2267]. shortly [863]. shot [478]. shuffle [2367]. SI [413]. Sibling [1252]. side [2006]. sided [976]. Sierpinski [869]. sieving [1088]. sigma [572]. SigMR [1163]. signal [989, 1232, 1409, 1975, 2032, 2062, 2274]. signals [148, 578, 1963, 2109]. signature [567, 911, 1163, 1527, 2214, 2249]. signature-based [2249]. signatures [92]. signcryption [966]. signed [1325]. significance [2258]. significance-aware [2258]. SIIS [1543]. silent [1617]. silicon [1157, 1174, 1197, 2317]. silicon-photonic [1197]. silver [69]. SIMD [528, 808, 841, 1458]. SIMD-parallel [528]. similar [49, 1573]. similarity [544, 594, 871, 929, 978, 1261, 1422, 1742, 1906, 1958, 1986, 2008, 2358]. simple [645, 991, 1239, 1592, 1615, 1799]. simplification [1399]. Simplified [303]. SimRank [1906]. Simulated [1294, 1506, 1666, 1762, 1775]. Simulating [4, 579, 1055, 1482]. Simulation [117, 122, 129, 163, 208, 253, 294, 354, 452, 575, 580, 600, 603, 925, 1070, 1324, 1382, 1397, 1443, 1444, 1668, 1771, 1793, 1800, 1848, 2029, 2066, 2117, 2124, 2188, 2219, 2229, 2320, 2351, 2372]. simulations [169, 484, 821, 1030, 1442, 1514, 1526, 2073, 2118, 2246]. Simulator [407, 819, 1826, 2055, 2205, 2375]. simulators [1384]. Simultaneous [639, 1408, 2184, 2185, 2210, 2248]. Sina [1146]. Single [339, 577, 764, 808, 1006, 1115, 1793, 1861]. single-core [764]. single-ended [1861]. single-GPU [1006]. single-mode [1793]. Single-tape [577]. single/multi [808, 1115]. single/multi-core [808, 1115]. sink [1471, 1472, 1881, 1954, 2331]. sink-targeted [1881]. sinks [1563]. sites [203, 1433]. situation [1564]. size [1234, 1415]. sized [1591]. sizes [570, 1935]. SkelCL [822]. Skeletal [606]. skeleton [1394]. skewed [1799, 1991]. skewness [2294]. Skylake [2364]. skyline [1368, 1645, 1646]. SLA [420, 507, 890, 1073, 1282, 1533, 1577, 2304]. SLA-aware [1577]. SLA-awareness [890]. SLA-based [507, 1533]. SLA-RALBA [2304]. slave [607]. sleep [1036]. Sleepy [2177]. Sleepy-LRU [2177]. Slice [1078]. Slice-based [1078]. slicing [854]. sliding [2187, 2363]. sliding-window [2187]. Slot [825]. slots [1040]. slotted [888]. SMACOF [2031]. small [222, 481, 746, 809, 1231, 1341, 1480, 2153, 2165, 2369]. small-cell [2369]. small-footprint [222]. small-scale [481]. small-world [809]. SmallClient [1892]. Smart [63, 430, 498, 626, 650, 659, 664, 674, 887, 971, 992, 995, 1084, 1091, 1131, 1210, 1221, 1294, 1298, 1453, 1456, 1458, 1463, 1464, 1466, 1467, 1837, 1845–1847, 1849, 1850, 1852, 1854, 1882, 1889, 2173, 2330]. smart-sensing [1889]. SmartData [2164]. SmartMic [977]. smartphone [977]. smartphone-based [977]. smartphones [2030]. SmartRank [1131]. smoking [294]. smoother [611]. SMP [162, 726]. SMP-NUMA [162]. snapshot [969]. SNSP [889]. SnW [856]. Social [561, 647, 650– 652, 657, 658, 708, 976, 1081, 1085, 1086, 1092, 1095, 1147, 1223, 1325, 1369, 1543, 1565, 1610, 1809, 1811, 1815, 2093, 2119, 2131]. social-balanced [1325]. social-networks [647]. socially [1089]. Soft [1281, 1532, 1590, 1631, 2047]. soft-hard [1532]. Software [56, 137, 223, 295, 432, 504, 523, 534, 614, 628, 719, 790, 826, 1098, 1107, 1374, 1515, 1556, 1728, 1859, 2002, 2072, 2095, 2179, 2180, 2184, 2185, 2190, 2216, 2222, 2237, 2287]. software-defined [1107, 1556, 2184, 2185, 2216, 2222, 2287]. solar [1225]. solid [568, 763, 928, 951, 1478]. solid-fluid [951]. solution [172, 223, 286, 378, 413, 596, 630, 636, 813, 1020, 1238, 1628, 1876, 1996, 2022, 2087, 2143]. solutions [323, 445, 873, 1241, 1353, 1380, 1417, 2211]. solve [156, 476, 1049]. solver [834, 1425, 1439, 1501, 1639, 2118, 2226]. solvers [440, 933, 1774, 1835, 2065]. Solving [140, 187, 237, 260, 261, 320, 371, 463, 534, 700, 720, 935, 937, 1006, 1164, 1266, 1317, 1427, 1431, 1498, 1667, 2226, 2253]. Some [831]. Sort [53, 142, 775]. sorting [1169, 1394, 1801, 1941]. Soundprism [2070]. source [149, 212, 312, 636, 989, 1192, 1229, 1384, 1398, 1429, 1445, 1462, 1508, 1965]. source-based [636]. source-channel [1462]. SP [1154, 2364]. SP-ChainMail [1154]. Space [388, 534, 751, 859, 1158, 1399, 1790, 2168]. space- [859]. space-time [1790]. spaces [971, 992, 1261, 1315, 1407, 1539, 1812]. spam [1610]. Spanish [408]. Spanning [113, 237, 488, 581, 616, 1031, 1389]. Spark [1584, 1601, 1629, 2115]. Spark-based [1584]. SPARQL [1117, 1163]. Sparse [130, 151, 162, 260, 261, 454, 475, 479, 694, 726, 834, 922, 1049, 1154, 1227, 1231, 1317, 1412, 1425, 1501, 1570–1572, 1589, 1735]. sparsity [2109]. Spatial [152, 379, 381, 552, 1548, 1572]. spatio [1807, 2007]. spatio-temporal [1807, 2007]. spatiotemporal [195]. Speaker [1291]. SPEC [212]. Special [37, 89, 302, 374, 456, 537, 565, 574, 660, 789, 900, 1060, 1140, 1207, 1339, 1456, 1709, 1837]. specific [226, 311, 1510, 1723, 1852]. Specification [259, 448]. specifications [485]. specifying [972]. spectral [1350]. spectrum [29, 1557, 1884, 2148]. speculative [744, 1601, 2191]. speech [1291, 2232]. speech-based [2232]. speed [282, 352, 359, 381, 1012, 1090, 1170, 1289, 1361, 1734, 1932, 2198, 2274, 2332]. speeding [766, 1115, 1334]. speedTrust [1763]. speedup [110, 1981]. SpExSim [2183]. SPHeRe [759]. spider [2144]. spike [2242]. spiking [2242]. spilling [893]. spilling-friendly [893]. spinal [2208]. SPIT [514]. splitting [404, 1408]. SpMT [1576]. SpMV [1589]. sponsored [885]. spoofing [1290]. sporadic [201]. sports [2273, 2276, 2374]. spot [1608]. Spray [856]. Spray-and-Wait [856]. spread [945, 2040]. spring [819, 1346, 1437]. square [166, 1159]. square-root [166]. Squares [1431]. SRAM [1798, 1861]. SRAM- [1798]. SRAMs [859]. SRP [2049]. SS [579, 2103]. SS-CCDN [2103]. SSD [176, 978, 1674]. SSD-based [176]. SSE [928]. sShield [746]. SSL [515]. Stab [1589]. Stability [1217, 1259, 1323, 1926, 2135]. stability-based [1217]. stability-oriented [1259]. stabilization [962]. stabilizing [1475]. stable [1448]. stack [1663]. stacked [1720]. Stackelberg [1787]. stage [102, 1015, 1035, 1522]. staged [1206]. stand [1383]. stand-alone [1383]. Standardisation [497]. star [251, 414]. State [78, 84, 214, 388, 491, 563, 568, 763, 988, 1111, 1118, 1133, 1374, 1825, 2086, 2136, 2289]. state-machine [1374]. State-of-the-art [563]. state-transition [1825]. static [56, 929, 2001, 2186]. static-dynamic [2186]. static-priority [2001]. stationary [1350]. Statistical [123, 284, 1127, 1142, 1350, 2026, 2094]. statistics [492]. status [217, 2121]. steady [1111]. stealing [1711, 1980]. steering [2135]. steering-by-wire [2135]. steganography [1842]. Stencil [386, 393, 1106, 1233, 1719, 1722, 1912, 2343, 2354, 2355]. stencil-based [2343]. stencils [1899]. step [947, 1133]. steps [1363]. stepwise [1206]. stereo [295, 972, 1956]. stereoscopic [152, 1148]. stiff [2361]. stigmergy [458]. stimulated [1793]. stimulation [1897]. stitch [2280]. Stochastic [107, 226, 237, 413, 1144, 1203, 1350, 1443, 1678, 1686, 1716, 1777, 1903, 2216]. stochasticity [1444]. stop [1417]. stop-and-restart [1417]. stopping [390]. Storage [1, 4, 88, 106, 444, 467, 531, 593, 722, 767, 958, 1022, 1026, 1117, 1156, 1225, 1230, 1269, 1326, 1341, 1358, 1359, 1361, 1434, 1509, 1544, 1609, 1638, 1655, 1663, 1676, 1691, 1743, 1812, 1922, 1981, 1983, 2011, 2103, 2149, 2215]. storage-based [2103]. store [539, 1296, 2338]. stored [763, 1697]. Storm [2305]. StoryCube [917]. storytelling [917]. straight [1634]. Straightforward [1380]. Strassen [368]. Strategic [871]. Strategies [280, 398, 470, 632, 692, 941, 948, 1183, 1570, 2082, 2206]. Strategy [72, 116, 184, 252, 270, 281, 389, 416, 431, 625, 647, 755, 811, 879, 913, 1073, 1230, 1282, 1341, 1611, 1719, 1896, 1951, 1962, 2204, 2209]. Streaker [1826]. Stream [247, 296, 322, 349, 524, 926, 1352, 1594, 1697, 1754, 1867, 1910, 2182, 2196, 2326]. stream-based [1594]. stream-parallel [296]. Streaming [39, 68, 349, 467, 533, 544, 586, 704, 743, 1578, 1971, 2187, 2298]. StreamPI [296]. streams [1463, 1519, 1776, 1827, 2259, 2279]. streets [827]. stride [1244]. string [1732, 1733, 2358]. strip [1827]. strip-aware [1827]. stripe [763, 1816]. stripe-based [1816]. striped [1827]. stroke [1891]. structural [594, 839, 1541, 1694]. structural-similarity-based [594]. structure [56, 168, 329, 531, 1116, 1227, 1647, 1668, 1741, 1955, 2017, 2062, 2159, 2160, 2326]. structure-free [1647]. structured [1076, 2181]. structures [1254]. STT [1720, 1798, 2285, 2306]. Study [95, 144, 198, 256, 295, 324, 364, 370, 450, 468, 563, 573, 580, 618, 643, 704, 801, 913, 925, 933, 1046, 1085, 1102, 1129, 1146, 1238, 1245, 1260, 1265, 1275, 1329, 1330, 1352, 1422, 1491, 1567, 1648, 1722, 1781, 1782, 1811, 1888, 1900, 1951, 1957, 2057, 2094, 2107, 2150, 2211, 2281, 2282, 2284, 2286, 2358, 2362]. studying [976]. style [2057]. suave [871]. sub [549, 1223, 1348, 2280]. sub-community [1223]. sub-ontology [549]. sub-patterns [1348]. sub-trajectory [2280]. subcarrier [690]. subgraph [1105, 1325, 1500]. Subjective [1000, 2162]. submission [1972]. suboptimal [2245]. subroutines [133]. subscribe [79]. subscription [845]. subscription-based [845]. subsequence [1346]. subsequent [877]. subspace [2012, 2052]. subspaces [982, 1794]. substitution [370]. subtree [80]. successive [2321]. sufficiently [114]. suffixes [1801]. suitability [2183]. suitable [533, 663, 1759]. suite [860, 1372]. sum [1717]. SUMIS [2047]. summarization [1882]. summarizing [387]. sums [2221]. Sunway [2176]. super [656, 1685, 1763]. super-peer [656]. super-resolution [1685]. supercomputer [1377, 1596]. supercomputers [55, 640, 1071, 1495, 1761, 2023]. Supercomputing [374, 392, 396, 543, 1041, 1099, 1198]. superpoints [429]. superscalar [1158]. supervised [2277]. supervising [1013]. supplementary [1068]. supplementing [1264]. supply [1487, 1581]. support [215, 218, 220, 348, 475, 977, 1008, 1060, 1100, 1140, 1187, 1188, 1272, 1286, 1702, 1760, 1782, 1956, 2238]. supported [231, 1448, 1460, 1653]. Supporting [97, 199, 465, 723, 917, 1346, 1513, 1548, 1578, 1815, 1834, 1860, 2181]. surface [330, 711, 800, 839]. surfaces [1070, 2030, 2157]. surrogate [1656]. surveillance [554, 887, 967, 1456– 1458, 1460, 1462–1464, 1466, 1467, 1469, 2275]. Survey [445, 511, 617, 749, 1004, 1150, 1183, 1528, 1534, 1679, 1744, 1800, 1925, 2220, 2230, 2290, 2311]. survivability [107, 693]. sustainability [2278]. sustainable [1356].

SVD [1572]. SVM [2270]. SW [2039]. swap [2165]. swarm [569, 631, 774, 775, 816, 1048, 1342, 1507, 1632, 1658, 1674, 1764, 1847, 2114]. switch [1681, 1704]. switch-centric [1681]. Switches [102, 896, 1114]. switching [61, 1122, 1853]. SX [1596]. Symbol [1310]. symbolic [700]. Symmetric [1174, 1366, 1480, 1523]. symmetry [1753]. Synchronization [51, 1051, 1093, 1160, 1361, 1466, 1467, 1661, 1687, 1702, 1979, 2194]. synchronizations [1719]. synchronized [1756]. synchronous [1474, 1605, 1775]. synergies [749]. synthesis [311, 1400, 1497, 1510, 2183]. synthesized [546]. Synthetic [1232]. System [24, 57, 66, 71, 79, 87, 93, 97, 99, 153, 165, 234, 266, 286, 301, 331, 413, 500, 501, 603, 620, 651, 680, 690, 694, 736, 742, 746, 788, 829, 872, 904, 909, 910, 967, 969, 1022, 1040, 1084, 1091, 1111, 1209, 1210, 1215, 1221, 1225, 1243, 1272, 1290, 1298, 1303, 1331, 1367, 1382, 1397, 1415, 1419, 1438, 1447, 1456, 1458, 1460, 1479, 1488, 1499, 1508, 1540, 1543, 1544, 1551, 1560, 1568, 1579, 1650, 1658, 1678, 1685, 1691, 1731, 1734, 1777, 1786, 1789, 1807, 1824, 1860, 1867, 1887, 1890, 1891, 1913, 1929, 1947, 2001, 2002, 2028, 2040, 2065, 2084, 2086, 2135, 2208, 2252, 2299, 2300, 2318]. system [6, 28, 40, 61, 108, 124, 162, 195, 218, 342, 414, 418, 427, 428, 446, 447, 520, 529, 592, 687, 778, 807, 853, 878, 911, 912, 934, 964, 977, 1006, 1057, 1273, 1295, 1372, 1455, 1457, 1546, 1562, 1593, 1599, 1652, 1672, 1707, 1725, 1740, 1874, 1978, 1980, 2093, 2124, 2145, 2189, 2199, 2202, 2234, 2275, 2278, 2282]. system-enriched [1887]. system-on-chip [87]. System-wide [1789]. systematic [804, 1275, 1743, 2019]. Systems [1, 5, 23, 44, 106, 135, 164, 178, 181, 184, 201, 203, 210, 219, 255, 260, 261, 267, 289, 298, 320, 321, 333, 335, 355, 362, 410, 470, 480, 483, 585, 589, 596, 606, 617, 684, 704, 747, 763, 784, 795, 798, 817, 819, 836, 843, 895, 899, 931, 946, 953, 962, 979, 1012, 1016, 1018, 1049, 1050, 1164, 1185, 1197, 1205, 1260, 1266, 1269, 1279, 1356, 1386, 1387, 1397, 1402, 1403, 1411, 1437, 1480, 1518, 1532, 1571, 1648, 1660, 1695, 1749, 1799, 1820, 1832, 1836, 1837, 1982, 1991, 2029, 2038, 2055, 2158, 2182, 2217, 2243, 2319, 2348, 2360]. systems [9, 11, 31, 46, 62, 63, 77, 83, 88, 110, 140, 176, 183, 189, 193, 206, 215, 221, 254, 259, 270, 271, 278, 290, 291, 305, 332, 365, 394, 431, 436, 453, 471, 477, 492, 531, 590, 600, 613, 625, 629, 643, 693, 710, 722, 729, 737, 789, 805, 822, 828, 887, 890, 930, 972, 1021, 1026, 1039, 1055, 1082, 1100, 1121, 1144, 1155, 1183, 1203, 1230, 1234, 1293, 1316, 1317, 1418, 1421, 1425, 1433, 1473, 1477, 1482, 1490, 1502, 1509, 1531, 1588, 1592, 1719, 1750, 1776, 1784, 1827, 1845, 1848, 1853, 1856, 1878, 1895, 1969, 2047, 2172, 2235, 2242, 2303, 2333, 2361, 2365]. systems [56, 69, 163, 883, 1058, 1162, 1498, 1620, 1789, 1847, 1965, 1981, 2006, 2193]. SZTS [1372].

Table [483, 1630, 1997, 1999]. tables [1530]. tabu [628, 1630, 2370]. tackling [1352]. tactical [1337]. tactics [1252]. tag [2210]. tagless [355]. tags [1465, 1669]. taint [983]. tame [17]. tamper [1336, 1842]. tampering [2141]. tangible [917]. Tango [1454]. tape [577]. target [10, 153, 559, 671, 1496, 1504]. targeted [1881]. Task [8, 19, 46, 118, 123, 201, 211, 242, 252, 327, 360, 457, 486, 519, 529, 647, 792, 795, 805, 995, 1002, 1024, 1047, 1050, 1056, 1071, 1100, 1134, 1158, 1168, 1285, 1502, 1518, 1533, 1570, 1590, 1682, 1710, 1764, 1765, 1784, 1973, 1994, 2003, 2004, 2078, 2101, 2128, 2362]. task-based [1710, 1765]. task-level [8]. task-temperature [327]. tasking [1024]. tasks [211, 242, 252, 276, 382, 394, 418, 661, 721, 1016, 1069, 1357, 1620, 1686, 1711, 1773, 1905, 2156, 2204, 2243]. tasks-to-processors [382]. Taxonomy [450, 965, 1108, 1528, 1710, 1876, 1993, 2244, 2290]. TB [856]. TB-SnW [856]. TBB [1713]. TCAMs [859]. TCP [236, 1170, 2287]. TDMA [438]. teaching [1245, 1435]. Teaching-Learning-Based [1435]. team [657, 701, 719]. technique [108, 139, 146, 214, 245, 311, 372, 691, 722, 859, 902, 1318, 1612, 1613, 1622, 1632, 1682, 1887, 1905, 1913, 1914, 2071, 2076, 2095, 2123, 2161, 2207, 2243, 2266]. techniques [3, 4, 28, 84, 207, 280, 292, 375, 505, 525, 834, 1004, 1317, 1370, 1481, 1537, 1558, 1590, 1725, 1743, 1853, 1875, 1957, 2010, 2043, 2220, 2235, 2244, 2313, 2371]. Technologies [27, 289, 620, 660, 900, 909, 1207, 1436, 1710, 1837, 2051, 2352]. technology [416, 551, 669, 1060, 1140, 1215, 1298, 1334, 1456, 1508, 1845, 1865, 1869, 2251, 2256, 2273, 2359, 2374]. TEL [2092]. tele [1544]. tele-health [1544]. telecommunication [1985]. telecommunications [1957]. telehealth [1351]. telemarketer [1291]. telepresence [1579]. temperature [327, 1111, 1749, 2243]. temperature-aware [2243]. tempered [2345]. template [676]. temporal [1468, 1807, 2007]. temporary [2006]. tenant [1890]. tensor [545, 2017, 2091]. Term [444, 1950, 2314]. terminals [2337]. terms [356, 608]. terrain [1395, 1423]. test [1005, 1298]. testbed [1856]. testing [1330, 1725, 1945, 1952]. tests [256]. tetrahedral [724]. text [1632, 1703, 2261]. textual [2083]. texture [258, 1573]. their [182]. theoretic [76, 1141, 1145, 1146, 1616]. theoretic-based [1616]. Theoretical [100, 1390, 1706, 1828]. Theory [89, 464, 486, 491, 847, 849, 1148, 1149, 1226, 1730, 1962, 1970, 2107, 2112, 2335]. theory-based [1149, 2112]. Thermal [274, 327, 496, 792, 1111, 1203, 1864, 2042]. thermal-aware [1864]. thin [1918]. thing [1850]. Things [451, 521, 708, 1213, 1214, 1326, 1463, 1470, 1580, 1670, 1795, 1808, 1837, 1839, 1843, 1845, 1847, 1848, 1853, 1855, 1861, 1927, 1943, 1959, 2197, 2346, 2347]. thread [443, 744, 797, 899, 1271, 1369, 1705, 1759, 2334]. thread-intensive [443]. thread-level [1759]. threaded [473, 1181, 1182, 1404]. threading [642, 2078, 2172]. threads [121]. Threats [1872, 1879, 1881]. Three [118, 370, 527, 665, 914, 1070, 1102, 1103, 1152, 1347, 1438, 1486, 1570, 1804, 1875, 2118, 2230, 2356]. Three-dimensional [118, 370, 527, 1070, 1102, 1103, 1152, 1570, 2118, 2230, 2356]. three-factor [1347, 1486]. three-party [665, 914]. three-phase [1438]. threshold [281, 792, 864, 1368, 1657]. threshold-based [281, 864, 1368, 1657]. thresholding [1468]. thresholds [2024]. throttling [1396, 2113]. throughput [698, 1271, 1374, 1405, 2242]. tier [194, 1097, 1683]. tier-based [194]. tiering [1057]. tile [630, 763, 873]. TILE64 [586]. tiled [336, 643]. tiled-CMPs [336]. tiles [2123]. tiling [475]. tilt [629]. Time [44, 54, 190, 219, 259, 286, 342, 385, 394, 411, 421, 423, 428, 496, 558, 578, 589, 692, 782, 817, 854, 862, 868, 877, 883, 889, 937, 962, 972, 989, 991, 995, 998, 1008, 1011, 1012, 1017, 1029, 1065, 1077, 1082, 1093, 1155, 1161, 1203, 1220, 1346, 1348, 1350, 1382, 1386, 1403, 1404, 1423, 1454, 1463, 1522, 1524, 1569, 1583, 1591, 1615, 1620, 1637, 1660, 1685, 1702, 1747, 1752, 1760, 1770, 1786, 1790, 1875, 1905, 1938, 1963, 1982, 2003, 2025, 2028, 2060, 2070, 2085, 2122, 2131, 2147, 2156, 2167, 2197, 2238, 2302, 2303, 2322, 2344]. time-aware [1077, 2131]. time-complexity [2344]. time-constrained [1065]. time-delay [962]. time-dependent [286]. Time-energy [2156]. time-evolving [2167]. time-invariant [937]. Time-of-arrival [1082]. Time-sensitivity-aware [2303]. time-series [877]. Time-shift [1770]. times [242, 1304]. TLS [1400]. TM [645, 1037]. Toeplitz [596]. Toffoli [2308]. Token [1592, 2298]. Token-based [1592, 2298]. tokens [764]. tolerance [74, 108, 333, 617, 632, 799, 881, 1162, 1224, 2143, 2333]. Tolerant [8, 47, 52, 102, 103, 233, 336, 345, 389, 638, 669, 698, 810, 833, 856, 1044, 1136, 1185, 1230, 1385, 1417, 1508, 1532, 1623, 1772, 1819, 1865, 1871, 1903, 2192, 2206, 2251, 2255, 2329, 2367]. tolerate [1385, 2363]. tolerating [982]. tool [826, 917, 927, 1013, 1131, 1393, 1921, 2072]. toolkit [641]. tools [40, 43, 155, 171, 595, 1384, 1945]. top [1368]. top- [1368]. topic [2131, 2264]. Topological [251, 353, 648]. topologies [827, 851, 1101, 1255, 1379, 1871]. Topology [235, 248, 249, 311, 317, 645, 753, 863, 980, 1044, 1157, 1189, 1248, 1285, 1485, 1495, 1529, 1539, 1779, 1992]. TOPSIS [2341]. Tor [689, 1875]. tori [1032]. Torus [70, 233, 248, 368, 527, 831, 1028, 1112, 1135, 1495, 1992]. torus-connected [233, 1495]. tournaments [955]. town [69]. TPR* [1973]. TPR*-Tree [1973]. trabeculation [2072]. trace [1290]. trace-based [1290]. traceroute [415]. traces [1610, 2358]. tracing [1297]. tracking [10, 146, 153, 558, 671, 887, 931, 983, 1504, 1549, 1734, 2093, 2146, 2227, 2273, 2374]. trade [1282, 1789, 2256]. trade-off [1282, 1789]. trade-offs [2256]. tradeoff [299, 1489]. Tradeoffs [101, 568]. trading [535, 536]. Traffic [9, 26, 49, 216, 391, 578, 746, 899, 1045, 1067, 1106, 1141, 1194, 1327, 1328, 1360, 1555, 1690, 1693, 1873, 1945, 1975, 2199]. trainable [1791]. training [1990, 2238, 2273, 2374]. trajectory [1973, 2280]. trajectory-based [1973]. transaction [269, 1293, 1696, 2214]. transactional [137, 435, 504, 784, 790, 1182, 1191, 1374, 1527, 1726, 2190]. transactions [1726]. transcoding [36, 2081]. Transfer [72, 131, 1264, 1274, 1701, 2021, 2143, 2219]. transference [706]. transfers [244, 402, 1405]. Transform [397, 470, 587, 953, 1152, 1593, 1634, 1729, 2053, 2064, 2226]. Transformation [317, 669, 1161, 2371]. transformations [1535]. transient [1324]. transition [1825, 1975]. transitions [1118]. translation [328, 626, 702, 1117]. translator [1087, 1302]. transmission [93, 349, 654, 690, 810, 956, 1462, 1971]. transmissions [883]. transparent [2166]. transport [603, 1427]. transportation [1372, 1703, 2107]. transports [1875]. Trasgo [165]. travel [1900]. traveler [2143]. traveling [1667]. travelling [1375]. traversal [898]. treatments [2042]. Tree [34, 50, 216, 230, 237, 249, 339, 488, 581, 597, 682, 1101, 1116, 1180, 1370, 1540, 1825, 1939, 1973, 2247, 2254, 2267]. Tree-based [1370, 1939]. tree-search [597]. Tree-turn [230]. Trees [80, 113, 142, 616, 993, 1031, 1045, 1378, 1389]. trend [1338]. trends [27, 357]. tri [1006]. tri-diagonal [1006]. triangle [869]. triangulated [2157]. triangulation [1175]. tridiagonal [1498, 2065]. Tridimensional [175]. trigger [1861]. trigger-based [1861]. trigonometric [2045]. Trilinos [1835]. triples [1296]. truly [2324]. truncation [2076]. Trust [41, 86, 509, 512, 517, 856, 884, 965, 1094, 1212, 1343, 1531, 1622, 1730, 1763, 1811, 1955, 2126, 2162, 2163]. Trust-based [856]. trusted [82, 83, 85, 912]. trustworthiness [2082]. TSM [976]. TSP [1353]. tsunami [1771, 1786]. tubular [1474]. tuned [1188]. Tuning [157, 530, 946, 1233, 1381, 1394, 1411, 1701, 1722, 2181, 2363, 2364]. Turing [577]. turn [230, 1555, 1819]. tutorial [1706]. TV [854, 971, 992]. Twitter [1809, 2277]. Two [218, 388, 483, 672, 800, 820, 966, 976, 1015, 1138, 1177, 1262, 1315, 1475, 1522, 1683, 1811, 1844, 1934, 1988, 1998]. Two-dimensional [218, 388, 820, 1988]. two-disjoint [1998]. two-hop [1262]. Two-level [483, 1315, 1844]. two-sided [976]. Two-stage [1015, 1522]. two-tier [1683]. Twofish [192]. Type [320, 880, 1723]. types [31, 1323, 2324].

U [66, 1091]. u-BabSang [66]. U-Health [1091]. UAV [968, 2227]. UAV-assisted [968]. UAVs [671]. uBench [608]. ubiquitous [69, 93, 95, 351, 489, 491, 494, 498, 537, 553, 1695, 2171, 2210]. ubiquity [621]. UHD [1685]. ULFM [1401]. ultra [1458, 1465, 1669, 1687]. ultra-HD [1458]. ultra-lightweight [1465, 1669, 1687]. Ultrafast [1492]. ultralightweight [1580]. ultrascale [1356]. UML [1161, 2371]. UML/MARTE [1161]. unannotated [2105]. unbalanced [418, 480]. uncertain [1370, 1637, 2253]. underlying [1272]. Understanding [309, 1022]. underwater [1932]. unequal [1755]. unfairness [1379]. unicast [362, 757]. Unified [329, 472, 525, 600, 742, 981, 1659, 2226, 2316]. uniform [700, 1027, 1315, 2345]. uniformization [1315]. Unit [525, 631, 1116, 1320, 1322, 1464, 2246]. unit-accelerated [1320]. units [132, 627, 712, 762, 814, 1107, 1480, 1571, 1744]. universal [911]. university [815]. unknown [1059]. Unleashing [2354]. Unmixing [934]. Unmixing-based [934]. unreliable [729]. unsafety [1044]. Unstructured [11, 41, 245, 246, 512, 724]. Unsupervised [50, 1632, 2009, 2252]. unsymmetrical [1501]. up-conversion [1553, 1685]. UPC [479, 952]. update [484, 503, 737, 1470, 2032, 2062]. updates [385]. Updating [912]. upfront [1995]. Uplink [654, 1808]. UPnP [97]. UPnP-based [97]. upon [662]. UPS [1225]. upward [1101]. urban [575, 578, 1217, 2199, 2330]. URL [1880]. usability [2083]. usage [1609, 1753, 1921, 1926, 1947]. use [291, 815, 820, 1085, 1330, 1510, 2076]. used [1192, 1275]. User [307, 457, 515, 674, 705, 754, 845, 908, 1039, 1051, 1061, 1177, 1249, 1345, 1392, 1810, 1884, 1923, 2096, 2234]. User-centric [1039, 1392]. user-friendly [754]. user-level [1810]. user-selectable [457]. users [220, 588, 1089, 1538, 2277]. Using [56, 58, 129, 156, 177, 188, 212, 214, 236, 250, 268, 283, 378, 410, 442, 486, 539, 546, 573, 589, 605, 658, 674, 712, 746, 775, 808, 834, 837, 908, 973, 1004, 1014, 1045, 1050, 1090, 1093, 1095, 1107, 1110, 1182, 1216, 1226, 1232, 1241, 1254, 1298, 1336, 1348, 1353, 1366, 1373, 1394, 1398, 1416, 1458, 1481, 1502, 1543, 1558, 1618, 1629, 1650, 1658, 1667, 1762, 1774, 1799, 1807, 1817, 1832, 1836, 1842, 1844, 1859, 1882, 1888, 1906, 1908, 1926, 1958, 1989, 1996, 2008, 2013, 2017, 2030, 2049, 2073, 2105, 2147, 2149, 2201, 2221, 2252, 2254, 2281, 2348, 2359, 2360, 2371]. using [28, 71, 216, 229, 242, 292, 311, 324, 464, 496, 516, 520, 572, 583, 646, 649, 676, 702, 707, 716, 724, 735, 743, 791, 796, 797, 803, 868, 871, 909, 936, 937, 950, 1010, 1017, 1057, 1103, 1116, 1144, 1170, 1245, 1247, 1263, 1315, 1327, 1328, 1401, 1415, 1433, 1457, 1468, 1488, 1492, 1503, 1509, 1531, 1540, 1547, 1561, 1570, 1590, 1604, 1624, 1628, 1652, 1666, 1678, 1721, 1755, 1777, 1816, 1853, 1873, 1875, 1900, 1973, 2051, 2069, 2081, 2090, 2101, 2103, 2133, 2136, 2144, 2150, 2161, 2195, 2241, 2247, 2266, 2309, 2312– 2314, 2325, 2335, 2336]. using [94, 107, 146, 193, 342, 376, 387, 428, 491, 636, 644, 756, 794, 810, 988, 1046, 1133, 1162, 1295, 1374, 1465, 1634, 1641, 1655, 1668, 1713, 1740, 1787, 1839, 1981, 2087, 2280, 2293, 2305]. UTFLA [1315]. utility [536, 797, 1039, 1190]. utility-oriented [1190]. utilization [248, 278, 286, 640, 766, 841, 948, 1135, 1700, 1925, 2114]. Utilizing [706, 1974, 2299].

v1.3 [1704]. vacation [1962]. Validation [131, 420, 1840]. validity [2312]. valuation [1636]. valuation-based [1636]. value [84, 1094, 2069]. value-based [1094]. VANET [1083, 1217, 1552]. VANETs [728, 2140]. vantage [2368]. VAP [2130]. variability [1509]. Variable [93, 506, 2078]. variable-precision [506]. variable-rate [93]. variance [417]. variant [2165]. variation [2081]. variations [570]. various [1427]. varying [1282, 1987]. VBF [1932]. VBF-based [1932]. VDS [1327, 1328]. vector [48, 162, 376, 454, 517, 970, 1115, 1188, 1286, 1596, 1666, 1735, 2238]. vectorization [1412, 1713]. vectorized [1769]. vectors [1044, 1549]. VectorTrust [517]. vehicle [558, 972, 1090, 1221, 1558, 1690, 1876, 2135]. Vehicles [1858]. vehicular [406, 728, 810, 1052, 1144, 1217, 1254, 1677, 1688, 1690, 1702, 1762]. ventricle [2072]. verbal [2083]. Verifiable [852, 1638, 1983]. verification [392, 448, 511, 584, 685, 725, 889, 1041, 1128, 2142, 2198, 2255]. verifier [911]. verify [255]. verifying [221]. versatile [55]. version [579, 2059, 2061, 2067]. versus [261, 386, 1000, 1417, 1444, 1829]. vertex [73, 630, 750, 2225]. vertex-centric [2225]. vertex-disjoint [73]. vertical [876, 1193]. very [1407, 1593]. via [98, 154, 187, 275, 304, 898, 899, 935, 1103, 1271, 1488, 1499, 1521, 1529, 1557, 1675, 1701, 1705, 1828, 1838, 1910, 2192, 2358]. vibration [928]. victim [1999]. Video [68, 152, 349, 554, 559, 586, 656, 754, 970, 979, 1092, 1099, 1462, 1464, 1578, 1685, 1971, 2059, 2081, 2103, 2227, 2273, 2275, 2276, 2298, 2374]. video-equipped [1099]. view [435, 453, 1971, 2154]. view-oriented [435]. ViMediaNet [1579]. vindictive [885]. violation [1982]. Virtex [1704]. Virtex-6 [1704]. Virtual [20, 86, 291, 292, 303, 384, 424, 426, 468, 681, 697, 703, 719, 721, 842, 846, 881, 891, 975, 990, 1108, 1135, 1187, 1204, 1241, 1247, 1252, 1288, 1505, 1506, 1673, 1756, 1782, 1783, 1901, 1919, 1920, 1924, 1929, 1960, 1968, 1979, 1995, 2034, 2099, 2100, 2130, 2173, 2248, 2262, 2286, 2290, 2293, 2309, 2323, 2336, 2341]. Virtualization [756, 881, 898, 1381, 1473, 1477, 1840, 1908, 1968, 2019]. virtualized [210, 543, 987, 1055, 1209, 1678, 1747, 1760, 1777, 1780, 1796, 1935, 2340]. Visible [675, 916]. Vision [295, 557]. Vision-based [557]. visual [147, 1364, 1800, 1959, 2083]. Visualization [195, 1026, 1395, 1423, 2064]. visualizations [340]. visualize [538, 555]. visually [1454]. VLIW [337]. VLSI [861]. VM [1199, 1450, 1489, 1612, 1613, 1616, 1862, 1870, 1925, 2101, 2201, 2240]. VM-to-hypervisor [1616]. VMBKS [1978]. VMDFS [1924]. VMM [680, 1870]. VMM-based [680]. VMs [1228]. VoIP [514]. volatile [1826, 2177]. volatility [442, 1259]. Volta [2316]. Voltage [1012, 1174, 1324, 1396]. voltage-frequency [1396]. volume [2, 724]. volumes [1154]. volunteer [46, 585, 1267, 1792]. voting [911]. VQ [1820, 1844]. VQ-based [1844]. vulnerabilities [1692]. vulnerability [791, 2130]. vulnerable [1952].

W2T [521]. Wait [856]. waiting [211]. Walker [765]. walking [2143]. walks [45]. WAN [1960]. warehouse [1625]. warehouses [774]. warehousing [1850]. warp [841]. warping [1403]. waste [1001]. Wasted [2113]. WatCache [2006]. Water [163, 603, 1357, 1397, 1438, 1536]. Water-level [1357]. watermarking [522, 675, 916]. Wave [2221]. waveform [2116]. wavelength [378, 883]. wavelength-time [883]. Wavelet [397, 470, 587, 953, 1102, 1152, 1729, 2064]. wavelet-based [1102]. way [1163]. WBANs [1787]. weak [1768]. weakly [47, 765]. wearable [1813, 1814, 1891]. weather [947, 2283]. Web [60, 159, 291, 521, 538, 555, 598, 621, 652, 793, 815, 872, 892, 1017, 1087, 1213, 1237, 1369, 1652, 1692, 1849, 1850, 1952, 2120, 2144, 2208, 2253, 2289]. web-based [2208]. WebCL [1087]. webinos [621]. WebRTC [1351]. website [2083]. WECPAR [994]. Weibo [1146]. weight [80, 523, 702, 999, 1288, 1956, 2088, 2107]. weight-constrained [80]. Weighted [682, 790, 1431, 2247, 2342]. weights [1790]. Weil [567]. well [116]. well-balanced [116]. wellness [1331, 1340, 2086]. whales [2292]. White [377]. White-light [377]. wide [967, 1789]. wide-area [967]. width [1042]. Wiedemann [1363]. WiMAX [520]. Wind [1412, 2073]. Window [2187, 2259]. Window-based [2259]. windows [666, 2363]. winner [1636]. winner-bid [1636]. wire [2135]. wired [402]. Wireless [6, 10, 14, 25, 30, 68, 119, 154, 194, 250, 288, 346, 398, 401, 438, 511, 516, 522, 544, 554, 560, 569, 571, 622, 623, 653, 690, 699, 736, 749, 752, 771, 830, 851, 907, 914, 963, 968, 1036, 1065, 1090, 1124, 1136, 1139, 1195, 1202, 1219, 1273, 1324, 1330, 1335, 1448, 1453, 1460, 1463, 1471, 1472, 1496, 1504, 1513, 1529, 1556, 1557, 1563, 1647, 1657, 1679, 1683, 1694, 1700, 1702, 1730, 1739, 1755, 1833, 1881, 1891, 1954, 1964, 1966, 2020, 2087, 2104, 2109, 2110, 2126, 2137, 2151, 2169, 2207, 2213, 2230, 2267, 2275, 2312, 2326, 2331]. wiring [1741]. Wisdom [521]. wise [1717, 1849]. WiseThrottling [1134]. without [462, 1283, 1370, 1741]. WLAN [1082, 1270]. WLANs [691]. WLS [1431]. word [5, 2083]. word-of-mouth [2083]. work [1703, 1711]. Workflow [13, 35, 57, 421, 423, 424, 431, 467, 685, 696, 718, 731, 775, 1002, 1047, 1150, 1168, 1307, 1358, 1449, 1507, 1522, 1541, 1595, 1662, 1949, 2016]. workflows [535, 1594, 1637, 1767, 1773, 1780, 1866, 1938, 2170, 2301]. Workload [169, 292, 327, 455, 480, 755, 816, 952, 1022, 1057, 1282, 1627, 1641, 2006, 2223, 2289]. workload-aware [292, 1627, 2006]. workload-driven [455]. workloads [60, 1384, 1633, 1720, 1980]. workstations [81]. world [809, 2015, 2037]. worm [416]. wormhole [368, 923, 1650]. wormhole-routed [368]. worms [12]. write [1576, 2299]. write-friendly [2299]. written [1771]. WSC [2120]. WSN [673, 2178]. WSNs [493, 593, 1640, 2327].

X [248, 2254]. X-torus [248]. X-tree [2254]. Xdraw [1772]. XenoCluster [131]. Xeon [1311, 1570, 1641, 1644, 1774, 1990, 1996, 2037, 2322]. XML [727, 1451, 1519, 2168, 2326]. XP [2237]. XQuery [1279]. XtratuM [446, 447]. XtratuM/PPC [446, 447]. XY [231]. XYZ [2020]. XYZ-planar [2020].

YARN [1577]. years [2019]. yoking [1738].

Zero [380, 1371]. Zero-energy [380]. zeroknowledge [1371]. Zeroing [1935]. ZigBee [391, 547]. zones [1703].

References

Carretero:2010:SSS

[1] Jesus Carretero and J. Daniel Garcia. Scalable storage systems and high-performance applications. The Journal of Supercomputing, 51(1): 1–2, January 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=1&spage=1.

Agrawal:2010:EID

[2] Anupam Agrawal, Josef Kohout, Gordon J. Clapworthy, Nigel J. B. Mc-Farlane, Feng Dong, et al. Enabling the interactive display of large medical volume datasets by multiresolution bricking. The Journal of Supercomputing, 51(1):3–19, January 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=1&spage=3.

Lee:2010:PFR

[3] Joahyung Lee and Inbum Jung. Parallel failure recovery techniques in cluster-based media servers. The Journal of Supercomputing, 51(1):20– 39, January 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=1&spage=20.

Nunez:2010:NTS

[4] Alberto Núñez, Javier Fernández, Jose D. Garcia, Félix Garcia, and Jesús Carretero. New techniques for simulating high performance MPI applications on large storage networks. The Journal of Supercomputing, 51 (1):40–57, January 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=1&spage=40.

Drews:2010:SPW

[5] Frank Drews, Jens Lichtenberg, and Lonnie Welch. Scalable parallel word search in multicore/multiprocessor systems. The Journal of Supercomputing, 51(1):58–75, January 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=1&spage=58.

Diaz:2010:SSN

[6] Soledad Escolar Díaz, Florin Isaila, Alejandro Calderón Mateos, Luis Miguel Sanchez García, and David E. Singh. SENFIS: a Sensor Node File System for increasing the scalability and reliability of Wireless Sensor Networks applications. The Journal of Supercomputing, 51(1):76– 93, January 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=51&issue=1&spage=76.

Li:2010:HPC

[7] Keqiu Li, Geyong Min, Laurence Yang, and Yongquan Wei. High performance computing and communications. The Journal of Supercomputing, 51(2):95–96, February 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=2&spage=95.

Wu:2010:ATL

[8] Yongwei Wu, Yulai Yuan, Guangwen Yang, and Weimin Zheng. An adaptive task-level fault-tolerant approach to Grid. The Journal of Supercomputing, 51(2):97–114, February 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=2&spage=97.

Wu:2010:AMN

[9] Yulei Wu, Geyong Min, Mohamed Ould-Khaoua, Hao Yin, and Lan Wang. Analytical modelling of networks in multicomputer systems under bursty and batch arrival traffic. The Journal of Supercomputing, 51 (2):115–130, February 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=2&spage=115.

Guo:2010:DEE

[10] Minyi Guo, Elizabeth Olule, Guojun Wang, and Song Guo. Designing energy efficient target tracking protocol with quality monitoring in wireless sensor networks. The Journal of Supercomputing, 51(2): 131–148, February 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=51&issue=2&spage=131.

Qu:2010:SFS

[11] Wenyu Qu, Wanlei Zhou, and Masaru Kitsuregawa. Sharable file searching in unstructured peer-to-peer systems. The Journal of Supercomputing, 51 (2):149–166, February 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=51&issue=2&spage=149.

Fan:2010:DAP

[12] Xiang Fan and Yang Xiang. Defending against the propagation of active worms. The Journal of Supercomputing, 51(2):167– 200, February 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=51&issue=2&spage=167.

Cao:2010:DED

[13] Haijun Cao, Hai Jin, Xiaoxin Wu, Song Wu, and Xuanhua Shi. DAGMap: efficient and dependable scheduling of DAG workflow job in Grid. The Journal of Supercomputing, 51(2): 201–223, February 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=2&spage=201.

Yang:2010:RHR

[14] Panlong Yang and Guihai Chen. Rematch: a highly reliable scheduling algorithm on heterogeneous wireless mesh network. The Journal of Supercomputing, 51(2):224– 240, February 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=2&spage=224.

Huang:2010:P

[15] Zhiyi Huang, John Hine, Laurent Lefèvre, Tony McGregor, Yi Pan, et al. Preface. The Journal of Supercomputing, 51(3):241– 243, March 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=3&spage=241.

Leiserson:2010:CCP

[16] Charles E. Leiserson. The Cilk++ concurrency platform. The Journal of Supercomputing, 51(3):244– 257, March 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=51&issue=3&spage=244.

Leung:2010:DRT

[17] K. Leung, Z. Huang, Q. Huang, and P. Werstein. Data race: tame the beast. The Journal of Supercomputing, 51(3):258–278, March 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=51&issue=3&spage=258.

Tanabe:2010:EMN

[18] Noboru Tanabe, Hirotaka Hakozaki, Hiroshi Ando, Yasunori Dohi, Zhengzhe Luo, et al. An enhancer of memory and network for applications with large-capacity data and noncontinuous data accessing. The Journal of Supercomputing, 51(3):279– 309, March 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=51&issue=3&spage=279.

Shahul:2010:STG

[19] Ahmed Zaki Semar Shahul and Oliver Sinnen. Scheduling task graphs optimally with A∗. The Journal of Supercomputing, 51(3):310– 332, March 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=51&issue=3&spage=310.

Zou:2010:LVO

[20] Yongqiang Zou, Li Zha, Xiaoning Wang, Haojie Zhou, and Peixu Li. A layered Virtual Organization architecture for grid. The Journal of Supercomputing, 51(3):333– 351, March 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=3&spage=333.

Lefevre:2010:DEE

[21] Laurent Lefèvre and Anne-Cécile Orgerie. Designing and evaluating an energy efficient Cloud. The Journal of Supercomputing, 51(3):352– 373, March 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=3&spage=352.

Shorfuzzaman:2010:APD

[22] Mohammad Shorfuzzaman, Peter Graham, and Rasit Eskicioglu. Adaptive popularity-driven replica placement in hierarchical data grids. The Journal of Supercomputing, 51(3):374– 392, March 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=3&spage=374.

Ebadi:2010:FFC

[23] Toktam Ebadi, Maryam Purvis, and Martin Purvis. A framework for facilitating cooperation in multi-agent systems. The Journal of Supercomputing, 51(3):393–417, March 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=51&issue=3&spage=393.

Casey:2010:LAC

[24] John Casey and Wanlei Zhou. A locality aware cache diffusion system. The Journal of Supercomputing, 52(1):1–22, April 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=1&spage=1.

Zeng:2010:CCM

[25] Yuanyuan Zeng, Cormac J. Sreenan, Naixue Xiong, Laurence T. Yang, and Jong Hyuk Park. Connectivity and coverage maintenance in wireless sensor networks. The Journal of Supercomputing, 52(1):23–46, April 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=1&spage=23.

Isazadeh:2010:TDE

[26] Ayaz Isazadeh and Mohsen Heydarian. Traffic distribution for end-to-end QoS routing with multicast multichannel services. The Journal of Supercomputing, 52(1):47–81, April 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=1&spage=47.

Jie:2010:AAI

[27] Wei Jie, Junaid Arshad, and Pascal Ekin. Authentication and authorization infrastructure for Grids — issues, technologies, trends and experiences. The Journal of Supercomputing, 52(1):82–96, April 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=1&spage=82.

Sadik:2010:MHA

[28] Sarmad Sadik, Alade Rahman, Arshad Ali, H. Farooq Ahmad, and Hiroki Suguri. Modeling high assurance agent-based Earthquake Management System using formal techniques. The Journal of Supercomputing, 52(2):97–118, May 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=2&spage=97.

Numrich:2010:CES

[29] Robert W. Numrich. The computational energy spectrum of a program as it executes. The Journal of Supercomputing, 52(2):119– 134, May 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=2&spage=119.

Myoupo:2010:RCA

[30] Jean Frédéric Myoupo, Aboubecrine Ould Cheikhna, and Idrissa Sow.

A randomized clustering of anonymous wireless ad hoc networks with an application to the initialization problem. The Journal of Supercomputing, 52(2):135–148, May 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=2&spage=135.

Sharifi:2010:DFI

[31] Mohsen Sharifi, Seyedeh Leili Mirtaheri, and Ehsan Mousavi Khaneghah. A dynamic framework for integrated management of all types of resources in P2P systems. The Journal of Supercomputing, 52(2):149– 170, May 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=2&spage=149.

Shih:2010:PBD

[32] Wen-Chung Shih, Chao-Tung Yang, and Shian-Shyong Tseng. Performancebased data distribution for data mining applications on grid computing environments. The Journal of Supercomputing, 52(2):171– 198, May 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=2&spage=171.

Yang:2010:NBA

[33] Chao-Tung Yang, Fang-Yie Leu, and Sung-Yi Chen. Network bandwidthaware job scheduling with dynamic information model for Grid resource brokers. The Journal of Supercomputing, 52(3):199–223, June 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=3&spage=199.

Mahafzah:2010:HDP

[34] Basel A. Mahafzah and Bashira A. Jaradat. The hybrid dynamic parallel scheduling algorithm for load balancing on Chained–Cubic Tree interconnection networks. The Journal of Supercomputing, 52(3):224– 252, June 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=3&spage=224.

Chin:2010:ASS

[35] Sung Ho Chin, Taeweon Suh, and Heon Chang Yu. Adaptive service scheduling for workflow applications in Service–Oriented Grid. The Journal of Supercomputing, 52(3): 253–283, June 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=52&issue=3&spage=253.

Li:2010:CMO

[36] Keqiu Li, Yanming Shen, Kai Lin, and Wenyu Qu. Coordinated multimedia object replacement in transcoding proxies. The Journal of Supercomputing, 52(3):284– 302, June 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=52&issue=3&spage=284.

Sarbazi-Azad:2010:SIN

[37] H. Sarbazi-Azad, A. Shahrabi, and H. Beigy. Special issue on networkbased high performance computing. The Journal of Supercomputing, 53 (1):1–4, July 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=53&issue=1&spage=1.

Ababneh:2010:AJS

[38] Ismail Ababneh, Saad Bani-Mohammad, and Mohamed Ould-Khaoua. An adaptive job scheduling scheme for mesh-connected multicomputers. The Journal of Supercomputing, 53(1): 5–25, July 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=53&issue=1&spage=5.

Liu:2010:EAL

[39] Shyh Chang Liu and Tsung Hung Chen. Enabling adaptive live streaming in P2P multipath networks. The Journal of Supercomputing, 53(1): 26–44, July 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=53&issue=1&spage=26.

Xhafa:2010:EPM

[40] Fatos Xhafa, Leonard Barolli, Santi Caballé, and Raul Fernández. Efficient peerGroup management in JXTA-overlay P2P system for developing groupware tools. The Journal of Supercomputing, 53(1): 45–65, July 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=45.

Mashayekhi:2010:CST

[41] Hoda Mashayekhi and Jafar Habibi. Combining search and trust models in unstructured peer-to-peer networks. The Journal of Supercomputing, 53(1):66–85, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=66.

Flauzac:2010:CMP

[42] Oliver Flauzac, Michael Krajecki, and Luiz-Angelo Steffenel. CONFIIT: a middleware for peer-to-peer computing. The Journal of Supercomputing, 53(1):86–102, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=86.

Batista:2010:PAA

[43] Daniel M. Batista, Luciano J. Chaves, Nelson L. S. da Fonseca, and Artur Ziviani. Performance analysis of available bandwidth estimation tools for grid networks. The Journal of Supercomputing, 53(1):103–121, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=103.

Du:2010:RPM

[44] Bing Du and Chun Ruan. Robust performance modelling and scheduling of distributed real-time systems. The Journal of Supercomputing, 53 (1):122–137, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=122.

Randles:2010:BRW

[45] M. Randles, O. Abu-Rahmeh, P. Johnson, and A. Taleb-Bendiab. Biased random walks on resource network graphs for load balancing. The Journal of Supercomputing, 53(1): 138–162, July 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=138.

Lee:2010:RTS

[46] Young Choon Lee, Albert Y. Zomaya, and Howard Jay Siegel. Robust task scheduling for volunteer computing systems. The Journal of Supercomputing, 53(1):163–181, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=163.

Xu:2010:DFT

[47] Zhenyu Xu, James Wang, and Pradip K. Srimani. Distributed fault tolerant computation of weakly connected dominating set in ad hoc networks. The Journal of Supercomputing, 53(1):182–195, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=182.

Yassein:2010:NPB

[48] Muneer Bani Yassein, Mustafa Bani Khalaf, and Ahmed Y. Al-Dubai. A new probabilistic broadcasting scheme for mobile ad hoc on-demand distance vector (AODV) routed networks. The Journal of Supercomputing, 53 (1):196–211, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=196.

Abu-Tair:2010:AMA

[49] Mamun Abu-Tair, Geyong Min, Qiang Ni, and Hong Liu. An adaptive medium access control scheme for mobile ad hoc networks under self-similar traffic. The Journal of Supercomputing, 53(1):212–230, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=1&spage=212.

Yasami:2010:NUC

[50] Yasser Yasami and Saadat Pour Mozaffari. A novel unsupervised classification approach for network anomaly detection by k-means clustering and ID3 decision tree learning methods. The Journal of Supercomputing, 53 (1):231–245, July 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=53&issue=1&spage=231.

Abellan:2010:CBS

[51] José L. Abellán, Juan Fernández, and Manuel E. Acacio. Characterizing the basic synchronization and communication operations in dual Cell-based blades through CellStats. The Journal of Supercomputing, 53 (2):247–268, August 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=53&issue=2&spage=247.

Charr:2010:DFT

[52] Jean-Claude Charr, Raphaël Couturier, and David Laiymani. A decentralized and fault tolerant convergence detection algorithm for asynchronous iterative algorithms. The Journal of Supercomputing, 53(2):269– 292, August 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=53&issue=2&spage=269.

Rashid:2010:AEP

[53] Layali Rashid, Wessam M. Hassanein, and Moustafa A. Hammad. Analyzing and enhancing the parallel sort operation on multithreaded architectures. The Journal of Supercomputing, 53(2):293–312, August 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=2&spage=293.

Cho:2010:BMR

[54] Haengrae Cho and Kook yeol Yoo. Buffer management in a real-time shared disks cluster. The Journal of Supercomputing, 53(2):313– 328, August 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=2&spage=313.

Li:2010:MVF

[55] Yamin Li, Shietung Peng, and Wanming Chu. Metacube — a versatile family of interconnection networks for extremely large-scale supercomputers. The Journal of Supercomputing, 53 (2):329–351, August 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=2&spage=329.

Zhang:2010:UCD

[56] Haohua Zhang, Hai Zhao, Wei Cai, Jie Liu, and Wanlei Zhou. Using the k-core decomposition to analyze the static structure of large-scale software systems. The Journal of Supercomputing, 53(2):352–369, August 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=2&spage=352.

Cao:2010:SQB

[57] Haijun Cao, Hai Jin, Xiaoxin Wu, and Song Wu. ServiceFlow: QoSbased hybrid service-oriented grid workflow system. The Journal of Supercomputing, 53(3):371–393, September 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=3&spage=371.

Pedersen:2010:UPI

[58] Leif Pedersen and Hassan Reza. Using Pit to improve security in low-level programs. The Journal of Supercomputing, 53(3):394– 410, September 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=3&spage=394.

Yang:2010:FRM

[59] Chao-Tung Yang, Chun-Pin Fu, and Ching-Hsien Hsu. File replication, maintenance, and consistency management services in data grids. The Journal of Supercomputing, 53(3): 411–439, September 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=3&spage=411.

Sharifian:2010:ABL

[60] Saeed Sharifian, Seyed A. Motamedi, and Mohammad K. Akbari. An approximation-based loadbalancing algorithm with admission control for cluster web servers with dynamic workloads. The Journal of Supercomputing, 53(3):440– 463, September 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=53&issue=3&spage=440.

Zhao:2010:CRM

[61] Jun Zhao and Xiaohan Sun. Contention resolution mechanisms for asynchronous optical packet switching based high performance computing system. The Journal of Supercomputing, 53(3):464–474, September 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0920-8542&volume=53&
issue=3&spage=464.

Lee:2010:ISF

[62] Deok Gyu Lee, Sajid Hussain, and Alan Chin-Chen Chang. Intelligent systems for future generation communications. The Journal of Supercomputing, 54(1):1–3, October 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=1&spage=1.

Kawsar:2010:DIF

[63] Fahim Kawsar, Tatsuo Nakajima, Jong Hyuk Park, and Sang-Soo Yeo. Design and implementation of a framework for building distributed smart object systems. The Journal of Supercomputing, 54(1):4– 28, October 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=1&spage=4.

Cheng:2010:HML

[64] Bo-Chao Cheng, Yi-An Tsai, Guo-Tan Liao, and Eui-Seok Byeon. HMM machine learning and inference for Activities of Daily Living recognition. The Journal of Supercomputing, 54 (1):29–42, October 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=1&spage=29.

Lai:2010:CAM

[65] Chin-Feng Lai, Yueh-Min Huang, and Han-Chieh Chao. A contextaware multi-model remote controller for electronic home devices. The Journal of Supercomputing, 54(1):43– 60, October 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=1&spage=43.

Oh:2010:UBC

[66] Yoosoo Oh, Ahyoung Choi, and Woontack Woo. u-BabSang: a contextaware food recommendation system. The Journal of Supercomputing, 54 (1):61–81, October 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=1&spage=61.

Park:2010:PEK

[67] Jong Hyuk Park. Privacy-enhanced Key Recovery in mobile communication environments. The Journal of Supercomputing, 54(1):82– 93, October 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=1&spage=82.

Shu:2010:CAC

[68] Lei Shu, Yan Zhang, Zhiwen Yu, Laurence T. Yang, Manfred Hauswirth, and Naixue Xiong. Context-aware cross-layer optimized video streaming in wireless multimedia sensor networks. The Journal of Supercomputing, 54 (1):94–121, October 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=1&spage=94.

Yu:2010:PDE

[69] Yun Seop Yu, Seong Ho Choi, Hyung-Kun Park, Changhoon Lee, Young-Sik Jeong, and Sang-Hoon Kim. A power-, delay- and emergency-efficient protocol of ubiquitous sensor network systems for silver town applications. The Journal of Supercomputing, 54 (1):122–137, October 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=1&spage=122.

Lucas:2010:PAF

[70] Keny T. Lucas and Prasanta K. Jana. Parallel algorithms for finding polynomial roots on OTIS-torus. The Journal of Supercomputing, 54(2): 139–153, November 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=2&spage=139.

Muppavarapu:2010:RBA

[71] Vineela Muppavarapu, Anil L. Pereira, and Soon M. Chung. Role-based access control for a Grid system using OGSA-DAI and Shibboleth. The Journal of Supercomputing, 54(2): 154–179, November 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=2&spage=154.

Yang:2010:IDA

[72] Chao-Tung Yang, Shih-Yu Wang, and William Cheng-Chung Chu. Implementation of a dynamic adjustment strategy for parallel file transfer in co-allocation data grids. The Journal of Supercomputing, 54(2):

180–205, November 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=2&spage=180.

Zhou:2010:CVD

[73] Shuming Zhou, Wenjun Xiao, and Behrooz Parhami. Construction of vertex-disjoint paths in alternating group networks. The Journal of Supercomputing, 54(2):206– 228, November 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=2&spage=206.

Chen:2010:FHH

[74] Y-Chuang Chen, Yong-Zen Huang, Lih-Hsing Hsu, and Jimmy J. M. Tan. A family of Hamiltonian and Hamiltonian connected graphs with fault tolerance. The Journal of Supercomputing, 54(2):229– 238, November 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=2&spage=229.

Shih:2010:MIH

[75] Yuan-Kang Shih, Hui-Chun Chuang, Shin-Shin Kao, and Jimmy J. M. Tan. Mutually independent Hamiltonian cycles in dual-cubes. The Journal of Supercomputing, 54(2): 239–251, November 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=2&spage=239.

Wei:2010:GTM

[76] Guiyi Wei, Athanasios V. Vasilakos, Yao Zheng, and Naixue Xiong. A gametheoretic method of fair resource allocation for cloud computing services. The Journal of Supercomputing, 54(2): 252–269, November 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=2&spage=252.

Li:2010:FHS

[77] Keqin Li. Fast and highly scalable parallel computations for fundamental matrix problems on distributed memory systems. The Journal of Supercomputing, 54(3):271– 297, December 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=3&spage=271.

Sehgal:2010:SOC

[78] Vivek Kumar Sehgal and Durg Singh Chauhan. State observer controller design for packets flow control in networks-on-chip. The Journal of Supercomputing, 54(3):298– 329, December 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=54&issue=3&spage=298.

Hassan:2010:DFE

[79] Mohammad Mehedi Hassan, Biao Song, and Eui-Nam Huh. A dynamic and fast event matching algorithm for a content-based publish/subscribe information dissemination system in Sensor-Grid. The Journal of Supercomputing, 54(3): 330–365, December 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=3&spage=330.

Hsieh:2010:WCM

[80] Sun-Yuan Hsieh and Ting-Yu Chou. The weight-constrained maximumdensity subtree problem and related problems in trees. The Journal of Supercomputing, 54(3):366– 380, December 2010. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=54&issue=3&spage=366.

Shah:2010:OJP

[81] Syed Munir Hussain Shah, Kalim Qureshi, and Haroon Rasheed. Optimal job packing, a backfill scheduling optimization for a cluster of workstations. The Journal of Supercomputing, 54(3):381–399, December 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0920-8542&volume=54&
issue=3&spage=381.

Li:2011:GEH

[82] Keqiu Li, Hai Jin, and Jingwei Jin. Guest editorial: high performance trusted computing. The Journal of Supercomputing, 55(1): 1–3, January 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=1&spage=1.

Militello:2011:EAP

[83] C. Militello, V. Conti, S. Vitabile, and F. Sorbello. Embedded access points for trusted data and resources access in HPC systems. The Journal of Supercomputing, 55(1):4– 27, January 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=1&spage=4.

Cebrian:2011:LED

[84] Juan M. Cebrián, Juan L. Aragón, José M. García, and Stefanos Kaxiras. Leakage-efficient design of value predictors through state and nonstate preserving techniques. The Journal of Supercomputing, 55(1):28– 50, January 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=1&spage=28.

Wu:2011:ACT

[85] Yongwei Wu, Chen Gang, Jia Liu, Rui Fang, Xiaomeng Huang, Guangwen Yang, and Weimin Zheng. Automatically constructing trusted cluster computing environment. The Journal of Supercomputing, 55(1):51– 68, January 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=1&spage=51.

Zou:2011:BAT

[86] Deqing Zou, Shangxin Du, Weide Zheng, and Hai Jin. Building Automated Trust Negotiation architecture in virtual computing environment. The Journal of Supercomputing, 55 (1):69–85, January 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=1&spage=69.

Chang:2011:RNC

[87] Kuei-Chung Chang. Reliable networkon-chip design for multi-core systemon-chip. The Journal of Supercomputing, 55(1):86–102, January 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0920-8542&volume=55&
issue=1&spage=86.

Xiao:2011:REH

[88] Nong Xiao, Tao Chen, and Fang Liu. RSEDP: an effective hybrid data placement algorithm for largescale storage systems. The Journal of Supercomputing, 55(1):103– 122, January 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=1&spage=103.

Kim:2011:SIT

[89] Tai-Hoon Kim, Omer F. Rana, Juan Tourino, and Isaac Woungang. Special issue on “Theory and practice of high-performance computing, communications, and security”. The Journal of Supercomputing, 55(2): 123–125, February 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=2&spage=123.

Taboada:2011:DEJ

[90] Guillermo L. Taboada, Sabela Ramos, Juan Touriño, and Ramón Doallo. Design of efficient Java message-passing collectives on multi-core clusters. The Journal of Supercomputing, 55(2): 126–154, February 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=2&spage=126.

Moon:2011:ABM

[91] Jong Sik Moon, Jong Hyuk Park, and Im-Yeong Lee. Admissible bilinear map-based key management protocol for HPCCS in heterogeneous network. The Journal of Supercomputing, 55 (2):155–172, February 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=55&issue=2&spage=155.

Tso:2011:ESC

[92] Raylin Tso, Xun Yi, and Xinyi Huang. Efficient and short certificateless signatures secure against realistic adversaries. The Journal of Supercomputing, 55(2):173– 191, February 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=2&spage=173.

Lee:2011:NDV

[93] Yang Sun Lee and Sang-Soo Yeo. A novel design of variable-rate RS encoder for ubiquitous high performance multimedia service in Gbps transmission system. The Journal of Supercomputing, 55(2):192– 206, February 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=2&spage=192.

Toegl:2011:AIL

[94] Ronald Toegl and Michael Hutter. An approach to introducing locality in remote attestation using near field communications. The Journal of Supercomputing, 55(2):207– 227, February 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=2&spage=207.

Chang:2011:SEE

[95] Hangbae Chang. The study on end-to-end security for ubiquitous commerce. The Journal of Supercomputing, 55 (2):228–245, February 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=2&spage=228.

Tran:2011:PBA

[96] Minh-Triet Tran, Thanh-Trung Nguyen, Anh-Duc Duong, and Isao Echizen. Pool-based anonymous communication framework for high-performance computing. The Journal of Supercomputing, 55(2):246–268, February 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0920-8542&volume=55&
                                        issue=2&spage=246.

Lai:2011:PUB

[97] Chin-Feng Lai, Sung-Yen Chang, Yueh-Min Huang, Jong Hyuk Park, and Han-Chieh Chao. A portable UPnP-based high performance content sharing system for supporting multimedia devices. The Journal of Supercomputing, 55(2):269– 283, February 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=2&spage=269.

Smith:2011:SMC

[98] Matthew Smith, Christian Schridde, Björn Agel, and Bernd Freisleben.

Secure mobile communication via identity-based cryptography and serveraided computations. The Journal of Supercomputing, 55(2):284– 306, February 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=2&spage=284.

Jeong:2011:ERS

[99] Cheol Ho Jeong and Kwang Seon Ahn. Efficient RNTS system for privacy of banking off-line customer. The Journal of Supercomputing, 55 (2):307–319, February 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=2&spage=307.

Khan:2011:MNG

[100] Samee Ullah Khan. Mosaic-Net: a game theoretical method for selection and allocation of replicas in ad hoc networks. The Journal of Supercomputing, 55(3):321–366, March 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=55&issue=3&spage=321.

Li:2011:TBE

[101] Chunlin Li and Layuan Li. Tradeoffs between energy consumption and QoS in mobile grid. The Journal of Supercomputing, 55(3):367– 399, March 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=3&spage=367.

Nitin:2011:DFT

[102] Nitin, Shruti Garhwal, and Neha Srivastava. Designing a fault-tolerant fully-chained combining switches multistage interconnection network with disjoint paths. The Journal of Supercomputing, 55(3):400–431, March 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=3&spage=400.

Park:2011:DSR

[103] Jae-Hyun Park. The deflection selfrouting Delta network: a dynamically fault-tolerant high-radix multistage interconnection network. The Journal of Supercomputing, 55(3):432– 447, March 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=55&issue=3&spage=432.

Tan:2011:APR

[104] Guangming Tan, Vugranam C. Sreedhar, and Guang R. Gao. Analysis and performance results of computing betweenness centrality on IBM Cyclops64. The Journal of Supercomputing, 56(1):1–24, April 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=56&issue=1&spage=1.

Kim:2011:RBP

[105] Seonggun Kim, Hwansoo Han, and Kwang-Moo Choe. Region-based parallelization of irregular reductions on explicitly managed memory hierarchies. The Journal of Supercomputing, 56(1):25–55, April 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=1&spage=25.

Chakraborty:2011:CAC

[106] Abhirup Chakraborty and Ajit Singh. Cost-aware caching schemes in heterogeneous storage systems. The Journal of Supercomputing, 56(1): 56–78, April 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=1&spage=56.

Wang:2011:MSA

[107] Yuanzhuo Wang, Chuang Lin, Peter D. Ungsunan, and Xiaomeng Huang. Modeling and survivability analysis of service composition using Stochastic Petri Nets. The Journal of Supercomputing, 56(1): 79–105, April 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=1&spage=79.

Qureshi:2011:HFT

[108] Kalim Qureshi, Fiaz Gul Khan, Paul Manuel, and Babar Nazir. A hybrid fault tolerance technique in grid computing system. The Journal of Supercomputing, 56(1):106– 128, April 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=1&spage=106.

Chang:2011:FPD

[109] Weng-Long Chang, Shu-Chien Huang, Kawuu Weicheng Lin, and Michael (Shan-Hui) Ho. Fast parallel DNAbased algorithms for molecular computation: discrete logarithm. The Journal of Supercomputing, 56(2): 129–163, May 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=2&spage=129.

Yang:2011:ESM

[110] Xuejun Yang, Jing Du, and Zhiyuan Wang. An effective speedup metric for measuring productivity in large-scale parallel computer systems. The Journal of Supercomputing, 56 (2):164–181, May 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=2&spage=164.

Lu:2011:PCP

[111] You Lu and Jun-Ming Xu. Panconnectivity of Cartesian product graphs. The Journal of Supercomputing, 56 (2):182–189, May 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=2&spage=182.

Lee:2011:MNO

[112] Jinkwan Lee, Jiyoung Song, Sangjoon Park, Hyunjoo Mun, Jongchan Lee, Youngsong Mun, and Byunggi Kim. MLP network for optimal MR decision in a large-scale nesting mobile networks. The Journal of Supercomputing, 56(2):190–211, May 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=2&spage=190.

Kim:2011:OIS

[113] Jong-Seok Kim, Hyeong-Ok Lee, Eddie Cheng, and László Lipták. Optimal independent spanning trees on odd graphs. The Journal of Supercomputing, 56(2):212–225, May 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=2&spage=212.

Bulic:2011:AMF

[114] Patricio Bulić and Tomaz Dobravec. An approximate method for filtering out data dependencies with a sufficiently large distance between memory references. The Journal of Supercomputing, 56(2):226– 244, May 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=2&spage=226.

Zhang:2011:IJS

[115] Junwei Zhang, Bu-Sung Lee, Xueyan Tang, and Chai-Kiat Yeo. Improving job scheduling performance with parallel access to replicas in Data Grid environment. The Journal of Supercomputing, 56(3):245– 269, June 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=56&issue=3&spage=245.

Yang:2011:CWB

[116] Chao-Tung Yang, Kuan-Chou Lai, and Hao-Yu Tung. On construction of a well-balanced allocation strategy for heterogeneous multi-cluster computing environments. The Journal of Supercomputing, 56(3):270– 299, June 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=56&issue=3&spage=270.

Linford:2011:SHP

[117] John C. Linford and Adrian Sandu. Scalable heterogeneous parallelism for atmospheric modeling and simulation. The Journal of Supercomputing, 56 (3):300–327, June 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=56&issue=3&spage=300.

Bargi:2011:TMT

[118] Ava Bargi and Hamid Sarbazi-Azad. Task migration in threedimensional meshes. The Journal of Supercomputing, 56(3):328– 352, June 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=3&spage=328.

Chen:2011:DMH

[119] Min Chen, Meikang Qiu, Linxia Liao, Jongan Park, and Jianhua Ma. Distributed multi-hop cooperative communication in dense wireless sensor networks. The Journal of Supercomputing, 56(3):353– 369, June 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=56&issue=3&spage=353.

Thor:2011:P

[120] Miroslaw Thor and Marek Tudruj. Preface. The Journal of Supercomputing, 57(1):1–4, July 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=1&spage=1.

Forsell:2011:MTP

[121] M. Forsell and V. Leppänen. A moving threads processor architecture MTPA. The Journal of Supercomputing, 57(1):5–19, July 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=1&spage=5.

Peng:2011:EHP

[122] Liu Peng, Manaschai Kunaseth, Hikmet Dursun, Ken ichi Nomura, Weiqiang Wang, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. Exploiting hierarchical parallelisms for molecular dynamics simulation on multicore clusters. The Journal of Supercomputing, 57(1):20– 33, July 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=1&spage=20.

Al-Qawasmeh:2011:SMQ

[123] Abdulla M. Al-Qawasmeh, Anthony A. Maciejewski, Haonan Wang, Jay Smith, Howard Jay Siegel, and Jerry Potter. Statistical measures for quantifying task and machine heterogeneities. The Journal of Supercomputing, 57 (1):34–50, July 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=1&spage=34.

Tudruj:2011:MCS

[124] Marek Tudruj, Lukasz Masko, and Miroslaw Thor. Multi-CMP system with data communication on the fly. The Journal of Supercomputing, 57(1):51–64, July 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=1&spage=51.

Aziz:2011:PES

[125] Abdul Aziz and Hesham El-Rewini. Power efficient scheduling heuristics for energy conservation in computational grids. The Journal of Supercomputing, 57(1):65–80, July 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=1&spage=65.

Pasquali:2011:MLS

[126] Marco Pasquali, Ranieri Baraglia, Gabriele Capannini, Laura Ricci, and Domenico Laforenza. A multi-level scheduler for batch jobs on grids. The Journal of Supercomputing, 57 (1):81–98, July 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=1&spage=81.

Cascon:2011:IIN

[127] Pablo Cascón, Julio Ortega, Yan Luo, Eric Murray, Antonio Díaz, and Ignacio Rojas. Improving IPS by network processors. The Journal of Supercomputing, 57(1):99–108, July 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=1&spage=99.

Bandini:2011:CBS

[128] Stefania Bandini, Andrea Bonomi, Giuseppe Vizzari, and Vito Acconci. CA-based self-organizing environments the case of a configurable adaptive illumination facility. The Journal of Supercomputing, 57(2):109– 120, August 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=2&spage=109.

Bandman:2011:UCA

[129] Olga Bandman. Using cellular automata for porous media simulation. The Journal of Supercomputing, 57 (2):121–131, August 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=2&spage=121.

Trinitis:2011:SMO

[130] Carsten Trinitis, Tilman Küstner, Josef Weidendorfer, and Jasmin Smajic. Sparse matrix operations on several multi-core architectures. The Journal of Supercomputing, 57(2):132– 140, August 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=2&spage=132.

Walters:2011:VCP

[131] Jesse D. Walters, Thomas B. Bair, Terry A. Braun, Todd E. Scheetz, John P. Robinson, and Thomas L. Casavant. Validation of computational prediction of horizontal gene transfer events — XenoCluster. The Journal of Supercomputing, 57(2):141– 150, August 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=2&spage=141.

Schellmann:2011:PMI

[132] Maraike Schellmann, Sergei Gorlatch, Dominik Meiländer, Thomas Kösters, Klaus Schäfers, Frank Wübbeling, and Martin Burger. Parallel medical image reconstruction: from graphics processing units (GPU) to Grids. The Journal of Supercomputing, 57 (2):151–160, August 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=2&spage=151.

Kireev:2011:FNA

[133] S. Kireev and V. Malyshkin. Fragmentation of numerical algorithms for parallel subroutines library. The Journal of Supercomputing, 57(2):161– 171, August 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=2&spage=161.

Kozlov:2011:DDE

[134] Konstantin Kozlov and Alexander Samsonov. DEEP — differential evolution entirely parallel method for gene regulatory networks. The Journal of Supercomputing, 57(2):172– 178, August 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=2&spage=172.

Biardzki:2011:DMB

[135] Christoph Biardzki and Thomas Ludwig. DMetabench — a metadata benchmark for distributed file systems. The Journal of Supercomputing, 57 (2):179–188, August 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=2&spage=179.

Xia:2011:PEP

[136] Yinglong Xia and Viktor K. Prasanna. Parallel evidence propagation on multicore processors. The Journal of Supercomputing, 57(2):189– 202, August 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=2&spage=189.

Imbs:2011:STM

[137] Damien Imbs and Michel Raynal. Software transactional memories: an approach for multicore programming. The Journal of Supercomputing, 57 (2):203–215, August 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=2&spage=203.

Szaban:2011:IQB

[138] Miroslaw Szaban and Franciszek Seredynski. Improving quality of DES Sboxes by cellular automata-based Sboxes. The Journal of Supercomputing, 57(2):216–226, August 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=2&spage=216.

Arora:2011:TNI

[139] Ritu Arora, Purushotham Bangalore, and Marjan Mernik. A technique for non-invasive applicationlevel checkpointing. The Journal of Supercomputing, 57(3):227– 255, September 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=3&spage=227.

Lai:2011:CRM

[140] Guoming Lai and Xiaola Lin. Condition for relaxed Monte Carlo method of solving systems of linear equations. The Journal of Supercomputing, 57(3): 256–264, September 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=57&issue=3&spage=256.

Qureshi:2011:EGA

[141] Kalim Qureshi, Attiqa Rehman, and Paul Manuel. Enhanced GridSim architecture with load balancing. The Journal of Supercomputing, 57(3): 265–275, September 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=3&spage=265.

Rakesh:2011:AMS

[142] Nitin Rakesh and Nitin. Analysis of multi-sort algorithm on multi-Mesh of Trees (MMT) architecture. The Journal of Supercomputing, 57(3): 276–313, September 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=3&spage=276.

Abderazek:2011:NIL

[143] Ben Abdallah Abderazek, Masashi Masuda, Arquimedes Canedo, and Kenichi Kuroda. Natural instruction level parallelism-aware compiler for high-performance QueueCore processor architecture. The Journal of Supercomputing, 57(3):314– 338, September 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=3&spage=314.

Banicescu:2011:PSH

[144] Ioana Banicescu, Hyeona Lim, Ricolindo L. Cariño, and Seongjai Kim. A parameter study of a hybrid Laplacian mean-curvature flow denoising model. The Journal of Supercomputing, 57(3): 339–356, September 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=57&issue=3&spage=339.

He:2011:F

[145] Matthew He and Zhengbing Hu. Foreword. The Journal of Supercomputing, 58(1):1–3, October 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=1.

Zhang:2011:MBO

[146] Jungen Zhang, Hongbing Ji, and Cheng Ouyang. Multitarget bearings-only tracking using fuzzy clustering technique and Gaussian particle filter. The Journal of Supercomputing, 58 (1):4–19, October 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=4.

Huang:2011:RIE

[147] Chaobing Huang, Quan Liu, and Shengsheng Yu. Regions of interest extraction from color image based on visual saliency. The Journal of Supercomputing, 58(1):20– 33, October 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=20.

Qu:2011:NNC

[148] Xiaomei Qu, Jie Zhou, and Yingting Luo. A new noise-compensated estimation scheme for multichannel autoregressive signals from noisy observations. The Journal of Supercomputing, 58(1):34–49, October 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=34.

Lv:2011:IIB

[149] Yingda Lv, Xuanjing Shen, and Haipeng Chen. An improved image blind identification based on inconsistency in light source direction. The Journal of Supercomputing, 58 (1):50–67, October 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=50.

Ding:2011:RME

[150] Yong Ding and Xiaolang Yan. A robust motion estimation with centerbiased diamond search and its parallel architecture for motioncompensated de-interlace. The Journal of Supercomputing, 58(1):68– 83, October 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=68.

Tang:2011:FRB

[151] Hengliang Tang, Yanfeng Sun, Baocai Yin, and Yun Ge. 3D face recognition based on sparse representation. The Journal of Supercomputing, 58 (1):84–95, October 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=84.

Zhu:2011:SEC

[152] Linlin Zhu, Yan Zhao, Shigang Wang, and Hexin Chen. Spatial error concealment for stereoscopic video coding based on pixel matching. The Journal of Supercomputing, 58(1):96– 105, October 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=96.

Chen:2011:RTM

[153] Hui Chen, Chongzhao Han, and Yuchun Zhang. Research on tracking of maneuvering multi-target based on bionics for IRST system. The Journal of Supercomputing, 58(1): 106–121, October 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=106.

Zuo:2011:CMW

[154] Ke Zuo, Huaimin Wang, Quanyuan Wu, and Dongmin Hu. Connectivity model of wireless networks via dependency links random graphs. The Journal of Supercomputing, 58(1): 122–141, October 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=1&spage=122.

Ranilla:2011:HPC

[155] J. Ranilla, E. S. Quintana, and J. Vigo-Aguiar. High performance computing tools in science and engineering. The Journal of Supercomputing, 58(2): 143–144, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=2&spage=143.

Marques:2011:UDC

[156] M. Marqués, G. Quintana-Ortí, E. S. Quintana-Ortí, and R. van de Geijn. Using desktop computers to solve largescale dense linear algebra problems. The Journal of Supercomputing, 58(2): 145–150, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=2&spage=145.

Martinez:2011:ATI

[157] J. A. Martínez, E. M. Garzón, A. Plaza, and I. García. Automatic tuning of iterative computation on heterogeneous multiprocessors with ADITHE. The Journal of Supercomputing, 58(2): 151–159, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=2&spage=151.

Galiano:2011:PNP

[158] Vicente Galiano, Héctor Migallón, Violeta Migallón, and José Penadés. Parallel nonlinear preconditioners on multicore architectures. The Journal of Supercomputing, 58(2):160– 167, November 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=160.

Santos:2011:WSB

[159] A. Santos, F. Almeida, V. Blanco, and J. C. Castillo. Web services based scheduling in OpenCF. The Journal of Supercomputing, 58(2): 168–176, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=168.

Sanjurjo:2011:OMC

[160] J. R. Sanjurjo, M. Amor, M. Bóo, R. Doallo, and J. Casares. Optimizing Monte Carlo radiosity on graphics hardware. The Journal of Supercomputing, 58(2):177– 185, November 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=177.

Orobitg:2011:EPP

[161] Miquel Orobitg, Fernando Guirado, Cedric Notredame, and Fernando Cores. Exploiting parallelism on progressive alignment methods. The Journal of Supercomputing, 58(2): 186–194, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=2&spage=186.

Pichel:2011:AES

[162] Juan C. Pichel, Juan A. Lorenzo, Dora B. Heras, Jose C. Cabaleiro, and Tomás F. Pena. Analyzing the execution of sparse matrix-vector product on the Finisterrae SMP-NUMA system. The Journal of Supercomputing, 58(2): 195–205, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=2&spage=195.

delaAsuncion:2011:SOL

[163] Marc de la Asunción, José M. Mantas, and Manuel J. Castro. Simulation of one-layer shallow water systems on multicore and CUDA architectures. The Journal of Supercomputing, 58(2): 206–214, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=2&spage=206.

Alonso:2011:NEM

[164] P. Alonso, R. Cortina, F. J. Martínez-Zaldívar, and J. Ranilla. Neville elimination on multi- and many-core systems: OpenMP, MPI and CUDA. The Journal of Supercomputing, 58(2): 215–225, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=2&spage=215.

Gonzalez-Escribano:2011:TNP

[165] Arturo González-Escribano and Diego R. Llanos. Trasgo: a nested-parallel programming system. The Journal of Supercomputing, 58(2):226– 234, November 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=226.

Martinez-Zaldivar:2011:HPS

[166] F. J. Martínez-Zaldívar, A. M. Vidal-Maciá, and D. Giménez. Heterogeneous pipelined square-root Kalman Filter algorithm for the MMSE–OSIC problem. The Journal of Supercomputing, 58(2): 235–243, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=235.

Pedraza:2011:GAB

[167] César Pedraza, Javier Castillo, José I. Martínez, Pablo Huerta, Jose L. Bosque, and Javier Cano. Genetic algorithm for Boolean minimization in an FPGA cluster. The Journal of Supercomputing, 58(2):244– 252, November 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=244.

Calvo:2011:CPM

[168] J. C. Calvo, J. Ortega, and M. Anguita. Comparison of parallel multi-objective approaches to protein structure prediction. The Journal of Supercomputing, 58(2):253– 260, November 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=253.

Vigueras:2011:WBD

[169] Guillermo Vigueras, Miguel Lozano, and Juan M. Orduña. Workload balancing in distributed crowd simulations: the partitioning method. The Journal of Supercomputing, 58(2): 261–269, November 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=261.

Redondo:2011:PEA

[170] J. L. Redondo, I. García, and P. M. Ortigosa. Parallel evolutionary algorithms based on shared memory programming approaches. The Journal of Supercomputing, 58(2):270– 279, November 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=2&spage=270.

Quintana-Orti:2011:HPC

[171] Enrique Quintana-Ortí, Jose Ranilla, and Jesús Vigo-Aguiar. High performance computing tools in science and engineering II. The Journal of Supercomputing, 58(3):281– 282, December 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=281.

Lopez-Portugues:2011:GSF

[172] Miguel López-Portugués, Jesús A. López-Fernández, Alberto Rodríguez-Campa, and José Ranilla. A GPGPU solution of the FMM near interactions for acoustic scattering problems. The Journal of Supercomputing, 58(3): 283–291, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=283.

Almeida:2011:PSM

[173] Francisco Almeida, Domingo Giménez, and Jose J. López-Espín. A parameterized shared-memory scheme for parameterized metaheuristics. The Journal of Supercomputing, 58(3): 292–301, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=292.

Cascon:2011:ANA

[174] Pablo Cascón, Andrés Ortiz, Julio Ortega, Antonio F. Díaz, and Ignacio Rojas. Accelerating network applications by distributed interfaces on heterogeneous multiprocessor architectures. The Journal of Supercomputing, 58(3): 302–313, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=3&spage=302.

Martinez-Zaldivar:2011:TBM

[175] F. J. Martínez-Zaldívar, A. M. Vidal-Maciá, Alberto Gonzalez, and Vicenç Almenar. Tridimensional block multiword LDPC decoding on GPUs. The Journal of Supercomputing, 58(3): 314–322, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=3&spage=314.

Prada:2011:PSA

[176] Laura Prada, Javier Garcia, J. Daniel Garcia, and Jesus Carretero. Power saving-aware prefetching for SSDbased systems. The Journal of Supercomputing, 58(3):323–331, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0920-8542&volume=58&
                                        issue=3&spage=323.

Martinez:2011:UAA

[177] Diego R. Martínez, Julio L. Albín, Tomás F. Pena, José C. Cabaleiro, Francisco F. Rivera, and Vicente Blanco. Using accurate AIC-based performance models to improve the scheduling of parallel applications. The Journal of Supercomputing, 58(3): 332–340, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=332.

Barri:2011:MMH

[178] I. Barri, C. Roig, F. Giné, and F. Solsona. Mapping MMOFPS over heterogeneous distributed systems. The Journal of Supercomputing, 58(3): 341–348, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=341.

Reyes:2011:ACG

[179] Ruyman Reyes and Francisco de Sande. Automatic code generation for GPUs in llc. The Journal of Supercomputing, 58 (3):349–356, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=349.

Padron:2011:PHR

[180] Emilio J. Padrón, Margarita Amor, Montserrat Bóo, Gabriel Rodríguez, and Ramón Doallo. Parallel hierarchical radiosity on hybrid platforms. The Journal of Supercomputing, 58(3): 357–366, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=357.

Bosque:2011:ESH

[181] Jose L. Bosque, Oscar D. Robles, Pablo Toharia, and Luis Pastor. Evaluating scalability in heterogeneous systems.

The Journal of Supercomputing, 58(3): 367–375, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=367.

Sanjuan-Estrada:2011:API

[182] J. F. Sanjuan-Estrada, L. G. Casado, and I. García. Adaptive parallel interval branch and bound algorithms based on their performance for multicore architectures. The Journal of Supercomputing, 58(3):376– 384, December 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=376.

Martinez:2011:ALB

[183] J. A. Martínez, F. Almeida, E. M. Garzón, A. Acosta, and V. Blanco. Adaptive load balancing of iterative computation on heterogeneous nondedicated systems. The Journal of Supercomputing, 58(3):385– 393, December 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=385.

Blanco:2011:MJC

[184] Héctor Blanco, Josep Lluís Lérida, Fernando Cores, and Fernando Guirado. Multiple job co-allocation strategy for heterogeneous multi-cluster systems based on linear programming. The Journal of Supercomputing, 58(3):

394–402, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=394.

Valero:2011:GBI

[185] Pedro Valero, José L. Sánchez, Diego Cazorla, and Enrique Arias. A GPUbased implementation of the MRF algorithm in ITK package. The Journal of Supercomputing, 58(3): 403–410, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=403.

Molero:2011:FAD

[186] J. M. Molero, A. Paz, E. M. Garzón, J. A. Martínez, A. Plaza, and I. García. Fast anomaly detection in hyperspectral images with RX method on heterogeneous clusters. The Journal of Supercomputing, 58(3):411– 419, December 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=58&issue=3&spage=411.

Redondo:2011:SFL

[187] J. L. Redondo, J. Fernández, I. García, and P. M. Ortigosa. Solving the facility location and design (1|1)-centroid problem via parallel algorithms. The Journal of Supercomputing, 58(3): 420–428, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=3&spage=420.

Ezzatti:2011:UGP

[188] P. Ezzatti, E. S. Quintana-Ortí, and A. Remón. Using graphics processors to accelerate the computation of the matrix inverse. The Journal of Supercomputing, 58(3):429– 437, December 2011. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=3&spage=429.

Migallon:2011:PPL

[189] Héctor Migallón, Violeta Migallón, and José Penadés. A Parallel Python library for nonlinear systems. The Journal of Supercomputing, 58(3): 438–448, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=3&spage=438.

Belloch:2011:RTM

[190] Jose A. Belloch, Alberto Gonzalez, F. J. Martínez-Zaldívar, and Antonio M. Vidal. Real-time massive convolution for audio applications on GPU massive convolution on GPU. The Journal of Supercomputing, 58(3): 449–457, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=58&issue=3&spage=449.

Sabbaghi-Nadooshan:2012:DBN

[191] Reza Sabbaghi-Nadooshan, Mehdi Modarressi, and Hamid Sarbazi-Azad. The 2D digraph-based NoCs: attractive alternatives to the 2D mesh NoCs. The Journal of Supercomputing, 59 (1):1–21, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=1.

Majzoub:2012:MRH

[192] Sohaib Majzoub and Hassan Diab. MorphoSys reconfigurable hardware for cryptography: the Twofish case. The Journal of Supercomputing, 59 (1):22–41, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007%2Fs11227-010-0413-3.

Wu:2012:DPL

[193] Chao-Chin Wu, Chao-Tung Yang, Kuan-Chou Lai, and Po-Hsun Chiu. Designing parallel loop self-scheduling schemes using the hybrid MPI and OpenMP programming model for multi-core grid systems. The Journal of Supercomputing, 59(1):42– 60, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=42.

Mavromoustakis:2012:TBA

[194] Constandinos X. Mavromoustakis and Helen D. Karatza. A tier-based asynchronous scheduling scheme for delay constrained energy efficient connectivity in asymmetrical wireless devices. The Journal of Supercomputing, 59(1):61–82, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=61.

Zhu:2012:PSA

[195] Mengxia Zhu, Guangxing Wang, and Tonny Oyana. Parallel spatiotemporal autocorrelation and visualization system for large-scale remotely sensed images. The Journal of Supercomputing, 59(1):83–103, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=83.

Falzon:2012:ELS

[196] Geoffrey Falzon and Maozhen Li. Enhancing list scheduling heuristics for dependent job scheduling in grid computing environments. The Journal of Supercomputing, 59(1):104– 130, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=104.

Sehrish:2012:RFS

[197] Saba Sehrish and Jun Wang. Reduced Function Set Abstraction (RFSA) for MPI-IO. The Journal of Supercomputing, 59(1):131– 146, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=131.

Nunez:2012:CSF

[198] Juan Núñez and Ángel F. Tenorio. A computational study of a family of nilpotent Lie algebras. The Journal of Supercomputing, 59(1): 147–155, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=147.

Arrieta-Salinas:2012:IRF

[199] Itziar Arrieta-Salinas, José Enrique Armendáriz-Iñigo, José Ramón Juárez-Rodríguez, and José Ramón González de Mendívil. An implementation of a replicated file server supporting the crash-recovery failure model. The Journal of Supercomputing, 59 (1):156–202, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=156.

Letaifa:2012:RPR

[200] Asma Ben Letaifa and Sami Tabbane. Reconfiguration process and routing management for service platforms. The Journal of Supercomputing, 59(1):203–226, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=203.

Ba:2012:PSS

[201] Wei Ba, Dabo Zhang, Qi Li, and Wei Wang. The partitioned scheduling of sporadic task systems on multiprocessors. The Journal of Supercomputing, 59(1):227–245, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=1&spage=227.

Cheng:2012:NAG

[202] Eddie Cheng, Ke Qiu, and Zhizhang Shen. A note on the alternating group network. The Journal of Supercomputing, 59(1):246– 248, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=1&spage=246.

Hababeh:2012:INS

[203] Ismail Hababeh. Improving network systems performance by clustering distributed database sites. The Journal of Supercomputing, 59(1): 249–267, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=1&spage=249.

Jiang:2012:DPN

[204] Fuu-Cheng Jiang, Chu-Hsing Lin, Der-Chen Huang, and Chao-Tung Yang. Dual paths node-disjoint routing for data salvation in mobile ad hoc. The Journal of Supercomputing, 59 (1):268–296, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=268.

Chan:2012:MPB

[205] Edward Chan, Yilin Wang, Wenzhong Li, and Sanglu Lu. Movement prediction based cooperative caching for location dependent information service in mobile ad hoc networks. The Journal of Supercomputing, 59 (1):297–322, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=297.

Lindberg:2012:CAE

[206] Peder Lindberg, James Leingang, Daniel Lysaker, Samee Ullah Khan, and Juan Li. Comparison and analysis of eight scheduling heuristics for the optimization of energy consumption and makespan in largescale distributed systems. The Journal of Supercomputing, 59(1):323– 360, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=323.

Filgueira:2012:DCD

[207] Rosa Filgueira, Jesús Carretero, David E. Singh, Alejandro Calderón, and Alberto Núñez. Dynamic–CoMPI: dynamic optimization techniques for MPI parallel applications. The Journal of Supercomputing, 59(1):361– 391, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=1&spage=361.

Muszala:2012:NLI

[208] Stefan P. Muszala, Gita Alaghband, James Hack, and Daniel Connors. Natural Load Indices (NLI) for scientific simulation. The Journal of Supercomputing, 59(1):392– 413, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=1&spage=392.

Yang:2012:PBD

[209] Chao-Tung Yang, Wen-Chung Shih, and Lung-Hsing Cheng. Performancebased dynamic loop scheduling in heterogeneous computing environments. The Journal of Supercomputing, 59 (1):414–442, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=1&spage=414.

Jang:2012:LON

[210] Jae-Wan Jang, Euiseong Seo, Heeseung Jo, and Jin-Soo Kim. A low-overhead networking mechanism for virtualized high-performance computing systems. The Journal of Supercomputing, 59 (1):443–468, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=443.

Parsa:2012:TDA

[211] Saeed Parsa and Reza Entezari-Maleki. Task dispatching approach to reduce the number of waiting tasks in grid environments. The Journal of Supercomputing, 59(1):469– 485, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=469.

Aldea:2012:USC

[212] Sergio Aldea, Diego R. Llanos, and Arturo González-Escribano. Using SPEC CPU2006 to evaluate the sequential and parallel code generated by commercial and open-source compilers. The Journal of Supercomputing, 59(1):486–498, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=486.

Li:2012:RSS

[213] Chunlin Li and Layuan Li. A resource selection scheme for QoS satisfaction and load balancing in ad hoc grid. The Journal of Supercomputing, 59(1):499–525, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=499.

Al-Dayaa:2012:RLT

[214] H. S. Al-Dayaa and D. B. Megherbi. Reinforcement learning technique using agent state occurrence frequency with analysis of knowledge sharing on the agent’s learning process in multiagent environments. The Journal of Supercomputing, 59(1):526– 547, January 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=526.

Sharifi:2012:PID

[215] Mohsen Sharifi, Ehsan Mousavi Khaneghah, Morteza Kashyian, and Seyedeh Leili Mirtaheri. A platform independent distributed IPC mechanism in support of programming heterogeneous distributed systems. The Journal of Supercomputing, 59(1): 548–567, January 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=1&spage=548.

Nitin:2012:CAT

[216] Nitin and Durg Singh Chauhan. Comparative analysis of traffic patterns on k-ary n-tree using adaptive algorithms based on Burton Normal Form. The Journal of Supercomputing, 59 (2):569–588, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=569.

Liu:2012:NLM

[217] Jun Liu. A novel load metric with enhanced ability of distinguishing different load status. The Journal of Supercomputing, 59(2):589– 609, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=589.

Wang:2012:AOS

[218] Fei Wang and Jack Jean. Architecture and operating system support for two-dimensional runtime partial reconfiguration. The Journal of Supercomputing, 59(2):610– 635, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=610.

Fazlali:2012:EDM

[219] Mahmood Fazlali, Ali Zakerolhosseini, and Georgi Gaydadjiev. Efficient datapath merging for the overhead reduction of run-time reconfigurable systems. The Journal of Supercomputing, 59(2):636– 657, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=636.

Ryu:2012:OFH

[220] Seonggeun Ryu, Kyunghye Lee, and Youngsong Mun. Optimized fast handover scheme in Mobile IPv6 networks to support mobile users for cloud computing. The Journal of Supercomputing, 59(2):658– 675, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=658.

Kim:2012:ESV

[221] Soon Seok Kim, Deok Gyu Lee, and Jong Hyuk Park. Efficient scheme of verifying integrity of application binaries in embedded operating systems. The Journal of Supercomputing, 59 (2):676–692, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=676.

Thibault:2012:AIF

[222] Julien C. Thibault and Inanc Senocak. Accelerating incompressible flow computations with a Pthreads–CUDA implementation on small-footprint multi-GPU platforms. The Journal of Supercomputing, 59(2):693– 719, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=693.

Dashtbozorgi:2012:HPS

[223] Mahdi Dashtbozorgi and Mohammad Abdollahi Azgomi. A highperformance and scalable multicore aware software solution for network monitoring. The Journal of Supercomputing, 59(2):720– 743, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=720.

Chizari:2012:EMN

[224] Hassan Chizari, Majid Hosseini, Shaharuddin Salleh, Shukor Abd Razak, and Abdul Hanan Abdullah. EF-MPR, a new energy efficient multi-point relay selection algorithm for MANET. The Journal of Supercomputing, 59 (2):744–761, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=744.

Beg:2012:PNC

[225] Azam Beg. On pedagogy of nanometric circuit reliability. The Journal of Supercomputing, 59(2):762– 778, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=762.

Nitin:2012:SCA

[226] Nitin and Durg Singh Chauhan. Stochastic communication for applicationspecific Networks-on-Chip. The Journal of Supercomputing, 59(2):779– 810, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=2&spage=779.

Gavrilova:2012:DCN

[227] Marina Gavrilova and Kushan Ahmadian. On-demand chaotic neural network for broadcast scheduling problem. The Journal of Supercomputing, 59 (2):811–829, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=2&spage=811.

Alachiotis:2012:DLM

[228] Nicolaos Alachiotis, Vasileios I. Kelefouras, George S. Athanasiou, Harris E. Michail, Angeliki S. Kritikakou, and Costas E. Goutis. A data locality methodology for matrix-matrix multiplication algorithm. The Journal of Supercomputing, 59(2):830– 851, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=2&spage=830.

Luo:2012:PMM

[229] Congnan Luo and Soon M. Chung. Parallel mining of maximal sequential patterns using multiple samples. The Journal of Supercomputing, 59 (2):852–881, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=2&spage=852.

Zhou:2012:TTR

[230] Jiazheng Zhou and Yeh-Ching Chung. Tree-turn routing: an efficient deadlockfree routing algorithm for irregular networks. The Journal of Supercomputing, 59(2):882–900, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=0920-8542&volume=59&
issue=2&spage=882.

Zhou:2012:HSM

[231] Jiazheng Zhou, Shen-En Liu, and Yeh-Ching Chung. A hardware supported multicast scheme based on XY routing for 2-D mesh InfiniBand networks. The Journal of Supercomputing, 59 (2):901–919, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=901.

Jana:2012:OME

[232] Prasanta K. Jana and Dheeresh K. Mallick. OTIS-MOT: an efficient interconnection network for parallel processing. The Journal of Supercomputing, 59 (2):920–940, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=920.

Safaei:2012:AMC

[233] F. Safaei and A. Khonsari. Analytical modeling and comparison of faulttolerant message flow control mechanisms in torus-connected networks.

The Journal of Supercomputing, 59 (2):941–974, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=941.

Moschakis:2012:EGS

[234] Ioannis A. Moschakis and Helen D. Karatza. Evaluation of gang scheduling performance and cost in a cloud computing system. The Journal of Supercomputing, 59(2):975– 992, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=975.

Liu:2012:QNS

[235] Dan Liu and Yuan-Da Cao. Queuing network of scale free topology: on modelling large scale network. The Journal of Supercomputing, 59(2): 993–1018, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=993.

Kim:2012:TCI

[236] Jaehyun Kim, Hyunseung Choo, and Dongsoo S. Kim. TCP CAE: an improved congestion control using comparative ACK-based estimator. The Journal of Supercomputing, 59(2): 1019–1034, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=1019.

Torkestani:2012:LAB

[237] Javad Akbari Torkestani and Mohammad Reza Meybodi. A learning automata-based heuristic algorithm for solving the minimum spanning tree problem in stochastic graphs. The Journal of Supercomputing, 59(2): 1035–1054, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=1035.

Lee:2012:ECL

[238] Kyunghye Lee, Seonggeun Ryu, and Youngsong Mun. An enhanced crosslayer fast handover scheme for mobile IPv6 in the IEEE 802.16e networks. The Journal of Supercomputing, 59(2): 1055–1070, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=1055.

Kanal:2012:PAI

[239] M. E. Kanal. Parallel algorithm on inversion for adjacent pentadiagonal matrices with MPI. The Journal of Supercomputing, 59(2):1071– 1078, February 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=2&spage=1071.

Arora:2012:RLA

[240] Ritu Arora, Purushotham Bangalore, and Marjan Mernik. Raising the level of abstraction for developing message passing applications. The Journal of Supercomputing, 59(2): 1079–1100, February 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=2&spage=1079.

Tang:2012:AME

[241] Jie Tang, Shaoshan Liu, Zhimin Gu, Xiao-Feng Li, and Jean-Luc Gaudiot. Achieving middleware execution efficiency: hardware-assisted garbage collection operations. The Journal of Supercomputing, 59(3): 1101–1119, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1101.

Min-Allah:2012:OTE

[242] Nasro Min-Allah, Samee Ullah Khan, and Wang Yongji. Optimal task execution times for periodic tasks using nonlinear constrained optimization. The Journal of Supercomputing, 59 (3):1120–1138, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1120.

Erdil:2012:DGL

[243] D. Cenk Erdil and Michael J. Lewis. Dynamic grid load sharing with adaptive dissemination protocols. The Journal of Supercomputing, 59(3): 1139–1166, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1139.

Li:2012:APR

[244] Yan Li, Sanjay Ranka, and Sartaj Sahni. In-advance path reservation for file transfers in e-science applications. The Journal of Supercomputing, 59(3):1167–1187, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1167.

Megherbi:2012:HCR

[245] Dalila B. Megherbi and Vikram Malayia. A hybrid cognitive/reactive intelligent agent autonomous path planning technique in a networkeddistributed unstructured environment for reinforcement learning. The Journal of Supercomputing, 59(3):1188– 1217, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1188.

Zhou:2012:UMP

[246] Min Zhou, Onkar Sahni, Ting Xie, Mark S. Shephard, and Kenneth E. Jansen. Unstructured mesh partition improvement for implicit finite element at extreme scale. The Journal of Supercomputing, 59(3):1218– 1228, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1218.

Wang:2012:OMS

[247] Li Wang, Jingling Xue, and Xuejun Yang. Optimizing modulo scheduling to achieve reuse and concurrency for stream processors. The Journal of Supercomputing, 59(3):1229– 1251, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1229.

Nitin:2012:DPA

[248] Nitin, Rajan Vaish, and Utkarsh Shrivastava. On a deadlock and performance analysis of ALBR and DAR algorithm on X-torus topology by optimal utilization of cross links and minimal lookups. The Journal of Supercomputing, 59(3):1252– 1288, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1252.

Choi:2012:DHR

[249] Sung Chune Choi and Hee Yong Youn. Dynamic hybrid replication effectively combining tree and grid topology. The Journal of Supercomputing, 59 (3):1289–1311, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1289.

Jiang:2012:LEW

[250] Fuu-Cheng Jiang, Der-Chen Huang, Chao-Tung Yang, and Fang-Yi Leu. Lifetime elongation for wireless sensor network using queue-based approaches. The Journal of Supercomputing, 59 (3):1312–1335, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1312.

Kim:2012:TPF

[251] Jong-Seok Kim, Sung Won Kim, Eddie Cheng, and László Lipták. Topological properties of folded hyper-star networks. The Journal of Supercomputing, 59(3):1336–1347, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1336.

Qureshi:2012:TPS

[252] Kalim Qureshi, Babar Majeed, Jawad Haider Kazmi, and Sajjad Ahmed Madani. Task partitioning, scheduling and load balancing strategy for mixed nature of tasks. The Journal of Supercomputing, 59(3):1348–1359, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1348.

Nimmagadda:2012:CSM

[253] Venkata Krishna Nimmagadda, Ali Akoglu, Salim Hariri, and Talal Moukabary. Cardiac simulation on multi-GPU platform. The Journal of Supercomputing, 59(3):1360– 1378, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1360.

Wu:2012:EFP

[254] Chao-Chin Wu, Lien-Fu Lai, and Yu-Shuo Chang. Extending Fuzzy-CLIPS for parallelizing data-dependent fuzzy expert systems. The Journal of Supercomputing, 59(3):1379– 1395, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1379.

Babamir:2012:CFR

[255] Seyed Morteza Babamir. Constructing formal rules to verify message communication in distributed systems. The Journal of Supercomputing, 59 (3):1396–1418, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1396.

Min-Allah:2012:CSR

[256] Nasro Min-Allah, Samee Ullah Khan, Nasir Ghani, Juan Li, Lizhe Wang, et al. A comparative study of rate monotonic schedulability tests. The Journal of Supercomputing, 59(3): 1419–1430, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1419.

Afgan:2012:SPJ

[257] Enis Afgan, Purushotham Bangalore, and Tibor Skala. Scheduling and planning job execution of loosely coupled applications. The Journal of Supercomputing, 59(3): 1431–1454, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1431.

Shahbahrami:2012:PIG

[258] Asadollah Shahbahrami, Tuan Anh Pham, and Koen Bertels. Parallel implementation of Gray Level Cooccurrence Matrices and Haralick texture features on Cell architecture. The Journal of Supercomputing, 59 (3):1455–1477, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1455.

Mahfoudhi:2012:CSR

[259] Adel Mahfoudhi, Yessine Hadj Kacem, Walid Karamti, and Mohamed Abid. Compositional specification of real time embedded systems by priority time Petri nets. The Journal of Supercomputing, 59(3):1478– 1503, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1478.

Couturier:2012:SSS

[260] Raphaël Couturier and Stéphane Domas. Sparse systems solving on GPUs with GMRES. The Journal of Supercomputing, 59(3):1504– 1516, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1504.

Jezequel:2012:SLS

[261] Fabienne Jezequel, Raphaël Couturier, and Christophe Denis. Solving large sparse linear systems in a grid environment: the GREMLINS code versus the PETSc library. The Journal of Supercomputing, 59(3):1517– 1532, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=59&issue=3&spage=1517.

Serrano:2012:CEH

[262] Mónica Serrano, Julio Sahuquillo, Salvador Petit, Houcine Hassan, and José Duato. A cost-effective heuristic to schedule local and remote memory in cluster computers. The Journal of Supercomputing, 59(3):1533– 1551, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1533.

Nam:2012:ADC

[263] Beomseok Nam and Alan Sussman. Analyzing design choices for distributed multidimensional indexing. The Journal of Supercomputing, 59 (3):1552–1576, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1552.

Qin:2012:THP

[264] Zheng Qin, Fei Chen, Qiang Wang, Alex X. Liu, and Zhiguang Qin. Towards high performance security policy evaluation. The Journal of Supercomputing, 59(3):1577– 1595, March 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1577.

Carrion:2012:DMA

[265] I. Marín Carrión, E. Arias Antúnez, M. M. Artigao Castillo, and J. J. Miralles Canals. A distributed memory architecture implementation of the False Nearest Neighbors method based on distribution of dimensions. The Journal of Supercomputing, 59 (3):1596–1618, March 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=59&issue=3&spage=1596.

Hsu:2012:ECO

[266] Ching-Hsien Hsu and Peter Sloot. Editorial: Communication optimization for scalable parallel system. The Journal of Supercomputing, 60 (1):1–3, April 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=1&spage=1.

Guo:2012:CFD

[267] Minyi Guo, Weng-Long Chang, Bo Jiang, Shu-Chien Huang, Sien-Tang Tsai, et al. Communicationfree data alignment for arrays with exponential references in parallelizing compilers for scalable parallel systems. The Journal of Supercomputing, 60(1):4–30, April 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=1&spage=4.

Wu:2012:UHM

[268] Chao-Chin Wu, Lien-Fu Lai, Chao-Tung Yang, and Po-Hsun Chiu. Using hybrid MPI and OpenMP programming to optimize communications in parallel loop self-scheduling schemes for multicore PC clusters. The Journal of Supercomputing, 60 (1):31–61, April 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=1&spage=31.

Tang:2012:CAE

[269] Feilong Tang and Minglu Li. Contextadaptive and energy-efficient mobile transaction management in pervasive environments. The Journal of Supercomputing, 60(1):62– 86, April 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=1&spage=62.

Massetto:2012:NSB

[270] Francisco Isidro Massetto, Liria Matsumoto Sato, and Kuan-Ching Li. A novel strategy for building interoperable MPI environment in heterogeneous high performance systems. The Journal of Supercomputing, 60 (1):87–116, April 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=1&spage=87.

Taboada:2012:FMS

[271] Guillermo L. Taboada, Juan Touriño, and Ramón Doallo. F-MPJ: scalable Java message-passing communications on parallel systems. The Journal of Supercomputing, 60(1): 117–140, April 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=1&spage=117.

Tu:2012:PAO

[272] Bibo Tu, Jianping Fan, Jianfeng Zhan, and Xiaofang Zhao. Performance analysis and optimization of MPI collective operations on multi-core clusters. The Journal of Supercomputing, 60(1):141–162, April 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=1&spage=141.

Khan:2012:EEH

[273] Samee Ullah Khan, Pascal Bouvry, and Thomas Engel. Energyefficient high-performance parallel and distributed computing. The Journal of Supercomputing, 60(2):163– 164, May 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=2&spage=163.

Lee:2012:PTM

[274] Eun Kyung Lee, Indraneel Kulkarni, Dario Pompili, and Manish Parashar. Proactive thermal management in green datacenters. The Journal of Supercomputing, 60(2): 165–195, May 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=2&spage=165.

Noori:2012:IPE

[275] Hamid Noori, Farhad Mehdipour, Koji Inoue, and Kazuaki Murakami. Improving performance and energy efficiency of embedded processors via post-fabrication instruction set customization. The Journal of Supercomputing, 60(2):196– 222, May 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=2&spage=196.

Li:2012:EES

[276] Keqin Li. Energy efficient scheduling of parallel tasks on multiprocessor computers. The Journal of Supercomputing, 60(2):223–247, May 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=2&spage=223.

Kdouh:2012:RAP

[277] Wael Kdouh and Hesham El-Rewini. Reliability-aware platform optimization for 3D chip multi-processors. The Journal of Supercomputing, 60 (2):248–267, May 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=2&spage=248.

Lee:2012:EEU

[278] Young Choon Lee and Albert Y. Zomaya. Energy efficient utilization of resources in cloud computing systems. The Journal of Supercomputing, 60(2):268–280, May 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=2&spage=268.

Yang:2012:P

[279] Chao-Tung Yang, Kuan-Chou Lai, Mitsuhisa Sato, and Tzung-Shi Chen. Preface. The Journal of Supercomputing, 60(3):281–283, June 2012. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=3&spage=281.

Hsu:2012:ESS

[280] Ching-Hsien Hsu and Shih-Chang Chen. Efficient selection strategies towards processor reordering techniques for improving data locality in heterogeneous clusters. The Journal of Supercomputing, 60(3):284– 300, June 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=3&spage=284.

Bsoul:2012:TBD

[281] Mohammad Bsoul, Ahmad Al-Khasawneh, Yousef Kilani, and Ibrahim Obeidat. A threshold-based dynamic data replication strategy. The Journal of Supercomputing, 60(3):301– 310, June 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=3&spage=301.

Shih:2012:IGP

[282] Po-Chi Shih, Kuo-Chan Huang, and Yeh-Ching Chung. Improving grid performance through processor allocation considering both speed heterogeneity and resource fragmentation. The Journal of Supercomputing, 60 (3):311–337, June 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=3&spage=311.

Chang:2012:SDE

[283] Yue-Shan Chang and Hsiang-Tai Cheng. A scientific data extraction architecture using classified metadata. The Journal of Supercomputing, 60 (3):338–359, June 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=3&spage=338.

Qu:2012:SBM

[284] Wenyu Qu, Keqiu Li, Masaru Kitsuregawa, and Weilian Xue. Statistical behaviors of mobile agents in network routing. The Journal of Supercomputing, 60(3):360– 388, June 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=60&issue=3&spage=360.

Lesage:2012:HCM

[285] Jean-Denis Lesage and Bruno Raffin. A hierarchical component model for large parallel interactive applications. The Journal of Supercomputing, 60(3):389–409, June 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=3&spage=389.

Akpan:2012:HOC

[286] Okon H. Akpan. On a high-order compact scheme and its utilization in parallel solution of a time-dependent system on a distributed memory processor. The Journal of Supercomputing, 60(3):410–419, June 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=3&spage=410.

Luo:2012:DMR

[287] Junzhou Luo, Zhiang Wu, Jiuxin Cao, and Tian Tian. Dynamic multi-resource advance reservation in grid environment. The Journal of Supercomputing, 60(3):420– 436, June 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=3&spage=420.

Yuanyuan:2012:IAM

[288] Zeng Yuanyuan, Naixue Xiong, Jong Hyuk Park, and Laurence T. Yang. An interference-aware multichannel media access control protocol for wireless sensor networks. The Journal of Supercomputing, 60 (3):437–460, June 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=60&issue=3&spage=437.

Hsu:2012:EET

[289] Ching-Hsien Hsu. Editorial: enabling technologies for programming extreme scale systems. The Journal of Supercomputing, 61(1): 1–5, July 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=1.

Kim:2012:DFA

[290] Byoung Uk Kim. Data flow analysis for anomaly detection and identification toward resiliency in extreme scale systems. The Journal of Supercomputing, 61(1):6–26, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=6.

Shi:2012:TSW

[291] Xuanhua Shi, Hai Jin, Hongbo Jiang, Xiaodong Pan, Dachuan Huang, and Bo Yu. Toward scalable Web systems on multicore clusters: making use of virtual machines. The Journal of Supercomputing, 61(1): 27–45, July 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=27.

Sharifi:2012:PED

[292] Mohsen Sharifi, Hadi Salimi, and Mahsa Najafzadeh. Power-efficient distributed scheduling of virtual machines using workload-aware consolidation techniques. The Journal of Supercomputing, 61(1):46– 66, July 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=46.

Wang:2012:SAC

[293] Wei-Jen Wang, Yue-Shan Chang, Cheng-Hui Wu, and Wei-Xiang Kang. A self-adaptive computing framework for parallel maximum likelihood evaluation. The Journal of Supercomputing, 61(1):67–83, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=67.

Yuan:2012:PCS

[294] Zhiyong Yuan, Weixin Si, Xiangyun Liao, Zhaoliang Duan, Yihua Ding, and Jianhui Zhao. Parallel computing of 3D smoking simulation based on OpenCL heterogeneous platform. The Journal of Supercomputing, 61 (1):84–102, July 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=84.

Li:2012:CSS

[295] Jia-Jhe Li, Chung-Kai Chen, Tung-Yu Wu, and Jenq Kuen Lee. Case study: stereo vision experiments with multi-core software API on embedded MPSoC environments. The Journal of Supercomputing, 61(1): 103–117, July 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=103.

Hong:2012:SSP

[296] Jingun Hong, Kirak Hong, Bernd Burgstaller, and Johann Blieberger. StreamPI: a stream-parallel programming extension for object-oriented programming languages. The Journal of Supercomputing, 61(1):118– 140, July 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=118.

Exposito:2012:DSJ

[297] Roberto R. Expósito, Guillermo L. Taboada, Juan Touriño, and Ramón Doallo. Design of scalable Java message-passing communications over InfiniBand. The Journal of Supercomputing, 61(1):141–165, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=1&spage=141.

Fanyang:2012:SAK

[298] Fanyang, Naixue Xiong, and Jong Hyuk Park. A self-adaptive K selection mechanism for re-authentication load balancing in large-scale systems. The Journal of Supercomputing, 61 (1):166–188, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=1&spage=166.

Li:2012:OCM

[299] Keqin Li. Optimal configuration of a multicore server processor for managing the power and performance tradeoff. The Journal of Supercomputing, 61(1):189–214, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=1&spage=189.

Cazalas:2012:LCS

[300] Jonathan Cazalas and Ratan Guha. Leveraging computation sharing and parallel processing in location-dependent query processing. The Journal of Supercomputing, 61(1):215–234, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0920-8542&volume=61&
                                        issue=1&spage=215.

Malyshkin:2012:OMP

[301] Victor Malyshkin and Vladislav Perepelkin. Optimization methods of parallel execution of numerical programs in the LuNA fragmented programming system. The Journal of Supercomputing, 61(1):235–248, July 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=1&spage=235.

Rong:2012:PSI

[302] Chunming Rong, Frode Eika Sandnes, and Rajkumar Buyya. Preface to special issue on Advances in Cloud Computing. The Journal of Supercomputing, 61(2):249– 250, August 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=2&spage=249.

Begnum:2012:SCO

[303] Kyrre Begnum. Simplified cloudoriented virtual machine management with MLN. The Journal of Supercomputing, 61(2):251– 266, August 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=2&spage=251.

Mowbray:2012:EPC

[304] Miranda Mowbray, Siani Pearson, and Yun Shen. Enhancing privacy in cloud computing via policy-based obfuscation. The Journal of Supercomputing, 61(2):267–291, August 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=2&spage=267.

Ranjan:2012:CLM

[305] Rajiv Ranjan, Aaron Harwood, and Rajkumar Buyya. Coordinated load management in peer-to-peer coupled federated grid systems. The Journal of Supercomputing, 61(2):292– 316, August 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=2&spage=292.

Liu:2012:ICD

[306] Jin Liu, Jing Zhou, Junfeng Wang, Feng Zhang, and Fei Liu. Irregular community discovery for cloud service improvement. The Journal of Supercomputing, 61(2):317– 336, August 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=2&spage=317.

Zhao:2012:RDM

[307] Gansen Zhao, Chunming Rong, Martin Gilje Jaatun, and Frode Eika Sandnes. Reference deployment models for eliminating user concerns on cloud security. The Journal of Supercomputing, 61(2):337– 352, August 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=2&spage=337.

Wu:2012:CIN

[308] Xiaoxin Wu, Wei Wang, Ben Lin, and Kai Miao. Composable IO: a novel resource sharing platform in personal Clouds. The Journal of Supercomputing, 61(2):353– 370, August 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=2&spage=353.

Jung:2012:TUO

[309] Jean-Pierre Jung and Ibrahima Sakho. Towards understanding optimal MIMD queueless routing of arbitrary permutations on hypercubes. The Journal of Supercomputing, 61(3):371– 393, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=371.

Zhou:2012:PCC

[310] Guoliang Zhou and Hong Chen. Parallel cube computation on modern CPUs and GPUs. The Journal of Supercomputing, 61(3):394– 417, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=394.

Lai:2012:FAA

[311] G. Lai and X. Lin. Floorplanaware application-specific networkon-chip topology synthesis using genetic algorithm technique. The Journal of Supercomputing, 61(3):418– 437, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=418.

Dang:2012:DDM

[312] Viet-Hung Dang, Sungyoung Lee, and Young-Koo Lee. A distributed design for multiple moving source positioning. The Journal of Supercomputing, 61(3): 438–462, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=438.

Tchendji:2012:ECG

[313] Vianney Kengne Tchendji and Jean Frédéric Myoupo. An efficient coarse-grain multicomputer algorithm for the minimum cost parenthesizing problem. The Journal of Supercomputing, 61(3): 463–480, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=463.

Seba:2012:ABC

[314] H. Seba, S. Lagraa, and H. Kheddouci. Alliance-based clustering scheme for group key management in mobile ad hoc networks. The Journal of Supercomputing, 61(3): 481–501, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=481.

Khan:2012:GPB

[315] Samee Ullah Khan and Nasro Min-Allah. A goal programming based energy efficient resource allocation in data centers. The Journal of Supercomputing, 61(3):502–519, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=502.

Niemi:2012:MBS

[316] Tapio Niemi and Ari-Pekka Hameri. Memory-based scheduling of scientific computing clusters. The Journal of Supercomputing, 61(3):520– 544, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=520.

Chen:2012:ABP

[317] Uei-Ren Chen and Woei Lin. Aligning block permutation methods for topology transformation on computational grids. The Journal of Supercomputing, 61(3):545–559, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=545.

Su:2012:MIH

[318] Hsun Su, Shih-Yan Chen, and Shin-Shin Kao. Mutually independent Hamiltonian cycles in alternating group graphs. The Journal of Supercomputing, 61(3):560– 571, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=560.

Mimaroglu:2012:ADC

[319] Selim Mimaroglu, Murat Yagci, and Dan A. Simovici. Approximative distance computation by random hashing. The Journal of Supercomputing, 61 (3):572–589, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=572.

Gravvanis:2012:SFD

[320] G. A. Gravvanis, C. K. Filelis-Papadopoulos, and K. M. Giannoutakis. Solving finite difference linear systems on GPUs: CUDA based parallel explicit preconditioned biconjugate conjugate gradient type methods. The Journal of Supercomputing, 61(3): 590–604, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=590.

Duh:2012:FPD

[321] Dyi-Rong Duh, Chien-Hong Chen, and Keh-Ning Chang. A fast pessimistic diagnosis algorithm for generalized hypercube multicomputer systems. The Journal of Supercomputing, 61(3): 605–618, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=605.

Safaei:2012:DSO

[322] Ali A. Safaei and Mostafa S. Haghjoo. Dispatching stream operators in parallel execution of continuous queries. The Journal of Supercomputing, 61(3): 619–641, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=619.

Chang:2012:MSR

[323] Weng-Long Chang, Kawuu Weicheng Lin, Ju-Chin Chen, Chih-Chiang Wang, Lai Chin Lu, Minyi Guo, and Michael (Shan-Hui) Ho. Molecular solutions of the RSA public-key cryptosystem on a DNA-based computer. The Journal of Supercomputing, 61(3): 642–672, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=642.

Pallipuram:2012:CSG

[324] Vivek K. Pallipuram, Mohammad Bhuiyan, and Melissa C. Smith. A comparative study of GPU programming models and architectures using neural networks. The Journal of Supercomputing, 61(3):673– 718, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=673.

Shieh:2012:PAR

[325] Wann-Yun Shieh and Bo-Syun Wang. Power-aware register assignment for large register file design. The Journal of Supercomputing, 61(3):719– 742, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=719.

Healy:2012:AME

[326] Philip D. Healy and John P. Morrison. ARC: a metacomputing environment for clusters augmented with reconfigurable hardware. The Journal of Supercomputing, 61(3):743– 779, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=743.

Wang:2012:TAW

[327] Lizhe Wang, Samee U. Khan, and Jai Dayal. Thermal aware workload placement with task-temperature profiles in a data center. The Journal of Supercomputing, 61(3):780– 803, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=780.

Chen:2012:AMI

[328] Wei Chen, Dan Chen, and Zhiying Wang. An approach to minimizing the interpretation overhead in Dynamic Binary Translation. The Journal of Supercomputing, 61(3): 804–825, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=804.

Xia:2012:UAA

[329] Fei Xia, Yong Dou, and Guoqing Jin. The unified accelerator architecture for RNA secondary structure prediction on FPGA. The Journal of Supercomputing, 61(3):826– 855, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=826.

Cheng:2012:SAA

[330] Eddie Cheng, Ke Qiu, and Zhizhang Shen. On the surface area of the augmented cubes. The Journal of Supercomputing, 61(3):856– 868, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=856.

Niu:2012:GCB

[331] Yu Niu, Brian J. d’Auriol, and Sungyoung Lee. General criteriabased clustering method for multinode computing system. The Journal of Supercomputing, 61(3):869– 893, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=869.

Liu:2012:MRP

[332] Shaoshan Liu, Richard Neil Pittman, Alessandro Forin, and Jean-Luc Gaudiot. Minimizing the runtime partial reconfiguration overheads in reconfigurable systems. The Journal of Supercomputing, 61(3):894– 911, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=894.

Barshan:2012:IAP

[333] Maryam Barshan, Mahmood Fathy, and Saleh Yousefi. Improving the availability of P2P-based network management systems by provisioning fault tolerance property. The Journal of Supercomputing, 61(3): 912–934, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=912.

Motallebi:2012:DLO

[334] Hassan Motallebi and Saeed Parsa. Data locality optimization of interference graphs based on polyhedral computations. The Journal of Supercomputing, 61(3):935– 965, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=935.

Cha:2012:RCC

[335] Kwangho Cha and Seungryoul Maeng. Reducing communication costs in collective I/O in multi-core cluster systems with non-exclusive scheduling. The Journal of Supercomputing, 61(3): 966–996, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=966.

Sanchez:2012:FTA

[336] Daniel Sánchez, Juan L. Aragón, and José M. García. A faulttolerant architecture for parallel applications in tiled-CMPs. The Journal of Supercomputing, 61(3):997– 1023, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=997.

Wu:2012:ISM

[337] Chung-Ju Wu, Yu-Te Lin, and Jenq-Kuen Lee. Instruction scheduling methods and phase ordering framework for VLIW DSP processors with distributed register files. The Journal of Supercomputing, 61(3):1024– 1047, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=1024.

Lin:2012:PLE

[338] Jing Lin and Xiaola Lin. Power and latency efficient mechanism: a seamless bridge between buffered and bufferless routing in on-chip network. The Journal of Supercomputing, 61(3): 1048–1067, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=1048.

Suresh:2012:SND

[339] S. Suresh, H. J. Kim, Cui Run, and T. G. Robertazzi. Scheduling nonlinear divisible loads in a single level tree network. The Journal of Supercomputing, 61(3):1068– 1088, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=61&issue=3&spage=1068.

dAuriol:2012:SV

[340] Brian J. d’Auriol. Serviceable visualizations. The Journal of Supercomputing, 61(3):1089–1115, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0920-8542&volume=61&
                                        issue=3&spage=1089.

Kuo:2012:HDS

[341] Chin-Lin Kuo, Ming-Jeng Yang, Yao-Ming Chang, and Yao-Ming Yeh. High diagnosability of a sequential diagnosis algorithm in hypercubes under the PMC model. The Journal of Supercomputing, 61(3):1116– 1134, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=1116.

Terzopoulos:2012:PER

[342] George Terzopoulos and Helen Karatza. Performance evaluation of a realtime grid system using power-saving capable processors. The Journal of Supercomputing, 61(3):1135–1153, September 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=1135.

Zhang:2012:SML

[343] Yang Zhang and Weixing Ji. A scalable method-level parallel library and its improvement. The Journal of Supercomputing, 61(3):1154– 1167, September 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=61&issue=3&spage=1154.

Kim:2012:GEA

[344] SooKyun Kim, Johnnes Arreymbi, and Chia-Chen Lin. Guest editorial: Advances in digital and multimedia convergence. The Journal of Supercomputing, 62(1):1–3, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=1.

Misra:2012:LAB

[345] Sudip Misra, P. Venkata Krishna, Akhil Bhiwal, Amardeep Singh Chawla, Bernd E. Wolfinger, and Changhoon Lee. A learning automata-based fault-tolerant routing algorithm for mobile ad hoc networks. The Journal of Supercomputing, 62(1):4– 23, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=4.

Wang:2012:RRN

[346] Hwang-Cheng Wang, Isaac Woungang, Jia-Bao Lin, Fang-Chang Kuo, and Kuo-Chang Ting. Revisiting relative neighborhood graph-based broadcasting algorithms for multimedia ad hoc wireless networks. The Journal of Supercomputing, 62(1):24– 41, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=24.

Lee:2012:SCI

[347] Song-Hee Lee, Nam-Sup Park, and Jin-Young Choi. Secure communication in IPTV broadcasting. The Journal of Supercomputing, 62(1):42– 52, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=42.

Lu:2012:OAS

[348] Wenhuan Lu, Naixue Xiong, and Doo-Soon Park. An ontological approach to support legal information modeling. The Journal of Supercomputing, 62(1):53–67, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=53.

Cheng:2012:ITQ

[349] Rung-Shiang Cheng, Cheng-Han Lin, Jiann-Liang Chen, and Han-Chieh Chao. Improving transmission quality of MPEG video stream by SCTP multistreaming and differential RED mechanisms. The Journal of Supercomputing, 62(1):68–83, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=68.

Butt:2012:LLQ

[350] Muhammad Rizwan Butt, Ali Hammad Akbar, Ki-Hyung Kim, Muhammad Mohsin Javed, Chae-Seong Lim, and Quratulain Taj. LABILE: Link quAlity-based lexIcaL routing mEtric for reactive routing protocols in IEEE 802.15.4 networks. The Journal of Supercomputing, 62(1):84– 104, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=84.

Park:2012:APO

[351] Jong Hyuk Park. An authentication protocol offering service anonymity of mobile device in ubiquitous environment. The Journal of Supercomputing, 62(1):105–117, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=105.

Seo:2012:FSH

[352] Sekwang Seo, Sang-Soo Yeo, and Young-Sik Jeong. FSH scheme for high-speed handover and anti-MITM on mobile computing. The Journal of Supercomputing, 62(1):118– 133, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=118.

Al-Sadi:2012:TPE

[353] Jehad Al-Sadi. Topological properties of the Extended OTIS-n-Cube interconnection network. The Journal of Supercomputing, 62(1):134– 149, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=134.

Simms:2012:PSD

[354] Andrew M. Simms and Valerie Daggett. Protein simulation data in the relational model. The Journal of Supercomputing, 62(1):150– 173, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=150.

Chen:2012:TCD

[355] Ching-Wen Chen and Chang-Jung Ku. A tagless cache design for power saving in embedded systems. The Journal of Supercomputing, 62 (1):174–198, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=174.

Kanal:2012:MMC

[356] M. E. Kanal and M. Demiralp. A modified method of calculating High Dimensional Model Representation (HDMR) Terms for parallelization with MPI and CUDA. The Journal of Supercomputing, 62(1):199– 213, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=199.

Kas:2012:TCD

[357] Miray Kas. Toward on-chip datacenters: a perspective on general trends and on-chip particulars. The Journal of Supercomputing, 62(1): 214–226, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=214.

Wang:2012:NDN

[358] Zhe Wang, Tao Li, Naixue Xiong, and Yi Pan. A novel dynamic network data replication scheme based on historical access record and proactive deletion. The Journal of Supercomputing, 62(1):227–250, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=227. See

erratum [373].

Sharma:2012:FEE

[359] Rohit Sharma, T. Chakravarty, and K. Choi. Fast and efficient extraction algorithm for high-speed interconnects with arbitrary boundaries. The Journal of Supercomputing, 62 (1):251–264, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=251.

Tosun:2012:ERA

[360] Suleyman Tosun. Energy- and reliability-aware task scheduling onto heterogeneous MPSoC architectures. The Journal of Supercomputing, 62 (1):265–289, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=265.

Falzon:2012:EGA

[361] Geoffrey Falzon and Maozhen Li. Enhancing genetic algorithms for dependent job scheduling in grid computing environments. The Journal of Supercomputing, 62(1):290– 314, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=290.

Heydarian:2012:HPO

[362] Mohsen Heydarian. A high performance optimal dynamic routing algorithm with unicast multichannel QoS guarantee in communication systems. The Journal of Supercomputing, 62 (1):315–344, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=315.

Ding:2012:PCC

[363] Zuohua Ding, Hui Shen, and Jianwen Cao. Parallel computation of continuous Petri nets based on hypergraph partitioning. The Journal of Supercomputing, 62(1):345– 377, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=345.

Green:2012:CFO

[364] Robert C. Green II, Lingfeng Wang, Mansoor Alam, and Richard A. Formato. Central force optimization on a GPU: a case study in high performance metaheuristics. The Journal of Supercomputing, 62(1):378– 398, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=378.

Wu:2012:PEE

[365] Chao-Chin Wu, Lien-Fu Lai, Liang-Tsung Huang, and MingLung Chen. Performance evaluation of enhancement of the layered self-scheduling approach for heterogeneous multicore cluster systems. The Journal of Supercomputing, 62(1):399– 430, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=399.

Pervez:2012:SSH

[366] Zeeshan Pervez, Asad Masood Khattak, Sungyoung Lee, and Young-Koo Lee. SAPDS: self-healing attributebased privacy aware data sharing in cloud. The Journal of Supercomputing, 62(1):431–460, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=431.

Briceno:2012:CIA

[367] Luis Diego Briceño, Howard Jay Siegel, Anthony A. Maciejewski, and Mohana Oltikar. Characterization of the iterative application of makespan heuristics on non-makespan machines in a heterogeneous parallel and distributed environment. The Journal of Supercomputing, 62(1):461– 485, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=461.

Baransel:2012:PIS

[368] Cesur Baransel and Kayhan M. Imre. A parallel implementation of Strassen’s matrix multiplication algorithm for wormhole-routed all-port 2D torus networks. The Journal of Supercomputing, 62(1):486–509, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=486.

Fang:2012:AMC

[369] Zhen Fang, Lixin Zhang, John B. Carter, Sally A. McKee, Ali Ibrahim, Michael A. Parker, and Xiaowei Jiang. Active memory controller. The Journal of Supercomputing, 62(1):510– 549, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=510.

Aksari:2012:FBS

[370] Yigitcan Aksari and Harun Artuner. Forward and back substitution algorithms on GPU: a case study on modified incomplete Cholesky Preconditioner for three-dimensional finite difference method. The Journal of Supercomputing, 62(1):550– 572, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=550.

Touzene:2012:NPB

[371] Abderezak Touzene. A new parallel block aggregated algorithm for solving Markov chains. The Journal of Supercomputing, 62(1):573– 587, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=573.

Al-Dayaa:2012:TML

[372] H. S. Al-Dayaa and D. B. Megherbi. Towards a Multiple-Lookahead-Levels agent reinforcement-learning technique and its implementation in integrated circuits. The Journal of Supercomputing, 62(1):588– 615, October 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=1&spage=588.

Wang:2012:END

[373] Zhe Wang, Tao Li, Naixue Xiong, and Yi Pan. Erratum to: “A novel dynamic network data replication scheme based on historical access record and proactive deletion”. The Journal of Supercomputing, 62 (1):616, October 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=1&spage=616. See

[358].

Dolev:2012:OSI

[374] Shlomi Dolev, Tobias Haist, and Mihai Oltean. Optical supercomputing: introduction to special issue. The Journal of Supercomputing, 62(2): 617–619, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=617.

Fey:2012:OMT

[375] D. Fey, M. Schneider, J. Jahns, and H. Knuppertz. Optical multiplexing techniques for photonic Clos networks in High Performance Computing Architectures. The Journal of Supercomputing, 62(2):620– 632, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=620.

Tamir:2012:PDC

[376] Dan E. Tamir, Natan T. Shaked, Wilhelmus J. Geerts, and Shlomi Dolev. Parallel decomposition of combinatorial optimization problems using electro-optical vector by matrix multiplication architecture. The Journal of Supercomputing, 62(2): 633–655, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=633.

Haist:2012:WLI

[377] Tobias Haist and Wolfgang Osten. White-light interferometric method for secure key distribution. The Journal of Supercomputing, 62(2): 656–662, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=656.

Goliaei:2012:OSS

[378] Sama Goliaei and Saeed Jalili. An optical solution to the 3-SAT problem using wavelength based selectors. The Journal of Supercomputing, 62(2): 663–672, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=663.

Zlotnik:2012:OSI

[379] Alexander Zlotnik, Melania Paturzo, Pietro Ferraro, and Zeev Zalevsky. Optical spatial image processor based on aliasing of pseudo-periodic sampling. The Journal of Supercomputing, 62(2): 673–680, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=673.

Caulfield:2012:ZEO

[380] H. John Caulfield, Andrey Zavalin, and Lei Qian. Zero-energy optical logic: can it be practical? The Journal of Supercomputing, 62(2): 681–688, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=681.

Li:2012:OSL

[381] Xiujian Li, Wenhua Hu, Hualiang Zhang, Yongming Nie, Jiankun Yang, and Junbo Yang. Operation speed limited by the electric properties of the photorefractive spatial light modulator. The Journal of Supercomputing, 62 (2):689–697, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=689.

Almeida:2012:FAM

[382] Francisco Almeida, Javier Cuenca, Domingo Giménez, Antonio Llanes-Castro, and Juan-Pedro Martínez-Gallar. A framework for the application of metaheuristics to tasksto-processors assignation problems. The Journal of Supercomputing, 62(2): 698–723, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=698.

Li:2012:EPN

[383] Cheng Hua Li, Iker Gondra, and Lijun Liu. An efficient parallel neural network-based multi-instance learning algorithm. The Journal of Supercomputing, 62(2):724– 740, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=724.

Deboosere:2012:ERM

[384] Lien Deboosere, Bert Vankeirsbilck, Pieter Simoens, Filip De Turck, Bart Dhoedt, and Piet Demeester. Efficient resource management for virtual desktop cloud computing. The Journal of Supercomputing, 62(2): 741–767, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=741.

Chtepen:2012:OET

[385] Maria Chtepen, Filip H. A. Claeys, Bart Dhoedt, Filip De Turck, Jan Fostier, Piet Demeester, and Peter A. Vanrolleghem. Online execution time prediction for computationally intensive applications with periodic progress updates. The Journal of Supercomputing, 62(2):768– 786, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=768.

Cecilia:2012:SCH

[386] José M. Cecilia, José L. Abellán, Juan Fernández, Manuel E. Acacio, José M. García, and Manuel Ujaldón. Stencil computations on heterogeneous platforms for the Jacobi method: GPUs versus Cell BE. The Journal of Supercomputing, 62(2): 787–803, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=787.

Syed:2012:FAD

[387] Raheel Hassan Syed, Jasmina Pazardzievska, and Julien Bourgeois. Fast attack detection using correlation and summarizing of security alerts in grid computing networks. The Journal of Supercomputing, 62(2):804– 827, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=804.

Guo:2012:SSR

[388] Yang Guo, Wanxia Qu, Long Zhang, and Weixia Xu. State space reduction in modeling checking parameterized cache coherence protocol by two-dimensional abstraction. The Journal of Supercomputing, 62(2): 828–854, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=828.

Nazir:2012:RBF

[389] Babar Nazir, Kalim Qureshi, and Paul Manuel. Replication based fault tolerant job scheduling strategy for economy driven grid. The Journal of Supercomputing, 62(2): 855–873, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=855.

Cheng:2012:IAE

[390] Chien-Fu Cheng and Kuo-Tang Tsai. From immediate agreement to eventual agreement: early stopping agreement protocol for dynamic networks with malicious faulty processors. The Journal of Supercomputing, 62(2): 874–894, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=874.

Jiang:2012:TLA

[391] Fuu-Cheng Jiang, Hsiang-Wei Wu, and Chao-Tung Yang. Traffic load analysis and its application to enhancing longevity on IEEE 802.15.4/ ZigBee sensor network. The Journal of Supercomputing, 62(2):895– 915, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=895.

Avila-George:2012:SGC

[392] Himer Avila-George, Jose Torres-Jimenez, Nelson Rangel-Valdez, Abel Carrión, and Vicente Hernández. Supercomputing and grid computing on the verification of covering arrays. The Journal of Supercomputing, 62(2): 916–945, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=916. See

erratum [1041].

Dursun:2012:HPO

[393] Hikmet Dursun, Manaschai Kunaseth, Ken ichi Nomura, Jacqueline Chame, Robert F. Lucas, Chun Chen, Mary Hall, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. Hierarchical parallelization and optimization of high-order stencil computations on multicore clusters. The Journal of Supercomputing, 62(2):946– 966, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=946.

Wang:2012:EET

[394] Hwang-Cheng Wang, Isaac Woungang, Cheng-Wen Yao, Alagan Anpalagan, and Mohammad S. Obaidat. Energy-efficient tasks scheduling algorithm for real-time multiprocessor embedded systems. The Journal of Supercomputing, 62(2): 967–988, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=967.

Li:2012:ORP

[395] Chunlin Li and La Yuan Li. Optimal resource provisioning for cloud computing environment. The Journal of Supercomputing, 62(2):989– 1022, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=989.

Davis:2012:PSE

[396] Neal E. Davis, Robert W. Robey, Charles R. Ferenbaugh, David Nicholaeff, and Dennis P. Trujillo. Paradigmatic shifts for exascale supercomputing. The Journal of Supercomputing, 62(2):1023–1044, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0920-8542&volume=62&
                                        issue=2&spage=1023.

Shahbahrami:2012:AAD

[397] Asadollah Shahbahrami. Algorithms and architectures for 2D discrete wavelet transform. The Journal of Supercomputing, 62(2):1045– 1064, November 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0920-8542&
                                        volume=62&issue=2&spage=1045.

Li:2012:PAC

[398] Wenzhong Li, Edward Chan, Daoxu Chen, and Sanglu Lu. Performance analysis of cache consistency strategies for multi-hop wireless networks. The Journal of Supercomputing, 62(2): 1065–1090, November 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&
volume=62&issue=2&spage=1065.

Khan:2012:GN

[399] Samee Ullah Khan, Sherali Zeadally, Pascal Bouvry, and Naveen Chilamkurti. Green networks. The Journal of Supercomputing, 62(3):1091– 1092, December 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0640-2;   http:
//link.springer.com/content/pdf/
10.1007/s11227-011-0640-2.

Zeadally:2012:EEN

[400] Sherali Zeadally, Samee Ullah Khan, and Naveen Chilamkurti. Energyefficient networking: past, present, and future. The Journal of Supercomputing, 62(3):1093–1118, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0632-2.

Yen:2012:NOB

[401] Hong-Hsu Yen, Steven S. W. Lee, and Florence G. H. Yap. A novel optimization-based bandwidth-aware minimum power multicast routing algorithm in green wireless networks. The Journal of Supercomputing, 62(3): 1119–1138, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0629-x.

Orgerie:2012:EEB

[402] Anne-Cécile Orgerie, Laurent Lefèvre, and Isabelle Guérin-Lassous. Energyefficient bandwidth reservation for bulk data transfers in dedicated wired networks. The Journal of Supercomputing, 62(3):1139–1166, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0603-7.

Hlavacs:2012:EEP

[403] Helmut Hlavacs, Roman Weidlich, and Thomas Treutner. Energy efficient peer-to-peer file sharing. The Journal of Supercomputing, 62(3):1167– 1188, December 2012. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0602-8.

Hamza:2012:CDS

[404] Haitham S. Hamza. Convert-and-deliver: a scalable multicast optical cross-connect with reduced power splitting fan-out. The Journal of Supercomputing, 62(3):1189–1212, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0565-9.

Ruiz:2012:OED

[405] Patricia Ruiz, Bernabé Dorronsoro, Giorgio Valentini l Frédéric Pinel, and Pascal Bouvry. Optimisation of the enhanced distance based broadcasting protocol for MANETs. The Journal of Supercomputing, 62(3):1213–1240, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0564-x.

Misra:2012:LEE

[406] Sudip Misra, P. Venkata Krishna, and V. Saritha. LACAV: an energyefficient channel assignment mechanism for vehicular ad hoc networks. The Journal of Supercomputing, 62(3): 1241–1262, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0552-1.

Kliazovich:2012:GPL

[407] Dzmitry Kliazovich, Pascal Bouvry, and Samee Ullah Khan. GreenCloud: a packet-level simulator of energyaware cloud computing data centers. The Journal of Supercomputing, 62(3): 1263–1283, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0504-1.

Garcia-Dorado:2012:NME

[408] José Luis García-Dorado, Eduardo Magaña, Pedro Reviriego, Daniel Morató Mikel Izal, Juan Antonio Maestro, Javier Aracil, and Jorge E. López de Vergara. Network monitoring for energy efficiency in large-scale networks: the case of the Spanish Academic Network. The Journal of Supercomputing, 62(3):1284–1304, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0643-z.

Espinosa:2012:AIM

[409] A. Espinosa, P. Hernandez, J. C. Moure, J. Protasio, and A. Ripoll. Analysis and improvement of mapreduce data distribution in read mapping applications. The Journal of Supercomputing, 62(3):1305–1317, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0792-8.

Chang:2012:IPM

[410] Kuei-Chung Chang, Ing-Ming Liao, and Chiu-Han Liao. Improving performance of multi-core NUCA coherent systems using NoC-assisted mechanisms. The Journal of Supercomputing, 62(3):1318–1337, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0793-7.

Cho:2012:EMR

[411] Haengrae Cho. Energy management for a real-time shared disk cluster. The Journal of Supercomputing, 62(3): 1338–1361, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0794-6.

Choi:2012:DHD

[412] Hyun-Hwa Choi, Mi-Young Lee, and Kyu-Chul Lee. Distributed high dimensional indexing for k-NN search. The Journal of Supercomputing, 62(3): 1362–1384, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0800-z.

Arif:2012:ESS

[413] Samiur Arif and Stephan Olariu. Efficient solution of a stochastic SI epidemic system. The Journal of Supercomputing, 62(3):1385–1403, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0802-x.

Wang:2012:SNA

[414] Chao Wang, Xi Li, Junneng Zhang, Xuehai Zhou, and Aili Wang. A star network approach in heterogeneous multiprocessors system on chip. The Journal of Supercomputing, 62(3): 1404–1424, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0810-x.

Ahmad:2012:ALS

[415] Mohammad Z. Ahmad and Ratan Guha. Analysis of large scale traceroute datasets in Internet routing overlays by parallel computation. The Journal of Supercomputing, 62(3):1425–1450, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0811-9.

Zheng:2012:SFL

[416] Xufei Zheng, Tao Li, and Yonghui Fang. Strategy of fast and light-load cloud-based proactive benign worm countermeasure technology to contain worm propagation. The Journal of Supercomputing, 62(3):1451–1479, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0812-8.

Fatone:2012:POP

[417] Lorella Fatone, Marco Giacinti, Francesca Mariani, Maria Cristina Recchioni, and Francesco Zirilli. Parallel option pricing on GPU: barrier options and realized variance options. The Journal of Supercomputing, 62(3): 1480–1501, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0813-7.

Tanase:2012:DUD

[418] Cristian Andy Tanase and Vasile Gheorghita Gaitan. Dynamic, unbalanced distribution of tasks on a PS3 cluster system for double precision calculation. The Journal of Supercomputing, 62(3): 1502–1518, December 2012. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0814-6;   http:
//link.springer.com/content/pdf/
10.1007/s11227-012-0814-6.

Dou:2013:EMO

[419] Wanchun Dou, Lianyong Qi, Xuyun Zhang, and Jinjun Chen. An evaluation method of outsourcing services for developing an elastic cloud platform. The Journal of Supercomputing, 63(1):1–23, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0491-2.

UlHaq:2013:RBV

[420] Irfan Ul Haq, Adrian Paschke, Erich Schikuta, and Harold Boley. Rulebased validation of SLA choreographies. The Journal of Supercomputing, 63(1):24–45, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0492-1.

Czarnul:2013:MRT

[421] Pawel Czarnul. Modeling, run-time optimization and execution of distributed workflow applications in the JEE-based BeesyCluster environment. The Journal of Supercomputing, 63 (1):46–71, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0499-7;   http:
//link.springer.com/content/pdf/
10.1007/s11227-010-0499-7.

Cao:2013:SPO

[422] Jian Cao, Jie Wang, Haiyan Zhao, and Xiaohan Sun. A service process optimization method based on model refinement. The Journal of Supercomputing, 63(1):72–88, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-010-
                                        0513-0.

Yu:2013:HAW

[423] Yang Yu, Ting Xie, and Xiaoyan Wang. A handling algorithm for workflow time exception based on history logs. The Journal of Supercomputing, 63 (1):89–106, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-010-0543-7.

Davidrajuh:2013:DWB

[424] Reggie Davidrajuh. Distributed workflow based approach for eliminating redundancy in virtual enterprising. The Journal of Supercomputing, 63 (1):107–125, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-010-0544-6.

Ren:2013:BDG

[425] Kaijun Ren, Junqiang Song, Min Zhu, and Nong Xiao. A bargaining-driven global QoS adjustment approach for optimizing service composition execution path. The Journal of Supercomputing, 63(1):126–149, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-011-
                                        0645-x.

Xiao:2013:DDV

[426] Nong Xiao, Bin Chen, and Zhiping Cai. A demand-driven virtual disk prefetch mechanism for seamless mobility of personal computing environments. The Journal of Supercomputing, 63(1):150–170, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0495-y.

Ren:2013:ECR

[427] Yizhi Ren, Mingchu Li, Yang Xiang, Yongrui Cui, and Kouichi Sakurai. Evolution of cooperation in reputation system by group-based scheme. The Journal of Supercomputing, 63 (1):171–190, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0498-8.

Vrba:2013:NRT

[428] Zeljko Vrba, Pål Halvorsen, Carsten Griwodz, Paul Beskow, Håvard Espeland, and Dag Johansen. The Nornir run-time system for parallel programs using Kahn process networks on multicore machines — a flexible alternative to MapReduce. The Journal of Supercomputing, 63(1):191–217, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0503-2;   http://link.springer.
com/content/pdf/10.1007/s11227-
010-0503-2.

Liu:2013:DST

[429] Weijiang Liu, Wenyu Qu, Xiaona He, and Zhaobin Liu. Detecting superpoints through a reversible counting Bloom filter. The Journal of Supercomputing, 63(1):218–234, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-010-
                                        0511-2.

Yoon:2013:RBB

[430] Eun-Jun Yoon and Kee-Young Yoo. Robust biometrics-based multi-server authentication with key agreement scheme for smart cards on elliptic curve cryptosystem. The Journal of Supercomputing, 63(1):235–255, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-010-
                                        0512-1.

Wu:2013:MOH

[431] Zhangjun Wu, Xiao Liu, Zhiwei Ni, Dong Yuan, and Yun Yang. A market-oriented hierarchical scheduling strategy in cloud workflow systems. The Journal of Supercomputing, 63(1):256–293, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0578-4.

Egi:2013:IPS

[432] Norbert Egi, Gianluca Iannaccone, Maziar Manesh, Laurent Mathy, and Sylvia Ratnasamy. Improved parallelism and scheduling in multi-core software routers. The Journal of Supercomputing, 63(1):294–322, January 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-011-
                                        0579-3.

Vishnu:2013:GEI

[433] Abhinav Vishnu, Pavan Balaji, and Yong Chen. Guest Editors’ introduction. The Journal of Supercomputing, 63(2):323–325, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0743-4;   http://link.springer.
com/content/pdf/10.1007/s11227-
012-0743-4.

Young:2013:DEC

[434] B. Dalton Young, Jonathan Apodaca, Luis Diego Briceño, Jay Smith, Sudeep Pasricha, Anthony A. Maciejewski, Howard Jay Siegel, Bhavesh Khemka, Shirish Bahirat, Adrian Ramirez, and Yong Zou. Deadline and energy constrained dynamic resource allocation in a heterogeneous computing environment. The Journal of Supercomputing, 63(2):326–347, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0740-7.

Leung:2013:RAC

[435] Kai-Cheung Leung, Yawen Chen, and Zhiyi Huang. Restricted admission control in view-oriented transactional memory. The Journal of Supercomputing, 63(2):348–366, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0733-y.

Tang:2013:MDJ

[436] Wei Tang, Narayan Desai, Venkatram Vishwanath, Daniel Buettner, and Zhiling Lan. Multi-domain job coscheduling for leadership computing systems. The Journal of Supercomputing, 63(2):367–384, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0741-6.

Berka:2013:CPC

[437] Tobias Berka, Giorgos Kollias, Helge Hagenauer, Marian Vajteršic, and Ananth Grama. Concurrent programming constructs for parallel MPI applications. The Journal of Supercomputing, 63(2):385–406, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-011-
                                        0739-5.

Cheng:2013:DSA

[438] Hongju Cheng, Naixue Xiong, Larence T. Yang, and Young-Sik Jeong. Distributed scheduling algorithms for channel access in TDMA wireless mesh networks. The Journal of Supercomputing, 63(2):407–430, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-008-
                                        0244-7.

Jin:2013:MAB

[439] Yingwei Jin, Wenyu Qu, Yong Zhang, and Yong Wang. A mobile agentbased routing model for grid computing. The Journal of Supercomputing, 63 (2):431–442, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0616-2.

Li:2013:GAP

[440] Ruipeng Li and Yousef Saad. GPUaccelerated preconditioned iterative linear solvers. The Journal of Supercomputing, 63(2):443–466, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0825-3.

Javadi:2013:EPF

[441] Bahman Javadi, Parimala Thulasiraman, and Rajkumar Buyya. Enhancing performance of failure-prone clusters by adaptive provisioning of cloud resources. The Journal of Supercomputing, 63(2):467–489, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0826-2.

Cai:2013:FLS

[442] Xianggao Cai, Guoming Lai, and Xiaola Lin. Forecasting large scale conditional volatility and covariance using neural network on GPU. The Journal of Supercomputing, 63(2): 490–507, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0827-1.

Yang:2013:ATI

[443] Qianming Yang, Mei Wen, Nan Wu, and Chunyuan Zhang. Accelerating thread-intensive and explicit memory management programs with dynamic partial reconfiguration. The Journal of Supercomputing, 63(2):508–537, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0828-0.

Pervez:2013:PAS

[444] Zeeshan Pervez, Ammar Ahmad Awan, Asad Masood Khattak, Sungyoung Lee, and Eui-Nam Huh. Privacy-aware searching with oblivious term matching for cloud storage. The Journal of Supercomputing, 63(2):538–560, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0829-z.

Modi:2013:SSI

[445] Chirag Modi, Dhiren Patel, Bhavesh Borisaniya, Avi Patel, and Muttukrishnan Rajarajan. A survey on security issues and solutions at different layers of Cloud computing. The Journal of Supercomputing, 63(2):561–592, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0831-5.

Zhou:2013:XPH

[446] Rui Zhou, Qingguo Zhou, Yong Sheng, and Kuan-Ching Li. XtratuM/PPC: a hypervisor for partitioned system on PowerPC processors. The Journal of Supercomputing, 63(2):593–610, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-

0833-3. See erratum [447].

Zhou:2013:EXP

[447] Rui Zhou, Qingguo Zhou, Yong Sheng, and Kuan-Ching Li. Erratum to: XtratuM/PPC: a hypervisor for partitioned system on PowerPC processors. The Journal of Supercomputing, 63(2):611, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0838-y;   http:
//link.springer.com/content/pdf/
10.1007/s11227-012-0838-y.  See

[446].

Babamir:2013:SVR

[448] Seyed Morteza Babamir. Specification and verification of reliability in dispatching multicast messages. The Journal of Supercomputing, 63(2): 612–635, February 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0834-2.

Khan:2013:GCC

[449] Samee Ullah Khan, Lizhe Wang, Laurence T. Yang, and Feng Xia. Green computing and communications. The Journal of Supercomputing, 63 (3):637–638, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.

                                        1007/s11227-011-0718-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-011-0718-x.

Wang:2013:RPM

[450] Lizhe Wang and Samee U. Khan. Review of performance metrics for green data centers: a taxonomy study. The Journal of Supercomputing, 63 (3):639–656, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0704-3.

Chen:2013:LMI

[451] Zhikui Chen, Feng Xia, Tao Huang, Fanyu Bu, and Haozhe Wang. A localization method for the Internet of Things. The Journal of Supercomputing, 63(3):657–674, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-011-
                                        0693-2.

Chen:2013:MPM

[452] Dan Chen, Lizhe Wang, Mingwei Tian, Jian Tian, Shuaiting Wang, Congcong Bian, and Xiaoli Li. Massively parallel modelling & simulation of large crowd with GPGPU. The Journal of Supercomputing, 63(3):675–690, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-011-
                                        0675-4.

Vishnu:2013:DEE

[453] Abhinav Vishnu, Shuaiwen Song, Andres Marquez, Kevin Barker, Darren Kerbyson, Kirk Cameron, and Pavan Balaji. Designing energy efficient communication runtime systems: a view from PGAS models. The Journal of Supercomputing, 63(3):691–709, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0699-9.

Xu:2013:PMO

[454] Shiming Xu, Wei Xue, and Hai Xiang Lin. Performance modeling and optimization of sparse matrix-vector multiplication on NVIDIA CUDA platform. The Journal of Supercomputing, 63(3):710–721, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0626-0;   http:
//link.springer.com/content/pdf/
10.1007/s11227-011-0626-0.

Guabtni:2013:WDA

[455] Adnene Guabtni, Rajiv Ranjan, and Fethi A. Rabhi. A workload-driven approach to database query processing in the cloud. The Journal of Supercomputing, 63(3):722–736, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0717-y.

Danoy:2013:ESI

[456] Grégoire Danoy. Editorial: special issue on parallel nature-inspired optimization. The Journal of Supercomputing, 63(3):737–739, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0820-8;   http://link.springer.
com/content/pdf/10.1007/s11227-
012-0820-8.

deAraujoMacedo:2013:MBS

[457] Emerson de Araujo Macedo, Alba Cristina Magalhaes Alves de Melo, Gerson Henrique Pfitscher, and Azzedine Boukerche. Multiple biological sequence alignment in heterogeneous multicore clusters with user-selectable task allocation policies. The Journal of Supercomputing, 63(3):740–756, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0768-8.

Korosec:2013:MCI

[458] Peter Korosec, Marian Vajtersic, Jurij Silc, and Rade Kutil. Multicore implementation of the differential ant-stigmergy algorithm for numerical optimization. The Journal of Supercomputing, 63(3):757–772, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0772-z.

Cecilia:2013:EGP

[459] José M. Cecilia, Andy Nisbet, Martyn Amos, José M. García, and Manuel Ujaldón. Enhancing GPU parallelism in nature-inspired algorithms. The Journal of Supercomputing, 63 (3):773–789, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0770-1.

Yoo:2013:CAK

[460] Paul D. Yoo and Albert Y. Zomaya. Combining analytic kernel models for energy-efficient data modeling and classification. The Journal of Supercomputing, 63(3):790–799, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0776-8.

Piwonska:2013:LCA

[461] Anna Piwonska, Franciszek Seredynski, and Miroslaw Szaban. Learning cellular automata rules for binary classification problem. The Journal of Supercomputing, 63(3):800–815, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0767-9.

Dorronsoro:2013:CGA

[462] Bernabé Dorronsoro and Pascal Bouvry. Cellular genetic algorithms without additional parameters. The Journal of Supercomputing, 63(3):816–835, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0773-y.

Khouadjia:2013:MEC

[463] Mostepha R. Khouadjia, El-Ghazali Talbi, Laetitia Jourdan, Briseida Sarasola, and Enrique Alba. Multienvironmental cooperative parallel metaheuristics for solving dynamic optimization problems. The Journal of Supercomputing, 63(3):836–853, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0774-x.

Seredynski:2013:ADC

[464] Marcin Seredynski and Pascal Bouvry. Analysing the development of cooperation in MANETs using evolutionary game theory. The Journal of Supercomputing, 63(3):854–870, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0769-7.

Heydarian:2013:NHP

[465] Mohsen Heydarian and Reza Hadi Mogavi. A new high performance approach: merging optimal multicast sessions for supporting multisource routing. The Journal of Supercomputing, 63(3):871–896, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0835-1.

Goude:2013:AFM

[466] Anders Goude and Stefan Engblom. Adaptive fast multipole methods on the GPU. The Journal of Supercomputing, 63(3):897–918, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0836-0;   http:
//link.springer.com/content/pdf/
10.1007/s11227-012-0836-0.

Czarnul:2013:MDI

[467] Pawel Czarnul. A model, design, and implementation of an efficient multithreaded workflow execution engine with data streaming, caching, and storage constraints. The Journal of Supercomputing, 63(3):919–945, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0837-z;   http://link.springer.
com/content/pdf/10.1007/s11227-
012-0837-z.

Shiraz:2013:SVM

[468] Muhammad Shiraz, Saeid Abolfazli, Zohreh Sanaei, and Abdullah Gani. A study on virtual machine deployment for application outsourcing in mobile cloud computing. The Journal of Supercomputing, 63(3):946–964, March 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0846-y.

Cortina:2013:PHP

[469] Raquel Cortina, F. J. Martínez-Zaldívar, Antonio M. Vidal, and Jesús Vigo-Aguiar. Preface to high performance computing applied to computational problems in science and engineering. The Journal of Supercomputing, 64(1):1–3, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0817-3;   http:
//link.springer.com/content/pdf/
10.1007/s11227-012-0817-3.

Galiano:2013:PSD

[470] V. Galiano, O. López, M. P. Malumbres, and H. Migallón. Parallel strategies for 2D Discrete Wavelet Transform in shared memory systems and GPUs. The Journal of Supercomputing, 64(1):4–16, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0750-5.

Lopez-Portugues:2013:PFD

[471] Miguel López-Portugués, Jesús A. López-Fernández, José Ranilla, R. G. Ayestarán, and Fernando Las-Heras. Parallelization of the FMM on distributed-memory GPGPU systems for acoustic-scattering prediction. The Journal of Supercomputing, 64(1): 17–27, April 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0786-6.

Frances:2013:DUF

[472] Jorge Francés, Sergio Bleda, Sergi Gallego, Cristian Neipp, Andrés Márquez, Inmaculada Pascual, and Augusto Beléndez. Development of a unified FDTD–FEM library for electromagnetic analysis with CPU and GPU computing. The Journal of Supercomputing, 64(1):28–37, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0803-9.

Herrera:2013:TAQ

[473] Juan F. R. Herrera, Leocadio G. Casado, Eligius M. T. Hendrix, and Inmaculada García. A threaded approach of the quadratic bi-blending algorithm. The Journal of Supercomputing, 64(1):38–48, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0783-9;   http:
//link.springer.com/content/pdf/
10.1007/s11227-012-0783-9.

Ortega:2013:BGM

[474] G. Ortega, E. M. Garzón, F. Vázquez, and I. García. The BiConjugate gradient method on GPUs. The Journal of Supercomputing, 64(1):49–58, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0761-2.

Fresno:2013:EHT

[475] Javier Fresno, Arturo Gonzalez-Escribano, and Diego R. Llanos. Extending a hierarchical tiling arrays library to support sparse data partitioning. The Journal of Supercomputing, 64(1):59–68, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0757-y.

Montanola:2013:PAC

[476] Alberto Montañola, Concepció Roig, Fernando Guirado, Porfidio Hernández, and Cedric Notredame. Performance analysis of computational approaches to solve Multiple Sequence Alignment. The Journal of Supercomputing, 64(1):69–78, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0751-4.

Rodriguez-Sanchez:2013:HAI

[477] Rafael Rodríguez-Sánchez, José Luis Martínez, Gerardo Fernández-Escribano, José Manuel Claver, and José L. Sánchez. H.264/AVC interprediction for heterogeneous computing systems. The Journal of Supercomputing, 64(1):79–88, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0782-x.

Toharia:2013:SSB

[478] Pablo Toharia, Oscar D. Robles, Jose L. Bosque, and Angel Rodríguez. Scalable shot boundary detection. The Journal of Supercomputing, 64 (1):89–99, April 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0784-8.

Gonzalez-Dominguez:2013:PES

[479] Jorge González-Domínguez, Óscar García-López, Guillermo L. Taboada, María J. Martín, and Juan Touriño. Performance evaluation of sparse matrix products in UPC. The Journal of Supercomputing, 64(1):100–109, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0796-4.

Bosque:2013:ASP

[480] Jose L. Bosque, Oscar D. Robles, Pablo Toharia, and Luis Pastor. Analyzing scalability of parallel systems with unbalanced workload. The Journal of Supercomputing, 64(1):110–119, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0765-y.

Lobeiras:2013:IMA

[481] J. Lobeiras, M. Amor, and R. Doallo. Influence of memory access patterns to small-scale FFT performance. The Journal of Supercomputing, 64(1): 120–131, April 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0807-5.

Alvarez-Bermejo:2013:SSK

[482] J. A. Álvarez-Bermejo, N. Antequera, R. García-Rubio, and J. A. López-Ramos. A scalable server for key distribution and its application to accounting. The Journal of Supercomputing, 64(1):132–143, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0787-5.

Diaz:2013:TLH

[483] Antonio F. Díaz, Mancia Anguita, Hugo E. Camacho, Erik Nieto, and Julio Ortega. Two-level hash/table approach for metadata management in distributed file systems. The Journal of Supercomputing, 64(1):144–155, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11227-012-
                                        0801-y.

Vigueras:2013:RCU

[484] Guillermo Vigueras, Juan M. Orduña, and Miguel Lozano. A Read–Copy Update based parallel server for distributed crowd simulations. The Journal of Supercomputing, 64(1):156–166, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0766-x.

Acosta:2013:LSP

[485] Alejandro Acosta, Francisco Almeida, and Ignacio Peláez. From LaTEX specifications to parallel codes. The Journal of Supercomputing, 64(1):167–176, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0797-3.

Abdeyazdan:2013:TGP

[486] Marjan Abdeyazdan, Saeed Parsa, and Amir Masoud Rahmani. Task graph pre-scheduling, using Nash equilibrium in game theory. The Journal of Supercomputing, 64(1):177–203, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0845-z.

Mansouri:2013:JSD

[487] Najme Mansouri and Gholam Hosein Dastghaibyfard. Job scheduling and dynamic data replication in data grid environment. The Journal of Supercomputing, 64(1):204–225, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0850-2.

AkbariTorkestani:2013:DCM

[488] Javad Akbari Torkestani. Degree constrained minimum spanning tree problem: a learning automata approach. The Journal of Supercomputing, 64 (1):226–249, April 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0851-1.

Hussain:2013:ACN

[489] Sajid Hussain, Firdous Kausar, Laurence T. Yang, Farrukh A. Khan, and Jong Hyuk Park. Advances in communication networks for pervasive and ubiquitous applications. The Journal of Supercomputing, 64(2):251–255, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0927-6.

Zhang:2013:DCC

[490] Daqiang Zhang, Min Chen, Hongyu Huang, and Minyi Guo. Decentralized checking of context inconsistency in pervasive computing environments. The Journal of Supercomputing, 64 (2):256–273, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0661-x.

Xiong:2013:ABD

[491] Wei Xiong, Naixue Xiong, Laurence T. Yang, Jong Hyuk Park, Hanping Hu, and Qian Wang. An anomaly-based detection in ubiquitous network using the equilibrium state of the catastrophe theory. The Journal of Supercomputing, 64(2):274–294, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0644-y.

Loni:2013:AFS

[492] Zia M. Loni and Noor M. Khan. Analysis of fading statistics in cellular mobile communication systems. The Journal of Supercomputing, 64(2):295–309, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0799-1.

Naderan:2013:PDB

[493] Marjan Naderan, Mehdi Dehghan, and Hossein Pedram. Primal and dualbased algorithms for sensing range adjustment in WSNs. The Journal of Supercomputing, 64(2):310–330, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-011-
                                        0702-5.

Choi:2013:MCC

[494] Min Choi, Jonghyuk Park, and Young-Sik Jeong. Mobile cloud computing framework for a pervasive and ubiquitous environment. The Journal of Supercomputing, 64(2):331–356, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0681-6.

Papapostolou:2013:HEA

[495] Apostolia Papapostolou and Hakima Chaouchi. Handoff with energy awareness for future pervasive environments. The Journal of Supercomputing, 64 (2):357–382, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0742-5.

Liu:2013:TCN

[496] Hang Liu, Eun Kyung Lee, Dario Pompili, and Xiangwei Kong. Thermal camera networks for large datacenters using real-time thermal monitoring mechanism. The Journal of Supercomputing, 64(2):383–408, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0781-y.

Ramrekha:2013:SAA

[497] Tipu Arvind Ramrekha, Emmanouil Panaousis, and Christos Politis. Standardisation advancements in the area of routing for mobile ad-hoc networks. The Journal of Supercomputing, 64(2):409–434, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0705-2.

Vaidya:2013:SCM

[498] Binod Vaidya, Dimitrios Makrakis, and Hussein Mouftah. Secure communication mechanism for ubiquitous Smart grid infrastructure. The Journal of Supercomputing, 64(2):435–455, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0674-5.

Bueno-Delgado:2013:MLB

[499] M. V. Bueno-Delgado and P. Pavón-Mariño. A maximum likelihood-based distributed protocol for passive RFID dense reader environments. The Journal of Supercomputing, 64(2):456–476, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0779-5.

Cho:2013:SDC

[500] Do-Eun Cho, Byoung-Soo Koh, and Sang-Soo Yeo. Secure D-CAS system for digital contents downloading services. The Journal of Supercomputing, 64(2):477–491, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0719-9.

Park:2013:HPN

[501] Sang Oh Park, Yang Sun Lee, and Sung Jo Kim. A high performance NAND array file system based on multiple NAND flash memories. The Journal of Supercomputing, 64(2):492–506, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0714-1.

Park:2013:ARB

[502] Jang Woo Park, Dae Heon Park, and Changhoon Lee. Angle and ranging based localization method for ad hoc network. The Journal of Supercomputing, 64(2):507–521, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0700-7.

Cho:2013:DLM

[503] Jaeik Cho, Taeshik Shon, Ken Choi, and Jongsub Moon. Dynamic learning model update of hybrid-classifiers for intrusion detection. The Journal of Supercomputing, 64(2):522–526, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0698-x.

Atoofian:2013:IPS

[504] Ehsan Atoofian. Improving performance of software transactional memory through contention locality. The Journal of Supercomputing, 64 (2):527–547, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0854-y.

Lira:2013:RTD

[505] Javier Lira, Carlos Molina, Ryan N. Rakvic, and Antonio González. Replacement techniques for dynamic NUCA cache designs on CMPs. The Journal of Supercomputing, 64(2): 548–579, May 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0859-6.

Lei:2013:FIE

[506] Yuanwu Lei, Yong Dou, Yazhuo Dong, Jie Zhou, and Fei Xia. FPGA implementation of an exact dot product and its application in variable-precision floating-point arithmetic. The Journal of Supercomputing, 64(2):580–605, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0860-0.

Son:2013:SBC

[507] Seokho Son, Gihun Jung, and Sung Chan Jun. An SLA-based cloud computing that facilitates resource allocation in the distributed data centers of a cloud provider. The Journal of Supercomputing, 64(2):606–637, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0861-z.

Tinetti:2013:RFL

[508] Fernando G. Tinetti, Mariano Méndez, and Armando De Giusti. Restructuring Fortran legacy applications for parallel computing in multiprocessors. The Journal of Supercomputing, 64 (2):638–659, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0863-x.

Wang:2013:TSP

[509] Guojun Wang, Wanlei Zhou, and Laurence T. Yang. Trust, security and privacy for pervasive applications. The Journal of Supercomputing, 64 (3):661–663, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-013-0953-4.pdf.

Good:2013:HAE

[510] Tim Good and Mohammed Benaissa. A holistic approach examining RFID design for security and privacy. The Journal of Supercomputing, 64(3):664–684, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0497-9.

Zeng:2013:SLL

[511] Yingpei Zeng, Jiannong Cao, Jue Hong, Shigeng Zhang, and Li Xie. Secure localization and location verification in wireless sensor networks: a survey. The Journal of Supercomputing, 64(3):685–701, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0501-4.

Li:2013:RDN

[512] Ruidong Li and Jie Li. Requirements and design for neutral trust management framework in unstructured networks. The Journal of Supercomputing, 64(3):702–716, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0502-3.

Pongaliur:2013:DCE

[513] Kanthakumar Pongaliur, Li Xiao, and Alex X. Liu. Dynamic camouflage event based malicious node detection architecture. The Journal of Supercomputing, 64(3):717–743, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0508-x.

Wang:2013:ARB

[514] Fei Wang, Fu Rong Wang, Benxiong Huang, and Laurence T. Yang. ADVS: a reputation-based model on filtering SPIT over P2P–VoIP networks. The Journal of Supercomputing, 64(3):744–761, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0545-5.

Qi:2013:URA

[515] Fang Qi, Zhe Tang, Guojun Wang, and Jie Wu. User requirements-aware security ranking in SSL protocol. The Journal of Supercomputing, 64(3):762–776, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0546-4.

Prathapani:2013:DBA

[516] Anoosha Prathapani, Lakshmi Santhanam, and Dharma P. Agrawal. Detection of blackhole attack in a Wireless Mesh Network using intelligent honeypot agents. The Journal of Supercomputing, 64(3):777–804, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0547-3.

Zhao:2013:VTV

[517] Huanyu Zhao and Xiaolin Li. VectorTrust: trust vector aggregation scheme for trust management in peerto-peer networks. The Journal of Supercomputing, 64(3):805–829, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0576-6.

Kou:2013:HPN

[518] Gang Kou. High performance networked computing in media, services and information management. The Journal of Supercomputing, 64 (3):830–834, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-013-0942-7.pdf.

Ergu:2013:AHP

[519] Daji Ergu, Gang Kou, Yi Peng, Yong Shi, and Yu Shi. The analytic hierarchy process: task scheduling and resource allocation in cloud computing environment. The Journal of Supercomputing, 64(3):835–848, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0625-1.

Sadri:2013:ISS

[520] Yasser Sadri and Sohrab Khanmohamadi. An intelligent scheduling system using fuzzy logic controller for management of services in WiMAX networks. The Journal of Supercomputing, 64(3):849–861, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-010-
                                        0523-y.

Zhong:2013:RCP

[521] Ning Zhong, Jian Hua Ma, Run He Huang, Ji Ming Liu, Yi Yu Yao, Yao Xue Zhang, and Jian Hui Chen. Research challenges and perspectives on Wisdom Web of Things (W2T). The Journal of Supercomputing, 64 (3):862–882, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-010-0518-8.

Wang:2013:CRA

[522] Honggang Wang. Communicationresource-aware adaptive watermarking for multimedia authentication in wireless multimedia sensor networks. The Journal of Supercomputing, 64 (3):883–897, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-010-0500-5.

Wu:2013:LCM

[523] Dengsheng Wu, Jianping Li, and Yong Liang. Linear combination of multiple case-based reasoning with optimized weight for software effort estimation. The Journal of Supercomputing, 64(3):898–918, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0525-9.

You:2013:QAP

[524] Kun You, Bin Tang, Zhuzhong Qian, Sanglu Lu, and Daoxu Chen. QoSaware placement of stream processing service. The Journal of Supercomputing, 64(3):919–941, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0548-2.

Jian:2013:PDM

[525] Liheng Jian, Cheng Wang, Ying Liu, Shenshen Liang, Weidong Yi, and Yong Shi. Parallel data mining techniques on Graphics Processing Unit with Compute Unified Device Architecture (CUDA). The Journal of Supercomputing, 64(3):942–967, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0672-7.

Song:2013:RRO

[526] Guanghua Song, Bowei Yang, Zhixing Wu, Junna Chuai, and Yao Zheng. ROIN: reputation-oriented inverted indexing for the P2P network. The Journal of Supercomputing, 64(3):968–986, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0878-y.

Seo:2013:TDP

[527] Jung hyun Seo. Three-dimensional Petersen-torus network: a fixed-degree network for massively parallel computers. The Journal of Supercomputing, 64(3):987–1007, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0716-z.

Huang:2013:SSP

[528] Xiaohuang Huang, Christopher I. Rodrigues, Stephen Jones, Ian Buck, and Wen mei Hwu. Scalable SIMD-parallel memory allocation for many-core machines. The Journal of Supercomputing, 64(3):1008–1020, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0680-7.

Xie:2013:EAO

[529] Bin Xie, Tianzhou Chen, Wei Hu, Xingsheng Tang, and Dazhou Wang. An energy-aware online task mapping algorithm in NoC-based system. The Journal of Supercomputing, 64 (3):1021–1037, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0678-1.

Chen:2013:HBA

[530] Quan Chen, Minyi Guo, Qianni Deng, Long Zheng, Song Guo, and Yao Shen. HAT: history-based auto-tuning MapReduce in heterogeneous environments. The Journal of Supercomputing, 64(3):1038–1054, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0682-5.

Li:2013:EIM

[531] Guohui Li, Pei Zhao, Ling Yuan, and Sheng Gao. Efficient implementation of a multi-dimensional index structure over flash memory storage systems. The Journal of Supercomputing, 64(3):1055–1074, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0679-0.

Wang:2013:SES

[532] Guojun Wang, Qiushuang Du, Wei Zhou, and Qin Liu. A scalable encryption scheme for multi-privileged group communications. The Journal of Supercomputing, 64(3):1075–1091, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0683-4.

Yuan:2013:SSP

[533] Xiaoqun Yuan, Hao Yin, Geyong Min, Xuening Liu, Wen Hui, and Guangxi Zhu. A suitable server placement for peer-to-peer live streaming. The Journal of Supercomputing, 64 (3):1092–1107, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0685-2.

Liu:2013:CSA

[534] Jin Liu, Jing Zhou, Junfeng Wang, Xue Chen, and Hainan Zhou. Cloud service: automatic construction and evolution of software process problemsolving resource space. The Journal of Supercomputing, 64(3):1108–1132, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0489-9.

Li:2013:EWG

[535] Yinan Li, Asim YarKhan, Jack Dongarra, Keith Seymour, and Aurèlie Hurault. Enabling workflows in GridSolve: request sequencing and service trading. The Journal of Supercomputing, 64(3):1133–1152, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0549-1.

Garg:2013:MME

[536] Saurabh Kumar Garg, Christian Vecchiola, and Rajkumar Buyya. Mandi: a market exchange for trading utility and cloud computing services. The Journal of Supercomputing, 64(3):1153–1174, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0568-6.

Kim:2013:ESS

[537] Tai hoon Kim and Sabah Mohammed. Editorial of special section on multimedia applications and ubiquitous computing. The Journal of Supercomputing, 65(1):1–3, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-013-0945-4.pdf.

Zhang:2013:DVA

[538] Meng Zhang, Byeong Ho Kang, and Quan Bai. Discover and visualize association rules from sensor observations on the Web. The Journal of Supercomputing, 65(1):4–15, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0697-y.

Kim:2013:RAA

[539] Jognwoo Kim, Sanggil Kang, Yujin Lim, and Hak-Man Kim. Recommendation algorithm of the app store by using semantic relations between apps. The Journal of Supercomputing, 65(1):16–26, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0701-6.

Sung:2013:CIC

[540] Kihoon Sung, Hee-Kyung Kong, and Taehan Kim. Convergence indicator: the case of cloud computing. The Journal of Supercomputing, 65(1):27– 37, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0706-1.

Yamauchi:2013:MAB

[541] Toshihiro Yamauchi, Takayuki Hara, and Hideo Taniguchi. A mechanism for achieving a bound on execution performance of process group to limit CPU abuse. The Journal of Supercomputing, 65(1):38–60, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0707-0.

Kim:2013:TDS

[542] Tai hoon Kim and Sabah Mohammed. Toward designing a secure biosurveillance cloud. The Journal of Supercomputing, 65(1):61–70, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://link.springer.com/
article/10.1007/s11227-011-0709-
y.

Um:2013:DVS

[543] Jung ho Um, Hoon Choi, Sa kwang Song, Sung pil Choi, Hwa mook Yoon, Hanmin Jung, and Tai hoon Kim. Development of a virtualized supercomputing environment for genomic analysis. The Journal of Supercomputing, 65(1):71–85, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0752-3.

Lee:2013:SMS

[544] Chongdeuk Lee. Streaming media service based on fuzzy similarity in wireless mobile networks. The Journal of Supercomputing, 65(1):86–105, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0778-6.

Jeong:2013:FOG

[545] Seungdo Jeong and Jungwon Cho. A framework for online gait recognition based on multilinear tensor analysis. The Journal of Supercomputing, 65(1):106–121, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0785-7.

Jeong:2013:GRU

[546] Seungdo Jeong, Tai hoon Kim, and Jungwon Cho. Gait recognition using description of shape synthesized by planar homography. The Journal of Supercomputing, 65(1):122–135, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0897-8.

Wu:2013:ISI

[547] Chia-Ming Wu, Ruay-Shiung Chang, Pu-I Lee, and Jei-Hsiang Yen. An innovative scheme for increasing connectivity and life of ZigBee networks. The Journal of Supercomputing, 65 (1):136–153, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0696-z.

Ranjan:2013:PPS

[548] Rajiv Ranjan and Liang Zhao. Peerto-peer service provisioning in cloud computing environments. The Journal of Supercomputing, 65(1):154–184, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0710-5.

Flahive:2013:OSO

[549] Andrew Flahive, David Taniar, and Wenny Rahayu. Ontology as a Service (OaaS): a case for sub-ontology merging on the cloud. The Journal of Supercomputing, 65(1):185–216, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0711-4.

Pedersen:2013:DMD

[550] Torben Bach Pedersen, Dennis Pedersen, and Karsten Riis. On-demand multidimensional data integration: toward a semantic foundation for cloud intelligence. The Journal of Supercomputing, 65(1):217–257, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0712-3.

Park:2013:MTP

[551] Jong Hyuk Park, Zhiwen Yu, and Liang Zhou. Multimedia technology for pervasive computing environment. The Journal of Supercomputing, 65(1):258–261, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0944-5.

Nakariyakul:2013:FSA

[552] Songyot Nakariyakul. Fast spatial averaging: an efficient algorithm for 2D mean filtering. The Journal of Supercomputing, 65(1):262–273, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0638-9.

Rho:2013:BSG

[553] Seungmin Rho and Sang-Soo Yeo. Bridging the semantic gap in multimedia emotion/mood recognition for ubiquitous computing environment. The Journal of Supercomputing, 65 (1):274–286, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0447-6.

Chen:2013:ELB

[554] Min Chen, Sergio González, Huasong Cao, Yan Zhang, and Son T. Vuong. Enabling low bit-rate and reliable video surveillance over practical wireless sensor network. The Journal of Supercomputing, 65(1):287–300, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0475-2.

Koo:2013:NAV

[555] Bonhyun Koo, Yang Sun Lee, and Taeshik Shon. A novel approach to visualize Web anomaly attacks in pervasive computing environment. The Journal of Supercomputing, 65(1):301–316, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0520-1.

Zhang:2013:SMD

[556] Hongguang Zhang, Hang Nguyen, Eduardo Martínez Graciá, Pedro Antonio Tudela Solano, Daqing Zhang, Noël Crespi, and Bin Guo. Scalable multimedia delivery with QoS management in pervasive computing environment. The Journal of Supercomputing, 65(1):317–335, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0581-9.

Kim:2013:VBA

[557] DoHyung Kim, Jaeyeon Lee, Ho-Sub Yoon, and Jaehong Kim. Visionbased arm gesture recognition for a long-range human-robot interaction. The Journal of Supercomputing, 65 (1):336–352, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0541-9.

Chang:2013:RTV

[558] J.-K. Chang, Seungteak Ryoo, and Heuiseok Lim. Real-time vehicle tracking mechanism with license plate recognition from road images. The Journal of Supercomputing, 65(1):353–364, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0580-x.

Kim:2013:LMC

[559] Yong hwan Kim, Youn-Hee Han, and Young-Sik Jeong. Lifetime maximization considering target coverage and connectivity in directional image/ video sensor networks. The Journal of Supercomputing, 65(1):365–382, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0646-9.

Cui:2013:LBA

[560] Yong Cui, Tianze Ma, Jiangchuan Liu, and Sajal Das. Load-balanced AP association in multi-hop wireless mesh networks. The Journal of Supercomputing, 65(1):383–409, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0519-7.

Li:2013:SDS

[561] Jianguo Li, Gansen Zhao, Chunming Rong, and Yong Tang. Semantic description of scholar-oriented social network cloud. The Journal of Supercomputing, 65(1):410–425, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-010-
0550-8.

Yang:2013:PEC

[562] Bo Yang, Feng Tan, and Yuan-Shun Dai. Performance evaluation of cloud service considering fault recovery. The Journal of Supercomputing, 65(1):426–444, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0551-2.

Jing:2013:SAR

[563] Si-Yuan Jing, Shahzad Ali, Kun She, and Yi Zhong. State-of-the-art research study for green cloud computing. The Journal of Supercomputing, 65(1):445–468, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0722-1.

Seitkulov:2013:NMS

[564] Yerzhan N. Seitkulov. New methods of secure outsourcing of scientific computations. The Journal of Supercomputing, 65(1):469–482, July 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0809-3.

Shon:2013:ESS

[565] Taeshik Shon, Shiuh-Jeng Wang, Lei Shu, and Liudong Xing. Editorial of special section on advanced in high performance, algorithm, and framework for future computing. The Journal of Supercomputing, 65(2):483, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
013-0963-2.pdf.

Choi:2013:RRB

[566] Min Choi, Jong Hyuk Park, and Young-Sik Jeong. Revisiting reorder buffer architecture for next generation high performance computing. The Journal of Supercomputing, 65 (2):484–495, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0734-x.

Chen:2013:SPS

[567] Bing-Chang Chen and Her-Tyan Yeh. Secure proxy signature schemes from the Weil pairing. The Journal of Supercomputing, 65(2):496–506, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0760-3.

Lee:2013:MGP

[568] Yangsun Lee, Leonard Barolli, and Seung-Ho Lim. Mapping granularity and performance tradeoffs for solid state drive. The Journal of Supercomputing, 65(2):507–523, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0798-2.

Woungang:2013:ASI

[569] Isaac Woungang, Sanjay Kumar Dhurandher, Lakshaya Agnani, Ankit Mahendru, and Alagan Anpalagan. An ant-swarm inspired dynamic multiresolution data dissemination protocol for wireless sensor networks. The Journal of Supercomputing, 65(2):524–542, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0804-8.

Papadakis:2013:LCG

[570] Harris Papadakis, Costas Panagiotakis, and Paraskevi Fragopoulou. Locating communities on graphs with variations in community sizes. The Journal of Supercomputing, 65(2):543–561, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0806-6.

Kim:2013:AIN

[571] Eui-Jik Kim, Sungkwan Youm, Taeshik Shon, and Chul-Hee Kang. Asynchronous inter-network interference avoidance for wireless body area networks. The Journal of Supercomputing, 65(2):562–579, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0840-4.

Lim:2013:ENR

[572] Hye-Youn Lim and Dae-Seong Kang. Efficient noise reduction in images using directional modified sigma filter. The Journal of Supercomputing, 65 (2):580–592, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0844-0.

Cho:2013:SFE

[573] Jae-Hyuk Cho, Kil-Woo Lee, Hong-Min Son, and Hyun-Sik Kim. A study on framework for effective R&D performance analysis of Korea using the Bayesian network and pairwise comparison of AHP. The Journal of Supercomputing, 65(2):593–611, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0876-0.

DAlotto:2013:SSC

[574] Lou D’Alotto, James F. Nystrom, and William Spataro. Special section on Cellular Automata. The Journal of Supercomputing, 65(2):612–613, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
013-0961-4.pdf.

Blecic:2013:CAS

[575] Ivan Blecic, Arnaldo Cecchini, and Giuseppe A. Trunfio. Cellular automata simulation of urban dynamics through GPGPU. The Journal of Supercomputing, 65(2):614–629, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0913-z.

DAmbrosio:2013:EAG

[576] Donato D’Ambrosio, Giuseppe Filippone, Davide Marocco, Rocco Rongo, and William Spataro. Efficient application of GPGPU for lava flow hazard mapping. The Journal of Supercomputing, 65(2):630–644, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0949-0.

Sergeyev:2013:STM

[577] Yaroslav D. Sergeyev and Alfredo Garro. Single-tape and multi-tape Turing machines through the lens of the Grossone methodology. The Journal of Supercomputing, 65(2):645–663, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0894-y.

Kalogeropoulos:2013:CAF

[578] G. Kalogeropoulos, G. C. Sirakoulis, and I. Karafyllidis. Cellular automata on FPGA for real-time urban traffic signals control. The Journal of Supercomputing, 65(2):664–681, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0952-5.

Avolio:2013:SSN

[579] Maria Vittoria Avolio, Salvatore Di Gregorio, Valeria Lupiano, and Paolo Mazzanti. SCIDDICA–SS_{3}: a new version of cellular automata model for simulating fast moving landslides. The Journal of Supercomputing, 65 (2):682–696, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0948-1.

Stafiej:2013:CPP

[580] Janusz Stafiej, Dung di Caprio, and Lukasz Bartosik. Corrosion-passivation processes in a cellular automata based simulation study. The Journal of Supercomputing, 65(2):697–709, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        013-0933-8.pdf.

Babaali:2013:NDG

[581] P. Babaali, E. Carta-Gerardino, and C. Knaplund. The number of DFAs for a given spanning tree. The Journal of Supercomputing, 65(2):710–722, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0957-0.

Keshavarz-Kohjerdi:2013:EPA

[582] Fatemeh Keshavarz-Kohjerdi and Alireza Bagheri. An efficient parallel algorithm for the longest path problem in meshes. The Journal of Supercomputing, 65(2):723–741, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0852-0.

Prakash:2013:NSM

[583] Shiv Prakash and Deo Prakash Vidyarthi. A novel scheduling model for computational grid using quantum genetic algorithm. The Journal of Supercomputing, 65(2):742–770, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0864-9.

Kapoor:2013:DFV

[584] Hemangee K. Kapoor, Praveen Kanakala, Malti Verma, and Shirshendu Das. Design and formal verification of a hierarchical cache coherence protocol for NoC based multiprocessors. The Journal of Supercomputing, 65(2):771–796, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0865-8.

Ghafarian:2013:PAL

[585] Toktam Ghafarian, Hossein Deldari, and Bahman Javadi. A proximityaware load balancing in peer-to-peer-based volunteer computing systems. The Journal of Supercomputing, 65 (2):797–822, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0866-7.

Lin:2013:EPP

[586] Xuan-Yi Lin, Kuan-Chou Lai, Kuan-Ching Li, and Yeh-Ching Chung. Efficient programming paradigm for video streaming processing on TILE64 platform. The Journal of Supercomputing, 65(2):823–847, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0867-6.

Galiano:2013:FWT

[587] V. Galiano, O. López-Granado, M. P. Malumbres, and H. Migallón. Fast 3D wavelet transform on multicore and many-core computing platforms. The Journal of Supercomputing, 65 (2):848–865, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0868-0.

Li:2013:ERA

[588] Chunlin Li and Layuan Li. Efficient resource allocation for optimizing objectives of cloud users, IaaS provider and SaaS provider in cloud environment. The Journal of Supercomputing, 65(2):866–885, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0869-z.

Choi:2013:ESS

[589] Hong Jun Choi, Dong Oh Son, Seung Gu Kang, and Jong Myon Kim. An efficient scheduling scheme using estimated execution time for heterogeneous computing systems. The Journal of Supercomputing, 65(2):886–902, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0870-6.

Lee:2013:ABA

[590] Wonjun Lee, Anna Squicciarini, and Elisa Bertino. Agent-based accountable grid computing systems. The Journal of Supercomputing, 65(2):903–929, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0871-5.

Lorenzo-Castillo:2013:FDP

[591] Juan A. Lorenzo-Castillo and Juan C. Pichel. A flexible and dynamic page migration infrastructure based on hardware counters. The Journal of Supercomputing, 65(2):930–948, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0872-4.

Picazo-Sanchez:2013:CRS

[592] Pablo Picazo-Sanchez, Lara Ortiz-Martin, and Pedro Peris-Lopez. Cryptanalysis of the RNTS system. The Journal of Supercomputing, 65(2):949– 960, August 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0873-3.

Boloorchi:2013:EES

[593] A. T. Boloorchi and M. H. Samadzadeh. Energy-efficient and secure in-network storage and retrieval for WSNs: an adaptive approach. The Journal of Supercomputing, 65(2):961–977, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0874-2.

Cesnovar:2013:GIS

[594] Rok Cesnovar, Vladimir Risojević, and Zdenka Babić. A GPU implementation of a structural-similarity-based aerial-image classification. The Journal of Supercomputing, 65(2):978–996, August 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0875-1.

Almeida:2013:HPC

[595] Francisco Almeida and Jesús Vigo-Aguiar. High performance computing tools in science and engineering. The Journal of Supercomputing, 65(3): 997–998, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-013-0989-5.pdf.

Alonso:2013:MSB

[596] Pedro Alonso, Daniel Argüelles, José Ranilla, and Antonio M. Vidal. A multicore solution to Block-Toeplitz linear systems of equations. The Journal of Supercomputing, 65(3):999– 1009, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0824-4.

Ramiro:2013:MIF

[597] Carla Ramiro, Sandra Roger, Alberto Gonzalez, Vicenc Almenar, and Antonio M. Vidal. Multicore implementation of a fixed-complexity tree-search detector for MIMO communications. The Journal of Supercomputing, 65(3): 1010–1019, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0839-x.

Ortiz:2013:LBI

[598] Andrés Ortiz, Julio Ortega, Antonio F. Díaz, and Mancia Anguita. Leveraging bandwidth improvements to Web servers through enhanced network interfaces. The Journal of Supercomputing, 65(3):1020–1036, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0841-3.

Naranjo:2013:FDA

[599] J. A. M. Naranjo, F. Cores, L. G. Casado, and F. Guirado. Fully distributed authentication with locality exploitation for the CoDiP2P peer-to-peer computing platform. The Journal of Supercomputing, 65(3):1037– 1049, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0842-2.

Teijeiro:2013:PSB

[600] Carlos Teijeiro, Godehard Sutmann, Guillermo L. Taboada, and Juan Touriño. Parallel simulation of Brownian dynamics on shared memory systems with OpenMP and Unified Parallel C. The Journal of Supercomputing, 65(3):1050–1062, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0843-1.

Reyes:2013:PEO

[601] Ruymán Reyes, Iván López, Juan J. Fumero, and Francisco de Sande. A preliminary evaluation of OpenACC implementations. The Journal of Supercomputing, 65(3):1063– 1075, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0853-z.

Orobitg:2013:IMS

[602] Miquel Orobitg, Fernando Cores, Fernando Guirado, Concepció Roig, and Cedric Notredame. Improving multiple sequence alignment biological accuracy through genetic algorithms. The Journal of Supercomputing, 65(3): 1076–1088, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0856-9.

Gonzalez:2013:NSP

[603] Carlos H. González, Basilio B. Fraguela, Diego Andrade, José A. García, and Manuel J. Castro. Numerical simulation of pollutant transport in a shallow-water system on the Cell heterogeneous processor. The Journal of Supercomputing, 65(3): 1089–1103, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0862-y.

Bosque:2013:LIL

[604] Jose Luis Bosque, Pablo Toharia, Oscar D. Robles, and Luis Pastor. A load index and load balancing algorithm for heterogeneous clusters. The Journal of Supercomputing, 65(3): 1104–1113, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0881-3.

Dufrechu:2013:ALL

[605] Ernesto Dufrechu, Pablo Ezzatti, Enrique S. Quintana-Ortí, and Alfredo Remón. Accelerating the Lyapack library using GPUs. The Journal of Supercomputing, 65(3):1114– 1124, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0889-8.

Acosta:2013:SBP

[606] Alejandro Acosta and Francisco Almeida. Skeletal based programming for dynamic programming on Multi-GPU systems. The Journal of Supercomputing, 65(3):1125–1136, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0895-x.

Almeida:2013:MEC

[607] F. Almeida, V. Blanco, A. Cabrera, and J. Ruiz. Modeling energy consumption for master-slave applications. The Journal of Supercomputing, 65(3): 1137–1149, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0914-y.

Torres:2013:UEI

[608] Yuri Torres, Arturo Gonzalez-Escribano, and Diego R. Llanos. uBench: exposing the impact of CUDA block geometry in terms of performance. The Journal of Supercomputing, 65(3): 1150–1163, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0921-z.

Alvarez-Bermejo:2013:HAM

[609] J. A. Álvarez-Bermejo, N. Antequera, and J. A. López-Ramos. Hierarchical approaches for multicast based on Euclid’s algorithm. The Journal of Supercomputing, 65(3):1164–1178, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0923-x.

Fernandez:2013:HMP

[610] Víctor Fernández, Juan M. Orduña, and Pedro Morillo. How mobile phones perform in collaborative augmented reality (CAR) applications. The Journal of Supercomputing, 65(3): 1179–1191, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0925-8.

Lee:2013:NNA

[611] Yang Sun Lee, Dong Kyoo Kim, and Leonard Barolli. Network numerical analysis for the smoother and the lagged joint-process estimator. The Journal of Supercomputing, 65(3): 1192–1204, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0753-2.

Fortin:2013:ADD

[612] Pierre Fortin and Jean-Luc Lamotte. An (almost) direct deployment of the Fast Multipole Method on the Cell processor. The Journal of Supercomputing, 65(3):1205–1222, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0877-z.

Vasupongayya:2013:EGO

[613] S. Vasupongayya and A. Prasitsupparote. Extending goal-oriented parallel computer job scheduling policies to heterogeneous systems. The Journal of Supercomputing, 65(3):1223– 1242, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0879-x.

Bushehrian:2013:DOS

[614] Omid Bushehrian and Reza Ghanbari Baghnavi. Deployment optimization of software objects by design-level delay estimation. The Journal of Supercomputing, 65(3):1243–1263, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0880-4.

Yang:2013:CDB

[615] Ming-Chien Yang. Conditional diagnosability of balanced hypercubes under the MM^{∗} model. The Journal of Supercomputing, 65(3):1264– 1278, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0882-2.

Cheng:2013:PCI

[616] Baolei Cheng, Jianxi Fan, Xiaohua Jia, and Juncheng Jia. Parallel construction of independent spanning trees and an application in diagnosis on Möbius cubes. The Journal of Supercomputing, 65(3):1279–1301, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0883-1.

Egwutuoha:2013:SFT

[617] Ifeanyi P. Egwutuoha, David Levy, Bran Selic, and Shiping Chen. A survey of fault tolerance mechanisms and checkpoint/restart implementations for high performance computing systems. The Journal of Supercomputing, 65(3): 1302–1326, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0884-0.

Jimenez:2013:BCA

[618] Jesús Jiménez and Juan Ruiz de Miras. Box-counting algorithm on GPU and multi-core CPU: an OpenCL crossplatform study. The Journal of Supercomputing, 65(3):1327–1352, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0885-z.

Bui:2013:NCI

[619] Alain Bui, Simon Clavière, and Devan Sohier. Nested clusters with intercluster routing. The Journal of Supercomputing, 65(3):1353–1382, September 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0886-y.

Park:2013:ATA

[620] Sang Oh Park, Bin Xiao, Victor Leung, and Young-Sik Jeong. Advanced technologies and applications for Highly–Reliable Cyber Physical System (HRCPS). The Journal of Supercomputing, 66(1):1–3, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        013-1027-3.pdf.

Desruelle:2013:MDA

[621] Heiko Desruelle, Simon Isenberg, John Lyle, and Frank Gielen. Multi-device application middleware: leveraging the ubiquity of the Web with webinos. The Journal of Supercomputing, 66(1):4–20, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0901-3.

Kim:2013:CLO

[622] Wooseong Kim, Joon-Sang Park, and Sanghyun Ahn. Cross-layer optimization for wireless multihop multicast networks. The Journal of Supercomputing, 66(1):21–34, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0929-4.

Huang:2013:HRB

[623] Chen-Che Huang, Tsun-Tse Huang, Jiun-Long Huang, and Lo-Yao Yeh. Hierarchical role-based data dissemination in wireless sensor networks. The Journal of Supercomputing, 66 (1):35–56, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0955-2.

Jeong:2013:ARM

[624] Young-Sik Jeong, Hyun-Woo Kim, and Haeng Jin Jang. Adaptive resource management scheme for monitoring of CPS. The Journal of Supercomputing, 66(1):57–69, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0970-3.

Kim:2013:ESN

[625] Yong hwan Kim, Chan-Myung Kim, Youn-Hee Han, Young-Sik Jeong, and Doo-Soon Park. An efficient strategy of nonuniform sensor deployment in cyber physical systems. The Journal of Supercomputing, 66(1):70–80, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0977-9.

Kwon:2013:RDA

[626] Se Jin Kwon, Hyung-Ju Cho, Sungsoo Kim, and Tae-Sun Chung. Random data-aware flash translation layer for NAND flash-based smart devices. The Journal of Supercomputing, 66 (1):81–93, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0979-7.

Zhang:2013:AFI

[627] Fan Zhang, Yan Zhang, and Jason D. Bakos. Accelerating frequent itemset mining on graphics processing units. The Journal of Supercomputing, 66 (1):94–117, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0887-x.

Wu:2013:EHT

[628] Jigang Wu, Pu Wang, Siew-Kei Lam, and Thambipillai Srikanthan. Efficient heuristic and tabu search for hardware/ software partitioning. The Journal of Supercomputing, 66(1):118–134, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0888-9.

Tabik:2013:OTO

[629] S. Tabik, A. Villegas, E. L. Zapata, and L. F. Romero. Optimal tilt and orientation maps: a multi-algorithm approach for heterogeneous multicore-GPU systems. The Journal of Supercomputing, 66(1):135–147, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0891-1.

Wu:2013:MSM

[630] Fan Wu, Kenli Li, Ahmed Sallam, and Xu Zhou. A molecular solution for minimum vertex cover problem in tile assembly model. The Journal of Supercomputing, 66(1):148–169, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0892-0.

Sharma:2013:NPS

[631] Bhanu Sharma, Ruppa K. Thulasiram, and Parimala Thulasiraman. Normalized particle swarm optimization for complex chooser option pricing on graphics processing unit. The Journal of Supercomputing, 66(1):170–192, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0893-z.

Sun:2013:AME

[632] Dawei Sun, Guiran Chang, Changsheng Miao, and Xingwei Wang. Analyzing, modeling and evaluating dynamic adaptive fault tolerance strategies in cloud computing environments. The Journal of Supercomputing, 66 (1):193–228, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0898-7.

Chung:2013:DAR

[633] Wu-Chun Chung, Chin-Jung Hsu, Kuan-Chou Lai, Kuan-Ching Li, and Yeh-Ching Chung. Direction-aware resource discovery in large-scale distributed computing environments. The Journal of Supercomputing, 66(1): 229–248, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0899-6.

Yuan:2013:DCF

[634] Fengkai Yuan and Zhenzhou Ji. DP&TB: a coherence filtering protocol for many-core chip multiprocessors. The Journal of Supercomputing, 66(1):249–261, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0900-4.

Aron:2013:QBR

[635] Rajni Aron and Inderveer Chana. QoS based resource provisioning and scheduling in grids. The Journal of Supercomputing, 66(1):262–283, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0903-1.

Villar:2013:ISQ

[636] Juan A. Villar, Pedro J. García, Francisco J. Alfaro, José L. Sánchez, and Francisco J. Quiles. An integrated solution for QoS provision and congestion management in high-performance interconnection networks using deterministic source-based routing. The Journal of Supercomputing, 66(1):284–304, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0904-0.

Huang:2013:EAL

[637] Tian Huang, Yongxin Zhu, Meikang Qiu, Xiaojing Yin, and Xu Wang. Extending Amdahl’s law and Gustafson’s law by evaluating interconnections on multi-core processors. The Journal of Supercomputing, 66(1):305–319, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0908-9.

Wu:2013:OMD

[638] Yahui Wu, Su Deng, and Hongbin Huang. Optimal management of dynamic information in delay tolerant networks. The Journal of Supercomputing, 66(1):320–338, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0909-8.

Mahafzah:2013:PAM

[639] Basel A. Mahafzah. Performance assessment of multithreaded quicksort algorithm on simultaneous multithreaded architecture. The Journal of Supercomputing, 66(1):339–363, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0910-2.

Chai:2013:REU

[640] Jun Chai, Huayou Su, Mei Wen, Xing Cai, and Nan Wu. Resource-efficient utilization of CPU/GPU-based heterogeneous supercomputers for Bayesian phylogenetic inference. The Journal of Supercomputing, 66(1):364–380, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0911-1.

Li:2013:COM

[641] Hung-Fu Li, Tyng-Yeu Liang, and Jun-Yao Chiu. A compound OpenMP/MPI program development toolkit for hybrid CPU/GPU clusters. The Journal of Supercomputing, 66(1):381–405, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0912-0.

Kunaseth:2013:ASD

[642] Manaschai Kunaseth, David F. Richards, and James N. Glosli. Analysis of scalable data-privatization threading algorithms for hybrid MPI/OpenMP parallelization of molecular dynamics. The Journal of Supercomputing, 66 (1):406–430, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0915-x.

Munir:2013:HPO

[643] Arslan Munir, Farinaz Koushanfar, and Ann Gordon-Ross. High-performance optimizations on tiled many-core embedded systems: a matrix multiplication case study. The Journal of Supercomputing, 66(1):431–487, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0916-9.

Soryani:2013:IIN

[644] Mohsen Soryani, Morteza Analoui, and Ghobad Zarrinchian. Improving internode communications in multi-core clusters using a contention-free process mapping algorithm. The Journal of Supercomputing, 66(1):488–513, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0918-7.

Wang:2013:TNS

[645] Xinyu Wang, Dong Xiang, and Zhigang Yu. TM: a new and simple topology for interconnection networks. The Journal of Supercomputing, 66 (1):514–538, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0922-y.

Slagter:2013:IPM

[646] Kenn Slagter, Ching-Hsien Hsu, and Yeh-Ching Chung. An improved partitioning mechanism for optimizing massive data analysis using MapReduce. The Journal of Supercomputing, 66 (1):539–555, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0924-9.

Shi:2013:SMC

[647] Zhiyuan Shi and Fei Hao. A strategy of multi-criteria decision-making task ranking in social-networks. The Journal of Supercomputing, 66(1): 556–571, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0934-7.

Azizi:2013:TPH

[648] Sadoon Azizi, Farshad Safaei, and Naser Hashemi. On the topological properties of HyperX. The Journal of Supercomputing, 66(1):572–593, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11227-013-
0935-6.

Myung:2013:EIO

[649] Jaeseok Myung and Sang goo Lee. Exploiting inter-operation parallelism for matrix chain multiplication using MapReduce. The Journal of Supercomputing, 66(1):594–609, October 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0936-5.

Kim:2013:SNH

[650] Jongsung Kim, Philip S. Yu, and Nasrullah Memon. Social network and high performance in smart communications. The Journal of Supercomputing, 66(2): 611–613, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-013-1029-1.pdf.

Park:2013:QBM

[651] James (Jong Huk) Park and Hwa-Young Jeong. The QoS-based MCDM system for SaaS ERP applications with Social Network. The Journal of Supercomputing, 66(2):614–632, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0832-4.

Sohn:2013:DFM

[652] Jong-Soo Sohn and In-Jeong Chung. Dynamic FOAF management method for social networks in the Social Web environment. The Journal of Supercomputing, 66(2):633–648, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0847-x.

Peng:2013:EEC

[653] Yuexing Peng, Yonghui Li, Lei Shu, and Wenbo Wang. An energy-efficient clustered distributed coding for largescale wireless sensor networks. The Journal of Supercomputing, 66(2): 649–669, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0848-9.

Chen:2013:URR

[654] Yen Wen Chen and Shaoh Chen Ke. A uplink radio resource allocation scheme for localized SC–FDMA transmission in LTE network. The Journal of Supercomputing, 66(2):670–685, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0849-8.

Wang:2013:DBA

[655] Xibin Wang, Xia Xie, Hai Jin, Xuanhua Shi, Wenzhi Cao, and Xijiang Ke. A disk bandwidth allocation mechanism with priority. The Journal of Supercomputing, 66(2):686–699, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0857-8.

Moon:2013:CAO

[656] Yong-Hyuk Moon, Jeong-Nyeo Kim, and Chan-Hyun Youn. Churn-aware optimal layer scheduling scheme for scalable video distribution in superpeer overlay networks. The Journal of Supercomputing, 66(2):700– 720, November 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0858-7.

Juang:2013:EAT

[657] Ming-Chin Juang, Chen-Che Huang, and Jiun-Long Huang. Efficient algorithms for team formation with a leader in social networks. The Journal of Supercomputing, 66(2):721–737, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0907-x.

Fatima:2013:MMO

[658] Iram Fatima, Muhammad Fahim, Young-Koo Lee, and Sungyoung Lee. MODM: multi-objective diffusion model for dynamic social networks using evolutionary algorithm. The Journal of Supercomputing, 66(2): 738–759, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0947-2.

Fatima:2013:AES

[659] Iram Fatima, Muhammad Fahim, Young-Koo Lee, and Sungyoung Lee. Analysis and effects of smart home dataset characteristics for daily life activity recognition. The Journal of Supercomputing, 66(2):760–780, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0978-8.

Yeo:2013:ESS

[660] Sangsoo Yeo and Shiuh-Jeng Wang. Editorial of special section on hybrid information security technologies — Part I. The Journal of Supercomputing, 66(2):781–782, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
013-1030-8.pdf.

Wang:2013:ASP

[661] Wei-Jen Wang, Yue-Shan Chang, Win-Tsung Lo, and Yi-Kang Lee. Adaptive scheduling for parallel tasks with QoS satisfaction for hybrid cloud environments. The Journal of Supercomputing, 66(2):783–811, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0890-2.

Hsu:2013:RIB

[662] Fu-Hau Hsu, Min-Hao Wu, Shiuh-Jeng Wang, and Chia-Ling Huang. Reversibility of image with balanced fidelity and capacity upon pixels differencing expansion. The Journal of Supercomputing, 66(2):812–828, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11227-013-
                                        0896-9.

Jeong:2013:CBC

[663] Kitae Jeong. Cryptanalysis of block cipher Piccolo suitable for cloud computing. The Journal of Supercomputing, 66 (2):829–840, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0902-2.

Fan:2013:DIP

[664] Chun-I Fan, Shi-Yuan Huang, and William Artan. Design and implementation of privacy preserving billing protocol for smart grid. The Journal of Supercomputing, 66(2):841–862, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0905-z.

Tso:2013:SAI

[665] Raylin Tso. Security analysis and improvements of a communicationefficient three-party password authenticated key exchange protocol. The Journal of Supercomputing, 66(2): 863–874, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0917-8.

Bang:2013:DBD

[666] Jewan Bang, Changhoon Lee, Sangjin Lee, and Kyungho Lee. Damaged backup data recovery method for Windows mobile. The Journal of Supercomputing, 66(2):875–887, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0919-6.

Hwang:2013:ADR

[667] Ren-Junn Hwang and Yu-Kai Hsiao. An anonymous distributed routing protocol in mobile ad-hoc networks. The Journal of Supercomputing, 66(2): 888–906, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0920-0.

Chen:2013:RPP

[668] Tung-Shou Chen, Wei-Bin Lee, Jeanne Chen, Yuan-Hung Kao, and Pei-Wen Hou. Reversible privacy preserving data mining: a combination of difference expansion and privacy preserving. The Journal of Supercomputing, 66 (2):907–917, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0926-7.

Lim:2013:NAC

[669] Jungmin Lim, Yongki Kim, Dongyoung Koo, Soojin Lee, Seokjoo Doo, and Hyunsoo Yoon. A novel Adaptive Cluster Transformation (ACT)- based intrusion tolerant architecture for hybrid information technology. The Journal of Supercomputing, 66(2): 918–935, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0928-5.

Han:2013:RCS

[670] Tae Youn Han and Mun-Kyu Lee. Reordering computation sequences for memory-efficient binary field multiplication. The Journal of Supercomputing, 66(2):936–949, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0930-y.

Choi:2013:IDS

[671] Hyo Hyun Choi, Su Hyun Nam, Taeshik Shon, and Myungwhan Choi. Information delivery scheme of micro UAVs having limited communication range during tracking the moving target. The Journal of Supercomputing, 66 (2):950–972, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0931-x.

Chou:2013:TIB

[672] Chih-Ho Chou, Kuo-Yu Tsai, and Chung-Fu Lu. Two ID-based authenticated schemes with key agreement for mobile environments. The Journal of Supercomputing, 66(2): 973–988, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0962-3.

Lin:2013:EAE

[673] Chi Lin and Guowei Wu. Enhancing the attacking efficiency of the node capture attack in WSN: a matrix approach. The Journal of Supercomputing, 66(2):989–1007, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0965-0.

Chen:2013:TSE

[674] Te-Yu Chen, Cheng-Chi Lee, Min-Shiang Hwang, and Jinn-Ke Jan. Towards secure and efficient user authentication scheme using smart card for multi-server environments. The Journal of Supercomputing, 66(2): 1008–1032, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0966-z.

Weng:2013:VWI

[675] Chi-Yao Weng, Yu Hong Zhang, Li Chun Lin, and Shiuh-Jeng Wang. Visible watermarking images in high quality of data hiding. The Journal of Supercomputing, 66(2):1033– 1048, November 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0969-9.

Nourian:2013:PAI

[676] Arash Nourian and Muthucumaru Maheswaran. Privacy aware image template matching in clouds using ambient data. The Journal of Supercomputing, 66(2):1049–1070, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0976-x.

Vaidya:2013:SRM

[677] Binod Vaidya, Dimitrios Makrakis, and Hussein Mouftah. Secure and robust multipath routings for advanced metering infrastructure. The Journal of Supercomputing, 66(2):1071– 1092, November 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1009-5.

Chang:2013:IIB

[678] Ya-Ting Chang, Cheng-Ta Huang, and Chin-Feng Lee. Image interpolating based data hiding in conjunction with pixel-shifting of histogram. The Journal of Supercomputing, 66(2): 1093–1110, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1016-6.

Chen:2013:CRC

[679] Hsing-Chung (Jack) Chen and Marsha Anjanette Violetta. Contract RBAC in cloud computing. The Journal of Supercomputing, 66(2): 1111–1131, November 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1017-5.

Jin:2013:VBI

[680] Hai Jin, Guofu Xiang, Deqing Zou, Song Wu, Feng Zhao, Min Li, and Weide Zheng. A VMM-based intrusion prevention system in cloud computing environment. The Journal of Supercomputing, 66(3):1133–1151, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0608-2.

Shi:2013:AGC

[681] Xuanhua Shi, Hai Jin, Song Wu, Wei Zhu, and Li Qi. Adapting grid computing environments dependable with virtual machines: design, implementation, and evaluations. The Journal of Supercomputing, 66(3):1152–1166, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0664-7.

Zhong:2013:NCR

[682] Ming Zhong, Yaoxue Zhang, Yuezhi Zhou, Laurence Tianruo Yang, Pengwei Tian, and Linkai Weng. A novel component retrieval method based on weighted facet tree. The Journal of Supercomputing, 66(3):1167–1177, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-011-
0665-6.

Wang:2013:DSH

[683] Lizhe Wang, Tobias Kurze, Jie Tao, Marcel Kunze, and Gregor von Laszewski. On-demand service hosting on production grid infrastructures. The Journal of Supercomputing, 66(3): 1178–1193, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0666-5.

Gil:2013:DCS

[684] Joon-Min Gil, Jong Hyuk Park, and Young-Sik Jeong. Data center selection based on neuro-fuzzy inference systems in cloud computing environments. The Journal of Supercomputing, 66(3): 1194–1214, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0667-4.

Cao:2013:PNB

[685] Haijun Cao, Hai Jin, Song Wu, and Shadi Ibrahim. Petri net based Grid workflow verification and optimization. The Journal of Supercomputing, 66(3): 1215–1230, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-011-0668-3.

Allenotor:2013:FGQ

[686] David Allenotor and Ruppa K. Thulasiram. A fuzzy Grid–QoS framework for obtaining higher grid resources availability. The Journal of Supercomputing, 66(3):1231–1242, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-011-
                                        0728-8.

Tan:2013:IPA

[687] Zhiyong Tan, Duo Liu, Xuejun Zhuo, Yiqi Dai, and Laurence T. Yang. Implementation and performance analysis of multilevel security system in pervasive computing environment. The Journal of Supercomputing, 66(3): 1243–1259, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-011-0732-z.

Ding:2013:SDC

[688] Zhiming Ding, Jiajie Xu, and Qi Yang. SeaCloudDM: a database cluster framework for managing and querying massive heterogeneous sensor sampling data. The Journal of Supercomputing, 66(3):1260–1284, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0762-1.

Yang:2013:HBT

[689] Ming Yang, Junzhou Luo, Lu Zhang, Xiaogang Wang, and Xinwen Fu. How to block Tor’s hidden bridges: detecting methods and countermeasures. The Journal of Supercomputing, 66(3): 1285–1305, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0788-4.

Liu:2013:SMB

[690] Jiang Liu, Mianxiong Dong, Laurence T. Yang, and Shigeru Shimamoto. A subcarrier modulation based optical wireless communication system employing transmission diversity. The Journal of Supercomputing, 66(3):1306–1319, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-012-
0795-5.

Wanalertlak:2013:SFH

[691] Weetit Wanalertlak, Ben Lee, Chansu Yu, Myungchul Kim, Seung-Min Park, and Won-Tae Kim. Scanless fast handoff technique based on global Path– Cache for WLANs. The Journal of Supercomputing, 66(3):1320–1349, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0805-7.

Adabi:2013:NSC

[692] Sepideh Adabi, Ali Movaghar, Amir Masoud Rahmani, and Hamid Beigy. Negotiation strategies considering market, time and behavior functions for resource allocation in computational grid. The Journal of Supercomputing, 66(3):1350–1389, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-012-
                                        0808-4.

Park:2013:CSR

[693] Joon S. Park, Pratheep Chandramohan, Avinash T. Suresh, Joseph V. Giordano, and Kevin A. Kwiat. Component survivability at runtime for mission-critical distributed systems. The Journal of Supercomputing, 66(3): 1390–1417, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-012-0818-2.

Guan:2013:SSA

[694] Hu Guan, Huakang Li, Cheng-Zhong Xu, and Minyi Guo. Semi-sparse algorithm based on multi-layer optimization for recommender system. The Journal of Supercomputing, 66(3): 1418–1437, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0830-6.

Cano:2013:HPE

[695] Alberto Cano, José María Luna, and Sebastián Ventura. High performance evaluation of evolutionary-mined association rules on GPUs. The Journal of Supercomputing, 66(3):1438–1461, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0937-4.

Cao:2013:DWM

[696] Fei Cao and Michelle M. Zhu. Distributed workflow mapping algorithm for maximized reliability under end-to-end delay constraint. The Journal of Supercomputing, 66(3):1462–1488, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0938-3.

Kim:2013:VMC

[697] Shin gyu Kim, Hyeonsang Eom, and Heon Y. Yeom. Virtual machine consolidation based on interference modeling. The Journal of Supercomputing, 66(3): 1489–1506, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0939-2.

Ahmed:2013:ADH

[698] Akram Ben Ahmed and Abderazek Ben Abdallah. Architecture and design of high-throughput, low-latency, and fault-tolerant routing algorithm for 3D-network-on-chip (3D-NoC). The Journal of Supercomputing, 66(3): 1507–1532, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0940-9.

Mohamadi:2013:LAB

[699] Hosein Mohamadi, Abdul Samad Ismail, Shaharuddin Salleh, and Ali Nodhei. Learning automata-based algorithms for finding cover sets in wireless sensor networks. The Journal of Supercomputing, 66(3):1533–1552, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0941-8.

Chen:2013:SSR

[700] Yunliang Chen, Dan Chen, Samee U. Khan, Jianzhong Huang, and Changsheng Xie. Solving symbolic regression problems with uniform designaided gene expression programming. The Journal of Supercomputing, 66(3): 1553–1575, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0943-6.

Gonzalez-Alvarez:2013:PCT

[701] David L. González-Álvarez and Miguel A. Vega-Rodríguez. A parallel cooperative team of multiobjective evolutionary algorithms for motif discovery. The Journal of Supercomputing, 66(3): 1576–1612, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0951-6.

Lee:2013:LWK

[702] Dongwoo Lee, Inhyuk Kim, Jeehong Kim, Hyung Kook Jun, Won Tae Kim, Sangwon Lee, and Young Ik Eom. Light-weight kernel instrumentation framework using dynamic binary translation. The Journal of Supercomputing, 66(3):1613–1628, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0954-3.

Jeong:2013:AVM

[703] Jinkyu Jeong, Sung-Hun Kim, Hwanju Kim, Joonwon Lee, and Euiseong Seo. Analysis of virtual machine livemigration as a method for powercapping. The Journal of Supercomputing, 66(3):1629–1655, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0956-1.

Gao:2013:MSP

[704] Guoqiang Gao, Ruixuan Li, Weijun Xiao, and Zhiyong Xu. Measurement study on P2P streaming systems. The Journal of Supercomputing, 66(3): 1656–1686, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0959-y.

Khan:2013:EDC

[705] Abdul Nasir Khan, M. L. Mat Kiah, Sajjad A. Madani, Atta ur Rehman Khan, and Mazhar Ali. Enhanced dynamic credential generation scheme for protection of user identity in mobile-cloud computing. The Journal of Supercomputing, 66(3): 1687–1706, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0967-y.

Amir:2013:ICP

[706] Hossein Amir and Hadi Shahriar Shahhoseini. Improving CompactMatrix phase in gang scheduling by changing transference condition and utilizing exchange. The Journal of Supercomputing, 66(3):1707–1728, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0971-2.

Song:2013:OHR

[707] Biao Song, Wei Tang, Tien-Dung Nguyen, Mohammad Mehedi Hassan, and Eui Nam Huh. An optimized hybrid remote display protocol using GPU-assisted M-JPEG encoding and novel high-motion detection algorithm. The Journal of Supercomputing, 66(3): 1729–1748, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0972-1.

Barthwal:2013:FOC

[708] Romil Barthwal, Sudip Misra, and Mohammad S. Obaidat. Finding overlapping communities in a complex network of social linkages and Internet of Things. The Journal of Supercomputing, 66(3):1749–1772, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0973-0.

Tian:2013:OPS

[709] Wenhong Tian, Qin Xiong, and Jun Cao. An online parallel scheduling method with application to energyefficiency in cloud computing. The Journal of Supercomputing, 66(3): 1773–1790, December 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0974-z.

Khaneghah:2014:AAM

[710] Ehsan Mousavi Khaneghah and Mohsen Sharifi. AMRC: an algebraic model for reconfiguration of high performance cluster computing systems at runtime. The Journal of Supercomputing, 67(1):1–30, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0982-z.

Shahhoseini:2014:NSL

[711] Hadi Shahriar Shahhoseini, Ehsan Saleh Kandzi, and Morteza Mollajafari. Nonflat surface level pyramid: a high connectivity multidimensional interconnection network. The Journal of Supercomputing, 67(1):31–46, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0983-y.

Chandar:2014:COO

[712] Dominic D. J. Chandar, Jayanarayanan Sitaraman, and Dimitri Mavriplis. CU++: an object oriented framework for computational fluid dynamics applications using graphics processing units. The Journal of Supercomputing, 67(1):47–68, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0985-9.

Ziaee:2014:HAD

[713] Mohsen Ziaee. A heuristic algorithm for the distributed and flexible jobshop scheduling problem. The Journal of Supercomputing, 67(1):69–83, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0986-8.

Shiraz:2014:IRP

[714] Muhammad Shiraz, Ejaz Ahmed, Abdullah Gani, and Qi Han. Investigation on runtime partitioning of elastic mobile applications for mobile cloud computing. The Journal of Supercomputing, 67(1):84–103, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0988-6.

Azarpeyvand:2014:AMR

[715] Ali Azarpeyvand, Mostafa E. Salehi, and Sied Mehdi Fakhraie. An analytical method for reliability aware instruction set extension. The Journal of Supercomputing, 67(1):104–130, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0990-z.

Reina:2014:IDP

[716] D. G. Reina, S. L. Toral, P. Johnson, and F. Barrero. Improving discovery phase of reactive ad hoc routing protocols using Jaccard distance. The Journal of Supercomputing, 67 (1):131–152, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0992-x.

Wang:2014:OPC

[717] Haifeng Wang and Qingkui Chen. Optimization power consumption model of reliability-aware GPU clusters. The Journal of Supercomputing, 67(1): 153–174, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0993-9.

Adabi:2014:BLF

[718] Sahar Adabi, Ali Movaghar, and Amir Masoud Rahmani. Bi-level fuzzy based advanced reservation of Cloud workflow applications on distributed Grid resources. The Journal of Supercomputing, 67(1):175–218, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0994-8.

Ergu:2014:FSS

[719] Daji Ergu and Yi Peng. A framework for SaaS software packages evaluation and selection with virtual team and BOCR of analytic network process. The Journal of Supercomputing, 67(1):219–238, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0995-7.

Touzene:2014:NPA

[720] Abderezak Touzene. A new parallel algorithm for solving large-scale Markov chains. The Journal of Supercomputing, 67(1):239–253, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0997-5.

Yu:2014:MPP

[721] Chao Yu, Leihua Qin, and Jingli Zhou. A multicore periodical preemption virtual machine scheduling scheme to improve the performance of computational tasks. The Journal of Supercomputing, 67(1):254–276, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0998-4.

Kim:2014:HCP

[722] Youngjae Kim, Aayush Gupta, Bhuvan Urgaonkar, Piotr Berman, and Anand Sivasubramaniam. HybridPlan: a capacity planning technique for projecting storage requirements in hybrid storage systems. The Journal of Supercomputing, 67(1):277–303, January 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0999-3.

Chang:2014:SFB

[723] Kuei-Chung Chang, Chen-Yu Chen, Chin-Sheng Yu, and Ching-Wen Chen. Supporting faulty banks in NUCA by NoC assisted remapping mechanisms. The Journal of Supercomputing, 67 (2):305–323, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1001-0.

Okuyan:2014:DVR

[724] Erhan Okuyan and Ugur Güdükbay. Direct volume rendering of unstructured tetrahedral meshes using CUDA and OpenMP. The Journal of Supercomputing, 67(2):324–344, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1004-x.

Rezaee:2014:FPA

[725] Ali Rezaee, Amir Masoud Rahmani, Ali Movaghar, and Mohammad Teshnehlab. Formal process algebraic modeling, verification, and analysis of an abstract Fuzzy Inference Cloud Service. The Journal of Supercomputing, 67 (2):345–383, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1005-9.

Filelis-Papadopoulos:2014:PMA

[726] Christos K. Filelis-Papadopoulos and George A. Gravvanis. Parallel multigrid algorithms based on generic approximate sparse inverses: an SMP approach. The Journal of Supercomputing, 67(2):384–407, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        1006-8.

Choi:2014:PLM

[727] Hyebong Choi, Kyong-Ha Lee, and Yoon-Joon Lee. Parallel labeling of massive XML data with MapReduce. The Journal of Supercomputing, 67 (2):408–437, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1008-6.

Ali:2014:SDA

[728] G. G. Md. Nawaz Ali, Edward Chan, and Wenzhong Li. On scheduling data access with cooperative load balancing in vehicular ad hoc networks (VANETs). The Journal of Supercomputing, 67(2):438–468, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        1011-y.

Park:2014:QBM

[729] Sung-Hoon Park and Seon-Hyong Lee. Quorum-based mutual exclusion in asynchronous distributed systems with unreliable failure detectors. The Journal of Supercomputing, 67 (2):469–484, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1012-x.

Bossard:2014:PDP

[730] Antoine Bossard and Keiichi Kaneko. k-pairwise disjoint paths routing in perfect hierarchical hypercubes. The Journal of Supercomputing, 67(2): 485–495, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1013-9.

Kianpisheh:2014:GWQ

[731] Somayeh Kianpisheh and Nasrolah Moghadam Charkari. A grid workflow Quality-of-Service estimation based on resource availability prediction. The Journal of Supercomputing, 67(2):496–527, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1014-8.

Niemeyer:2014:RPC

[732] Kyle E. Niemeyer and Chih-Jen Sung. Recent progress and challenges in exploiting graphics processors in computational fluid dynamics. The Journal of Supercomputing, 67(2):528–564, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1015-7.

Savadi:2014:MLP

[733] Abdorreza Savadi and Hossein Deldari. Measurement of the latency parameters of the multi-BSP model: a multicore benchmarking approach. The Journal of Supercomputing, 67 (2):565–584, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1018-4.

Chopra:2014:SAS

[734] Inderpreet Chopra and Maninder Singh. SHAPE — an approach for selfhealing and self-protection in complex distributed networks. The Journal of Supercomputing, 67(2):585–613, February 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        1019-3.

Park:2014:FRI

[735] Jeong Soo Park, Ki Seok Bae, Yong Je Choi, and Doo Ho Choi. A faultresistant implementation of AES using differential bytes between input and output. The Journal of Supercomputing, 67(3):615–634, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0950-7.

Huang:2014:SWC

[736] Yi-Li Huang, Fang-Yie Leu, Ilsun You, and Yao-Kuo Sun. A secure wireless communication system integrating RSA, Diffie–Hellman PKDS, intelligent protection-key chains and a Data Connection Core in a 4G environment. The Journal of Supercomputing, 67(3):635–652, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0958-z.

Xu:2014:NQN

[737] Guangwei Xu, Cheng Shao, Yu Han, and Kangbin Yim. New quasi-Newton iterative learning control scheme based on rank-one update for nonlinear systems. The Journal of Supercomputing, 67(3):653–670, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-0960-5.

Tian:2014:DFS

[738] Haibo Tian, Xiaofeng Chen, and Willy Susilo. Deniability and forward secrecy of one-round authenticated key exchange. The Journal of Supercomputing, 67(3):671–690, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0968-x.

Castiglione:2014:SFS

[739] Aniello Castiglione, Luigi Catuogno, and Aniello Del Sorbo. A secure file sharing service for distributed computing environments. The Journal of Supercomputing, 67(3):691–710, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0975-y.

Choi:2014:OBA

[740] Chang Choi, Junho Choi, and Pankoo Kim. Ontology-based access control model for security policy reasoning in cloud computing. The Journal of Supercomputing, 67(3):711–722, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0980-1.

Liu:2014:NCA

[741] Yuhua Liu, Jianzhi Jin, Yi Zhang, and Cui Xu. A new clustering algorithm based on data field in complex networks. The Journal of Supercomputing, 67(3):723–737, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0984-x.

Lee:2014:USE

[742] Chanhee Lee, Jonghwa Kim, Seong je Cho, and Jongmoo Choi. Unified security enhancement framework for the Android operating system. The Journal of Supercomputing, 67(3):738–756, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0991-y.

Lee:2014:GBR

[743] Jang Ho Lee. Grid-based recording and replay architecture in hybrid remote experiment using distributed streaming network. The Journal of Supercomputing, 67(3):757–777, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0996-6.

Liu:2014:TPA

[744] Bin Liu, Yinliang Zhao, Yuxiang Li, Yanjun Sun, and Boqin Feng. A thread partitioning approach for speculative multithreading. The Journal of Supercomputing, 67(3):778–805, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1000-1.

Jiao:2014:ECA

[745] Zhu-Qing Jiao, Ling Zou, Yin Cao, Nong Qian, and Zheng-Hua Ma. Effective connectivity analysis of fMRI data based on network motifs. The Journal of Supercomputing, 67(3):806–819, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1010-z.

Kang:2014:SSD

[746] Ho-Seok Kang and Sung-Ryul Kim. sShield: small DDoS defense system using RIP-based traffic deflection in autonomous system. The Journal of Supercomputing, 67(3):820–836, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1031-7.

Diaz:2014:SLC

[747] Cesar O. Diaz, Johnatan E. Pecero, and Pascal Bouvry. Scalable, low complexity, and fast greedy scheduling heuristics for highly heterogeneous distributed computing systems. The Journal of Supercomputing, 67(3):837–853, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1038-0.

Li:2014:IEF

[748] Chenghua Li, Man Lin, Laurence T. Yang, and Chen Ding. Integrating the enriched feature with machine learning algorithms for human movement and fall detection. The Journal of Supercomputing, 67(3):854–865, March 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1056-y.

Rawat:2014:WSN

[749] Priyanka Rawat, Kamal Deep Singh, Hakima Chaouchi, and Jean Marie Bonnin. Wireless sensor networks: a survey on recent developments and potential synergies. The Journal of Supercomputing, 68(1):1–48, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1021-9.

Mo:2014:NPA

[750] Zeyao Mo, Aiqing Zhang, and Zhang Yang. A new parallel algorithm for vertex priorities of data flow acyclic digraphs. The Journal of Supercomputing, 68(1):49–64, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1022-8.

Thompson:2014:CIC

[751] Elizabeth A. Thompson and Timothy R. Anderson. A CUDA implementation of the Continuous Space Language Model. The Journal of Supercomputing, 68(1):65–86, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1023-7.

Kuila:2014:ASL

[752] Pratyay Kuila and Prasanta K. Jana. Approximation schemes for load balanced clustering in wireless sensor networks. The Journal of Supercomputing, 68(1):87–105, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1024-6.

Reza:2014:LAS

[753] Akram Reza, Hamid Sarbazi-Azad, Ahmad Khademzadeh, Hesam Shabani, and Behrad Niazmand. A loss aware scalable topology for photonic on chip interconnection networks. The Journal of Supercomputing, 68 (1):106–135, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1026-4.

Lee:2014:VBD

[754] Yong-Ju Lee. Video block device for user-friendly delivery in IaaS clouds. The Journal of Supercomputing, 68 (1):136–156, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1032-6.

Foglia:2014:WIE

[755] Pierfrancesco Foglia and Manuel Comparetti. A workload independent energy reduction strategy for D-NUCA caches. The Journal of Supercomputing, 68(1):157–182, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        1033-5.

Yang:2014:IGV

[756] Chao-Tung Yang, Jung-Chun Liu, Hsien-Yi Wang, and Ching-Hsien Hsu. Implementation of GPU virtualization using PCI pass-through mechanism. The Journal of Supercomputing, 68(1):183–213, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1034-4.

Heydarian:2014:DDU

[757] Mohsen Heydarian. Dynamic distributed unicast routing: optimal incremental paths. The Journal of Supercomputing, 68(1):214–244, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        1035-3.

Tu:2014:DRP

[758] Manghui Tu and I-Ling Yen. Distributed replica placement algorithms for correlated data. The Journal of Supercomputing, 68(1):245–273, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1036-2.

Amin:2014:S

[759] Muhammad Bilal Amin, Rabia Batool, Wajahat Ali Khan, Sungyoung Lee, and Eui-Nam Huh. SPHeRe. The Journal of Supercomputing, 68(1):274–301, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1037-1.

Lu:2014:HHL

[760] Huazhong Lü and Heping Zhang. Hyper-Hamiltonian laceability of balanced hypercubes. The Journal of Supercomputing, 68(1):302–314, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1040-6.

Borovska:2014:AFC

[761] Plamenka Borovska and Dragi Kimovski. Adaptive flow control in high-performance interconnection networks. The Journal of Supercomputing, 68(1):315–338, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1041-5.

Alvarado:2014:MIS

[762] Rigo Alvarado, Juan J. Tapia, and Julio C. Rolón. Medical image segmentation with deformable models on graphics processing units. The Journal of Supercomputing, 68(1):339–364, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1042-4.

Carcenac:2014:TAS

[763] Manuel Carcenac. From tile algorithm to stripe algorithm: a CUBLAS-based parallel implementation on GPUs of Gauss method for the resolution of extremely large dense linear systems stored on an array of solid state devices. The Journal of Supercomputing, 68(1):365–413, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1043-3.

Cebrian:2014:MPC

[764] Juan M. Cebrián, Daniel Sánchez, Juan L. Aragón, and Stefanos Kaxiras. Managing power constraints in a single-core scenario through power tokens. The Journal of Supercomputing, 68(1):414–442, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1044-2.

Leu:2014:PRS

[765] Jenq-Shiou Leu, Wei-Hsiang Lin, and Jheng-Huei Chen. P2P resource searching with Cloning Random Walker assisted by Weakly Connected Dominating Set. The Journal of Supercomputing, 68(1):443–458, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1046-0.

Kelefouras:2014:MSE

[766] Vasilios Kelefouras, Angeliki Kritikakou, and Costas Goutis. A methodology for speeding up edge and line detection algorithms focusing on memory architecture utilization. The Journal of Supercomputing, 68(1):459–487, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1049-x.

Lai:2014:TFL

[767] Wei Kuang Lai, Yi-Uan Chen, Tin-Yu Wu, and Mohammad S. Obaidat. Towards a framework for large-scale multimedia data storage and processing on Hadoop platform. The Journal of Supercomputing, 68(1):488–507, April 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1050-4.

Somasundaram:2014:SEC

[768] Thamarai Selvi Somasundaram, Kannan Govindarajan, Usha Kiruthika, and Rajkumar Buyya. Semanticenabled CARE Resource Broker (Se-CRB) for managing grid and cloud environment. The Journal of Supercomputing, 68(2):509–556, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1047-z.

Sentis:2014:DEP

[769] Josep M. Sentís, Francesc Solsona, Damià Castellà, and Josep Rius. Dis-CoP2P: an efficient P2P computing overlay. The Journal of Supercomputing, 68(2):557–573, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1052-2.

Abdeyazdan:2014:DCB

[770] Marjan Abdeyazdan. Data clustering based on hybrid K-harmonic means and modifier imperialist competitive algorithm. The Journal of Supercomputing, 68(2):574–598, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1053-1.

Ahmadi:2014:ERA

[771] Ali Ahmadi, Mohammad Shojafar, Seyede Fatemeh Hajeforosh, Mehdi Dehghan, and Mukesh Singhal. An efficient routing algorithm to preserve kcoverage in wireless sensor networks. The Journal of Supercomputing, 68 (2):599–623, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1054-0.

Khan:2014:IPR

[772] Abdul Nasir Khan, M. L. Mat Kiah, Sajjad A. Madani, Mazhar Ali, Atta ur Rehman Khan, and Shahaboddin Shamshirband. Incremental proxy reencryption scheme for mobile cloud computing environment. The Journal of Supercomputing, 68(2):624–651, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1055-z.

Xiong:2014:NSM

[773] Huanliang Xiong, Guosun Zeng, Yuan Zeng, Wei Wang, and Canghai Wu. A novel scalability metric about isoarea of performance for parallel computing. The Journal of Supercomputing, 68(2):652–671, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1057-x.

Toumi:2014:PSO

[774] Lyazid Toumi, Abdelouahab Moussaoui, and Ahmet Ugur. Particle swarm optimization for bitmap join indexes selection problem in data warehouses. The Journal of Supercomputing, 68(2):672–708, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1058-9.

Garg:2014:MOW

[775] Ritu Garg and Awadhesh Kumar Singh. Multi-objective workflow grid scheduling using ε-fuzzy dominance sort based discrete particle swarm optimization. The Journal of Supercomputing, 68(2):709–732, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1059-8.

Zhao:2014:PED

[776] Yue Zhao, Kenji Yoshigoe, and Mengjun Xie. Pre-execution data prefetching with I/O scheduling. The Journal of Supercomputing, 68(2): 733–752, May 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1060-2.

Khan:2014:PFA

[777] Atta ur Rehman Khan, Mazliza Othman, Mazhar Ali, Abdul Nasir Khan, and Sajjad Ahmad Madani. Pirax: framework for application piracy control in mobile cloud environment. The Journal of Supercomputing, 68 (2):753–776, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1061-1.

Sundriyal:2014:ARF

[778] Vaibhav Sundriyal, Masha Sosonkina, and Zhao Zhang. Automatic runtime frequency-scaling system for energy savings in parallel applications. The Journal of Supercomputing, 68 (2):777–797, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1062-0.

Li:2014:RAA

[779] Bo Li, Yijian Pei, Hao Wu, and Bin Shen. Resource availabilityaware advance reservation for parallel jobs with deadlines. The Journal of Supercomputing, 68(2):798–819, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1067-8.

Zhu:2014:ENB

[780] Lei Zhu, Hai Jin, Ran Zheng, and Xiaowen Feng. Effective naive Bayes nearest neighbor based image classification on GPU. The Journal of Supercomputing, 68(2):820–848, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1068-7.

Yan:2014:NDC

[781] Jili Yan, Guoming Lai, and Xiaola Lin. A novel distributed congestion control for bufferless network-on-chip. The Journal of Supercomputing, 68(2):849–866, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1069-6.

Terzopoulos:2014:EER

[782] George Terzopoulos and Helen Karatza. Energy-efficient real-time heterogeneous cluster scheduling with node replacement due to failures. The Journal of Supercomputing, 68(2):867–889, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1070-0.

Lee:2014:DMA

[783] Jaehwan Lee, Pete Keleher, and Alan Sussman. Decentralized multiattribute range search for resource discovery and load balancing. The Journal of Supercomputing, 68(2):890–913, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1071-z.

Gaona:2014:SDS

[784] Epifanio Gaona, J. Rubén Titos-Gil, Juan Fernández, and Manuel E. Acacio. Selective dynamic serialization for reducing energy consumption in hardware transactional memory systems. The Journal of Supercomputing, 68(2):914–934, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1072-y.

Huang:2014:AIM

[785] Jian Huang, Mei Yang, Jian guo Hao, and Jianxing Gong. An algorithm for interest management in High Level Architecture. The Journal of Supercomputing, 68(2):935–947, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1074-9.

Yazdanbakhsh:2014:CPI

[786] Amir Yazdanbakhsh, Mostafa E. Salehi, and Sied Mehdi Fakhraie. Customized pipeline and instruction set architecture for embedded processing engines. The Journal of Supercomputing, 68(2):948–977, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1075-8.

Shiraz:2014:LAS

[787] Muhammad Shiraz and Abdullah Gani. A lightweight active service migration framework for computational offloading in mobile cloud computing. The Journal of Supercomputing, 68 (2):978–995, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1076-7.

Li:2014:EDF

[788] Xiuqiao Li, Limin Xiao, Meikang Qiu, Bin Dong, and Li Ruan. Enabling dynamic file I/O path selection at runtime for parallel file system. The Journal of Supercomputing, 68(2):996–1021, May 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1077-6.

Rahmani:2014:SSA

[789] Amir-Mohammad Rahmani, Pasi Liljeberg, Juha Plosila, and Hannu Tenhunen. Special section on advances in methods for adaptive multicore systems. The Journal of Supercomputing, 68(3):1023–1026, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-014-1217-7.pdf.

Ansari:2014:WAC

[790] Mohammad Ansari. Weighted adaptive concurrency control for software transactional memory. The Journal of Supercomputing, 68(3):1027–1047, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1138-5.

Maghsoudloo:2014:CVM

[791] Mohammad Maghsoudloo and Hamid R. Zarandi. Cache vulnerability mitigation using an adaptive cache coherence protocol. The Journal of Supercomputing, 68(3):1048–1067, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1139-4.

Salami:2014:PTM

[792] Bagher Salami, Mohammadreza Baharani, and Hamid Noori. Proactive task migration with a self-adjusting migration threshold for dynamic thermal management of multi-core processors. The Journal of Supercomputing, 68(3):1068–1087, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1140-y.

Piga:2014:AGP

[793] Leonardo Piga, Reinaldo A. Bergamaschi, Mauricio Breternitz, and Sandro Rigo. Adaptive global power optimization for Web servers. The Journal of Supercomputing, 68(3):1088–1112, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1141-x.

Utrera:2014:SPJ

[794] Gladys Utrera, Julita Corbalan, and Jesús Labarta. Scheduling parallel jobs on multicore clusters using CPU oversubscription. The Journal of Supercomputing, 68(3):1113–1140, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1142-9.

Holmbacka:2014:TMM

[795] Simon Holmbacka, Mohammad Fattah, Wictor Lund, Amir-Mohammad Rahmani, Sébastien Lafond, and Johan Lilius. A task migration mechanism for distributed many-core operating systems. The Journal of Supercomputing, 68(3):1141–1162, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1144-7.

Marowka:2014:MES

[796] Ami Marowka. Maximizing energy saving of dual-architecture processors using DVFS. The Journal of Supercomputing, 68(3):1163–1183, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1147-4.

Moore:2014:BUA

[797] Ryan W. Moore and Bruce R. Childers. Building and using application utility models to dynamically choose thread counts. The Journal of Supercomputing, 68(3):1184–1213, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1148-3.

Farahnakian:2014:ALB

[798] F. Farahnakian, M. Ebrahimi, M. Daneshtalab, P. Liljeberg, and J. Plosila. Adaptive load balancing in learningbased approaches for many-core embedded systems. The Journal of Supercomputing, 68(3):1214–1234, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1166-1.

Guo:2014:FTH

[799] Litao Guo and Xiaofeng Guo. Fault tolerance of hypercubes and folded hypercubes. The Journal of Supercomputing, 68(3):1235–1240, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1078-5.

Cheng:2014:DLT

[800] Eddie Cheng, Ke Qiu, and Zhizhang Shen. Deriving length two path centered surface area for the arrangement graph: a generating function approach. The Journal of Supercomputing, 68 (3):1241–1264, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1085-1.

Vijayalakshmi:2014:ASR

[801] Saravanan Vijayalakshmi, Alagan Anpalagan, D. P. Kothari, Isaac Woungang, and Mohammad S. Obaidat.

An analytical study of resource division and its impact on power and performance of multi-core processors. The Journal of Supercomputing, 68 (3):1265–1279, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1086-0.

Miedes:2014:IBM

[802] Emili Miedes and Francesc D. Muñoz-Escoí. Improving the benefits of multicast prioritization algorithms. The Journal of Supercomputing, 68 (3):1280–1301, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1087-z.

Naderan-Tahan:2014:APU

[803] Mahmood Naderan-Tahan and Hamid Sarbazi-Azad. Adaptive prefetching using global history buffer in multicore processors. The Journal of Supercomputing, 68(3):1302–1320, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1088-y.

Durao:2014:SRC

[804] Frederico Durao, Jose Fernando S. Carvalho, Anderson Fonseka, and Vinicius Cardoso Garcia. A systematic review on cloud computing. The Journal of Supercomputing, 68(3):1321–1346, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1089-x.

Mei:2014:RAS

[805] Jing Mei, Kenli Li, and Keqin Li. A resource-aware scheduling algorithm with reduced task duplication on heterogeneous computing systems. The Journal of Supercomputing, 68 (3):1347–1377, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1090-4.

Aldea:2014:BCA

[806] Sergio Aldea, Diego R. Llanos, and Arturo Gonzalez-Escribano. The BonaFide C Analyzer: automatic looplevel characterization and coverage measurement. The Journal of Supercomputing, 68(3):1378–1401, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1091-3.

Yang:2014:NMB

[807] Qiangpeng Yang, Chenglei Peng, He Zhao, Yao Yu, Yu Zhou, Ziqiang Wang, and Sidan Du. A new method based on PSR and EA–GMDH for host load prediction in cloud computing system. The Journal of Supercomputing, 68(3):1402–1417, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1097-x.

Kelefouras:2014:MMM

[808] Vasilios Kelefouras, Angeliki Kritikakou, and Costas Goutis. A matrix– matrix multiplication methodology for single/multi-core architectures using SIMD. The Journal of Supercomputing, 68(3):1418–1440, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1098-9.

Boton-Fernandez:2014:SAR

[809] María Botón-Fernández, Francisco Prieto-Castrillo, and Miguel A. Vega-Rodríguez. A self-adaptive resources selection model through a small-world based heuristic. The Journal of Supercomputing, 68(3):1441–1461, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1100-6.

Tang:2014:CTC

[810] Xiaolan Tang, Juhua Pu, Kun Ma, and Zhang Xiong. Cooperative transmission control scheme using erasure coding for vehicular delay-tolerant networks. The Journal of Supercomputing, 68(3):1462–1486, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1101-5.

Khan:2014:EGS

[811] Kashif Hesham Khan, Kalim Qureshi, and Mostafa Abd-El-Barr. An efficient grid scheduling strategy for data parallel applications. The Journal of Supercomputing, 68(3):1487–1502, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1114-0.

Park:2014:BPM

[812] Changyong Park, NamYeong Kwon, Honguk Woo, and Hyunseung Choo. Buffering in proxy mobile IPv6: implementation and analysis. The Journal of Supercomputing, 68(3):1503–1520, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1122-0.

Gong:2014:EPS

[813] Chunye Gong, Weimin Bao, Guojian Tang, Bo Yang, and Jie Liu. An efficient parallel solution for Caputo fractional reaction-diffusion equation. The Journal of Supercomputing, 68 (3):1521–1537, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1123-z.

Arab:2014:MCB

[814] Mohammad Norouzi Arab and Mohsen Sharifi. A model for communication between resource discovery and load balancing units in computing environments. The Journal of Supercomputing, 68(3):1538–1555, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1124-y.

Ros:2014:CBA

[815] Salvador Ros, Agustín C. Caminero, Roberto Hernández, Antonio Robles-Gómez, and Llanos Tobarra. Cloudbased architecture for web applications with load forecasting mechanism: a use case on the e-learning services of a distant university. The Journal of Supercomputing, 68(3):1556–1578, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1125-x.

Netjinda:2014:COS

[816] Nuttapong Netjinda, Booncharoen Sirinaovakul, and Tiranee Achalakul. Cost optimal scheduling in IaaS for dependent workload with particle swarm optimization. The Journal of Supercomputing, 68(3):1579–1603, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1126-9.

Karamti:2014:SAB

[817] Walid Karamti and Adel Mahfoudhi. Scheduling analysis based on model checking for multiprocessor real-time systems. The Journal of Supercomputing, 68(3):1604–1629, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://link.springer.com/
article/10.1007/s11227-014-1127-
8.

Laosooksathit:2014:RAP

[818] Supada Laosooksathit, Raja Nassar, Chokchai Leangsuksun, and Mihaela Paun. Reliability-aware performance model for optimal GPU-enabled cluster environment. The Journal of Supercomputing, 68(3):1630–1651, June 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11227-014-
                                        1128-7.

Campos:2014:GBH

[819] Ricardo Silva Campos, Marcelo Lobosco, and Rodrigo Weber dos Santos. A GPU-based heart simulator with mass-spring systems and cellular automaton. The Journal of Supercomputing, 69(1):1–8, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1199-5.

Skaruz:2014:TDP

[820] Jarosław Skaruz, Franciszek Seredynski, and Anna Piwonska. Twodimensional patterns and images reconstruction with use of cellular automata. The Journal of Supercomputing, 69(1):9–16, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-014-1214-x.pdf.

Nobile:2014:GAS

[821] Marco S. Nobile, Paolo Cazzaniga, Daniela Besozzi, and Giancarlo Mauri. GPU-accelerated simulations of massaction kinetics models with cupSODA. The Journal of Supercomputing, 69 (1):17–24, July 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1208-8.

Steuwer:2014:SHL

[822] Michel Steuwer and Sergei Gorlatch. SkelCL: a high-level extension of OpenCL for multi-GPU systems. The Journal of Supercomputing, 69(1):25– 33, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1213-y.

Zhou:2014:MCI

[823] Shijie Zhou, Yun R. Qu, and Viktor K. Prasanna. Multi-core implementation of decomposition-based packet classification algorithms. The Journal of Supercomputing, 69(1):34–42, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1205-y.

Basile:2014:FFS

[824] Davide Basile, Pierpaolo Degano, and Gian-Luigi Ferrari. A formal framework for secure and complying services. The Journal of Supercomputing, 69(1):43–52, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1211-0.

Toporkov:2014:SSA

[825] Victor Toporkov, Anna Toporkova, Alexey Tselishchev, and Dmitry Yemelyanov. Slot selection algorithms in distributed computing. The Journal of Supercomputing, 69(1):53–60, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1210-1.

Clarke:2014:FST

[826] David Clarke, Ziming Zhong, Vladimir Rychkov, and Alexey Lastovetsky. Fu-PerMod: a software tool for the optimization of data-parallel applications on heterogeneous platforms. The Journal of Supercomputing, 69(1):61–69, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1207-9.

Hoffmann:2014:AAC

[827] Rolf Hoffmann and Dominique Désérable. All-to-all communication with cellular automata agents in 2D grids: topologies, streets and performances. The Journal of Supercomputing, 69(1):70– 80, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1206-x.

Kopysov:2014:SHI

[828] Sergey Kopysov, Igor Kuzmin, Nikita Nedozhogin, Alexander Novikov, and Yulia Sagdeeva. Scalable hybrid implementation of the Schur complement method for multi-GPU systems. The Journal of Supercomputing, 69(1):81– 88, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1209-7.

Malyshkin:2014:PIL

[829] Victor E. Malyshkin and Vladislav A. Perepelkin. The PIC implementation in LuNA system of fragmented programming. The Journal of Supercomputing, 69(1):89–97, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1216-8.

Afzal:2014:LAL

[830] Samira Afzal and Hamid Beigy. A localization algorithm for large scale mobile wireless sensor networks: a learning approach. The Journal of Supercomputing, 69(1):98–120, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1129-6.

Kim:2014:SPA

[831] Jong-Seok Kim, Sung Won Kim, Ke Qiu, and Hyeong-Ok Lee. Some properties and algorithms for the hyper-torus network. The Journal of Supercomputing, 69(1):121–138, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1130-0.

Di:2014:CMC

[832] Sheng Di, Derrick Kondo, and Franck Cappello. Characterizing and modeling cloud applications/jobs on a Google data center. The Journal of Supercomputing, 69(1):139–160, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1131-z.

Bistouni:2014:IEG

[833] Fathollah Bistouni and Mohsen Jahanshahi. Improved extra group network: a new fault-tolerant multistage interconnection network. The Journal of Supercomputing, 69(1):161–199, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1132-y.

Khodja:2014:PSL

[834] Lilia Ziane Khodja, Raphaël Couturier, Arnaud Giersch, and Jacques M. Bahi. Parallel sparse linear solver with GM-RES method using minimization techniques of communications for GPU clusters. The Journal of Supercomputing, 69(1):200–224, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1143-8.

Chang:2014:PIC

[835] Hong-Yi Chang, Shih-Chang Huang, and Chih-Chun Lai. A personalized IPTV channel-recommendation mechanism based on the MapReduce framework. The Journal of Supercomputing, 69(1):225–247, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1145-6.

Cecilia:2014:ESP

[836] José M. Cecilia, José M. García, Ginés D. Guerrero, and Manuel Ujaldón. Evaluating the SAT problem on P systems for different highperformance architectures. The Journal of Supercomputing, 69(1):248–272, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1150-9.

Cano:2014:SCD

[837] Alberto Cano, Sebastián Ventura, and Krzysztof J. Cios. Scalable CAIM discretization on multiple GPUs using concurrent kernels. The Journal of Supercomputing, 69(1):273–292, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1151-8.

Bossard:2014:DPH

[838] Antoine Bossard. The decycling problem in hierarchical cubic networks. The Journal of Supercomputing, 69 (1):293–305, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1152-7.

Klavzar:2014:ADS

[839] Sandi Klavzar and Meijie Ma. Average distance, surface area, and other structural properties of exchanged hypercubes. The Journal of Supercomputing, 69(1):306–317, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1153-6.

Guerrero:2014:CEP

[840] Ginés D. Guerrero, José M. Cecilia, Antonio Llanes, José M. García, Martyn Amos, and Manuel Ujaldón. Comparative evaluation of platforms for parallel Ant Colony Optimization. The Journal of Supercomputing, 69 (1):318–329, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1154-5.

Choi:2014:CWE

[841] Hong Jun Choi, Dong Oh Son, Jong Myon Kim, and Cheol Hong Kim. Concurrent warp execution: improving performance of GPU-likely SIMD architecture by increasing resource utilization. The Journal of Supercomputing, 69(1):330–356, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1155-4.

Liu:2014:PAC

[842] Xiaodong Liu, Weiqin Tong, Xiaoli Zhi, Fu ZhiRen, and Liao WenZhao. Performance analysis of cloud computing services considering resources sharing among virtual machines. The Journal of Supercomputing, 69(1):357–374, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1156-3.

Jo:2014:ODE

[843] Heeseung Jo, Seung-Tae Hong, Jae-Woo Chang, and Dong Hoon Choi. Offloading data encryption to GPU in database systems. The Journal of Supercomputing, 69(1):375–394, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11227-014-
1159-0.

Farash:2014:SEI

[844] Mohammad Sabzinejad Farash and Mahmoud Ahmadian Attari. A secure and efficient identity-based authenticated key exchange protocol for mobile client-server networks. The Journal of Supercomputing, 69(1):395–411, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1170-5.

Vankeirsbilck:2014:USB

[845] Bert Vankeirsbilck, Lien Deboosere, Pieter Simoens, Piet Demeester, Filip De Turck, and Bart Dhoedt. User subscription-based resource management for Desktop-as-a-Service platforms. The Journal of Supercomputing, 69(1):412–428, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1171-4.

Cao:2014:EAH

[846] Zhibo Cao and Shoubin Dong. An energy-aware heuristic framework for virtual machine consolidation in Cloud computing. The Journal of Supercomputing, 69(1):429–451, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1172-3.

Sun:2014:PCM

[847] Zhanquan Sun, Geoffrey Fox, Weidong Gu, and Zhao Li. A parallel clustering method combined information bottleneck theory and centroid-based clustering. The Journal of Supercomputing, 69(1):452–467, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1174-1.

Rajkumar:2014:DDG

[848] S. Rajkumar and Neeraj Kumar Goyal. Design of 4-disjoint gamma interconnection network layouts and reliability analysis of gamma interconnection networks. The Journal of Supercomputing, 69(1):468–491, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1175-0.

Vilaplana:2014:QTM

[849] Jordi Vilaplana, Francesc Solsona, Ivan Teixidó, Jordi Mateo, Francesc Abella, and Josep Rius. A queuing theory model for cloud computing. The Journal of Supercomputing, 69(1):492–507, July 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1177-y.

Shen:2014:P

[850] Hong Shen and Shaohua Tang. Preface. The Journal of Supercomputing, 69(2):509–511, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-014-1244-4.pdf.

Tian:2014:MNL

[851] Hui Tian, Hong Shen, and Yingpeng Sang. Maximizing network lifetime in wireless sensor networks with regular topologies. The Journal of Supercomputing, 69(2):512–527, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
0987-7.

Xu:2014:VCA

[852] Lingling Xu and Shaohua Tang. Verifiable computation with access control in cloud computing. The Journal of Supercomputing, 69(2):528–546, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1039-z.

Zhang:2014:LDP

[853] Ludan Zhang, Yi Liu, Rui Wang, and Depei Qian. Lightweight dynamic partitioning for last-level cache of multicore processor on real system. The Journal of Supercomputing, 69 (2):547–560, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1092-2.

Fujita:2014:ASB

[854] Satoshi Fujita. Approximation scheme for burst scheduling with minimum overhead in time slicing mobile TV. The Journal of Supercomputing, 69 (2):561–575, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1093-1.

HoseinyFarahabady:2014:RAS

[855] MohammadReza HoseinyFarahabady, Young Choon Lee, and Albert Y. Zomaya. Randomized approximation scheme for resource allocation in hybrid-cloud environment. The Journal of Supercomputing, 69(2):576–592, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1094-0.

Al-Hinai:2014:TST

[856] Aysha Al-Hinai, Haibo Zhang, Yawen Chen, and Yidong Li. TB-SnW: Trust-based Spray-and-Wait routing for delay-tolerant networks. The Journal of Supercomputing, 69(2):593–609, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1095-z.

Wu:2014:PRA

[857] Jigang Wu, Guiyuan Jiang, Yuze Shen, Siew-Kei Lam, Jizhou Sun, and Thambipillai Srikanthan. Parallel reconfiguration algorithms for mesh-connected processor arrays. The Journal of Supercomputing, 69(2):610–628, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1096-y.

Gava:2014:BAF

[858] Frédéric Gava, Franck Pommereau, and Michaël Guedj. A BSP algorithm for on-the-fly checking CTL* formulas on security protocols. The Journal of Supercomputing, 69(2):629–672, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1099-8.

Shen:2014:SPE

[859] Ransheng Shen, Xianfeng Li, and Hui Li. A space- and power-efficient multi-match packet classification technique combining TCAMs and SRAMs. The Journal of Supercomputing, 69 (2):673–692, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1109-x.

Yan:2014:OMB

[860] Xin Yan, Xiaohua Shi, Lina Wang, and Haiyan Yang. An OpenCL microbenchmark suite for GPUs and CPUs. The Journal of Supercomputing, 69 (2):693–713, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1112-2.

Zhu:2014:ALM

[861] Wenxing Zhu, Jianli Chen, and Weiguo Li. An augmented Lagrangian method for VLSI global placement. The Journal of Supercomputing, 69(2):714–738, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1113-1.

Teng:2014:NRT

[862] Fei Teng, Frédéric Magoulès, Lei Yu, and Tianrui Li. A novel real-time scheduling algorithm and performance analysis of a MapReduce-based cloud. The Journal of Supercomputing, 69 (2):739–765, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1115-z.

Furhad:2014:SCM

[863] Md. Hasan Furhad and Jong-Myon Kim. A shortly connected mesh topology for high performance and energy efficient network-on-chip architectures. The Journal of Supercomputing, 69 (2):766–792, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1178-x.

Lee:2014:JPT

[864] Taewhi Lee, Hye-Chan Bae, and Hyoung-Joo Kim. Join processing with threshold-based filtering in MapReduce. The Journal of Supercomputing, 69(2):793–813, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1179-9.

Abbas:2014:PEM

[865] Mostafa M. Abbas, Hazem M. Bahig, Mohamed Abouelhoda, and M. M. Mohie-Eldin. Parallelizing exact motif finding algorithms on multi-core. The Journal of Supercomputing, 69 (2):814–826, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1180-3.

Entezari-Maleki:2014:CPA

[866] Reza Entezari-Maleki, Ali Mohammadkhan, Heon Young Yeom, and Ali Movaghar. Combined performance and availability analysis of distributed resources in grid computing. The Journal of Supercomputing, 69(2):827–844, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1184-z.

Hadian:2014:HPP

[867] Ali Hadian and Saeed Shahrivari. High performance parallel k-means clustering for disk-resident datasets on multi-core CPUs. The Journal of Supercomputing, 69(2):845–863, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1185-y.

Li:2014:RTO

[868] Liang Li, Endong Wang, Xingjun Zhang, Kang Yan, Tao Ju, and Xiaoshe Dong. A run-time optimization approach for reducing data movements using locality-aware searching. The Journal of Supercomputing, 69 (2):864–886, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1186-x.

Qi:2014:STB

[869] Han Qi, Muhammad Shiraz, Abdullah Gani, Md Whaiduzzaman, and Suleman Khan. Sierpinski triangle based data center architecture in cloud computing. The Journal of Supercomputing, 69(2):887–907, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1187-9.

Lakhlef:2014:EME

[870] Hicham Lakhlef, Hakim Mabed, and Julien Bourgeois. An energy and memory-efficient distributed selfreconfiguration for modular sensor/ robot networks. The Journal of Supercomputing, 69(2):908–929, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1196-8.

Lakshminarayanan:2014:SSP

[871] Mahalakshmi Lakshminarayanan, William F. Acosta, Robert C. Green II, and Vijay Devabhaktuni. Strategic and suave processing for performing similarity joins using MapReduce. The Journal of Supercomputing, 69(2):930–954, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1197-7.

Javanmardi:2014:PNA

[872] Saeed Javanmardi, Mohammad Shojafar, Shahdad Shariatmadari, Jemal H. Abawajy, and Mukesh Singhal. PGSW-OS: a novel approach for resource management in a semantic web operating system based on a P2P grid architecture. The Journal of Supercomputing, 69(2):955–975, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1221-y.

Zhou:2014:MSM

[873] Xu Zhou, YanTao Zhou, KenLi Li, Ahmed Sallam, and Keqin Li. Molecular solutions for minimum and exact cover problems in the tile assembly model. The Journal of Supercomputing, 69(2):976–1005, August 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1222-x.

Yen:2014:CAT

[874] Neil Y. Yen, Cho-Li Wang, Sajid Hussain, and Jong Hyuk Park. Computational awareness towards green environments. The Journal of Supercomputing, 69(3):1007–1012, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1281-z;   http://link.springer.
com/content/pdf/10.1007/s11227-
014-1281-z.pdf.

Lee:2014:PAB

[875] Hwa Min Lee, Young-Sik Jeong, and Haeng Jin Jang. Performance analysis based resource allocation for green cloud computing. The Journal of Supercomputing, 69(3):1013– 1026, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1020-x.

Lee:2014:PEV

[876] Wooyeob Lee, Wontae Kim, and Inwhee Joe. A power-efficient vertical handover with MIH-based network scanning through consistency check. The Journal of Supercomputing, 69(3): 1027–1038, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1003-y.

Ihm:2014:EDB

[877] Sun-Young Ihm, Aziz Nasridinov, and Jeong-Hoon Lee. Efficient dualitybased subsequent matching on timeseries data in green computing. The Journal of Supercomputing, 69(3): 1039–1053, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1028-2.

Park:2014:CCB

[878] Jong Hyuk Park and Hwa Young Jeong. Cloud computing-based jam management for a manufacturing system in a Green IT environment. The Journal of Supercomputing, 69(3): 1054–1067, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-1007-7.

Yu:2014:ECR

[879] Junyang Yu, Zhigang Hu, Neal N. Xiong, He Liu, and Zhou Zhou. An energy conservation replica placement strategy for Dynamo. The Journal of Supercomputing, 69(3):1068– 1086, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1219-5.

Enokido:2014:EES

[880] Tomoya Enokido, Ailixier Aikebaier, and Makoto Takizawa. Energy-efficient server clusters to perform communication type application processes. The Journal of Supercomputing, 69(3): 1087–1102, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1025-5.

Yang:2014:ICV

[881] Chao-Tung Yang, Jung-Chun Liu, and Ching-Hsien Hsu. On improvement of cloud virtual machine availability with virtualization fault tolerance mechanism. The Journal of Supercomputing, 69(3):1103–1122, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1045-1.

Wong:2014:PHD

[882] Kok-Seng Wong and Myung Ho Kim. On private Hamming distance computation. The Journal of Supercomputing, 69(3):1123–1138, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1063-z.

Yen:2014:HAD

[883] Chih-Ta Yen, Hsu-Chih Cheng, and Ing-Jr Ding. Hybrid analog/digital wavelength-time optical CDMA systems in radio-over-fiber transmissions. The Journal of Supercomputing, 69(3): 1139–1153, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0964-1.

Li:2014:RLB

[884] Weimin Li, Mengke Yao, Xiaokang Zhou, and Shoji Nishimura. Recommendation of location-based services based on composite measures of trust degree. The Journal of Supercomputing, 69(3):1154–1165, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1084-2.

Tsung:2014:CVB

[885] Chen-Kun Tsung, Hann-Jang Ho, and Sing-Ling Lee. Correcting vindictive bidding behaviors in sponsored search auctions. The Journal of Supercomputing, 69(3):1166–1182, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        1002-z.

Dhurandher:2014:GGA

[886] Sanjay K. Dhurandher, Deepak Kumar Sharma, and Isaac Woungang. GAER: genetic algorithmbased energy-efficient routing protocol for infrastructure-less opportunistic networks. The Journal of Supercomputing, 69(3):1183–1214, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1195-9.

Park:2014:ACS

[887] Su-Wan Park, Deok Gyu Lee, Jong Wook Han, and JeongNyeo Kim. Automated and coupled services of advanced smart surveillance systems toward green IT: tracking, retrieval and digital evidence. The Journal of Supercomputing, 69(3):1215–1234, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1164-3.

Wang:2014:GIE

[888] Hui Wang, Shengliang Xiao, Feiyu Lin, and Tianchu Yang. Group improved enhanced dynamic frame slotted ALOHA anti-collision algorithm. The Journal of Supercomputing, 69(3): 1235–1253, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1189-7.

Zhou:2014:DFV

[889] Rui Zhou, Chanjuan Li, Rong Min, Qi Yu, and Fei Gu. On design and formal verification of SNSP: a novel real-time communication protocol for safety-critical applications. The Journal of Supercomputing, 69(3): 1254–1283, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1157-2.

Ouyang:2014:OCP

[890] Yen-Chieh Ouyang, Yi-Ju Chiang, and Ching-Hsien Hsu. An optimal control policy to realize green cloud systems with SLA-awareness. The Journal of Supercomputing, 69(3):1284– 1310, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1190-1.

Choi:2014:AHP

[891] Min-Hyung Choi, Mohammed Bahni Alquzi, and Min Hong. Assessment of human perceptual sensitivity to physically non-conforming motion in virtual environments. The Journal of Supercomputing, 69(3):1311–1323, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1169-y.

Choi:2014:EIW

[892] Min Choi, Ikkyun Kim, and Jong Hyuk Park. An enhanced integrity of web contents through mobile cloud environments. The Journal of Supercomputing, 69(3):1324–1341, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1278-7.

Lu:2014:ASF

[893] Chia-Han Lu, Wen-Li Shih, Chung-Ju Wu, and Jenq Kuen Lee. Achieving spilling-friendly register file assignment for highly distributed register files. The Journal of Supercomputing, 69(3): 1342–1362, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1181-2.

Dai:2014:AOM

[894] Yuan Dai, Dongjian He, Yong Fang, and Long Yang. Accelerating 2D orthogonal matching pursuit algorithm on GPU. The Journal of Supercomputing, 69(3):1363–1381, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1188-8.

Chang:2014:ICR

[895] Junsheng Chang, Zhengbin Pang, Weixia Xu, Huaimin Wang, and Gang Yin. An incentive compatible reputation mechanism for P2P systems. The Journal of Supercomputing, 69(3): 1382–1409, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1204-z.

Villar:2014:FCM

[896] Juan A. Villar, Francisco J. Andújar, Francisco J. Alfaro, José L. Sánchez, and José Duato. Formalization and configuration methodology for highradix combined switches. The Journal of Supercomputing, 69(3):1410– 1444, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1223-9.

Horri:2014:NRA

[897] Abbas Horri, Mohammad Sadegh Mozafari, and Gholamhossein Dastghaibyfard. Novel resource allocation algorithms to performance and energy efficiency in cloud computing. The Journal of Supercomputing, 69(3): 1445–1461, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1224-8.

Yan:2014:EFG

[898] Jie Yan, Guangming Tan, and Ninghui Sun. Exploiting fine-grained parallelism in graph traversal algorithms via lock virtualization on multi-core architecture. The Journal of Supercomputing, 69(3):1462–1490, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1239-1.

Fu:2014:OMA

[899] Weiwei Fu, Tianzhou Chen, Chao Wang, and Li Liu. Optimizing memory access traffic via runtime thread migration for on-chip distributed memory systems. The Journal of Supercomputing, 69(3):1491–1516, September 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1240-8.

Yeo:2014:ESS

[900] Sangsoo Yeo and Shiuh-Jeng Wang. Editorial of special section on Hybrid Information Security Technologies: part II. The Journal of Supercomputing, 70(1):1–2, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1284-9;   http:
//link.springer.com/content/pdf/
10.1007/s11227-014-1284-9.pdf.

Kim:2014:SCC

[901] Hyongsoon Kim and Eunyoung Lee. Secure component composition with modular behavioral properties. The Journal of Supercomputing, 70(1):3– 19, October 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1283-x.

Lee:2014:IDF

[902] Seokjun Lee and Taeshik Shon. Improved deleted file recovery technique for Ext2/3 filesystem. The Journal of Supercomputing, 70(1):20–30, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1282-y.

Luo:2014:ARP

[903] Jia Ning Luo and Ming Hour Yang. An anonymous e-rental protocol based on ID-based cryptography and NFC. The Journal of Supercomputing, 70 (1):31–53, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1051-3.

Chen:2014:DAS

[904] Chin-Ling Chen, Yu-Yi Chen, Cheng-Chi Lee, and Cheng-Hsun Wu. Design and analysis of a secure and effective emergency system for mountaineering events. The Journal of Supercomputing, 70(1):54–74, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1066-9.

Chou:2014:EMA

[905] Jue-Sam Chou. An efficient mutual authentication RFID scheme based on elliptic curve cryptography. The Journal of Supercomputing, 70(1):75–94, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1073-x.

Lee:2014:BCP

[906] Changhoon Lee. Biclique cryptanalysis of PRESENT-80 and PRESENT-128. The Journal of Supercomputing, 70(1):95–103, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1103-3.

Jabbar:2014:MCD

[907] Sohail Jabbar, Abid Ali Minhas, Anand Paul, and Seungmin Rho. Multilayer cluster designing algorithm for lifetime improvement of wireless sensor networks. The Journal of Supercomputing, 70(1):104–132, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1108-y.

Hsieh:2014:AMU

[908] Wen-Bin Hsieh and Jenq-Shiou Leu. An anonymous mobile user authentication protocol using self-certified public keys based on multi-server architectures. The Journal of Supercomputing, 70(1):133–148, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1135-8.

Majore:2014:SRE

[909] Sekie Amanuel Majore, Hyunguk Yoo, and Taeshik Shon. Secure and reliable electronic record management system using digital forensic technologies. The Journal of Supercomputing, 70(1):149–165, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1137-6.

Ahn:2014:SEH

[910] Sang Un Ahn, Il Yeon Yeo, and Sang Oh Park. Secure and efficient high-performance PROOF-based cluster system for high-energy physics. The Journal of Supercomputing, 70 (1):166–176, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1146-5.

Zuo:2014:DAS

[911] Liming Zuo, Neeraj Kumar, Hang Tu, and Ankit Singh. Detection and analysis of secure intelligent universal designated verifier signature scheme for electronic voting system. The Journal of Supercomputing, 70(1):177–199, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1149-2.

Yoon:2014:UTC

[912] Heejun Yoon, Il-Yeon Yeo, and Jeong Heon Kim. Updating the trusted connection of re-organized computing resource under the automated system management platform. The Journal of Supercomputing, 70(1):200–210, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1158-1.

Lee:2014:SSS

[913] Chang-Moo Lee and Hangbae Chang. A study on security strategy in ICT convergence environment. The Journal of Supercomputing, 70(1):211–223, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1194-x.

Tu:2014:EPB

[914] Hang Tu, Neeraj Kumar, Debiao He, and Jongsung Kim. An efficient password-based three-party authenticated multiple key exchange protocol for wireless mobile networks.

The Journal of Supercomputing, 70 (1):224–235, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1198-6.

Koo:2014:CRB

[915] Bonwook Koo, Dongyoung Roh, and Daesung Kwon. Converting random bits into random numbers. The Journal of Supercomputing, 70(1):236–246, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1202-1.

Hsu:2014:VWR

[916] Fu-Hau Hsu, Min-Hao Wu, Cheng-Hsing Yang, and Shiuh-Jeng Wang. Visible watermarking with reversibility of multimedia images for ownership declarations. The Journal of Supercomputing, 70(1):247–268, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1258-y.

Wang:2014:SSC

[917] Danli Wang, Liang He, and Keqin Dou. StoryCube: supporting children’s storytelling with a tangible tool. The Journal of Supercomputing, 70 (1):269–283, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-012-0855-x.

Yang:2014:HPD

[918] Luobin Yang, Steve C. Chiu, and Wei-Keng Liao. High performance data clustering: a comparative analysis of performance for GPU, RASC, MPI, and OpenMP implementations. The Journal of Supercomputing, 70 (1):284–300, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-013-0906-y.

Zydek:2014:MCL

[919] Dawid Zydek, Grzegorz Chmaj, and Steve Chiu. Modeling computational limitations in h-Phy and Overlay– NoC architectures. The Journal of Supercomputing, 70(1):301–320, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-013-
                                        0932-9.

Che:2014:MPC

[920] Yonggang Che, Lilun Zhang, Yongxian Wang, Chuanfu Xu, Wei Liu, and Zhenghua Wang. Microarchitectural performance comparison of Intel Knights Corner and Intel Sandy Bridge with CFD applications. The Journal of Supercomputing, 70(1):321–348, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1245-3.

Karanikolaou:2014:PSE

[921] E. M. Karanikolaou, E. I. Milovanović, I. Z. Milovanović, and M. P. Bekakos. Performance scalability and energy consumption on distributed and manycore platforms. The Journal of Supercomputing, 70(1):349–364, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1248-0.

Gravvanis:2014:DGA

[922] George A. Gravvanis and Christos K. Filelis-Papadopoulos. Distributed generic approximate sparse inverses. The Journal of Supercomputing, 70 (1):365–384, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1249-z.

Hu:2014:POE

[923] Su Hu, Wenzheng Xu, Jing Lin, and Xiaola Lin. Probabilistic odd-even: an adaptive wormhole routing algorithm for 2D mesh network-on-chip. The Journal of Supercomputing, 70 (1):385–407, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1250-6.

Yang:2014:CCS

[924] Xiaoyu Yang, David Wallom, Simon Waddington, Jianwu Wang, Arif Shaon, Brian Matthews, Michael Wilson, Yike Guo, Li Guo, Jon D. Blower, Athanasios V. Vasilakos, Kecheng Liu, and Philip Kershaw. Cloud computing in e-science: research challenges and opportunities. The Journal of Supercomputing, 70(1):408–464, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1251-5.

Saravanan:2014:CSS

[925] Vijayalakshmi Saravanan, Alagan Anpalagan, D. P. Kothari, Isaac Woungang, and Mohammad S. Obaidat. A comparative simulation study on the power-performance of multi-core architecture. The Journal of Supercomputing, 70(1):465–487, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1263-1.

Su:2014:ECG

[926] Yan Su, Feng Shi, Shahnawaz Talpur, Jin Wei, and Hai Tan. Exploiting controlled-grained parallelism in message-driven stream programs. The Journal of Supercomputing, 70(1): 488–509, October 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1264-0.

Ranilla:2014:HPC

[927] J. Ranilla, E. M. Garzón, and J. Vigo-Aguiar. High performance computing: an essential tool for science and engineering breakthroughs. The Journal of Supercomputing, 70(2): 511–513, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1279-6;   http:
//link.springer.com/content/pdf/
10.1007/s11227-014-1279-6.pdf.

Frances:2014:PAS

[928] Jorge Francés, Sergio Bleda, and Andrés Márquez. Performance analysis of SSE and AVX instructions in multi-core CPUs and GPU computing on FDTD scheme for solid and fluid vibration problems. The Journal of Supercomputing, 70(2):514–526, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1065-x.

Uribe-Paredes:2014:TES

[929] Roberto Uribe-Paredes, Diego Cazorla, and Enrique Arias. Towards an efficient static scheduling scheme for delivering queries to heterogeneous clusters in the similarity search problem. The Journal of Supercomputing, 70(2): 527–540, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1079-4.

Ramiro:2014:GII

[930] Carla Ramiro, M. Ángeles Simarro, and F. J. Martínez-Zaldívar. A GPU implementation of an iterative receiver for energy saving MIMO ID– BICM systems. The Journal of Supercomputing, 70(2):541–551, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1081-x.

Fernandez:2014:CPE

[931] Víctor Fernández, Juan M. Orduña, and Pedro Morillo. Comparative performance evaluation of CAR systems based on mobile phones and feature tracking. The Journal of Supercomputing, 70(2):552–563, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1082-9.

Alonso:2014:PAN

[932] P. Alonso, V. M. García, and F. J. Martínez-Zaldívar. Parallel approach to NNMF on multicore architecture. The Journal of Supercomputing, 70(2): 564–576, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1083-8.

Tabik:2014:PEK

[933] S. Tabik, G. Ortega, and E. M. Garzón. Performance evaluation of kernel fusion BLAS routines on the GPU: iterative solvers as case study. The Journal of Supercomputing, 70(2): 577–587, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1102-4.

Sevilla:2014:UBC

[934] Jorge Sevilla, Sergio Bernabe, and Antonio Plaza. Unmixing-based content retrieval system for remotely sensed hyperspectral imagery on GPUs. The Journal of Supercomputing, 70(2): 588–599, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1104-2.

Arrondo:2014:SLF

[935] A. G. Arrondo, J. L. Redondo, J. Fernández, and P. M. Ortigosa. Solving a leader-follower facility problem via parallel evolutionary approaches. The Journal of Supercomputing, 70(2): 600–611, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1106-0.

Lopez-Portugues:2014:ANS

[936] M. López-Portugués, J. A. López-Fernández, and N. Díaz-Gracia. Aircraft noise scattering prediction using different accelerator architectures. The Journal of Supercomputing, 70(2): 612–622, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1107-z.

Peinado:2014:STI

[937] Jesús Peinado, Pedro Alonso, and Javier Ibáñez. Solving time-invariant differential matrix Riccati equations using GPGPU computing. The Journal of Supercomputing, 70(2):623– 636, November 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1111-3.

Arnal:2014:PRE

[938] Josep Arnal, Héctor Migallón, and Violeta Migallón. Parallel relaxed and extrapolated algorithms for computing PageRank. The Journal of Supercomputing, 70(2):637–648, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1118-9.

Acosta:2014:ATM

[939] Alejandro Acosta and Francisco Almeida. Android^{T M} development and performance analysis. The Journal of Supercomputing, 70(2):649– 659, November 2014. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1119-8.

Cores:2014:MAL

[940] Iván Cores, Gabriel Rodríguez, and María J. Martín. In-memory application-level checkpoint-based migration for MPI programs. The Journal of Supercomputing, 70(2): 660–670, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1120-2.

Pinol:2014:PSA

[941] P. Piñol, H. Migallón, O. López-Granado, and M. P. Malumbres. Parallel strategies analysis over the HEVC encoder. The Journal of Supercomputing, 70(2):671–683, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1121-1.

Garcia-Martinez:2014:GIH

[942] J. M. García-Martínez, E. M. Garzón, and P. M. Ortigosa. A GPU implementation of a hybrid evolutionary algorithm: GPuEGO. The Journal of Supercomputing, 70(2):684–695, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1136-7.

Lorenzo:2014:DRM

[943] O. G. Lorenzo, T. F. Pena, J. C. Cabaleiro, and J. C. Pichel. 3DyRM: a dynamic roofline model including memory latency information. The Journal of Supercomputing, 70(2): 696–708, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1163-4.

Bermejo:2014:DPM

[944] J. A. Alvarez Bermejo, M. A. Lodroman, and J. A. Lopez-Ramos. A decentralized protocol for mobile control access. The Journal of Supercomputing, 70(2):709–720, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1165-2.

Brun:2014:EMM

[945] Carlos Brun, Tomàs Margalef, Ana Cortés, and Anna Sikora. Enhancing multi-model forest fire spread prediction by exploiting multi-core parallelism. The Journal of Supercomputing, 70(2):721–732, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1168-z.

Boratto:2014:ART

[946] Murilo Boratto, Pedro Alonso, and Domingo Gimenéz. Automatic routine tuning to represent landform attributes on multicore and multi-GPU systems. The Journal of Supercomputing, 70(2): 733–745, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1191-0.

Silva:2014:ASF

[947] Juan Pablo Silva, José Hagopian, Marcel Burdiat, Ernesto Dufrechou, Martín Pedemonte, Alejandro Gutiérrez, Gabriel Cazes, and Pablo Ezzatti. Another step to the full GPU implementation of the weather research and forecasting model. The Journal of Supercomputing, 70(2):746–755, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1193-y.

Navarro:2014:SMU

[948] Angeles Navarro, Antonio Vilches, Francisco Corbera, and Rafael Asenjo. Strategies for maximizing utilization on multi-CPU and multi-GPU heterogeneous architectures. The Journal of Supercomputing, 70(2):756–771, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1200-3.

Jozsa:2014:HPL

[949] Csaba M. Józsa, Fernando Domene, Antonio M. Vidal, Gema Piñero, and Alberto González. High performance lattice reduction on heterogeneous computing platform. The Journal of Supercomputing, 70(2): 772–785, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1201-2.

Ortega-Arranz:2014:OAI

[950] Hector Ortega-Arranz, Yuri Torres, Arturo Gonzalez-Escribano, and Diego R. Llanos. Optimizing an APSP implementation for NVIDIA GPUs using kernel characterization criteria. The Journal of Supercomputing, 70(2): 786–798, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1212-z.

Valero-Lara:2014:ASF

[951] Pedro Valero-Lara. Accelerating solidfluid interaction based on the immersed boundary method on multicore and GPU architectures. The Journal of Supercomputing, 70(2):799–815, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1262-2.

Gonzalez-Dominguez:2014:AAW

[952] Jorge González-Domínguez, Osni A. Marques, María J. Martín, and Juan Touriño. A 2D algorithm with asymmetric workload for the UPC conjugate gradient method. The Journal of Supercomputing, 70(2):816–829, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1300-0.

Bernabe:2014:IAE

[953] Gregorio Bernabé, Javier Cuenca, Luis Pedro García, and Domingo Giménez. Improving an autotuning engine for 3D Fast Wavelet Transform on manycore systems. The Journal of Supercomputing, 70(2):830–844, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1302-y.

Pahlavan:2014:PRH

[954] Ali Pahlavan, Mahmoud Momtazpour, and Maziar Goudarzi. Power reduction in HPC data centers: a joint server placement and chassis consolidation approach. The Journal of Supercomputing, 70(2):845–879, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1265-z.

Gonzalez-Alvarez:2014:POH

[955] David L. González-Álvarez, Miguel A. Vega-Rodríguez, and Álvaro Rubio-Largo. Parallelizing and optimizing a hybrid differential evolution with Pareto tournaments for discovering motifs in DNA sequences. The Journal of Supercomputing, 70(2): 880–905, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1266-y.

Gholizadeh:2014:OPD

[956] Somayyeh Gholizadeh, Azizol Abdullah, Mohamed Othman, Zurina Mohd Hanapi, and Mohsen Heydarian. Optimized performance data transmission in Mobile IP networks. The Journal of Supercomputing, 70(2): 906–929, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1267-x.

Lai:2014:NHC

[957] Guoming Lai, Dehui Yuan, and Shenyun Yang. A new hybrid combinatorial genetic algorithm for multidimensional knapsack problems. The Journal of Supercomputing, 70(2): 930–945, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1268-9.

Khan:2014:BBB

[958] Abdul Nasir Khan, M. L. Mat Kiah, Mazhar Ali, Sajjad A. Madani, Atta ur Rehman Khan, and Shahaboddin Shamshirband. BSS: block-based sharing scheme for secure data storage services in mobile cloud environment. The Journal of Supercomputing, 70(2): 946–976, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1269-8.

Rahnama:2014:TIP

[959] Behnam Rahnama. Towards improved parallelism through order reduction of accessing data in nD matrices.

The Journal of Supercomputing, 70(2): 977–986, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1271-1.

Farash:2014:CIE

[960] Mohammad Sabzinejad Farash. Cryptanalysis and improvement of an efficient mutual authentication RFID scheme based on elliptic curve cryptography. The Journal of Supercomputing, 70(2):987–1001, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1272-0.

Farash:2014:ECC

[961] Mohammad Sabzinejad Farash and Mahmoud Ahmadian Attari. An efficient client–client password-based authentication scheme with provable security. The Journal of Supercomputing, 70(2):1002–1022, November 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1273-z.

Chen:2014:RRC

[962] Lanping Chen, Zhengzhi Han, and Zhenghua Ma. Research on robust control and exponential stabilization for large scale impulsive hybrid network systems with time-delay. The Journal of Supercomputing, 70(3): 1023–1035, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-013-1064-y.

ElBouabidi:2014:DAS

[963] Imen El Bouabidi, Faouzi Zarai, and Mohammad S. Obaidat. Design and analysis of secure host-based mobility protocol for wireless heterogeneous networks. The Journal of Supercomputing, 70(3):1036–1050, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-013-
1080-y.

Yan:2014:PPR

[964] Zheng Yan, Yu Chen, and Yue Shen. PerContRep: a practical reputation system for pervasive content services. The Journal of Supercomputing, 70(3): 1051–1074, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1116-y.

Perez:2014:TTR

[965] Juan M. Marin Perez and Jorge Bernal Bernabe. Taxonomy of trust relationships in authorization domains for cloud computing. The Journal of Supercomputing, 70(3):1075–1099, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1117-x.

Yeh:2014:ITP

[966] Jyh haw Yeh. The insecurity of two proxy signcryption schemes: proxy credential forgery attack and how to prevent it. The Journal of Supercomputing, 70(3):1100–1119, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1134-9.

Nieminen:2014:RAD

[967] Mikko Nieminen, Nikolay Tcholtchev, and Ina Schieferdecker. Robust architecture for distributed intelligence in an IP-based mobile wide-area surveillance system. The Journal of Supercomputing, 70(3):1120–1141, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1160-7;   http://link.springer.
com/content/pdf/10.1007/s11227-
014-1160-7.pdf.

Dong:2014:UAD

[968] Mianxiong Dong, Kaoru Ota, Man Lin, Zunyi Tang, and Suguo Du. UAVassisted data gathering in wireless sensor networks. The Journal of Supercomputing, 70(3):1142–1155, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1161-6.

Luo:2014:PES

[969] C. Luo, H. Okamura, and T. Dohi. Performance evaluation of snapshot isolation in distributed database system under failure-prone environment. The Journal of Supercomputing, 70(3): 1156–1179, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1162-5.

Liu:2014:HAV

[970] Xingang Liu, Wenjie Yang, and Zhixin Shen. H.264/AVC video error concealment algorithm by employing motion vector recovery under cloud computing environment. The Journal of Supercomputing, 70(3):1180–1199, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1167-0.

Saleemi:2014:ESS

[971] M. Mohsin Saleemi and Johan Lilius. Exploiting smart spaces for interactive TV applications development. The Journal of Supercomputing, 70(3): 1200–1217, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1183-0. See erra-

tum [992].

Santos:2014:DSR

[972] André C. Santos, João M. P. Cardoso, and Pedro C. Diniz. A DSL for specifying run-time adaptations for embedded systems: an application to vehicle stereo navigation. The Journal of Supercomputing, 70(3):1218–1248, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1192-z.

Cui:2014:OBD

[973] Xiaoli Cui, Pingfei Zhu, Xin Yang, Keqiu Li, and Changqing Ji. Optimized big data K-means clustering using MapReduce. The Journal of Supercomputing, 70(3):1249–1259, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1225-7.

Zhu:2014:PEA

[974] Liang Zhu, Hai Jin, Xiaofei Liao, and Jianhui Yue. Performance-energy adaptation of parallel programs in pervasive computing. The Journal of Supercomputing, 70(3):1260–1278, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1226-6.

Tang:2014:DFS

[975] Zhuo Tang, Yanqing Mo, Kenli Li, and Keqin Li. Dynamic forecast scheduling algorithm for virtual machine placement in cloud computing environment. The Journal of Supercomputing, 70(3): 1279–1296, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1227-5.

Wang:2014:SBM

[976] Yufeng Wang, Jing Tang, Qun Jin, and Jianhua Ma. On studying business models in mobile social networks based on two-sided market (TSM). The Journal of Supercomputing, 70(3): 1297–1317, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1228-4.

Xu:2014:SSB

[977] Huang Xu, Zhiwen Yu, Zhu Wang, and Hongbo Ni. SmartMic: a smartphonebased meeting support system. The Journal of Supercomputing, 70(3): 1318–1330, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1229-3.

Chen:2014:EAL

[978] Zhijian Chen, Wenhai Luo, Dan Wu, Xiang Huang, and Jian He. Exploiting application-level similarity to improve SSD cache performance in Hadoop. The Journal of Supercomputing, 70(3): 1331–1344, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1230-x.

Dai:2014:CAA

[979] Jie Dai, Yu Zhao, Yunhuai Liu, Li Qi, and Chuanping Hu. Cloud-assisted analysis for energy efficiency in intelligent video systems. The Journal of Supercomputing, 70(3):1345–1364, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1231-9.

Zhang:2014:DCN

[980] Xu Zhang, Hai Wang, Qingyuan Gong, and Xin Wang. Decluster: a complex network model-based data center network topology. The Journal of Supercomputing, 70(3):1365–1382, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1232-8.

Xia:2014:MUD

[981] Wenfeng Xia, Yonggang Wen, Haiyong Xie, and Bin Liu. μDC2: unified data collection for data centers. The Journal of Supercomputing, 70(3): 1383–1404, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1233-7.

Zhang:2014:LFL

[982] Mingwu Zhang, Chunzhi Wang, and Kirill Morozov. LR-FEAD: leakagetolerating and attribute-hiding functional encryption mechanism with delegation in affine subspaces. The Journal of Supercomputing, 70(3): 1405–1432, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1234-6.

Yuan:2014:CET

[983] Jinfeng Yuan, Weizhong Qiang, Hai Jin, and Deqing Zou. CloudTaint: an elastic taint tracking framework for malware detection in the cloud. The Journal of Supercomputing, 70(3): 1433–1450, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1235-5.

Rauber:2014:EMM

[984] Thomas Rauber, Gudula Rünger, Michael Schwind, Haibin Xu, and Simon Melzner. Energy measurement, modeling, and prediction for processors with frequency scaling. The Journal of Supercomputing, 70(3):1451–1476, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1236-4.

Tu:2014:ESM

[985] Renlong Tu, Xin Wang, and Yue Yang. Energy-saving model for SDN data centers. The Journal of Supercomputing, 70(3):1477–1495, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1237-3.

Wu:2014:PCM

[986] Song Wu, Yaqiong Peng, Hai Jin, and Jun Zhang. Peacock: a customizable MapReduce for multicore platform. The Journal of Supercomputing, 70(3): 1496–1513, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1238-2.

Takouna:2014:MRO

[987] Ibrahim Takouna, Kai Sachs, and Christoph Meinel. Multiperiod robust optimization for proactive resource provisioning in virtualized data centers. The Journal of Supercomputing, 70(3): 1514–1536, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1246-2.

Wang:2014:ESK

[988] Zhi Wang, Jinsong Han, Wei Xi, and Jizhong Zhao. Efficient and secure key extraction using channel state information. The Journal of Supercomputing, 70(3):1537–1554, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1247-1.

Pani:2014:RTB

[989] Danilo Pani, Alessandro Pani, and Luigi Raffo. Real-time blind audio source separation: performance assessment on an advanced digital signal processor. The Journal of Supercomputing, 70(3):1555–1576, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1252-4.

Xie:2014:DIP

[990] Xia Xie, Wenzhi Cao, Hai Jin, Xijiang Ke, and Shuwen Luo. Design and implementation of process-aware predictive scheduling scheme for virtual machine. The Journal of Supercomputing, 70(3):1577–1587, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1254-2.

Basanta-Val:2014:SDG

[991] Pablo Basanta-Val and Marisol García-Valls. A simple distributed garbage collector for distributed real-time Java. The Journal of Supercomputing, 70(3): 1588–1616, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1259-x.

Saleemi:2014:EES

[992] M. Mohsin Saleemi, Natalia Díaz Rodríguez, and Johan Lilius. Erratum to: “Exploiting smart spaces for interactive TV applications development”. The Journal of Supercomputing, 70(3):1617, December 2014. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
014-1296-5.pdf. See [971].

Liu:2015:ECB

[993] Zhao Liu, Jianxi Fan, and Xiaohua Jia. Embedding complete binary trees into parity cubes. The Journal of Supercomputing, 71(1):1–27, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1274-y.

El-Boghdadi:2015:CPW

[994] Hatem M. El-Boghdadi. On the computational power of WECPAR. The Journal of Supercomputing, 71 (1):28–44, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1275-x.

Hosseinimotlagh:2015:SSE

[995] Seyedmehdi Hosseinimotlagh, Farshad Khunjush, and Rashidaldin Samadzadeh. SEATS: smart energyaware task scheduling in real-time cloud computing. The Journal of Supercomputing, 71(1):45–66, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1276-9.

Valls:2015:PCE

[996] Joan J. Valls, Alberto Ros, Julio Sahuquillo, and Maria E. Gomez. PS-Cache: an energy-efficient cache design for chip multiprocessors. The Journal of Supercomputing, 71(1):67–86, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1288-5.

Hasanzadeh:2015:DOG

[997] Mohammad Hasanzadeh and Mohammad Reza Meybodi. Distributed optimization grid resource discovery. The Journal of Supercomputing, 71 (1):87–120, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1289-4.

Daryanavard:2015:FPA

[998] H. Daryanavard, M. Eshghi, and A. Jahanian. A fast placement algorithm for embedded just-in-time reconfigurable extensible processing platform.

The Journal of Supercomputing, 71 (1):121–143, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1290-y.

Cheng:2015:SCP

[999] Chien-Fu Cheng and Hsien-Chun Liao. The k-set consensus problem with weight consideration. The Journal of Supercomputing, 71(1):144–161, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1291-x.

Pallipuram:2015:SVO

[1000] Vivek K. Pallipuram, Melissa C. Smith, Nilim Sarma, Ranajeet Anand, Edwin Weill, and Karan Sapra. Subjective versus objective: classifying analytical models for productive heterogeneous performance prediction. The Journal of Supercomputing, 71(1):162–201, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1292-9.

Lynar:2015:BGQ

[1001] Timothy M. Lynar and Mark D.

Nelson. Blue Gene/Q defragmentation for energy waste minimisation. The Journal of Supercomputing, 71 (1):202–216, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1293-8.

Huang:2015:TRA

[1002] Kuo-Chan Huang, Ying-Lin Tsai, and Hsiao-Ching Liu. Task ranking and allocation in list-based workflow scheduling on parallel computing platform. The Journal of Supercomputing, 71 (1):217–240, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1294-7.

Singh:2015:QQA

[1003] Sukhpal Singh and Inderveer Chana. QRSF: QoS-aware resource scheduling framework in cloud computing. The Journal of Supercomputing, 71 (1):241–292, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1295-6.

Ahmed:2015:SGS

[1004] Munib Ahmed, Ishfaq Ahmad, and Mohammad Saad Ahmad. A survey of genome sequence assembly techniques and algorithms using highperformance computing. The Journal of Supercomputing, 71(1):293–339, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1297-4.

Zhao:2015:IND

[1005] Jie Zhao, Rongcai Zhao, Xi Chen, and Bo Zhao. An improved nonlinear data dependence test. The Journal of Supercomputing, 71(1):340–368, January 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1298-3.

Zhao:2015:EST

[1006] Di Zhao and Jinhang Yu. Efficiently solving tri-diagonal system by chunked cyclic reduction and single-GPU shared memory. The Journal of Supercomputing, 71(2):369–390, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1299-2.

Ma:2015:EEP

[1007] Yung-Cheng Ma, Wen-Shih Chao, and Tse-An Liu. Enabling energyproportional computing on instructionlevel parallel processors. The Journal of Supercomputing, 71(2):391–447, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1301-z.

Cordeschi:2015:EEA

[1008] Nicola Cordeschi, Mohammad Shojafar, Danilo Amendola, and Enzo Baccarelli. Energy-efficient adaptive networked datacenters for the QoS support of real-time applications. The Journal of Supercomputing, 71(2): 448–478, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1305-8.

Dominguez:2015:MFI

[1009] Carlos Domínguez, Houcine Hassan, Alfons Crespo, and José Albaladejo. Multicore and FPGA implementations of emotional-based agent architectures. The Journal of Supercomputing, 71 (2):479–507, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1307-6.

Parsa:2015:MFI

[1010] Saeed Parsa and Mehdi Sakhaei-nia. Modeling flow information of loops using compositional condition of controls. The Journal of Supercomputing, 71 (2):508–536, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1308-5.

Asanya:2015:DPQ

[1011] Charles Asanya and Ratan Guha. Direct private query in location-based services with GPU run time analysis. The Journal of Supercomputing, 71 (2):537–573, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1309-4.

Han:2015:DSS

[1012] Sangchul Han, Minkyu Park, Xuefeng Piao, and Moonju Park. A dual speed scheme for dynamic voltage scaling on real-time multiprocessor systems. The Journal of Supercomputing, 71 (2):574–590, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1310-y.

Vilaplana:2015:HPC

[1013] Jordi Vilaplana, Francesc Solsona, Francesc Abella, Josep Cuadrado, Ivan Teixidó, Jordi Mateo, and Josep Rius. H-PC: a cloud computing tool for supervising hypertensive patients. The Journal of Supercomputing, 71 (2):591–612, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1312-9.

Cebrian-Marquez:2015:AHU

[1014] Gabriel Cebrián-Márquez, José Luis Hernández-Losada, José Luis Martínez, Pedro Cuenca, Minhao Tang, and Jiangtao Wen. Accelerating HEVC using heterogeneous platforms. The Journal of Supercomputing, 71(2): 613–628, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1313-8.

Jiang:2015:TSD

[1015] Jincheng Jiang and Lixin Wu. Twostage distributed parallel algorithm with message passing interface for maximum flow problem. The Journal of Supercomputing, 71(2):629–647, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
014-1314-7.pdf.

Iturriaga:2015:PLS

[1016] Santiago Iturriaga, Sergio Nesmachnow, Francisco Luna, and Enrique Alba. A parallel local search in CPU/GPU for scheduling independent tasks on large heterogeneous computing systems. The Journal of Supercomputing, 71(2):648–672, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1315-6.

Nourikhah:2015:MPM

[1017] Hossein Nourikhah, Mohammad Kazem Akbari, and Mohammad Kalantari. Modeling and predicting measured response time of cloud-based web services using long-memory time series. The Journal of Supercomputing, 71 (2):673–696, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1317-4.

Bistouni:2015:SCN

[1018] Fathollah Bistouni and Mohsen Jahanshahi. Scalable crossbar network: a non-blocking interconnection network for large-scale systems. The Journal of Supercomputing, 71(2):697–728, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1319-2.

Castillo:2015:FAM

[1019] Emilio Castillo, Cristóbal Camarero, Ana Borrego, and Jose Luis Bosque. Financial applications on multi-CPU and multi-GPU architectures. The Journal of Supercomputing, 71(2): 729–739, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1316-5.

Benner:2015:ELS

[1020] Peter Benner, Alfredo Remón, Ernesto Dufrechou, Pablo Ezzatti, and Enrique S. Quintana-Ortí. Extending lyapack for the solution of band Lyapunov equations on hybrid CPU– GPU platforms. The Journal of Supercomputing, 71(2):740–750, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1322-7.

Ramiro:2015:MHP

[1021] Carla Ramiro, Antonio M. Vidal, and Alberto Gonzalez. MIMOPack: a high-performance computing library for MIMO communication systems. The Journal of Supercomputing, 71 (2):751–760, February 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1328-1.

Kim:2015:UWC

[1022] Youngjae Kim and Raghul Gunasekaran. Understanding I/O workload characteristics of a peta-scale storage system. The Journal of Supercomputing, 71(3):761–780, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11227-014-
                                        1321-8.

Ros:2015:DDL

[1023] Alberto Ros and Manuel E. Acacio. DASC-DIR: a low-overhead coherence directory for many-core processors. The Journal of Supercomputing, 71 (3):781–807, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1325-4.

Wang:2015:AST

[1024] Chun-Kun Wang and Peng-Sheng Chen. Automatic scoping of task clauses for the OpenMP tasking model. The Journal of Supercomputing, 71 (3):808–823, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1326-3.

Norouzi:2015:RAA

[1025] Mohammad Norouzi and Ali Jannesari. Resource and application-aware resource discovery in computing environments. The Journal of Supercomputing, 71(3):824–839, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1327-2.

Kim:2015:IPP

[1026] Seong Jo Kim, Yuanrui Zhang, Seung Woo Son, Mahmut Kandemir, Wei keng Liao, Rajeev Thakur, and Alok Choudhary. IOPro: a parallel I/O profiling and visualization framework for high-performance storage systems. The Journal of Supercomputing, 71(3):840–870, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1329-0.

Beigy:2015:LAB

[1027] Hamid Beigy and M. R. Meybodi. A learning automata-based adaptive uniform fractional guard channel algorithm. The Journal of Supercomputing, 71(3):871–893, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1330-7.

Kim:2015:NPT

[1028] Jong-Seok Kim, Hyeong-Ok Lee, Mihye Kim, and Sung Won Kim. The new Petersen-torus networks. The Journal of Supercomputing, 71(3):894–908, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1342-3.

Bampis:2015:RTI

[1029] L. Bampis, C. Iakovidou, S. A.

Chatzichristofis, Y. S. Boutalis, and A. Amanatiadis. Real-time indexing for large image databases: color and edge directivity descriptor on GPU. The Journal of Supercomputing, 71 (3):909–937, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1343-2.

Duran:2015:SOB

[1030] Ahmet Duran, M. Serdar Celebi, Senol Piskin, and Mehmet Tuncel. Scalability of OpenFOAM for bio-medical flow simulations. The Journal of Supercomputing, 71(3):938–951, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1344-1.

Yang:2015:FPS

[1031] Jinn-Shyong Yang, Meng-Ru Wu, Jou-Ming Chang, and Yu-Huei Chang. A fully parallelized scheme of constructing independent spanning trees on Möbius cubes. The Journal of Supercomputing, 71(3):952–965, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1346-z.

Pascual:2015:LAP

[1032] Jose A. Pascual, Jose Miguel-Alonso, and Jose A. Lozano. Locality-aware policies to improve job scheduling on 3D tori. The Journal of Supercomputing, 71(3):966–994, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1347-y.

Tosun:2015:AMA

[1033] Suleyman Tosun, Ozcan Ozturk, Erencan Ozkan, and Meltem Ozen. Application mapping algorithms for meshbased network-on-chip architectures. The Journal of Supercomputing, 71 (3):995–1017, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1348-x.

Naserian:2015:CAJ

[1034] Elahe Naserian, Seyed Mohammad Ghoreyshi, Hossein Shafiei, Payam Mousavi, and Ahmad Khonsari. Cooling aware job migration for reducing cost in cloud environment. The Journal of Supercomputing, 71(3):1018–1037, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1349-9.

Amiri-Zarandi:2015:PEG

[1035] Mohammad Amiri-Zarandi, Farshad Safaei, and Milad Roozikhar. Performance evaluation of generic multistage interconnection networks with blocking and back-pressure mechanism. The Journal of Supercomputing, 71 (3):1038–1066, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1350-3.

Zhang:2015:NSS

[1036] Jing Zhang, Li Xu, Shuming Zhou, and Xiucai Ye. A novel sleep scheduling scheme in green wireless sensor networks. The Journal of Supercomputing, 71(3):1067–1094, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1354-z.

Chen:2015:PST

[1037] Chia-Jung Chen and Rong-Guey Chang. A priority scheduling for TM pathologies. The Journal of Supercomputing, 71(3):1095–1115, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1356-x.

Garcia:2015:FAR

[1038] Marina García, Enrique Vallejo, Ramón Beivide, Cristóbal Camarero, Mateo Valero, Germán Rodríguez, and Cyriel Minkenberg. On-the-fly adaptive routing for dragonfly interconnection networks. The Journal of Supercomputing, 71(3):1116–1142, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1357-9.

Kianfar:2015:NMA

[1039] K. Kianfar, G. Moslehi, and R. Yahyapour. A novel metaheuristic algorithm and utility function for QoS based scheduling in user-centric grid systems. The Journal of Supercomputing, 71(3): 1143–1162, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1358-8.

Cocana-Fernandez:2015:EEA

[1040] Alberto Cocaña-Fernández, Jose Ranilla, and Luciano Sánchez. Energy-efficient allocation of computing node slots in HPC clusters through parameter learning and hybrid genetic fuzzy system modeling. The Journal of Supercomputing, 71(3):1163–1174, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1320-9.

Avila-George:2015:ESG

[1041] Himer Avila-George, Jose Torres-Jimenez, Nelson Rangel-Valdez, Abel Carrión, and Vicente Hernández. Erratum to: Supercomputing and grid computing on the verification of covering arrays. The Journal of Supercomputing, 71(3):1175–1176, March 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
015-1389-9.pdf. See [392].

El-Boghdadi:2015:DWR

[1042] Hatem M. El-Boghdadi. Dynamicwidth reconfigurable parallel prefix circuits. The Journal of Supercomputing, 71(4):1177–1195, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1270-2.

Chunlin:2015:CEA

[1043] Li Chunlin and Li LaYuan. Cost and energy aware service provisioning for mobile client in cloud computing environment. The Journal of Supercomputing, 71(4):1196–1223, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1345-0.

Azizi:2015:FTR

[1044] Sadoon Azizi, Farshad Safaei, and Milad Roozikhar. A fault-tolerant routing algorithm in HyperX topology based on unsafety vectors. The Journal of Supercomputing, 71(4):1224–1248, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1355-y.

Farouk:2015:CEC

[1045] Abeer Farouk and Hatem M. El-Boghdadi. A cost-efficient congestion management methodology for fattrees using traffic pattern detection. The Journal of Supercomputing, 71 (4):1249–1276, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1359-7.

Zarrabi:2015:GSA

[1046] Amirreza Zarrabi, Khairulmizam Samsudin, and Ettikan K. Karuppiah. Gravitational search algorithm using CUDA: a case study in highperformance metaheuristics. The Journal of Supercomputing, 71(4):1277– 1296, April 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1360-1.

Chen:2015:AMW

[1047] Wei Chen, Young Choon Lee, Alan Fekete, and Albert Y. Zomaya. Adaptive multiple-workflow scheduling with task rearrangement. The Journal of Supercomputing, 71(4):1297–1317, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11227-014-
1361-0.

Djenouri:2015:GBB

[1048] Youcef Djenouri, Ahcene Bendjoudi, Malika Mehdi, Nadia Nouali-Taboudjemat, and Zineb Habbas. GPU-based bees swarm optimization for association rules mining. The Journal of Supercomputing, 71(4): 1318–1344, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1366-8.

Couturier:2015:SMA

[1049] Raphaël Couturier and Ziane Khodja Lilia. A scalable multisplitting algorithm to solve large sparse linear systems. The Journal of Supercomputing, 71(4):1345–1356, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1367-7.

Jiang:2015:TSG

[1050] Yi-Syuan Jiang and Wei-Mei Chen.

Task scheduling for grid computing systems using a genetic algorithm. The Journal of Supercomputing, 71 (4):1357–1377, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1368-6.

Ahmad:2015:OUM

[1051] Mahmood Ahmad, Zeeshan Pervez, Taechoong Cheong, and Sungyoung Lee. Oblivious user management for cloud-based data synchronization. The Journal of Supercomputing, 71 (4):1378–1400, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1369-5.

Arkian:2015:CBV

[1052] Hamid Reza Arkian, Reza Ebrahimi Atani, Abolfazl Diyanat, and Atefe Pourkhalili. A cluster-based vehicular cloud architecture with learning-based resource management. The Journal of Supercomputing, 71(4):1401–1426, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1370-z.

Aron:2015:HHA

[1053] Rajni Aron, Inderveer Chana, and Ajith Abraham. A hyper-heuristic approach for resource provisioningbased scheduling in grid environment. The Journal of Supercomputing, 71 (4):1427–1450, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1373-9.

deBlanche:2015:ACM

[1054] Andreas de Blanche and Thomas Lundqvist. Addressing characterization methods for memory contention aware co-scheduling. The Journal of Supercomputing, 71(4):1451–1483, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1374-8.

Nikounia:2015:GMG

[1055] Seyed Hossein Nikounia and Siamak Mohammadi. Gem5v: a modified gem5 for simulating virtualized systems. The Journal of Supercomputing, 71(4):1484–1504, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1375-7.

Panda:2015:ETS

[1056] Sanjaya K. Panda and Prasanta K.

Jana. Efficient task scheduling algorithms for heterogeneous multi-cloud environment. The Journal of Supercomputing, 71(4):1505–1533, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1376-6.

Salkhordeh:2015:OSL

[1057] Reza Salkhordeh, Hossein Asadi, and Shahriar Ebrahimi. Operating system level data tiering using online workload characterization. The Journal of Supercomputing, 71(4):1534–1562, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1377-0.

Zhang:2015:EGC

[1058] Tao Zhang, Jingjie Zhang, Wei Shu, Min-You Wu, and Xiaoyao Liang. Efficient graph computation on hybrid CPU and GPU systems. The Journal of Supercomputing, 71(4):1563–1586, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1378-z.

Taheri:2015:BCU

[1059] Erfan Taheri and Mohammad Izadi. Byzantine consensus for unknown dynamic networks. The Journal of Supercomputing, 71(4):1587–1603, April 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1379-y.

Li:2015:SSS

[1060] Kin Fun Li and Wendy Rahayu. Special section on support technology and architecture for networked and distributed applications in big data era. The Journal of Supercomputing, 71(5):1605–1606, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-

015-1424-x.pdf. See erratum [1140].

Xhafa:2015:PAA

[1061] Fatos Xhafa, Jianglang Feng, Yinghui Zhang, Xiaofeng Chen, and Jin Li. Privacy-aware attribute-based PHR sharing with user accountability in cloud computing. The Journal of Supercomputing, 71(5):1607–1619, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1253-3.

Palmieri:2015:EOD

[1062] Francesco Palmieri, Sergio Ricciardi, Ugo Fiore, Massimo Ficco, and Aniello Castiglione. Energy-oriented denial of service attacks: an emerging menace for large cloud infrastructures. The Journal of Supercomputing, 71 (5):1620–1641, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1242-6.

Duolikun:2015:EED

[1063] Dilawaer Duolikun, Tomoya Enokido, Ailixier Aikebaier, and Makoto Takizawa. Energy-efficient dynamic clusters of servers. The Journal of Supercomputing, 71(5):1642–1656, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1261-3.

Lin:2015:AJC

[1064] Jia-Chun Lin, Fang-Yie Leu, and Ying-Ping Chen. Analyzing job completion reliability and job energy consumption for a heterogeneous MapReduce cluster under different intermediatedata replication policies. The Journal of Supercomputing, 71(5):1657–1677, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1286-7.

Abid:2015:NDA

[1065] Besem Abid, Tien Trung Nguyen, and Hamida Seba. New data aggregation approach for time-constrained wireless sensor networks. The Journal of Supercomputing, 71(5):1678–1693, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1241-7.

Dhurandher:2015:ERB

[1066] Sanjay K. Dhurandher, Deepak Kumar Sharma, Isaac Woungang, and Aakanksha Saini. Efficient routing based on past information to predict the future location for message passing in infrastructure-less opportunistic networks. The Journal of Supercomputing, 71(5):1694–1711, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1243-5.

Stenico:2015:MNT

[1067] Jeferson Wilian de Godoy Stênico and Lee Luan Ling. Modern network traffic modeling based on binomial multiplicative cascades. The Journal of Supercomputing, 71(5):1712–1735, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1255-1.

Wang:2015:AFD

[1068] Youwei Wang, Can Ma, Weiping Wang, and Dan Meng. An approach of fast data manipulation in HDFS with supplementary mechanisms. The Journal of Supercomputing, 71(5):1736–1753, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1287-6.

Pop:2015:DSA

[1069] Florin Pop, Ciprian Dobre, Valentin Cristea, Nik Bessis, Fatos Xhafa, and Leonard Barolli. Deadline scheduling for aperiodic tasks in inter-cloud environments: a new approach to resource management. The Journal of Supercomputing, 71(5):1754–1765, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1285-8.

Nakata:2015:PBP

[1070] Susumu Nakata and Yasuaki Sakamoto.

Particle-based parallel fluid simulation in three-dimensional scene with implicit surfaces. The Journal of Supercomputing, 71(5):1766–1775, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1323-6.

Wu:2015:HTM

[1071] Jingjin Wu, Xuanxing Xiong, and Zhiling Lan. Hierarchical task mapping for parallel applications on supercomputers. The Journal of Supercomputing, 71(5):1776–1802, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1324-5.

Alvarez-Bermejo:2015:PAS

[1072] J. A. Álvarez-Bermejo, J. M. Arrufat, and J. A. López-Ramos. A parallel algorithm for secure multicast. The Journal of Supercomputing, 71 (5):1803–1816, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1336-1.

Vilaplana:2015:SPS

[1073] Jordi Vilaplana, Jordi Mateo, Ivan Teixidó, Francesc Solsona, Francesc Giné, and Concepció Roig. An SLA and power-saving scheduling consolidation strategy for shared and heterogeneous clouds. The Journal of Supercomputing, 71(5):1817–1832, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1351-2.

Llados:2015:RAM

[1074] Jordi Lladós, Fernando Guirado, Fernando Cores, Josep Lluís Lérida, and Cedric Notredame. Recovering accuracy methods for scalable consistency library. The Journal of Supercomputing, 71(5):1833–1845, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
014-1362-z.pdf.

Diaz-Gracia:2015:INH

[1075] N. Díaz-Gracia, A. Cocaña-Fernández, M. Alonso-González, F. J. Martínez-Zaldívar, R. Cortina, V. M. García-Mollá, P. Alonso, J. Ranilla, and A. M. Vidal. Improving NNMFPACK with heterogeneous and efficient kernels for β-divergence metrics. The Journal of Supercomputing, 71(5):1846–1856, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1363-y.

Portales:2015:PMD

[1076] Cristina Portalés, Juan M. Orduña, and Pedro Morillo. Parallelization of a method for dense 3D object reconstruction in structured light scanning. The Journal of Supercomputing, 71 (5):1857–1868, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1364-x.

Artes:2015:ECE

[1077] Tomàs Artés, Andrés Cencerrado, Ana Cortés, and Tomàs Margalef. Enhancing computational efficiency on forest fire forecasting by time-aware genetic algorithms. The Journal of Supercomputing, 71(5):1869–1881, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1365-9.

Pinol:2015:SBP

[1078] P. Piñol, H. Migallón, O. López-Granado, and M. P. Malumbres. Slicebased parallel approach for HEVC encoder. The Journal of Supercomputing, 71(5):1882–1892, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1371-y.

Leon:2015:EPP

[1079] G. León, J. M. Molero, E. M. Garzón, I. García, A. Plaza, and E. S. Quintana-Ortí. Exploring the performancepower-energy balance of low-power multicore and manycore architectures for anomaly detection in remote sensing. The Journal of Supercomputing, 71(5):1893–1906, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1372-x.

Aci:2015:HCC

[1080] Çigdem Inan Aci and Mehmet Fatih Akay. A hybrid congestion control algorithm for broadcast-based architectures with multiple input queues. The Journal of Supercomputing, 71 (5):1907–1931, May 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1384-1.

Jun:2015:SMA

[1081] Sanghoon Jun, Daehoon Kim, Mina Jeon, Seungmin Rho, and Eenjun Hwang. Social mix: automatic music recommendation and mixing scheme based on social network analysis. The Journal of Supercomputing, 71 (6):1933–1954, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1182-1.

Kim:2015:TAE

[1082] Dong Kyoo Kim and Yang Sun Lee. Time-of-arrival estimation through WLAN physical layer systems. The Journal of Supercomputing, 71(6): 1955–1974, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1203-0.

Tsai:2015:IPA

[1083] Ming-Fong Tsai, Po-Ching Wang, Ce-Kuen Shieh, Wen-Shyang Hwang, Naveen Chilamkurti, Seungmin Rho, and Yang Sun Lee. Improving positioning accuracy for VANET in real city environments. The Journal of Supercomputing, 71(6):1975–1995, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1215-9.

Lee:2015:AGD

[1084] Kwang-Soo Lee and Hyuk-Jun Kwon.

Analyzing the geographic distribution of major medical equipment with smart geographic system. The Journal of Supercomputing, 71(6):1996–2019, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1218-6.

Lee:2015:IUL

[1085] Kyuhwan Lee, Sungpyo Hong, and Hangbae Chang. ICT use level and performance evaluation on the social network environment: Korea case study. The Journal of Supercomputing, 71 (6):2020–2034, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1220-z.

Jiang:2015:FHR

[1086] Feng Jiang, Seungmin Rho, Bo-Wei Chen, Xiaodan Du, and Debin Zhao. Face hallucination and recognition in social network services. The Journal of Supercomputing, 71(6):2035–2049, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1257-z.

Cho:2015:OAO

[1087] Myeongjin Cho, Youngsun Han, Minseong Kim, and Seon Wook Kim. O2WebCL: an automatic OpenCLto-WebCL translator for high performance web computing. The Journal of Supercomputing, 71(6):2050–2065, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1260-4.

Chen:2015:PMD

[1088] Yong Chen, Yin Lu, Prathamesh Amritkar, Rajeev Thakur, and Yu Zhuang. Performance model-directed data sieving for high-performance I/O. The Journal of Supercomputing, 71(6): 2066–2090, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1277-8.

Seo:2015:DAA

[1089] Sanghyun Seo, Hakjeon Bang, and Hun-Joo Lee. Delay analysis for ad hoc multiplayer game of socially clustered mobile users. The Journal of Supercomputing, 71(6):2091–2100, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1304-9.

Kim:2015:PVS

[1090] Do-Hyun Kim, Kyoung-Ho Choi, Ki-Joune Li, and Yang-Sun Lee. Performance of vehicle speed estimation using wireless sensor networks: a regionbased approach. The Journal of Supercomputing, 71(6):2101–2120, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1306-7.

Kim:2015:UHS

[1091] Jin Kim and Sang Oh Park. U-Health Smart system architecture and ontology model. The Journal of Supercomputing, 71(6):2121–2137, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1334-3.

Ji:2015:HEV

[1092] Xiangyang Ji, Sam Kwong, Bo-Wei Chen, and Seungmin Rho. Highefficient video compression for social multimedia distribution. The Journal of Supercomputing, 71(6):2138–2152, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1340-5.

Hashmi:2015:UPL

[1093] S. M. Usman Hashmi, Imran Shafi, Jamil Ahmad, Anand Paul, and Sang Oh Park. Using physical layer clock recovery to augment application layer time synchronization. The Journal of Supercomputing, 71 (6):2153–2176, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1388-x.

Kim:2015:TVB

[1094] Sejin Kim, Hyerim Kim, and Sung Jo Kim. A trust value-based scheme for efficient contents sharing over mobile P2P networks. The Journal of Supercomputing, 71(6):2177–2189, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1393-0.

Huang:2015:URM

[1095] Lin Huang, Shangguang Wang, Ching-Hsien Hsu, Juanjuan Zhang, and Fangchun Yang. Using reputation measurement to defend mobile social networks against malicious feedback ratings. The Journal of Supercomputing, 71(6):2190–2203, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1432-x.

Vinas:2015:DAM

[1096] Moisés Viñas, Zeki Bozkus, Basilio B. Fraguela, Diego Andrade, and Ramón Doallo. Developing adaptive multidevice applications with the Heterogeneous Programming Library. The Journal of Supercomputing, 71(6): 2204–2220, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1352-1.

Beltran:2015:APM

[1097] Marta Beltrán. Automatic provisioning of multi-tier applications in cloud computing environments. The Journal of Supercomputing, 71(6):2221–2250, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1380-5.

Jiang:2015:AAG

[1098] Guiyuan Jiang, Jigang Wu, Siew-Kei Lam, Thambipillai Srikanthan, and Jizhou Sun. Algorithmic aspects of graph reduction for hardware/software partitioning. The Journal of Supercomputing, 71(6):2251–2274, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1381-4.

Seo:2015:OMC

[1099] Junsang Seo, Myeongsu Kang, Cheol Hong Kim, and Jong-Myon Kim. An optimal many-core model-based supercomputing for accelerating video-equipped fire detection. The Journal of Supercomputing, 71(6):2275–2308, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1382-3.

Khan:2015:AST

[1100] Behram Khan, Daniel Goodman, Salman Khan, Will Toms, Paolo Faraboschi, Mikel Luján, and Ian Watson. Architectural support for task scheduling: hardware scheduling for dataflow on NUMA systems. The Journal of Supercomputing, 71(6):2309–2338, June 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1383-2.

Gomez:2015:HBA

[1101] C. Gómez, F. Gilabert, M. E. Gómez, P. López, and J. Duato. A HoLblocking aware mechanism for selecting the upward path in fat-tree topologies. The Journal of Supercomputing, 71(7):2339–2364, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1303-x.

Yildirim:2015:CSP

[1102] Ahmet Artu Yildirim and Dan Watson. A comparative study of the parallel wavelet-based clustering algorithm on three-dimensional dataset. The Journal of Supercomputing, 71(7):2365–2380, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1385-0.

Li:2015:AMR

[1103] Jiansen Li, Jianqi Sun, Ying Song, and Jun Zhao. Accelerating MRI reconstruction via three-dimensional dual-dictionary learning using CUDA.

The Journal of Supercomputing, 71 (7):2381–2396, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1386-z.

Toloo:2015:EEL

[1104] Mehdi Toloo, Ameneh Zandi, and Ali Emrouznejad. Evaluation efficiency of large-scale data set with negative data: an artificial neural network approach. The Journal of Supercomputing, 71 (7):2397–2411, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1387-y.

Shahrivari:2015:HPP

[1105] Saeed Shahrivari and Saeed Jalili.

High-performance parallel frequent subgraph discovery. The Journal of Supercomputing, 71(7):2412–2432, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1391-2.

Su:2015:AGP

[1106] Huayou Su, Xing Cai, Mei Wen, and Chunyuan Zhang. An analytical GPU performance model for 3D stencil computations from the angle of data traffic. The Journal of Supercomputing, 71(7):2433–2453, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1392-1.

Kim:2015:FEE

[1107] Jaeyoung Kim, Myeongsu Kang, Md Shohidul Islam, Cheol-Hong Kim, and Jong-Myon Kim. A fast and energy-efficient Hamming decoder for software-defined radio using graphics processing units. The Journal of Supercomputing, 71(7):2454–2472, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1396-x.

Ahmad:2015:VMM

[1108] Raja Wasim Ahmad, Abdullah Gani, Siti Hafizah Ab. Hamid, Muhammad Shiraz, Feng Xia, and Sajjad A. Madani. Virtual machine migration in cloud data centers: a review, taxonomy, and open research issues. The Journal of Supercomputing, 71 (7):2473–2515, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1400-5.

Meneses:2015:CCA

[1109] Esteban Meneses and Laxmikant V. Kalé. Camel: collective-aware message logging. The Journal of Supercomputing, 71(7):2516–2538, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1402-3.

Elsayed:2015:NPE

[1110] Essam Elsayed and Hatem M. El-Boghdadi. A novel power-efficient multi-operand digit-multiplier using reconfiguration and clock gating. The Journal of Supercomputing, 71(7): 2539–2564, July 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1403-2.

Fan:2015:ECP

[1111] Ming Fan, Rong Rong, Shuo Liu, and Gang Quan. Energy calculation for periodic multi-core scheduling in system thermal steady state with consideration of leakage and temperature dependency. The Journal of Supercomputing, 71(7):2565–2584, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1405-0.

Touzene:2015:AAB

[1112] Abderezak Touzene and Khaled Day.

All-to-all broadcasting in torus Network on Chip. The Journal of Supercomputing, 71(7):2585–2596, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1406-z.

Stojanovic:2015:DMI

[1113] Igor Z. Stojanovic, Milica D. Jovanovic, and Goran Lj. Djordjevic. Dual-mode inter-router communication channel for deflection-routed networks-on-chip. The Journal of Supercomputing, 71 (7):2597–2613, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1407-y.

Villar:2015:OCC

[1114] Juan A. Villar, Francisco J. Andujar, Francisco J. Alfaro, Jose L. Sanchez, and Jose Duato. Optimizing the configuration of combined highradix switches. The Journal of Supercomputing, 71(7):2614–2643, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1408-x.

Kelefouras:2015:MSM

[1115] Vasilios Kelefouras, Angeliki Kritikakou, Elissavet Papadima, and Costas Goutis. A methodology for speeding up matrix vector multiplication for single/multi-core architectures. The Journal of Supercomputing, 71(7):2644–2667, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1409-9.

Kotiyal:2015:RLB

[1116] Saurabh Kotiyal, Himanshu Thapliyal, and Nagarajan Ranganathan. Reversible logic based multiplication computing unit using binary tree data structure. The Journal of Supercomputing, 71(7):2668–2693, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1410-3.

Karim:2015:SSO

[1117] Naila Karim, Khalid Latif, Zahid Anwar, Sharifullah Khan, and Amir Hayat. Storage schema and ontologyindependent SPARQL to HiveQL translation. The Journal of Supercomputing, 71(7):2694–2719, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1411-2.

Lai:2015:LAD

[1118] Zhiquan Lai, King Tin Lam, Cho-Li Wang, and Jinshu Su. Latencyaware DVFS for efficient power state transitions on many-core architectures. The Journal of Supercomputing, 71 (7):2720–2747, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1415-y.

Wang:2015:DHR

[1119] Zhanjie Wang and Xianxian Su.

Dynamically hierarchical resourceallocation algorithm in cloud computing environment. The Journal of Supercomputing, 71(7):2748–2766, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1416-x.

Chen:2015:DPC

[1120] Hon-Chan Chen, Tzu-Liang Kung, and Li-Yen Hsu. 2-disjoint-path-coverable panconnectedness of crossed cubes. The Journal of Supercomputing, 71 (7):2767–2782, July 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1417-9.

Sarbazi-Azad:2015:AMS

[1121] H. Sarbazi-Azad, N. Bagherzadeh, and G. Jaberipour. Advances in multicore systems architectures. The Journal of Supercomputing, 71(8):2783–2786, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        015-1487-8.pdf.

Pakdaman:2015:ICP

[1122] Farhad Pakdaman, Abbas Mazloumi, and Mehdi Modarressi. Integrated circuit-packet switching NoC with efficient circuit setup mechanism. The Journal of Supercomputing, 71(8): 2787–2807, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-014-1337-0.

Falahati:2015:PEP

[1123] Hajar Falahati, Shaahin Hessabi, Mania Abdi, and Amirali Baniasadi. Power-efficient prefetching on GPG-PUs. The Journal of Supercomputing, 71(8):2808–2829, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s11227-014-1331-
                                        6.

Hu:2015:DAM

[1124] Wen-Hsiang Hu, Chifeng Wang, and Nader Bagherzadeh. Design and analysis of a mesh-based wireless networkon-chip. The Journal of Supercomputing, 71(8):2830–2846, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1341-4.

Valls:2015:PDS

[1125] Joan J. Valls, Alberto Ros, Julio Sahuquillo, and María E. Gómez. PS directory: a scalable multilevel directory cache for CMPs. The Journal of Supercomputing, 71(8):2847–2876, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1332-5.

Daneshtalab:2015:ODA

[1126] Masoud Daneshtalab, Masoumeh Ebrahimi, Sergei Dytckov, and Juha Plosila. In-order delivery approach for 2D and 3D NoCs. The Journal of Supercomputing, 71(8):2877–2899, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1339-y.

Karami:2015:SPA

[1127] Ali Karami, Farshad Khunjush, and Seyyed Ali Mirsoleimani. A statistical performance analyzer framework for OpenCL kernels on Nvidia GPUs. The Journal of Supercomputing, 71(8): 2900–2921, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1338-z.

Wu:2015:DHD

[1128] Zhendong Wu, Kai Lu, Xiaoping Wang, Xu Zhou, and Chen Chen. Detecting harmful data races through parallel verification. The Journal of Supercomputing, 71(8):2922–2943, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1418-8.

Fe:2015:EON

[1129] Jorge D. Fe, Ramón J. Aliaga, and Rafael Gadea-Gironés. Evolutionary optimization of neural networks with heterogeneous computation: study and implementation. The Journal of Supercomputing, 71(8):2944–2962, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1419-7.

Chen:2015:FEC

[1130] Sui Chen, Greg Bronevetsky, Bin Li, Marc Casas Guix, and Lu Peng. A framework for evaluating comprehensive fault resilience mechanisms in numerical programs. The Journal of Supercomputing, 71(8):2963–2984, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1422-z.

Silva:2015:SSS

[1131] Francisco Airton Silva, Paulo Maciel, and Rubens Matos. SmartRank: a smart scheduling tool for mobile cloud computing. The Journal of Supercomputing, 71(8):2985–3008, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11227-015-
1423-y.

Li:2015:HAH

[1132] Bo Li, Yijian Pei, Hao Wu, and Bin Shen. Heuristics to allocate highperformance cloudlets for computation offloading in mobile ad hoc clouds. The Journal of Supercomputing, 71 (8):3009–3036, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1425-9.

Yang:2015:MSA

[1133] Qiangpeng Yang, Yu Zhou, Yao Yu, Jie Yuan, Xianglei Xing, and Sidan Du. Multi-step-ahead host load prediction using autoencoder and echo state networks in cloud computing. The Journal of Supercomputing, 71 (8):3037–3053, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1426-8.

Lv:2015:WNA

[1134] Fang Lv, Lei Liu, Hui min Cui, Lei Wang, Ying Liu, Xiao bing Feng, and Pen-Chung Yew. WiseThrottling: a new asynchronous task scheduler for mitigating I/O bottleneck in large-scale datacenter servers. The Journal of Supercomputing, 71(8):3054–3093, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1427-7.

Yu:2015:BVC

[1135] Zhigang Yu, Dong Xiang, and Xinyu Wang. Balancing virtual channel utilization for deadlock-free routing in torus networks. The Journal of Supercomputing, 71(8):3094–3115, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1428-6.

Dehghani:2015:FTH

[1136] Abbas Dehghani and Kamal Jamshidi. A fault-tolerant hierarchical hybrid mesh-based wireless network-on-chip architecture for multicore platforms. The Journal of Supercomputing, 71 (8):3116–3148, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1430-z.

Al-Ayyoub:2015:GBI

[1137] Mahmoud Al-Ayyoub, Ansam M. Abu-Dalo, Yaser Jararweh, Moath Jarrah, and Mohammad Al Sa’d. A GPUbased implementations of the fuzzy C-means algorithms for medical image segmentation. The Journal of Supercomputing, 71(8):3149–3162, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1431-y.

Arshad:2015:SAI

[1138] Hamed Arshad and Morteza Nikooghadam. Security analysis and improvement of two authentication and key agreement schemes for session initiation protocol. The Journal of Supercomputing, 71 (8):3163–3180, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1434-8.

Lin:2015:MDN

[1139] Chi Lin, Guowei Wu, Chang Wu Yu, and Lin Yao. Maximizing destructiveness of node capture attack in wireless sensor networks. The Journal of Supercomputing, 71(8):3181–3212, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1435-7.

Li:2015:ESS

[1140] Kin Fun Li and Wenny Rahayu. Erratum to: Special section on support technology and architecture for networked and distributed applications in big data era. The Journal of Supercomputing, 71(8):3213, August 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
015-1442-8.pdf. See [1060].

Chen:2015:GTA

[1141] Daniel Bo-Wei Chen, Wen Ji, and Yong Liu. Game theoretic analysis for large-scale networks and traffic data. The Journal of Supercomputing, 71(9): 3215–3216, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1500-2;   http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-015-1500-2.pdf.

Chen:2015:DBR

[1142] Xin Chen, Yuan Si, and Xudong Xiang. Delay-bounded resource allocation for femtocells exploiting the statistical multiplexing gain. The Journal of Supercomputing, 71(9):3217– 3236, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1494-9.

Xiao:2015:PCR

[1143] Liang Xiao, Yan Li, Jinliang Liu, and Yifeng Zhao. Power control with reinforcement learning in cooperative cognitive radio networks against jamming. The Journal of Supercomputing, 71(9): 3237–3257, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1420-1.

Kumar:2015:OCD

[1144] Neeraj Kumar, Rasmeet Singh Bali, and Rahat Iqbal. Optimized clustering for data dissemination using stochastic coalition game in vehicular cyber-physical systems. The Journal of Supercomputing, 71(9):3258– 3287, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1436-6.

Wang:2015:CGT

[1145] Zeng Wang, Bo Hu, Xin Wang, and Shanzhi Chen. Cooperative gametheoretic power control with a balancing factor in large-scale LTE networks: an energy efficiency perspective. The Journal of Supercomputing, 71(9): 3288–3300, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1454-4.

Wang:2015:SGT

[1146] Ru Wang and Wandong Cai. A sequential game-theoretic study of the retweeting behavior in Sina Weibo. The Journal of Supercomputing, 71(9): 3301–3319, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1456-2.

Wang:2015:PPB

[1147] Yuanzhuo Wang, Jingyuan Li, Qiang Liu, and Yan Ren. Prediction of purchase behaviors across heterogeneous social networks. The Journal of Supercomputing, 71(9):3320–3336, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1495-8;   http://link.springer.
com/content/pdf/10.1007/s11227-
015-1495-8.pdf.

Jiang:2015:GTB

[1148] Feng Jiang, K. Bharanitharan, Shovan Barma, and Hailiang Wang. Game theory based no-reference perceptual quality assessment for stereoscopic images. The Journal of Supercomputing, 71(9): 3337–3352, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1412-1.

Liu:2015:GTB

[1149] Shaohui Liu, Anand Paul, Guochao Zhang, and Gwanggil Jeon. A game theory-based block image compression method in encryption domain. The Journal of Supercomputing, 71(9): 3353–3372, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1413-0.

Wu:2015:WSC

[1150] Fuhui Wu, Qingbo Wu, and Yusong Tan. Workflow scheduling in cloud: a survey. The Journal of Supercomputing, 71(9):3373–3418, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1438-4.

Lotfi-Kamran:2015:PPG

[1151] Pejman Lotfi-Kamran. Per-packet global congestion estimation for fast packet delivery in networks-on-chip. The Journal of Supercomputing, 71(9): 3419–3439, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1439-3.

Zhao:2015:FFB

[1152] Di Zhao. Fast filter bank convolution for three-dimensional wavelet transform by shared memory on mobile GPU computing. The Journal of Supercomputing, 71(9):3440– 3455, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1443-7.

Chrust:2015:ALF

[1153] Marcin Chrust, Eric Laurendeau, and Luc Ostiguy. Accelerating low-fidelity aerodynamic codes on multi- and many-core architectures. The Journal of Supercomputing, 71(9):3456– 3481, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1444-6.

Rodriguez:2015:SCG

[1154] Alejandro Rodríguez, Alejandro León, and Germán Arroyo. SP-ChainMail: a GPU-based sparse parallel Chain-Mail algorithm for deforming medical volumes. The Journal of Supercomputing, 71(9):3482–3499, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1445-5.

Kim:2015:OSC

[1155] Ikjoon Kim, Jidong Zhai, Yan Li, and Wenguang Chen. Optimizing seam carving on multi-GPU systems for real-time content-aware image resizing. The Journal of Supercomputing, 71(9): 3500–3524, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1446-4.

Moon:2015:OHM

[1156] Sangwhan Moon, Jaehwan Lee, Xiling Sun, and Yang suk Kee. Optimizing the Hadoop MapReduce Framework with high-performance storage devices. The Journal of Supercomputing, 71(9): 3525–3548, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1447-3.

Modarressi:2015:LDS

[1157] Mehdi Modarressi and Hamid Sarbazi-Azad. Leveraging dark silicon to optimize networks-on-chip topology. The Journal of Supercomputing, 71(9): 3549–3566, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1448-2.

Yazdanpanah:2015:DSE

[1158] Fahimeh Yazdanpanah and Mohammad Alaei. Design space exploration of hardware task superscalar architecture. The Journal of Supercomputing, 71(9): 3567–3592, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1449-1.

Chen:2015:ALS

[1159] Ching-Wen Chen, Kuan-Lin Huang, and Yuh-Dauh Lyuu. Accelerating the least-square Monte Carlo method with parallel computing. The Journal of Supercomputing, 71(9):3593– 3608, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1451-7.

Stankovic:2015:SAP

[1160] Vladimir V. Stankovic and Nebojsa Z.

Milenkovic. Synchronization algorithm for predictors for SDRAM memories. The Journal of Supercomputing, 71(9): 3609–3636, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1452-6.

Mahfoudhi:2015:TPR

[1161] Adel Mahfoudhi and Walid Karamti.

Transformation process of RTS scheduling analysis requirements from UML/ MARTE to dynamic priority time Petri nets. The Journal of Supercomputing, 71(10):3637–3667, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1455-3.

Zhu:2015:OFT

[1162] Lei Zhu, Jianhua Gu, Yunlan Wang, and Tianhai Zhao. Optimizing the fault-tolerance overheads of HPC systems using prediction and multiple proactive actions. The Journal of Supercomputing, 71(10):3668–3694, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1458-0.

Ahn:2015:SMB

[1163] Jinhyun Ahn, Dong-Hyuk Im, and Hong-Gee Kim. SigMR: MapReducebased SPARQL query processing by signature encoding and multi-way join. The Journal of Supercomputing, 71(10):3695–3725, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1459-z.

Jimenez:2015:EEM

[1164] Ernesto Jiménez, Sergio Arévalo, Carlos Herrera, and Jian Tang. Eventual election of multiple leaders for solving consensus in anonymous systems. The Journal of Supercomputing, 71(10): 3726–3743, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1460-6.

Li:2015:OPM

[1165] Keqin Li. Optimal partitioning of a multicore server processor. The Journal of Supercomputing, 71(10): 3744–3769, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1463-3.

Noghondar:2015:LCL

[1166] Amir Fadakar Noghondar and Midia Reshadi. A low-cost and latency bypass channel-based on-chip network. The Journal of Supercomputing, 71(10): 3770–3786, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1466-0.

Thompson:2015:PCI

[1167] Elizabeth Thompson, Nathan Clem, and David A. Peter. Parallel CUDA implementation of conflict detection for application to airspace deconfliction. The Journal of Supercomputing, 71(10):3787–3810, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1467-z.

Tsai:2015:ADC

[1168] Ying-Lin Tsai, Hsiao-Ching Liu, and Kuo-Chan Huang. Adaptive dual-criteria task group allocation for clustering-based multi-workflow scheduling on parallel computing platform. The Journal of Supercomputing, 71(10):3811–3831, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1469-x.

Seo:2015:ASP

[1169] Jung-Hyun Seo, Jong-Seok Kim, and Hyeong-Ok Lee. An algorithm for sorting pancake by restricted reversals. The Journal of Supercomputing, 71(10): 3832–3850, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1473-1.

Roquero:2015:HST

[1170] Paula Roquero, Javier Ramos, Victor Moreno, and Iván González. Highspeed TCP flow record extraction using GPUs. The Journal of Supercomputing, 71(10):3851–3876, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1478-9.

Guyeux:2015:ECS

[1171] Christophe Guyeux, Raphaël Couturier, and Pierre-Cyrille Héam. Efficient and cryptographically secure generation of chaotic pseudorandom numbers on GPU. The Journal of Supercomputing, 71(10):3877–3903, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1479-8.

Chaturvedi:2015:AMR

[1172] Nitin Chaturvedi, Arun Subramaniyan, and S. Gurunarayanan. An adaptive migration-replication scheme (AMR) for shared cache in chip multiprocessors. The Journal of Supercomputing, 71(10):3904–3933, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1482-0.

Filipovic:2015:OCC

[1173] Jirí Filipovic, Matús Madzin, Jan Fousek, and Ludek Matyska. Optimizing CUDA code by kernel fusion: application on BLAS. The Journal of Supercomputing, 71(10):3934–3957, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1483-z.

Nejatollahi:2015:VSD

[1174] Hamid Nejatollahi and Mostafa E.

Salehi. Voltage scaling and dark silicon in symmetric multicore processors. The Journal of Supercomputing, 71(10): 3958–3973, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1486-9.

Mbock:2015:RBM

[1175] Etienne Aubin Mbe Mbock. Reconfiguration based model for matrix triangulation and hardware device creation concept. The Journal of Supercomputing, 71(10):3974–3990, October 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1493-x.

Hasanov:2015:HAO

[1176] Khalid Hasanov, Jean-Noël Quintin, and Alexey Lastovetsky. Hierarchical approach to optimization of parallel matrix multiplication on largescale platforms. The Journal of Supercomputing, 71(11):3991–4014, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1133-x.

Shi:2015:CTS

[1177] Run hua Shi, Hong Zhong, and Shun Zhang. Comments on two schemes of identity-based user authentication and key agreement for mobile clientserver networks. The Journal of Supercomputing, 71(11):4015–4018, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1496-7.

Bistouni:2015:FBR

[1178] Fathollah Bistouni and Mohsen Jahanshahi. Formulating broadcast reliability equations on multilayer multistage interconnection networks. The Journal of Supercomputing, 71(11): 4019–4041, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1502-0.

Liu:2015:QMC

[1179] Xiaodong Liu, Songyang Li, and Weiqin Tong. A queuing model considering resources sharing for cloud service performance. The Journal of Supercomputing, 71(11):4042–4055, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1503-z.

Fang:2015:KAM

[1180] Juan Fang, Lu Yu, Sitong Liu, Jiajia Lu, and Tan Chen. KL GA: an application mapping algorithm for meshof-tree (MoT) architecture in networkon-chip design. The Journal of Supercomputing, 71(11):4056–4071, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1504-y;   http://link.springer.
com/content/pdf/10.1007/s11227-
015-1504-y.pdf.

Bai:2015:SPA

[1181] Xiuxiu Bai, Endong Wang, Xiaoshe Dong, and Xingjun Zhang. A scalability prediction approach for multithreaded applications on manycore processors. The Journal of Supercomputing, 71(11):4072–4094, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1505-x.

Ghosh:2015:NCC

[1182] Ammlan Ghosh, Rituparna Chaki, and Nabendu Chaki. A new concurrency control mechanism for multithreaded environment using transactional memory. The Journal of Supercomputing, 71(11):4095–4115, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1507-8;   http://link.springer.
com/content/pdf/10.1007/s11227-
015-1507-8.pdf.

Tos:2015:DRS

[1183] Uras Tos, Riad Mokadem, Abdelkader Hameurlain, and Tolga Ayav. Dynamic replication strategies in data grid systems: a survey. The Journal of Supercomputing, 71(11):4116–4140, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1508-7.

Jin:2015:GAP

[1184] Jing Jin, Xianggao Cai, Guoming Lai, and Xiaola Lin. GPUaccelerated parallel algorithms for linear rankSVM. The Journal of Supercomputing, 71(11):4141–4171, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1509-6.

Jimenez:2015:FTB

[1185] Ernesto Jiménez, Sergio Arévalo, and Jian Tang. Fault-tolerant broadcast in anonymous systems. The Journal of Supercomputing, 71(11): 4172–4191, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1512-y.

Dummler:2015:IBP

[1186] Jörg Dümmler and Sebastian Egerland. Interval-based performance modeling for the all-pairs-shortest-path problem on GPUs. The Journal of Supercomputing, 71(11):4192–4214, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1514-9.

Souravlas:2015:SAR

[1187] Stavros Souravlas and Manos Roumeliotis. Scheduling array redistribution with virtual channel support. The Journal of Supercomputing, 71(11): 4215–4234, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1519-4.

Barati:2015:HHB

[1188] Masoud Barati and Saeed Sharifian.

A hybrid heuristic-based tuned support vector regression model for cloud load prediction. The Journal of Supercomputing, 71(11):4235–4259, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1520-y.

Johari:2015:MBR

[1189] Suchi Johari and Vivek Kumar Sehgal.

Master-based routing algorithm and communication-based cluster topology for 2D NoC. The Journal of Supercomputing, 71(11):4260–4286, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1521-x.

Abawajy:2015:SLA

[1190] Jemal Abawajy, Mohd Farhan Fudzee, and Mohammad Mehedi Hassan. Service level agreement management framework for utility-oriented computing platforms. The Journal of Supercomputing, 71(11):4287–4303, November 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1526-5.

Gaona:2015:FEC

[1191] Epifanio Gaona, José L. Abellán, and Manuel E. Acacio. Fast and efficient commits for lazy–lazy hardware transactional memory. The Journal of Supercomputing, 71(12):4305–4326, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1523-8.

Borkar:2015:RFU

[1192] Meenal Borkar and Nitin. Realizing frequently used permutations on gamma interconnection network’s family networks with the help of alternate source. The Journal of Supercomputing, 71(12):4327–4351, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1527-4.

Kirsal:2015:MAV

[1193] Yonal Kirsal, Enver Ever, and Altan Kocyigit. Modelling and analysis of vertical handover in highly mobile environments. The Journal of Supercomputing, 71(12):4352–4380, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1528-3.

Hussain:2015:BTD

[1194] Zaid Hussain. Better traffic distribution one-to-all broadcast in higher dimensional Gaussian networks. The Journal of Supercomputing, 71(12): 4381–4399, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1532-7.

Messai:2015:LKM

[1195] Mohamed-Lamine Messai, Hamida Seba, and Makhlouf Aliouat. A lightweight key management scheme for wireless sensor networks. The Journal of Supercomputing, 71(12): 4400–4422, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1534-5.

Malekimajd:2015:MLG

[1196] Marzieh Malekimajd and Ali Movaghar.

Minimizing latency in geo-distributed clouds. The Journal of Supercomputing, 71(12):4423–4445, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1538-1.

Ahmed:2015:HSP

[1197] Achraf Ben Ahmed and Abderazek Ben Abdallah. Hybrid silicon-photonic network-on-chip for future generations of high-performance many-core systems. The Journal of Supercomputing, 71(12):4446–4475, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1539-0.

Fernandez-Gonzalez:2015:HRF

[1198] Álvaro Fernández-González and Rafael Rosillo. Historical review and future challenges in supercomputing and networks of scientific communication. The Journal of Supercomputing, 71(12): 4476–4503, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1544-3.

Filiposka:2015:CBV

[1199] Sonja Filiposka, Anastas Mishev, and Carlos Juiz. Community-based VM placement framework. The Journal of Supercomputing, 71(12):4504– 4528, December 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1546-1.

Leslie:2015:RRA

[1200] Luke M. Leslie, Young Choon Lee, and Albert Y. Zomaya. RAMP: reliability-aware elastic instance provisioning for profit maximization. The Journal of Supercomputing, 71(12): 4529–4554, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1548-z.

Raei:2015:APM

[1201] Hassan Raei and Nasser Yazdani. Analytical performance model for mobile network operator cloud. The Journal of Supercomputing, 71(12): 4555–4577, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1551-4.

Idrees:2015:DLC

[1202] Ali Kadhum Idrees, Karine Deschinkel, and Michel Salomon. Distributed lifetime coverage optimization protocol in wireless sensor networks. The Journal of Supercomputing, 71(12): 4578–4593, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1558-x.

Mohaqeqi:2015:TAS

[1203] Morteza Mohaqeqi and Mehdi Kargahi. Thermal analysis of stochastic DVFS-enabled multicore real-time systems. The Journal of Supercomputing, 71(12):4594–4622, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1562-1.

Kim:2015:PMS

[1204] Changhyeon Kim, Changho Jeon, Wonjoo Lee, and Sungil Yang. A parallel migration scheme for fast virtual machine relocation on a cloud cluster. The Journal of Supercomputing, 71(12):4623–4645, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1563-0.

Dastgeer:2015:PAC

[1205] Usman Dastgeer and Christoph Kessler. Performance-aware composition framework for GPU-based systems. The Journal of Supercomputing, 71(12):4646–4662, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1105-1.

Salmito:2015:SAD

[1206] Tiago Salmito, Ana Lúcia de Moura, and Noemi Rodriguez. A stepwise approach to developing staged applications. The Journal of Supercomputing, 71(12):4663–4679, December 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-014-
                                        1110-4.

Lee:2016:ESS

[1207] Changhoon Lee, Kyusuk Han, and Juan Li. Editorial: A special section on “emerging platform technologies”. The Journal of Supercomputing, 72(1):1–4, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-015-1595-5.pdf.

Jiang:2016:OFB

[1208] Feng Jiang, Bo-Wei Chen, Seungmin Rho, Wen Ji, Liqiang Pan, Hongwei Guo, and Debin Zhao. Optimal filter based on scale-invariance generation of natural images. The Journal of Supercomputing, 72(1):5–23, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1398-8.

Kim:2016:NSA

[1209] Sang-Kon Kim, Seung-Young Ma, and Jongsub Moon. A novel secure architecture of the virtualized server system. The Journal of Supercomputing, 72(1):24–37, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1401-4.

Kim:2016:SKS

[1210] Jin Kim, Sung Yun Yu, and Sang Oh Park. Smart knowledge sharing system for cyberinfrastructure. The Journal of Supercomputing, 72(1):38–45, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1440-x.

Kim:2016:NAD

[1211] Eunyoung Kim and Jongsub Moon. A new approach to deploying private mobile network exploits. The Journal of Supercomputing, 72(1):46–57, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1461-5.

Jabbar:2016:TMS

[1212] Sohail Jabbar, Kashif Naseer, Moneeb Gohar, Seungmin Rho, and Hangbae Chang. Trust model at service layer of cloud computing for educational institutes. The Journal of Supercomputing, 72(1):58–83, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1488-7.

Yu:2016:AIT

[1213] Jaehak Yu, Hyo-Chan Bang, Hosung Lee, and Yang Sun Lee. Adaptive Internet of Things and Web of Things convergence platform for Internet of reality services. The Journal of Supercomputing, 72(1):84–102, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1489-6.

Kang:2016:SEC

[1214] Minkyung Kang, Onechul Na, and Hangbae Chang. Security experts’ capability design for future Internet of Things platform. The Journal of Supercomputing, 72(1):103–119, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1490-0.

Hsu:2016:DCH

[1215] Fu-Hau Hsu, Min-Hao Wu, Syun-Cheng Ou, and Shiuh-Jeng Wang. Data concealments with high privacy in new technology file system. The Journal of Supercomputing, 72 (1):120–140, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1492-y.

Kao:2016:CIC

[1216] Da-Yu Kao. Cybercrime investigation countermeasure using createdaccessed-modified model in cloud computing environments. The Journal of Supercomputing, 72(1):141–160, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1516-7.

Kwon:2016:NSB

[1217] Jung-Hyok Kwon, Hyun Soo Chang, Taeshik Shon, Jai-Jin Jung, and Eui-Jik Kim. Neighbor stability-based VANET clustering for urban vehicular environments. The Journal of Supercomputing, 72(1):161–176, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1517-6.

Kim:2016:MAS

[1218] Soon Seok Kim. Mutual authentication scheme between biosensor device and data manager in healthcare environment. The Journal of Supercomputing, 72(1):177–184, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1536-3.

Seo:2016:CBS

[1219] Sanghyun Seo, Hakjeon Bang, and Hunjoo Lee. Coloring-based scheduling for interactive game application with wireless body area networks. The Journal of Supercomputing, 72 (1):185–195, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1540-7.

Kang:2016:RTM

[1220] Hyeongseok Kang, Dohyeon Kim, Jeongnam Kang, and Kanghee Kim. Real-time motion control on Android platform. The Journal of Supercomputing, 72(1):196–213, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1542-5.

Cho:2016:NSA

[1221] Kyumin Cho, Hyunwoo Lim, Taeshik Shon, and Seungjoo Kim. A novel security architecture of electronic vehicle system for smart grid communication. The Journal of Supercomputing, 72(1):214–231, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1552-3.

Cho:2016:ADS

[1222] Haehyun Cho, Jongsu Lim, Hyunki Kim, and Jeong Hyun Yi. Antidebugging scheme for protecting mobile apps on Android platform. The Journal of Supercomputing, 72(1): 232–246, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1559-9.

Elmisery:2016:PEM

[1223] Ahmed M. Elmisery, Seungmin Rho, and Dmitri Botvich. Privacy-enhanced middleware for location-based subcommunity discovery in implicit social groups. The Journal of Supercomputing, 72(1):247–274, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1574-x.

Nikolic:2016:IFT

[1224] Tatjana R. Nikolic, Goran S. Nikolic, Goran Lj. Djordjevic, and Mile K. Stojcev. Improving fault-tolerance capability of on-chip binary CDMA bus. The Journal of Supercomputing, 72 (1):275–294, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1513-x.

Liu:2016:RUA

[1225] Longjun Liu, Hongbin Sun, Chao Li, Yang Hu, Jingmin Xin, Nanning Zheng, and Tao Li. RE-UPS: an adaptive distributed energy storage system for dynamically managing solar energy in green datacenters. The Journal of Supercomputing, 72(1):295–316, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1529-2.

Baranwal:2016:ACC

[1226] Gaurav Baranwal and Deo Prakash Vidyarthi. Admission control in cloud computing using game theory. The Journal of Supercomputing, 72 (1):317–346, January 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1565-y.

Soysal:2016:SMA

[1227] Ömer M. Soysal, Eera Gupta, and Harisha Donepudi. A sparse memory allocation data structure for sequential and parallel association rule mining. The Journal of Supercomputing, 72 (2):347–370, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1566-x.

Wu:2016:OGA

[1228] Xiaohong Wu, Yonggen Gu, Jie Tao, Guoqiang Li, Prem Prakash Jayaraman, Daniel Sun, Rajiv Ranjan, Albert Zomaya, and Jingti Han. An online greedy allocation of VMs with non-increasing reservations in clouds. The Journal of Supercomputing, 72 (2):371–390, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1567-9.

Duy:2016:HDF

[1229] Truong Vinh Truong Duy and Taisuke Ozaki. Hybrid and 4-D FFT implementations of an open-source parallel FFT package OpenFFT. The Journal of Supercomputing, 72(2):391–416, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1568-8.

Quezada-Naquid:2016:RPA

[1230] Moisés Quezada-Naquid, Ricardo Marcelín-Jiménez, J. L. Gonzalez-Compeán, and Jesus Carretero Perez. RS-Pooling: an adaptive data distribution strategy for fault-tolerant and large-scale storage systems. The Journal of Supercomputing, 72(2): 417–437, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1569-7.

Akleylek:2016:SPM

[1231] Sedat Akleylek, Erdem Alkim, and Zaliha Yüce Tok. Sparse polynomial multiplication for lattice-based cryptography with small complexity. The Journal of Supercomputing, 72 (2):438–450, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1570-1.

Denham:2016:SAR

[1232] Mónica Denham, Javier Areta, and Fernando G. Tinetti. Synthetic aperture radar signal processing in parallel using GPGPU. The Journal of Supercomputing, 72(2):451–467, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1572-z.

Cheikh:2016:TFS

[1233] Taieb Lamine Ben Cheikh, Alexandra Aguiar, Sofiene Tahar, and Gabriela Nicolescu. Tuning framework for stencil computation in heterogeneous parallel platforms. The Journal of Supercomputing, 72(2):468–502, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1575-9.

Marino:2016:LLC

[1234] Mario D. Marino and Kuan-Ching Li.

Last level cache size heterogeneity in embedded systems. The Journal of Supercomputing, 72(2):503–544, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1576-8.

Lopez-Novoa:2016:KDE

[1235] Unai Lopez-Novoa, Alexander Mendiburu, and Jose Miguel-Alonso. Kernel density estimation in accelerators. The Journal of Supercomputing, 72(2): 545–566, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1577-7.

Yan:2016:EGC

[1236] Jili Yan. Enhanced global congestion awareness (EGCA) for load balance in networks-on-chip. The Journal of Supercomputing, 72(2):567–587, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1583-9.

Sheikhi:2016:PFL

[1237] Sanaz Sheikhi and Seyed Morteza Babamir. A predictive framework for load balancing clustered web servers. The Journal of Supercomputing, 72 (2):588–611, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1584-8.

Fernandez-Pascual:2016:DSC

[1238] Ricardo Fernández-Pascual, Alberto Ros, and Manuel E. Acacio. Are distributed sharing codes a solution to the scalability problem of coherence directories in manycores? An evaluation study. The Journal of Supercomputing, 72(2):612–638, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1596-4.

Yu:2016:CFS

[1239] Zhigang Yu, Xinyu Wang, and Kele Shen. Conditional forwarding: simple flow control to increase adaptivity for fully adaptive routing algorithms. The Journal of Supercomputing, 72 (2):639–653, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1597-3.

Aviles-Gonzalez:2016:BOI

[1240] Ana Avilés-González, Juan Piernas, and Pilar González-Férez. Batching operations to improve the performance of a distributed metadata service. The Journal of Supercomputing, 72 (2):654–687, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1602-x.

Arianyan:2016:NHC

[1241] Ehsan Arianyan, Hassan Taheri, and Saeed Sharifian. Novel heuristics for consolidation of virtual machines in cloud data centers using multi-criteria resource management solutions. The Journal of Supercomputing, 72(2): 688–717, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1603-9.

Hijaz:2016:LAD

[1242] Farrukh Hijaz, Qingchuan Shi, George Kurian, Srinivas Devadas, and Omer Khan. Locality-aware data replication in the last-level cache for large scale multicores. The Journal of Supercomputing, 72(2):718–752, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1608-4.

dAuriol:2016:AOL

[1243] Brian J. d’Auriol. All-optical Linear Array with a Reconfigurable Pipelined Bus System (OLARPBS) optical bus parallel computing model. The Journal of Supercomputing, 72(2):753– 769, February 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1611-9.

Yu:2016:DDS

[1244] Hairong Yu, Guohui Li, and LihChyun Shu. DS_{spirit}: a data dependence and stride reference patterns profiling infrastructure. The Journal of Supercomputing, 72(2):770–788, February 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1612-8.

Pan:2016:SPE

[1245] Wen-Tsao Pan, Chiung-En Huang, and Chiung-Lin Chiu. Study on the performance evaluation of online teaching using the quantile regression analysis and artificial neural network. The Journal of Supercomputing, 72(3):789–803, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1599-1.

Kelefouras:2016:HPM

[1246] Vasilios Kelefouras, A. Kritikakou, Iosif Mporas, and Vasilios Kolonias. A high-performance matrix-matrix multiplication methodology for CPU and GPU architectures. The Journal of Supercomputing, 72(3):804–844, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1613-7.

Kononenko:2016:AEC

[1247] Kirill Kononenko. An approach to error correction in program code using dynamic optimization in a virtual execution environment. The Journal of Supercomputing, 72(3):845–873, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1616-4.

Azizi:2016:HEN

[1248] Sadoon Azizi, Naser Hashemi, and Ahmad Khonsari. HHS: an efficient network topology for large-scale data centers. The Journal of Supercomputing, 72(3):874–899, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-015-1617-3.

Luo:2016:SAU

[1249] Min Luo, Jingyin Zhang, Debiao He, and Jian Shen. Security analysis of a user registration approach. The Journal of Supercomputing, 72(3):900–903, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-015-
                                        1619-1.

Imre:2016:DMR

[1250] Kayhan M. Imre. Dual-mode routing approach for photonic network on chip platforms. The Journal of Supercomputing, 72(3):904–925, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1620-3.

Singh:2016:RPS

[1251] Sukhpal Singh and Inderveer Chana. Resource provisioning and scheduling in clouds: QoS perspective. The Journal of Supercomputing, 72(3):926–960, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1626-x.

OLoughlin:2016:SVM

[1252] John O’Loughlin and Lee Gillam. Sibling virtual machine co-location confirmation and avoidance tactics for public infrastructure clouds. The Journal of Supercomputing, 72(3):961–984, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
016-1627-9.pdf.

Convolbo:2016:CAD

[1253] Moïse W. Convolbo and Jerry Chou.

Cost-aware DAG scheduling algorithms for minimizing execution cost on cloud resources. The Journal of Supercomputing, 72(3):985–1012, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1637-7.

Balouchzahi:2016:EIB

[1254] Nik-Mohammad Balouchzahi, Mahmood Fathy, and Ahmad Akbari. An efficient infrastructure based service discovery in vehicular networks using P2P structures. The Journal of Supercomputing, 72(3):1013–1034, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1638-6.

Penaranda:2016:ADI

[1255] Roberto Peñaranda, Crispín Gómez, María E. Gómez, Pedro López, and Jose Duato. The k-ary n-direct sindirect family of topologies for largescale interconnection networks. The Journal of Supercomputing, 72(3): 1035–1062, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1640-z.

Ranokphanuwat:2016:PPM

[1256] Ratthaslip Ranokphanuwat and Surin Kittitornkun. Parallel Partition and Merge QuickSort (PPMQSort) on multicore CPUs. The Journal of Supercomputing, 72(3):1063–1091, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1641-y.

Thapliyal:2016:DPN

[1257] Himanshu Thapliyal, Carson Labrado, and Ke Chen. Design procedures and NML cost analysis of reversible barrel shifters optimizing garbage and ancilla lines. The Journal of Supercomputing, 72(3):1092–1124, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1644-8. See erra-

tum [1258].

Thapliyal:2016:EDP

[1258] Himanshu Thapliyal, Carson Labrado, and Ke Chen. Erratum to: Design procedures and NML cost analysis of reversible barrel shifters optimizing garbage and ancilla lines. The Journal of Supercomputing, 72(3):1125, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        016-1685-z.pdf. See [1257].

Wang:2016:LVP

[1259] Fei Wang, Xiaofeng Gao, and Guihai Chen. Lowering the volatility: a practical cache allocation prediction and stability-oriented co-runner scheduling algorithms. The Journal of Supercomputing, 72(3):1126–1151, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1645-7.

Guo:2016:PSI

[1260] Dingding Guo, Yu-Kwong Kwok, Xin Jin, and Jian Deng. A performance study of incentive schemes in peer-to-peer file-sharing systems. The Journal of Supercomputing, 72(3):1152–1178, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1648-4.

Liu:2016:EMA

[1261] Wen Liu, Yanming Shen, and Peng Wang. An efficient MapReduce algorithm for similarity join in metric spaces. The Journal of Supercomputing, 72(3):1179–1200, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1651-9.

Phoummavong:2016:LAR

[1262] Phonepadith Phoummavong, Keisuke Utsu, Chee Onn Chow, and Hiroshi Ishii. Location-aided route discovery mechanism based on two-hop neighbor information for ad hoc network. The Journal of Supercomputing, 72 (3):1201–1214, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1656-4.

Kobayashi:2016:ASC

[1263] Kei Kobayashi, Yosuke Totani, Keisuke Utsu, and Hiroshi Ishii. Achieving secure communication over MANET using secret sharing schemes. The Journal of Supercomputing, 72(3): 1215–1225, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1657-3.

Kobayashi:2016:PLD

[1264] Kei Kobayashi, Yosuke Totani, Keisuke Utsu, and Hiroshi Ishii. A proposal on location data supplementing information transfer method over MANET. The Journal of Supercomputing, 72 (3):1226–1236, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1658-2.

Imaizumi:2016:SEF

[1265] Naoya Imaizumi, Kei Kobayashi, Keisuke Utsu, and Hiroshi Ishii. A study on effective flooding over MANET based on exchange of neighbor information. The Journal of Supercomputing, 72(3):1237–1245, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1659-1.

Abdollahi:2016:ICO

[1266] Mahdi Abdollahi, Asgarali Bouyer, and Davoud Abdollahi. Improved cuckoo optimization algorithm for solving systems of nonlinear equations. The Journal of Supercomputing, 72 (3):1246–1269, March 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1660-8.

Saad:2016:SOV

[1267] Walid Saad, Heithem Abbes, Christophe Cérin, and Mohamed Jemni. A selforganized volunteer cloud for e-science. The Journal of Supercomputing, 72 (4):1271–1290, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1564-z.

Hernandez-Becerril:2016:GIS

[1268] Rogelio Adrian Hernandez-Becerril, Ariana Guadalupe Bucio-Ramirez, Mariko Nakano-Miyatake, Hector Perez-Meana, and Marco Pedro Ramirez-Tachiquin. A GPU implementation of secret sharing scheme based on cellular automata. The Journal of Supercomputing, 72(4):1291–1311, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1646-6.

Ahn:2016:DEC

[1269] Hoo-Young Ahn, Kyong-Ha Lee, and Yoon-Joon Lee. Dynamic erasure coding decision for modern blockoriented distributed storage systems. The Journal of Supercomputing, 72 (4):1312–1341, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1661-7.

Kang:2016:SBR

[1270] Saehoon Kang and Wonyong Yoon. SDN-based resource allocation for heterogeneous LTE and WLAN multiradio networks. The Journal of Supercomputing, 72(4):1342–1362, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1662-6.

Pouyan:2016:RAT

[1271] Fatemeh Pouyan, Ali Azarpeyvand, Saeed Safari, and Sied Mehdi Fakhraie. Reliability aware throughput management of chip multi-processor architecture via thread migration. The Journal of Supercomputing, 72(4):1363–1380, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1665-3.

Manaka:2016:CSU

[1272] Ayami Manaka, Hiroshi Ishii, Chee Onn Chow, and Keisuke Utsu. Concepts, services, and underlying network for a daily life support system for citizens. The Journal of Supercomputing, 72(4):1381–1398, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1667-1.

Utsu:2016:BBI

[1273] Keisuke Utsu, Chee Onn Chow, Hiroaki Nishikawa, and Hiroshi Ishii. Broadcast-based information sharing system (BBISS) on wireless ad hoc communication environment. The Journal of Supercomputing, 72(4): 1399–1421, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1668-0.

Totani:2016:EBB

[1274] Yosuke Totani, Kei Kobayashi, Keisuke Utsu, and Hiroshi Ishii. An efficient broadcast-based information transfer method based on location data over MANET. The Journal of Supercomputing, 72(4):1422–1430, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1669-z.

Silva:2016:BAU

[1275] Francisco Airton Silva, Germano Zaicaner, Eder Quesado, Matheus Dornelas, Bruno Silva, and Paulo Maciel. Benchmark applications used in mobile cloud computing research: a systematic mapping study. The Journal of Supercomputing, 72(4):1431–1452, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1674-2.

Torrents:2016:FPC

[1276] Martí Torrents, Raul Martínez, and Carlos Molina. Facing prefetching challenges in distributed shared memories for CMPs. The Journal of Supercomputing, 72(4):1453–1476, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11227-016-
                                        1675-1.

Jayashree:2016:AIG

[1277] H. V. Jayashree, Himanshu Thapliyal, Hamid R. Arabnia, and V. K. Agrawal. Ancilla-input and garbage-output optimized design of a reversible quantum integer multiplier. The Journal of Supercomputing, 72(4):1477–1493, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1676-0.

Anagnostopoulos:2016:HBD

[1278] I. Anagnostopoulos, S. Zeadally, and E. Exposito. Handling big data: research challenges and future directions. The Journal of Supercomputing, 72 (4):1494–1516, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1677-z.

Chen:2016:APX

[1279] Rongxin Chen, Husheng Liao, Zongyue Wang, and Hang Su. Automatic parallelization of XQuery programs on multi-core systems. The Journal of Supercomputing, 72(4):1517–1548, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1679-x.

Sundriyal:2016:JFS

[1280] Vaibhav Sundriyal and Masha Sosonkina. Joint frequency scaling of processor and DRAM. The Journal of Supercomputing, 72(4):1549–1569, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1680-4.

Chen:2016:SER

[1281] Sui Chen, Greg Bronevetsky, Lu Peng, Bin Li, and Xin Fu. Soft error resilience in Big Data kernels through modular analysis. The Journal of Supercomputing, 72(4):1570–1596, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1682-2.

Son:2016:ATS

[1282] Seokho Son, Dong-Jae Kang, Seyoung Phillip Huh, Won-Young Kim, and Wan Choi. Adaptive trade-off strategy for bargaining-based multiobjective SLA establishment under varying cloud workload. The Journal of Supercomputing, 72(4):1597–1622, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1686-y.

Irshad:2016:EAM

[1283] Azeem Irshad, Muhammad Sher, Shehzad Ashraf Chaudhary, Husnain Naqvi, and Mohammad Sabzinejad Farash. An efficient and anonymous multi-server authenticated key agreement based on chaotic map without engaging Registration Centre. The Journal of Supercomputing, 72 (4):1623–1644, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1688-9.

Jiang:2016:DLM

[1284] Yunyun Jiang, Yi Yang, Tian Xiao, Tianwei Sheng, and Wenguang Chen. DRDDR: a lightweight method to detect data races in Linux kernel. The Journal of Supercomputing, 72 (4):1645–1659, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1691-1.

Rafie:2016:PET

[1285] Mahnaz Rafie, Ahmad Khademzadeh, Akram Reza, and Midia Reshadi. Performance evaluation of task migration in contiguous allocation for mesh interconnection topology. The Journal of Supercomputing, 72(4):1660–1677, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1692-0.

He:2016:ELS

[1286] Xinyu He, Bo-Wei Chen, Wen Ji, Seungmin Rho, and Sun-Yuan Kung. Erratum to: Large-scale image colorization based on divide-and-conquer support vector machines. The Journal of Supercomputing, 72(4):1678, April 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        016-1699-6.pdf.

Do:2016:NBP

[1287] Cong Thuan Do, Hong Jun Choi, Dong Oh Son, Jong Myon Kim, and Cheol Hong Kim. NTB branch predictor: dynamic branch predictor for high-performance embedded processors. The Journal of Supercomputing, 72(5):1679–1693, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1280-0.

Choi:2016:RWC

[1288] Cheol-Rim Choi, Hwa-Young Jeong, Jong Hyuk Park, Haeng Jin Jang, and Young-Sik Jeong. Relative weight comparison between virtual key factors of cloud computing with analytic network process. The Journal of Supercomputing, 72(5):1694–1714, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1311-x.

Lee:2016:HAS

[1289] Seoung-Hyeon Lee, Young-Hyuk Kim, Jae-Kwang Lee, and Deok Gyu Lee. Hybrid app security protocol for high speed mobile communication. The Journal of Supercomputing, 72(5): 1715–1739, May 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1318-3.

Moon:2016:RRT

[1290] Daesung Moon, Jae Dong Lee, Young-Sik Jeong, and Jong Hyuk Park.

RTNSS: a routing trace-based network security system for preventing ARP spoofing attacks. The Journal of Supercomputing, 72(5):1740–1756, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-014-
1353-0.

Jung:2016:SDM

[1291] Dahae Jung, Min-Kyoung Bae, Man Yong Choi, Eui Chul Lee, and Jinoo Joung. Speaker diarization method of telemarketer and client for improving speech dictation performance. The Journal of Supercomputing, 72(5):1757–1769, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1470-4.

Peng:2016:DHE

[1292] Junjie Peng, Rong Shen, and Xianshun Ping. Design of a high-efficient MSD adder. The Journal of Supercomputing, 72(5):1770–1784, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1484-y.

Marino:2016:ISM

[1293] Mario D. Marino and Kuan-Ching Li. Implications of shallower memory controller transaction queues in scalable memory systems. The Journal of Supercomputing, 72(5):1785–1798, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1485-x.

You:2016:SSP

[1294] Shingchern D. You and Yu-Chu Lin.

Simulated smart phone recordings for audio identification. The Journal of Supercomputing, 72(5):1799–1812, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1533-6.

Veeraraghavan:2016:APS

[1295] Prakash Veeraraghavan, Sadiq Almuairfi, and Naveen Chilamkurti. Anonymous paperless secure payment system using clouds. The Journal of Supercomputing, 72(5):1813–1824, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1560-3.

Um:2016:DRS

[1296] Jung-Ho Um, Seungwoo Lee, Tae-Hong Kim, Chang-Hoo Jeong, Sa-Kwang Song, and Hanmin Jung. Distributed RDF store for efficient searching billions of triples based on Hadoop. The Journal of Supercomputing, 72 (5):1825–1840, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1670-6.

Fujii:2016:EDF

[1297] Shota Fujii, Masaya Sato, Toshihiro Yamauchi, and Hideo Taniguchi. Evaluation and design of function for tracing diffusion of classified information for file operations with KVM. The Journal of Supercomputing, 72 (5):1841–1861, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1671-5.

Huh:2016:AMI

[1298] Jun-Ho Huh, Sugarbayar Otgonchimeg, and Kyungryong Seo. Advanced metering infrastructure design and test bed experiment using intelligent agents: focusing on the PLC network base technology for Smart Grid system. The Journal of Supercomputing, 72 (5):1862–1877, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1672-4.

Han:2016:HPC

[1299] Seok-Hyeon Han and Gangman Yi. High performance clustering algorithm for analysis of protein family clusters. The Journal of Supercomputing, 72 (5):1878–1896, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1706-y.

Jahani:2016:CSR

[1300] Arezoo Jahani and Leyli Mohammad Khanli. Cloud service ranking as a multi objective optimization problem. The Journal of Supercomputing, 72 (5):1897–1926, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1690-2. See erra-

tum [1301].

Jahani:2016:ECS

[1301] Arezoo Jahani and Leyli Mohammad Khanli. Erratum to: Cloud service ranking as a multi objective optimization problem. The Journal of Supercomputing, 72(5):1927, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
016-1709-8.pdf. See [1300].

Li:2016:JJM

[1302] Bing Li, Junbo Zhang, Ning Yu, and Yi Pan. J2M: a Java to MapReduce translator for cloud computing. The Journal of Supercomputing, 72 (5):1928–1945, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1695-x.

Oyama:2016:EAO

[1303] Yoshihiro Oyama, Shun Ishiguro, Jun Murakami, Shin Sasaki, Ryo Matsumiya, and Osamu Tatebe. Experimental analysis of operating system jitter caused by page reclaim. The Journal of Supercomputing, 72 (5):1946–1972, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1703-1.

Marendic:2016:NMR

[1304] P. Marendic, J. Lemeire, D. Vucinic, and P. Schelkens. A novel MPI reduction algorithm resilient to imbalances in process arrival times. The Journal of Supercomputing, 72(5):1973–2013, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1707-x.

Belyi:2016:CAR

[1305] Eugene Belyi, Philippe J. Giabbanelli, Indravadan Patel, Naga Harish Balabhadrapathruni, Aymen Ben Abdallah, Wedyan Hameed, and Vijay K. Mago. Combining association rule mining and network analysis for pharmacosurveillance. The Journal of Supercomputing, 72(5):2014–2034, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1714-y.

Wang:2016:NHA

[1306] Xinyu Wang, Haikuo Liu, and Zhigang Yu. A novel heuristic algorithm for IP block mapping onto meshbased networks-on-chip. The Journal of Supercomputing, 72(5):2035–2058, May 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1719-6.

Tang:2016:OMW

[1307] Zhuo Tang, Min Liu, Almoalmi Ammar, Kenli Li, and Keqin Li. An optimized MapReduce workflow scheduling algorithm for heterogeneous computing. The Journal of Supercomputing, 72(6):2059–2079, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1335-2.

Wang:2016:EPC

[1308] Zhongya Wang, Ying Liu, and Steve Chiu. An efficient parallel collaborative filtering algorithm on multi-GPU platform. The Journal of Supercomputing, 72(6):2080–2094, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-014-1333-4.

He:2016:OMD

[1309] Jinrong He, Di Wu, Naixue Xiong, and Chuansheng Wu. Orthogonal margin discriminant projection for dimensionality reduction. The Journal of Supercomputing, 72(6):2095–2110, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1453-5.

Liu:2016:DEM

[1310] Ying Liu, Haixin Zheng, Renliang Zhao, and Liheng Jian. Design and evaluation of multi-GPU enabled Multiple Symbol Detection algorithm. The Journal of Supercomputing, 72 (6):2111–2131, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1475-z.

Atahary:2016:PMD

[1311] Tanvir Atahary, Tarek M. Taha, and Scott Douglass. Parallelized mining of domain knowledge on GPGPU and Xeon Phi clusters. The Journal of Supercomputing, 72(6):2132–2156, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1712-0.

Nasiri:2016:IPA

[1312] Mahdi Nasiri and Behrouz Minaei. Increasing prediction accuracy in collaborative filtering with initialized factor matrices. The Journal of Supercomputing, 72(6):2157–2169, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1717-8.

Lubas:2016:CAB

[1313] Robert Lubaś, Jarosław Was, and Jakub Porzycki. Cellular Automata as the basis of effective and realistic agent-based models of crowd behavior. The Journal of Supercomputing, 72(6):2170–2196, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1718-7.

Li:2016:ESS

[1314] Chunlin Li, Liu Yanpei, and Luo Youlong. Efficient service selection approach for mobile devices in mobile cloud. The Journal of Supercomputing, 72(6):2197–2220, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1720-0.

Mahjoub:2016:UUN

[1315] Shabnam Mahjoub and Hakimeh Vojoudi. The UTFLA: uniformization of non-uniform iteration spaces in twolevel perfect nested loops using SFLA. The Journal of Supercomputing, 72 (6):2221–2234, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1725-8.

Masarat:2016:MPR

[1316] Saman Masarat, Saeed Sharifian, and Hassan Taheri. Modified parallel random forest for intrusion detection systems. The Journal of Supercomputing, 72(6):2235–2258, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1727-6.

Makaratzis:2016:PMR

[1317] Antonios T. Makaratzis, Christos K.

Filelis-Papadopoulos, and George A. Gravvanis. Parallel multilevel recursive approximate inverse techniques for solving general sparse linear systems. The Journal of Supercomputing, 72 (6):2259–2282, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1728-5.

Pang:2016:CAB

[1318] Ke Pang, Virginie Fresse, and Suying Yao. Communication-aware branch and bound with cluster-based latencyconstraint mapping technique on network-on-chip. The Journal of Supercomputing, 72(6):2283–2309, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11227-016-
                                        1732-9.

Rajkumar:2016:MIN

[1319] S. Rajkumar and Neeraj Kumar Goyal. Multistage interconnection networks reliability analysis. The Journal of Supercomputing, 72(6):2310–2350, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1734-7.

Dai:2016:GPU

[1320] Yuan Dai, Yong Fang, Long Yang, and Gwanggil Jeon. Graphics processing unit-accelerated joint-bitplane belief propagation algorithm in DSC. The Journal of Supercomputing, 72 (6):2351–2375, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1736-5.

Tian:2016:HOA

[1321] Wenhong Tian, Guozhong Li, Wutong Yang, and Rajkumar Buyya. HScheduler: an optimal approach to minimize the makespan of multiple MapReduce jobs. The Journal of Supercomputing, 72(6):2376–2393, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1737-4.

Zhou:2016:OPI

[1322] Yi Zhou, Fazhi He, and Yimin Qiu. Optimization of parallel iterated local search algorithms on graphics processing unit. The Journal of Supercomputing, 72(6):2394–2416, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1738-3.

Zhou:2016:SPC

[1323] Haohao Zhou, Su Deng, and Hongbin Huang. Stability property of clouds and cooperative scheduling policies on multiple types of resources in cloud computing. The Journal of Supercomputing, 72(6):2417–2436, June 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1741-8.

Gu:2016:CST

[1324] Chengxi Gu, G. U. Caidong, and Fan Ling. Calculation and simulation of transient optimal voltage output point in wireless sensor networks. The Journal of Supercomputing, 72 (7):2767–2781, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1605-7.

Hao:2016:ISB

[1325] Fei Hao, Doo-Soon Park, Zheng Pei, HwaMin Lee, and Young-Sik Jeong. Identifying the social-balanced densest subgraph from signed social networks. The Journal of Supercomputing, 72(7):2782–2795, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-015-1606-6.

Xu:2016:BLS

[1326] Quanqing Xu, Khin Mi Mi Aung, Yongqing Zhu, and Khai Leong Yong. Building a large-scale object-based active storage platform for data analytics in the Internet of Things. The Journal of Supercomputing, 72(7):2796–2814, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11227-016-
                                        1621-2.

Park:2016:HTA

[1327] Seong hun Park, Sung min Kim, and Young guk Ha. Highway traffic accident prediction using VDS big data analysis. The Journal of Supercomputing, 72(7):2815–2831, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11227-016-1624-z. See erra-

tum [1328].

Park:2016:EHT

[1328] Seong hun Park, Sung min Kim, and Young guk Ha. Erratum to: Highway traffic accident prediction using VDS big data analysis. The Journal of Supercomputing, 72(7):2832, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        016-1655-5.pdf. See [1327].

Wang:2016:SBE

[1329] BeiLie Wang, Hui Liu, and Jie Song. SaaS-based enterprise application integration approach and case study. The Journal of Supercomputing, 72 (7):2833–2847, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1625-y.

Sohn:2016:FSU

[1330] Surgwon Sohn. Feasibility study on the use of wireless accelerometers in the experimental modal testing. The Journal of Supercomputing, 72(7):2848–2859, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1628-8.

Im:2016:RBA

[1331] Hyungjin Im, Jeong Kyu Lee, and Jong Hyuk Park. A rhythm-based access control system for secure wellness services. The Journal of Supercomputing, 72(7):2860–2873, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1642-x.

Lin:2016:NCM

[1332] Yun Lin, Can Wang, Chunguang Ma, Zheng Dou, and Xuefei Ma. A new combination method for multisensor conflict information. The Journal of Supercomputing, 72(7):2874–2890, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1681-3.

Zhou:2016:RCM

[1333] Wei Zhou and Su Yu. Research on the communication method of mobile network shadow fading based on interference alignment algorithm. The Journal of Supercomputing, 72(7):2891–2909, July 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1705-z.

Huang:2016:ASP

[1334] Tzu-Chi Huang, Ce-Kuen Shieh, Naveen Chilamkurti, Ming-Fong Tsai, and Seungmin Rho. Architecture for speeding up program execution with cloud technology. The Journal of Supercomputing, 72(9):3601– 3618, September 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-016-1715-x.

Phuc:2016:SAS

[1335] Tran Song Dat Phuc and Changhoon Lee. Security analysis of SDDObased block cipher for wireless sensor network. The Journal of Supercomputing, 72(9):3619–3628, September 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-015-
1589-3.

Cho:2016:MAT

[1336] Haehyun Cho, Jiwoong Bang, Myeongju Ji, and Jeong Hyun Yi. Mobile application tamper detection scheme using dynamic code injection against repackaging attacks. The Journal of Supercomputing, 72(9):3629–3645, September 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s11227-016-
1763-2.

Song:2016:CBS

[1337] Seheon Song, Changhoon Lee, SangIl Lee, and JaeHyun Park. Capabilitybased semantic matching for dynamic resource allocation in tactical edge environment. The Journal of Supercomputing, 72(9):3646–3662, September 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1765-0.

Kim:2016:STA

[1338] Mucheol Kim. Scientific trend analysis and curation with Korean R&D information. The Journal of Supercomputing, 72(9):3663–3673, September 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11227-016-
1831-7.

Mohammed:2016:BDA

[1339] Sabah Mohammed and Tai Hoon Kim. Big data applications for healthcare: preface to special issue. The Journal of Supercomputing, 72(10): 3675–3676, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-016-1853-1.pdf.

Ahmad:2016:HFN

[1340] Mahmood Ahmad, Muhammad Bilal Amin, Shujaat Hussain, Byeong Ho Kang, Taechoong Cheong, and Sungyoung Lee. Health Fog: a novel framework for health and wellness applications. The Journal of Supercomputing, 72(10):3677–3695, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

He:2016:OSH

[1341] Hui He, Zhonghui Du, Weizhe Zhang, and Allen Chen. Optimization strategy of Hadoop small file storage for big data in healthcare. The Journal of Supercomputing, 72(10):3696–3707, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2016:ICP

[1342] Jinyan Li, Simon Fong, Sabah Mohammed, and Jinan Fiaidhi. Improving the classification performance of biological imbalanced datasets by swarm optimization algorithms. The Journal of Supercomputing, 72(10):3708– 3728, October 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Albuquerque:2016:LIS

[1343] Robson de Oliveira Albuquerque, Luis Javier García Villalba, Ana Lucila Sandoval Orozco, Rafael Timóteo de Sousa Júnior, and Tai-Hoon Kim. Leveraging information security and computational trust for cybersecurity. The Journal of Supercomputing, 72(10):3729–3763, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fong:2016:RAM

[1344] Simon Fong, Xi Wang, Qiwen Xu, Raymond Wong, Jinan Fiaidhi, and Sabah Mohammed. Recent advances in metaheuristic algorithms: Does the Makara dragon exist? The Journal of Supercomputing, 72(10):3764–3786, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Portela:2016:DUR

[1345] Javier Portela, Luis Javier García Villalba, Alejandra Guadalupe Silva Trujillo, Ana Lucila Sandoval Orozco, and Tai-Hoon Kim. Disclosing user relationships in email networks. The Journal of Supercomputing, 72(10):3787– 3800, October 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gong:2016:NSE

[1346] Xueyuan Gong, Simon Fong, Jonathan H.

Chan, and Sabah Mohammed. NSPRING: the SPRING extension for subsequence matching of time series supporting normalization. The Journal of Supercomputing, 72(10):3801–3825, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jiang:2016:PPT

[1347] Qi Jiang, Muhammad Khurram Khan, Xiang Lu, Jianfeng Ma, and Debiao He. A privacy preserving three-factor authentication protocol for e-Health clouds. The Journal of Supercomputing, 72(10):3826–3849, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See comments [1486].

Gong:2016:DSP

[1348] Xueyuan Gong, Simon Fong, Raymond K. Wong, Sabah Mohammed, Jinan Fiaidhi, and Athanasios V. Vasilakos. Discovering sub-patterns from time series using a normalized crossmatch algorithm. The Journal of Supercomputing, 72(10):3850–3867, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Brito:2016:GEB

[1349] Ricardo Brito, Simon Fong, Kyungeun Cho, Wei Song, Raymond Wong, Sabah Mohammed, and Jinan Fiaidhi. GPU-enabled back-propagation artificial neural network for digit recognition in parallel. The Journal of Supercomputing, 72(10):3868–3886, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fong:2016:TSP

[1350] Simon Fong, Kyungeun Cho, Osama Mohammed, Jinan Fiaidhi, and Sabah Mohammed. A time series preprocessing methodology with statistical and spectral analysis for classifying non-stationary stochastic biosignals. The Journal of Supercomputing, 72(10):3887–3908, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ma:2016:ESL

[1351] Linh Van Ma, Jisue Kim, Sanghyun Park, Jinsul Kim, and Jonghyeon Jang. An efficient Session Weight load balancing and scheduling methodology for high-quality telehealth care service based on WebRTC. The Journal of Supercomputing, 72(10):3909– 3926, October 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fong:2016:IMT

[1352] Simon Fong, Kexing Liu, Kyungeun Cho, Raymond Wong, Sabah Mohammed, and Jinan Fiaidhi. Improvised methods for tackling big data stream mining challenges: case study of human activity recognition. The Journal of Supercomputing, 72(10):3927– 3959, October 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Deb:2016:FAS

[1353] Suash Deb, Simon Fong, Zhonghuan Tian, Raymond K. Wong, Sabah Mohammed, and Jinan Fiaidhi. Finding approximate solutions of NP-hard optimization and TSP problems using elephant search algorithm. The Journal of Supercomputing, 72(10):3960– 3992, October 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Brito:2016:TIR

[1354] Ricardo Brito, Simon Fong, Kyungeun Cho, Wei Song, Raymond Wong, Sabah Mohammed, and Jinan Fiaidhi. Towards implementation of residualfeedback GMDH neural network on parallel GPU memory guided by a regression curve. The Journal of Supercomputing, 72(10):3993–4020, October 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chhieng:2016:APB

[1355] Van M. Chhieng, Raymond K. Wong, Simon Fong, and Sabah Mohammed. Autonomous path based data acquisition in sensor networks. The Journal of Supercomputing, 72(10):4021– 4042, October 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Carretero:2016:ISU

[1356] Jesus Carretero, Javier Garcia-Blas, and Raimondas Ciegis. Introduction to sustainable ultrascale computing systems and applications. The Journal of Supercomputing, 72(11): 4043–4046, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-016-1822-8.pdf.

Dietze:2016:WLS

[1357] Robert Dietze, Michael Hofmann, and Gudula Rünger. Water-level scheduling for parallel tasks in compute-intensive application components. The Journal of Supercomputing, 72(11):4047– 4068, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Duro:2016:EMS

[1358] Francisco Rodrigo Duro, Fabrizio Marozzo, Javier Garcia Blas, Domenico Talia, and Paolo Trunfio. Exploiting in-memory storage for improving workflow executions in cloud platforms. The Journal of Supercomputing, 72(11): 4069–4088, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Llopis:2016:AEC

[1359] Pablo Llopis, Manuel F. Dolz, Javier Garcia Blas, Florin Isaila, Mohammad Reza Heidari, and Michael Kuhn. Analyzing the energy consumption of the storage data path. The Journal of Supercomputing, 72(11):4089– 4106, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Morla:2016:HPN

[1360] Ricardo Morla, Pedro Gonçalves, and Jorge G. Barbosa. High-performance network traffic analysis for continuous batch intrusion detection. The Journal of Supercomputing, 72(11):4107– 4128, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gonzalez-Ferez:2016:MNS

[1361] Pilar González-Férez and Angelos Bilas. Mitigation of NUMA and synchronization effects in high-speed network storage over raw Ethernet. The Journal of Supercomputing, 72(11):4129– 4159, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gong:2016:NPG

[1362] Jing Gong, Stefano Markidis, Erwin Laure, Matthew Otten, Paul Fischer, and Misun Min. Nekbone performance on GPUs with OpenACC and CUDA Fortran implementations. The Journal of Supercomputing, 72(11):4160– 4180, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhou:2016:PMH

[1363] Tong Zhou and Jingfei Jiang. Performance modeling of hyper-scale custom machine for the principal steps in block Wiedemann algorithm. The Journal of Supercomputing, 72(11):4181– 4203, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ciegis:2016:ADP

[1364] Raimondas Ciegis, Vadimas Starikovicius, Natalija Tumanova, and Minvydas Ragulskis. Application of distributed parallel computing for dynamic visual cryptography. The Journal of Supercomputing, 72(11):4204– 4220, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Stamatovic:2016:CAL

[1365] Biljana Stamatovic and Roman Trobec. Cellular automata labeling of connected components in n-dimensional binary lattices. The Journal of Supercomputing, 72(11):4221–4232, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jho:2016:SSE

[1366] Nam-Su Jho, Ku-Young Chang, Dowon Hong, and Changho Seo. Symmetric searchable encryption with efficient range query using multi-layered linked chains. The Journal of Supercomputing, 72(11):4233–4246, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Islam:2016:DIA

[1367] Md Saiful Islam, Jung-Chul Lee, and Uipil Chong. Design and implementation of an automated monitoring system. The Journal of Supercomputing, 72(11):4247–4261, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Son:2016:ACS

[1368] Yunsik Son, Sun-Young Ihm, Aziz Nasridinov, and Young-Ho Park. Adaptive convex skyline: a threshold-based project partitioned layer-based index for efficient-processing top-k queries in entrepreneurship applications. The Journal of Supercomputing, 72(11): 4262–4275, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-016-1826-4.pdf.

Faisal:2016:NFS

[1369] Ch. Muhammad Shahzad Faisal, Ali Daud, Faisal Imran, and Seungmin Rho. A novel framework for social web forums’ thread ranking based on semantics and post quality features. The Journal of Supercomputing, 72(11):4276–4295, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2016:ATB

[1370] Gangin Lee, Unil Yun, and Kyung-Min Lee. Analysis of tree-based uncertain frequent pattern mining techniques without pattern losses. The Journal of Supercomputing, 72(11): 4296–4318, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kang:2016:DSA

[1371] Jungho Kang, Geunil Park, and Jong Hyuk Park. Design of secure authentication scheme between devices based on zero-knowledge proofs in home automation service environments. The Journal of Supercomputing, 72(11):4319–4336, November 2016.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xiong:2016:STS

[1372] Wen Xiong, Zhibin Yu, Lieven Eeckhout, Zhengdong Bei, Fan Zhang, and Chengzhong Xu. ShenZhen transportation system (SZTS): a novel big data benchmark suite. The Journal of Supercomputing, 72(11):4337– 4364, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cheng:2016:AMM

[1373] Chi-Hsiang Cheng and Wei-Mei Chen. Application mapping onto mesh-based network-on-chip using constructive heuristic algorithms. The Journal of Supercomputing, 72(11):4365–4378, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2016:HTS

[1374] Wenbing Zhao, William Yang, Honglei Zhang, Jack Yang, Xiong Luo, Yueqin Zhu, Mary Yang, and Chaomin Luo. High-throughput state-machine replication using software transactional memory. The Journal of Supercomputing, 72(11):4379–4398, November 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kang:2016:GBP

[1375] Semin Kang, Sung-Soo Kim, Jongho Won, and Young-Min Kang. GPUbased parallel genetic approach to large-scale travelling salesman problem. The Journal of Supercomputing, 72(11):4399–4414, November 2016.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Escudero-Sahuquillo:2016:HPI

[1376] Jesús Escudero-Sahuquillo and Pedro Javier Garcia. High-performance interconnection networks in the Exascale and Big-Data Era. The Journal of Supercomputing, 72(12):4415– 4417, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-016-1893-6.pdf.

Vigneras:2016:BRA

[1377] Pierre Vignéras and Jean-Noël Quintin.

The BXI routing architecture for exascale supercomputer. The Journal of Supercomputing, 72(12):4418– 4437, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zahid:2016:CNR

[1378] Feroz Zahid, Ernst Gunnar Gran, Bartosz Bogdański, Bjørn Dag Johnsen, Tor Skeie, and Evangelos Tasoulas. Compact network reconfiguration in fat-trees. The Journal of Supercomputing, 72(12):4438–4467, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fuentes:2016:NUD

[1379] Pablo Fuentes, Enrique Vallejo, Cristóbal Camarero, Ramón Beivide, and Mateo Valero. Network unfairness in dragonfly topologies. The Journal of Supercomputing, 72(12):4468– 4496, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yebenes:2016:SSR

[1380] Pedro Yébenes, Jesus Escudero-Sahuquillo, Pedro J. García, and Francisco J. Quiles. Straightforward solutions to reduce HoL blocking in different Dragonfly fully-connected interconnection patterns. The Journal of Supercomputing, 72(12):4497– 4519, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Reano:2016:TRG

[1381] Carlos Reaño and Federico Silla. Tuning remote GPU virtualization for InfiniBand networks. The Journal of Supercomputing, 72(12):4520–4545, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Colombo:2016:ODC

[1382] Tommaso Colombo, Holger Fröning, Pedro Javier Garcìa, and Wainer Vandelli. Optimizing the data-collection time of a large-scale data-acquisition system through a simulation framework. The Journal of Supercomputing, 72(12):4546–4572, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shankar:2016:CBS

[1383] Dipti Shankar, Xiaoyi Lu, Md. Wasi ur Rahman, Nusrat Islam, and Dhabaleswar K. Panda. Characterizing and benchmarking stand-alone Hadoop MapReduce on modern HPC clusters. The Journal of Supercomputing, 72(12):4573–4600, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Andujar:2016:OSF

[1384] Francisco J. Andújar, Juan A. Villar, Francisco J. Alfaro, José L. Sánchez, and Jesus Escudero-Sahuquillo. An open-source family of tools to reproduce MPI-based workloads in interconnection network simulators. The Journal of Supercomputing, 72(12):4601– 4628, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Akbar:2016:EFT

[1385] Reza Akbar, Ali Asghar Etedalpour, and Farshad Safaei. An efficient faulttolerant routing algorithm in NoCs to tolerate permanent faults. The Journal of Supercomputing, 72(12):4629– 4650, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alrashed:2016:ESC

[1386] Saleh Alrashed, Jamal Alhiyafi, Aamir Shafi, and Nasro Min-Allah. An efficient schedulability condition for nonpreemptive real-time systems at common scheduling points. The Journal of Supercomputing, 72(12):4651– 4661, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hukerikar:2016:RRO

[1387] Saurabh Hukerikar and Robert F. Lucas. Rolex: resilience-oriented language extensions for extreme-scale systems. The Journal of Supercomputing, 72(12):4662–4695, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Guerra:2016:PCS

[1388] Aníbal Guerra, Jaime Lotero, and Sebastián Isaza. Performance comparison of sequential and parallel compression applications for DNA raw data. The Journal of Supercomputing, 72(12):4696–4717, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

AlBdaiwi:2016:EDN

[1389] Bader AlBdaiwi, Zaid Hussain, Anton Cerny, and Robert Aldred. Edgedisjoint node-independent spanning trees in dense Gaussian networks. The Journal of Supercomputing, 72(12): 4718–4736, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nezarat:2016:GTM

[1390] Amin Nezarat and Gh. Dastghaibyfard. A game theoretical model for profit maximization resource allocation in cloud environment with budget and deadline constraints. The Journal of Supercomputing, 72(12):4737– 4770, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dauwe:2016:HNP

[1391] Daniel Dauwe, Eric Jonardi, Ryan D. Friese, Sudeep Pasricha, Anthony A. Maciejewski, David A. Bader, and Howard Jay Siegel. HPC node performance and energy modeling with the co-location of applications. The Journal of Supercomputing, 72(12):4771– 4809, December 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2016:UCP

[1392] Sang-Min Park, Young-Gab Kim, and Doo-Kwon Baik. User-centric product recommendation on heterogeneous IoT device platform. The Journal of Supercomputing, 72(12):4810–4825, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alonso:2017:HPC

[1393] Pedro Alonso, José Ranilla, and Jesús Vigo-Aguiar. High-performance computing: the essential tool and the essential challenge. The Journal of Supercomputing, 73(1):1–3, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
016-1922-5.pdf.

Dieguez:2017:BBG

[1394] Adrián P. Diéguez, Margarita Amor, and Ramón Doallo. BPLG-BMCS: GPU-sorting algorithm using a tuning skeleton library. The Journal of Supercomputing, 73(1):4–16, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Olanda:2017:IHD

[1395] Ricardo Olanda, Mariano Pérez, Juan M. Orduña, and Silvia Rueda. Improving hybrid distributed architectures for interactive terrain visualization. The Journal of Supercomputing, 73(1):17–28, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aliaga:2017:ACT

[1396] José I. Aliaga, María Barreda, M. Asunción Castaño, Manuel F. Dolz, and Enrique S. Quintana-Ortí. Adapting concurrency throttling and voltagefrequency scaling for dense eigensolvers. The Journal of Supercomputing, 73(1):29–43, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alvarruiz:2017:IPW

[1397] Fernando Alvarruiz, Fernando Martínez Alzamora, and Antonio M. Vidal. Improving the performance of water distribution systems’ simulation on multicore systems. The Journal of Supercomputing, 73(1):44–56, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lopez-Portugues:2017:UHC

[1398] M. López-Portugués, J. A. López-Fernández, José Ranilla, R. G. Ayestarán, and F. Las-Heras. Using heterogeneous computing for scattering prediction in scenarios with several source configurations. The Journal of Supercomputing, 73(1):57–74, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Benito-Picazo:2017:RSS

[1399] F. Benito-Picazo, P. Cordero, M. Enciso, and A. Mora. Reducing the search space by closure and simplification paradigms. The Journal of Supercomputing, 73(1):75–87, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aldea:2017:BAS

[1400] Sergio Aldea, Diego R. Llanos, and Arturo Gonzalez-Escribano. BFCA+: automatic synthesis of parallel code with TLS capabilities. The Journal of Supercomputing, 73(1):88–99, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Losada:2017:RMA

[1401] Nuria Losada, Iván Cores, María J.

Martín, and Patricia González. Resilient MPI applications using an application-level checkpointing framework and ULFM. The Journal of Supercomputing, 73(1):100–113, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garzon:2017:AOE

[1402] E. M. Garzón, J. J. Moreno, and J. A.

Martínez. An approach to optimise the energy efficiency of iterative computation on integrated GPU–CPU systems. The Journal of Supercomputing, 73(1): 114–125, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alonso:2017:POT

[1403] Pedro Alonso, Raquel Cortina, F. J.

Rodríguez-Serrano, P. Vera-Candeas, M. Alonso-González, and José Ranilla. Parallel online time warping for realtime audio-to-score alignment in multicore systems. The Journal of Supercomputing, 73(1):126–138, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Catalan:2017:TEM

[1404] Sandra Catalán, Francisco D. Igual, Rafael Mayo, Rafael Rodríguez-Sánchez, and Enrique S. Quintana-Ortí. Time and energy modeling of a high-performance multi-threaded Cholesky factorization. The Journal of Supercomputing, 73(1):139–151, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Palacios:2017:HTM

[1405] Raúl H. Palacios, Antonio F. Díaz, Mancia Anguita, Julio Ortega, and Cristina Rodríguez-Quintana. Highthroughput multi-multicast transfers in data center networks. The Journal of Supercomputing, 73(1):152–163, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Migallon:2017:DMP

[1406] H. Migallón, V. Galiano, P. Piñol, O. López-Granado, and M. P. Malumbres. Distributed memory parallel approaches for HEVC encoder. The Journal of Supercomputing, 73(1):164– 175, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Artigas-Fuentes:2017:AVH

[1407] F. J. Artigas-Fuentes and J. M. Badía. Accessing very high dimensional spaces in parallel. The Journal of Supercomputing, 73(1):176–189, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Diaz-Honrubia:2017:CSA

[1408] A. J. Diaz-Honrubia, J. De Praeter, G. Van Wallendael, J. L. Martinez, P. Cuenca, J. M. Puerta, and J. A. Gamez. CTU splitting algorithm for H.264/AVC and HEVC simultaneous encoding. The Journal of Supercomputing, 73(1):190–202, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Belloch:2017:AMC

[1409] Jose A. Belloch, Fran J. Alventosa, Pedro Alonso, Enrique S. Quintana-Ortí, and Antonio M. Vidal. Accelerating multi-channel filtering of audio signal on ARM processors. The Journal of Supercomputing, 73(1):203–214, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia-Saiz:2017:CBK

[1410] Diego García-Saiz, Marta Zorrilla, and José Luis Bosque. A clustering-based knowledge discovery process for data centre infrastructure management. The Journal of Supercomputing, 73(1):215– 226, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Boratto:2017:ATP

[1411] Murilo Boratto, Pedro Alonso, Domingo Giménez, and Alexey Lastovetsky. Automatic tuning to performance modelling of matrix polynomials on multicore and multi-GPU systems. The Journal of Supercomputing, 73(1):227– 239, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sanjuan:2017:AVD

[1412] Gemma Sanjuan, Carles Tena, Tomàs Margalef, and Ana Cortés. Applying vectorization of diagonal sparse matrix to accelerate wind field calculation. The Journal of Supercomputing, 73(1):240–258, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-016-1696-9.pdf.

Cruz:2017:HPC

[1413] N. C. Cruz, J. L. Redondo, M. Berenguel, J. D. Álvarez, A. Becerra-Teron, and P. M. Ortigosa. High performance computing for the heliostat field layout evaluation. The Journal of Supercomputing, 73(1):259–276, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Diaz-Honrubia:2017:FHS

[1414] A. J. Diaz-Honrubia, J. L. Martinez, and P. Cuenca. A fast hybrid scalable H.264/AVC and HEVC encoder. The Journal of Supercomputing, 73(1): 277–290, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lopez-Fernandez:2017:IFP

[1415] J. A. López-Fernández, M. López-Portugués, and José Ranilla. Improving the FMM performance using optimal group size on heterogeneous system architectures. The Journal of Supercomputing, 73(1):291–301, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moreno:2017:ULP

[1416] J. J. Moreno, G. Ortega, E. Filatovas, J. A. Martínez, and Ester M. Garzón. Using low-power platforms for Evolutionary Multi-Objective Optimization algorithms. The Journal of Supercomputing, 73(1):302–315, January 2017.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Losada:2017:ARV

[1417] Nuria Losada, María J. Martín, and Patricia González. Assessing resilient versus stop-and-restart fault-tolerant solutions in MPI applications. The Journal of Supercomputing, 73(1):316– 329, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Perez:2017:EEL

[1418] Borja Pérez, Esteban Stafford, José Luis Bosque, and Ramón Beivide. Energy efficiency of load balancing for dataparallel applications in heterogeneous systems. The Journal of Supercomputing, 73(1):330–342, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alonso:2017:EMA

[1419] Pedro Alonso, P. Vera-Candeas, Raquel Cortina, and José Ranilla. An efficient musical accompaniment parallel system for mobile devices. The Journal of Supercomputing, 73(1):343–353, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gabaldon:2017:BMO

[1420] Eloi Gabaldon, Josep Lluis Lerida, Fernando Guirado, and Jordi Planes. Blacklist muti-objective genetic algorithm for energy saving in heterogeneous environments. The Journal of Supercomputing, 73(1):354–369, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/content/pdf/10.1007/s11227-
                                        016-1866-9.pdf.

Ortega:2017:APM

[1421] G. Ortega, A. M. Puertas, and E. M. Garzón. Accelerating the problem of microrheology in colloidal systems on a GPU. The Journal of Supercomputing, 73(1):370–383, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Uribe-Paredes:2017:ESP

[1422] R. Uribe-Paredes, E. Arias, D. Cazorla, and J. L. Sánchez. An efficiency study of a pivot-based algorithm for similarity search on a heterogeneous platform. The Journal of Supercomputing, 73(1): 384–401, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gonzalez:2017:CDM

[1423] Cesar González, Mariano Pérez, and Juan M. Orduña. Combining displacement mapping methods on the GPU for real-time terrain visualization. The Journal of Supercomputing, 73(1): 402–413, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cebrian-Marquez:2017:IIP

[1424] Gabriel Cebrián-Márquez, José Luis Martínez, and Pedro Cuenca. Inter and intra pre-analysis algorithm for HEVC. The Journal of Supercomputing, 73(1): 414–432, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lin:2017:JSS

[1425] Shaozhong Lin and Zhiqiang Xie. A Jacobi PCG solver for sparse linear systems on multi-GPU cluster. The Journal of Supercomputing, 73(1):433– 454, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Galiano:2017:GBH

[1426] V. Galiano, H. Migallón, V. Herranz, P. Piñol, O. López-Granado, and M. P. Malumbres. GPU-based HEVC intraprediction module. The Journal of Supercomputing, 73(1):455–468, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bondarenco:2017:CVS

[1427] Marcelo Bondarenco, Pablo Gamazo, and Pablo Ezzatti. A comparison of various schemes for solving the transport equation in many-core platforms. The Journal of Supercomputing, 73(1): 469–481, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jimenez:2017:EIM

[1428] Luis Ignacio Jiménez, Javier Plaza, and Antonio Plaza. Efficient implementation of morphological index for building/shadow extraction from remotely sensed images. The Journal of Supercomputing, 73(1):482–494, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia-Lucas:2017:PPE

[1429] David García-Lucas, Gabriel Cebrián-Márquez, and Pedro Cuenca. Parallelization and performance evaluation of open-source HEVC codecs. The Journal of Supercomputing, 73(1):495– 513, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Haut:2017:CIK

[1430] Juan Mario Haut, Mercedes Paoletti, Javier Plaza, and Antonio Plaza. Cloud implementation of the K-means algorithm for hyperspectral image analysis. The Journal of Supercomputing, 73(1): 514–529, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Belloch:2017:SWL

[1431] Jose A. Belloch, Balázs Bank, Francisco D. Igual, Enrique S. Quintana-Ortí, and Antonio M. Vidal. Solving Weighted Least Squares (WLS) problems on ARM-based architectures. The Journal of Supercomputing, 73(1): 530–542, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Migallon:2017:PAF

[1432] H. Migallón, P. Piñol, O. López-Granado, V. Galiano, and M. P. Malumbres. Performance analysis of frame partitioning in parallel HEVC encoders. The Journal of Supercomputing, 73(1):543–556, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lorenzo:2017:LSD

[1433] Oscar G. Lorenzo, Jorge Martínez, David L. Vilariño, Tomás F. Pena, José C. Cabaleiro, and Francisco F. Rivera. Landing sites detection using LiDAR data on manycore systems.

The Journal of Supercomputing, 73(1): 557–575, January 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Palacios:2017:ERD

[1434] Raúl H. Palacios, Cristina Rodríguez-Quintana, Antonio F. Díaz, Mancia Anguita, and Julio Ortega. Evaluation of redundant data storage in clusters based on multi-multicast and local storage. The Journal of Supercomputing, 73(1):576–590, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cruz:2017:PTL

[1435] N. C. Cruz, J. L. Redondo, J. D.

Álvarez, M. Berenguel, and P. M. Ortigosa. A parallel Teaching-Learning-Based Optimization procedure for automatic heliostat aiming. The Journal of Supercomputing, 73(1):591–606, January 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Malyshkin:2017:PCT

[1436] Victor Malyshkin. Parallel computing technologies 2016. The Journal of Supercomputing, 73(2):607–608, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
016-1843-3.pdf.

Campos:2017:MPS

[1437] Ricardo Silva Campos, Bernardo Martins Rocha, Marcelo Lobosco, and Rodrigo Weber dos Santos. Multilevel parallelism scheme in a genetic algorithm applied to cardiac models with massspring systems. The Journal of Supercomputing, 73(2):609–623, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Akhmed-Zaki:2017:ITD

[1438] Darkhan Akhmed-Zaki, Danil Lebedev, and Vladislav Perepelkin. Implementation of a three dimensional three-phase fluid flow (“oil-water-gas”) numerical model in LuNA fragmented programming system. The Journal of Supercomputing, 73(2):624–630, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Menshov:2017:HSI

[1439] Igor Menshov and Pavel Pavlukhin. Highly scalable implementation of an implicit matrix-free solver for gas dynamics on GPU-accelerated clusters. The Journal of Supercomputing, 73(2): 631–638, February 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Borisenko:2017:GPB

[1440] Andrey Borisenko, Michael Haidl, and Sergei Gorlatch. A GPU parallelization of branch-and-bound for multiproduct batch plants optimization. The Journal of Supercomputing, 73(2):639– 651, February 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2017:ODD

[1441] Ren Chen, Shreyas G. Singapura, and Viktor K. Prasanna. Optimal dynamic data layouts for 2D FFT on 3D memory integrated FPGA. The Journal of Supercomputing, 73(2):652–663, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rojek:2017:PMM

[1442] Krzysztof Rojek and Roman Wyrzykowski.

Performance modeling of 3D MP-DATA simulations on GPU cluster. The Journal of Supercomputing, 73 (2):664–675, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-016-1774-z.pdf.

Tangherloni:2017:GSS

[1443] Andrea Tangherloni, Marco S. Nobile, Paolo Cazzaniga, Daniela Besozzi, and Giancarlo Mauri. Gillespie’s Stochastic Simulation Algorithm on MIC coprocessors. The Journal of Supercomputing, 73(2):676–686, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bandman:2017:PEV

[1444] Olga Bandman. Parallelization efficiency versus stochasticity in simulation reaction-diffusion by cellular automata. The Journal of Supercomputing, 73(2):687–699, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ozelim:2017:IDF

[1445] Luan Carlos de S. M. Ozelim, André Luís B. Cavalcante, and Jan M. Baetens. On the iota-delta function: a link between cellular automata and partial differential equations for modeling advection-dispersion from a constant source. The Journal of Supercomputing, 73(2):700–712, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hasanov:2017:HRC

[1446] Khalid Hasanov and Alexey Lastovetsky. Hierarchical redesign of classic MPI reduction algorithms. The Journal of Supercomputing, 73(2):713– 725, February 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Malyshkin:2017:SDD

[1447] Victor Malyshkin, Vladislav Perepelkin, and Georgy Schukin. Scalable distributed data allocation in LuNA fragmented programming system. The Journal of Supercomputing, 73(2):726– 732, February 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Naranjo:2017:PSP

[1448] Paola G. Vinueza Naranjo, Mohammad Shojafar, Habib Mostafaei, Zahra Pooranian, and Enzo Baccarelli. P-SEP: a prolong stable election routing algorithm for energy-limited heterogeneous fog-supported wireless sensor networks. The Journal of Supercomputing, 73(2):733–755, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Deldari:2017:CDC

[1449] Arash Deldari, Mahmoud Naghibzadeh, and Saeid Abrishami. CCA: a deadlineconstrained workflow scheduling algorithm for multicore resources on the cloud. The Journal of Supercomputing, 73(2):756–781, February 2017.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Teng:2017:EEV

[1450] Fei Teng, Lei Yu, Tianrui Li, Danting Deng, and Frédéric Magoulès. Energy efficiency of VM consolidation in IaaS clouds. The Journal of Supercomputing, 73(2):782–809, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahn:2017:DPA

[1451] Jinhyun Ahn, Dong-Hyuk Im, Taewhi Lee, and Hong-Gee Kim. A dynamic and parallel approach for repetitive prime labeling of XML with MapReduce. The Journal of Supercomputing, 73(2):810–836, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2017:MIH

[1452] Kai-Siou Wu, Yi-Chun Wang, and Justie Su-Tzu Juan. Mutually independent hamiltonianicity of Cartesian product graphs. The Journal of Supercomputing, 73(2):837–865, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nejad:2017:PRR

[1453] Hossein Mohammadi Nejad, Naser Movahhedinia, and Mohammad Reza Khayyambashi. Provisioning required reliability of wireless data communication in smart grid neighborhood area networks. The Journal of Supercomputing, 73(2):866–886, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jafri:2017:GAR

[1454] Rabia Jafri. A GPU-accelerated realtime contextual awareness application for the visually impaired on Google’s Project Tango device. The Journal of Supercomputing, 73(2):887–899, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

dAuriol:2017:HBF

[1455] Brian J. d’Auriol. High-bandwidth flexible interconnections in the alloptical linear array with a reconfigurable pipelined bus system (OLARPBS) optical conduit parallel computing model. The Journal of Supercomputing, 73(2):900–922, February 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2017:SSE

[1456] Brian Kim, Kostas Psannis, and Harish Bhaskar. Special section on emerging multimedia technology for smart surveillance system with IoT environment. The Journal of Supercomputing, 73(3):923–925, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-016-1939-9.pdf.

Park:2017:SAA

[1457] Hwin Dol Park, Ok-Gee Min, and Yong-Ju Lee. Scalable architecture for an automated surveillance system using edge computing. The Journal of Supercomputing, 73(3):926–939, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jun:2017:DUH

[1458] Dongsan Jun, Sung-Chang Lim, Jinho Lee, Hahyun Lee, Jongho Kim, Jungwon Kang, Jinwook Seok, Younhee Kim, Soon heung Jung, Hui-Yong Kim, and Jin Soo Choi. Development of an ultra-HD HEVC encoder using SIMD implementation and fast encoding schemes for smart surveillance system. The Journal of Supercomputing, 73(3):940–960, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chatterji:2017:AFE

[1459] Sanjay Chatterji, Nitish Varshney, and Ranjan Kumar Rahul. Aspect-FrameNet: a frameNet extension for analysis of sentiments around product aspects. The Journal of Supercomputing, 73(3):961–972, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alsmirat:2017:ISC

[1460] Mohammad A. Alsmirat, Yaser Jararweh, Islam Obaidat, and Brij B. Gupta. Internet of surveillance: a cloud supported large-scale wireless surveillance system. The Journal of Supercomputing, 73(3):973–992, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Batalla:2017:EMO

[1461] Jordi Mongay Batalla, Constandinos X. Mavromoustakis, George Mastorakis, Daniel Négru, and Eugen Borcoci. Evolutionary multiobjective optimization algorithm for multimedia delivery in critical applications through Content-Aware Networks. The Journal of Supercomputing, 73(3):993–1016, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        016-1731-x.pdf.

Jung:2017:ERS

[1462] Tae jun Jung, Yo-Won Jeong, Hong rae Lee, and Kwang deok Seo. Errorresilient surveillance video transmission based on a practical joint sourcechannel distortion computing model. The Journal of Supercomputing, 73(3): 1017–1043, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kokkonis:2017:RTW

[1463] George Kokkonis, Kostas E. Psannis, Manos Roumeliotis, and Dan Schonfeld. Real-time wireless multisensory smart surveillance with 3D-HEVC streams for Internet-of-Things (IoT). The Journal of Supercomputing, 73(3): 1044–1062, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2017:FCU

[1464] Byung-Gyu Kim. Fast coding unit (CU) determination algorithm for highefficiency video coding (HEVC) in smart surveillance application. The Journal of Supercomputing, 73(3): 1063–1084, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tewari:2017:CNU

[1465] Aakanksha Tewari and B. B. Gupta. Cryptanalysis of a novel ultralightweight mutual authentication protocol for IoT devices using RFID tags.

The Journal of Supercomputing, 73(3): 1085–1102, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2017:GBS

[1466] Donghyeok Lee and Namje Park.

Geocasting-based synchronization of Almanac on the maritime cloud for distributed smart surveillance. The Journal of Supercomputing, 73(3):1103– 1118, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [1467].

Lee:2017:EGB

[1467] Donghyeok Lee and Namje Park. Erratum to: Geocasting-based synchronization of Almanac on the maritime cloud for distributed smart surveillance. The Journal of Supercomputing, 73(3):1119, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-1967-0.pdf. See [1466].

Paul:2017:MOD

[1468] Nihal Paul, Ashish Singh, Abhishek Midya, Partha Pratim Roy, and Debi Prosad Dogra. Moving object detection using modified temporal differencing and local fuzzy thresholding. The Journal of Supercomputing, 73(3): 1120–1139, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2017:NCS

[1469] Kwangho Kim, InJung Kim, and Jongin Lim. National cyber security enhancement scheme for intelligent surveillance capacity with public IoT environment. The Journal of Supercomputing, 73(3):1140–1151, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2017:BBS

[1470] Boohyung Lee and Jong-Hyouk Lee. Blockchain-based secure firmware update for embedded devices in an Internet of Things environment. The Journal of Supercomputing, 73(3):1152– 1167, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sharma:2017:RBR

[1471] Suraj Sharma, Deepak Puthal, Sanjay Kumar Jena, Albert Y. Zomaya, and Rajiv Ranjan. Rendezvous based routing protocol for wireless sensor networks with mobile sink. The Journal of Supercomputing, 73(3):1168– 1188, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [1472].

Sharma:2017:ERB

[1472] Suraj Sharma, Deepak Puthal, Sanjay Kumar Jena, Albert Y. Zomaya, and Rajiv Ranjan. Erratum to: Rendezvous based routing protocol for wireless sensor networks with mobile sink. The Journal of Supercomputing, 73(3):1189–1191, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        016-1851-3.pdf. See [1471].

Modi:2017:VLS

[1473] Chirag N. Modi and Kamatchi Acha. Virtualization layer security challenges and intrusion detection/prevention systems in cloud computing: a comprehensive review. The Journal of Supercomputing, 73(3):1192–1234, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rizk-Allah:2017:NFF

[1474] Rizk M. Rizk-Allah, Ragab A. El-Sehiemy, Suash Deb, and Gai-Ge Wang. A novel fruit fly framework for multi-objective shape design of tubular linear synchronous motor. The Journal of Supercomputing, 73(3): 1235–1256, March 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karaata:2017:OAS

[1475] Mehmet Hakan Karaata and Fawaz M.

Alazemi. An optimal adaptive stabilizing algorithm for two edge-disjoint paths problem. The Journal of Supercomputing, 73(3):1257–1273, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Raei:2017:APM

[1476] Hassan Raei and Nasser Yazdani. Analytical performance models for resource allocation schemes of cloudlet in mobile cloud computing. The Journal of Supercomputing, 73(3):1274–1305, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2017:PEH

[1477] Kihong Lee, DongWoo Lee, Sungkil Lee, and Young Ik Eom. Power-efficient and high-performance block I/O framework for mobile virtualization systems. The Journal of Supercomputing, 73(4):

1307–1321, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sardroud:2017:ECP

[1478] Asghar Asgharian Sardroud and Alireza Bagheri. Embedding cycles and paths on solid grid graphs. The Journal of Supercomputing, 73(4):1322–1336, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cha:2017:AMR

[1479] Myung-Hoon Cha, Dong-Oh Kim, Hong-Yeon Kim, and Young-Kyun Kim. Adaptive metadata rebalance in exascale file system. The Journal of Supercomputing, 73(4):1337–1359, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abbas-Turki:2017:RSR

[1480] L. A. Abbas-Turki and S. Graillat. Resolving small random symmetric linear systems on graphics processing units. The Journal of Supercomputing, 73(4): 1360–1386, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karimi:2017:QAS

[1481] Mohammad Bagher Karimi, Ayaz Isazadeh, and Amir Masoud Rahmani. QoS-aware service composition in cloud computing using data mining techniques and genetic algorithm. The Journal of Supercomputing, 73(4): 1387–1415, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Whalen:2017:SDC

[1482] Shannon Whalen, Arijit Ghosh, and Shivakumar Sastry. Simulating distributed and coordinated conveying systems. The Journal of Supercomputing, 73(4):1416–1437, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sagharichian:2017:IIP

[1483] Masoud Sagharichian and Hassan Naderi. Intelligent and independent processes for overcoming big graphs. The Journal of Supercomputing, 73(4): 1438–1466, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cattaneo:2017:EEA

[1484] Giuseppe Cattaneo, Umberto Ferraro Petrillo, Raffaele Giancarlo, and Gianluca Roscigno. An effective extension of the applicability of alignment-free biological sequence comparison algorithms with Hadoop. The Journal of Supercomputing, 73(4):1467–1483, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Azizi:2017:FHP

[1485] Sadoon Azizi, Naser Hashemi, and Ahmad Khonsari. A flexible and highperformance data center network topology. The Journal of Supercomputing, 73(4):1484–1503, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Irshad:2017:CPP

[1486] Azeem Irshad and Shehzad Ashraf Chaudhry. Comments on “A privacy preserving three-factor authentication protocol for e-health clouds”.

The Journal of Supercomputing, 73(4): 1504–1508, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See [1347].

Li:2017:CDS

[1487] Peng Li, Tin-Yu Wu, Xin-Ming Li, Hong Luo, and Mohammad S. Obaidat. Constructing data supply chain based on layered PROV. The Journal of Supercomputing, 73(4):1509–1531, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2017:CBM

[1488] Chen Li, Wei Wei, Jiaxue Li, and Wei Song. A cloud-based monitoring system via face recognition using Gabor and CS–LBP features. The Journal of Supercomputing, 73(4):1532–1546, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
016-1840-6.pdf.

Nadjar:2017:LDA

[1489] Ali Nadjar, Saeid Abrishami, and Hossein Deldari. Load dispersion-aware VM placement in favor of energyperformance tradeoff. The Journal of Supercomputing, 73(4):1547–1566, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Meyer:2017:MFR

[1490] Michael Meyer, Yuichi Okuyama, and Abderazek Ben Abdallah. Microring fault-resilient photonic networkon-chip for reliable high-performance many-core systems. The Journal of Supercomputing, 73(4):1567–1599, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Momtazpour:2017:TDG

[1491] Mahmoud Momtazpour. Towards designing a green data center farm for Internet services: Iran’s case study. The Journal of Supercomputing, 73(4): 1600–1628, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kouetcha:2017:USP

[1492] Daniella Nguemalieu Kouetcha, Hamidréza Ramézani, and Nathalie Cohaut. Ultrafast scalable parallel algorithm for the radial distribution function histogramming using MPI maps. The Journal of Supercomputing, 73(4): 1629–1653, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2017:EEN

[1493] Ji Wu, Dezun Dong, Xiangke Liao, and Li Wang. Energy-efficient NoC with multi-granularity power optimization. The Journal of Supercomputing, 73(4): 1654–1671, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Roy:2017:AAC

[1494] Deepsubhra Guha Roy, Debashis De, Anwesha Mukherjee, and Rajkumar Buyya. Application-aware cloudlet selection for computation offloading in multi-cloudlet environment. The Journal of Supercomputing, 73(4):1672– 1690, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2017:TMI

[1495] Jingjin Wu, Xuanxing Xiong, Eduardo Berrocal, Jia Wang, and Zhiling Lan. Topology mapping of irregular parallel applications on torus-connected supercomputers. The Journal of Supercomputing, 73(4):1691–1714, April 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shahbazian:2017:DCT

[1496] Reza Shahbazian and Seyed Ali Ghorashi. Distributed cooperative target detection and localization in decentralized wireless sensor networks. The Journal of Supercomputing, 73(4): 1715–1732, April 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khan:2017:ASQ

[1497] Mozammel H. A. Khan, Himanshu Thapliyal, and Edgard Munoz-Coreas. Automatic synthesis of quaternary quantum circuits. The Journal of Supercomputing, 73(5):1733–1759, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2017:PSM

[1498] Wangdong Yang, Kenli Li, and Keqin Li. A parallel solving method for block-tridiagonal equations on CPU– GPU heterogeneous computing systems. The Journal of Supercomputing, 73(5):1760–1781, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2017:CDP

[1499] Jinjing Li, Qingkui Chen, and Bocheng Liu. Classification and disease probability prediction via machine learning programming based on multi-GPU cluster MapReduce system. The Journal of Supercomputing, 73(5):1782– 1809, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Feher:2017:DSI

[1500] Péter Fehér, Márk Asztalos, Tamás Vajk, Tamás Mészáros, and László Lengyel. Detecting subgraph isomorphism with MapReduce. The Journal of Supercomputing, 73(5):1810–1851, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Celebi:2017:ISS

[1501] M. Serdar Celebi, Ahmet Duran, Figen Oztoprak, Mehmet Tuncel, and Bora Akaydin. On the improvement of a scalable sparse direct solver for unsymmetrical linear equations. The Journal of Supercomputing, 73(5):1852–1904, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khan:2017:TSH

[1502] Minhaj Ahmad Khan. Task scheduling for heterogeneous systems using an incremental approach. The Journal of Supercomputing, 73(5):1905–1928, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shehab:2017:ACI

[1503] Mohammed Shehab, Mahmoud Al-Ayyoub, Yaser Jararweh, and Moath Jarrah. Accelerating compute-intensive image segmentation algorithms using GPUs. The Journal of Supercomputing, 73(5):1929–1951, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Darabkh:2017:ICA

[1504] Khalid A. Darabkh, Wijdan Y. Albtoush, and Iyad F. Jafar. Improved clustering algorithms for target tracking in wireless sensor networks. The Journal of Supercomputing, 73(5): 1952–1977, May 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kommeri:2017:EED

[1505] Jukka Kommeri, Tapio Niemi, and Jukka K. Nurminen. Energy efficiency of dynamic management of virtual cluster with heterogeneous hardware. The Journal of Supercomputing, 73(5):1978–2000, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rajabzadeh:2017:EAF

[1506] Mehdi Rajabzadeh and Abolfazl Toroghi Haghighat. Energy-aware framework with Markov chain-based parallel simulated annealing algorithm for dynamic management of virtual machines in cloud data centers. The Journal of Supercomputing, 73(5):2001–2017, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ambursa:2017:PSO

[1507] Faruku Umar Ambursa, Rohaya Latip, Azizol Abdullah, and Shamala Subramaniam. A particle swarm optimization and min-max-based workflow scheduling algorithm with QoS satisfaction for service-oriented grids. The Journal of Supercomputing, 73(5):

2018–2051, May 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2017:HDT

[1508] Xigao Li and Lin Qian. A hybrid disaster-tolerant model with DDF technology for MooseFS open-source distributed file system. The Journal of Supercomputing, 73(5):2052–2068, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Son:2017:RVU

[1509] Seung Woo Son, Saba Sehrish, Wei keng Liao, Ron Oldfield, and Alok Choudhary. Reducing I/O variability using dynamic I/O path characterization in petascale storage systems. The Journal of Supercomputing, 73(5): 2069–2097, May 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vardi:2017:HCA

[1510] Fatemeh Vardi, Ahmad Khadem-Zadeh, and Midia Reshadi. A heuristic clustering approach to use caseaware application-specific network-on-chip synthesis. The Journal of Supercomputing, 73(5):2098–2129, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ever:2017:PAC

[1511] Enver Ever. Performability analysis of cloud computing centers with large numbers of servers. The Journal of Supercomputing, 73(5):2130–2156, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jin:2017:DED

[1512] Jing Jin, Guoming Lai, Xiaola Lin, and Xianggao Cai. DLRankSVM: an efficient distributed algorithm for linear RankSVM. The Journal of Supercomputing, 73(5):2157–2186, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kiani:2017:MMA

[1513] Vahdaneh Kiani and Midia Reshadi. Mapping multiple applications onto 3D NoC-based MPSoCs supporting wireless links. The Journal of Supercomputing, 73(5):2187–2213, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Komosinski:2017:MCE

[1514] Maciej Komosinski and Szymon Ulatowski. Multithreaded computing in evolutionary design and in artificial life simulations. The Journal of Supercomputing, 73(5):2214–2228, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        016-1923-4.pdf.

Mendez:2017:CDD

[1515] Mariano Méndez and Fernando G. Tinetti. Change-driven development for scientific software. The Journal of Supercomputing, 73(5):2229–2257, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shen:2017:PMB

[1516] Chao Shen, Weiqin Tong, Jenq-Neng Hwang, and Qiang Gao. Performance modeling of big data applications in the cloud centers. The Journal of Supercomputing, 73(5):2258–2283, May 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gonzalez-Alvarez:2017:HMO

[1517] David L. González-Álvarez, Miguel A.

Vega-Rodríguez, and Álvaro Rubio-Largo. A hybrid MPI/OpenMP parallel implementation of NSGA–II for finding patterns in protein sequences. The Journal of Supercomputing, 73(6): 2285–2312, June 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

AlEbrahim:2017:TSH

[1518] Shaikhah AlEbrahim and Imtiaz Ahmad. Task scheduling for heterogeneous computing systems. The Journal of Supercomputing, 73(6):2313–2338, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2017:MQP

[1519] Soo-Hyung Kim, Kyong-Ha Lee, and Yoon-Joon Lee. Multi-query processing of XML data streams on multicore. The Journal of Supercomputing, 73(6): 2339–2368, June 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fanfakh:2017:ECR

[1520] Ahmed Fanfakh, Jean-Claude Charr, Raphaël Couturier, and Arnaud Giersch. Energy consumption reduction for asynchronous message-passing applications. The Journal of Supercomputing, 73(6):2369–2401, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Choi:2017:MCR

[1521] I. Stephen Choi and Donald Yeung. Multi-cache resizing via greedy coordinate descent. The Journal of Supercomputing, 73(6):2402–2429, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khorsand:2017:AAT

[1522] Reihaneh Khorsand, Faramarz Safi-Esfahani, Naser Nematbakhsh, and Mehran Mohsenzade. ATSDS: adaptive two-stage deadline-constrained workflow scheduling considering run-time circumstances in cloud computing environments. The Journal of Supercomputing, 73(6):2430–2455, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Granado-Criado:2017:HCH

[1523] José M. Granado-Criado and Miguel A. Vega-Rodríguez. Hardware coprocessors for high-performance symmetric cryptography. The Journal of Supercomputing, 73(6):2456–2482, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yamamoto:2017:EPT

[1524] Hiroshi Yamamoto, Yutaro Kuriyama, and Hiroshi Ishii. Estimating the processing time of a model of cloud computing. The Journal of Supercomputing, 73(6):2483–2493, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yamamoto:2017:QMI

[1525] Hiroshi Yamamoto, Yusuke Hiraide, and Hiroshi Ishii. A quantitative measure of the information leaked from queries to search engines and a scheme to reduce it. The Journal of Supercomputing, 73(6):2494–2505, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2017:PMO

[1526] Dali Li, Chuanfu Xu, Bin Cheng, Min Xiong, Xiang Gao, and Xiaogang Deng. Performance modeling and optimization of parallel LU–SGS on many-core processors for 3D high-order CFD simulations. The Journal of Supercomputing, 73(6):2506–2524, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Quislant:2017:LID

[1527] Ricardo Quislant, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Leveraging irrevocability to deal with signature saturation in hardware transactional memory. The Journal of Supercomputing, 73(6):2525–2557, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bhatia:2017:DSM

[1528] Tarunpreet Bhatia and A. K. Verma.

Data security in mobile cloud computing paradigm: a survey, taxonomy and open research issues. The Journal of Supercomputing, 73(6):2558–2631, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khalil:2017:REE

[1529] Enan A. Khalil and Suat Ozdemir. Reliable and energy efficient topology control in probabilistic Wireless Sensor Networks via multi-objective optimization. The Journal of Supercomputing, 73(6):2632–2656, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Won:2017:MMA

[1530] Heesun Won, Minh Chau Nguyen, Myeong-Seon Gil, Yang-Sae Moon, and Kyu-Young Whang. Moving metadata from ad hoc files to database tables for robust, highly available, and scalable HDFS. The Journal of Supercomputing, 73(6):2657–2681, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Siadat:2017:IFF

[1531] Safieh Siadat, Amir Masoud Rahmani, and Hamidreza Navid. Identifying fake feedback in cloud trust management systems using feedback evaluation component and Bayesian game model. The Journal of Supercomputing, 73(6):2682–2704, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dang:2017:LOS

[1532] Khanh N. Dang, Michael Meyer, Yuichi Okuyama, and Abderazek Ben Abdallah. A low-overhead soft-hard fault-tolerant architecture, design and management scheme for reliable highperformance many-core 3D-NoC systems. The Journal of Supercomputing, 73(6):2705–2729, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Panda:2017:SBT

[1533] Sanjaya K. Panda and Prasanta K. Jana. SLA-based task scheduling algorithms for heterogeneous multi-cloud environment. The Journal of Supercomputing, 73(6):2730–2762, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ghorbel:2017:PCC

[1534] Amal Ghorbel, Mahmoud Ghorbel, and Mohamed Jmaiel. Privacy in cloud computing environments: a survey and research challenges. The Journal of Supercomputing, 73(6):2763–2800, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hasan:2017:ERH

[1535] K. M. Azharul Hasan and Md Abu Hanif Shaikh. Efficient representation of higher-dimensional arrays by dimension transformations. The Journal of Supercomputing, 73(6):2801–2822, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2017:DHP

[1536] Peng Zhang, Chao Yang, Chungang Chen, Xingliang Li, Xueshun Shen, and Feng Xiao. Development of a hybrid parallel MCV-based high-order global shallow-water model. The Journal of Supercomputing, 73(6):2823–2842, June 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hung:2017:ANE

[1537] Jason C. Hung and Gangman Yi. Advances in next era cloud-empowered computing and techniques. The Journal of Supercomputing, 73(7):2843– 2850, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-017-2099-2.pdf.

Kim:2017:DHG

[1538] Yeji Kim, Seoungjae Cho, Simon Fong, Yong Woon Park, and Kyungeun Cho. Design of hand gesture interaction framework on clouds for multiple users. The Journal of Supercomputing, 73(7): 2851–2866, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2017:SPP

[1539] Yan Li, Jong-Hyuk Park, and Byeong-Seok Shin. A shortest path planning algorithm for cloud computing environment based on multi-access point topology analysis for complex indoor spaces. The Journal of Supercomputing, 73(7): 2867–2880, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moon:2017:DII

[1540] Daesung Moon, Hyungjin Im, Ikkyun Kim, and Jong Hyuk Park. DTB-IDS: an intrusion detection system based on decision tree using behavior analysis for preventing APT attacks. The Journal of Supercomputing, 73(7): 2881–2895, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fakhfakh:2017:DSC

[1541] Fairouz Fakhfakh, Hatem Hadj Kacem, and Ahmed Hadj Kacem. Dealing with structural changes on provisioning resources for deadline-constrained workflow. The Journal of Supercomputing, 73(7):2896–2918, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2017:SBD

[1542] Yong hwan Kim, Hyun kyo Lim, Kyoung han Kim, and Youn-Hee Han. A SDN-based distributed mobility management in LTE/EPC network. The Journal of Supercomputing, 73(7): 2919–2933, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jo:2017:SII

[1543] Youngho Jo, Jincheol Woo, Heajin Kim, and Mincheol Whang. Social identification-issuing system (SIIS) using micro-movement in social lifelogging. The Journal of Supercomputing, 73(7):2934–2948, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2017:ECS

[1544] Longbin Chen, Meikang Qiu, Wenyun Dai, and Houcine Hassan. An efficient cloud storage system for tele-health services. The Journal of Supercomputing, 73(7):2949–2965, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chang:2017:NCA

[1545] Min-Kuan Chang, Yu-Wei Chan, Hsiao-Ping Tsai, Ting-Chen Chen, and Min-Han Chuang. Node connectivity analysis in cloud-assisted IoT environments. The Journal of Supercomputing, 73(7):2966–2986, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2017:DCO

[1546] Shih-Jung Wu, Rui-Dong Chiang, and Zheng-Hong Ji. Development of a Chinese opinion-mining system for application to Internet online forums. The Journal of Supercomputing, 73(7): 2987–3001, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pei:2017:PCS

[1547] Yan Pei. Principal component selection using interactive evolutionary computation. The Journal of Supercomputing, 73(7):3002–3020, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Won:2017:NRE

[1548] Chee Sun Won and Seung-Won Jung. Near-reversible efficient image resizing for devices supporting different spatial resolutions. The Journal of Supercomputing, 73(7):3021–3037, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2017:GTC

[1549] Byoung Cheul Kim, Daejune Ko, Unsoo Jang, Hyuksoo Han, and Eui Chul Lee. 3D Gaze tracking by combining eye- and facial-gaze vectors. The Journal of Supercomputing, 73(7): 3038–3052, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2017:NAF

[1550] Zhen Li, Haiqing Pan, Wenhao Liu, Fei Xu, Zigang Cao, and Gang Xiong. A network attack forensic platform against HTTP evasive behavior.

The Journal of Supercomputing, 73(7): 3053–3064, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2017:CCA

[1551] Jeong Kyu Lee, Seo Yeon Moon, and Jong Hyuk Park. CloudRPS: a cloud analysis based enhanced ransomware prevention system. The Journal of Supercomputing, 73(7):3065–3084, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2017:SPM

[1552] Wei-Chen Wu. A secret push messaging service in VANET clouds. The Journal of Supercomputing, 73(7):3085–3097, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moon:2017:LCM

[1553] Hae-Min Moon, Chang Ho Seo, and Sung Bum Pan. A low-cost media quality enhancement resolution upconversion for mobile cloud. The Journal of Supercomputing, 73(7):3098– 3111, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhuang:2017:LBB

[1554] Yan Zhuang, Simon Fong, Meng Yuan, Yunsick Sung, Kyungeun Cho, and Raymond K. Wong. Location-based big data analytics for guessing the next Foursquare check-ins. The Journal of Supercomputing, 73(7):3112–3127, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhuang:2017:PNT

[1555] Yan Zhuang, Simon Fong, Meng Yuan, Yunsick Sung, Kyungeun Cho, and Raymond K. Wong. Predicting the next turn at road junction from big traffic data. The Journal of Supercomputing, 73(7):3128–3148, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chang:2017:NSD

[1556] Shih-Hao Chang and Hung Mao-Sheng. A novel software-defined wireless network architecture to improve ship area network performance. The Journal of Supercomputing, 73(7):3149–3160, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Feng:2017:ADS

[1557] Qingsong Feng, Yabin Zhang, Chao Li, Zheng Dou, and Jin Wang. Anomaly detection of spectrum in wireless communication via deep auto-encoders. The Journal of Supercomputing, 73(7): 3161–3178, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cho:2017:BDP

[1558] Wonhee Cho and Eunmi Choi. Big data pre-processing methods with vehicle driving data using MapReduce techniques. The Journal of Supercomputing, 73(7):3179–3195, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2017:DRA

[1559] Yi-Hsuan Lee, Kuo-Chan Huang, Meng-Ru Shieh, and Kuan-Chou Lai.

Distributed resource allocation in federated clouds. The Journal of Supercomputing, 73(7):3196–3211, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chang:2017:ACF

[1560] Jieh-Ren Chang, You-Shyang Chen, Hong-Wun Lin, and Hwai-Tsu Hu. An advanced computing in fuzzy rulebased preprocessing design of image filters’ system for removing impulse noises. The Journal of Supercomputing, 73(7):3212–3228, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sharma:2017:ECR

[1561] Vishal Sharma, Kathiravan Srinivasan, Rajesh Kumar, Han-Chieh Chao, and Kai-Lung Hua. Efficient cooperative relaying in flying ad hoc networks using fuzzy-bee colony optimization. The Journal of Supercomputing, 73(7): 3229–3259, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yong:2017:IMS

[1562] Binbin Yong, Gaofeng Zhang, Huaming Chen, and Qingguo Zhou. Intelligent monitor system based on cloud and convolutional neural networks. The Journal of Supercomputing, 73(7): 3260–3276, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2017:EEC

[1563] Jin Wang, Jiayi Cao, Sai Ji, and Jong Hyuk Park. Energy-efficient cluster-based dynamic routes adjustment approach for wireless sensor networks with mobile sinks. The Journal of Supercomputing, 73(7):3277–3290, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2017:MLB

[1564] Deok Gyu Lee. A multi-level behavior network-based dangerous situation recognition method in cloud computing environments. The Journal of Supercomputing, 73(7):3291–3306, July 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2017:NAM

[1565] Jong Hyuk Park, Yunsick Sung, Pradip Kumar Sharma, Young-Sik Jeong, and Gangman Yi. Novel assessment method for accessing private data in social network security services. The Journal of Supercomputing, 73(7): 3307–3325, July 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hung:2017:EMC

[1566] Che-Lun Hung, Frédéric Magoulès, Meikang Qiu, Robert C. Hsu, and Chun-Yuan Lin. Embedded multicore computing and applications. The Journal of Supercomputing, 73(8): 3327–3332, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-017-2107-6.pdf.

Lin:2017:CSM

[1567] Ying-Chin Lin, Chun-Yao Wang, and Jing-Yun Zeng. A case study on mathematical expression recognition to GPU. The Journal of Supercomputing, 73(8):

3333–3343, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hu:2017:EED

[1568] Wei Hu, Hong Guo, Kai Zhang, Jun Liu, Xiaoming Liu, and Qingsong Shi. An energy-efficient design of microkernel-based on-chip OS for NOC-based manycore system. The Journal of Supercomputing, 73(8): 3344–3365, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Paulino:2017:CCR

[1569] Hervé Paulino and Nuno Delgado.

Cache-conscious run-time decomposition of data parallel computations. The Journal of Supercomputing, 73(8): 3366–3390, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lin:2017:CTD

[1570] Chun-Yuan Lin, Huang Ting Yen, and Che-Lun Hung. Compressing threedimensional sparse arrays using interand intra-task parallelization strategies on Intel Xeon and Xeon Phi. The Journal of Supercomputing, 73(8): 3391–3410, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahamed:2017:EIJ

[1571] Abal-Kassim Cheik Ahamed and Frédéric Magoulès. Efficient implementation of Jacobi iterative method for large sparse linear systems on graphic processing units. The Journal of Supercomputing, 73(8):3411–3432, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Song:2017:GIS

[1572] Weijing Song, Ze Deng, Lizhe Wang, Bo Du, Peng Liu, and Ke Lu. G-IK-SVD: parallel IK–SVD on GPUs for sparse representation of spatial big data. The Journal of Supercomputing, 73(8):3433–3450, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Junior:2017:IID

[1573] José Raniery Ferreira Junior, Marcelo Costa Oliveira, and Paulo Mazzoncini de Azevedo-Marques. Integrating 3D image descriptors of margin sharpness and texture on a GPU-optimized similar pulmonary nodule retrieval engine. The Journal of Supercomputing, 73(8): 3451–3467, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Magoules:2017:JAC

[1574] Frédéric Magoulès and Guillaume Gbikpi-Benissan. JACK: an asynchronous communication kernel library for iterative algorithms. The Journal of Supercomputing, 73(8):3468–3487, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Peng:2017:RAC

[1575] Junjie Peng, Jinbao Chen, Xiaofei Zhi, Meikang Qiu, and Xiaolan Xie. Research on application classification method in cloud computing environment. The Journal of Supercomputing, 73(8):3488–3507, August 2017.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2017:SWB

[1576] Ming Chen, John Ye, Tianzhou Chen, and Hongjun Dai. Shared write buffer to boost applications on SpMT architecture. The Journal of Supercomputing, 73(8):3508–3525, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cai:2017:SAE

[1577] Xiaojun Cai, Feng Li, Ping Li, Lei Ju, and Zhiping Jia. SLA-aware energyefficient scheduling scheme for Hadoop YARN. The Journal of Supercomputing, 73(8):3526–3546, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2017:SHQ

[1578] Longbin Chen, Meikang Qiu, Wenyun Dai, and Ning Jiang. Supporting highquality video streaming with SDNbased CDNs. The Journal of Supercomputing, 73(8):3547–3561, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lu:2017:VES

[1579] Xiang Lu, Linlin Tu, Xinyun Zhou, Naixue Xiong, and Limin Sun. ViMediaNet: an emulation system for interactive multimedia based telepresence services. The Journal of Supercomputing, 73(8):3562–3578, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Safkhani:2017:PSD

[1580] Masoumeh Safkhani and Nasour Bagheri. Passive secret disclosure attack on an ultralightweight authentication protocol for Internet of Things. The Journal of Supercomputing, 73(8): 3579–3585, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mostafaeipour:2017:IFR

[1581] Ali Mostafaeipour, Mojtaba Qolipour, and Hamidreza Eslami. Implementing fuzzy rank function model for a new supply chain risk management. The Journal of Supercomputing, 73(8): 3586–3602, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Carcenac:2017:GPS

[1582] Manuel Carcenac, Soydan Redif, and Server Kasap. GPU parallelization of the sequential matrix diagonalization algorithm and its application to high-dimensional data. The Journal of Supercomputing, 73(8):3603–3634, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Queudet:2017:KRT

[1583] Audrey Queudet, Nadine Abdallah, and Maryline Chetto. KTS: a realtime mapping algorithm for NoC-based many-cores. The Journal of Supercomputing, 73(8):3635–3651, August 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sethi:2017:HSB

[1584] Krishan Kumar Sethi and Dharavath Ramesh. HFIM: a Spark-based hybrid frequent itemset mining algorithm for big data processing. The Journal of Supercomputing, 73(8):3652– 3668, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cheng:2017:RCA

[1585] Eddie Cheng, Ke Qiu, and Zhizhang Shen. On the restricted connectivity of the arrangement graph. The Journal of Supercomputing, 73(8):3669– 3682, August 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chunlin:2017:ERP

[1586] Li Chunlin, Zhou Min, and Luo Youlong. Elastic resource provisioning in hybrid mobile cloud for computationally intensive mobile applications. The Journal of Supercomputing, 73(9): 3683–3714, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2017:MOP

[1587] Shuai Zhang, Zhao Wang, Ying Peng, Bertil Schmidt, and Weiguo Liu. Mapping of option pricing algorithms onto heterogeneous many-core architectures. The Journal of Supercomputing, 73(9): 3715–3737, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moreno:2017:HLB

[1588] Andreu Moreno, Anna Sikora, Eduardo César, Joan Sorribes, and Tomàs Margalef. HeDPM: load balancing of linear pipeline applications on heterogeneous systems. The Journal of Supercomputing, 73(9):3738–3760, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/content/pdf/10.1007/s11227-
                                        017-1971-4.pdf.

Al-Mouhamed:2017:SBS

[1589] Mayez A. Al-Mouhamed and Ayaz H. Khan. SpMV and BiCG–Stab optimization for a class of hepta-diagonal-sparse matrices on GPU. The Journal of Supercomputing, 73(9):3761– 3795, September 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rashidi:2017:CDS

[1590] Shima Rashidi and Saeed Sharifian. Cloudlet dynamic server selection policy for mobile task off-loading in mobile cloud computing using soft computing techniques. The Journal of Supercomputing, 73(9):3796–3820, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Keshavarz-Kohjerdi:2017:LTA

[1591] Fatemeh Keshavarz-Kohjerdi and Alireza Bagheri. A linear-time algorithm for finding Hamiltonian (s, t)- paths in odd-sized rectangular grid graphs with a rectangular hole. The Journal of Supercomputing, 73(9): 3821–3860, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Neamatollahi:2017:STB

[1592] Peyman Neamatollahi, Yasser Sedaghat, and Mahmoud Naghibzadeh. A simple token-based algorithm for the mutual exclusion problem in distributed systems. The Journal of Supercomputing, 73(9):3861–3878, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vega:2017:LTL

[1593] Carlos Vega, Paula Roquero, Rafael Leira, Ivan Gonzalez, and Javier Aracil. Loginson: a transform and load system for very large-scale log analysis in large IT infrastructures. The Journal of Supercomputing, 73(9): 3879–3900, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmad:2017:ODI

[1594] Saima Gulzar Ahmad, Chee Sun Liew, M. Mustafa Rafique, and Ehsan Ullah Munir. Optimization of data-intensive workflows in stream-based data processing models. The Journal of Supercomputing, 73(9):3901–3923, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Casas:2017:PDS

[1595] Israel Casas, Javid Taheri, Rajiv Ranjan, and Albert Y. Zomaya. PSO-DS: a scheduling engine for scientific workflow managers. The Journal of Supercomputing, 73(9):3924–3947, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Egawa:2017:PMV

[1596] Ryusuke Egawa, Kazuhiko Komatsu, Shintaro Momose, Yoko Isobe, Akihiro Musa, Hiroyuki Takizawa, and Hiroaki Kobayashi. Potential of a modern vector supercomputer for practical applications: performance evaluation of SX–ACE. The Journal of Supercomputing, 73(9):3948–3976, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        017-1993-y.pdf.

Vasudevan:2017:PBD

[1597] Meera Vasudevan, Yu-Chu Tian, Maolin Tang, Erhan Kozan, and Weizhe Zhang. Profile-based dynamic application assignment with a repairing genetic algorithm for greener data centers. The Journal of Supercomputing, 73(9):3977–3998, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Luo:2017:DPR

[1598] Peng Luo, Deqing Zou, Hai Jin, Yajuan Du, and Jinan Shen. A dynamic predictive race detector for C/C++ programs. The Journal of Supercomputing, 73(9):3999–4019, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Youn:2017:CCB

[1599] Young-Sun Youn, Su-Kyung Yoon, and Shin-Dug Kim. Cloud computing burst system (CCBS): for exa-scale computing system. The Journal of Supercomputing, 73(9):4020–4041, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ziafat:2017:MOS

[1600] Hassan Ziafat and Seyed Morteza Babamir. A method for the optimum selection of datacenters in geographically distributed clouds. The Journal of Supercomputing, 73(9):4042– 4081, September 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2017:SPD

[1601] Zhoukai Wang, Yinliang Zhao, Yang Liu, Zhong Chen, Cuocuo Lv, and Yuxiang Li. A speculative parallel decompression algorithm on Apache Spark. The Journal of Supercomputing, 73(9): 4082–4111, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Farshin:2017:MSM

[1602] Alireza Farshin and Saeed Sharifian.

MAP-SDN: a metaheuristic assignment and provisioning SDN framework for cloud datacenters. The Journal of Supercomputing, 73(9):4112– 4136, September 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hu:2017:EGC

[1603] Sensen Hu, Feng Shi, Weixing Ji, Xu Chen, and Shahnawaz Talpur. Exploring grouped coherence for clustered hierarchical cache. The Journal of Supercomputing, 73(9):4137– 4157, September 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Reza:2017:LCP

[1604] Akram Reza and Mahnaz Rafie. Limited contiguous processor allocation mechanism in the mesh-connected multiprocessors using compaction. The Journal of Supercomputing, 73(9): 4158–4189, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Siddique:2017:IAH

[1605] Kamran Siddique, Zahid Akhtar, Yangwoo Kim, Young-Sik Jeong, and Edward J. Yoon. Investigating Apache Hama: a bulk synchronous parallel computing framework. The Journal of Supercomputing, 73(9):4190– 4205, September 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2017:DEM

[1606] Wenbing Zhao and Mary Q. Yang. Dependability enhancing mechanisms for integrated clinical environments. The Journal of Supercomputing, 73(10):4207–4220, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s11227-017-

2003-0.pdf. See correction [1802].

AlBdaiwi:2017:GBG

[1607] Bader F. AlBdaiwi and Hosam M. F. AboElFotoh. A GPU-based genetic algorithm for the p-median problem. The Journal of Supercomputing, 73(10): 4221–4244, October 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wan:2017:DBS

[1608] Jianxiong Wan, Xiang Gui, and Ran Zhang. Dynamic bidding in spot market for profit maximization in the public cloud. The Journal of Supercomputing, 73(10):4245–4274, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anada:2017:CGS

[1609] Hiroaki Anada, Junpei Kawamoto, Chenyutao Ke, Kirill Morozov, and Kouichi Sakurai. Cross-group secret sharing scheme for secure usage of cloud storage over different providers and regions. The Journal of Supercomputing, 73(10):4275–4301, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mihaescu:2017:DAS

[1610] Marian Cristian Mihaescu, Paul Stefan Popescu, and Elvira Popescu. Data analysis on social media traces for detection of “spam” and “don’t care” learners. The Journal of Supercomputing, 73(10):4302–4323, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pan:2017:EAC

[1611] Shaoming Pan, Yanwen Chong, Zhengquan Xu, and Xicheng Tan. An enhanced active caching strategy for data-intensive computations in distributed GIS. The Journal of Supercomputing, 73(10):4324–4346, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-2012-z.pdf.

Fard:2017:DVC

[1612] Seyed Yahya Zahedi Fard, Mohamad Reza Ahmadi, and Sahar Adabi. A dynamic VM consolidation technique for QoS and energy consumption in cloud environment. The Journal of Supercomputing, 73(10):4347– 4368, October 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [1613].

Fard:2017:EDV

[1613] Seyed Yahya Zahedi Fard, Mohamad Reza Ahmadi, and Sahar Adabi. Erratum to: A dynamic VM consolidation technique for QoS and energy consumption in cloud environment. The Journal of Supercomputing, 73(10):4369–4372, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-2079-6.pdf. See [1612].

Rojek:2017:MPC

[1614] Krzysztof Rojek, Enrique S. Quintana-Ortí, and Roman Wyrzykowski. Modeling power consumption of 3D MP-DATA and the CG method on ARM and Intel multicore architectures. The Journal of Supercomputing, 73(10): 4373–4389, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-017-2020-z.pdf.

Hofinger:2017:MPO

[1615] Siegfried Höfinger and Ernst Haunschmid. Modelling parallel overhead from simple run-time records. The Journal of Supercomputing, 73(10):4390– 4406, October 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-017-2023-9.pdf.

Nezarat:2017:GTB

[1616] Amin Nezarat and Yaser Shams. A game theoretic-based distributed detection method for VM-to-hypervisor attacks in cloud environment. The Journal of Supercomputing, 73(10): 4407–4427, October 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fernandez-Pascual:2017:SIE

[1617] Ricardo Fernández-Pascual, Alberto Ros, and Manuel E. Acacio. To be silent or not: on the impact of evictions of clean data in cache-coherent multicores. The Journal of Supercomputing, 73(10):4428–4443, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Elghamrawy:2017:PFC

[1618] Sally M. Elghamrawy and Aboul Ella Hassanien. A partitioning framework for Cassandra NoSQL database using Rendezvous hashing. The Journal of Supercomputing, 73(10):4444– 4465, October 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmed:2017:FAS

[1619] Ejaz Ahmed, Anjum Naveed, Siti Hafizah Ab Hamid, Abdullah Gani, and Khaled Salah. Formal analysis of seamless application execution in mobile cloud computing. The Journal of Supercomputing, 73(10):4466– 4492, October 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zeng:2017:ITS

[1620] Guosun Zeng and Wenjuan Liu. An iso-time scaling method for big data tasks executing on parallel computing systems. The Journal of Supercomputing, 73(10):4493–4516, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhou:2017:PDS

[1621] Wei Zhou, Zhanxiu Cai, Bo Lian, Jincai Wang, and Jianping Ma. Protein database search of hybrid alignment algorithm based on GPU parallel acceleration. The Journal of Supercomputing, 73(10):4517–4534, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Somu:2017:RSB

[1622] Nivethitha Somu, Kannan Kirthivasan, and V. S. Shankar Sriram. A rough setbased hypergraph trust measure parameter selection technique for cloud service selection. The Journal of Supercomputing, 73(10):4535–4559, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Habibian:2017:FTR

[1623] Hossein Habibian and Ahmad Patooghy. Fault-tolerant routing methodology for hypercube and cubeconnected cycles interconnection networks. The Journal of Supercomputing, 73(10):4560–4579, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rani:2017:CBA

[1624] Sita Rani and O. P. Gupta. CLUS GPU-BLASTP: accelerated protein sequence alignment using GPU-enabled cluster. The Journal of Supercomputing, 73(10):4580–4595, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Barkhordari:2017:AMB

[1625] Mohammadhossein Barkhordari and Mahdi Niamanesh. Atrak: a MapReduce-based data warehouse for big data. The Journal of Supercomputing, 73(10):4596–4610, October 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Barrientos:2017:GBE

[1626] Ricardo J. Barrientos, Fabricio Millaguir, José L. Sánchez, and Enrique Arias. GPU-based exhaustive algorithms processing kNN queries. The Journal of Supercomputing, 73(10): 4611–4634, October 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2017:RPP

[1627] Ju-Won Park and Eunhye Kim. Runtime prediction of parallel applications with workload-aware clustering. The Journal of Supercomputing, 73(11): 4635–4651, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2017:PSC

[1628] Rakesh Ranjan Kumar, Siba Mishra, and Chiranjeev Kumar. Prioritizing the solution of cloud service selection using integrated MCDM methods under fuzzy environment. The Journal of Supercomputing, 73(11):4652– 4682, November 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alemi:2017:CUS

[1629] Mehdi Alemi, Hassan Haghighi, and Saeed Shahrivari. CCFinder: using Spark to find clustering coefficient in big graphs. The Journal of Supercomputing, 73(11):4683–4710, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wei:2017:RPT

[1630] Kai-Cheng Wei, Xue Sun, Hsun Chu, and Chao-Chin Wu. Reconstructing permutation table to improve the Tabu Search for the PFSP on GPU. The Journal of Supercomputing, 73(11): 4711–4738, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

LeCompte:2017:SER

[1631] Travis LeCompte, Walker Legrand, Sui Chen, and Lu Peng. Soft error resilience of Big Data kernels through algorithmic approaches. The Journal of Supercomputing, 73(11):4739– 4772, November 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abualigah:2017:UTF

[1632] Laith Mohammad Abualigah and Ahamad Tajudin Khader. Unsupervised text feature selection technique based on hybrid particle swarm optimization algorithm with genetic operators for the text clustering. The Journal of Supercomputing, 73(11):4773– 4795, November 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hanani:2017:MPS

[1633] Ali Hanani, Amir Masoud Rahmani, and Amir Sahafi. A multi-parameter scheduling method of dynamic workloads for big data calculation in cloud computing. The Journal of Supercomputing, 73(11):4796–4822, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yam-Uicab:2017:FHT

[1634] R. Yam-Uicab, J. L. Lopez-Martinez, J. A. Trejo-Sanchez, H. Hidalgo-Silva, and S. Gonzalez-Segura. A fast Hough transform algorithm for straight lines detection in an image using GPU parallel computing with CUDA-C. The Journal of Supercomputing, 73(11): 4823–4842, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Valinataj:2017:NPP

[1635] Mojtaba Valinataj. Novel paritypreserving reversible logic array multipliers. The Journal of Supercomputing, 73(11):4843–4867, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Salehan:2017:OVB

[1636] Alireza Salehan, Hossein Deldari, and Saeid Abrishami. An online valuationbased sealed winner-bid auction game for resource allocation and pricing in clouds. The Journal of Supercomputing, 73(11):4868–4905, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2017:RTW

[1637] Huangke Chen, Jianghan Zhu, Zhenshi Zhang, Manhao Ma, and Xin Shen. Real-time workflows oriented online scheduling in uncertain cloud environment. The Journal of Supercomputing, 73(11):4906–4922, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xu:2017:SAP

[1638] Zhiyan Xu, Libing Wu, Debiao He, and Muhammad Khurram Khan. Security analysis of a publicly verifiable data possession scheme for remote storage. The Journal of Supercomputing, 73(11):4923–4930, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mandikas:2017:PMS

[1639] Vassilios G. Mandikas and Emmanuel N. Mathioudakis. A parallel multigrid solver for incompressible flows on computing architectures with accelerators. The Journal of Supercomputing, 73(11):4931–4956, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mostafaei:2017:BCW

[1640] Habib Mostafaei, Mohammad Shojafar, Bahman Zaher, and Mukesh Singhal. Barrier coverage of WSNs with the imperialist competitive algorithm. The Journal of Supercomputing, 73(11): 4957–4980, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2017:IWB

[1641] Chao-Tung Yang, Chao-Wei Huang, and Shuo-Tsung Chen. Improvement of workload balancing using parallel loop self-scheduling on Intel Xeon Phi. The Journal of Supercomputing, 73(11):4981–5005, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Al-Yatama:2017:MAA

[1642] Anwar Al-Yatama, Imtiaz Ahmad, and Naelah Al-Dabbous. Memory allocation algorithm for cloud services. The Journal of Supercomputing, 73(11): 5006–5033, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aghaei:2017:CEF

[1643] Babak Aghaei, Kambiz Badie, Ahmad Khademzadeh, and Midia Reshadi. The cost-effective fault detection and fault location approach for communication channels in NoC. The Journal of Supercomputing, 73(11):5034– 5052, November 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tirado:2017:EEX

[1644] Felipe Tirado, Ricardo J. Barrientos, Paulo González, and Marco Mora. Efficient exploitation of the Xeon Phi architecture for the Ant Colony Optimization (ACO) metaheuristic. The Journal of Supercomputing, 73(11): 5053–5070, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2017:ECF

[1645] Boseon Yu, Wonik Choi, and Ling Liu.

Exploring correlation for fast skyline computation. The Journal of Supercomputing, 73(11):5071–5102, November 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [1646].

Yu:2017:EEC

[1646] Boseon Yu, Wonik Choi, and Ling Liu. Erratum to: Exploring correlation for fast skyline computation. The Journal of Supercomputing, 73(11):5103, November 2017.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://link.springer.com/
                                        content/pdf/10.1007/s11227-017-

2086-7.pdf. See [1645].

Koupaee:2017:SSF

[1647] Mahnaz Koupaee, Mohammad Reza Kangavari, and Mohammad Javad Amiri. Scalable structure-free data fusion on wireless sensor networks. The Journal of Supercomputing, 73(12): 5105–5124, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Green:2017:CSM

[1648] Robert C. Green II and Vishakha Agrawal. A case study in multi-core parallelism for the reliability evaluation of composite power systems. The Journal of Supercomputing, 73(12):5125– 5149, December 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2017:RSA

[1649] Chunlin Li, Liye Zhu, Yanpei Liu, and Youlong Luo. Resource scheduling approach for multimedia cloud content management. The Journal of Supercomputing, 73(12):5150–5172, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jamali:2017:DDA

[1650] Shahram Jamali and Reza Fotohi. DAWA: Defending against wormhole attack in MANETs by using fuzzy logic and artificial immune system. The Journal of Supercomputing, 73(12): 5173–5196, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kyziropoulos:2017:PSM

[1651] P. E. Kyziropoulos, C. K. Filelis-Papadopoulos, G. A. Gravvanis, and C. Efthymiopoulos. A parallel self mesh-adaptive N-body method based on approximate inverses. The Journal of Supercomputing, 73(12):5197– 5220, December 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Salmanian:2017:OWS

[1652] Zolfaghar Salmanian, Habib Izadkhah, and Ayaz Isazadeh. Optimizing web server RAM performance using birthdeath process queuing system: scalable memory issue. The Journal of Supercomputing, 73(12):5221–5238, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shojafar:2017:FBD

[1653] Mohammad Shojafar, Zahra Pooranian, Paola G. Vinueza Naranjo, and Enzo Baccarelli. FLAPS: bandwidth and delay-efficient distributed data searching in fog-supported P2P content delivery networks. The Journal of Supercomputing, 73(12):5239– 5260, December 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

ElKafhali:2017:EDS

[1654] Said El Kafhali and Khaled Salah.

Efficient and dynamic scaling of fog nodes for IoT devices. The Journal of Supercomputing, 73(12):5261– 5284, December 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xu:2017:SEP

[1655] Zhiyan Xu, Libing Wu, Muhammad Khurram Khan, Kim-Kwang Raymond Choo, and Debiao He. A secure and efficient public auditing scheme using RSA algorithm for cloud storage. The Journal of Supercomputing, 73(12):5285–5309, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Eslami:2017:NSP

[1656] Ghazaleh Eslami and Abolfazl Toroghi Haghighat. A new surrogate placement algorithm for cloud-based content delivery networks. The Journal of Supercomputing, 73(12):5310– 5331, December 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Darabkh:2017:CDC

[1657] Khalid A. Darabkh, Wala’a S. Al-Rawashdeh, Raed T. Al-Zubi, and Sharhabeel H. Alnabelsi. C-DTB-CHR: centralized density- and threshold-based cluster head replacement protocols for wireless sensor networks. The Journal of Supercomputing, 73(12):5332–5353, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abdeyazdan:2017:NMI

[1658] Marjan Abdeyazdan. A new method for the informed discovery of resources in the grid system using particle swarm optimization algorithm (RDT PSO). The Journal of Supercomputing, 73 (12):5354–5377, December 2017. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jarzabek:2017:PEU

[1659] Lukasz Jarzabek and Pawel Czarnul. Performance evaluation of unified memory and dynamic parallelism for selected parallel CUDA applications. The Journal of Supercomputing, 73(12):5378–5401, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-2091-x.pdf.

Alrashed:2017:RPC

[1660] Saleh Alrashed. Reducing power consumption of non-preemptive real-time systems. The Journal of Supercomputing, 73(12):5402–5413, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2017:EBM

[1661] Licheng Yu, Yulong Pei, Tianzhou Chen, Xueqing Lou, Minghui Wu, and Tiefei Zhang. Enable back memory and global synchronization on LLC buffer. The Journal of Supercomputing, 73(12):5414–5439, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2017:GBW

[1662] Madhu Sudan Kumar, Indrajeet Gupta, Sanjaya K. Panda, and Prasanta K. Jana. Granularity-based workflow scheduling algorithm for cloud computing. The Journal of Supercomputing, 73(12):5440–5464, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Llopis:2017:MBE

[1663] Pablo Llopis, Florin Isaila, Javier Garcia Blas, and Jesus Carretero. Modelbased energy-aware data movement optimization in the storage I/O stack. The Journal of Supercomputing, 73 (12):5465–5495, December 2017. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vahidifar:2017:LAR

[1664] Samira Vahidifar and Midia Reshadi. Loss-aware routing algorithm for photonic networks on chip. The Journal of Supercomputing, 73(12):5496– 5514, December 2017. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [1665].

Vahidifar:2017:ELA

[1665] Samira Vahidifar and Midia Reshadi. Erratum to: Loss-aware routing algorithm for photonic networks on chip. The Journal of Supercomputing, 73(12):5515, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        017-2108-5.pdf. See [1664].

Sait:2017:OMD

[1666] Sadiq M. Sait and Kh. Shahzada Shahid. Optimal multi-dimensional vector bin packing using simulated evolution. The Journal of Supercomputing, 73(12):5516–5538, December 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Al-Adwan:2018:STS

[1667] Aryaf Al-Adwan, Basel A. Mahafzah, and Ahmad Sharieh. Solving traveling salesman problem using parallel repetitive nearest neighbor algorithm on OTIS–Hypercube and OTIS– Mesh optoelectronic architectures. The Journal of Supercomputing, 74(1):1– 36, January 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zigon:2018:ISF

[1668] Bob Zigon, Luoding Zhu, and Fengguang Song. Interactive 3D simulation for fluid-structure interactions using dual coupled GPUs. The Journal of Supercomputing, 74(1):37–64, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:SNU

[1669] King-Hang Wang, Chien-Ming Chen, Weicheng Fang, and Tsu-Yang Wu. On the security of a new ultra-lightweight authentication protocol in IoT environment for RFID tags. The Journal of Supercomputing, 74(1):65–70, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rostampour:2018:SLG

[1670] Samad Rostampour, Nasour Bagheri, Mehdi Hosseinzadeh, and Ahmad Khademzadeh. A scalable and lightweight grouping proof protocol for Internet of Things applications. The Journal of Supercomputing, 74(1): 71–86, January 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Salami:2018:GQM

[1671] Esther Salamí, Cristina Barrado, Antonia Gallardo, and Enric Pastor. General queuing model for optimal seamless delivery of payload processing in multi-core processors. The Journal of Supercomputing, 74(1):87–104, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-2109-4.pdf.

Piao:2018:RSA

[1672] Zhegao Piao, Seong Joon Yoo, Yeong Hyeon Gu, Jaechun No, Zhiyan Jiang, and Helin Yin. Recommender system architecture based on Mahout and a main memory database. The Journal of Supercomputing, 74(1):105– 121, January 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gupta:2018:RAV

[1673] Madnesh K. Gupta and Tarachand Amgoth. Resource-aware virtual machine placement algorithm for IaaS cloud. The Journal of Supercomputing, 74(1):122–140, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:NBM

[1674] Yi Li Wang, Kyung Tae Kim, Byungjun Lee, and Hee Yong Youn. A novel buffer management scheme based on particle swarm optimization for SSD. The Journal of Supercomputing, 74(1):141–159, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmadzadeh:2018:HPE

[1675] Armin Ahmadzadeh, Omid Hajihassani, and Saeid Gorgin. A highperformance and energy-efficient exhaustive key search approach via GPU on DES-like cryptosystems. The Journal of Supercomputing, 74(1):160– 182, January 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Meng:2018:CAD

[1676] Xianfu Meng. A churn-aware durable data storage scheme in hybrid P2P networks. The Journal of Supercomputing, 74(1):183–204, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Niari:2018:EEC

[1677] Ali Kazemi Niari, Reza Berangi, and Mahmood Fathy. ECCN: an extended CCN architecture to improve data access in vehicular content-centric network. The Journal of Supercomputing, 74(1):205–221, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nguyen:2018:CEA

[1678] Tuan Anh Nguyen, Dugki Min, and Eunmi Choi. A comprehensive evaluation of availability and operational cost for a virtualized server system using stochastic reward nets. The Journal of Supercomputing, 74(1):222–276, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [1777].

Rostami:2018:SCH

[1679] Ali Shokouhi Rostami, Marzieh Badkoobe, Farahnaz Mohanna, Hengameh keshavarz, Ali Asghar Rahmani Hosseinabadi, and Arun Kumar Sangaiah. Survey on clustering in heterogeneous and homogeneous wireless sensor networks. The Journal of Supercomputing, 74(1):277–323, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bahig:2018:FOP

[1680] Hazem M. Bahig. A fast optimal parallel algorithm for a short addition chain. The Journal of Supercomputing, 74(1): 324–333, January 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Popoola:2018:ECS

[1681] Olusogo Popoola and Bernardi Pranggono. On energy consumption of switch-centric data center networks. The Journal of Supercomputing, 74 (1):334–369, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-017-2132-5.pdf.

Mishra:2018:ATA

[1682] Sambit Kumar Mishra, Deepak Puthal, Bibhudatta Sahoo, Sajay Kumar Jena, and Mohammad S. Obaidat. An adaptive task allocation technique for green cloud computing. The Journal of Supercomputing, 74(1):370–385, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Randhawa:2018:EEL

[1683] Sukhchandan Randhawa and Sushma Jain. Energy-efficient load balancing scheme for two-tier communication in wireless sensor networks. The Journal of Supercomputing, 74(1):386– 416, January 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fareghzadeh:2018:DPI

[1684] Nafiseh Fareghzadeh, Mir Ali Seyyedi, and Mehran Mohsenzadeh. Dynamic performance isolation management for cloud computing services. The Journal of Supercomputing, 74(1):417–455, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:GBR

[1685] Dae Yeol Lee, Jooyoung Lee, Ji-Hoon Choi, Jong-Ok Kim, Hui Yong Kim, and Jin Soo Choi. GPU-based realtime super-resolution system for highquality UHD video up-conversion. The Journal of Supercomputing, 74(1):456– 484, January 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2018:ECS

[1686] Keqin Li. Energy constrained scheduling of stochastic tasks. The Journal of Supercomputing, 74(1):485–508, January 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aghili:2018:ISA

[1687] Seyed Farhad Aghili, Maede Ashouri-Talouki, and Hamid Mala. DoS, impersonation and de-synchronization attacks against an ultra-lightweight RFID mutual authentication protocol for IoT. The Journal of Supercomputing, 74(1):509–525, January 2018.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2018:FAH

[1688] James J. Park. Fusion algorithms and high-performance applications for vehicular cloud computing. The Journal of Supercomputing, 74(3):995–1000, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2270-4.pdf.

Xu:2018:SKS

[1689] Li Xu, Chi-Yao Weng, Lun-Pin Yuan, Mu-En Wu, Raylin Tso, and Hung-Min Sun. A shareable keyword search over encrypted data in cloud computing. The Journal of Supercomputing, 74(3): 1001–1023, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nkenyereye:2018:SVT

[1690] Lewis Nkenyereye, Youngho Park, and Kyung-Hyune Rhee. Secure vehicle traffic data dissemination and analysis protocol in vehicular cloud computing. The Journal of Supercomputing, 74(3): 1024–1044, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2018:EES

[1691] Longbin Chen, Meikang Qiu, Jeungeun Song, Zenggang Xiong, and Houcine Hassan. E2FS: an elastic storage system for cloud computing. The Journal of Supercomputing, 74(3):1045– 1060, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Munoz:2018:AFR

[1692] Fernando Román Muñoz and Luis Javier García Villalba. An algorithm to find relationships between web vulnerabilities. The Journal of Supercomputing, 74(3):1061–1089, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lim:2018:EDM

[1693] JongBeom Lim, Young Sik Jeong, Doo-Soon Park, and HwaMin Lee. An efficient distributed mutual exclusion algorithm for intersection traffic control. The Journal of Supercomputing, 74(3): 1090–1107, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ji:2018:CSD

[1694] Sai Ji, Chang Tan, Ping Yang, Ya-Jie Sun, Desheng Fu, and Jin Wang. Compressive sampling and data fusionbased structural damage monitoring in wireless sensor network. The Journal of Supercomputing, 74(3):1108– 1131, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2018:APM

[1695] You-Shyang Chen, Chienwen Wu, Heng-Hsing Chu, Chien-Ku Lin, and Huan-Ming Chuang. Analysis of performance measures in cloud-based ubiquitous SaaS CRM project systems. The Journal of Supercomputing, 74(3): 1132–1156, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ramadoss:2018:NIT

[1696] Ravikumar Ramadoss, N. M. Elango, Satheesh Abimannan, and Ching-Hsien Hsu. Non-intrusive transaction aware filtering during enterprise application modernization. The Journal of Supercomputing, 74(3):1157–1181, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:RDI

[1697] Keon Myung Lee, Kyung Mi Lee, and Sang Ho Lee. Remote data integrity check for remotely acquired and stored stream data. The Journal of Supercomputing, 74(3):1182–1201, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hao:2018:CNF

[1698] Fei Hao and Doo-Soon Park. cSketch: a novel framework for capturing cliques from big graph. The Journal of Supercomputing, 74(3):1202–1214, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:LBS

[1699] Hye-Young Kim. A load balancing scheme with Loadbot in IoT networks. The Journal of Supercomputing, 74(3): 1215–1226, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Min:2018:CLD

[1700] Seokhong Min, Youngsik Jeong, and Jungho Kang. Cross-layer design and performance analysis for maximizing the network utilization of wireless mesh networks in cloud computing. The Journal of Supercomputing, 74(3):

1227–1254, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:MEE

[1701] Won-Hyuk Lee, Jong-Seon Park, Seung-Hae Kim, Jin-Hyung Park, and Joon-Min Gil. A method for enhancing end-to-end transfer efficiency via performance tuning factors on dedicated circuit networks with a public cloud platform. The Journal of Supercomputing, 74(3):1255–1266, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tan:2018:EDF

[1702] Chang Tan, Sai Ji, Ziyuan Gui, Jian Shen, De-Sheng Fu, and Jin Wang. An effective data fusion-based routing algorithm with time synchronization support for vehicular wireless sensor networks. The Journal of Supercomputing, 74(3):1267–1282, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2018:ABT

[1703] Shin Hyoung Park, Jienki Synn, Oh Hoon Kwon, and Yunsick Sung. Apriori-based text mining method for the advancement of the transportation management plan in expressway work zones. The Journal of Supercomputing, 74(3):1283–1298, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yazdinejad:2018:EDH

[1704] Abbas Yazdinejad, Ali Bohlooli, and Kamal Jamshidi. Efficient design and hardware implementation of the Open-Flow v1.3 Switch on the Virtex-6 FPGA ML605. The Journal of Supercomputing, 74(3):1299–1320, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khaleghzadeh:2018:HMT

[1705] Hamidreza Khaleghzadeh, Hossein Deldari, Ravi Reddy, and Alexey Lastovetsky. Hierarchical multicore thread mapping via estimation of remote communication. The Journal of Supercomputing, 74(3):1321–1340, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dolbeau:2018:TPF

[1706] Romain Dolbeau. Theoretical peak FLOPS per instruction set: a tutorial. The Journal of Supercomputing, 74(3): 1341–1377, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [?].

Chung:2018:AMD

[1707] Manhyun Chung, Woogeun Ahn, Byunggil Min, Jungtaek Seo, and Jongsub Moon. An analytical method for developing appropriate protection profiles of Instrumentation & Control System for nuclear power plants. The Journal of Supercomputing, 74(3): 1378–1393, March 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kwon:2018:EIP

[1708] Jihoon Kwon, Seog Chung Seo, and Seokhie Hong. An efficient implementation of pairing-based cryptography on MSP430 processor. The Journal of Supercomputing, 74(3):1394–1417, March 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [1751].

Marowka:2018:SSP

[1709] Ami Marowka and Przemysław Stpiczyński.

Special section on parallel programming. The Journal of Supercomputing, 74(4):1419–1421, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-018-2278-9.pdf.

Thoman:2018:TTB

[1710] Peter Thoman, Kiril Dichev, Thomas Heller, Roman Iakymchuk, Xavier Aguilar, Khalid Hasanov, Philipp Gschwandtner, Pierre Lemarinier, Stefano Markidis, Herbert Jordan, Thomas Fahringer, Kostas Katrinis, Erwin Laure, and Dimitrios S. Nikolopoulos. A taxonomy of taskbased parallel programming technologies for high-performance computing. The Journal of Supercomputing, 74 (4):1422–1434, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-018-2238-4.pdf.

Posner:2018:HWS

[1711] Jonas Posner and Claudia Fohry. Hybrid work stealing of locality-flexible and cancelable tasks for the APGAS library. The Journal of Supercomputing, 74(4):1435–1448, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Marowka:2018:PAH

[1712] Ami Marowka. Python accelerators for high-performance computing.

The Journal of Supercomputing, 74(4): 1449–1460, April 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Stpiczynski:2018:LBV

[1713] Przemysław Stpiczyński. Languagebased vectorization and parallelization using intrinsics, OpenMP, TBB and Cilk Plus. The Journal of Supercomputing, 74(4):1461–1472, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        017-2231-3.pdf.

Blocker:2018:PPC

[1714] Christopher Blöcker and Ulrich Hoffmann. Pardis: a process calculus for parallel and distributed programming in Haskell. The Journal of Supercomputing, 74(4):1473–1484, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Batko:2018:AMA

[1715] Paweł Batko and Marcin Kuta. Actor model of Anemone functional language. The Journal of Supercomputing, 74(4):1485–1496, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-017-2233-1.pdf.

Bylina:2018:PSB

[1716] Jarosław Bylina. Parallelization of stochastic bounds for Markov chains on multicore and manycore platforms. The Journal of Supercomputing, 74 (4):1497–1509, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-018-2235-7.pdf.

Tokura:2018:AOC

[1717] Hiroki Tokura, Toru Fujita, Koji Nakano, Yasuaki Ito, and Jacir L. Bordim. Almost optimal column-wise prefix-sum computation on the GPU. The Journal of Supercomputing, 74(4): 1510–1521, April 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Halver:2018:FPM

[1718] Rene Halver, Wilhelm Homberg, and Godehard Sutmann. Function portability of molecular dynamics on heterogeneous parallel architectures with OpenCL. The Journal of Supercomputing, 74(4):1522–1533, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Szustak:2018:SDF

[1719] Lukasz Szustak. Strategy for dataflow synchronizations in stencil parallel computations on multi-/manycore systems. The Journal of Supercomputing, 74(4):1534–1546, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-018-2239-3.pdf.

Nasri:2018:EES

[1720] Adnan Nasri, Mahmood Fathy, and Ali Broumandnia. An energy-efficient 3D-stacked STT–RAM cache architecture for cloud processors: the effect on emerging scale-out workloads. The Journal of Supercomputing, 74(4):

1547–1561, April 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sinha:2018:HMB

[1721] Ankita Sinha and Prasanta K. Jana. A hybrid MapReduce-based k-means clustering using genetic algorithm for distributed datasets. The Journal of Supercomputing, 74(4):1562–1579, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tabik:2018:TAI

[1722] S. Tabik, M. Peemen, and L. F. Romero. A tuning approach for iterative multiple 3d stencil pipeline on GPUs: Anisotropic Nonlinear Diffusion algorithm as case study. The Journal of Supercomputing, 74(4):1580–1608, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Atoofian:2018:DTS

[1723] Ehsan Atoofian and Sean Rea. Datatype specific cache compression in GPGPUs. The Journal of Supercomputing, 74(4):1609–1635, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Seo:2018:HPN

[1724] Jung-Hyun Seo, Jong-Seok Kim, Hyung Jae Chang, and Hyeong-Ok Lee. The hierarchical Petersen network: a new interconnection network with fixed degree. The Journal of Supercomputing, 74(4):1636–1654, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2018:KDF

[1725] Jie Zhao and Rongcai Zhao. K-DT:

a formal system for the evaluation of linear data dependence testing techniques. The Journal of Supercomputing, 74(4):1655–1675, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Quislant:2018:PTL

[1726] Ricardo Quislant, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Privatizing transactions for Lee’s algorithm in commercial hardware transactional memory. The Journal of Supercomputing, 74(4):1676–1694, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Braeken:2018:EAA

[1727] An Braeken, Pardeep Kumar, Madhusanka Liyanage, and Ta Thi Kim Hue. An efficient anonymous authentication protocol in multiple server communication networks (EAAM). The Journal of Supercomputing, 74(4):1695–1714, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khaneghah:2018:MMC

[1728] Ehsan Mousavi Khaneghah, Nosratollah Shadnoush, and Amir Hossein Ghobakhlou. A mathematical model to calculate real cost/performance in software distributed shared memory on computing environments. The Journal of Supercomputing, 74(4):1715–1764, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bernabe:2018:PIF

[1729] Gregorio Bernabé, Raúl Hernández, and Manuel E. Acacio. Parallel implementations of the 3D fast wavelet transform on a Raspberry Pi 2 cluster. The Journal of Supercomputing, 74(4): 1765–1778, April 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2018:NTM

[1730] Wei Zhang, Shiwei Zhu, Jian Tang, and Naixue Xiong. A novel trust management scheme based on Dempster– Shafer evidence theory for malicious nodes detection in wireless sensor networks. The Journal of Supercomputing, 74(4):1779–1801, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cecilia:2018:EMP

[1731] José M. Cecilia, José-Matías Cutillas-Lozano, Domingo Giménez, and Baldomero Imbernón. Exploiting multilevel parallelism on a many-core system for the application of hyperheuristics to a molecular docking problem. The Journal of Supercomputing, 74(5): 1803–1814, May 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ho:2018:NAF

[1732] ThienLuan Ho, Seung-Rohk Oh, and HyunJin Kim. New algorithms for fixed-length approximate string matching and approximate circular string matching under the Hamming distance. The Journal of Supercomputing, 74(5): 1815–1834, May 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [1733].

Ho:2018:CNA

[1733] ThienLuan Ho, Seung-Rohk Oh, and HyunJin Kim. Correction to: New algorithms for fixed-length approximate string matching and approximate circular string matching under the Hamming distance. The Journal of Supercomputing, 74(5):1835, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2324-7.pdf. See [1732].

Mompean:2018:DAH

[1734] Juan Mompeán, Juan L. Aragón, Pedro M. Prieto, and Pablo Artal. Design of an accurate and high-speed binocular pupil tracking system based on GPGPUs. The Journal of Supercomputing, 74(5):1836–1862, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-2193-5.pdf.

Fousek:2018:ESM

[1735] Jan Fousek. Efficient sparse matrixdelayed vector multiplication for discretized neural field model. The Journal of Supercomputing, 74(5):1863– 1884, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:PBC

[1736] Minki Kim and Haengrae Cho.

Popularity-based covering sets for energy proportionality in shared-nothing clusters. The Journal of Supercomputing, 74(5):1885–1910, May 2018.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Forsell:2018:RMM

[1737] Martti Forsell, Jussi Roivainen, and Ville Leppänen. REPLICA MB-TAC: multithreaded dual-mode processor. The Journal of Supercomputing, 74 (5):1911–1933, May 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bagheri:2018:ILC

[1738] Nasour Bagheri, Masoumeh Safkhani, Mojtaba Eslamnezhad Namin, and Samad Rostampour. An improved low-cost yoking proof protocol based on Kazahaya’s flaws. The Journal of Supercomputing, 74(5):1934–1948, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Idrees:2018:MDL

[1739] Ali Kadhum Idrees, Karine Deschinkel, Michel Salomon, and Raphaël Couturier. Multiround distributed lifetime coverage optimization protocol in wireless sensor networks. The Journal of Supercomputing, 74(5):1949–1972, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vincent:2018:KAA

[1740] O. R. Vincent and O. M. Lawal. A key agreement authentication protocol using an improved parallel Pollard rho for electronic payment system. The Journal of Supercomputing, 74(5): 1973–1993, May 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Heikalabad:2018:FAS

[1741] Saeed Rasouli Heikalabad, Mazaher Naji Asfestani, and Mehdi Hosseinzadeh. A full adder structure without cross-wiring in quantum-dot cellular automata with energy dissipation analysis. The Journal of Supercomputing, 74(5):1994–2005, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gil-Costa:2018:GBC

[1742] Veronica Gil-Costa, Mauricio Marin, Carolina Bonacic, and Roberto Solar. A graph-based cache for large-scale similarity search engines. The Journal of Supercomputing, 74(5):2006–2034, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kaur:2018:DDT

[1743] Ravneet Kaur, Inderveer Chana, and Jhilik Bhattacharya. Data deduplication techniques for efficient cloud storage management: a systematic review. The Journal of Supercomputing, 74(5): 2035–2085, May 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tran:2018:SGP

[1744] Ha-Nguyen Tran and Erik Cambria. A survey of graph processing on graphics processing units. The Journal of Supercomputing, 74(5):2086–2115, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vakilinia:2018:PCR

[1745] Shahin Vakilinia and Mohamed Cheriet.

Preemptive cloud resource allocation modeling of processing jobs. The Journal of Supercomputing, 74(5):2116– 2150, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rahnama:2018:LBH

[1746] Amir Hossein Akhavan Rahnama, Mehdi Toloo, and Nezer Jacob Zaidenberg. An LP-based hyperparameter optimization model for language modeling. The Journal of Supercomputing, 74 (5):2151–2160, May 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Scarpiniti:2018:EPH

[1747] Michele Scarpiniti, Enzo Baccarelli, Paola G. Vinueza Naranjo, and Aurelio Uncini. Energy performance of heuristics and meta-heuristics for real-time joint resource scaling and consolidation in virtualized networked data centers. The Journal of Supercomputing, 74(5): 2161–2198, May 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tamizharasan:2018:ALD

[1748] P. S. Tamizharasan and N. Ramasubramanian. Analysis of large deviations behavior of multi-GPU memory access in deep learning. The Journal of Supercomputing, 74(5):2199–2212, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abdi:2018:HHS

[1749] Athena Abdi and Hamid R. Zarandi. HYSTERY: a hybrid scheduling and mapping approach to optimize temperature, energy consumption and lifetime reliability of heterogeneous multiprocessor systems. The Journal of Supercomputing, 74(5):2213–2238, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2018:EPB

[1750] SungHoon Park, SuChang Yoo, and BoKyoung Kim. An election protocol based on group membership detection algorithm in mobile ad hoc distributed systems. The Journal of Supercomputing, 74(5):2239–2253, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kwon:2018:CEI

[1751] Jihoon Kwon, Seog Chung Seo, and Seokhie Hong. Correction to: An efficient implementation of pairingbased cryptography on MSP430 processor. The Journal of Supercomputing, 74(5):2254, May 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s11227-018-

2320-y.pdf. See [1708].

Aydin:2018:RTP

[1752] Semra Aydin, Refik Samet, and Omer Faruk Bay. Real-time parallel image processing applications on multicore CPUs with OpenMP and GPGPU with CUDA. The Journal of Supercomputing, 74(6):2255–2275, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Stewart:2018:IDU

[1753] Iain A. Stewart and Alejandro Erickson. The influence of datacenter usage on symmetry in datacenter network design. The Journal of Supercomputing, 74(6):2276–2313, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-017-2217-1.pdf.

Neves:2018:SDP

[1754] Nuno Neves, Pedro Tomás, and Nuno Roma. Stream data prefetcher for the GPU memory interface. The Journal of Supercomputing, 74(6):2314–2328, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Neamatollahi:2018:DUC

[1755] Peyman Neamatollahi and Mahmoud Naghibzadeh. Distributed unequal clustering algorithm in large-scale wireless sensor networks using fuzzy logic. The Journal of Supercomputing, 74(6): 2329–2352, June 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khorandi:2018:NCO

[1756] Sina Mahmoodi Khorandi and Mohsen Sharifi. Non-clairvoyant online scheduling of synchronized jobs on virtual clusters. The Journal of Supercomputing, 74(6):2353–2384, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Arastouie:2018:SAR

[1757] Narges Arastouie and Masoud Sabaei. Self-adaptive risk-aware routing in opportunistic network. The Journal of Supercomputing, 74(6):2385–2411, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mukherjee:2018:CLP

[1758] Anwesha Mukherjee, Priti Deb, Debashis De, and Rajkumar Buyya. C2OF2N: a low power cooperative code offloading method for femtolet-based fog network. The Journal of Supercomputing, 74(6):2412–2448, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Andoh:2018:TLP

[1759] Yoshimichi Andoh, Soichiro Suzuki, Satoshi Ohshima, Tatsuya Sakashita, Masao Ogino, Takahiro Katagiri, Noriyuki Yoshii, and Susumu Okazaki. A thread-level parallelization of pairwise additive potential and force calculations suitable for current manycore architectures. The Journal of Supercomputing, 74(6):2449–2469, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Naranjo:2018:DEE

[1760] Paola G. Vinueza Naranjo, Enzo Baccarelli, and Michele Scarpiniti. Design and energy-efficient resource management of virtualized networked fog architectures for the real-time support of IoT applications. The Journal of Supercomputing, 74(6):2470–2507, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:NRB

[1761] Jun Wang, Dezhi Han, and Ruijun Wang. A new rule-based poweraware job scheduler for supercomputers. The Journal of Supercomputing, 74(6):2508–2527, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bagherlou:2018:RPV

[1762] Hosein Bagherlou and Ali Ghaffari. A routing protocol for vehicular ad hoc networks using simulated annealing algorithm and neural networks. The Journal of Supercomputing, 74(6): 2528–2552, June 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Meng:2018:SSP

[1763] Xianfu Meng. speedTrust: a super peer-guaranteed trust model in hybrid P2P networks. The Journal of Supercomputing, 74(6):2553–2580, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Torabi:2018:DTS

[1764] Shadi Torabi and Faramarz Safi-Esfahani. A dynamic task scheduling framework based on chicken swarm and improved raven roosting optimization methods in cloud computing. The Journal of Supercomputing, 74(6): 2581–2626, June 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Caminal:2018:PEE

[1765] Helena Caminal, Diego Caballero, Juan M. Cebrián, Roger Ferrer, Marc Casas, Miquel Moretó, Xavier Martorell, and Mateo Valero. Performance and energy effects on task-based parallelized applications. The Journal of Supercomputing, 74(6):2627–2637, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2018:PPC

[1766] Hon-Chan Chen. The panpositionable panconnectedness of crossed cubes.

The Journal of Supercomputing, 74(6): 2638–2655, June 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Viil:2018:FAP

[1767] Jaagup Viil and Satish Narayana Srirama. Framework for automated partitioning and execution of scientific workflows in the cloud. The Journal of Supercomputing, 74(6):2656–2683, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Luo:2018:DPD

[1768] Peng Luo, Deqing Zou, Hai Jin, Yajuan Du, Long Zheng, and Jinan Shen. DigHR: precise dynamic detection of hidden races with weak causal relation analysis. The Journal of Supercomputing, 74(6):2684–2704, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

AlHasib:2018:VMA

[1769] Abdullah Al Hasib, Juan M. Cebrian, and Lasse Natvig. A vectorized k-means algorithm for compressed datasets: design and experimental analysis. The Journal of Supercomputing, 74(6):2705–2728, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tsai:2018:TSR

[1770] Hong-Bin Tsai and Chin-Laung Lei.

Time-shift replacement algorithm for main memory performance optimization. The Journal of Supercomputing, 74(6):2729–2746, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kono:2018:EOW

[1771] Fumiya Kono, Naohito Nakasato, Kensaku Hayashi, Alexander Vazhenin, and Stanislav Sedukhin. Evaluations of OpenCL-written tsunami simulation on FPGA and comparison with GPU implementation. The Journal of Supercomputing, 74(6):2747–2775, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dou:2018:FTC

[1772] Wanfeng Dou and Yanan Li. A faulttolerant computing method for Xdraw parallel algorithm. The Journal of Supercomputing, 74(6):2776–2800, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abdi:2018:CMB

[1773] Somayeh Abdi, Latif PourKarimi, Mahmood Ahmadi, and Farzad Zargari. Cost minimization for bagof-tasks workflows in a federation of clouds. The Journal of Supercomputing, 74(6):2801–2822, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Coronado-Barrientos:2018:IPI

[1774] Edoardo Coronado-Barrientos, Guillermo Indalecio, and Antonio García-Loureiro. Improving performance of iterative solvers with the AXC format using the Intel Xeon Phi. The Journal of Supercomputing, 74(6):2823–2840, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Goncalves-e-Silva:2018:PSA

[1775] Kayo Gonçalves e Silva, Daniel Aloise, and Samuel Xavier de Souza. Parallel synchronous and asynchronous coupled simulated annealing. The Journal of Supercomputing, 74(6):2841–2869, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:AAP

[1776] Sangmin Lee, Soon J. Hyun, Hong-Yeon Kim, and Young-Kyun Kim. APS: adaptable prefetching scheme to different running environments for concurrent read streams in distributed file systems. The Journal of Supercomputing, 74(6):2870–2902, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nguyen:2018:CCE

[1777] Tuan Anh Nguyen, Dugki Min, and Eunmi Choi. Correction to: A comprehensive evaluation of availability and operational cost for a virtualized server system using stochastic reward nets. The Journal of Supercomputing, 74(6):2903, June 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2428-0.pdf. See [1678].

Su:2018:HED

[1778] Nan Su, Huaxi Gu, Kun Wang, Xiaoshan Yu, and Bowen Zhang. A highly efficient dynamic router for application-oriented network on chip. The Journal of Supercomputing, 74(7): 2905–2915, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2018:JOM

[1779] Yonghui Li, Kun Wang, Huaxi Gu, Yintang Yang, Nan Su, Yawen Chen, and Haibo Zhang. A joint optimization method for NoC topology generation. The Journal of Supercomputing, 74(7): 2916–2934, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qu:2018:IEE

[1780] Xilong Qu, Peng Xiao, and Lirong Huang. Improving the energy efficiency and performance of data-intensive workflows in virtualized clouds. The Journal of Supercomputing, 74(7): 2935–2955, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Guerrero:2018:ROC

[1781] Carlos Guerrero, Isaac Lera, and Carlos Juiz. Resource optimization of container orchestration: a case study in multi-cloud microservices-based applications. The Journal of Supercomputing, 74(7):2956–2983, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Riahi:2018:MOD

[1782] Montassar Riahi and Saoussen Krichen. A multi-objective decision support framework for virtual machine placement in cloud data centers: a real case study. The Journal of Supercomputing, 74(7):2984–3015, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xu:2018:IAV

[1783] Heyang Xu, Yang Liu, Wei Wei, and Wenqiang Zhang. Incentiveaware virtual machine scheduling in cloud computing. The Journal of Supercomputing, 74(7):3016–3038, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2349-y.pdf.

Maurya:2018:BTS

[1784] Ashish Kumar Maurya and Anil Kumar Tripathi. On benchmarking task scheduling algorithms for heterogeneous computing systems. The Journal of Supercomputing, 74(7):3039–3070, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Proficz:2018:IAR

[1785] Jerzy Proficz. Improving all-reduce collective operations for imbalanced process arrival patterns. The Journal of Supercomputing, 74(7):3071–3092, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2356-z.pdf.

Musa:2018:RTT

[1786] Akihiro Musa, Osamu Watanabe, Hiroshi Matsuoka, Hiroaki Hokari, Takuya Inoue, Yoichi Murashima, Yusaku Ohta, Ryota Hino, Shunichi Koshimura, and Hiroaki Kobayashi. Real-time tsunami inundation forecast system for tsunami disaster prevention and mitigation. The Journal of Supercomputing, 74(7):3093–3113, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.

                                        com/content/pdf/10.1007/s11227-
                                        018-2363-0.pdf.

Wang:2018:PBC

[1787] Jingxian Wang, Yongmei Sun, and Yuefeng Ji. Priority-based capacity and power allocation in co-located WBANs using Stackelberg and bargaining games. The Journal of Supercomputing, 74(7):3114–3147, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anandhi:2018:ARR

[1788] S. Anandhi, R. Anitha, and Venkatasamy Sureshkumar. An automatic RFID reader-to-reader delegation protocol for SCM in cloud computing environment. The Journal of Supercomputing, 74(7): 3148–3167, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2018:SWT

[1789] Li Yu, Zhou Zhou, Yuping Fan, Michael E. Papka, and Zhiling Lan. System-wide trade-off modeling of performance, power, and resilience on petascale systems. The Journal of Supercomputing, 74(7):3168–3192, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hasanikhah:2018:EIS

[1790] Narjes Hasanikhah, Siavash Amin-Nejad, Ghafar Darvish, and M. R. Moniri. Efficient implementation of space-time adaptive processing for adaptive weights calculation based on floating point FPGAs. The Journal of Supercomputing, 74(7):3193–3210, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Richter:2018:BLA

[1791] Edward Richter, Spencer Valancius, Josiah McClanahan, John Mixter, and Ali Akoglu. Balancing the learning ability and memory demand of a perceptron-based dynamically trainable neural network. The Journal of Supercomputing, 74(7):3211–3235, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kijsipongse:2018:HGC

[1792] Ekasit Kijsipongse, Apivadee Piyatumrong, and Suriya U-ruekolan. A hybrid GPU cluster and volunteer computing platform for scalable deep learning. The Journal of Supercomputing, 74 (7):3236–3263, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sanchez-Lara:2018:SID

[1793] R. Sanchez-Lara, J. A. Trejo-Sanchez, J. L. Lopez-Martinez, and J. A. Alvarez-Chavez. Simulation of an inelastic dispersive phenomenon: stimulated Brillouin scattering in a singlemode fiber segment through parallelism. The Journal of Supercomputing, 74(7):3264–3277, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Miyata:2018:HSA

[1794] Takafumi Miyata. A heuristic search algorithm based on subspaces for PageRank computation. The Journal of Supercomputing, 74(7):3278–3294, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Said:2018:DPE

[1795] Omar Said and Amr Tolba. Design and performance evaluation of mixed multicast architecture for Internet of Things environment. The Journal of Supercomputing, 74(7):3295–3328, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dezhabad:2018:LBD

[1796] Naghmeh Dezhabad and Saeed Sharifian. Learning-based dynamic scalable load-balanced firewall as a service in network function-virtualized cloud computing environments. The Journal of Supercomputing, 74(7):3329–3358, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abderrahim:2018:BBD

[1797] Wiem Abderrahim and Zied Choukair. Brokerage-based dependability integration in cloud computing services. The Journal of Supercomputing, 74(7): 3359–3387, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gao:2018:SSR

[1798] Lan Gao, Rui Wang, Yunlong Xu, Hailong Yang, Zhongzhi Luan, Depei Qian, Han Zhang, and Jihong Cai. SRAM- and STT–RAM-based hybrid, shared last-level cache for onchip CPU–GPU heterogeneous architectures. The Journal of Supercomputing, 74(7):3388–3414, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gavagsaz:2018:LBR

[1799] Elaheh Gavagsaz, Ali Rezaee, and Hamid Haj Seyyed Javadi. Load balancing in reducers for skewed data in MapReduce systems by using scalable simple random sampling. The Journal of Supercomputing, 74(7):3415–3440, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Navarro-Hinojosa:2018:PBV

[1800] Octavio Navarro-Hinojosa, Sergio Ruiz-Loza, and Moisés Alencastre-Miranda. Physically based visual simulation of the Lattice Boltzmann method on the GPU: a survey. The Journal of Supercomputing, 74(7): 3441–3467, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lao:2018:FIS

[1801] Bin Lao, Ge Nong, Wai Hong Chan, and Yi Pan. Fast induced sorting suffixes on a multicore machine. The Journal of Supercomputing, 74(7): 3468–3485, July 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2018:CDE

[1802] Wenbing Zhao and Mary Q. Yang.

Correction to: Dependability enhancing mechanisms for integrated clinical environments. The Journal of Supercomputing, 74(7):3486–3487, July 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2410-x.pdf. See [1606].

Lee:2018:AMM

[1803] Taejin Lee, Bomin Choi, Youngsang Shin, and Jin Kwak. Automatic malware mutant detection and group classification based on the n-gram and clustering coefficient. The Journal of Supercomputing, 74(8):3489–3503, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        015-1594-6.pdf.

Chaudhry:2018:IRB

[1804] Shehzad Ashraf Chaudhry, Husnain Naqvi, Mohammad Sabzinejad Farash, Taeshik Shon, and Muhammad Sher. An improved and robust biometricsbased three factor authentication scheme for multiserver environments. The Journal of Supercomputing, 74(8): 3504–3520, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Choi:2018:SIF

[1805] Jongseok Choi, Youngjin In, Changjun Park, Seonhee Seok, Hwajeong Seo, and Howon Kim. Secure IoT framework and 2D architecture for end-to-end security. The Journal of Supercomputing, 74(8):3521–3535, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [1806].

Choi:2018:ESI

[1806] Jongseok Choi, Youngjin In, Changjun Park, Seonhee Seok, Hwajeong Seo, and Howon Kim. Erratum to: Secure IoT framework and 2D architecture for End-To-End security. The Journal of Supercomputing, 74(8):3536, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
016-1931-4.pdf. See [1805].

Choi:2018:CEM

[1807] Kwanghoon Choi, Sung-Hwa Lim, and Jai-Hoon Kim. Cost-effective monitoring algorithm for cyber-physical system platform using combined spatiotemporal model. The Journal of Supercomputing, 74(8):3537–3548, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:USM

[1808] Tae-Yoon Kim and Eui-Jik Kim. Uplink scheduling of MU–MIMO gateway for massive data acquisition in Internet of Things. The Journal of Supercomputing, 74(8):3549–3563, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:TNE

[1809] Yongsung Kim, Eenjun Hwang, and Seungmin Rho. Twitter news-in-education platform for social, collaborative, and flipped learning. The Journal of Supercomputing, 74(8):3564– 3582, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:DIU

[1810] Dongwoo Kim, Sangwho Kim, and Jaecheol Ryou. Design and implementation of user-level dynamic binary instrumentation on ARM architecture. The Journal of Supercomputing, 74(8):

3583–3595, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Salehan:2018:SES

[1811] Mohammad Salehan, Dan J. Kim, and Chulmo Koo. A study of the effect of social trust, trust in social networking services, and sharing attitude, on two dimensions of personal information sharing behavior. The Journal of Supercomputing, 74(8):3596– 3619, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Youn:2018:DAH

[1812] Taek-Young Youn, Nam-Su Jho, and Ku-Young Chang. Design of additive homomorphic encryption with multiple message spaces for secure and practical storage services over encrypted data. The Journal of Supercomputing, 74(8): 3620–3638, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Choi:2018:DLC

[1813] Younggeun Choi, Kyounghwan Yoo, Shin Jin Kang, Beomjoo Seo, and Soo Kyun Kim. Development of a lowcost wearable sensing glove with multiple inertial sensors and a light and fast orientation estimation algorithm. The Journal of Supercomputing, 74(8): 3639–3652, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See erratum [1814].

Choi:2018:EDL

[1814] Younggeun Choi, Kyounghwan Yoo, Shin Jin Kang, Beomjoo Seo, and Soo Kyun Kim. Erratum to: Development of a low-cost wearable sensing glove with multiple inertial sensors and a light and fast orientation estimation algorithm. The Journal of Supercomputing, 74(8):3653, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-2021-y.pdf. See [1813].

Kim:2018:CSN

[1815] Mucheol Kim and Sangyong Han. Cognitive social network analysis for supporting the reliable decision-making process. The Journal of Supercomputing, 74(8):3654–3665, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Loghman:2018:FDE

[1816] Maziar Loghman, Joohee Kim, and Kyuwon Choi. Fast depth estimation using semi-global matching and adaptive stripe-based optimization. The Journal of Supercomputing, 74(8):3666–3684, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khan:2018:APS

[1817] Imran Khan, Shehzad Ashraf Chaudhry, Muhammad Sher, Javed I. Khan, and Muhammad Khurram Khan. An anonymous and provably secure biometric-based authentication scheme using chaotic maps for accessing medical drop box data. The Journal of Supercomputing, 74(8):3685–3703, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jo:2018:DFS

[1818] WooYeon Jo, Hyunsoo Chang, and Taeshik Shon. Digital forensic science approach by file recovery research. The Journal of Supercomputing, 74(8): 3704–3725, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xie:2018:LFT

[1819] Ruilian Xie, Jueping Cai, Xin Xin, and Bo Yang. LBFT: a fault-tolerant routing algorithm for load-balancing network-on-chip based on odd–even turn model. The Journal of Supercomputing, 74(8):3726–3747, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Huang:2018:RRI

[1820] Cheng-Ta Huang, Cheng-Hsing Yang, Wei-Jen Wang, Shiau-Rung Tsui, and Shiuh-Jeng Wang. Raw reversibility of information hiding on the basis of VQ systems. The Journal of Supercomputing, 74(8):3748–3777, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moon:2018:FPC

[1821] Jihoon Moon, Jinwoong Park, Eenjun Hwang, and Sanghoon Jun. Forecasting power consumption for higher educational institutions based on machine learning. The Journal of Supercomputing, 74(8):3778–3800, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wong:2018:TFI

[1822] Kok-Seng Wong and Myung Ho Kim. Toward a fair indictment for sealedbid auction with self-enforcing privacy.

The Journal of Supercomputing, 74(8): 3801–3819, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Raji:2018:RFC

[1823] Mohsen Raji and Behnam Ghavami.

Redressing fork constraints in nanoscale quasi-delay-insensitive asynchronous pipelines. The Journal of Supercomputing, 74(8):3820–3840, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kang:2018:PCB

[1824] Dongwann Kang and Sanghyun Seo.

Photo collage-based photograph display system on mobile computing platform. The Journal of Supercomputing, 74(8):3841–3854, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2018:RAG

[1825] Je-Ho Park, Young B. Park, and Soo-Kyung Choi. Representation and automatic generation of state-transition mapping tree. The Journal of Supercomputing, 74(8):3855–3874, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hu:2018:NSF

[1826] Danqi Hu, Fang Lv, Chenxi Wang, Hui-Min Cui, Lei Wang, Ying Liu, and Xiao-Bing Feng. NVM Streaker: a fast and reconfigurable performance simulator for non-volatile memory-based memory architecture. The Journal of Supercomputing, 74(8):3875–3903, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:FBA

[1827] Sangmin Lee, Soon J. Hyun, Hong-Yeon Kim, and Young-Kyun Kim. Fair bandwidth allocating and strip-aware prefetching for concurrent read streams and striped RAIDs in distributed file systems. The Journal of Supercomputing, 74(8):3904–3932, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fedoruk:2018:DRJ

[1828] John Fedoruk, Byron Schmuland, Julia Johnson, and Giseon Heo. Dimensionality reduction via the Johnson– Lindenstrauss Lemma: theoretical and empirical bounds on embedding dimension. The Journal of Supercomputing, 74(8):3933–3949, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shooshtarian:2018:SVS

[1829] L. Shooshtarian, F. Safaei, and A. Tizghadam. Scaling-up versus scaling-out networking in data centers: a comparative robustness analysis. The Journal of Supercomputing, 74 (8):3950–3974, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Baydoun:2018:CGP

[1830] Mohammed Baydoun, Hassan Ghaziri, and Mohammed Al-Husseini. CPU and GPU parallelized kernel K-means. The Journal of Supercomputing, 74(8): 3975–3998, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:EPA

[1831] Yuzhu Wang, Jinrong Jiang, Junqiang Zhang, Juanxiong He, He Zhang, Xuebin Chi, and Tianxiang Yue. An efficient parallel algorithm for the coupling of global climate models and regional climate models on a large-scale multicore cluster. The Journal of Supercomputing, 74(8):3999–4018, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gomez-Martin:2018:ORP

[1832] César Gómez-Martín and Miguel A.

Vega-Rodríguez. Optimization of resources in parallel systems using a multiobjective artificial bee colony algorithm. The Journal of Supercomputing, 74(8):4019–4036, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2018:QQG

[1833] Long Chen, Jigang Wu, Gangqiang Zhou, and Longjie Ma. QUICK: QoSguaranteed efficient cloudlet placement in wireless metropolitan area networks. The Journal of Supercomputing, 74(8): 4037–4059, August 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pinheiro:2018:PPS

[1834] Thiago Felipe da Silva Pinheiro, Francisco Airton Silva, Iure Fé, Sokol Kosta, and Paulo Maciel. Performance prediction for supporting mobile applications’ offloading. The Journal of Supercomputing, 74(8):4060–4103, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kotteda:2018:PPI

[1835] V. M. Krushnarao Kotteda, Vinod Kumar, and William Spotz. Performance of preconditioned iterative solvers in MFiX–Trilinos for fluidized beds. The Journal of Supercomputing, 74(8):4104–4126, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bordel:2018:PEC

[1836] Borja Bordel, Ramón Alcarria, Diego Sánchez de Rivera, and Tomás Robles. Process execution in cyber–physical systems using cloud and cyber–physical Internet services. The Journal of Supercomputing, 74(8):4127–4169, August 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jaeik:2018:ESS

[1837] Cho Jaeik, Naveen Chilamkurti, and S. J. Wang. Editorial of special section on enabling technologies for industrial and smart sensor Internet of Things systems. The Journal of Supercomputing, 74(9):4171–4172, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-2568-2.pdf.

Agrawal:2018:RLR

[1838] Megha Agrawal, Tarun Kumar Bansal, Donghoon Chang, Amit Kumar Chauhan, Seokhie Hong, Jinkeon Kang, and Somitra Kumar Sanadhya. RCB: leakage-resilient authenticated encryption via re-keying. The Journal of Supercomputing, 74(9):4173– 4198, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tekchandani:2018:SCC

[1839] Rajkumar Tekchandani, Rajesh Bhatia, and Maninder Singh. Semantic code clone detection for Internet of Things applications using reaching definition and liveness analysis. The Journal of Supercomputing, 74(9): 4199–4226, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Batalla:2018:VVP

[1840] Jordi Mongay Batalla, Konrad Sienkiewicz, Waldemar Latoszek, Piotr Krawiec, Constandinos X. Mavromoustakis, and George Mastorakis. Validation of virtualization platforms for I-IoT purposes. The Journal of Supercomputing, 74(9): 4227–4241, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-016-1844-2.pdf.

Kim:2018:IIC

[1841] Jihyun Kim, Yonghun Jeon, and Howon Kim. The intelligent IoT common service platform architecture and service implementation. The Journal of Supercomputing, 74(9):4242– 4260, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:ARD

[1842] Sung Ryoung Kim, Jeong Nyeo Kim, Sung Tae Kim, Sunwoo Shin, and Jeong Hyun Yi. Anti-reversible dynamic tamper detection scheme using distributed image steganography for IoT applications. The Journal of Supercomputing, 74(9):4261– 4280, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alamr:2018:SEB

[1843] Amjad Ali Alamr, Firdous Kausar, Jongsung Kim, and Changho Seo. A secure ECC-based RFID mutual authentication protocol for Internet of Things. The Journal of Supercomputing, 74(9):4281–4294, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Huang:2018:VBD

[1844] Cheng-Ta Huang, Min-Yi Tsai, Li-Chiun Lin, Wei-Jen Wang, and Shiuh-Jeng Wang. VQ-based data hiding in IoT networks using two-level encoding with adaptive pixel replacements. The Journal of Supercomputing, 74(9): 4295–4314, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tsai:2018:SSR

[1845] Ming-Fong Tsai, Ye Chin Kiong, and Ang Sinn. Smart service relying on Internet of Things technology in parking systems. The Journal of Supercomputing, 74(9):4315–4338, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:CBS

[1846] Shiyong Wang, Jiafu Wan, Muhammad Imran, Di Li, and Chunhua Zhang. Cloud-based smart manufacturing for personalized candy packing application. The Journal of Supercomputing, 74(9):4339–4357, September 2018.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yeh:2018:NPS

[1847] Wei-Chang Yeh and Jsen-Shung Lin.

New parallel swarm algorithm for smart sensor systems redundancy allocation problems in the Internet of Things. The Journal of Supercomputing, 74(9):4358–4384, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:FIF

[1848] Hwi-Ho Lee, Jung-Hyok Kwon, and Eui-Jik Kim. FS-IIoTSim: a flexible and scalable simulation framework for performance evaluation of industrial Internet of Things systems. The Journal of Supercomputing, 74(9): 4385–4402, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2018:WWO

[1849] Jaehak Yu, NamKyung Lee, Cheol-Sig Pyo, and Yang Sun Lee. WISE: web of object architecture on IoT environment for smart home and building energy management. The Journal of Supercomputing, 74(9):4403– 4418, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jabbar:2018:RBI

[1850] Sohail Jabbar, Murad Khan, Bhagya Nathali Silva, and Kijun Han. A REST-based industrial web of things’ framework for smart warehousing. The Journal of Supercomputing, 74(9):4419– 4433, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:FCB

[1851] Sung Jin Kim and Taeshik Shon. Field classification-based novel fuzzing case generation for ICS protocols. The Journal of Supercomputing, 74(9):4434– 4450, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2018:HPF

[1852] Ching-Han Chen, Ming-Yi Lin, and Xing-Chen Guo. High-performance Fieldbus application-specific integrated circuit design for industrial smart sensor networks. The Journal of Supercomputing, 74(9):4451–4469, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lin:2018:ANN

[1853] Jia-Yang Lin, Yi-Ting Hsieh, Trong Nghia Le, and Wen-Long Chin. AnyNoC: new network on a chip switching using the shared-memory and output-queue techniques for complex Internet of Things systems. The Journal of Supercomputing, 74(9):4470–4480, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chang:2018:PEF

[1854] Hangbae Chang. Performance evaluation framework design for smart sensor business. The Journal of Supercomputing, 74(9):4481–4496, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Koo:2018:EHL

[1855] ChangJin Koo and JeongYeon Kim.

Enforcing high-level security policies for Internet of Things. The Journal of Supercomputing, 74(9):4497– 4505, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:DIC

[1856] Seokcheol Lee, Seokjun Lee, Hyunguk Yoo, Sungmoon Kwon, and Taeshik Shon. Design and implementation of cybersecurity testbed for industrial IoT systems. The Journal of Supercomputing, 74(9):4506–4520, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lin:2018:AGK

[1857] Tzu-Wei Lin and Chien-Lung Hsu.

Anonymous group key agreement protocol for multi-server and mobile environments based on Chebyshev chaotic maps. The Journal of Supercomputing, 74(9):4521–4541, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aadil:2018:CAI

[1858] Farhan Aadil, Waleed Ahsan, Zahoor Ur Rehman, Peer Azmat Shah, Seungmin Rho, and Irfan Mehmood. Clustering algorithm for Internet of Vehicles (IoV) based on dragonfly optimizer (CAVDO). The Journal of Supercomputing, 74(9):4542–4567, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kalsoom:2018:DRB

[1859] Anum Kalsoom, Muazzam Maqsood, Mustansar Ali Ghazanfar, Farhan Aadil, and Seungmin Rho. A dimensionality reduction-based efficient software fault prediction using Fisher linear discriminant analysis (FLDA). The Journal of Supercomputing, 74(9): 4568–4602, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:MFB

[1860] Keonsoo Lee, Yang Sun Lee, and Yunyoung Nam. A model of FSMbased planner and dialogue supporting system for emergency call services. The Journal of Supercomputing, 74(9): 4603–4612, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sanapala:2018:STB

[1861] Kishore Sanapala, Sakthivel R, and Sang-Soo Yeo. Schmitt trigger-based single-ended 7T SRAM cell for Internet of Things (IoT) applications. The Journal of Supercomputing, 74(9): 4613–4622, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kamran:2018:QAV

[1862] Kamran and Babar Nazir. QoS-aware VM placement and migration for hybrid cloud infrastructure. The Journal of Supercomputing, 74(9):4623– 4646, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khoroush:2018:AMH

[1863] Somayeh Khoroush, Midia Reshadi, and Ahmad Khademzadeh. Application mapping in hybrid photonic networks-on-chip for reducing insertion loss. The Journal of Supercomputing, 74(9):4647–4671, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Meyer:2018:SPT

[1864] Michael Meyer, Yuichi Okuyama, and Abderazek Ben Abdallah. SAFT-PHENIC: a thermal-aware microring fault-resilient photonic NoC. The Journal of Supercomputing, 74(9):4672– 4695, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmadpour:2018:NFT

[1865] Seyed-Sajad Ahmadpour and Mohammad Mosleh. A novel fault-tolerant multiplexer in quantum-dot cellular automata technology. The Journal of Supercomputing, 74(9):4696–4716, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [2251].

Mohammadi:2018:ILP

[1866] Somayeh Mohammadi, Hossein Pedram, and Latif PourKarimi. Integer linear programming-based cost optimization for scheduling scientific workflows in multi-cloud environments. The Journal of Supercomputing, 74(9): 4717–4745, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:ACP

[1867] Sung Min Lee, Su-Kyung Yoon, Jeong-Geun Kim, and Shin-Dug Kim. Adaptive correlated prefetch with large-scale hybrid memory system for stream processing. The Journal of Supercomputing, 74(9):4746–4770, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Irshad:2018:ESD

[1868] Azeem Irshad, Husnain Naqvi, Shehzad Ashraf Chaudhry, Shouket Raheem, Saru Kumari, Ambrina Kanwal, and Muhammad Usman. An efficient and secure design of multi-server authenticated key agreement protocol. The Journal of Supercomputing, 74(9): 4771–4797, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:NDF

[1869] Lei Wang and Guangjun Xie. Novel designs of full adder in quantum-dot cellular automata technology. The Journal of Supercomputing, 74(9):4798– 4816, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Torquato:2018:MAP

[1870] Matheus Torquato, I M Umesh, and Paulo Maciel. Models for availability and power consumption evaluation of a private cloud with VMM rejuvenation enabled by VM Live Migration. The Journal of Supercomputing, 74(9): 4817–4841, September 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tosun:2018:EAP

[1871] Suleyman Tosun and Vahid Babaei Ajabshir. Energy-aware partitioning of fault-tolerant irregular topologies for 3D network-on-chips. The Journal of Supercomputing, 74(9):4842– 4863, September 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chaudhry:2018:TCI

[1872] Junaid Chaudhry, Al-Sakib Khan Pathan, Mubashir Husain Rehmani, and Ali Kashif Bashir. Threats to critical infrastructure from AI and human intelligence. The Journal of Supercomputing, 74(10):4865–4866, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2614-0.pdf.

Shafiq:2018:MLA

[1873] Muhammad Shafiq, Xiangzhan Yu, Ali Kashif Bashir, Hassan Nazeer Chaudhry, and Dawei Wang. A machine learning approach for feature selection traffic classification using security analysis. The Journal of Supercomputing, 74(10):4867–4892, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2018:EEC

[1874] Wencheng Yang, Song Wang, Guanglou Zheng, Junaid Chaudhry, and Craig Valli. ECB4CI: an enhanced cancelable biometric system for securing critical infrastructures. The Journal of Supercomputing, 74(10):4893– 4909, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Soleimani:2018:RTI

[1875] Mohammad Hassan Mojtahed Soleimani, Muharram Mansoorizadeh, and Mohammad Nassiri. Real-time identification of three Tor pluggable transports using machine learning techniques. The Journal of Supercomputing, 74(10):4910–4927, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sedjelmaci:2018:CSM

[1876] Hichem Sedjelmaci and Sidi Mohamed Senouci. Cyber security methods for aerial vehicle networks: taxonomy, challenges and solution. The Journal of Supercomputing, 74(10):4928– 4944, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sifah:2018:CBB

[1877] Emmanuel Boateng Sifah, Qi Xia, Kwame Opuni-Boachie Obour Agyekum, Sandro Amofa, Jianbin Gao, Ruidong Chen, Hu Xia, James C. Gee, Xiaojiang Du, and Mohsen Guizani. Chainbased big data access control infrastructure. The Journal of Supercomputing, 74(10):4945–4964, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nur:2018:ICA

[1878] Abdullah Yasin Nur and Mehmet Engin Tozal. Identifying critical autonomous systems in the Internet. The Journal of Supercomputing, 74(10): 4965–4985, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ghafir:2018:STC

[1879] Ibrahim Ghafir, Jibran Saleem, Mohammad Hammoudeh, Hanan Faour, Vaclav Prenosil, Sardar Jaf, Sohail Jabbar, and Thar Baker. Security threats to critical infrastructure: the human factor. The Journal of Supercomputing, 74(10):4986–5002, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2337-2.pdf.

Hussain:2018:TOB

[1880] Mubashar Hussain, Mansoor Ahmed, Hasan Ali Khattak, Muhammad Imran, Abid Khan, Sadia Din, Awais Ahmad, Gwanggil Jeon, and Alavalapati Goutham Reddy. Towards ontologybased multilingual URL filtering: a big data problem. The Journal of Supercomputing, 74(10):5003–5021, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ubaid:2018:SSC

[1881] Saqib Ubaid, M. Farrukh Shafeeq, Majid Hussain, Ali Hammad Akbar, Abdelrahman Abuarqoub, M. Sultan Zia, and Beenish Abbas. SCOUT: a sink camouflage and concealed data delivery paradigm for circumvention of sinktargeted cyber threats in wireless sensor networks. The Journal of Supercomputing, 74(10):5022–5040, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmed:2018:IPM

[1882] Mohiuddin Ahmed and Abu S. S. M. Barkat Ullah. Infrequent pattern mining in smart healthcare environment using data summarization. The Journal of Supercomputing, 74(10): 5041–5059, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khan:2018:FID

[1883] Rashid Masood Khan, Waseem Iqbal, Muhammad Faisal Amjad, Haider Abbas, Hammad Afzal, Abdul Rauf, and Maruf Pasha. Forensic investigation to detect forgeries in ASF files of contemporary IP cameras. The Journal of Supercomputing, 74(10):5060– 5081, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mirza:2018:CCB

[1884] Muhammad Ayzed Mirza, Mudassar Ahmad, Muhammad Asif Habib, Nasir Mahmood, C. M. Nadeem Faisal, and Usman Ahmad. CDCSS: cluster-based distributed cooperative spectrum sensing model against primary user emulation (PUE) cyber attacks. The Journal of Supercomputing, 74(10):5082– 5098, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Malomo:2018:NGC

[1885] Olumide O. Malomo, Danda B. Rawat, and Moses Garuba. Next-generation cybersecurity through a blockchainenabled federated cloud framework. The Journal of Supercomputing, 74 (10):5099–5126, October 2018. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anjum:2018:EAP

[1886] Adeel Anjum, Naveed Ahmad, Saif U. R. Malik, Samiya Zubair, and Basit Shahzad. An efficient approach for publishing microdata for multiple sensitive attributes. The Journal of Supercomputing, 74(10):5127–5155, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mehmood:2018:NMN

[1887] Amjad Mehmood, Mithun Mukherjee, Syed Hassan Ahmed, Houbing Song, and Khalid Mahmood Malik. NBC-MAIDS: Naïve Bayesian classification technique in multi-agent systemenriched IDS for securing IoT against DDoS attacks. The Journal of Supercomputing, 74(10):5156–5170, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ibrahim:2018:SRL

[1888] Ahmed Ibrahim, Craig Valli, Ian McAteer, and Junaid Chaudhry. A security review of local government using NIST CSF: a case study. The Journal of Supercomputing, 74(10):5171–5186, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2479-2.pdf.  See correction

[2281].

Al-Turjman:2018:CSS

[1889] Fadi Al-Turjman and Sinem Alturjman. Confidential smart-sensing framework in the IoT era. The Journal of Supercomputing, 74(10):5187– 5198, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hawedi:2018:MTI

[1890] Mohamed Hawedi, Chamseddine Talhi, and Hanifa Boucheneb. Multi-tenant intrusion detection system for public cloud (MTIDS). The Journal of Supercomputing, 74(10):5199–5230, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Min:2018:LCW

[1891] Se-Dong Min, Chang-Won Wang, Hwa-Min Lee, and Bong-Keun Jung. A low cost wearable wireless sensing system for paretic hand management after stroke. The Journal of Supercomputing, 74(10):5231–5240, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Siddiqa:2018:MSI

[1892] Aisha Siddiqa, Ahmad Karim, and Victor Chang. Modeling SmallClient indexing framework for big data analytics. The Journal of Supercomputing, 74(10):5241–5262, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chunlin:2018:MQS

[1893] Li Chunlin, Tang Jianhang, and Luo Youlong. Multi-queue scheduling of heterogeneous jobs in hybrid geodistributed cloud environment. The Journal of Supercomputing, 74(10): 5263–5292, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yoon:2018:DDN

[1894] Su-Kyung Yoon, Young-Sun Youn, Jeong-Geun Kim, and Shin-Dug Kim. Design of DRAM–NAND flash hybrid main memory and Q-learning-based prefetching method. The Journal of Supercomputing, 74(10):5293– 5313, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Malekimajd:2018:OFC

[1895] M. Malekimajd, D. Ardagna, M. Ciavotta, E. Gianniti, M. Passacantando, and A. M. Rizzi. An optimization framework for the capacity allocation and admission control of MapReduce jobs in cloud systems. The Journal of Supercomputing, 74(10):5314–5348, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mansouri:2018:HDR

[1896] N. Mansouri and M. M. Javidi. A hybrid data replication strategy with fuzzy-based deletion for heterogeneous cloud data centers. The Journal of Supercomputing, 74(10):5349–5372, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bounouni:2018:ABP

[1897] Mahdi Bounouni and Louiza Bouallouche-Medjkoune. Acknowledgment-based punishment and stimulation scheme for mobile ad hoc network. The Journal of Supercomputing, 74(10):5373– 5398, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khalid:2018:OLB

[1898] Yasir Noman Khalid, Muhammad Aleem, Radu Prodan, Muhammad Azhar Iqbal, and Muhammad Arshad Islam. E-OSched: a load balancing scheduler for heterogeneous multicores. The Journal of Supercomputing, 74(10):5399–5431, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Seyfari:2018:OIN

[1899] Yousef Seyfari, Shahriar Lotfi, and Jaber Karimpour. Optimizing internest data locality in imperfect stencils based on loop blocking. The Journal of Supercomputing, 74(10):5432– 5460, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mostafaeipour:2018:PAT

[1900] Ali Mostafaeipour, Alireza Goli, and Mojtaba Qolipour. Prediction of air travel demand using a hybrid artificial neural network (ANN) with Bat and Firefly algorithms: a case study. The Journal of Supercomputing, 74(10): 5461–5484, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khansari:2018:VSS

[1901] Mina Emami Khansari, Saeed Sharifian, and Seyed Ahmad Motamedi. Virtual sensor as a service: a new multicriteria QoS-aware cloud service composition for IoT applications. The Journal of Supercomputing, 74(10):5485– 5512, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alyatama:2018:CMA

[1902] Anwar Alyatama, Asmaa Alsumait, and Maryam Alotaibi. Continuous memory allocation model for cloud services. The Journal of Supercomputing, 74(10):5513–5538, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Thuan:2018:SLF

[1903] Bui Thi Thuan, Lam Boi Ngoc, and Keiichi Kaneko. A stochastic link-fault-tolerant routing algorithm in folded hypercubes. The Journal of Supercomputing, 74(10):5539–5557, October 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cheong:2018:CBN

[1904] Se-Hang Cheong and Yain-Whar Si.

CWBound: boundary node detection algorithm for complex non-convex mobile ad hoc networks. The Journal of Supercomputing, 74(10):5558– 5577, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Safari:2018:PDC

[1905] Monireh Safari and Reihaneh Khorsand. PL-DVFS: combining poweraware list-based scheduling algorithm with DVFS technique for real-time tasks in cloud computing. The Journal of Supercomputing, 74(10):5578– 5600, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

He:2018:INB

[1906] Chaobo He, Xiang Fei, Hanchao Li, Yong Tang, Hai Liu, and Shuangyin Liu. Improving NMF-based community discovery using distributed robust nonnegative matrix factorization with Sim-Rank similarity measure. The Journal of Supercomputing, 74(10):5601– 5624, October 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia:2018:RPH

[1907] J. Daniel García, Kevin Hammond, and Lutz Schubert. Reengineering for parallelism in heterogeneous parallel platforms. The Journal of Supercomputing, 74(11):5625–5627, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2651-8.pdf.

Castello:2018:EIR

[1908] Adrián Castelló, Antonio J. Peña, Rafael Mayo, Judit Planas, Enrique S. Quintana-Ortí, and Pavan Balaji. Exploring the interoperability of remote GPGPU virtualization using rCUDA and directive-based programming models. The Journal of Supercomputing, 74(11):5628–5642, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2018:MGM

[1909] Lu Li and Christoph Kessler. MeterPU: a generic measurement abstraction API. The Journal of Supercomputing, 74(11):5643–5658, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Danelutto:2018:DSP

[1910] Marco Danelutto, Tiziano De Matteis, Gabriele Mencagli, and Massimo Torquati. Data stream processing via code annotations. The Journal of Supercomputing, 74(11):5659– 5673, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Astorga:2018:ADP

[1911] David del Rio Astorga, Rafael Sotomayor, Luis Miguel Sanchez, Javier Garcia Blas, Alejandro Calderon, and Javier Fernandez. Assessing and discovering parallelism in C++ code for heterogeneous platforms. The Journal of Supercomputing, 74(11):5674– 5689, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aldinucci:2018:PPI

[1912] M. Aldinucci, M. Danelutto, M. Drocco, P. Kilpatrick, C. Misale, G. Peretti Pezzi, and M. Torquati. A parallel pattern for iterative stencil + reduce. The Journal of Supercomputing, 74(11):5690–5705, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kang:2018:SEL

[1913] Byungseok Kang and Hyunseung Choo. An SDN-enhanced loadbalancing technique in the cloud system. The Journal of Supercomputing, 74(11):5706–5729, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bui:2018:EFD

[1914] Dinh-Mao Bui, Thien Huynh-The, and Sungyoung Lee. Early fault detection in IaaS cloud computing based on fuzzy logic and prediction technique. The Journal of Supercomputing, 74(11):5730–5745, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ali:2018:DIA

[1915] Toqeer Ali, Roslan Ismail, Shahrulniza Musa, Mohammad Nauman, and Sohail Khan. Design and implementation of an attestation protocol for measured dynamic behavior. The Journal of Supercomputing, 74(11):5746– 5773, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rizvi:2018:SEF

[1916] Syed Rizvi, Jungwoo Ryoo, John Kissell, William Aiken, and Yuhong Liu. A security evaluation framework for cloud security auditing. The Journal of Supercomputing, 74(11):5774– 5796, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2018:ECM

[1917] Yuanlong Yu, Zhifan Ye, Xianghan Zheng, and Chunming Rong. An efficient cascaded method for network intrusion detection based on extreme learning machines. The Journal of Supercomputing, 74(11):5797– 5812, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ibrahim:2018:ABA

[1918] Maged Hamada Ibrahim, Saru Kumari, Ashok Kumar Das, and Vanga Odelu. Attribute-based authentication on the cloud for thin clients. The Journal of Supercomputing, 74(11):5813– 5845, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Luo:2018:IPN

[1919] Gangyi Luo, Zhuzhong Qian, Mianxiong Dong, Kaoru Ota, and Sanglu Lu. Improving performance by network-aware virtual machine clustering and consolidation. The Journal of Supercomputing, 74(11):5846– 5864, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Song:2018:GAH

[1920] Tae-Geon Song, Mehdi Pirahandeh, Cheong-Jin Ahn, and Deok-Hwan Kim. GPU-accelerated high-performance encoding and decoding of hierarchical RAID in virtual machines. The Journal of Supercomputing, 74(11):5865– 5888, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tseng:2018:CBA

[1921] Chun-Hsiung Tseng and Yung-Hui Chen. A camera-based attention level assessment tool designed for classroom usage. The Journal of Supercomputing, 74(11):5889–5902, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jin:2018:FDC

[1922] Hao Jin, Ke Zhou, and Yan Luo. A framework with data-centric accountability and auditability for cloud storage. The Journal of Supercomputing, 74(11):5903–5926, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mo:2018:RUA

[1923] Jiaqing Mo, Zhongwang Hu, and Yuhua Lin. Remote user authentication and key agreement for mobile client-server environments on elliptic curve cryptography. The Journal of Supercomputing, 74(11):5927– 5943, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shojaei:2018:VVM

[1924] Kiamars Shojaei, Faramarz Safi-Esfahani, and Saeed Ayat. VMDFS: virtual machine dynamic frequency scaling framework in cloud computing. The Journal of Supercomputing, 74(11):5944–5979, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mishra:2018:SOU

[1925] Ashish Kumar Mishra, Brajesh Kumar Umrao, and Dharmendra K. Yadav. A survey on optimal utilization of preemptible VM instances in cloud computing. The Journal of Supercomputing, 74(11):5980–6032, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gupta:2018:JFS

[1926] Shaifu Gupta, A. D. Dileep, and Timothy A. Gonsalves. A joint feature selection framework for multivariate resource usage prediction in cloud servers using stability and prediction performance. The Journal of Supercomputing, 74(11):6033–6068, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Roy:2018:IMT

[1927] Samarjit Roy, Dhiman Sarkar, Sourav Hati, and Debashis De. Internet of Music Things: an edge computing paradigm for opportunistic crowdsensing. The Journal of Supercomputing, 74(11):6069–6101, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rezaei-Ravari:2018:RCA

[1928] Maryam Rezaei-Ravari and Vahid Sattari-Naeini. Reliable congestionaware path prediction mechanism in 2D NoCs based on EFuNN. The Journal of Supercomputing, 74(11):6102– 6125, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hui:2018:VMA

[1929] Yang Hui. A virtual machine anomaly detection system for cloud computing infrastructure. The Journal of Supercomputing, 74(11):6126–6134, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2018:ABF

[1930] Feng Zhang, Heng Lin, Jidong Zhai, Jie Cheng, Dingyi Xiang, Jizhong Li, Yunpeng Chai, and Xiaoyong Du. An adaptive breadth-first search algorithm on integrated architectures. The Journal of Supercomputing, 74(11):6135– 6155, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2018:PPA

[1931] Libing Wu, Jing Wang, Sherali Zeadally, and Debiao He. Privacypreserving auditing scheme for shared data in public clouds. The Journal of Supercomputing, 74(11):6156– 6183, November 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See comment [2373].

Khosravi:2018:DRC

[1932] Mohammad Reza Khosravi, Hamid Basri, Habib Rostami, and Sadegh Samadi. Distributed random cooperation for VBF-based routing in highspeed dense underwater acoustic sensor networks. The Journal of Supercomputing, 74(11):6184–6200, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bock:2018:BML

[1933] Nicolas Bock, Christian F. A. Negre, Susan M. Mniszewski, Jamaludin Mohd-Yusof, Bálint Aradi, Jean-Luc Fattebert, Daniel Osei-Kuffuor, Timothy C. Germann, and Anders M. N. Niklasson. The Basic Matrix Library (BML) for quantum chemistry. The Journal of Supercomputing, 74(11): 6201–6219, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Safkhani:2018:ISD

[1934] Masoumeh Safkhani and Mahyar Shariat. Implementation of secret disclosure attack against two IoT lightweight authentication protocols. The Journal of Supercomputing, 74 (11):6220–6235, November 2018. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gad:2018:ZMD

[1935] Ramy Gad, Simon Pickartz, Tim Süß, Lars Nagel, Stefan Lankes, Antonello Monti, and André Brinkmann. Zeroing memory deallocator to reduce checkpoint sizes in virtualized HPC environments. The Journal of Supercomputing, 74(11):6236–6257, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abutaleb:2018:REQ

[1936] M. M. Abutaleb. Robust and efficient QCA cell-based nanostructures of elementary reversible logic gates. The Journal of Supercomputing, 74(11): 6258–6274, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qi:2018:DME

[1937] Hongyuan Qi, Jinchen Xu, and Shaozhong Guo. Detection of the maximum error of mathematical functions. The Journal of Supercomputing, 74(11):6275–6290, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2018:CER

[1938] Huangke Chen, Jianghan Zhu, Guohua Wu, and Lisu Huo. Cost-efficient reactive scheduling for real-time workflows in clouds. The Journal of Supercomputing, 74(11):6291–6309, November 2018.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tavizi:2018:TBR

[1939] Amir Tavizi and Ali Ghaffari. Treebased reliable and energy-aware multicast routing protocol for mobile ad hoc networks. The Journal of Supercomputing, 74(11):6310–6332, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hersonsky:2018:ACM

[1940] Sa’ar Hersonsky. Approximation of conformal mappings and novel applications to shape recognition of planar domains. The Journal of Supercomputing, 74(11):6333–6368, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abdel-hafeez:2018:CFS

[1941] Saleh Abdel-hafeez, Ann Gordon-Ross, and Samer Abubaker. A comparisonfree sorting algorithm on CPUs and GPUs. The Journal of Supercomputing, 74(11):6369–6400, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2018:FCI

[1942] D. S. Park. Future computing with IoT and cloud computing. The Journal of Supercomputing, 74(12): 6401–6407, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-018-2652-7.pdf.

Meng:2018:DDC

[1943] Yi Meng and Chen QingKui. DC-SACA: distributed constraint serviceaware collaborative access algorithm based on large-scale access to the Internet of Things. The Journal of Supercomputing, 74(12):6408–6427, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumari:2018:SAS

[1944] Saru Kumari, Marimuthu Karuppiah, Ashok Kumar Das, Xiong Li, Fan Wu, and Neeraj Kumar. A secure authentication scheme based on elliptic curve cryptography for IoT and cloud servers. The Journal of Supercomputing, 74(12):6428–6453, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Munoz:2018:ATP

[1945] Fernando Román Muñoz, Esteban Alejandro Armas Vega, and Luis Javier García Villalba. Analyzing the traffic of penetration testing tools with an IDS. The Journal of Supercomputing, 74(12):6454–6469, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aslanpour:2018:RPC

[1946] Mohammad Sadegh Aslanpour, Seyed Ebrahim Dashti, Mostafa Ghobaei-Arani, and Ali Asghar Rahmanian. Resource provisioning for cloud applications: a 3-d, provident and flexible approach. The Journal of Supercomputing, 74(12): 6470–6501, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Baek:2018:EEG

[1947] Dusan Baek, Jae-Hyeon Park, and Jung-Won Lee. An energy efficiency grading system for mobile applications based on usage patterns. The Journal of Supercomputing, 74(12):6502– 6515, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:DIF

[1948] Hye-Young Kim. A design and implementation of a framework for games in IoT. The Journal of Supercomputing, 74(12):6516–6528, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        017-1973-2.pdf.

Chen:2018:CWS

[1949] Long Chen and Xiaoping Li. Cloud workflow scheduling with hybrid resource provisioning. The Journal of Supercomputing, 74(12):6529–6553, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Song:2018:HLP

[1950] Binbin Song, Yao Yu, Yu Zhou, Ziqiang Wang, and Sidan Du. Host load prediction with long short-term memory in cloud computing. The Journal of Supercomputing, 74(12):6554– 6568, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chaabouni:2018:EMS

[1951] Taha Chaabouni and Maher Khemakhem. Energy management strategy in cloud computing: a perspective study. The Journal of Supercomputing, 74(12):6569–6597, December 2018.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Munoz:2018:EVW

[1952] Fernando Román Muñoz, Iván Israel Sabido Cortes, and Luis Javier García Villalba. Enlargement of vulnerable web applications for testing. The Journal of Supercomputing, 74(12):6598– 6617, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:MFR

[1953] Dohyun Kim, Yunho Lee, and Sangjin Lee. Mobile forensic reference set (MFReS) and mobile forensic investigation for Android devices. The Journal of Supercomputing, 74(12):6618– 6632, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:IAC

[1954] Jin Wang, Jiayi Cao, R. Simon Sherratt, and Jong Hyuk Park. An improved ant colony optimization-based approach with mobile sink for wireless sensor networks. The Journal of Supercomputing, 74(12):6633– 6645, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:SMD

[1955] Yun kyung Lee, Jeong nyeo Kim, Kyung-Soo Lim, and Hyunsoo Yoon. Secure mobile device structure for trust IoT. The Journal of Supercomputing, 74(12):6646–6664, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ok:2018:MAS

[1956] Seung-Ho Ok, Jae Hoon Shim, and Byungin Moon. Modified adaptive support weight and disparity search range estimation schemes for stereo matching processors. The Journal of Supercomputing, 74(12):6665–6690, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:IST

[1957] Sung-Kil Kim and Junseok Oh. Information science techniques for investigating research areas: a case study in telecommunications policy. The Journal of Supercomputing, 74(12):6691– 6718, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cho:2018:DDF

[1958] Chaeho Cho, Kwangsik Chung, and Yoojae Won. Detection of damaged files and measurement of similarity to originals using entropy graph characteristics. The Journal of Supercomputing, 74(12):6719–6728, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Oh:2018:CDL

[1959] Seon Ho Oh, Geon-Woo Kim, and Kyung-Soo Lim. Compact deep learned feature-based face recognition for Visual Internet of Things. The Journal of Supercomputing, 74(12):6729– 6741, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:LIG

[1960] Dongkyun Kim, Yong-Hwan Kim, Ki-Hyun Kim, Joo-Beom Kim, Gi-Seong You, and Joon-Min Gil. Logically isolated group network for virtual convergence environment over SD–WAN. The Journal of Supercomputing, 74(12): 6742–6752, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Oh:2018:DFL

[1961] Seon Ho Oh, Seung-Wan Han, Bum-Suk Choi, Geon-Woo Kim, and Kyung-Soo Lim. Deep feature learning for person re-identification in a large-scale crowdsourced environment. The Journal of Supercomputing, 74(12):6753– 6765, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chunxia:2018:ESS

[1962] Yin Chunxia and Jin Shunfu. An energy-saving strategy based on multiserver vacation queuing theory in cloud data center. The Journal of Supercomputing, 74(12):6766–6784, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:SRT

[1963] Jeong-Hwan Kim, Kang-Hwi Lee, Jeong-Whan Lee, and Kyeong-Seop Kim. Semi-real-time removal of baseline fluctuations in electrocardiogram (ECG) signals by an infinite impulse response low-pass filter (IIR-LPF). The Journal of Supercomputing, 74(12): 6785–6793, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2018:DMK

[1964] Dazhi Li, Minglu Li, and Jianhua Liu. A dynamic multiple-keys gamebased industrial wireless sensor-cloud authentication scheme. The Journal of Supercomputing, 74(12):6794– 6814, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhu:2018:MOO

[1965] Min Zhu, Dengyin Zhang, and Jin Wang. Multi-objective optimization design for multi-source multicasting MIMO AF relay systems. The Journal of Supercomputing, 74(12):6815– 6830, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Borujeni:2018:FBE

[1966] Elham Mirzavand Borujeni, Dadmehr Rahbari, and Mohsen Nickray. Fogbased energy-efficient routing protocol for wireless sensor networks. The Journal of Supercomputing, 74(12):6831– 6858, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:ISC

[1967] Daewon Lee and HwaMin Lee. IoT service classification and clustering for integration of IoT service platforms. The Journal of Supercomputing, 74(12): 6859–6875, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2018:CVG

[1968] Chao-Tung Yang, Shuo-Tsung Chen, Yu-Sheng Lo, Endah Kristiani, and Yu-Wei Chan. On construction of a virtual GPU cluster with InfiniBand and 10 Gb Ethernet virtualization. The Journal of Supercomputing, 74(12):6876– 6897, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2018:ABM

[1969] JongHyuk Lee, Weidong Shi, and Joon-Min Gil. Accelerated bulk memory operations on heterogeneous multi-core systems. The Journal of Supercomputing, 74(12):6898–6922, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2018:EDP

[1970] Xiaofan Wang, Lei Wang, Shengji Li, and Jin Wang. An event-driven plan recognition algorithm based on intuitionistic fuzzy theory. The Journal of Supercomputing, 74(12):6923– 6938, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ma:2018:ETM

[1971] Linh Van Ma, Gwanghyun Yu, Jin-Young Kim, Yonggwan Won, and Jinsul Kim. An efficient transmission method based on HEVC multi-view adaptive video streaming over P2P network in NFV. The Journal of Supercomputing, 74(12):6939–6959, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jung:2018:JSM

[1972] Daeyong Jung, Myungil Kim, Jungha Lee, Han-Yee Kim, and Daewon Lee. Job submission and monitoring management in integrated computing environment for finite element analysis. The Journal of Supercomputing, 74(12):6960–6975, December 2018.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2018:PAT

[1973] Yan Li and Byeong-Seok Shin. Privacyaware task data management using TPR*-Tree for trajectory-based crowdsourcing. The Journal of Supercomputing, 74(12):6976–6987, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2018:ULP

[1974] Sangho Park, Hyunjin Kim, and Jaecheol Ryou. Utilizing a lightweight PKI mechanism to guarantee a secure service in a cloud environment. The Journal of Supercomputing, 74(12): 6988–7002, December 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yao:2018:PFB

[1975] Jiao Yao, Kaimin Zhang, Yaxuan Dai, and Jin Wang. Power function-based signal recovery transition optimization model of emergency traffic. The Journal of Supercomputing, 74(12):7003– 7023, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2018:CCC

[1976] Nam Yong Kim, Jung Hyun Ryu, Byoung Wook Kwon, Yi Pan, and Jong Hyuk Park. CF-CloudOrch: container fog node-based cloud orchestration for IoT networks. The Journal of Supercomputing, 74(12):7024– 7045, December 2018. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qiang:2019:PSC

[1977] Weizhong Qiang. Performance and security in cloud computing. The Journal of Supercomputing, 75(1):1–3, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2671-4.pdf.

Shi:2019:VSM

[1978] Xiang Shi, Xiaofei Liao, Dayang Zheng, Hai Jin, and Haikun Liu. VMBKS: a shared memory cache system based on booting kernel in cloud. The Journal of Supercomputing, 75(1):4–19, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2019:LAV

[1979] Chao Yu, Leihua Qin, and Jingli Zhou.

A lock-aware virtual machine scheduling scheme for synchronization performance. The Journal of Supercomputing, 75(1):20–32, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tan:2019:RSR

[1980] Yusong Tan, Fuhui Wu, Qingbo Wu, and Xiangke Liao. Resource stealing: a resource multiplexing method for mix workloads in cloud system. The Journal of Supercomputing, 75(1): 33–49, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

You:2019:TDS

[1981] Pengfei You, Zhen Huang, Yuxing Peng, Changjian Wang, and Guofeng Yan. Towards a delivery scheme for speedup of data backup in distributed storage systems using erasure codes. The Journal of Supercomputing, 75(1): 50–64, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hwang:2019:PVR

[1982] Gwan-Hwan Hwang and Yi-Ling Yuan. Proof of violation for response time auditing in cloud systems. The Journal of Supercomputing, 75(1):65–76, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tang:2019:NPV

[1983] Chun ming Tang and Xiao jun Zhang. A new publicly verifiable data possession on remote storage. The Journal of Supercomputing, 75(1):77–91, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:MAP

[1984] Aiping Li, Shuang Tan, and Yan Jia. A method for achieving provable data integrity in cloud computing. The Journal of Supercomputing, 75(1):92– 108, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shen:2019:DDC

[1985] Jinan Shen, Deqing Zou, Hai Jin, Bin Yuan, and Weiqi Dai. A domaindivided configurable security model for cloud computing-based telecommunication services. The Journal of Supercomputing, 75(1):109–122, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:EPS

[1986] Seunghee Kim, Hongyeon Kim, and Jun-Ki Min. An efficient parallel similarity matrix construction on MapReduce for collaborative filtering. The Journal of Supercomputing, 75(1):123– 141, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:ADA

[1987] Jeong-Hun Kim, Jong-Hyeok Choi, Kwan-Hee Yoo, and Aziz Nasridinov. AA-DBSCAN: an approximate adaptive DBSCAN for finding clusters with varying densities. The Journal of Supercomputing, 75(1):142–169, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2019:PBO

[1988] Ki Yong Lee and Young-Kyoon Suh.

A pattern-based outlier region detection method for two-dimensional arrays. The Journal of Supercomputing, 75(1):170–188, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:DCC

[1989] Minjong Kim and Suyoung Chi. Detection of centerline crossing in abnormal driving using CapsNet. The Journal of Supercomputing, 75(1):189– 196, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Viebke:2019:CPS

[1990] André Viebke, Suejb Memeti, Sabri Pllana, and Ajith Abraham. CHAOS: a parallelization scheme for training convolutional neural networks on Intel Xeon Phi. The Journal of Supercomputing, 75(1):197–227, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
017-1994-x.pdf.

Gavagsaz:2019:LBJ

[1991] Elaheh Gavagsaz, Ali Rezaee, and Hamid Haj Seyyed Javadi. Load balancing in join algorithms for skewed data in MapReduce systems. The Journal of Supercomputing, 75(1):228– 254, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xie:2019:MTN

[1992] Peibo Xie, Huaxi Gu, Kun Wang, Xiaoshan Yu, and Shangqi Ma. Meshof-torus: a new topology for servercentric data center networks. The Journal of Supercomputing, 75(1):255– 271, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fareghzadeh:2019:THP

[1993] Nafiseh Fareghzadeh, Mir Ali Seyyedi, and Mehran Mohsenzadeh. Toward holistic performance management in clouds: taxonomy, challenges and opportunities. The Journal of Supercomputing, 75(1):272–313, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mukherjee:2019:MAT

[1994] Anwesha Mukherjee, Deepsubhra Guha Roy, and Debashis De. Mobility-aware task delegation model in mobile cloud computing. The Journal of Supercomputing, 75(1):314–339, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2019:RUC

[1995] Yan Zhao, Hongwei Liu, Yan Wang, Zhan Zhang, and Decheng Zuo. Reducing the upfront cost of private clouds with clairvoyant virtual machine placement. The Journal of Supercomputing, 75(1):340–369, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Campon:2019:FSE

[1996] J. L. Campon and L. Landesa. Fast solution of electromagnetic scattering problems using Xeon Phi coprocessors. The Journal of Supercomputing, 75(1): 370–383, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hatami:2019:HPA

[1997] Rashid Hatami and Hossein Bahramgiri. High-performance architecture for flow-table lookup in SDN on FPGA. The Journal of Supercomputing, 75(1): 384–399, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lu:2019:PMM

[1998] Huazhong Lü. Paired many-to-many two-disjoint path cover of balanced hypercubes with faulty edges. The Journal of Supercomputing, 75(1):400– 424, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bae:2019:DDT

[1999] Han Jun Bae and Lynn Choi. Dynamic directory table with victim cache: ondemand allocation of directory entries for active shared cache blocks. The Journal of Supercomputing, 75(1):425– 446, January 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kudithi:2019:HPE

[2000] Thirumalesu Kudithi and R. Sakthivel. High-performance ECC processor architecture design for IoT security applications. The Journal of Supercomputing, 75(1):447–474, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Min-Allah:2019:EOS

[2001] Nasro Min-Allah. Effect of ordered set on feasibility analysis of static-priority system. The Journal of Supercomputing, 75(1):475–487, January 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hsu:2019:ABD

[2002] Ching-Hsien Hsu, Geoffrey Fox, Geyong Min, and Sugam Sharma. Advances in big data programming, system software and HPC convergence. The Journal of Supercomputing, 75 (2):489–493, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-018-2706-x.pdf.

Zhang:2019:LDP

[2003] Weizhe Zhang, Yao Hu, Hui He, Yawei Liu, and Allen Chen. Linear and dynamic programming algorithms for real-time task scheduling with task duplication. The Journal of Supercomputing, 75(2):494–509, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yue:2019:DDS

[2004] Shasha Yue, Yan Ma, Lajiao Chen, Yuzhu Wang, and Weijing Song. Dynamic DAG scheduling for manytask computing of distributed ecohydrological model. The Journal of Supercomputing, 75(2):510–532, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nguyen:2019:PBM

[2005] Minh Chau Nguyen, Heesun Won, Siwoon Son, Myeong-Seon Gil, and Yang-Sae Moon. Prefetching-based metadata management in Advanced Multitenant Hadoop. The Journal of Supercomputing, 75(2):533–553, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yu:2019:WWA

[2006] Jie Yu, Guangming Liu, Wenrui Dong, and Xiaoyong Li. WatCache: a workload-aware temporary cache on the compute side of HPC systems. The Journal of Supercomputing, 75(2):554– 586, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sharma:2019:EPS

[2007] Sugam Sharma and Shashi Gadia. Expanding ParaSQL for spatio-temporal (big) data. The Journal of Supercomputing, 75(2):587–606, February 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:BEF

[2008] TaeGuen Kim, Yeo Reum Lee, Boo-Joong Kang, and Eul Gyu Im. Binary executable file similarity calculation using function matching. The Journal of Supercomputing, 75(2):607– 622, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dou:2019:UBP

[2009] Chenxiao Dou, Yi Cui, Daniel Sun, Raymond Wong, Muhammad Atif, Guoqiang Li, and Rajiv Ranjan. Unsupervised blocking and probabilistic parallelisation for record matching of distributed big data. The Journal of Supercomputing, 75(2):623–645, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2019:ACA

[2010] Rui Zhang, Wenguang Chen, Tse-Chuan Hsu, Hongji Yang, and Yeh-Ching Chung. ANG: a combination of apriori and graph computing techniques for frequent itemsets mining. The Journal of Supercomputing, 75(2): 646–661, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hu:2019:ACC

[2011] Cheng Hu and Yuhui Deng. Aggregating correlated cold data to minimize the performance degradation and power consumption of cold storage nodes. The Journal of Supercomputing, 75(2):662–687, February 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ding:2019:APC

[2012] Chuntao Ding and Shangguang Wang.

Appropriate points choosing for subspace learning over image classification. The Journal of Supercomputing, 75(2): 688–703, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jifara:2019:MID

[2013] Worku Jifara, Feng Jiang, Seungmin Rho, Maowei Cheng, and Shaohui Liu. Medical image denoising using convolutional neural network: a residual learning approach. The Journal of Supercomputing, 75(2):704–718, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wen:2019:CLM

[2014] Jianbo Wen, Wei Zhang, and Wanneng Shu. A cognitive learning model in distance education of higher education institutions based on chaos optimization in big data environment. The Journal of Supercomputing, 75(2): 719–731, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jiang:2019:PFG

[2015] Yu Jiang, Minghao Zhao, Chengquan Hu, Lili He, Hongtao Bai, and Jin Wang. A parallel FP-growth algorithm on World Ocean Atlas data with multicore CPU. The Journal of Supercomputing, 75(2):732–745, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rezaeian:2019:FMW

[2016] Amin Rezaeian, Mahmoud Naghibzadeh, and Dick H. J. Epema. Fair multipleworkflow scheduling with different quality-of-service goals. The Journal of Supercomputing, 75(2):746–769, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bo:2019:NFN

[2017] Li Bo, Junrui Lv, Xuegang Luo, Huajun Wang, and Sen Wang. A novel and fast nonlocal means denoising algorithm using a structure tensor. The Journal of Supercomputing, 75(2):770– 782, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fan:2019:EAE

[2018] Weibei Fan, Jianxi Fan, Cheng-Kuan Lin, Guijuan Wang, Baolei Cheng, and Ruchuan Wang. An efficient algorithm for embedding exchanged hypercubes into grids. The Journal of Supercomputing, 75(2):783–807, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bermejo:2019:VCS

[2019] Belen Bermejo, Carlos Juiz, and Carlos Guerrero. Virtualization and consolidation: a systematic review of the past 10 years of research on energy and performance. The Journal of Supercomputing, 75(2):808–836, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yazdanpanah:2019:EEP

[2020] Fahimeh Yazdanpanah, Raheel Afshar-Mazayejani, Mohammad Alaei, Amin Rezaei, and Masoud Daneshtalab. An energy-efficient partition-based XYZplanar routing algorithm for a wireless network-on-chip. The Journal of Supercomputing, 75(2):837–861, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:TLB

[2021] Wei Li, Shuai Ding, Yi Chen, Hao Wang, and Shanlin Yang. Transfer learning-based default prediction model for consumer credit in China. The Journal of Supercomputing, 75(2): 862–884, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ray:2019:TMP

[2022] Benay Kumar Ray, Avirup Saha, Sunirmal Khatua, and Sarbani Roy. Toward maximization of profit and quality of cloud federation: solution to cloud federation formation problem. The Journal of Supercomputing, 75(2): 885–929, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Morinigo:2019:MOC

[2023] José A. Moríñigo, Manuel Rodríguez-Pascual, and Rafael Mayo-García. On the modelling of optimal coordinated checkpoint period in supercomputers. The Journal of Supercomputing, 75(2): 930–954, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Qin:2019:OMT

[2024] Jun Qin, Xuanjing Shen, Fang Mei, and Zheng Fang. An Otsu multithresholds segmentation algorithm based on improved ACO. The Journal of Supercomputing, 75(2):955–967, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bahig:2019:NCT

[2025] Hazem M. Bahig. A new constanttime parallel algorithm for merging. The Journal of Supercomputing, 75(2): 968–983, February 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mwitondi:2019:SDF

[2026] Kassim S. Mwitondi, Farha A. Al-Kuwari, Raed A. Saeed, and Shahrzad Zargari. A statistical downscaling framework for environmental mapping. The Journal of Supercomputing, 75 (2):984–997, February 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-018-2624-y.pdf.

Llanos:2019:CMM

[2027] Diego R. Llanos and Jesús Vigo-Aguiar. Computational and mathematical models meet heterogeneous computing. The Journal of Supercomputing, 75(3):999–1000, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-2713-y.pdf.

Cabanas-Molero:2019:HHR

[2028] P. Cabañas-Molero, Raquel Cortina-Parajón, E. F. Combarro, Pedro Alonso, and F. J. Bris-Peñalver. HRe-MAS: hybrid real-time musical alignment system. The Journal of Supercomputing, 75(3):1001–1013, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bonchis:2019:HOP

[2029] C. Bonchis, E. Kaslik, and F. Rosu.

HPC optimal parallel communication algorithm for the simulation of fractional-order systems. The Journal of Supercomputing, 75(3):1014–1025, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alvarez-Bermejo:2019:EIB

[2030] J. A. Álvarez-Bermejo, D. P. Morales-Santos, E. Castillo-Morales, L. Parrilla, and J. A. López-Ramos. Efficient image-based analysis of fruit surfaces using CCD cameras and smartphones. The Journal of Supercomputing, 75(3): 1026–1037, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Orts:2019:IEE

[2031] F. Orts, E. Filatovas, G. Ortega, O. Kurasova, and E. M. Garzón. Improving the energy efficiency of SMA-COF for multidimensional scaling on modern architectures. The Journal of Supercomputing, 75(3):1038–1050, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alventosa:2019:FBU

[2032] Fran J. Alventosa, Pedro Alonso, Antonio M. Vidal, Gema Piñero, and Enrique S. Quintana-Ortí. Fast block QR update in digital signal processing. The Journal of Supercomputing, 75(3):

1051–1064, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garea:2019:CCB

[2033] Alberto S. Garea, Dora B. Heras, and Francisco Argüello. Caffe CNN-based classification of hyperspectral images on GPU. The Journal of Supercomputing, 75(3):1065–1077, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cocana-Fernandez:2019:IEE

[2034] Alberto Cocaña-Fernández, Julio Rodríguez-Soares, Luciano Sánchez, and José Ranilla. Improving the energy efficiency of virtual data centers in an IT service provider through proactive fuzzy rules-based multicriteria decision making. The Journal of Supercomputing, 75(3):1078–1093, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jimeno-Morenilla:2019:JOA

[2035] A. Jimeno-Morenilla, J. L. Sánchez-Romero, H. Migallón, and H. Mora-Mora. Jaya optimization algorithm with GPU acceleration. The Journal of Supercomputing, 75(3):1094– 1106, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Parrilla:2019:ECC

[2036] Luis Parrilla, José A. Álvarez-Bermejo, Encarnación Castillo, Juan A. López-Ramos, Diego P. Morales-Santos, and Antonio García. Elliptic curve cryptography hardware accelerator for highperformance secure servers. The Journal of Supercomputing, 75(3):1107– 1122, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nozal:2019:LBH

[2037] Raúl Nozal, Borja Perez, Jose Luis Bosque, and Ramón Beivide. Load balancing in a heterogeneous world: CPU– Xeon Phi co-execution of data-parallel kernels. The Journal of Supercomputing, 75(3):1123–1136, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Boratto:2019:EHP

[2038] Murilo Boratto, Pedro Alonso, Clicia Pinto, Pedro Melo, Marcos Barreto, and Spiros Denaxas. Exploring hybrid parallel systems for probabilistic record linkage. The Journal of Supercomputing, 75(3):1137–1149, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Paraschiv:2019:PIM

[2039] Elena Georgiana Paraschiv, Damián Ruiz-Coll, Maria Pantoja, and Gerardo Fernández-Escribano. Parallelization and improvement of the MDV–SW algorithm for HEVC intra-prediction coding. The Journal of Supercomputing, 75(3):1150–1162, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Farguell:2019:SMP

[2040] Angel Farguell, Ana Cortés, Tomàs Margalef, Josep R. Miró, and J. Mercader. Scalability of a multi-physics system for forest fire spread prediction in multi-core platforms. The Journal of Supercomputing, 75(3):1163–1174, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-2330-9.pdf.

Martinovic:2019:HSA

[2041] Tomás Martinovic and Georg Zitzlsberger. Highly scalable algorithm for computation of recurrence quantitative analysis. The Journal of Supercomputing, 75(3):1175–1186, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ferrandez:2019:HPC

[2042] M. R. Ferrández, S. Puertas-Martín, J. L. Redondo, B. Ivorra, A. M. Ramos, and P. M. Ortigosa. High-performance computing for the optimization of highpressure thermal treatments in food industry. The Journal of Supercomputing, 75(3):1187–1202, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia-Lucas:2019:AIM

[2043] David García-Lucas, Gabriel Cebrián-Márquez, Antonio Jesús Díaz-Honrubia, and Pedro Cuenca. Acceleration of the integer motion estimation in JEM through pre-analysis techniques. The Journal of Supercomputing, 75(3):1203–1214, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cebrian-Marquez:2019:HCP

[2044] Gabriel Cebrián-Márquez, Vicente Galiano, Héctor Migallón, José Luis Martínez, Pedro Cuenca, and Otoniel López-Granado. Heterogeneous CPU plus GPU approaches for HEVC. The Journal of Supercomputing, 75(3):

1215–1226, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alonso:2019:CMT

[2045] Pedro Alonso, Jesús Peinado, Javier Ibáñez, Jorge Sastre, and Emilio Defez. Computing matrix trigonometric functions with GPUs through Matlab. The Journal of Supercomputing, 75(3): 1227–1240, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Perez:2019:GBM

[2046] Mariano Pérez, Silvia Rueda, and Juan M. Orduña. Geometry-based methods for general non-planar perspective projections on curved displays. The Journal of Supercomputing, 75(3): 1241–1255, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ramiro:2019:PSS

[2047] Carla Ramiro, M. Ángeles Simarro, Alberto Gonzalez, and Antonio M. Vidal. Parallel SUMIS soft detector for large MIMO systems on multicore and GPU. The Journal of Supercomputing, 75(3): 1256–1267, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cruz:2019:DPG

[2048] N. C. Cruz, S. Salhi, J. L. Redondo, J. D. Álvarez, M. Berenguel, and P. M. Ortigosa. Design of a parallel genetic algorithm for continuous and pattern-free heliostat field optimization. The Journal of Supercomputing, 75(3):1268–1283, March 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Badia:2019:ASP

[2049] Jose M. Badía, Jose A. Belloch, Maximo Cobos, Francisco D. Igual, and Enrique S. Quintana-Ortí. Accelerating the SRP–PHAT algorithm on multi- and many-core platforms using OpenCL. The Journal of Supercomputing, 75(3):1284–1297, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sebastian:2019:AEP

[2050] Pablo San Juan Sebastián, Tuomas Virtanen, Victor M. Garcia-Molla, and Antonio M. Vidal. Analysis of an efficient parallel implementation of activeset Newton algorithm. The Journal of Supercomputing, 75(3):1298– 1309, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Llados:2019:OCB

[2051] Jordi Lladós, Fernando Cores, and Fernando Guirado. Optimization of consistency-based multiple sequence alignment using Big Data technologies. The Journal of Supercomputing, 75(3): 1310–1322, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Leon:2019:NEH

[2052] Germán León, Carlos González, Rafael Mayo, Daniel Mozos, and Enrique S. Quintana-Ortí. Noise estimation for hyperspectral subspace identification on FPGAs. The Journal of Supercomputing, 75(3):1323–1335, March 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mora:2019:HPA

[2053] H. Mora, M. T. Signes-Pont, A. Jimeno-Morenilla, and J. L. Sánchez-Romero. High-performance architecture for digital transform processing. The Journal of Supercomputing, 75(3):1336–1349, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Benito-Picazo:2019:MGA

[2054] F. Benito-Picazo, P. Cordero, M. Enciso, and A. Mora. Minimal generators, an affordable approach by means of massive computation. The Journal of Supercomputing, 75(3):1350– 1367, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cano:2019:PSM

[2055] José-Carlos Cano, Javier Cuenca, Domingo Giménez, Mariano Saura-Sánchez, and Pablo Segado-Cabezos. A parallel simulator for multibody systems based on group equations. The Journal of Supercomputing, 75(3): 1368–1381, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Afonso:2019:HPC

[2056] Sergio Afonso, Alejandro Acosta, and Francisco Almeida. High-performance code optimizations for mobile devices. The Journal of Supercomputing, 75(3): 1382–1395, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Whitney:2019:DEC

[2057] James Whitney, Chandler Gifford, and Maria Pantoja. Distributed execution of communicating sequential processstyle concurrency: Golang case study. The Journal of Supercomputing, 75(3): 1396–1409, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia:2019:PCC

[2058] Juan F. García and M. V. Carriegos. On parallel computation of centrality measures of graphs. The Journal of Supercomputing, 75(3):1410–1428, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lopez-Granado:2019:HSP

[2059] O. López-Granado, H. Migallón, M. Martínez-Rach, V. Galiano, M. P. Malumbres, and Glenn Van Wallendael. A highly scalable parallel encoder version of the emergent JEM video encoder. The Journal of Supercomputing, 75(3):1429–1442, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tomcala:2019:ATS

[2060] Jirí Tomcala. Acceleration of time series entropy algorithms. The Journal of Supercomputing, 75(3):1443– 1454, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aliaga:2019:EGV

[2061] José I. Aliaga, Ernesto Dufrechou, Pablo Ezzatti, and Enrique S. Quintana-Ortí. An efficient GPU version of the preconditioned GMRES method. The Journal of Supercomputing, 75(3):1455–1469, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dolz:2019:PSB

[2062] Manuel F. Dolz, Fran J. Alventosa, Pedro Alonso-Jordá, and Antonio M. Vidal. A pipeline structure for the block QR update in digital signal processing. The Journal of Supercomputing, 75(3): 1470–1482, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vila:2019:ESS

[2063] Sergi Vila, Fernando Guirado, Josep L.

Lerida, and Fernando Cores. Energysaving scheduling on IaaS HPC cloud environments based on a multiobjective genetic algorithm. The Journal of Supercomputing, 75(3):1483– 1495, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fernandez:2019:VDM

[2064] Lisardo Fernández, Mariano Pérez, and Juan M. Orduña. Visualization of DNA methylation results through a GPUbased parallelization of the wavelet transform. The Journal of Supercomputing, 75(3):1496–1509, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dieguez:2019:PPO

[2065] Adrián P. Diéguez, Margarita Amor, and Ramón Doallo. Parallel prefix operations on GPU: tridiagonal system solvers and scan operators. The Journal of Supercomputing, 75(3):

1510–1523, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moreno:2019:DBP

[2066] Andreu Moreno, Juan J. Rodríguez, Daniel Beltrán, Anna Sikora, Josep Jorba, and Eduardo César. Designing a benchmark for the performance evaluation of agent-based simulation applications on HPC. The Journal of Supercomputing, 75(3):1524–1550, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2688-8.pdf.

Fernandez-Fabeiro:2019:MDV

[2067] Jorge Fernández-Fabeiro, Álvaro Ordóñez, Arturo Gonzalez-Escribano, and Dora B. Heras. A multi-device version of the HYFMGPU algorithm for hyperspectral scenes registration. The Journal of Supercomputing, 75(3):1551–1564, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bascoy:2019:EAP

[2068] Pedro G. Bascoy, Pablo Quesada-Barriuso, Dora B. Heras, Francisco Argüello, Begüm Demir, and Lorenzo Bruzzone. Extended attribute profiles on GPU applied to hyperspectral image classification. The Journal of Supercomputing, 75(3):1565–1579, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ortega:2019:CAC

[2069] G. Ortega, E. M. T. Hendrix, and I. García. A CUDA approach to compute perishable inventory control policies using value iteration. The Journal of Supercomputing, 75(3):1580–1593, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2692-z.pdf.

Munoz-Montoro:2019:RTS

[2070] A. J. Muñoz-Montoro, J. Ranilla, P. Vera-Candeas, E. F. Combarro, and P. Alonso-Jordá. Real-time Soundprism. The Journal of Supercomputing, 75(3):1594–1609, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cabrera:2019:HTI

[2071] Alberto Cabrera, Alejandro Acosta, Francisco Almeida, and Vicente Blanco. A heuristic technique to improve energy efficiency with dynamic load balancing. The Journal of Supercomputing, 75(3):1610–1624, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bernabe:2019:SOS

[2072] Gregorio Bernabé, José D. Casanova, Javier Cuenca, and Josefa González-Carrillo. A self-optimized software tool for quantifying the degree of left ventricle hyper-trabeculation. The Journal of Supercomputing, 75(3):1625– 1640, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jimenez-Ruiz:2019:WFS

[2073] Alberto Jiménez-Ruiz, Miguel Cañas-Carretón, Gerardo Fernández-Escribano, Damián Ruiz-Coll, Sergio Martín-Martínez, and Emilio Gómez-Lázaro. Wind farm simulations based on a DFIG machine using parallel programming. The Journal of Supercomputing, 75(3):1641–1653, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rico-Gallego:2019:ACP

[2074] Juan A. Rico-Gallego, Juan C. Díaz-Martín, Carmen Calvo-Jurado, Sergio Moreno-Álvarez, and Juan L. García-Zapata. Analytical communication performance models as a metric in the partitioning of data-parallel kernels on heterogeneous platforms. The Journal of Supercomputing, 75(3):1654– 1669, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bermudez:2019:OLI

[2075] Aurelio Bermúdez, Francisco Montero, María T. López, Antonio Fernández-Caballero, and José L. Sánchez. Optimization of lateral interaction in accumulative computation on GPU-based platform. The Journal of Supercomputing, 75(3):1670–1685, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Belloch:2019:UMC

[2076] Jose A. Belloch, Adrian Amor-Martin, Daniel Garcia-Donoro, Francisco J. Martínez-Zaldívar, and Luis E. Garcia-Castillo. On the use of many-core machines for the acceleration of a mesh truncation technique for FEM. The Journal of Supercomputing, 75(3): 1686–1696, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Migallon:2019:MBM

[2077] H. Migallón, A. Jimeno-Morenilla, J. L.

Sánchez-Romero, H. Rico, and R. V. Rao. Multipopulation-based multilevel parallel enhanced Jaya algorithms. The Journal of Supercomputing, 75(3):1697–1716, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rey:2019:VIT

[2078] Antón Rey, Francisco D. Igual, and Manuel Prieto-Matías. Variable intratask threading for power-constrained performance and energy optimization in DAG scheduling. The Journal of Supercomputing, 75(3):1717–1731, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Guzman:2019:CCG

[2079] María Angélica Dávila Guzmán, Raúl Nozal, Rubén Gran Tejero, María Villarroya-Gaudó, Darío Suárez Gracia, and Jose Luis Bosque. Cooperative CPU, GPU, and FPGA heterogeneous execution with EngineCL. The Journal of Supercomputing, 75(3): 1732–1746, March 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shin:2019:HCC

[2080] Byeong-Seok Shin, Houcine Hassan, and Qun Jin. Human-computer cooperation for future computing. The Journal of Supercomputing, 75(4): 1747–1750, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-018-2661-6.pdf.

Kim:2019:EVQ

[2081] Hyoungseok Kim and Joonseok Park. Efficient video quality assessment for on-demand video transcoding using intensity variation analysis. The Journal of Supercomputing, 75(4):1751–1765, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ma:2019:OMA

[2082] Xixi Ma, Qun Jin, Julong Pan, and Yufeng Wang. Optimization modeling and analysis of trustworthiness determination strategies for service discovery of MSNP. The Journal of Supercomputing, 75(4):1766–1782, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Stojmenovic:2019:ITV

[2083] Milica Stojmenovic, Robert Biddle, John Grundy, and Vivienne Farrell. The influence of textual and verbal word-of-mouth on website usability and visual appeal. The Journal of Supercomputing, 75(4):1783–1830, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Huh:2019:LBM

[2084] Jun-Ho Huh and Tae-Jung Kim. A location-based mobile health care facility search system for senior citizens. The Journal of Supercomputing, 75(4): 1831–1848, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dominguez:2019:HCC

[2085] Carlos Domínguez, Juan-Miguel Martínez, Jose V. Busquets-Mataix, and Houcine Hassan. Human-computer cooperation platform for developing real-time robotic applications. The Journal of Supercomputing, 75(4):1849–1868, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia:2019:SMW

[2086] Laura García, Lorena Parra, Oscar Romero, and Jaime Lloret. System for monitoring the wellness state of people in domestic environments employing emoticon-based HCI. The Journal of Supercomputing, 75(4):1869–1893, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2019:OSI

[2087] Hui Zhao, Meikang Qiu, Keke Gai, and Xin He. Optimal solution to intelligent multi-channel wireless communications using dynamic programming. The Journal of Supercomputing, 75(4): 1894–1908, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chae:2019:GAB

[2088] Jeongsook Chae, Yong Jin, Mingyun Wen, Weiqiang Zhang, Yunsick Sung, and Kyungeun Cho. Genetic algorithm-based adaptive weight decision method for motion estimation framework. The Journal of Supercomputing, 75(4):1909–1921, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2019:MSM

[2089] Jin Liu, Chenkai Gu, Jin Wang, Geumran Youn, and Jeong-Uk Kim. Multiscale multi-class conditional generative adversarial network for handwritten character generation. The Journal of Supercomputing, 75(4):1922–1940, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2019:IPI

[2090] YoungKyung Lee, Jinho Song, and Yoojae Won. Improving personal information detection using OCR feature recognition rate. The Journal of Supercomputing, 75(4):1941–1952, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hao:2019:LSC

[2091] Fei Hao, Doo-Soon Park, Xiaoyan Yin, Xiaoming Wang, and Vilakone Phonexay. A location-sensitive overthe-counter medicines recommender based on tensor decomposition. The Journal of Supercomputing, 75(4): 1953–1970, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:AIM

[2092] Kyungrog Kim and Nammee Moon. Activity index model for self-regulated learning with learning analysis in a TEL environment. The Journal of Supercomputing, 75(4):1971–1989, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Song:2019:ETS

[2093] Hyejin Song and Nammee Moon. Eye-tracking and social behavior preference-based recommendation system. The Journal of Supercomputing, 75(4):1990–2006, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2019:DBA

[2094] Mu-Song Chen, Chi-Pan Hwang, Tze-Yee Ho, Hsuan-Fu Wang, Chih-Min Shih, Hsing-Yu Chen, and Wen Kai Liu. Driving behaviors analysis based on feature selection and statistical approach: a preliminary study. The Journal of Supercomputing, 75(4):2007– 2026, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:DCT

[2095] Doohwan Kim, Seungwoo Nam, and Jang-Eui Hong. A dynamic control technique to enhance the flexibility of software artifact reuse in large-scale repository. The Journal of Supercomputing, 75(4):2027–2057, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

An:2019:DIA

[2096] Yongli An, Ruihua Sun, Xinwen Wu, and Xiaochuan Sun. Distributed interference alignment algorithm in downlink multi-user cooperative networks. The Journal of Supercomputing, 75(4): 2058–2069, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Singh:2019:FCA

[2097] Simar Preet Singh, Anand Nayyar, Rajesh Kumar, and Anju Sharma. Fog computing: from architecture to edge computing and big data processing.

The Journal of Supercomputing, 75(4): 2070–2105, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hashemi:2019:DAE

[2098] Sara Hashemi, Mostafa Rahimi Azghadi, and Keivan Navi. Design and analysis of efficient QCA reversible adders. The Journal of Supercomputing, 75(4):2106–2125, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sayadnavard:2019:REA

[2099] Monireh H. Sayadnavard, Abolfazl Toroghi Haghighat, and Amir Masoud Rahmani. A reliable energy-aware approach for dynamic virtual machine consolidation in cloud data centers. The Journal of Supercomputing, 75(4): 2126–2147, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See [2100].

Sayadnavard:2019:CRE

[2100] Monireh H. Sayadnavard, Abolfazl Toroghi Haghighat, and Amir Masoud Rahmani. Correction to: A reliable energy-aware approach for dynamic virtual machine consolidation in cloud data centers. The Journal of Supercomputing, 75(4):2148, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-02733-1.pdf. See [2099].

Mishra:2019:IRR

[2101] Ashish Kumar Mishra, Dharmendra K. Yadav, Yogesh Kumar, and Naman Jain. Improving reliability and reducing cost of task execution on preemptible VM instances using machine learning approach. The Journal of Supercomputing, 75(4):2149–2180, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Maharajan:2019:MCI

[2102] K. Maharajan and B. Paramasivan.

Membrane computing inspired protocol to enhance security in cloud network. The Journal of Supercomputing, 75(4): 2181–2192, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sajithabanu:2019:DPP

[2103] S. Sajithabanu and S. R. Balasundaram. Direct push-pull or assisted push-pull? Toward optimal video content delivery using shared storage-based cloud CDN (SS-CCDN). The Journal of Supercomputing, 75(4): 2193–2220, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khan:2019:CSA

[2104] Farrukh Aslam Khan, Ashfaq Hussain Farooqi, and Abdelouahid Derhab. A comprehensive security analysis of LEACH++ clustering protocol for wireless sensor networks. The Journal of Supercomputing, 75(4):2221–2242, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kaity:2019:ANE

[2105] Mohammed Kaity and Vimala Balakrishnan. An automatic non-English sentiment lexicon builder using unannotated corpus. The Journal of Supercomputing, 75(4):2243–2268, April 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmad:2019:PQE

[2106] Awais Ahmad, Mudassar Ahmad, Muhammad Asif Habib, Shahzad Sarwar, Junaid Chaudhry, Muhammad Ahsan Latif, Saadat Hanif Dar, and Muhammad Shahid. Parallel query execution over encrypted data in database-as-a-service (DaaS). The Journal of Supercomputing, 75(4): 2269–2288, April 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gao:2019:SME

[2107] Song Gao, Xinwu Du, and Jumei Yue. Study on measurement error of iron ore pipeline transportation flow based on weight function theory of electromagnetic flow sensor. The Journal of Supercomputing, 75(5):2289–2303, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fan:2019:JCR

[2108] Xincan Fan, Kaiyang Liu, and Haibo Yi. Joint collaborative representation algorithm for face recognition. The Journal of Supercomputing, 75(5): 2304–2314, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Azarnia:2019:GCD

[2109] Ghanbar Azarnia, Mohammad Ali Tinati, and Tohid Yousefi Rezaii. Generic cooperative and distributed algorithm for recovery of signals with the same sparsity profile in wireless sensor networks: a non-convex approach. The Journal of Supercomputing, 75(5): 2315–2340, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sharma:2019:HCW

[2110] Deepak Sharma, Amritesh Ojha, and Amol P. Bhondekar. Heterogeneity consideration in wireless sensor networks routing algorithms: a review. The Journal of Supercomputing, 75(5): 2341–2394, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shehab:2019:HCS

[2111] Mohammad Shehab, Ahamad Tajudin Khader, Makhlouf Laouchedi, and Osama Ahmad Alomari. Hybridizing cuckoo search algorithm with bat algorithm for global numerical optimization. The Journal of Supercomputing, 75(5):2395–2422, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cai:2019:GTB

[2112] Xibiao Cai, Xiaoling Liu, and Zhi Qu. Game theory-based device-to-device network access algorithm for heterogeneous networks. The Journal of Supercomputing, 75(5):2423–2435, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Owahid:2019:WDP

[2113] Abdullah A. Owahid and Eugene B.

John. Wasted dynamic power and correlation to instruction set architecture for CPU throttling. The Journal of Supercomputing, 75(5):2436–2454, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Meshkati:2019:EAR

[2114] Jafar Meshkati and Faramarz Safi-Esfahani. Energy-aware resource utilization based on particle swarm optimization and artificial bee colony algorithms in cloud computing. The Journal of Supercomputing, 75(5):2455– 2496, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cheng:2019:FFE

[2115] Feng Cheng and Zhe Yang. FastMFDs: a fast, efficient algorithm for mining minimal functional dependencies from large-scale distributed data with Spark. The Journal of Supercomputing, 75(5): 2497–2517, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karimi:2019:OWD

[2116] Vahid Karimi, Reza Mohseni, and Sadegh Samadi. OFDM waveform design based on mutual information for cognitive radar applications. The Journal of Supercomputing, 75(5):2518– 2534, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Siavashi:2019:GEC

[2117] Ahmad Siavashi and Mahmoud Momtazpour. GPUCloudSim: an extension of CloudSim for modeling and simulation of GPUs in cloud data centers. The Journal of Supercomputing, 75(5): 2535–2561, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nguyen:2019:GPM

[2118] M. T. Nguyen, P. Castonguay, and E. Laurendeau. GPU parallelization of multigrid RANS solver for threedimensional aerodynamic simulations on multiblock grids. The Journal of Supercomputing, 75(5):2562–2583, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nawaz:2019:SAD

[2119] Asif Nawaz, Sohail Asghar, and Syed Husnain Abbas Naqvi. A segregational approach for determining aspect sentiments in social media analysis. The Journal of Supercomputing, 75(5): 2584–2602, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ghobaei-Arani:2019:LWL

[2120] Mostafa Ghobaei-Arani and Alireza Souri. LP-WSC: a linear programming approach for web service composition in geographically distributed cloud environments. The Journal of Supercomputing, 75(5):2603–2628, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2019:SSE

[2121] Jia Wu, Zhigang Chen, and Ming Zhao.

SECM: status estimation and cache management algorithm in opportunistic networks. The Journal of Supercomputing, 75(5):2629–2647, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Imanparast:2019:LTR

[2122] Mahdi Imanparast and Seyed Naser Hashemi. A linear time randomized approximation algorithm for Euclidean matching. The Journal of Supercomputing, 75(5):2648–2664, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bielecki:2019:ITG

[2123] Wlodzimierz Bielecki and Piotr Skotnicki. Insight into tiles generated by means of a correction technique. The Journal of Supercomputing, 75 (5):2665–2690, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-018-2678-x.pdf.

Poshtkohi:2019:PPS

[2124] Alireza Poshtkohi, M. B. Ghaznavi-Ghoushchi, and Kamyar Saghafi. PSML: parallel system modeling and simulation language for electronic system level. The Journal of Supercomputing, 75(5):2691–2724, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dokulil:2019:CMR

[2125] Jiri Dokulil. Consistency model for runtime objects in the Open Community Runtime. The Journal of Supercomputing, 75(5):2725–2760, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-2681-2.pdf.

Wang:2019:TES

[2126] Tianshu Wang, Kongfa Hu, Xichen Yang, Gongxuan Zhang, and Yongli Wang. A trust enhancement scheme for cluster-based wireless sensor networks. The Journal of Supercomputing, 75(5): 2761–2788, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vivek:2019:PFF

[2127] V. Vivek, R. Srinivasan, R. Elijah Blessing, and R. Dhanasekaran. Payload fragmentation framework for highperformance computing in cloud environment. The Journal of Supercomputing, 75(5):2789–2804, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:RAT

[2128] Chunlin Li, Jing Zhang, and Hengliang Tang. Replica-aware task scheduling and load balanced cache placement for delay reduction in multi-cloud environment. The Journal of Supercomputing, 75(5):2805–2836, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Asadi:2019:EIF

[2129] Ali Naghash Asadi, Mohammad Abdollahi Azgomi, and Reza Entezari-Maleki. Evaluation of the impacts of failures and resource heterogeneity on the power consumption and performance of IaaS clouds. The Journal of Supercomputing, 75(5):2837–2861, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Patil:2019:DEF

[2130] Rajendra Patil and Chirag Modi. Designing an efficient framework for vulnerability assessment and patching (VAP) in virtual environment of cloud computing. The Journal of Supercomputing, 75(5):2862–2889, May 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:ACE

[2131] Chunlin Li and Jingpan Bai. Automatic content extraction and timeaware topic clustering for large-scale social network on cloud platform. The Journal of Supercomputing, 75(5): 2890–2924, May 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jeong:2019:MLB

[2132] Young-Sik Jeong, Houcine Hassan, and Arun Kumar Sangaiah. Machine learning on big data for future computing. The Journal of Supercomputing, 75(6):2925–2929, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-019-02872-z.pdf.

Peng:2019:ADG

[2133] Lizhi Peng, Haibo Zhang, Houcine Hassan, Yuehui Chen, and Bo Yang. Accelerating data gravitation-based classification using GPU. The Journal of Supercomputing, 75(6):2930–2949, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hur:2019:EBP

[2134] Cheonghwan Hur and Sanggil Kang. Entropy-based pruning method for convolutional neural networks. The Journal of Supercomputing, 75(6): 2950–2963, June 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2019:MDO

[2135] Huipeng Chen, Xuanwei Chen, Chang Chen, Zhangming Peng, Youping Gong, and Guojin Chen. Multidisciplinary design optimization for vehicle handling stability of steering-by-wire system. The Journal of Supercomputing, 75(6):2964–2985, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Panwar:2019:YED

[2136] Sharaj Panwar, Paul Rad, Kim-Kwang Raymond Choo, and Mehdi Roopaei. Are you emotional or depressed? Learning about your emotional state from your music using machine learning. The Journal of Supercomputing, 75(6):2986–3009, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lin:2019:IIM

[2137] Yun Lin, Xiaolei Zhu, Zhigao Zheng, Zheng Dou, and Ruolin Zhou. The individual identification method of wireless device based on dimensionality reduction and machine learning. The Journal of Supercomputing, 75(6):3010–3027, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2019:SAI

[2138] Jooyoung Lee, Daesung Moon, Ikkyun Kim, and Youngseok Lee. A semantic approach to improving machine readability of a large-scale attack graph. The Journal of Supercomputing, 75(6): 3028–3045, June 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kaur:2019:CCB

[2139] Harmanjeet Kaur, Neeraj Kumar, and Shalini Batra. ClaMPP: a cloud-based multi-party privacy preserving classification scheme for distributed applications. The Journal of Supercomputing, 75(6):3046–3075, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2019:SCC

[2140] Pankaj Kumar, Saru Kumari, Vishnu Sharma, Xiong Li, Arun Kumar Sangaiah, and SK Hafizul Islam. Secure CLS and CL–AS schemes designed for VANETs. The Journal of Supercomputing, 75(6):3076–3098, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hu:2019:CAC

[2141] Chengyu Hu, Rupeng Yang, Pengtao Liu, Tong Li, and Fanyu Kong. A countermeasure against cryptographic key leakage in cloud: public-key encryption with continuous leakage and tampering resilience. The Journal of Supercomputing, 75(6):3099–3122, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Huh:2019:BBM

[2142] Jun-Ho Huh and Kyungryong Seo. Blockchain-based mobile fingerprint verification and automatic log-in platform for future computing. The Journal of Supercomputing, 75(6):3123– 3139, June 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ai:2019:SMT

[2143] Yi Ai, Zongping Li, and Mi Gan. A solution to measure traveler’s transfer tolerance for walking mode and dockless bike-sharing mode. The Journal of Supercomputing, 75(6):3140–3157, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Muneeswaran:2019:ASG

[2144] V. Muneeswaran and M. Pallikonda Rajasekaran. Automatic segmentation of gallbladder using bio-inspired algorithm based on a spider web construction model. The Journal of Supercomputing, 75(6):3158–3183, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Subramaniyaswamy:2019:ODP

[2145] V. Subramaniyaswamy, Gunasekaran Manogaran, R. Logesh, V. Vijayakumar, Naveen Chilamkurti, D. Malathi, and N. Senthilselvan. An ontologydriven personalized food recommendation in IoT-based healthcare system. The Journal of Supercomputing, 75(6): 3184–3216, June 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2019:OTM

[2146] Weiqiang Zhang, Seoungjae Cho, Jeongsook Chae, Yunsick Sung, and Kyungeun Cho. Object tracking method based on data computing. The Journal of Supercomputing, 75(6):

3217–3228, June 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cho:2019:RTR

[2147] Seoungjae Cho and Kyungeun Cho. Real-time 3D reconstruction method using massive multi-sensor data analysis and fusion. The Journal of Supercomputing, 75(6):3229–3248, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fu:2019:HMB

[2148] Shuang Fu, Guoyin Zhang, and Takeo Fujii. A heuristic method-based parallel cooperative spectrum sensing in heterogeneous network. The Journal of Supercomputing, 75(6):3249–3263, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bedi:2019:MES

[2149] Rajeev Kumar Bedi, Jaswinder Singh, and Sunil Kumar Gupta. MWC: an efficient and secure multi-cloud storage approach to leverage augmentation of multi-cloud storage services on mobile devices using fog computing. The Journal of Supercomputing, 75(6): 3264–3287, June 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:SPM

[2150] Juyoung Kim, Kwantae Cho, Yong-Kyun Kim, Kyung-Soo Lim, and Sang Uk Shin. Study on peak misdetection recovery of key exchange protocol using heartbeat. The Journal of Supercomputing, 75(6):3288–3301, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Diwakaran:2019:CPM

[2151] S. Diwakaran, B. Perumal, and K. Vimala Devi. A cluster prediction model-based data collection for energy efficient wireless sensor network. The Journal of Supercomputing, 75(6): 3302–3316, June 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sun:2019:IMD

[2152] Ruxia Sun, Lingfeng Shi, Chunyong Yin, and Jin Wang. An improved method in deep packet inspection based on regular expression. The Journal of Supercomputing, 75(6):3317–3333, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:BDF

[2153] Ha-Kyun Kim, Won-Hyun So, and Seung-Mo Je. A big data framework for network security of small and medium enterprises for future computing. The Journal of Supercomputing, 75(6):3334–3367, June 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vega-Rodriguez:2019:PCB

[2154] Miguel A. Vega-Rodríguez and Sergio Santander-Jiménez. Parallel computing in bioinformatics: a view from highperformance, heterogeneous, and cloud computing. The Journal of Supercomputing, 75(7):3369–3373, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vitali:2019:EOO

[2155] Emanuele Vitali, Davide Gadioli, Gianluca Palermo, Andrea Beccari, Carlo Cavazzoni, and Cristina Silvano. Exploiting OpenMP and OpenACC to accelerate a geometric approach to molecular docking in heterogeneous HPC nodes. The Journal of Supercomputing, 75(7):3374–3396, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Escobar:2019:TEA

[2156] Juan José Escobar, Julio Ortega, Antonio F. Díaz, Jesús González, and Miguel Damas. Time-energy analysis of multilevel parallelism in heterogeneous clusters: the case of EEG classification in BCI tasks. The Journal of Supercomputing, 75(7):3397–3425, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Daberdaku:2019:ACT

[2157] Sebastian Daberdaku. Accelerating the computation of triangulated molecular surfaces with OpenMP. The Journal of Supercomputing, 75(7):3426–3470, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gonzalez:2019:HPM

[2158] Patricia González, Pablo Argüeso-Alejandro, David R. Penas, Xoan C. Pardo, Julio Saez-Rodriguez, Julio R. Banga, and Ramón Doallo. Hybrid parallel multimethod hyperheuristic for mixed-integer dynamic optimization problems in computational systems biology. The Journal of Supercomputing, 75(7):3471–3498, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ferretti:2019:OCB

[2159] Marco Ferretti, Luigi Santangelo, and Mirto Musci. Optimized cloud-based scheduling for protein secondary structure analysis. The Journal of Supercomputing, 75(7):3499–3520, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [2160].

Ferretti:2019:COC

[2160] Marco Ferretti, Luigi Santangelo, and Mirto Musci. Correction to: Optimized cloud-based scheduling for protein secondary structure analysis. The Journal of Supercomputing, 75(7):3521, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02931-5.pdf. See [2159].

Nandini:2019:ESD

[2161] D. Usha Nandini and Ezil Sam Leni.

Efficient shadow detection by using PSO segmentation and regionbased boundary detection technique. The Journal of Supercomputing, 75(7): 3522–3533, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kurdi:2019:LTM

[2162] Heba Kurdi, Auhood Alfaries, Abeer Al-Anazi, Sara Alkharji, Maimona Addegaither, Lina Altoaimy, and Syed Hassan Ahmed. A lightweight trust management algorithm based on subjective logic for interconnected cloud computing environments. The Journal of Supercomputing, 75(7): 3534–3554, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hashemi:2019:DCT

[2163] Seyyed Yasser Hashemi and Fereidoon Shams Aliee. Dynamic and comprehensive trust model for IoT and its integration into RPL. The Journal of Supercomputing, 75(7):3555–3584, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Javan:2019:SFD

[2164] Morteza Sargolzaei Javan and Mohammad Kazem Akbari. SmartData 4.0: a formal description framework for big data. The Journal of Supercomputing, 75(7):3585–3620, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:DSC

[2165] Jong-Seok Kim, Donghyun Kim, Ke Qiu, and Hyeong-Ok Lee. The divide-and-swap cube: a new hypercube variant with small network cost. The Journal of Supercomputing, 75(7): 3621–3639, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gutierrez-Aguado:2019:TTE

[2166] Juan Gutiérrez-Aguado, Jose M. Claver, and Raúl Peña-Ortiz. Toward a transparent and efficient GPU cloudification architecture. The Journal of Supercomputing, 75(7):3640–3672, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ji:2019:LLC

[2167] Shuo Ji, Yinliang Zhao, and Xiaomei Zhao. A low-latency computing framework for time-evolving graphs. The Journal of Supercomputing, 75(7): 3673–3692, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khan:2019:TEX

[2168] Minhaj Ahmad Khan. Towards efficient XML parsing through minimization of JVM parameter space. The Journal of Supercomputing, 75(7): 3693–3711, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Singh:2019:MLC

[2169] Harmanpreet Singh and Damanpreet Singh. Multi-level clustering protocol for load-balanced and scalable clustering in large-scale wireless sensor networks. The Journal of Supercomputing, 75(7):3712–3739, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ijaz:2019:MLB

[2170] Samia Ijaz and Ehsan Ullah Munir.

MOPT: list-based heuristic for scheduling workflows in cloud environment. The Journal of Supercomputing, 75(7): 3740–3768, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Salehan:2019:OCA

[2171] Alireza Salehan, Hossein Deldari, and Saeid Abrishami. An online contextaware mechanism for computation offloading in ubiquitous and mobile cloud environments. The Journal of Supercomputing, 75(7):3769–3809, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2019:HLN

[2172] Shusen Wu, Xiaoshe Dong, Xingjun Zhang, and Zhengdong Zhu. NoT: a high-level no-threading parallel programming method for heterogeneous systems. The Journal of Supercomputing, 75(7):3810–3841, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nashaat:2019:SES

[2173] Heba Nashaat, Nesma Ashry, and Rawya Rizk. Smart elastic scheduling algorithm for virtual machine migration in cloud computing. The Journal of Supercomputing, 75(7):3842–3865, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ciglaric:2019:OLP

[2174] Tadej Ciglaric, Rok Cesnovar, and Erik Strumbelj. An OpenCL library for parallel random number generators. The Journal of Supercomputing, 75(7): 3866–3881, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kamalinasab:2019:CGC

[2175] Shokooh Kamalinasab, Faramarz Safi-Esfahani, and Majid Shahbazi. CRFF.GP: cloud runtime formulation framework based on genetic programming. The Journal of Supercomputing, 75(7):3882–3916, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhou:2019:MLO

[2176] Bei Zhou, Yongzhong Huang, Jinchen Xu, Shaozhong Guo, and Hongyuan Qi.

Memory latency optimizations for the elementary functions on the Sunway architecture. The Journal of Supercomputing, 75(7):3917–3944, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ghaemi:2019:SLE

[2177] Seyedeh Golsana Ghaemi, Iman Ahmadpour, Mehdi Ardebili, and Hamed Farbeh. Sleepy-LRU: extending the lifetime of non-volatile caches by reducing activity of age bits. The Journal of Supercomputing, 75(7):3945–3974, July 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ullah:2019:NDA

[2178] Ihsan Ullah and Hee Yong Youn.

A novel data aggregation scheme based on self-organized map for WSN. The Journal of Supercomputing, 75(7): 3975–3996, July 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Garcia:2019:RSH

[2179] J. Daniel Garcia. Refactoring software to heterogeneous parallel platforms. The Journal of Supercomputing, 75 (8):3997–4000, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-019-02942-2.pdf.

Ban:2019:PMP

[2180] Dénes Bán, Rudolf Ferenc, István Siket, Ákos Kiss, and Tibor Gyimóthy. Prediction models for performance, power, and energy efficiency of software executed on heterogeneous hardware. The Journal of Supercomputing, 75(8):

4001–4025, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gazzarri:2019:SSP

[2181] Leonardo Gazzarri and Marco Danelutto. Supporting structured parallel program design, development and tuning in FastFlow. The Journal of Supercomputing, 75(8):4026–4041, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Griebler:2019:SPO

[2182] Dalvan Griebler, Renato B. Hoffmann, Marco Danelutto, and Luiz G. Fernandes. Stream parallelism with ordered data constraints on multi-core systems. The Journal of Supercomputing, 75(8): 4042–4061, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Oppermann:2019:SAK

[2183] Julian Oppermann, Lukas Sommer, and Andreas Koch. SpExSim: assessing kernel suitability for C-based highlevel hardware synthesis. The Journal of Supercomputing, 75(8):4062– 4077, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nunez-Yanez:2019:SMS

[2184] Jose Nunez-Yanez, Sam Amiri, Mohammad Hosseinabady, Andrés Rodríguez, Rafael Asenjo, Angeles Navarro, Dario Suarez, and Ruben Gran. Simultaneous multiprocessing in a software-defined heterogeneous FPGA. The Journal of Supercomputing, 75(8):4078–4095, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2367-9.pdf.  See correction

[2185].

Nunez-Yanez:2019:CSM

[2185] Jose Nunez-Yanez, Sam Amiri, Mohammad Hosseinabady, Andrés Rodríguez, Rafael Asenjo, Angeles Navarro, Dario Suarez, and Ruben Gran. Correction to: Simultaneous multiprocessing in a software-defined heterogeneous FPGA. The Journal of Supercomputing, 75(8):4096–4097, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
018-2409-3.pdf. See [2184].

Astorga:2019:HSD

[2186] D. del Rio Astorga, Manuel F. Dolz, Javier Fernandez, and Javier Garcia Blas. Hybrid static-dynamic selection of implementation alternatives in heterogeneous environments. The Journal of Supercomputing, 75(8):4098– 4113, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Torquati:2019:DMA

[2187] M. Torquati, G. Mencagli, M. Drocco, M. Aldinucci, T. De Matteis, and M. Danelutto. On dynamic memory allocation in sliding-window parallel patterns for streaming analytics. The Journal of Supercomputing, 75(8): 4114–4131, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sang:2019:EIP

[2188] Janche Sang, Che-Rung Lee, Vernon Rego, and Chung-Ta King. Experiences with implementing parallel discreteevent simulation on GPU. The Journal of Supercomputing, 75(8):4132– 4149, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sinaei:2019:MOA

[2189] Sima Sinaei and Omid Fatemi. Multiobjective algorithms for the application mapping problem in heterogeneous multiprocessor embedded system design. The Journal of Supercomputing, 75(8):4150–4176, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Villegas:2019:TST

[2190] Alejandro Villegas, Angeles Navarro, Rafael Asenjo, and Oscar Plata. Toward a software transactional memory for heterogeneous CPU–GPU processors. The Journal of Supercomputing, 75(8):4177–4192, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:HSG

[2191] Yuxiang Li, Yinliang Zhao, Liyu Sun, and Mengjuan Shen. A hybrid sample generation approach in speculative multithreading. The Journal of Supercomputing, 75(8):4193–4225, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2019:TFT

[2192] Cheng Chen, Yunfei Du, Ke Zuo, Jianbin Fang, and Canqun Yang. Toward fault-tolerant hybrid programming over large-scale heterogeneous clusters via checkpointing/restart optimization. The Journal of Supercomputing, 75(8):4226–4247, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yi:2019:HPC

[2193] Gangman Yi and Vincenzo Loia. Highperformance computing systems and applications for AI. The Journal of Supercomputing, 75(8):4248–4251, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02937-z.pdf.

Jung:2019:EDS

[2194] Daeyong Jung, Daewon Lee, Myungil Kim, and Jaesung Kim. Efficient data synchronization method on integrated computing environment. The Journal of Supercomputing, 75(8):4252– 4266, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2019:ESM

[2195] Sangdo Lee and Jun-Ho Huh. An effective security measures for nuclear power plant using big data analysis approach. The Journal of Supercomputing, 75(8):4267–4294, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yin:2019:IDP

[2196] Chunyong Yin, Hongyi Wang, Xiang Yin, Ruxia Sun, and Jin Wang. Improved deep packet inspection in data stream detection. The Journal of Supercomputing, 75(8):4295–4308, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hwang:2019:NCM

[2197] Soyoung Hwang. A network clock model for time awareness in the Internet of Things and artificial intelligence applications. The Journal of Supercomputing, 75(8):4309–4328, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2019:HAP

[2198] Sokjoon Lee, Hwajeong Seo, Hyeokchan Kwon, and Hyunsoo Yoon. Hybrid approach of parallel implementation on CPU–GPU for high-speed ECDSA verification. The Journal of Supercomputing, 75(8):4329–4349, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2019:CRP

[2199] Shin Hyung Park, Shin Hyoung Park, Oh Hoon Kwon, and Yunsick Sung. Continuous risk profile and clusteringbased method for investigating the effect of the automated enforcement system on urban traffic collisions. The Journal of Supercomputing, 75(8): 4350–4371, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ryu:2019:BBD

[2200] Jung Hyun Ryu, Pradip Kumar Sharma, Jeong Hoon Jo, and Jong Hyuk Park. A blockchain-based decentralized efficient investigation framework for IoT digital forensics. The Journal of Supercomputing, 75(8):4372–4387, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Han:2019:EJM

[2201] Seok-Hyeon Han, Hyun-Woo Kim, and Young-Sik Jeong. An efficient job management of computing service using integrated idle VM resources for high-performance computing based on OpenStack. The Journal of Supercomputing, 75(8):4388–4407, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2019:EEC

[2202] Chao-Tung Yang, Shuo-Tsung Chen, Jung-Chun Liu, Yu-Wei Chan, Chien-Chih Chen, and Vinod Kumar Verma. An energy-efficient cloud system with novel dynamic resource allocation methods. The Journal of Supercomputing, 75(8):4408–4429, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tian:2019:ORM

[2203] Yifei Tian, Wei Song, Su Sun, Simon Fong, and Shuanghui Zou. 3D object recognition method with multiple feature extraction from LiDAR point clouds. The Journal of Supercomputing, 75(8):4430–4442, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hao:2019:EPS

[2204] Fei Hao, Huijuan Guo, Doo-Soon Park, and Jungho Kang. An efficient pricing strategy of sensing tasks for crowdphotographing. The Journal of Supercomputing, 75(8):4443–4458, August 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:SCM

[2205] Hyun-Woo Kim, Jungho Kang, and Young-Sik Jeong. Simulator considering modeling and performance evaluation for high-performance computing of collaborative-based mobile cloud infrastructure. The Journal of Supercomputing, 75(8):4459–4471, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction notice [2375].

Lee:2019:AFT

[2206] JongHyuk Lee and JoonMin Gil. Adaptive fault-tolerant scheduling strategies for mobile cloud computing. The Journal of Supercomputing, 75(8):4472– 4488, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lim:2019:DDP

[2207] Kwang Kyu Lim, JiSu Park, and Jin Gon Shon. Differential data processing technique to improve the performance of wireless sensor networks. The Journal of Supercomputing, 75(8): 4489–4504, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jeong:2019:WBO

[2208] Ji-Seong Jeong, Min-Ho Song, Sang-Ho Lee, Mihye Kim, Nakhoon Baek, and Kwan-Hee Yoo. A web-based 3D ontology navigation system for spinal disease diagnosis. The Journal of Supercomputing, 75(8):4505–4518, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fang:2019:MAL

[2209] Juan Fang, Xibei Zhang, Shijian Liu, and Zeqing Chang. Miss-aware LLC buffer management strategy based on heterogeneous multi-core. The Journal of Supercomputing, 75(8):4519–4528, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02763-3.pdf.

Kang:2019:LMA

[2210] Jungho Kang. Lightweight mutual authentication RFID protocol for secure multi-tag simultaneous authentication in ubiquitous environments. The Journal of Supercomputing, 75(8): 4529–4542, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Singh:2019:CSA

[2211] Saurabh Singh, Pradip Kumar Sharma, Seo Yeon Moon, Daesung Moon, and Jong Hyuk Park. A comprehensive study on APT attacks and countermeasures for future networks and communications: challenges and solutions. The Journal of Supercomputing, 75(8): 4543–4574, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yi:2019:DFD

[2212] Hayoon Yi, Yeongpil Cho, Yunheung Paek, and Kwangman Ko. DADE: a fast data anomaly detection engine for kernel integrity monitoring. The Journal of Supercomputing, 75(8): 4575–4600, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Feng:2019:SRS

[2213] Tung-Huang Feng, Neng-Yih Shih, and Min-Shiang Hwang. Safety relay selection algorithms based on fuzzy relationship for wireless sensor networks. The Journal of Supercomputing, 75(8): 4601–4616, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2019:ESR

[2214] Sung-Wook Park and Im-Yeong Lee. Enhanced signature RTD transaction scheme based on Chebyshev polynomial for mobile payments service in IoT device environment. The Journal of Supercomputing, 75(8):4617– 4637, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kim:2019:CAE

[2215] Dong-Oh Kim, Hong-Yeon Kim, Young-Kyun Kim, and Jeong-Joon Kim. Cost analysis of erasure coding for exa-scale storage. The Journal of Supercomputing, 75(8):4638–4656, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-2663-4.pdf.

Nguyen:2019:SRN

[2216] Tuan Anh Nguyen, Kihong Han, Dugki Min, Eunmi Choi, Tran Duc Thang, and Yun-Jeong Choi. A stochastic reward net-based assessment of reliability, availability and operational cost for a software-defined network infrastructure. The Journal of Supercomputing, 75(8):4657–4683, August 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ayoade:2019:SDP

[2217] Gbadebo Ayoade, Amir El-Ghamry, Vishal Karande, Latifur Khan, Mohammed Alrahmawy, and Magdi Zakria Rashad. Secure data processing for IoT middleware systems. The Journal of Supercomputing, 75(8):4684– 4709, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Xu:2019:AAN

[2218] Zhen Xu, Xia Zhao, Zhiying Wang, and Canqun Yang. Application-aware NoC management in GPUs multitasking. The Journal of Supercomputing, 75 (8):4710–4730, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jia:2019:PSM

[2219] Nan Jia, Yuan Huang, Jiapei Li, Haigang An, Xiaomin Jia, and Ruomei Wang. Parallel simulation model for heat and moisture transfer of clothed human body. The Journal of Supercomputing, 75(8):4731–4749, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khattar:2019:TEE

[2220] Nagma Khattar, Jagpreet Sidhu, and Jaiteg Singh. Toward energy-efficient cloud computing: a survey of dynamic power management and heuristicsbased optimization techniques. The Journal of Supercomputing, 75(8): 4750–4810, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Dinneen:2019:FCS

[2221] Michael J. Dinneen, Anuradha Mahasinghe, and Kai Liu. Finding the chromatic sums of graphs using a D-Wave quantum computer. The Journal of Supercomputing, 75(8):4811– 4828, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sahoo:2019:TSS

[2222] Kshira Sagar Sahoo, Sanjaya Kumar Panda, Sampa Sahoo, Bibhudatta Sahoo, and Ratnakar Dash. Toward secure software-defined networks against distributed denial of service attack. The Journal of Supercomputing, 75(8): 4829–4874, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2019:DBO

[2223] Sanghyun Park and Taeweon Suh. DQN-based OpenCL workload partition for performance optimization. The Journal of Supercomputing, 75(8): 4875–4893, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bhih:2019:DIA

[2224] Amhmed Bhih, Princy Johnson, and Martin Randles. Decentralized iterative approaches for community clustering in the networks. The Journal of Supercomputing, 75(8):4894–4917, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        019-02765-1.pdf.

Brighen:2019:LAM

[2225] Assia Brighen, Hachem Slimani, Abdelmounaam Rezgui, and Hamamache Kheddouci. Listing all maximal cliques in large graphs on vertex-centric model. The Journal of Supercomputing, 75(8): 4918–4946, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Grylonakis:2019:PUT

[2226] E. N. G. Grylonakis, G. A. Gravvanis, C. K. Filelis-Papadopoulos, and A. S. Fokas. A parallel unified transform solver based on domain decomposition for solving linear elliptic PDEs. The Journal of Supercomputing, 75(8): 4947–4985, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Micheal:2019:AOT

[2227] Ancy A. Micheal and K. Vani. Automatic object tracking in optimized UAV video. The Journal of Supercomputing, 75(8):4986–4999, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gopikrishnan:2019:HHA

[2228] S. Gopikrishnan, P. Priakanth, and Rolly Maulana Awangga. HSIR: hybrid architecture for sensor identification and registration for IoT applications. The Journal of Supercomputing, 75(8):5000–5018, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Waidyasooriya:2019:OBD

[2229] Hasitha Muthumala Waidyasooriya, Masanori Hariyama, Masamichi J.

Miyama, and Masayuki Ohzeki. OpenCL-based design of an FPGA accelerator for quantum annealing simulation. The Journal of Supercomputing, 75(8):5019–5039, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumari:2019:LTD

[2230] Jyoti Kumari, Prabhat Kumar, and Sunil Kumar Singh. Localization in three-dimensional wireless sensor networks: a survey. The Journal of Supercomputing, 75(8):5040–5083, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jensen:2019:PMP

[2231] Erik J. Jensen, Evan Coleman, and Masha Sosonkina. Predictive modeling of the performance of asynchronous iterative methods. The Journal of Supercomputing, 75(8):5084–5105, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Saravanan:2019:CCI

[2232] N. Saravanan and R. Sivaramakrishnan. Command and control of industrial manipulator through speechbased interfaces in Indic languages. The Journal of Supercomputing, 75(8): 5106–5117, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Oskouei:2019:DNR

[2233] Saeed Mirzajani Oskouei and Ali Ghaffari. Designing a new reversible ALU by QCA for reducing occupation area. The Journal of Supercomputing, 75(8): 5118–5144, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Perumal:2019:IFR

[2234] Sankar Pariserum Perumal, Ganapathy Sannasi, and Kannan Arputharaj. An intelligent fuzzy rule-based e-learning recommendation system for dynamic user interests. The Journal of Supercomputing, 75(8):5145–5160, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sivakumaran:2019:CCT

[2235] Krupa Sivakumaran and Arul Siromoney.

Cache control techniques to provide QoS on real systems. The Journal of Supercomputing, 75(8):5161–5188, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Du:2019:IPF

[2236] Zhenxin Du, Dezhi Han, and Kuan-Ching Li. Improving the performance of feature selection and data clustering with novel global search and eliteguided artificial bee colony algorithm. The Journal of Supercomputing, 75(8): 5189–5226, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ozcan-Top:2019:WEM

[2237] Özden Özcan-Top and Fergal McCaffery. To what extent the medical device software regulations can be achieved with agile software development methods? XP–DSDM-Scrum. The Journal of Supercomputing, 75(8): 5227–5260, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Birzhandi:2019:CCB

[2238] Pardis Birzhandi and Hee Yong Youn. CBCH (clustering-based convex hull) for reducing training time of support vector machine. The Journal of Supercomputing, 75(8):5261–5279, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Al-Betar:2019:IFP

[2239] Mohammed Azmi Al-Betar, Mohammed A. Awadallah, Iyad Abu Doush, Abdelaziz I. Hammouri, Majdi Mafarja, and Zaid Abdi Alkareem Alyasseri. Island flower pollination algorithm for global optimization. The Journal of Supercomputing, 75(8): 5280–5323, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Filiposka:2019:MHV

[2240] Sonja Filiposka, Anastas Mishev, and Katja Gilly. Multidimensional hierarchical VM migration management for HPC cloud environments. The Journal of Supercomputing, 75(8):5324– 5346, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Marimuthu:2019:OOA

[2241] Poorani Marimuthu, Varalakshmi Perumal, and Vaidehi Vijayakumar. OAFPM: optimized ANFIS using frequent pattern mining for activity recognition. The Journal of Supercomputing, 75(8):5347–5366, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vu:2019:APA

[2242] Huy-The Vu, Yuichi Okuyama, and Abderazek Ben Abdallah. Analytical performance assessment and highthroughput low-latency spike routing algorithm for spiking neural network systems. The Journal of Supercomputing, 75(8):5367–5397, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ekhtiyari:2019:TAE

[2243] Zohreh Ekhtiyari, Vahidreza Moghaddas, and Hakem Beitollahi. A temperature-aware and energy-efficient fuzzy technique to schedule tasks in heterogeneous MPSoC systems. The Journal of Supercomputing, 75(8): 5398–5419, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sebaa:2019:QOC

[2244] Abderrazak Sebaa and Abdelkamel Tari. Query optimization in cloud environments: challenges, taxonomy, and techniques. The Journal of Supercomputing, 75(8):5420–5450, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhu:2019:IRS

[2245] Xiaojun Zhu, Bing Chen, Muhui Shen, and Yanchao Zhao. Inapproximability results and suboptimal algorithms for minimum delay cache placement in campus networks with content-centric network routers. The Journal of Supercomputing, 75(8):5451–5474, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Esmaili:2019:ISB

[2246] Ehsan Esmaili, Ali Akoglu, Salim Hariri, and Talal Moukabary. Implementation of scalable bidomain-based 3D cardiac simulations on a graphics processing unit cluster. The Journal of Supercomputing, 75(8):5475– 5506, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shanmugam:2019:IFS

[2247] S. Shanmugam and J. Preethi. Improved feature selection and classification for rheumatoid arthritis disease using weighted decision tree approach (REACT). The Journal of Supercomputing, 75(8):5507–5519, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Farshin:2019:MKB

[2248] Alireza Farshin and Saeed Sharifian. A modified knowledge-based ant colony algorithm for virtual machine placement and simultaneous routing of NFV in distributed cloud architecture. The Journal of Supercomputing, 75(8): 5520–5550, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bahador:2019:HSB

[2249] Mohammad Bagher Bahador, Mahdi Abadi, and Asghar Tajoddin. HLMD: a signature-based approach to hardwarelevel behavioral malware detection and classification. The Journal of Supercomputing, 75(8):5551–5582, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2019:OOD

[2250] Fan Wang and Weisheng Zhao. Oneto-one disjoint path covers in hypercubes with faulty edges. The Journal of Supercomputing, 75(8):5583– 5595, August 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmadpour:2019:CNF

[2251] Seyed-Sajad Ahmadpour and Mohammad Mosleh. Correction to: A novel fault-tolerant multiplexer in quantumdot cellular automata technology. The Journal of Supercomputing, 75(8):5596, August 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02870-1.pdf. See [1865].

Choi:2019:ULA

[2252] Hyunseung Choi, Mintae Kim, Gyubok Lee, and Wooju Kim. Unsupervised learning approach for network intrusion detection system using autoencoders. The Journal of Supercomputing, 75(9):5597–5621, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Seghir:2019:IBM

[2253] Fateh Seghir, Abdallah Khababa, and Fouzi Semchedine. An interval-based multi-objective artificial bee colony algorithm for solving the web service composition under uncertain QoS. The Journal of Supercomputing, 75(9): 5622–5666, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abdulhassan:2019:CFB

[2254] A. A. Abdulhassan and M. Ahmadi. Cuckoo filter-based many-field packet classification using X-tree. The Journal of Supercomputing, 75(9):5667– 5687, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moghimizadeh:2019:NDF

[2255] Taban Moghimizadeh and Mohammad Mosleh. A novel design of faulttolerant RAM cell in quantum-dot cellular automata with physical verification. The Journal of Supercomputing, 75(9):5688–5716, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Armejach:2019:DTO

[2256] Adrià Armejach, Marc Casas, and Miquel Moretó. Design trade-offs for emerging HPC processors based on mobile market technology. The Journal of Supercomputing, 75(9):5717– 5740, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nogueira:2019:ACG

[2257] Bruno Nogueira, Eduardo Tavares, Jean Araujo, and Gustavo Callou. Accelerating continuous GRASP with a GPU. The Journal of Supercomputing, 75(9):5741–5759, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ahmadvand:2019:SSA

[2258] Hossein Ahmadvand and Maziar Goudarzi. SAIR: significance-aware approach to improve QoR of big data processing in case of budget constraint.

The Journal of Supercomputing, 75(9): 5760–5781, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2019:WBM

[2259] Wen Liu, Tuqian Zhang, and Junxia Liu. Window-based multiple continuous query algorithm for data streams. The Journal of Supercomputing, 75(9): 5782–5807, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2019:MCD

[2260] Xinsheng Zhang and Zhe Wang. A microcalcification cluster detection method based on deep learning and multi-scale feature fusion. The Journal of Supercomputing, 75(9):5808– 5830, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gao:2019:AID

[2261] Hongyi Gao, Xi Zeng, and Chunhua Yao. Application of improved distributed naive Bayesian algorithms in text classification. The Journal of Supercomputing, 75(9):5831– 5847, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Pal:2019:NPR

[2262] Souvik Pal, Raghvendra Kumar, Le Hoang Son, Krishnan Saravanan, Mohamed Abdel-Basset, Gunasekaran Manogaran, and Pham Huy Thong. Novel probabilistic resource migration algorithm for cross-cloud live migration of virtual machines in public cloud. The Journal of Supercomputing, 75(9):

5848–5865, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tembhare:2019:RBP

[2263] Akshay Tembhare, S. Sibi Chakkaravarthy, D. Sangeetha, V. Vaidehi, and M. Venkata Rathnam. Role-based policy to maintain privacy of patient health records in cloud. The Journal of Supercomputing, 75(9):5866– 5881, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wu:2019:MMI

[2264] Shaofei Wu, Jun Liu, and Lizhi Liu. Modeling method of Internet public information data mining based on probabilistic topic model. The Journal of Supercomputing, 75(9):5882– 5897, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yi:2019:MAP

[2265] Guohong Yi, Zhili Feng, Tiancan Mei, Pushan Li, Wang Jin, and Siyuan Chen. Multi-AGVs path planning based on improved ant colony algorithm. The Journal of Supercomputing, 75(9):5898–5913, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Selvam:2019:PRC

[2266] Blessy Selvam, S. Ravimaran, and Sheba Selvam. Parallelized root cause analysis using cause-related aspect formulation technique (CRAFT). The Journal of Supercomputing, 75(9): 5914–5929, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:2019:GCT

[2267] Guolong Chen, Zhenghua Xin, Han Li, Tonghai Zhu, Maodi Wang, Yaxiang Liu, and Sanqiang Wei. A greedy constructing tree algorithm for shortest path in perpetual wireless recharging wireless sensor network. The Journal of Supercomputing, 75(9):5930– 5945, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2019:ACK

[2268] Zhongru Wang and Binxing Fang. Application of combined kernel function artificial intelligence algorithm in mobile communication network security authentication mechanism. The Journal of Supercomputing, 75(9):5946– 5964, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [2269].

Wang:2019:CAC

[2269] Zhongru Wang and Binxing Fang.

Correction to: Application of combined kernel function artificial intelligence algorithm in mobile communication network security authentication mechanism. The Journal of Supercomputing, 75(9):5965, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02949-9.pdf. See [2268].

Gong:2019:RSE

[2270] Yunlu Gong and Lianguo Jia. Research on SVM environment performance of parallel computing based on large data set of machine learning. The Journal of Supercomputing, 75(9): 5966–5983, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ma:2019:LCA

[2271] Jingyan Ma and Kehong Zhang. Landmark community approximate optimal path query with pure linguistic attributes. The Journal of Supercomputing, 75(9):5984–6003, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hu:2019:APA

[2272] Jun Hu, Jun Fang, Yanhua Du, Zhe Liu, and Pengyang Ji. Application of PLS algorithm in discriminant analysis in multidimensional data mining. The Journal of Supercomputing, 75(9): 6004–6020, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hui:2019:MVT

[2273] Qiuli Hui. Motion video tracking technology in sports training based on mean-shift algorithm. The Journal of Supercomputing, 75(9):6021– 6037, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See retraction notice [2374].

Hanuman:2019:IMP

[2274] C. R. S. Hanuman, J. Kamala, and A. R. Aruna. Implementation of multiprecision floating point divider for high speed signal processing applications. The Journal of Supercomputing, 75(9): 6038–6054, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhang:2019:AWS

[2275] Zhicheng Zhang and Yan Zhang. Application of wireless sensor network in dynamic linkage video surveillance system based on Kalman filtering algorithm. The Journal of Supercomputing, 75(9):6055–6069, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhao:2019:ACA

[2276] Shifei Zhao. Application of a clustering algorithm in sports video image extraction and processing. The Journal of Supercomputing, 75(9):6070– 6084, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

BalaAnand:2019:EGB

[2277] M. BalaAnand, N. Karthikeyan, S. Karthik, R. Varatharajan, Gunasekaran Manogaran, and C. B. Sivaparthipan. An enhanced graph-based semi-supervised learning algorithm to detect fake users on Twitter. The Journal of Supercomputing, 75(9):6085– 6105, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Vignesh:2019:ORP

[2278] V. Vignesh and K. Premalatha. Optimal route path sustainability in military information system with reduced interference effect. The Journal of Supercomputing, 75(9):6106– 6117, September 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Park:2019:OAP

[2279] Cheong Hee Park. Outlier and anomaly pattern detection on data streams. The Journal of Supercomputing, 75(9): 6118–6128, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yoo:2019:IST

[2280] Jae-Jun Yoo, Woong-Kee Loh, and Kyu-Young Whang. Indexable subtrajectory matching using multisegment approximation: a partitionand-stitch framework. The Journal of Supercomputing, 75(9):6129–6157, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ibrahim:2019:CSR

[2281] Ahmed Ibrahim, Craig Valli, Ian McAteer, and Junaid Chaudhry. Correction to: A security review of local government using NIST CSF: a case study. The Journal of Supercomputing, 75(9):6158, September 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        019-02972-w.pdf. See [1888].

Xie:2019:SCR

[2282] Yuantao Xie, Wen Wang, Yabo Guo, and Juan Yang. Study on the country risk rating with distributed crawling system. The Journal of Supercomputing, 75(10):6159–6177, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Moreno:2019:HGO

[2283] R. Moreno, E. Arias, A. Navarro, and F. J. Tapiador. How good is the OpenPOWER architecture for highperformance CPU-oriented weather forecasting applications? The Journal of Supercomputing, 75(10):6178– 6193, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abdelhafez:2019:CBS

[2284] Amr Abdelhafez, Enrique Alba, and Gabriel Luque. A component-based study of energy consumption for sequential and parallel genetic algorithms. The Journal of Supercomputing, 75(10):6194–6219, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shen:2019:PLB

[2285] Fanfan Shen, Yanxiang He, Jun Zhang, and Chao Xu. Periodic learning-based region selection for energy-efficient MLC STT–RAM cache. The Journal of Supercomputing, 75(10):6220– 6238, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Motaki:2019:CSB

[2286] Saloua El Motaki, Ali Yahyaouy, Hamid Gualous, and Jalal Sabor. Comparative study between exact and metaheuristic approaches for virtual machine placement process as knapsack problem. The Journal of Supercomputing, 75(10):6239–6259, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Darmani:2019:QET

[2287] Yousef Darmani and Mehrdad Sangelaji. QoS-enabled TCP for software-defined networks: a combined scheduler-per-node approach. The Journal of Supercomputing, 75(10): 6260–6276, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Aljabri:2019:IPE

[2288] Naif Aljabri, Muhammad Al-Hashimi, Mostafa Saleh, and Osama Abulnaja. Investigating power efficiency of mergesort. The Journal of Supercomputing, 75(10):6277–6302, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02850-5.pdf.

Nguyen:2019:EES

[2289] Hoang Minh Nguyen, Gaurav Kalra, Tae Joon Jun, Sungpil Woo, and Daeyoung Kim. ESNemble: an Echo State Network-based ensemble for workload prediction and resource allocation of Web applications in the cloud. The Journal of Supercomputing, 75(10): 6303–6323, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hbaieb:2019:STV

[2290] Ameni Hbaieb, Mahdi Khemakhem, and Maher Ben Jemaa. A survey and taxonomy on virtual data center embedding. The Journal of Supercomputing, 75(10):6324–6360, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rosendo:2019:MAA

[2291] Daniel Rosendo, Demis Gomes, Guto Leoni Santos, Glauco Goncalves, Andre Moreira, Leylane Ferreira, Patricia Takako Endo, Judith Kelner, Djamel Sadok, Amardeep Mehta, and Mattias Wildeman. A methodology to assess the availability of nextgeneration data centers. The Journal of Supercomputing, 75(10):6361– 6385, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hemasian-Etefagh:2019:DSA

[2292] Farinaz Hemasian-Etefagh and Faramarz Safi-Esfahani. Dynamic scheduling applying new population grouping of whales meta-heuristic in cloud computing. The Journal of Supercomputing, 75(10):6386–6450, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tavakoli-Someh:2019:MOV

[2293] Sanaz Tavakoli-Someh and Mohammad Hossein Rezvani. Multi-objective virtual network function placement using NSGA–II meta-heuristic approach. The Journal of Supercomputing, 75 (10):6451–6487, October 2019. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Irandoost:2019:LAB

[2294] Mohammad Amin Irandoost, Amir Masoud Rahmani, and Saeed Setayeshi. Learning automata-based algorithms for MapReduce data skewness handling. The Journal of Supercomputing, 75(10):6488–6516, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rodriguez-Pascual:2019:JMH

[2295] Manuel Rodríguez-Pascual, Jiajun Cao, José A. Moríñigo, Gene Cooperman, and Rafael Mayo-García. Job migration in HPC clusters by means of checkpoint/restart. The Journal of Supercomputing, 75(10):6517–6541, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Puche:2019:FLP

[2296] José Puche, Salvador Petit, Julio Sahuquillo, and María Engracia Gómez. FOS: a low-power cache organization for multicores. The Journal of Supercomputing, 75(10):6542– 6573, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abbasi:2019:CAF

[2297] M. Abbasi and M. Rafiee. A calibrated asymptotic framework for analyzing packet classification algorithms on GPUs. The Journal of Supercomputing, 75(10):6574–6611, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

ImaniMehr:2019:TBI

[2298] Zahra ImaniMehr and Mehdi DehghanTakhtFooladi. Token-based incentive mechanism for peer-to-peer video streaming networks. The Journal of Supercomputing, 75(10):6612– 6631, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Choi:2019:WFA

[2299] Won Gi Choi and Sanghyun Park. A write-friendly approach to manage namespace of Hadoop distributed file system by utilizing nonvolatile memory. The Journal of Supercomputing, 75(10):6632–6662, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Arab:2019:IEM

[2300] Alireza Arab, Mohammad Javad Rostami, and Behnam Ghavami. An image encryption method based on chaos system and AES algorithm. The Journal of Supercomputing, 75(10): 6663–6682, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s11227-019-02878-7.pdf.

Mohammadi:2019:ILP

[2301] Somayeh Mohammadi, Latif PourKarimi, and Hossein Pedram. Integer linear programming-based multi-objective scheduling for scientific workflows in multi-cloud environments. The Journal of Supercomputing, 75(10):6683– 6709, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kelefouras:2019:MCC

[2302] Vasilios Kelefouras and Karim Djemame. A methodology correlating code optimizations with data memory accesses, execution time and energy consumption. The Journal of Supercomputing, 75(10):6710–6745, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lee:2019:TSA

[2303] Myoungjun Lee and Soontae Kim.

Time-sensitivity-aware shared cache architecture for multi-core embedded systems. The Journal of Supercomputing, 75(10):6746–6776, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hussain:2019:SRC

[2304] Altaf Hussain, Muhammad Aleem, Muhammad Azhar Iqbal, and Muhammad Arshad Islam. SLA-RALBA: costefficient and resource-aware load balancing algorithm for cloud computing. The Journal of Supercomputing, 75(10):6777–6803, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yang:2019:PIA

[2305] Seokwoo Yang, Siwoon Son, Mi-Jung Choi, and Yang-Sae Moon. Performance improvement of Apache Storm using InfiniBand RDMA. The Journal of Supercomputing, 75(10):6804– 6830, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khajekarimi:2019:EMS

[2306] Elyas Khajekarimi, Kamal Jamshidi, and Abbas Vafaei. Energy minimization in the STT–RAM-based highcapacity last-level caches. The Journal of Supercomputing, 75(10):6831– 6854, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jain:2019:SES

[2307] Anugrah Jain, Vijay Laxmi, Meenakshi Tripathi, Manoj Singh Gaur, and Rimpy Bishnoi. S2DIO: an extended scalable 2D mesh network-on-chip routing reconfiguration for efficient bypass of link failures. The Journal of Supercomputing, 75(10):6855– 6881, October 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Das:2019:NDR

[2308] Jadav Chandra Das and Debashis De.

Novel design of reversible priority encoder in quantum dot cellular automata based on Toffoli gate and Feynman gate. The Journal of Supercomputing, 75(10):6882–6903, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mohammadhosseini:2019:EEA

[2309] Mahdieh Mohammadhosseini, Abolfazl Toroghi Haghighat, and Ebrahim Mahdipour. An efficient energy-aware method for virtual machine placement in cloud data centers using the cultural algorithm. The Journal of Supercomputing, 75(10):6904–6933, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Maleki:2019:MIR

[2310] Neda Maleki, Amir Masoud Rahmani, and Mauro Conti. MapReduce: an infrastructure review and research insights. The Journal of Supercomputing, 75(10):6934–7002, October 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yokoyama:2019:SAP

[2311] Daniel Yokoyama, Bruno Schulze, Fábio Borges, and Giacomo Mc Evoy. The survey on ARM processors for HPC. The Journal of Supercomputing, 75(10):7003–7036, October 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sohrabi:2019:IPN

[2312] Mohammad Karim Sohrabi and Somayyeh Alimirzaee. Improving performance of node clustering in wireless sensor networks using meta-heuristic algorithms and a novel validity index. The Journal of Supercomputing, 75(11):7550–7572, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Silva:2019:ILP

[2313] Jefferson Silva, Márcio Kreutz, Monica Pereira, and Marjory Da Costa-Abreu. An investigation of latency prediction for NoC-based communication architectures using machine learning techniques. The Journal of Supercomputing, 75(11):7573–7591, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nguyen:2019:HLP

[2314] Hoang Minh Nguyen, Gaurav Kalra, and Daeyoung Kim. Host load prediction in cloud computing using long short-term memory encoder–decoder. The Journal of Supercomputing, 75 (11):7592–7605, November 2019. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gul:2019:CRB

[2315] Beenish Gul, Imran Ali Khan, Saad Mustafa, Osman Khalid, and Atta ur Rehman Khan. CPU–RAM-based energy-efficient resource allocation in clouds. The Journal of Supercomputing, 75(11):7606–7624, November 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Knap:2019:PEU

[2316] Marcin Knap and Paweł Czarnul. Performance evaluation of unified memory with prefetching and oversubscription for selected parallel CUDA applications on NVIDIA Pascal and Volta GPUs. The Journal of Supercomputing, 75(11):7625–7645, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11227-
019-02966-8.pdf.

Dadashi:2019:DLC

[2317] Sajed Dadashi, Akram Reza, Midia Reshadi, and Ahmad Khademzadeh. Decreasing latency considering power consumption issue in silicon interposerbased network-on-chip. The Journal of Supercomputing, 75(11):7646– 7664, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cha:2019:EMM

[2318] Myung-Hoon Cha, Sang-Min Lee, Hong-Yeon Kim, and Young-Kyun Kim. Effective metadata management in exascale file system. The Journal of Supercomputing, 75(11):7665– 7689, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Cheng:2019:HCP

[2319] Chien-Fu Cheng and Chung-Wei Huang. The harmonized consensus protocol in distributed systems. The Journal of Supercomputing, 75(11):7690– 7722, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zaheer:2019:LAP

[2320] Saad Zaheer, Asad Waqar Malik, Anis Ur Rahman, and Safdar Abbas Khan. Locality-aware process placement for parallel and distributed simulation in cloud data centers. The Journal of Supercomputing, 75(11):7723– 7745, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Indumathi:2019:NSF

[2321] G. Indumathi, V. P. M. B. Aarthi, and M. Ramesh. A novel semi-folded parallel successive cancellation-based polar decoder for optimal-register allocation. The Journal of Supercomputing, 75(11):7037–7052, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Fernandez:2019:ATS

[2322] Ivan Fernandez, Alejandro Villegas, Eladio Gutierrez, and Oscar Plata. Accelerating time series motif discovery in the Intel Xeon Phi KNL processor. The Journal of Supercomputing, 75(11):7053–7075, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Guo:2019:SSA

[2323] Wenxia Guo, Ping Kuang, Yaqiu Jiang, Xiang Xu, and Wenhong Tian. SAVE: self-adaptive consolidation of virtual machines for energy efficiency of CPUintensive applications in the cloud. The Journal of Supercomputing, 75(11): 7076–7100, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rodriguez-Gutiez:2019:TBL

[2324] Eduardo Rodriguez-Gutiez, Ana Moreton-Fernandez, Arturo Gonzalez-Escribano, and Diego R. Llanos. Toward a BLAS library truly portable across different accelerator types. The Journal of Supercomputing, 75(11):7101– 7124, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mukherjee:2019:IFE

[2325] Anwesha Mukherjee, Priti Deb, Debashis De, and Rajkumar Buyya. IoT-F2N: An energy-efficient architectural model for IoT using Femtolet-based fog network. The Journal of Supercomputing, 75(11):7125–7146, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shokri:2019:ESS

[2326] Madeh Shokri and Meghdad Mirabi.

An efficient stream structure for broadcasting the encrypted XML data in mobile wireless broadcast channels. The Journal of Supercomputing, 75(11): 7147–7173, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Masdari:2019:CCD

[2327] Mohammad Masdari, Saeid Barshande, and Suat Ozdemir. CDABC: chaotic discrete artificial bee colony algorithm for multi-level clustering in large-scale WSNs. The Journal of Supercomputing, 75(11):7174–7208, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:SRS

[2328] Chunlin Li, Jianhang Tang, and Youlong Luo. Scalable replica selection based on node service capability for improving data access performance in edge computing environment. The Journal of Supercomputing, 75(11):7209–7243, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karimi:2019:EFM

[2329] Shiva Karimi and Yousef Darmani. pepidemic forwarding method for heterogeneous delay-tolerant networks. The Journal of Supercomputing, 75 (11):7244–7264, November 2019. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

deOliveira:2019:LCD

[2330] Carlos Henrique Rodrigues de Oliveira, Ana Paula Ferreira Costa, Vitor Fonseca Thomaz, and Igor Amorim Silva. Low-cost deployment proposal to urban mobility in smart cities. The Journal of Supercomputing, 75(11):7265– 7289, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Karimi:2019:REC

[2331] A. Karimi and S. M. Amini. Reduction of energy consumption in wireless sensor networks based on predictable routes for multi-mobile sink. The Journal of Supercomputing, 75(11):7290– 7313, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kabirirad:2019:HSG

[2332] Saeideh Kabirirad, Mahmood Fazlali, and Ziba Eslami. High-speed GPU implementation of a secret sharing scheme based on cellular automata. The Journal of Supercomputing, 75(11):7314– 7336, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Thai:2019:HBF

[2333] Quang Tung Thai, Jong-Chul Yim, Tae-Whan Yoo, Hyun-Kyung Yoo, Ji-Young Kwak, and Sun-Me Kim. Hierarchical Byzantine fault-tolerance protocol for permissioned blockchain systems. The Journal of Supercomputing, 75(11):7337–7365, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:PPC

[2334] Yuxiang Li, Zhiyong Zhang, Lili Zhang, Danmei Niu, Changwei Zhao, Bin Song, and Liuke Liang. ProCTA: program characteristic-based thread partition approach. The Journal of Supercomputing, 75(11):7366–7390, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mohammadi:2019:NOA

[2335] Alireza Mohammadi and Mohammad Hossein Rezvani. A novel optimized approach for resource reservation in cloud computing using producer–consumer theory of microeconomics. The Journal of Supercomputing, 75(11):7391–7425, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Singh:2019:LBV

[2336] Sunil Kumar Singh and Prabhat Kumar. A load balancing virtual level routing (LBVLR) using mobile mule for large sensor networks. The Journal of Supercomputing, 75(11):7426– 7459, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2019:ABS

[2337] Feng Wang, Dingde Jiang, Hong Wen, and Houbing Song. Adaboost-based security level classification of mobile intelligent terminals. The Journal of Supercomputing, 75(11):7460– 7478, November 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Khatibi:2019:DDD

[2338] Elahe Khatibi and Seyedeh Leili Mirtaheri. A dynamic data dissemination mechanism for Cassandra NoSQL data store. The Journal of Supercomputing, 75(11):7479–7496, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kalantari:2019:LCC

[2339] Zeinab Kalantari, Mohammad Eshghi, Majid Mohammadi, and Somayeh Jassbi. Low-cost and compact design method for reversible sequential circuits. The Journal of Supercomputing, 75(11):7497–7519, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rathinaraja:2019:DRB

[2340] J. Rathinaraja, V. S. Ananthanarayana, and Anand Paul. Dynamic ranking-based MapReduce job scheduler to exploit heterogeneous performance in a virtualized environment. The Journal of Supercomputing, 75(11):7520–7549, November 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Shooshtarian:2019:MRE

[2341] L. Shooshtarian and F. Safaei. A maximally robustness embedding algorithm in virtual data centers with multi-attribute node ranking based on TOPSIS. The Journal of Supercomputing, 75(12):8059–8093, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Joghan:2019:WLP

[2342] Hamid Shahrivari Joghan, Alireza Bagheri, and Meysam Azad. Weighted label propagation based on local edge betweenness. The Journal of Supercomputing, 75(12):8094–8114, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Matsumoto:2019:IPE

[2343] Kazuya Matsumoto, Yasuhiro Idomura, Takuya Ina, Akie Mayumi, and Susumu Yamada. Implementation and performance evaluation of a communication-avoiding GMRES method for stencil-based code on GPU cluster. The Journal of Supercomputing, 75(12):8115–8146, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hagras:2019:LBT

[2344] Tarek Hagras, Asmaa Atef, and Yousef B. Mahdy. Lower-bound timecomplexity greening mechanism for duplication-based scheduling on largescale computing platforms. The Journal of Supercomputing, 75(12):8147– 8167, December 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Heris:2019:PCS

[2345] Mahdi Saedshoar Heris and Mohammad Javidi. A predictor-corrector scheme for the tempered fractional differential equations with uniform and non-uniform meshes. The Journal of Supercomputing, 75(12):8168–8206, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hendawi:2019:BLS

[2346] Abdeltawab Hendawi, Jayant Gupta, Jiayi Liu, Ankur Teredesai, Naveen Ramakrishnan, Mohak Shah, Shaker El-Sappagh, Kyung-Sup Kwak, and Mohamed Ali. Benchmarking largescale data management for Internet of Things. The Journal of Supercomputing, 75(12):8207–8230, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chikouche:2019:PPC

[2347] Noureddine Chikouche, Pierre-Louis Cayrel, El Hadji Modou Mboup, and Brice Odilon Boidje. A privacypreserving code-based authentication protocol for Internet of Things. The Journal of Supercomputing, 75(12): 8231–8261, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Jimenez:2019:HBC

[2348] Ernesto Jiménez, José Luis López-Presa, and Javier Martín-Rueda. Hybrid binary consensus in anonymous asynchronous systems using coins and failure detectors. The Journal of Supercomputing, 75(12):8262–8292, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2019:ABC

[2349] S. K. Pravin Kumar, M. G. Sumithra, and N. Saranya. Artificial bee colonybased fuzzy c means (ABC-FCM) segmentation algorithm and dimensionality reduction for leaf disease detection in bioinformatics. The Journal of Supercomputing, 75(12):8293– 8311, December 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kumar:2019:DCF

[2350] C. U. Om Kumar and Ponsy R.

K. Sathia Bhama. Detecting and confronting flash attacks from IoT botnets. The Journal of Supercomputing, 75(12):8312–8338, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2019:CFF

[2351] Kun Li, Shigang Li, Shan Huang, Yifeng Chen, and Yunquan Zhang. Correction to: FastNBL: fast neighbor lists establishment for molecular dynamics simulation based on bitwise operations. The Journal of Supercomputing, 75(12):8339–8340, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        019-02956-w.pdf. See [2372].

Malyshkin:2019:PCT

[2352] Victor E. Malyshkin. Parallel computing technologies 2018. The Journal of Supercomputing, 75(12):7747– 7749, December 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s11227-019-03014-1.pdf.

Bodei:2019:PCA

[2353] Chiara Bodei, Pierpaolo Degano, Gian-Luigi Ferrari, and Letterio Galletta. Programming in a context-aware language. The Journal of Supercomputing, 75(12):7750–7764, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Szustak:2019:UPC

[2354] Lukasz Szustak, Kamil Halbiniak, Roman Wyrzykowski, and Ondrej Jakl. Unleashing the performance of ccNUMA multiprocessor architectures in heterogeneous stencil computations. The Journal of Supercomputing, 75(12):7765–7777, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s11227-
                                        018-2460-0.pdf.

Perepelkina:2019:DLA

[2355] Anastasia Perepelkina, Vadim Levchenko, and Sergey Khilkov. The Diamond-Candy LRnLA algorithm: raising efficiency of the 3D cross-stencil schemes. The Journal of Supercomputing, 75 (12):7778–7789, December 2019. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kireeva:2019:PIT

[2356] A. E. Kireeva, K. K. Sabelfeld, E. N.

Gribov, and N. V. Maltseva. Parallel implementation of a three-dimensional cellular automaton model of the electrochemical oxidation of carbon “Ketjenblack EC-600JD”. The Journal of Supercomputing, 75(12):7790–7798, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Hoffmann:2019:DPF

[2357] Rolf Hoffmann and Dominique Désérable.

Domino pattern formation by cellular automata agents. The Journal of Supercomputing, 75(12):7799–7813, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Torres:2019:SST

[2358] Raul Torres, Julian M. Kunkel, Manuel F. Dolz, and Thomas Ludwig. A similarity study of I/O traces via string kernels. The Journal of Supercomputing, 75(12):7814–7826, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Akhmed-Zaki:2019:IMH

[2359] Darkhan Akhmed-Zaki, Danil Lebedev, and Vladislav Perepelkin. Implementation of a 3D model heat equation using fragmented programming technology. The Journal of Supercomputing, 75(12):7827–7832, December 2019.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kataev:2019:EUD

[2360] Nikita Kataev and Alexander Kolganov. The experience of using DVM and SAPFOR systems in semi automatic parallelization of an application for 3D modeling in geophysics. The Journal of Supercomputing, 75(12): 7833–7843, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Nobile:2019:GMP

[2361] Marco S. Nobile, Paolo Cazzaniga, Daniela Besozzi, and Giancarlo Mauri. ginSODA: massive parallel integration of stiff ODE systems on GPUs. The Journal of Supercomputing, 75(12): 7844–7856, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sukhoroslov:2019:ESS

[2362] Oleg Sukhoroslov, Alexey Nazarenko, and Roman Aleksandrov. An experimental study of scheduling algorithms for many-task applications. The Journal of Supercomputing, 75(12):7857– 7871, December 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Zhu:2019:TLB

[2363] Suxia Zhu, Zhigang Chen, and Guanglu Sun. Tuning lock-based multicore program based on sliding windows to tolerate data race. The Journal of Supercomputing, 75(12):7872– 7894, December 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Lim:2019:ATG

[2364] Roktaek Lim, Yeongha Lee, Raehyun Kim, Jaeyoung Choi, and Myungho Lee. Auto-tuning GEMM kernels on the Intel KNL and Intel Skylake–SP processors. The Journal of Supercomputing, 75(12):7895–7908, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Kholod:2019:FBP

[2365] Ivan Kholod, Andrey Shorov, Evgenii Titkov, and Sergei Gorlatch. A formally based parallelization of data mining algorithms for multi-core systems. The Journal of Supercomputing, 75(12):7909–7920, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Borisenko:2019:CGP

[2366] Andrey Borisenko and Sergei Gorlatch. Comparing GPU-parallelized metaheuristics to branch-and-bound for batch plants optimization. The Journal of Supercomputing, 75(12): 7921–7933, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Prakash:2019:DRA

[2367] Amit Prakash and Dilip Kumar Yadav.

Design and reliability analysis of faulttolerant shuffle exchange gamma logical neighborhood interconnection network. The Journal of Supercomputing, 75(12):7934–7951, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Watve:2019:DIV

[2368] Alok Watve, Sakti Pramanik, Sungwon Jung, and Chae Yong Lim. Dataindependent vantage point selection for range queries. The Journal of Supercomputing, 75(12):7952–7978, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Damodaran:2019:OLC

[2369] Selvam Paranche Damodaran, Vishvaksenan Kuttathati Srinivasan, and Kalidoss Rajakani. Optimized and lowcomplexity power allocation and beamforming with full duplex in massive MIMO and small-cell networks. The Journal of Supercomputing, 75(12): 7979–7993, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Alazzam:2019:HJS

[2370] Hadeel Alazzam, Esraa Alhenawi, and Rizik Al-Sayyed. A hybrid job scheduling algorithm based on Tabu and Harmony search algorithms. The Journal of Supercomputing, 75(12):7994– 8011, December 2019. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bernal:2019:ICA

[2371] Adrián Bernal, M. Emilia Cambronero, Alberto Núñez, Pablo C. Cañizares, and Valentín Valero. Improving cloud architectures using UML profiles and M2T transformation techniques. The Journal of Supercomputing, 75(12): 8012–8058, December 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Li:2020:FFN

[2372] Kun Li, Shigang Li, Shan Huang, Yifeng Chen, and Yunquan Zhang.

FastNBL: fast neighbor lists establishment for molecular dynamics simulation based on bitwise operations. The Journal of Supercomputing, 76(7): 5501–5520, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See correction [2351].

Bai:2020:CPP

[2373] Jianli Bai and Rong Hao. Comment on “Privacy-preserving public auditing for non-manager group shared data”. The Journal of Supercomputing, 76(7): 5563–5577, July 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). See [1931].

Hui:2023:RNM

[2374] Qiuli Hui. Retraction note: Motion video tracking technology in sports training based on mean–shift algorithm. The Journal of Supercomputing, 79(5):5836, March 2023. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04863-z.  See

[2273].

Kim:2023:RNS

[2375] Hyun-Woo Kim, Jungho Kang, and Young-Sik Jeong. Retraction note: Simulator considering modeling and performance evaluation for high-performance computing of collaborative-based mobile cloud infrastructure. The Journal of Supercomputing, 79(5): 5837, March 2023. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-022-04857-x.